A Bibliography of Publications in IEEE Transactions on Parallel and Distributed Systems: 1990–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 2.02

Title word cross-reference

(k + 1) [AEA97]. (m, k) [Ram99]. (N − 1) [LW95a]. 1.5 [LH05]. 2 [HY04, JKA07, ST99a, SY00, SJPS01, TSP^{+}08]. 3 [CYY00, DS05, WH03a, XPL04]. ^{EI} [RRRM09]. d [SV97]. K [LWJ06, BP98, CW00, Chi98, DAA97a, DMR01, HY01, HY04, HNO98c, KP99, KH97b, Kuo01, Li03, LWS04, LBS01, PSK99, PW99, PG07, RC95, SX08, SX09, TLM04, Wan98, YW03a]. L_{2} [WH01]. m [ME93]. M^{3} [BEK^{+}93]. speedup(n) [HM90]. N [CST02, OPZ99, Soh95, BP98, CW00, Chi98, DAA97a, HM90, KP99, PSK99, PW99, PG07, RC95, SX08, SX09, TLM04]. n^{2} [NS95b]. n × n [NS95b]. O((log log n)^{2})

[HNO98a]. O(1) [WH03a, XL08]. O(n) [LM06]. p [Wan04, WLZ08]. ±2^{b} [Nas93]. r [JJ07, Wan04]. wr [KH98].

-Arbiters [Kuo01]. -Ary [SX08, TLM04, BP98, CW00, Chi98, DAA97a, KP99, PSK99, PW99, PG07, RC95, SX09, Soh95]. -Centers [Wan04]. -Connectivity [LBS01]. -Coteries [HY01, HY04, KH97b, KH98]. -Cube [BP98, Chi98, PW99, TLM04, CW00, DAA97a, KP99, PSK99, PG07]. -Cubes [RC95, SX08, SX09]. -Dimensional [Li03, SV97, CST02]. -Dominating [Wan04]. -Exclusion [AEA97]. -Firm [Ram99]. -Fold [YW03a]. -level [ME93]. -Local [LWS04]. -Merge [HNO98c]. -Neighborhood [JJ07]. -Partition [HY04]. -resilient [LW95a]. -Self-Protection [WLZ08]. -Set [DMR01]. -Tree [Wan98].

1 [DM93]. 1-Hop [LJW^{+}07]. 1999 [Ano99g].

2 [GR90, KWOA05, MCH^{+}90]. 2-D [LMN94, TC95b, GR90]. 2004 [Ano05b]. 2008 [Ano08d]. 2009 [Ano09d]. 2D [SY98, TLGP97, YK98, YYS97]. 2D/3D [TLGP97, SY98].

3D [SY98, TLGP97]. 3PC [SK02].

5 [DCSM96].

802.11 [BCG04, JASA08, NK08, XLW^{+}06, ZL07b]. 802.11-Based [ZL07b]. 802.11e [XL04]. 802.15.4 [HPH08, MGZN07, MSM06]. 802.15.4-Based [MGZN07].

A* [MD97]. Ability [SM97, SZ95a]. Abstraction [WMB96]. Abstractions [AGL^{+}98]. Accelerating [CHJ^{+}07, SGTP08]. Accelerator [Ano09c]. Access [AMS97, ADD^{+}02, BRSR08, CWYZ09, DKKS04, Deb96, JGA08, KL02, LJZA04, LH94, Lop02, MR02, NW98, NKP^{+}96, Par01, RO99, RNKZ03, XL04, AM93, BC92, FC91, Geh93, GS91, LC94, KP93b]. accessed [Tho93]. Accesses [YY95, Har91]. Accessibility [KCW09, SSP^{+}09]. Accessible [FARH02]. Accountable [Ros03]. Accounting [BGMZ97]. Accuracy [HV07, HE92]. Accurate [KPBD09, ZL07b]. achievable [KH93]. Achieve [Gen00, TLM04]. Achieving [PS96a]. Acknowledgments [CH04b]. ACOM [CSC07]. acquisition [CR94]. across [ABJ^{+}93, RM90]. acting [MM96]. actions [RPW93]. Activation [CGL07]. Active [BKI06, KMW95, MR03, MBTPV06, MAJ^{+}07]. Activities [SH96]. Actor [AYA09, BBS^{+}09]. Actuator [KHM05, RE09]. acyclic [GY93]. Ad [ALW^{+}03, Ano04d, BK09, BMPP06, BS08, CLW03, CKWC08, DW04a, DW04b, DW06, DPH08, DAMK06, DB08, GJDA06, GYS05, GY07, GS03, ISRS06, JJ07, LCWW03, LWS04, LH06a, LWC^{+}09, LJW^{+}07, NO00b, OSRS06a, OSRS06b, PDH06, SLFW06, TR06, WY07, WO04, Wu02, WCDY06, WD06, WYD07, XP05, YWD08, Yi09, ZL07b]. Ada [SMBT90, STMD96]. Adapt [MTL95]. Adaptation [BES06, CMBAN08, KZN07, LLY04, dLCK^{+}05, JASA08]. Adaptive [BCCP04, BWC^{+}03, BG09, CO94, Chi00, CS02b, CCD^{+}09, DHB01, DWX09, DS03b, Dua95a, Dua95b, DP01, FFPF05, GCCC^{+}04, GLY07, GKK05, GPBS94, GS03, GKG06, HHL08, HP07, HY07, HJB^{+}09, HPH08, JNGS06, KIBW99, KA06, KHY09, KLC97, KS06, KSC03, KgCS04, KL02, Lan95, LB00a, LP07, LC99, LLH^{+}01, MTM02, MLSS07, OKSA01, PC07, PGDS94, PGBI03, QNR99, RCS01, RE09, RLD03, SKK01, SVM07, SCW07, TX08, TW98, TD01, TR04, TR06, TW00, VSD01, VS11, WCH^{+}08, Wu98, Wu00, XCZ04, YR06, ZPY06, DA93, Dua93, KK92, OL92, PGFS94, SH93, YTB92]. adaptive-hash [OL92]. Adaptive-Trail [QNR99]. Adaptively [YJZ97]. Adding [SB94a]. Additions [Ano05b]. Address [KAY^{+}06, QD05, SKS02]. addresses [Kop94]. Addressing [CDV^{+}06, DS05, NSZ02]. Adjacency [RC95]. Adjustable [JJ07]. Adjustment [ZMC03]. Admission [CS02b, HYP02, JXT^{+}04, LLY04, STY09, XHYL05]. Advance [RRX09]. Advanced [CE95, KP09, MAS08, PNZ^{+}02]. Advancements [BP96]. Advances [CMR07]. Affinity [AAD08, ML94, SL93c]. affordable [NE93]. against [CS05, LW09a, PZZ09, SX03, TC07]. Agent [LJW05, MX03, SSsLY03]. Agent-Based [LJW05, MX03, SSsLY03]. Agents [DS02]. Aggregate [CCSC09, CC03, CH08, CB03, DZH05]. Aggregates [CPX06, TCLY07]. Aggregating [BcFGM08]. Aggregation [TF01, ZPY06]. Aggressive [KGMB94]. Agreement [AKNR^{+}04, FMR07, SCY98, STW00, WCY95, WYWZ08, KA94]. AI [DM93]. aid [WG90]. Aided [JK99, SR91]. Algebra [CHC04, KCS^{+}99, AC93, EHJ94]. Algebraic [THT^{+}97, CWL92]. Algorithm [ACT^{+}97, AR97, Ano04c, AMP07, AB03, BCVCV05, BQF99, BT98, BS08, COP00, CS01a, CRS06, CGK04, CY95, CFW98, CD08, CY96c, DW04a, DA98, DTE07, DS05, DB08, DY05, Din01, EW97, EAF00, FE97, FG06a, FB01b, GMRC07, GW96a, GRY07, Gon03, GFG^{+}99, GRT97, GY07, HNO98a, HPT04, yH02, Hsi03, HALT95, HH95, HZ96, wJPP97, JK99, KKM08, KZ96, KR00, KM01, KA99, KC98, Lan95, LO95a, LH05, LM06, LT97, LL06a, LH03, LLWC09, LK00, LC02b, MM98a, MM98b, MS03, McK98, MBM98, MF96, NO97, NO98, OZ96, OB00, Pre99, RCS01, SRD04, SyFL99, SWC95, SSsLY03, SOM05, TW98, jTM96, UKY98, WCL97, WH03a, WR04, WLL^{+}07, WMN99, WYJ^{+}04, YJ97a, YJ97b, YR06, YC95, ZY07, ZH98, BCBzC92, BW94, BLO^{+}94, BP94, CC93b, CH92, CL94, FA94]. algorithm [GR90, HAR94, KSA94, LW95a, LG94, LK94, ME95, MC93, NZ95, NM92, NLM90, Omi90, OL92, Pan93, RST95, RJ94, Sin92, SY93, SCD97, SW92, SR94, Var93, VJ93, VJ94, WL91, WYTD93, WDY93, YD94a, You93, YC96]. Algorithm-Architecture [GMRC07]. Algorithm-Based [CD08, YJ97a, YJ97b, BP94, RJ94, VJ93, VJ94]. Algorithm-Hardware [ZY07]. algorithm-machine [SR94]. Algorithm-Specific [GW96a]. Algorithmic [EAK97, PR05b, PD99]. Algorithms [AF05, AFAGR97, AB99, AV96, ABK98, AD95, BT00, BCVCV05, BCVC05, BcFGM08, BKB96, BCL09, BBG^{+}95, BGOS98, BNO^{+}01, BC96, BCR98, BHK^{+}97, CLW03, ÇF99a, CYW08, CCY03, CC93a, CH04a, CBE93, Che96, CST02, CPhX04, CPX06, CK96, CBDW96, CFR99, DSO02, DÖ02, DVV07, DCF95, FYS05, FARH02, GVV09, GVGD95, GG94b, GG95, GW06, GKK97, HNO98b, HNO98c, HTPS02, Ian97, IB95, Jou03, JKA07, KABK03, KHWT95, KB03, KPK09, KSP09, LC95, Lee97, LL06b, LCB96, LPZ98, LRG99, Li07, Li08, LHSML95, LNO^{+}00, LCL03, LRS02, LH06b, LSVMW07, LWLN97, LZ05, MGZN07, NLW99, NS95a, PHKC09, PPR99, PPP04, PGFS94, RL98, Raj05, RKHM06, RK08, RJ99, Rav07, RLW^{+}07, RS97b, SKK01, SM97, SBF00, SZ02, SVM07, SX07, Sto97, SL01a, SSZ02, Sto04, SY00, SJPS01, TCR96, TR93]. Algorithms [Tse05, TNPK01, VV99, WKS01, WHW05, WLZ08, WH03b, XLPH06, XC01, XTL06, YF97, YKS03, YvdRC05, YD95, YMG03, YZC08, ZY04, ZCLC06, ZCXF09, ZP07, ZT01, ZW02, dCVGG02, AAG94, AC92, Aln94a, Aln95, AC93, AB91b, AIK91, BJS90, BDS94, Cap92, CARW93, CA93, CCCS90, Che95a, EHJ94, EG93, HMR94, IS90, JR93, wJNPS97, KCN90a, KCN90b, KK92, LK90, LWY93, LL94, MS91, Nas93, NGL94, OW91, OSZ92, PJC93, PDC94, RSS90, RWF94, Rao96, RJ90, SC94, SP93, SF92a, SC91, SMJ92, Tak93, TB94, UEA95, WC90, WW92, Zia93]. Aligned [TG99]. Alignment [CHC04, GAL01, LSVMW07]. Alignments [RA04, SA09]. Alive [MRT09]. All-Port [HÖ00, HK95, KLS00, jTM96, YW02]. All-To-All [SR98, SY98, BHK^{+}97, CCY96, FYP07, FH97, GP03, SS01, TG96, YW00, YW01, YW02, CYW94, LS94b]. all-to-many [RWF94]. Alleviate [KZN07]. Alleviating [BP98]. Allocating [Bil94, CT94, HJS^{+}06, HC97, KA96, Men05]. Allocation [AMSK04, CW00, CC99, CYY00, CML05, CXN06, CNT05, DP02, DD95, FLZ09, GBD07, GLV06, HÖ99, HP07, HPT04, HPH08, JLS02, JJ09, KY98, LC95, LKHL03, LJCL08, LWLN97, MEKOT03, MMJ03, PC07, PC05, RTS95, Ram95, RK08, SKJ07, SP95, SJ99, TF96b, VKS^{+}09, WMWL08, WFS09, WHC03, XQ08, YMP08, YLL^{+}07, YL08, YYS97, YD95, YL97, ZX04, ZWX06, ZW02, AM91, CD94, CO95, CS94, KDL91, KLDR94, Lat94, PJC93, SST94, WM93, ZS95b]. Allocations [XCZ02, XCZ04]. Allowing [KY97]. Almost [BP94, DNSC09]. Alternatives [SSP00, YV98, And90, DAF95]. Among [MAJ^{+}07, RPW93, WYWZ08, YA93]. Analysis [AEA97, AM93, AKSS04, AT07, Bak05, BKB96, BCL09, Bor00, CWLR09, CGK04, CHJL04, CPX06, CH08, CY00a, CH95, CYD98, CF94, DW04b, DYJ97, FHA06, Fei05, FYJ^{+}09, GRT97, KGKL08, KW08, KP09, LKK95, LP96, LCB96, Li07, LYW08, Li08, LR96, MM98b, MTL95, ON06, PP96, PF08, PK04, RLW^{+}07, RBSP02, SKJ07, SRT96, SST94, SV97, SRL98, SK95, SSsLY03, SM02, SMH02, TC06, TXL08, TL05, Tos07, TRS90, TKW98, TK96b, Var01, VMXQ04, VR05, WR04, WH98, XPL04, XTL06, YJ97a, YFM98, YJHG06, ZH99b, ADM92, AV94, AC92, AS92, BE92, BCJ90, BDS94, CH92, CTC93, DY93, HK91, KK93b, KGS94, KK92, KS93, LYZ90, ME92, ME93, MS94b, MRW92, MB92, MD96, Pad91, RB90, RM90, SMBT90, STMD96, SF92b, Tze93]. Analytic [LC04, SH93, SLEV03, Yi09]. Analytical [FCF00, HY99, MZA02, OKSA01, RAHM05, Sob96, SE98]. Analyzer [WHL95]. Analyzing [FLP^{+}07, MYA01, HMW93]. And-Parallel [PG01]. AND/OR [ZMM04]. Angle [NO97]. Angle-Restricted [NO97]. Annealed [GS95]. Annealing [CFW98, HM95, LL96, Soh95, BJS90, NZ95, WCF91]. Annual [Ano97a, Ano98a, Ano99b, Ano04a, Ano05b, Ano07a, Ano08a, Ano08d, Ano09d].

Anonymity [HL08, XXZ03, ZB09].

Anonymous

[JKR01, LHL^{+}08, WLHB08, YK96a, YK96b]. Answering [SMH02]. Antenna [LJZA04]. Antennas [DW06, JGA08]. Antiworm [CT07]. Any [CSC07]. Any-source [CSC07]. Anycast [JXT^{+}04, XJZZ00]. Aperiodic [GMM97]. APIs [dLCK^{+}05]. AppLeS [BWC^{+}03]. Application [AAS03, Agr98, BB05, BIWK00, DDV^{+}07, HDRS00, HP06, HALT95, KHM05, KPR05, LCWW03, OSS93, PHKC09, PK99a, QR07, STMD96, SkLC^{+}03, SSRV99, SCP02, SZ04, TSAL97, TS98, TSRS07, VSD01, VJA97, YM09, AM91, BCJ90, KK93a, MN92, SS90, XB93, You93]. Application-Centric [SCP02]. Application-dependent [OSS93]. Application-Driven [SSRV99, BCJ90]. application-oriented [MN92]. Application-Specific [HP06]. Applications [ASS95, BRS07, BCCP04, BKI06, BCF^{+}08, BM00b, BNO^{+}01, BES06, CLB08, CH04b, Che95b, CN02, CN04, CHJ^{+}07, CSR07, CG02a, CG02b, Din01, DÖ02, FB01a, FLP^{+}07, GN06, GB06, HÖD99, HNO98b, HCD97, KKC^{+}05, KKCB02a, KKCB02b, KR00, LCB00, LCGC07, LH93, LSZ09, LWS04, LP07, NO97, NSZ02, OZ96, PK95b, PM96, RNR^{+}03, Ram99, RJ96, Rob04, RRG07, RD09, VKS^{+}09, WC09, WCCR^{+}97, WH03b, WCDY06, XP07, XL96, GH93, HKM^{+}94, HB92, LO95b, MTSDA93, SA94, SSG91, TMTH96]. Applied [CDR98, SKB04, dSF03]. Approach [ASB02, ASS95, AAB^{+}00, BOC09, CS01b, CS02a, CWLR09, CT97, CLS04, DHP^{+}07, FO05, GLY07, GY95b, GMR98, GS08, HP03, JBW^{+}08, JZ04, KH97b, LQZ09, LZTY09, MRLD01, PK00, PD95, RGL05, RAHM05, SP03, SL09, SvVB05, SZ08, TCLY07, TC07, TGV08, TF01, TLGP97, TWH99, VKS^{+}09, WT98, WTCY95, WDY98, WYJ^{+}04, WCR09, YZZ00, YKS03, YM09, YXW03, ZFMS03, dBL98, dB98, CS90, KK93a, O’H91, SSG91, jTM97, YW93]. Approaches [MB07]. Approximate [BM00b, KPK09, LR96, THH08, Tse05, XTL08, KA94]. Approximating [BI95, yCM98]. Approximation [FH03, LH05]. Approximations [Gre98]. APTEEN [MZA02]. AQM [WLL^{+}07]. Arachne [DR98]. Arbiters [Kuo01, ZY07, TC93]. arbitrarily [EA93]. Arbitrary [AMS97, Bar98, VB96, VM04, WM95, LS94a]. Arbitration [MLSS07]. Architectural [KPBD09, SSP00, SKPS01]. Architecture [AGGD04, AGGD05, AAS03, AB03, BS96, DSY99, DZHG04, FV09, GMRC07, GM97, GSS06, ILL07, KH04, KW08, LCGC07, LK07, LWY96, LOSW99, LNOZ03, LLA^{+}06, MR03, MGA^{+}09, OC05, PABD^{+}99, SS08, SCL05, SSP02, Ste96, WCCR^{+}97, YWY08, YYX^{+}09, YKDV02, ZYKG07, ZN04, ZH07c, AS92, AG96, ABDZ94, BCJ90, CPA93, DFD93, Efe92, GP93, HISS94, Lee93, LWY93, MLL92, TC94, YZW94, ZA92]. Architectures [AFM02, AS96, CGM^{+}07, CF01, CBDW96, CG02a, CG02b, Din01, FPGAD08, FJY98, GR06, IT07, LBC03, MCG08, MYA01, OHRW99, RD98, SLEV03, SvAS04, TSG09, XZL05, YYS97, ZYC95, AM93, KSA94, OD93, OS94b, PLW96, RB90, RP94, SP93, SL93a, SRT94, SMS93, YD94b, ZY95, ZL96]. Area [CBD^{+}01, FARH02, SC05, Ant94, CAB93, CCJ02]. ARIMA [TR04]. Arithmetic [RSP02]. Arrangement [HCH99, LC01, BGM94]. Array [BFL^{+}01, CE95, CLPT02, CY00a, DSO02, DDP^{+}98, GWL97, GR06, HTPS02, HCYD01, KKC^{+}05, KP93b, KKC03, LHS03, LPZ98, LCL03, Par95, PPR99, RS97a, SK95, TCR96, TC95b, WHW05, XRY09, Cap92, GR94, JWC94, Lin93, O’H91, SC92, SA93]. Array-Intensive [KKC^{+}05]. Arrays [AKN95, CHC04, Che95b, CM95, Din01, GW96a, LHSML95, PK99a, RJ99, TKP00, TC95a, VMXQ04, WH01, YL96, vDSP96, GM94, LK90, Mar93, NJ94, SF92a, WC90, TL05]. Artificial [SZ03a, SSZ06]. Ary [SX08, TLM04, BP98, CW00, Chi98, DAA97a, KP99, PSK99, PW99, PG07, RC95, SG94, Soh95, SX09]. ASAP [GLY07]. ASCEND [AV96, Nas93]. ASCEND/DESCEND [AV96]. Aspects [AF05, ZJ03, MJ94, NSD93]. Assignable [PH05]. Assignment [AAB^{+}00, BPT03, BRTM09, CB00, CYD98, HTPS02, KGM97, KM02, KA99, LS97, Lee06, NYD09, SKS02, SZXS05, CNNS94, WW92]. Assignments [LO95a]. Assisted [AYA09, CF01, YLW07, ZH07a]. associated [CO94]. Association [BS08, JZ04, PPBSA97]. Associative [SDFV96, WM95]. Assumptions [MRT06]. Assurance [XHYL05]. Assuring [CWYZ09]. Asymmetric [CB00]. Asymptotical [LC02a]. Asymptotics [DF09]. Asynchronous [BCVCV05, BCVC05, BKB96, BCCP04, BBS^{+}09, CF99b, DMR01, FG01, GMRC07, GY95b, HHM^{+}00, HLH04, LL96, LT97, LCB96, LH01, MRT09, QR07, SW95, VM99, WDCK04, CF94, MLS94, MD96, MMSA94]. ATM [KS01]. atomic [KST94, LG90, RPW93]. Attached [MKR00, ZBJ^{+}05]. Attack [YWF^{+}09]. Attackers [LLY05]. Attacks [CS05, CHK07, CPM07, PZZ09, SL09, SX03, WS03, WCBX06, XZG09]. attribute [KG92]. Attributes [HSH^{+}99, PR05b]. Auction [CZLM09]. Auction-Based [CZLM09]. Auctions [CGM05, WLL08]. Augmented [ABC^{+}01a]. Authentication [LHL^{+}08]. Authorization [WRB09]. Auto [FO05]. Auto-Parallelizing [FO05]. Autocorrelated [ZMRS08]. Automata [JASA08, SZ02, SZ03b, SSZ06, TK96a]. Automata-Based [SZ02]. Automated [TC07]. Automatic [AKN95, BW96, EHP98, Fos91, GP92, KCS^{+}99, LL02, MSH00, PD00, RSP02, RR02, SK02, TR04, GB92, KKP91]. Autonomous [BQF99]. Availability [KH98, MJ98, MG09, RD09, TF96a, AT07, DMTB93]. Average [RMO^{+}95, SRT96, GG94b]. AVMON [MG09]. Avoidance [BPT03, CY06, FF98, YM09, Bir93]. Avoiding [WDY98, WCD08]. Aware [ACM08, AD08, Ano07c, Bar98, CJLN09, CNT05, GVV09, GDK09, KZN07, KSC03, Li08, LLGS09, LZR09, MROD07, PS08, RGK09, SY07, SX07, SP07, SGL06, SL01b, TX05, TGV08, VVR07, WS03, WWLS08, XQ08, YGE06, ZRS^{+}05, ZCLC06, ZMM04, ZH05]. Awareness [LXL^{+}05].

B [GM97]. B-Spline [GM97]. Back [AT01, KCD07, LLY05, SOM05]. Back-End [KCD07]. Back-Propagation [SOM05]. Backbone [BMPP06, SY97, WWL06, YWD08]. Backfilling [Fei05, MF01b, TEF07, ZFMS03]. Backoff [XLW^{+}06]. backpropagation [KSA94]. Backtracking [LC01, PG01, RK93]. Backup [MAJ^{+}07, ZJ99]. Bag [BCF^{+}08, Ros02]. Bag-of-Tasks [BCF^{+}08, Ros02]. Balance [PH05]. Balanced [AOB93, BBR07, CTS96, CHHC06, DPS96a, DPS96b, DP02, GZ06, HV07]. Balancing [BCVC05, BCCP04, BBR07, CT08, CK02, CCJ02, DHB01, DHP^{+}07, DB06, DvdMK09, GZ09, GB06, HC99b, JJ09, LRRV04, LC99, LJW05, Mit01, SVM07, SX07, SZ08, TP95, Tse09, WT98, Wu97b, ZRS^{+}05, ZS09, ZH05, ZT01, Bok93, GO93, GT93, LK94, Lin93, WLR93, ZMRS08]. Band [WNKS96]. Bandwidth [ACT06, BGMZ97, CS05, CKWC08, CS02b, DZHG04, GBD07, GLQL09, LKKS05, NE01, PC07, SY07, SSRV99, TCLY07, TSK06, TLGP97, US04, WCH^{+}08, WFS09, WLL08, YL07, ZX04, MS94b, ZS95b]. Bandwidth-Constrained [CKWC08, GBD07, WCH^{+}08]. Bandwidth-Efficient [YL07]. Bandwidth-Optimal [TLGP97]. Bank [BGMZ97, TSP^{+}08]. Banker [LM06]. Banyan [YJHG06, SF95, YN90, YA93]. Banyan-Based [YJHG06]. Banyan-hypercube [YN90]. Barrier [CS95, OS02, SH95a, SCL01, YK98, OD93]. Barriers [Sol02]. Base [PSK99]. Based [AFM02, AJ95, AEA97, AAD08, AA00, AMP07, BQF99, BA07, BGOS97, BES06, BOC09, BRTM09, CS01a, CB05, ÇA99, CCSC09, CBM^{+}07, CT97, CST02, CS05, CY06, CD08, CLY08b, CHC09, CJHG08, CGL07, CZLM09, CMDP09, CAZ04, CNT05, CMBAN08, DS96, DW04b, DP06, EKOAW02, EBS04, FYS05, FG06b, FIMR01, FT97, FYJ^{+}09, GBD07, GPST09, GVV09, GB06, HS99a, HY07, HJB^{+}09, HH08, HLL09, HS98b, HCC06, JZXX99, JJ09, Jou03, JKA07, KKM08, KZ96, KH04, KA06, KP01, KL99, KLH07, KCD07, KK03b, LNYY03, LDCO08, LZ08, LWY96, LMS04, LL06a, LL06b, LLSZ08, LC99, LJLN07, LCL03, LJW05, LS06, LW09c, LQZ09, LZTY09, MKR00, MGZN07, MGQS^{+}08, MKY^{+}09, MX03, MTK06, MAJ^{+}07, MRT06, MBM98, NGB^{+}05, NE01, NGM97, PC07, PPR95, QCC99, RS97b, RLD03, SS08, SF08, SD04, SKB04, SZ02]. Based [SJd^{+}09, SFP03, SSP00, SCO^{+}07, SP05, SC05, SCW07, Ste96, SCP02, SSZ02, Sto04, SvVB05, SDDY00, SSsLY03, Sun02, SS09, SX03, SS00, TJ08, TC04a, TC06, TC07, TCC07, TXL08, TF01, TAKB06, TN08, TPL96, TYK99, TF96b, Tze04, VM99, WC09, WCH^{+}08, WL08a, Wu98, Wu02, XZNX08, YJ97a, YJ97b, YK98, YKS03, YLW07, Yi09, YJHG06, ZYKG07, ZYC95, ZMMS08, ZWX06, ZL07b, ZLKK07, ZH05, ZH07c, ZJWX08, ZL05, ZCSY08, ZCO98, BW94, BP94, CR94, CH92, CTC93, DK92, DD95, DI95, FHRT93, GDI93, HMR94, JF94, LB94, MXEN94, MB92, NE93, RJ94, SMBT90, SSG91, VJ93, VJ94, YK92, DMTB93]. Baseline [YW05b]. Basic [CHB98, DCF95, NO98, WS98, YN00]. Basic-Cycle [CHB98]. basics [PK92]. Basis [CXP09]. Be [MRT06, SVP08]. Beacon [MSM06, XCZ08]. Beam [JGA08, LJZA04]. Behaved [BDL95]. Behavior [Bor00, CHL09, CB03, GY95b, HS99a, NN96, RD98, YJHG06, TMTH96]. Behavior-Based [HS99a]. Behavior-Level [GY95b]. Benchmark [HXA96, HWWX99, KHS07]. Benchmarking [MTSDA93]. Benchmarks [MM07, BE92, EHP98]. Benes [DC98, LO95a]. Best [HY07, KY98]. Best-Effort [HY07]. Best-Fit [KY98]. Between [MT97, PPR99, ZYC95, BCdSFL09, CJPW06, DAF95, EF96, GZ09, HWSH00, QCC99]. Beyond [PW95]. BGP [WZP^{+}03]. Biconnectivity [KR00]. Bicriteria [Tse09]. Bidirectional [DY05, SFP03]. bin [BW94]. Binary [AfAGR00, CCP95, Che95b, LC96b, LNO^{+}00, SF07, YR96, AM90, AM91, CL93, CO94, GM94, Pad91]. binding [RK94a]. Bioinformatics [ON06]. Biological [LSVMW07, YFM98]. Biology [AAB06, Ano05c, LS06]. Bipanconnectivity [SX09]. Bipancyclicity [SX09]. Bipartite [LXN07, YC96]. bipartite-permutation [YC96]. Bisector [WKS01]. Bit [ST99b, SDFV96]. Bit-Pattern [SDFV96]. Bitonic [LB00b]. BitTorrent [LYW08]. BitTorrent-Like [LYW08]. BLAST [ON06]. Block [ASS95, DDP^{+}98, EG93, Har91, JR96, LRG99, PPR99, PHP03, PD99, XRY09, KK93a, SMJ92]. Block-Cyclic [DDP^{+}98, LRG99, PPR99, PD99]. Blocking [HY99, MGA^{+}09, WP00, YJHG06]. Blocks [YN00]. Bluetooth [LSW04, TSK06]. BON [BBR07]. Boolean [CT97]. Boost [CW06].

Boosting [FLMD02a, FLMD02b].

Bootstrapping [MCL^{+}07]. Borrowing [EKOAW02]. Both [CBE93, TCS97]. Bottleneck [BP98]. Bound [BDvD98, GT02, EA93, YD94a]. boundaries [WF94]. Boundary [LCN^{+}07]. Bounded [CH09, CSR07, KWL^{+}09, LZ02, NSU97, HK91]. Bounded-Bypass [CH09]. Bounded-Collision [CSR07]. Bounded-Size [LZ02]. Bounding [LL98]. Bounds [AV96, BC95, HK06, LDG04, LMT98, RO99, VV99, XU01, GG94b, JR94, SRT94, TR93]. Branch [EAK95, MC95, UEA95, YD94a]. branch-and-bound [YD94a]. branch-and-combine [UEA95]. Branching [Lee95]. Breadth [SVP08]. Breadth-First [SVP08]. Break [JBW^{+}08]. Break-In [JBW^{+}08]. bridge [EF96]. Broadband [SA09]. Broadcast [BDD^{+}96, CCY96, DW04b, GP03, HK95, KH04, KLS00, MSMA90, MQ97, NOS99, NOZ02, SR98, SPS98, SLFW06, SPC^{+}02, TJ08, jTM96, THT^{+}97, XTL06, YW02, ZL05, CYW94, LS94b, LG90, jTM97, VB93, XUAS99]. Broadcast-Based [KH04]. Broadcast-Efficient [NOS99]. Broadcasting [BNH99, BBG^{+}95, CFKR98, DW06, HK98, ISRS06, LWS04, PC96, PS96b, SWC95, SSZ02, Sto04, TWH99, VB95, BLO^{+}94, CCCS90, LA93, MS92]. Broadcasts [BLMR05, VB96, ST93]. Broker [DZHG04]. Brokering [BGJ06]. Browsing [LA04]. Bruijn [BCH94, HW97]. BSR [Sto96, XUAS99, XU01]. buddy [LC91b]. Buffer [CY06, CCJ02, GLV06, Par01, VV99, YZC08, ZCL04, DY93, MS93]. Buffered [CCQ^{+}05, GLS07, LKK95, Mha09, MD96]. Buffering [LWY96, MLW06, ZY06]. Buffers [WHM09]. Building [BK09, HLL09, LXN07, YN00]. Built [CXP09, WS03]. Built-In [WS03]. Bulk [FH03, RRX09, YXW03]. bundled [BR94].

Bus [AV96, CG08, CS97b, DSO02, EAK97, FYS05, GP99a, HTPS02, KH97a, LP96, LPZ98, RMO^{+}95, THT^{+}97, TH01, WHW05, BIA^{+}97, Lee93, TV92, WC90, WS93]. Bus-Based [FYS05]. Bus-Networked [CG08]. Bused [Fid92]. Buses [Chu95, LOSW99, PZLS01, RS97a, WH01, GM94, LO95b, SP93]. Butterfly [HWSH00, WMN99, Tze93]. Bypass [CH09]. bypassing [AB94]. Byzantine [AMPR01, BCdSFL09, NT09, SCY98, WCY95].

C [Geh93, FO05, ZH99b]. C/C [Geh93]. Cache [CC03, CH04a, CY00a, CY00b, FPGAD08, GCCC^{+}04, HNY02, HKS^{+}07, KKGS01, MM07, MTL95, PNZ^{+}02, PPP04, PD95, PD00, PPR95, SSP^{+}09, SPC^{+}02, TCO01, VGSS01, WDCK04, WDY98, YZZ00, YZC08, ZCL04, AH91, JF94, LY93a, MB92, NGL94, SG93, SL93c, SF92b, YTB92]. Cache-Based [PPR95, JF94]. Cached [GS95]. Cacheminer [YZZ00]. Caches [WM95, WFP90]. Caching [BB08, ILL07, LSB^{+}07, LWY96, LA06, LAS04, SD04, SWH98, TCC05, WXLZ06, WH98, LWY93]. CAD [HB92]. Calculation [CHB98]. Call [Ano97d, Ano97b, Ano97c, Ano98c, Ano98b, Ano99c, Ano99d, Ano99e, Ano01b, Ano01c, Ano01d, Ano02b, Ano03c, Ano04b, Ano04c, Ano04d, Ano05c, Ano07c, Ano08c, Ano09c, Ano09b, HYP02, SFP03]. Call-Overflow [SFP03]. Can [LLY05, MRT06]. Can’t [LLY05]. Capability [ZY04, HISS94]. Capable [YKDV02]. Capacity [CSC07, SSP^{+}09, TSRS07, ZCLC06, ZL08, ZLP09, KK93b]. Capacity-Aware [ZCLC06]. Capacity-constrained [CSC07]. Capsules [Geh93]. Capture [CAZ04]. CAREL [SR91]. Carlo [You93]. Carrier [CLW03]. Carry [WYD07]. Case [AD98, Fei05, GRT97, LS06, TSJ07, XRY09, DI95]. cases [YK96a]. Categorization [PS08]. Causal [CGK04]. Causes [Fei05]. Cayley [CL97, DD01, VS96, WMN99]. CC [BIWK00, PGBI03, ZY95, AGGD05]. CC-NUMA [BIWK00, PGBI03, ZY95, AGGD05]. CCL [BBC^{+}95]. CD [BB08]. Cedar [TZ97]. Cell [Mha09, SZ03a, BJS90, KBD08, SA09, SVP08]. Cell/BE [SVP08]. Cellular [CS02b, HYP02, JLS02, NSZ02, SZ02, SFP03, SZ03b, SSZ06, XPL04]. Center [Wan98]. Centers [TGV08, Wan04]. Centralized [BCF^{+}08]. Centric [HJH02, PK00, SCP02, YWY08]. certain [BP94]. Certificate [JEG07]. Certificateless [STY09]. Certification [Ara08]. CGIN [Chu96]. Chain [LKHL03, Li07, TKP00, VM04]. Chains [CHC09, JEG07, LL96, MMSM06, HN93]. Challenges [Ano98b, LLY07, TL05, VMXQ04, WA99]. Chameleon [KIBW99]. Changes [BCQD07]. Changing [CH08, Lai00, VJA97]. Channel [BP98, BPT03, HTPS02, JLS02, KL02, MBW02, XL04, ZW02, Dal92]. Channel-Adaptive [KL02]. Channel-Assignment [HTPS02]. Channels [CS97b, GN96, HSH^{+}99, LSF^{+}09, SCK00, SD00b, TPL96, VSD01, ZS95a, Ahu93, DA93, SG94]. Characteristics [MTL95, NKP^{+}96]. Characterization [Bor00, BES06, CY95, KPBD09, KK03b, LJW05, MS99a, MM07, PW99, SCP02]. Characterizing [AD98, TMTH96, YK96a]. Checking [Qad03, TNPK01]. Checkpoint [Qua01, WCLF95]. Checkpointing [AT01, BQF99, CS98, CS01b, CS02a, CCD^{+}09, MS99a, PK92, PLP98, PS96c, QS03, SE98, TKW98, Tsa03, Vai99, WCLF95, KP93a, LNP94]. Checkpoints [CS01b, CS02a, MNS97]. Checks [ANKA99]. Chief [Bhu06b]. Chip [AGGD04, Ano03c, BB05, BJM^{+}05, HP06, KKC^{+}05, LM06, MKY^{+}09, PHKC09, PSGD05, PP05, WOT^{+}07, XL08].

Chip-Scale [BB05]. Chips [KAY^{+}06]. Chitra [ADM92]. Choice [FCF00]. Choices [Mit01]. Cholesky [KBD08]. Choose [KS08a]. Chord [SL09]. Chordal [Ano99f, PK99b, YCTW07]. Churn [SX07]. Churn-Resilient [SX07]. Circuit [AR97, CDR98, HALT95, PC96, PS96b, SJM09, SV97, Bok93, HC92]. Circuit-Switched [PC96, PS96b, Bok93]. Circuits [ZMP07]. Circular [FT97, HS98b, Tze93, WS93]. Circulation [IKOY02]. CLAM [GMR98]. Clarifications [ME93]. Class [IB95, RJ96, WL00, YW01, YW03b, YW04, ZCXF09, AB91b, BL91, CAB93, CI92, CNNS94, LC94, ME92, ME93, Nic92, OW91, Sch91, YD94a, Zia93]. classes [Nas93]. Classical [BS96, O’H91]. Classification [GR06, JWC94, Ksh03, KK03b, MS99a]. Classifier [KGKL08]. Classify [MR02]. Classifying [BOPZ04, XLW^{+}06]. Client [AFM02, CN02, CN04, ILL07, Rob04, TCC05, ICT93]. Client-Server [AFM02, ICT93]. Client-Side [TCC05]. Clients [dLCK^{+}05]. Cloaking [WLHB08]. Clock [EAK95, SS08, ZL07b, dB98, Arv94, OS94a, UEA95, YM95]. Clocking [EA93, PN95]. Clocks [Her00, MB92, TKT92]. Closed [Bar98]. Closed-Form [Bar98]. Closer [QD05]. Closest [WHW05]. Closure [TC95b, SC92, WC90]. Cluster [AAB^{+}00, FHBJ97, FG06b, GB06, HCC06, HJH02, JKR01, KB03, KLH07, KCD07, KWOA05, MSM06, NGB^{+}05, OXL06, RNR^{+}03, SC05, TSRS07, VVR07, XCZ02, ZSMF01, ZN04, ZJWX08, AT07]. Cluster-Based [FG06b, GB06, HCC06, KCD07, NGB^{+}05, ZJWX08]. Cluster/Grid [VVR07]. Clustered [AF05, BP96, CB05, HÖD99, PSGD05, SJd^{+}09, YGE06, ZRS^{+}05, ZH98]. Clusterer [WCR09]. Clustering [BMPP06, DAMK06, GRS99, HP03, KABK03, KB06, Raj05, RGL05, RS91b, SYC03, YYY09, GY93, PLW96]. Clusters [Ano04c, BP06, CdMB05, CRS06, CJPW06, DDV^{+}07, FYP07, FB01a, GKK05, JZ04, LLH^{+}01, LBS05, MAS^{+}07, RK08, SH95a, US04, XCZ04, XQ08, YKDV02]. CM [DCSM96]. CM-5 [DCSM96]. CMP [CASM07, FPGAD08, HKS^{+}07, IT07, SSP^{+}09]. CMPs [CHJ^{+}07]. coalescing [OD93]. Coallocation [BE07]. Coarse [AFAGR97, KL01, DAF95]. Coarse-Grained [AFAGR97, KL01, DAF95]. Coarsest [RL98]. Code [CK08, GAK03, MM07, Pre99]. Codes [CAZ04, CMBAN08, HT06, KLS00, LLL09, MQ97, WL08b, XB98, ZL96]. Codesign [ZY07]. Coding [AJ95, CJHG08, CZLM09, MJ98, WLL08, Kop94]. Coding-Based [AJ95]. Coherence [CLS05, CH04a, CH07, CY00a, CY00b, FPGAD08, GCCC^{+}04, GP99a, MM07, MTL95, PD95, PD00, SPC^{+}02, TF96a, LY93a, MB92, YTB92]. coherency [AH91, DY93]. Coherent [PNZ^{+}02]. Collaboration [KyK09, SGB08]. Collaborative [BRS07, CHK07, CL09, LZR09, MTM02, Sun02, SS09]. Collecting [KK93b]. Collection [Bar98, EVW07, GLY07, HV07, KMW95, LWP07, RKHM06, SNI02a, SNI02b, TX08, HM92, IT93]. Collection-Aware [Bar98]. Collective [BBC^{+}95, Kan01]. Collective-I [Kan01]. Collective-I/O [Kan01]. Collectives [VR05]. Collector [CRN09, MJ06]. Collision [CSR07, MLSS07, NO00a]. Collisions [WDY98]. Colored [JK99, BCBzC92, LR93]. Coloring [Hsi03, JBW^{+}08]. Column [LC96b, SP93]. Columns [BOPZ04]. COMA [ZY95]. combinations [SR94]. Combinatorial [HC99a, YGE06]. Combine [BNBH^{+}95, BDD^{+}96, EAK95, jTM97, UEA95]. Combined [AS99, MRT06, WSB09]. Combining [KGS94, LKK95, LS94a].

Commensurable [SS08]. Comment [Che07, CN04, HS98b, Rob04, SH97, TL05, Tho06, VS11]. Comments [CL97, Sto04, YMP08, YP98]. Commerce [ZWX06]. Commercial [Bor00]. Commit [HRG00]. Common [CLY08b, LL94]. Communication [AB99, AKNR^{+}04, ABK98, Ano04d, BBC^{+}95, BS96, BV05, BC99, CB05, CS94, DS03b, FYP07, FH97, GMR98, Gon03, Gon08, GDK09, GRT97, GS95, GSS96, HS99a, HSLA05, HMR99, HJB^{+}09, HWKH01, JVYA05, JKR01, KCRK00, KB03, KL99, KS03, KgCS04, LB00b, LNYY03, MFLX01, MX03, MJ94, NOZ02, OSRS06a, OSRS06b, PH04, QM97, Res97, RMC95, STY09, SK02, SH96, SS05, SWH98, Sto97, SY98, SDDY00, SS01, SS00, TSAL97, TTB^{+}00, TKW98, Tsa03, TG96, TG99, VRKL96, WCDY06, YW04, YMG03, ZS98, AS92, Ant94, BGM94, Bil94, GR90, Gup92, KSF94, LC91a, LR93, LN93, MXEN94, NZ95, RSV90, RWF94, SS94, SC93, TC93]. Communication-Aware [GDK09]. communication-efficient [LC91a]. Communication-free [CS94]. Communication-Induced [HMR99, TKW98, Tsa03]. Communications [BHK^{+}97, GT02, GBC^{+}07, HCYL06, Li03, SO95, SJM09, YL08, QM94]. Communicators [DFKS01]. Compact [MBW02]. Compaction [BOC09, TC98, NE93]. compaction-based [NE93]. Comparative [ZY95, ZYC95, ZWM99, DT94]. Comparator [CBE93]. Comparing [PBA03, AGE94]. Comparison [BMPP06, DvdMK09, Fan02a, Fan02b, GB00, MDL06, SZ03a, SPF99, Tos07, BL91]. Compartmentalized [Lee06]. Competitive [WH98]. competitors [ÖD96]. Compilation [Agr98, CKK^{+}04, KCRB03, PSC^{+}95, RSP02, SPF99, UZCZ97, PAM94]. Compile [AH91, ASS95, GS91, KA99, MTL95, OS02, RS91a, SL93a]. Compile-Time [ASS95, KA99, MTL95, AH91, GS91, RS91a, SL93a]. Compiled [YMG03, RK94b]. Compiler [BF04, CF01, CK08, CY00a, CY00b, FO05, Kan01, LCB00, McK98, NZP03, PNZ^{+}02, SJM09, SCO^{+}07, YLL^{+}07, YYX^{+}09, NSD93, TMTH96]. Compiler-Assisted [CF01]. Compiler-Directed [CK08, CY00b, Kan01, SCO^{+}07]. compiler-parallelized [TMTH96]. Compilers [Ano97d, Ano97b, Ano97c, FS00, HCYL06, BE92, CS94, GB92, LYZ90, SLY90, TN93b]. Compiling [KM91, LC91a, Pre99, RP94]. Complement [HWKH01]. Complete [CTS96, CW00, FO05, LC96b, SY00, SJPS01, TLGP97, CL93, FD94]. Completion [LL98]. Complex [JJ09]. Complexity [BBD00, CLS05, KA99, SKJ07, THW02, YC95, ZCXF09, AB91b, CARW93, KST94]. Component [KCK^{+}06, RGK09]. Component-Oriented [KCK^{+}06]. Composing [GN06]. Composite [ADD^{+}02, Kuo01, NL02, SF95]. Composition [HL09b, KKS07, PS08, RGK09]. Compositions [GvG06]. Comprehensive [LK07, YC93]. Compressed [EAF00]. Compression [KS06, MNM04, NLW99, YKP08]. Compressionless [KLC97]. Compressions [Kla98]. CompuP2P [GSS06]. Computation [BC06, BGO^{+}96, CKK^{+}04, CPX06, CH08, DGFHR03, GM97, JKR01, JB01, LHS03, MNS97, MSG07, NZP03, RJ05, SS96, Soh95, XH08, YFM98, CWL92, Efe92, GG94a, GR90, WCF91]. Computation-Efficient [XH08]. Computation/Compilation [CKK^{+}04]. Computational [ATML08, AAB06, Ano05c, BGJ06, BP06, FLZ09, KA09, LS06, RD09, SVM07, SZ08, VVR07, WBO^{+}01, XZNX08, wJNPS97]. Computationally [Ara08]. Computations [BW96, BGOS97, Chu95, GWL97, GRS99, KCRK00, LRRV04, LT00, MR06, NO98, PM96, SkLC^{+}03, YF97, YXW03, AMAM94, CNNS94, HE92, ML90, Nas93]. Compute [EK95, HNO98a]. Compute-Intensive [EK95]. Computer [BA97, BHL^{+}07, CV08, Chu95, GG95, JK99, RJ99, SR91, SP03, Var01, WS98, WS00, vDSP96, CPA93, Don91, GG94b, NLM90, SC93, YK92, BG90]. Computers [AGWFH97, AFAGR97, Ano97d, Ano97b, Ano97c, BBC^{+}95, GKS95, Lee97, Li08, MT97, PZLS01, SGTP08, SW96, YFJ^{+}01, ATG92, CCCS90, DK92, GK93, HISS94, HQL^{+}91, JS90, KK94, KDL91, KLDR94, SP93, SW95, WLR93]. Computing [AN94, ACM08, AAD08, AAB^{+}00, Ano01b, Ano01c, Ano01d, Ano09c, Ano09b, ABC01b, BNBH^{+}95, BWC^{+}03, BFL^{+}01, CS01b, CS02a, CW02b, CY96b, CK02, DÖ02, EBS02, FLP^{+}07, GBD07, GSS06, HHM^{+}00, HJH02, JKR01, KKS07, KB03, KL99, KSME08, KL02, LZ08, LLGS09, LSBS98, LBS05, LWN98, LMT98, MTM02, MX03, ON02, PS08, PC05, PS96c, Ros03, RD09, SRL98, SC05, SZ03a, SZ03b, TSAL97, TS98, TGV08, TAKB06, THW02, VB95, WNKS96, WOT^{+}07, WL00, YK96a, YK96b, YDW^{+}09, YK03, Zha03, ZS98, ZH07b, ZP07, ZW02, CO95, CYW94, DGB^{+}96, EA93, FA94, SR91]. Concept [CCJ02, KCN90a]. Concepts [LO95b]. concurrence [AB91b]. Concurrency [KWH02, FHRT93]. Concurrent [AG96, Ant94, EDO06, GDJ94, HISS94, KMW95, Pan93, XRR00, BCBzC92, CTC93, LNP94, TH93, VJ94, Geh93]. Condition [Dua95a, Dua96, VS11]. Conditional [HL09b, Lee95]. Conditions [NX95]. Conference [YW04]. Configurable [DDY99, RSP02, SY00]. Configuration [Add97, HDRS00]. configurations [LK94]. conflict [BR91]. conflict-free [BR91]. Conflicts [YD95]. Conformed [PSK99]. Congested [hKY08]. Congestion [BLD05, CSH00, CY06, FH97, GW06, KZN07, LSC95, SP05, TLM04, TR06]. Conjugate [GKS95]. Connected [AD95, CL00, CXP09, Chu95, CY96c, DW04a, GG95, KWL^{+}09, Kla98, LW95b, LWLN97, MBM98, PZLS01, TKP00, WCY95, WL00, Wu00, dCVGG02, CCCS90, CT94, CS92, EF96, GG94b, MC93, PN93, SP93, TC94]. Connecting [Add97]. Connection [AM06, NSZ02, AS92]. Connection-Limited [AM06]. Connectionless [CHA07]. Connective [KH97a]. Connectivity [AYA09, AD09, LBS01, LWXS06, SRZF04, Aln95]. Conquer [CPM07, LRTZ96]. Conscious [VKS^{+}09]. Consensus [DMR01, FIMR01, LC02a, MP91, NCV05, SCY96, TYK99, WCR09, AB91a, Fu97]. Consensus-Based [FIMR01]. Conservation [TSRS07]. Conservative [BT00, HN93, Nic92, WHL95]. Consideration [SH96]. Considerations [CY00b, KPC09, SZ95b, IC92]. Consistency [AK99a, CLS05, CH95, KKGS01, Lee91, LXL08, Qad03, TC04a, TC06, TCC07, TXL08, WDCK04, LH94]. Consistent [AJF96, GMS09, HMR99, HK06, MNS97, MG09, NX95, RS08, Vai99]. Constant [Aln94a, BM00a, BGOS98, CL97, Gen00, HALT95, wJNPS97, Sto96, WC90, Aln95, EA93, KS91, VS96, ZA92]. Constant-Time [BGOS98, Aln94a, Aln95]. Constrained [BKS03, BBD00, BGOS98, CKWC08, GBD07, GAG96, HÖ99, KHM05, KSME08, TX08, WCH^{+}08, ZLAV04, ZPY06, ANN95, AMAM94, CSC07, SS94, SL93a]. Constraints [AA00, BRS07, CKC08, GLV06, GLQL09, LT00, RC95, RSG06, ZL08, ZLP09]. Constructing [KPK09, KWL^{+}09, KWH03, KH97b, LS96, ST99b, WCL97]. Construction [AfAGR00, HY05, JVYA05, LCN^{+}07, PH96, TSK06, XP07, YWD08, Sch91, You93]. Constructions [AM99]. Constructive [DR94]. Consumption [BP98, KGKL08, KA09, ZS09]. Containing [LH03, WNKS96]. Contaminations [JBW^{+}08]. Content [CHA07, CLB08, CF08, KYB08, TX05, VR05, ZYKG07, ZH07c]. Content-Based [ZYKG07, ZH07c]. Contention [BGMZ97, DMKJ96, KP99, MF01a, SS05, ZYC95]. Contents [Ano00b, Ano00c, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, TC04b]. Context [HV07, SS09, YK03]. Context-Based [SS09]. Context-Sensitive [YK03]. Contextual [JJ09]. Continuous [BV05, Gon08, JN08, LL02, SBK02a, SBK02b, XRY09, HN93]. Continuous-Media [BV05, LL02]. Continuum [AD09]. Contributory [AKNR^{+}04]. Control [ASB02, ANKA99, BÖ98, BRSS08, BLD05, BG09, CWYZ09, CS02b, DDY99, DWX09, DF99, HYP02, JJ07, JXT^{+}04, KWH02, KL02, LJZA04, LLY04, LLY07, LWS04, LH06a, LH06b, Lop02, LWK05, LLA^{+}06, MGZN07, NW98, PK99a, Ram99, RLD03, RNKZ03, SRT96, SLFW06, TB93, TLM04, TSJ07, TK96a, WJLK07, WCH^{+}08, WD06, XHYL05, XL04, YXW03, ZL07a, Bir93, Dal92, FHRT93, NSD93, SS90]. Control-Based [RLD03, WCH^{+}08]. control-flow [NSD93]. Control-Theoretical [ASB02]. Controllable [RAHM05]. Controller [HY07, WOT^{+}07]. Controllers [CH07]. Controlling [TF01]. Conventional [KET06]. Convergence [BCVCV05, BKB96, HPT04, HH95, Kin06, dB98]. conversation [WF94, YK92]. Conversion [ZY04, ZY06]. Convex [BGO^{+}96, HNO98a, AD98]. Cooperating [CF95]. Cooperative [BB08, ILL07, KA09, KYB08, WRB09, WCDY06]. Coordinated [CS98, CLY08a]. COPACC [ILL07]. copies [AGE94, BL91]. CoProcessors [KSWR03]. copy [LG94]. copying [IT93]. CORBA [AFM02, FWDC^{+}00, LNYY03, MFLX01]. CORBA-Based [AFM02]. Core [CGM^{+}07, CHJ^{+}07, DW03, DZHG04, GS03, JZXX99, KCRK00, LRG99, MGDZ07, Wan98, WFS09]. Core-Based [JZXX99]. Corona [BBS^{+}09]. Correcting [KLS00, XB98]. Correction [Ano99g, Ano99f, Ano99h, CS02a, DPS96a, MBW02, MTM02]. Corrections [Sto04, ME93]. Correlated [MM07]. Correlation [LWP07, MAJ^{+}07, SLT03]. Corrupted [HZ97]. Coscheduling [FFPF05, SL06]. cosine [MM96]. Cost [AAB^{+}00, Chi98, CZLM09, GG09, GvG06, GMCB01, JLF03, KB03, LW09a, MLW06, MRLD01, MAS^{+}07, MKY^{+}09, OZ96, OC05, PS96c, Qua01, RvG02, Sar93, SWH98, TC04a, TC04b, WKS01, WWL06, XXZ03, YW05a, BL91]. Cost-Effective [JLF03, MRLD01, MAS^{+}07, YW05a]. Cost-Efficient [MKY^{+}09]. Cost-Optimal [OZ96, WKS01]. Costs [ABK98, KDW01, KM02, SRL98, SY98, TF96a, Bil94, Gup92]. Coterie [HY01, HY05, NM92]. Coteries [BI95, HY97, HY01, HY04, KH97b, KH98, IK93]. Counter [WS03, XLW^{+}06]. Counters [SY97]. Counting [GPST09]. Coupled [ADG^{+}08, LJLS09]. coupling [YD94b]. covariance [LH93]. Coverage [AD09, BSCB09, LWXS06, RLW^{+}07, WT08, XLPH06]. Covered [FG06b]. Covering [TF96b]. Covers [PKL06]. CPU [US04]. CPUs [SL06]. CRAP [KHWT95]. Crash [RCS01, VJA97]. CRCW [WH03a]. creation [MKH91]. Critical [AD09, HK06, Hol98, KA96, XTL06]. Critical-Path [KA96]. Cross [DAA97b, ZCXF09]. Cross-Layer [ZCXF09]. Crossbar [Mha09, WL00, TC93, YC93]. Crossbar-Connected [WL00]. Crossed [CSH00, Fan02a, Fan02b, FLJ05, Wan08, Efe92]. Cryptography [BRTM09, EP05]. Cube [BP98, CL00, Chi98, CY96c, HGC05, JVYA05, Kla98, LCRW98, PW99, PN93, SCL00, TLM04, TF96b, Wu98, CW00, DAA97a, Efe92, KP99, MC93, OC93, ÖD96, PSK99, PG07, SG94, SB94a, TC94, ZL96]. Cube-Based [Wu98]. Cube-Connected [CL00, CY96c, Kla98, MC93, TC94]. Cubes [CSH00, Fan98, Fan02a, Fan02b, FLJ05, FJL07, cFC98, Hsu93, HWSH00, JHK97, RC95, Sca99, SX08, Wan08, Wu97a, SX09]. Cubic [COP00, GD95, SP95, YP98]. Cubical [LW95b, Cap92, SC94]. Curves [GM97, PB96]. Customized [BJM^{+}05]. Customizing [HSH^{+}99]. Cut [CFKR98, Dua96, KP01, QNR99]. Cut-Through [CFKR98, Dua96, KP01, QNR99]. Cyber [Ano08c, TGV08]. Cyber-Physical [Ano08c, TGV08]. Cycle [CHB98, GW06, LH05, Ros02, RH04, ZKB08, SKF94]. Cycle-Stealing [Ros02]. Cycles [BT98, CL00, HCH99, Kla98, LW95b, MS03, Wan08, MC93, TC94, YM95]. Cyclic [DDP^{+}98, cFC98, HWSH00, LRG99, LW09b, MJRS06, PPR99, PD99, TG99]. Cyclic-Cubes [cFC98, HWSH00].

D [CYY00, DS05, GR90, JKA07, LMN94, ST99a, SY00, SJPS01, TSP^{+}08, TC95b, WH03a]. DaAgent [MX03]. Daemon [KY97]. DAG [BOC09, KLH07, KGS94, MLS94, WSG01]. Dags [CMR07]. Daisy [VM04]. DASH [LLJ^{+}93]. Data [AKN95, AMY09, AMS97, AM06, AKSS04, BcFGM08, BW96, BE98, CW02a, CHC04, CS97a, CL09, CY00a, CHB98, CJPW06, CN02, CN04, CGM05, CAZ04, CSR07, DYJ97, DGFHR03, EBS02, EDO06, EVW07, GAL01, GLY07, GLV06, GSS96, HV07, HQL^{+}91, HCYL06, HH95, HZ96, JLDC05, JVYA05, KK04, KCS^{+}99, KCW09, KAY^{+}06, KCPT96, KET06, LAV03, LC95, Lee97, LRG99, LSCZ07, LCL03, LRS02, LWP07, MY07, MNM04, MTL95, NZP03, NSD93, ON06, OXL06, PK99a, Par95, PHP03, PC05, PP96, PS03, PSC^{+}95, PPBSA97, PLT00, PK04, PW95, RKHM06, RSB97, RD98, Rob04, RJ05, Sah00a, SF08, SKB04, SkLC^{+}03, SvVB05, SPF99, TS98, TX08, TGV08, TF96a, TTB^{+}00, XCZ04, XL04, XRY09, XTL06, YKP08, ZS09, ZXZ^{+}09, ZH98, ZPY06, AB91a, CS94, DY93, EG93, GDJ94, GB92, HN90, KN95, KCN90a]. data [KCN90b, KGS94, LHS92, LYZ90, RS91a, RST95, SMS93, SB94b, TB93, TT94, WYTD93, WDY93, WT92, HSWB07]. Data-Driven [KET06, PK99a, ZXZ^{+}09]. Data-Flow [CS97a, CY00a, EG93]. Data-Gathering [ZS09]. Data-Intensive [ON06, OXL06, XCZ04]. Data-Parallel [GSS96, LC95, SPF99, HQL^{+}91]. Database [FCF00, ZBJ^{+}05, GD94, Omi90, TB93, Var93]. Databases [GLV06, HCY97, LC04, Men05, WH98, PK92]. Dataflow [BG90, AM93, Lee91, LHS92, PAM94]. Dataspace [SvVB05, CR90]. DAW [CT07]. dBCube [CAB93]. dbx [NE01]. DCMP [ZKB08]. DDoS [CS05, CHK07, LLY05, SX03, WS03]. Deadline [KGM97]. Deadlock [BC96, CBD^{+}01, DA93, Dua95a, Dua95b, Dua96, DP01, DLPP05, FF98, GFG^{+}99, JKA07, LMN94, LPD05, MRLD01, PPD03, RLD03, SP03, SP05, TW00, VS11, WP00, XL08, Bir93, Dua93, GPBS94, PGDS94, PGFS94, PN93, STMD96]. deadlock-and [GPBS94, PGDS94]. deadlock-avoidance [Bir93]. Deadlock-Free [BC96, CBD^{+}01, Dua95a, Dua95b, Dua96, DP01, DLPP05, JKA07, LPD05, PPD03, TW00, VS11, DA93, LMN94, Dua93, PGFS94, PN93]. Deadlocks [BCR98, PW99]. Deallocation [LPE^{+}99]. deBruijn [GP93]. Debugger [NE01]. debugging [GH93]. Decentralized [BCVCV05, BBR07, LC02a, RGL05, SVM07, SBK02a, SBK02b, WJLK07, WZZ09]. Deciding [Ost90]. decision [YK96b]. Declustering [SL93b, Tos07, GD94]. Decoding [Sto96, THH96]. Decomposed [CDR98]. Decomposition [AAD97, ÇA99, KGKL08, KR00, LK94, PLT00, SK02, WMB96, MS94b]. Decompositions [PD99]. Decoupling [GBC^{+}07]. Deep [GR06]. Deeply [ZMP07]. Defending [SX03]. Defense [CS05]. Deferred [DYJ97]. defined [MM96]. Deflection [BC95, FR96, Kuc01, RS97b]. Deflection-Routed [FR96]. Degradation [YJ97b, HW91]. Degree [CL97, EF95, HALT95, LSW04, WMN99, YV98, PN93, VS96]. Degrees [cFC98]. Delaunay [LCWW03, LSW04]. Delay [AH06, BRS07, BGMZ97, BC95, CS01a, DF09, Fu97, LLY04, LLA^{+}06, SJKC06, TYK99, TSJ07]. Delay-Optimal [CS01a, Fu97]. Delays [DHP^{+}07, GRT97, VRKL96, BGM94, BC92, RS94]. Delegated [Ara08]. Delivery [BV05, CLB08, DHN95, Gon08, LLX06, SL01a, TC04b, XHYL05]. Demand [CZLM09, HL09a, ILL07, JGA08, LZTY09, SKS02, WL08a, XTL06]. Demands [XCZ02]. Demonstration [GB92]. Denial [CPM07, SL09]. Denial-of-Service [CPM07, SL09]. Density [AD09]. Departure [CHL09]. Dependability [PPD03, DK92]. Dependable [Ano98c, ABC01b, HSH^{+}99, PABD^{+}99, SR99]. Dependence [BE98, PP96, PK04, TN93a, KKP91, LYZ90, SF92a, VJ93, WT92]. Dependences [PW95, XC01, KS91]. Dependency [CTC93, TKW98]. Dependent [CASM07, SP03, AT07, OSS93]. Deployment [CBM^{+}07, SKCL09, WT08, YLW07, YG08]. Depth [CS90, PWW00, FHRT93]. Depth-First [PWW00, CS90]. Derived [WL97]. Deriving [Abr97, XP07]. DESCEND [AV96, Nas93]. Description [QS03]. Design [ANKA99, AS96, ABS01, Ano04c, ACD^{+}09, BDD^{+}96, CRS06, CSR^{+}09, CJHG08, CV08, CY00b, CL05, CS03, Din06, FVR03, GVV09, GMCB01, GMR98, HCHM09, HP06, HY07, IC92, JKA07, KYD^{+}07, KCN90b, LB00a, kLCC^{+}06, LG08, LK04, LAS04, LLA^{+}06, MNM04, MB92, MCG08, MYA01, Pad91, Pak07, PGBI03, RB90, RLW^{+}07, SKJ07, SBF00, SVM07, SMBT90, SH94, SF09, SP07, SM02, TC95a, VJ94, WMXZ06, WF06, XPL04, YJ97a, YTB92, YN00, LKG92, TV92, WF94]. Design-Space [MCG08]. Designing [Ano98b, BP96, BC96, CCCS90, GWL97, KHWT95, THH96, WA99, WCR09, YK98]. Designs [TC95b, YW05a]. Detecting [HZ97, ISAZM09, MSM09, SM97, SWWJ08]. Detection [ANKA99, AMPR01, BCVCV05, BT98, CHK07, CK96, DTE07, DL02, FMG02, GW94, GW96b, HS99a, LT97, LLS06, LCN^{+}07, MS03, MSG07, NO00a, PK00, RLW^{+}07, RLD03, RNKZ03, TT01, XL08, XL96, ZLKK07, GMG96, HISS94, LW95a, TH93, VJ94]. Detectors [HHM^{+}00]. Determination [CH01, HMR99, KCS^{+}99, KL99]. Determining [HMW93, Tho93]. Deterministic [BRS97, CF95, KLH07, KWOA05, PF96, XZG09, XB98, AV94]. DEUCON [WJLK07]. Developing [GMS09, LPD05]. Development [TS98, Gab90]. Devices [CKK^{+}04]. DFT [GR90]. DHT [CSC07, LQZ09, ZH05]. DHT-Based [LQZ09, ZH05]. Diagnosabilities [CCC05]. Diagnosability [Fan98, Fan02a, Fan02b, HC09, HT07]. Diagnosis [CBE93, DC98, Fan02a, Fan02b, HALT95, KHM05, SS07, SB04, BP94, LS94c, Rao96, VJ94]. Diagonal [TLGP97, YFJ^{+}01]. Diagonal-Propagation [TLGP97].

Diagram [AD08, EW97]. Diameter [DAA97a, DAA00, EF95, MC93, TR93]. Diameters [KWL^{+}09]. Diamond [PK01]. DiCAS [WXLZ06]. Dictionary [NLW99, YL96, FC91]. Difference [EAF00, PR05b, PR05a, Kop94]. Different [KKCB02a, KKCB02b, BDS94]. differential [You93]. Differentiated [GRY07, LAS04, RAHM05, SY07, WFS09]. Differentiation [TJ08, XP05, ZX04, ZWX06]. differing [YA93]. Difficulty [CJLN09]. Difficulty-Aware [CJLN09]. DiffServ [LLY04]. DiffServ-Enabled [LLY04]. Diffusion [SKK01]. Digital [KKC03, LOSW99, SMJ92]. Dimension [BC99]. Dimension-Order [BC99]. Dimensional [AD09, yCM98, CWCC07, CST02, CC99, GW96a, LCRW98, LHS03, Li03, SV97, ZWX06, LC91b, SF92a]. Direct [BA07, DHN96, GY95a, MDL06]. Directed [BM00a, CK08, CY00b, GT02, Kan01, LPE^{+}99, SCO^{+}07, GY93]. direction [PKK93]. Directional [AJF96, DW06, GJDA06, YWD08]. directories [LY93a, SG93]. Directory [AGGD05]. Dirty [DYJ97]. discipline [ZLE91]. Disconnected [KKGS01]. Discovering [NT09]. Discovery [AOK09, AMH08, CHC09, DP06, MG09, YK03, ZZMN07]. Discrepancies [PM02]. Discrete [NL02, TSP^{+}08, XC04]. Discrete-Event [NL02]. Disjoint [KWH03, PKL06, YW03b, YW05b, YD95]. Disk [BSCB09, DP02, JO95, LL02, Par95, SCO^{+}07, TL05, VMXQ04, XRY09]. Diskless [PLP98]. Disks [MKR00]. Dispersal [JEG07]. display [IA95]. Dissemination [EVW07, Gon08, KMG03, MDSS09, RHM09, THH08, BFP96]. Distance [CPhX04, yH02, Hsi03, LHS03, SWWJ08, WH03a]. Distance-Hereditary [yH02, Hsi03]. Distinct [YK99]. Distributed [AD98, AS99, AKN95, AJ95, AEA97, Agr98, AK99a, ACM08, AJMJS03, AJF96, ABS01, AKSS04, Ano97d, Ano97b, Ano97c, Ano02a, Ano07c, Ano08c, Ara08, AMH08, AMP07, BQF99, BcFGM08, BRSS08, BBD00, BV05, BCF^{+}08, Bor00, BT98, BG09, CLW03, CS98, CS01a, CCY03, CG08, CC93a, CLJ^{+}04, lCL95, CT02, CPX06, CPM07, CT07, CH08, CWYZ09, CK96, CY96b, CK02, CS96, CLS04, CYD98, CF99b, DPN09, DA98, DPH08, DTE07, DHP^{+}07, DB06, DS02, Din06, DL02, EBS02, EP05, EDO06, EVW07, FHA06, FHRT93, FJY98, FI95, GB00, GAL01, GG09, GMS09, GY95b, GBD07, GFG^{+}99, GLV06, GY07, HDRS00, HY05, HHM^{+}00, HSH^{+}99, HKM^{+}94, HM95, HPT04, HCD97, HJH02, JR96, JNGS06, JS90, JXT^{+}04, JLS02, JW00, KMW95, KKGS01, KKM08, KHM05, KGM97, KH04, KR00]. Distributed [KPK09, KK93a, KL99, KCW09, KA05, KW08, LTZS06, LSB^{+}07, Lee97, Lee06, LJCL08, Li07, LC99, LCL03, LLL09, LJW^{+}07, LK00, Lop02, LC04, LWK05, LC02b, MZ05, MJ98, MNS97, MS03, MJRS06, MBTPV06, MMJ03, Men05, MG09, MOFD05, MROD07, MP97, NSU97, PAM95, PR05a, PN95, QD05, RvG02, RKHM06, RSB97, RGL05, RMO^{+}95, RGK09, RHM09, RBSP02, RLD03, RRFH98, SF08, SM97, SKS02, SKCL09, SBK02a, SBK02b, SH95a, SGB08, SCK00, SW96, SE98, SP05, SCW07, SvAS04, SJ99, SB04, SNI02a, SNI02b, SS09, SM02, SMH02, TCLY07, TF01, TSK06, TD01, TF96a, TM97, Tho06, TH06, TP95, Tse05, TT01, VVR07, WXLZ06, WWL06, WCBX06, WJLK07, WT08, WOT^{+}07, WH98, XHYL05, XL04, XLW^{+}06, XCZ08, XB98, XRR00, YF97, YW98, Zha03, ZS98, ZH98, ZPY06, ZKB08, ZJWX08]. Distributed [vDSP96, vdMDM07, ADM92, Arv94, BGM94, BIA^{+}97, Bil94, CR94, CO95, CY92, CH92, CYW94, CF94, Fu97, GW94, GG94a, GW96b, HMR94, IK93, KP93a, KK93b, KM91, Kum92, KH93, LW95a, LKG92, LY94, LY93b, MN92, MSMA90, MR92, MMSA94, OSS93, PJC93, PLW96, PK92, RS94, RS91a, RP94, SST94, SH93, SC93, SH94, SM94, SSG91, Sin92, SR91, SY93, SW92, Tho93, TKT92, Var93, VB93, WCS92, WS93, WM93, YJZ97, YK92, ZSLW92]. Distributed-Memory [DA98, RvG02, SST94]. Distributed-Parallel [MJ98]. Distributed-Shared-Memory [Bor00]. Distribution [AF05, Bar98, BGJ06, CHA07, CF08, CWCC07, CN02, CN04, DDV^{+}07, GAL01, GLQL09, KM02, KYB08, Lee97, Li03, MZ05, NZP03, Rob04, SF08, SvVB05, TC04a, TX05, VR05, WCD08, XH08, YM09, CV92, RS91a]. Distributions [LRG99, PSC^{+}95, TG99]. Distributive [CY96c]. Diverse [LG08]. Divide [CPM07, LRTZ96]. Divide-and-Conquer [CPM07]. Divisible [Bar98, BCL^{+}05, CG08, CWCC07, DW03, GKK05, Li03, SRL98, VM04, YvdRC05]. division [QM94]. DNS [WZP^{+}03]. DOACROSS [CY96a, CY99, KS91, XC01]. Document [Tse05]. Documentation [GM09]. Documents [BV05]. Doing [SF09]. Domain [BJM^{+}05, GMS09, Pak07, Pre99, PLT00, SK02, SKB04, SCP02, BGO^{+}97]. Domain-Based [SCP02]. Domain-Oriented [GMS09]. Domain-Specific [Pak07, Pre99, BGO^{+}97]. Domains [CHK07, ADM92]. Dominating [DW04a, KWL^{+}09, SSZ02, Sto04, Wan04, Wu02, WCDY06, jTM97]. Dominating-Set-Based [Wu02]. Domination [yH02]. Domino [LNOZ03]. Double [DY05, SZ95a]. Double-Loop [DY05]. Down [KP01, KDL91]. Down* [SRD04]. Downlink [MSM06]. Download [LA04, SJKC06]. DP [JKR01]. dQUOB [PS03]. DRAM [WHM09]. Draw [COP00]. Driven [BÖ98, CML05, KET06, LZTY09, PK99a, PPR95, RE09, RBSP02, SSRV99, SJKC06, SJ99, WR04, ZXZ^{+}09, BCJ90, HE92, HB92, NGL94]. DRP [GJDA06]. DSC [YG94]. DSDM [AMH08]. DSM [CH04a, LBS05, PBA03]. DSP [FO05, GR94, SZXS05]. Dual [CDV^{+}06, LSZ09, MGDZ07, OC05, RMO^{+}95, SCY96, BR91, CV92, KGM96, MP91]. Dual-Core [MGDZ07]. dual-network [CV92]. Dual-Objective [LSZ09]. Dual-Thread [OC05]. Duality [CMR07]. Duplicate [FRGJ07, MD97]. Duplication [AK98, BKS03, BOC09, CKC08]. Duplication-Based [BOC09]. Dynamic [AMP07, BCVC05, BM00a, CdMB05, CBD^{+}01, CO95, CRN09, CKC08, CHB98, CAZ04, DB08, DHP^{+}07, DB06, DvdMK09, DIM97, DLPP05, DMKJ96, HKL00, HV07, HCYL06, HW08, HS99b, JLS02, KKS07, KBC^{+}01, KSME08, KPC09, KA96, LW95b, LLY04, LCB96, Li08, LBS01, LLWC09, LPD05, MM98a, MM98b, MMJ03, OB00, PP96, PB96, PPD03, PS03, Pre99, RRRM09, RPW93, SKK01, STW00, SB04, SS00, TSG09, TC04b, THH08, TF96a, VB95, WL08a, WT98, WLL08, XCZ02, XZL05, XC01, YXW03, ZMC03, ZLP09, ZT01, AM93, GDI93, HK93, HLV94, Lee93, LC94, OSS93, Sin92, WLR93]. Dynamically [DDY99, TW98]. Dynamics [MZT08, SGTP08, YD94b].

e-Commerce [ZWX06]. E-Kernel [MS94a]. e-Transaction [QR07]. E-Transactions [FG01]. Ear [KR00]. Early [DGFHR03]. EDF [Bak05]. Edge [CSH00, FH97, HL09b, KWH03, RS08, SLH97, WY07, LR93]. edge-colored [LR93]. Edge-Disjoint [KWH03]. Edge-Fault [HL09b]. Editing [SS09]. Editor [ACM08, Bhu06b, Bhu06a, Bhu07a, Bhu07b, Bhu08, Bhu09b, Bhu09c, KMT91, Yew03, Yew04a, Yew04b, Yew05a, Yew05b]. Editor-in-Chief [Bhu06b]. Editorial [AAB06, Bhu06b, Bhu09a, CRS06, IT07, Law95, Law97, PP05, Sta98, Sta99, Sta00, Sta01, Sta02, SR99, Yew02, Yew06, Ano99g, GZ03, Zha03]. Editors [LL07, ON02, WA99, ZH99a]. Effect [CC03, CHL09, ZLE91]. Effective [CY96a, JLF03, JKA07, KM02, KA96, LLY05, MRLD01, MAS^{+}07, SRD04, THW02, WX07, YW05a, YL97, AN93, SH94]. Effectively [LSF^{+}09, OXL06]. Effectiveness [WCBX06, Sar93]. Effects [HWWX99, KSP09, WSNA95]. Efficiency [CW06, CTF09, LH06b, MGDZ07, MT97, RK93, TT94]. Efficient [ACT06, Ara08, AD95, AB03, BCVC05, BGBP01, BHJ02, BG09, BHK^{+}97, ÇF99a, CHA07, CF00, CCSC09, yCM98, CC03, CBE93, Che95a, Che95b, CW00, CT02, CPhX04, CY96b, CC98, CC99, CCD^{+}09, CHB98, CLS04, CMDP09, DW06, DZ04, DS94, DDV^{+}07, EBS02, EDO06, FJY98, FARH02, GPST09, GVV09, Gon03, GJDA06, GAK03, GW06, GLV06, GKG06, HÖ00, HHL08, HP06, yH02, HW97, Ian97, ISRS06, IB95, JZXX99, JTP^{+}08, JTC08, JB01, KABK03, KZ96, KSP02, KHWT95, KP01, KB06, KP93a, KYB08, Lee97, LDCO08, LWY96, LMS04, LYZ90, LPZ98, LRG99, LXL08, LWC^{+}09, LOSW99, LCL03, LH03, LNOZ03, LJW^{+}07, LWP07, LC02b, MGZN07, MY07, MB07, MZ05, MM98a, MS03, MKY^{+}09, MQ97, NO98, NOS99, NO00a, NOZ01, NOZ02, NSU97, Par95, PH96, PPR99, Par01, PM02, PDC94, Pre99, PH02, Raj05, RSS90]. Efficient [RE09, RJ90, SS96, STY09, SVP08, SJPL08, SO95, SZXS05, SJM09, SP95, SCP99, SPS98, SKPS01, ST93, SW92, TGV08, TSK06, TCR96, TD01, TS08, TC95a, TWH99, WW92, WHW05, WXLZ06, WWL06, WLZ08, WMWL08, WSG01, XUAS99, XL96, XH08, YL07, YLL^{+}07, YWD08, YK03, YV98, YYS97, YL96, YC96, ZH05, Zia93, dB98, AM91, CC93b, CCCS90, CAB93, Cor92, Gab90, KN95, LG94, LC91a, MS93, MM96].

Effort [HY07]. EIC [Bhu09a, Yew06]. EKMR [LCL03]. Elastic [KSP02]. Election [CC93a, DB08, DIM97, NO02, Sin96, YK99, AAG94]. Electrocardiogram [JNGS06]. Electronic [LZ05]. Element [LC99]. Elementary [ADD^{+}02, CHC04]. Elements [PKL06]. Eligibility [LMS04]. Eligibility-Based [LMS04]. Eliminate [PW95]. Eliminating [GP99a, NSD^{+}91]. Elimination [Agr98, ABK98, CY99, FRGJ07, MGA^{+}09, SSZ02, Sto04]. Elimination-Based [SSZ02, Sto04]. Embedded [BB05, CLS04, FDC00, GVV09, JNGS06, KHM05, KB06, KMW08, LA04, MZ05, YW98, ZBM09, Tak93]. Embedding [Ano99h, Avr99, BS96, FLJ05, GW06, GM94, HS97, JHK97, LC96b, LH05, LC01, SBS98, SX08, TCS97, Wan08, YR96, CARW93, CL93, MS94a]. Embeddings [FJL07, GS95, dBL98]. Emphasis [GMCB01]. Empirical [SLY90, DF97]. Employing [ADG06]. EMPOWER [ZN04]. Emulation [WLZN07, ZN04]. Emulations [OHRW99]. Enabled [BB08, CKK^{+}04, LLY04, MSM06]. Encoding [SPS98, THH96, RJ94]. Encoding/Decoding [THH96]. End [ASB02, JTC08, KCD07, KMW08, LWK05, SF07, SS07, WJLK07]. End-Host [SF07]. End-Systems [ASB02]. End-to-End [JTC08, KMW08, LWK05, SS07, WJLK07]. Energy [AD08, CHA07, CKK^{+}04, CTF09, DZ04, DKKS04, FHA06, FLP^{+}07, GVV09, GY07, ISRS06, KA09, KSME08, KMW08, LLTW08, LDCO08, LWC^{+}09, LRS02, LH06b, LWP07, MGZN07, MY07, MZ05, NO00a, NOZ01, NOZ02, SJPL08, SCO^{+}07, TM06, TGV08, TSK06, TSRS07, WMWL08, WCD08, YK03, YZC08, ZS09]. Energy-Aware [AD08, GVV09]. Energy-Efficient [DZ04, LDCO08, LWC^{+}09, LWP07, MGZN07, MY07, MZ05, NO00a, NOZ01, NOZ02, TGV08, WMWL08, YK03].

Energy-Limited [FHA06]. Energy-Oriented [YZC08]. Energy-Time [FLP^{+}07]. Enforced [BCdSFL09]. Enforcement [MTL95]. Enforcing [LW09a, TF96a]. Engine [SA09]. Engineering [SY07]. Enhance [OHRW99, XL04]. Enhanced [BGO^{+}98, BGOS97, HCHM09, KK03b, MZA02, SM03, BGO^{+}97, KS94]. Enhancing [AA09, CLY08b, CK96, LK07, RD09, ZH06]. Ensure [WT08]. Ensuring [KK03a, QR07]. Enumeration [BDL95]. Envelope [CW02b]. Environment [BA04, DS02, DvdMK09, Gon03, KKGS01, KWH02, LMT98, LC02b, MOFD05, MROD07, RRFH98, SGB08, SkLC^{+}03, CD94, DY93, GG94a, LHS92, RK94a, SM94]. Environments [AJF96, AKSS04, CLY08a, EDO06, EVW07, GRS99, GN06, HS99b, KA06, KW08, PF08, SVM07, SWH98, SB04, TCO01, WDCK04]. Epidemic [GKG06]. Epidemic-Style [GKG06]. Equations [BAH01, KBD08, MBM98, CARW93, You93]. equivalence [WY94]. EREW [Che95a, PDC94]. Errata [Ano02c]. Erratum [Ano99h]. Error [ANKA99, KLS00, MBW02, MTM02, SM97, WFP90, XB98, HISS94, JF94, TH93, VJ94]. Error-Correcting [KLS00, XB98]. Error-Detecting [SM97]. EST [KABK03]. Establishing [SCK00]. Establishment [ZS95a]. Estimates [MF01b, TEF07]. Estimation [KPR05, MRT09, SVM07, TSRS07, YW98]. Estimators [BCVC05]. Ethernet [BDS94, FYP07, KS03, KgCS04, WR04]. Ethernet-FDDI [BDS94]. Euclidean [CPhX04, LS96, LHS03, WH03a]. Evaluate [LZTY09]. Evaluating [ATML08, DAF95, HW08, JW00, MSH00, RFDS97]. Evaluation [ANKA99, ABS01, BT00, CLB08, CV92, CLJ^{+}04, DS96, FS00, Fei05, HS99a, HXA96, IT93, KKCB02a, KKCB02b, KWOA05, KHS07, LEH92, LJZA04, LB00a, LR97, LAS04, MKR00, MMSM06, NGM97, PP96, PPR95, PK04, QNR99, SFP03, SH96, SLEV03, SRD08, TCO01, VDS99, WM95, XTL06, YD94a, YZC08, ZYC95, AMAM94, BCBzC92, DF97, EMS90, HC92, HK93, ICT93, KG92, LG94, SH94, Var93, YC93, YD94b, ZY95]. evaluator [SR91]. Even [Chi00, cFC98, Pad91, RS90]. even-sized [Pad91]. Event [AJF96, CK96, NL02, XC04, ADM92, HMW93]. Eventual [MRT06, WCR09]. EveryWare [WBO^{+}01]. Evil [AS00]. Evolving [SZ03b]. Exact [AV96, HH95, PF96]. Example [Abr97, LBS05, PK95b, BCBzC92]. Exception [XRR00]. Exchange [DD98, DD01, SY00, SJPS01, TLGP97, YW00, YW01, BCH94, Pad91]. Exchanged [Che07, LHP05]. Exclusion [AEA97, AMP07, CS01a, CH09, FT97, HY05, HS98b, JK99, Jou03, KKM08, KM01, LK00, RRRM09, TYK99, BCBzC92, HMR94, IK93, NLM90, Sin92]. Executing [FB01a, GVGD95, WW92]. Execution [Abr97, AKSS04, CF00, CY96a, DHN96, DÖ02, HÖ99, HCF03, HCY97, KL01, KPR05, MGDZ07, MT97, PH02, TSAL97, WSB09, ClW91, KK93a, KM91, MLS94, RK94a, RK94b, RM90, Uht92, WCS92]. Executions [MJRS06]. Existing [dLCK^{+}05]. expanding [JS93]. Expansion [dBL98]. Expansive [CMR07]. Expected [WWWA09]. Experience [CSR^{+}09, DCSM96]. Experimental [BCJ90, Fei05, HS99a, KKCB02a, KKCB02b, NN96, PK04]. Experiments [GMR98]. Expiration [TC04a, TC06]. Expiration-Based [TC04a, TC06]. Explicit [YL08]. Exploit [RSP02, WX07, YZZ00]. Exploitation [PLT00]. Exploiting [AGGD04, AK98, CW06, HT06, LCB00, LL90, LWP07, MA01, Pre99, RSB97, RM90, RH00, TLM04, TT94]. Exploration [CL05, MCG08]. Exploring [CC03, CH04a, KYD^{+}07, PC05, SP07]. exponential [MM96]. Exposure [ZZMN07]. Expression [CT97]. Expression-Based [CT97]. Extended [DW04a, KP92, Sca99, Wu97a, Wu00, Wu02, WCDY06, YJ97a, ZMMS08, LH93, jTM97, VGGD94]. Extending [FPGAD08]. Extensible [Din06, RFDS97]. extension [FD94]. Extensions [UZCZ97]. Extraction [CTF09, JNGS06, GO93, GP92].

F [Ahu93]. F-channels [Ahu93]. Fabrics [HDF07, Tze04]. Factor [GZ09]. Factorization [FJY98, GKK97, KBD08]. Fail [CD08]. Fail-Stop [CD08]. Failure [DÖ02, FCF00, HS99a, HHM^{+}00, KHM05, LL02, PS96c, SCY96, WYWZ08, ZS95a, ZLKK07, MP91]. Failure-Detection [HS99a]. Failures [CD08, CS96, Par95, RCS01, Sin96, SS07, TCS97]. Fair [DVV07, HS08, IKOY02, KSP02, LMS04, LK00, MEKOT03, TB94]. Fairness [AMY09, JS98, Kar01]. Faithful [GG09]. False [KCRB03, PW95]. Families [TH01]. Family [BLD05, CL97, cFC98, GY95a, Kop96, Tak93, OSZ92, VS96, Zia94]. FAN [AV96]. FAN-IN [AV96]. Farewell [Bhu09a, Yew06]. Farm [WSC97]. Farms [DR98]. Fast [AD95, BC06, BLO^{+}94, CLPT02, DSO02, DCSM96, GVV09, Hsi03, JK99, LZ02, LO95a, LPZ98, MM96, PJC93, SP95, SZ04, TC98, WM93, WH03b, ZY07, ABDZ94, BCBzC92, CH92, ZA92]. FASTEST [KA99]. Fat [CMDP09, MKY^{+}09, RRRM09]. Fat-Tree [CMDP09]. Fault [AOK09, AB99, AM95, AMPR01, Ano98b, BMR99, BHL^{+}07, BC99, BCH94, CYW08, CL93, CLJ^{+}04, lCL95, CC01, CD08, CXP09, CC98, CCD^{+}09, DDY99, DC98, DAA97a, DAA00, DAMK06, DY05, Dua97, FD94, FPGAD08, FIMR01, GY95a, GMM97, GN96, GMCB01, HÖD99, HY99, HDF07, Her00, HCH99, HL09b, JZXX99, KIBW99, KH04, KLC97, KH97a, Lan95, LDCO08, LH06a, LLGS09, LHSML95, LH03, MGDZ07, MM98b, MJRS06, MBM98, OS94a, OS94b, PG07, RO99, RST95, RRRM09, SyFL99, SCP99, SB04, SDDY00, SNI02a, SNI02b, SLH97, TJ07, THH96, TL06, TB94, TCS97, TH01, VDS99, WC09, WMWL08, Wu98, WA99, Wu00, Xia01, YJ97a, YJ97b, YDW^{+}09, ZS98, dB98, AM91, BS95, BP94, CS90, Chu96, GMG96, KK93a, LG90, MN92, OC93, Rao96, RJ94, SB94a, SM94, Tze93, TC94, VJ93, VJ94, WF94]. fault [YZW94]. Fault-Aware [LLGS09]. Fault-Containing [LH03]. Fault-Free [HCH99]. Fault-Local [DAMK06]. Fault-Resilient [AOK09]. Fault-Tolerance [GMM97]. Fault-Tolerant [AB99, AM95, Ano98b, BMR99, BC99, CYW08, lCL95, CC01, CC98, CCD^{+}09, DDY99, DY05, Dua97, FIMR01, GY95a, GN96, GMCB01, HY99, JZXX99, KH04, KLC97, Lan95, LDCO08, LH06a, LHSML95, MM98b, MJRS06, MBM98, PG07, RO99, RRRM09, SCP99, SDDY00, SNI02a, SNI02b, THH96, TCS97, TH01, VDS99, Wu98, WA99, Wu00, Xia01, YDW^{+}09, ZS98, dB98, BCH94, CL93, FD94, OS94a, OS94b, RST95, TB94, BS95, CS90, KK93a, LG90, SM94, Tze93, VJ93, VJ94, WF94, YZW94]. Faults [CBE93, CC01, NT09, RCS01, SCY98, KA94]. Faulty [Ano99h, Avr99, CCP95, CT97, CH01, Fu05, GP99b, HCH99, JHK97, KY98, LC01, PKL06, SR98, SX08, TW00, YR96, TR93]. FC3D [RLD03]. FDDI [BDS94, KZ96, SZ95a, ZS95b]. FDDI-Based [KZ96]. FDDI-M [SZ95a]. Feasibility [WR04]. Feature [JNGS06, GO93]. Feedback [FZGC06, LWK05, LLA^{+}06, PC07, SC05, SS90]. Feedback-Based [PC07, SC05]. Feedforward [EAK97]. Fei [YYX^{+}09]. Ferry [ZH07c]. Fetching [WB98]. FFT [GK93, Har91, SBF00, TH93]. fiber [AAG94]. fiber-optic [AAG94]. Fibonacci [Hsu93, JHK97, Sca99, Wu97a]. File [FV09, FBD96, HY96, JO95, LYW08, MMJ03, NKP^{+}96, SJKC06, WX07, AGE94, BL91, KE90]. File-Access [NKP^{+}96]. Files [DP02, HZ97, KA06, PM02]. Filling [AB07]. Filter [LH93, TSP^{+}08]. Filtering [LKK02, LZR09, SX03, WH03b, SMJ92]. Filters [WH01]. Find [XZG09]. Finding [HNO98b, LH03, MNS97, Wan98, Wan04, CF94]. Fine [Ksh03, PKJ97, RH00, RH04, Sun02, ZF07, DAF95]. Fine-Grain [RH04, Sun02]. Fine-Grained [Ksh03, PKJ97, RH00, ZF07, DAF95]. Finite [GLS07, LKK95, LC99, SKB04, TK96a, MD96]. Finite-Buffered [GLS07, MD96]. Firewall [LG08, LG09]. Firm [Ram99]. First [BMR99, PWW00, SVP08, CS90]. First-Fit [BMR99]. Fit [BMR99, KY98]. Fixed [EF95, cFC98, OPZ99, WMWL08, PN93]. Fixed-Degrees [cFC98]. Fixed-Priority [WMWL08]. Flat [TC04b]. Flexible [DSY99, FFPF05, HKS^{+}07, LDCO08, SDFV96, TL06, XZG09, RFDS97]. Flexible-Schedule-Based [LDCO08]. Flexibly [PH05]. FlexiTP [LDCO08]. Flip [CBM^{+}07]. Flip-Based [CBM^{+}07]. Floating [ZP07]. Floating-Point [ZP07]. Flooding [DP06, LJW^{+}07, SL01a]. Flooding-Based [DP06]. Floods [SWWJ08]. Floor [BRSS08]. Flow [AAS03, ANKA99, BÖ98, BJM^{+}05, CS97a, CY00a, DDY99, DF99, FYJ^{+}09, LL06b, LMN95, RLD03, AN94, Bok93, Dal92, EG93, KGS94, MS94b, NSD93, SMS93, TB93]. Flow-Based [FYJ^{+}09, LL06b]. Flows [JXT^{+}04, LW09a, ZMRS08]. Floyd [MF96]. Fly [KS06, MRT09, PK00]. Fold [YW03a]. Folded [DCF95, ÖD96, EAL91, KS94]. Force [LW09c]. Forests [VRKL96]. Fork [Che01, LMT98, KS93, TRS90]. Fork-Join [LMT98, KS93, TRS90]. Fork/Join [Che01]. Form [Bar98, ME95]. Formal [GT02, PD00, WP00]. formalization [AH93]. Format [EBS02]. Formation [BMPP06, DW04a, LSW04, WWL06]. Formulation [PK01, KSA94]. Fortran [SLY90]. Fortran/HPF [UZCZ97]. Forward [Dua96, MTM02, WYD07]. Forwarding [BSCB09, KCD07, WCBX06, WLHB08, YL08, KCPT96]. Four [CL97, CH95, WMN99, AH93, VS96]. Fourier [FA94, ZA92]. FPGAs [ZMP07]. Fragment [MMJ03, SY93]. fragmentation [NSD^{+}91, YW93]. Fragments [Men05]. Framework [Agr99, BCCP04, BF04, BC96, CAZ04, DLS09, FS00, GAL01, GAG96, GSS96, KCS^{+}99, KCRK00, KCRB03, KLC97, KyK09, KPBD09, LK07, LL07, LWP07, MAS08, MYA01, PNZ^{+}02, PK95a, RSB97, SBF00, SKCL09, SA94, THH08, Yi09, YR06, ZCO98, vDSP96, EHJ94]. Free [BC96, CBD^{+}01, Dua95a, Dua95b, Dua96, DP01, DLPP05, GPST09, GY09, HCH99, JKA07, Kuc01, LYW08, LPD05, Mic04, MRT06, PPD03, SGB08, SL01a, TW00, VS11, WWWA09, BR91, CS94, DA93, Dua93, GPBS94, HM92, LMN94, PGDS94, PGFS94, PN93, SC93]. Free-Riding [LYW08]. frequency [ADM92]. FRoots [TL06]. FT [RRRM09]. FTPA [YDW^{+}09]. Full [CCP95, FRGL09, MT97, PS96a, RO99, LC94]. Full-Information [FRGL09]. Fully [LBS01, MBTPV06, PGFS94, RLD03, TW98, vdMDM07]. Function [WR04]. Function-Driven [WR04]. Functional [AGWFH97, CE95, PAM95, YA93, GP92, MR94]. Functions [HHM^{+}00, LBS05, GG94a, MM96]. Fundamental [DZH05, Sah00b]. Fusing [FZVT98]. Fusing-Restricted [FZVT98]. Fusion [MA97, SvVB05, JWC94].

G [XPL04]. Gabriel [WY07]. Game [BHL^{+}07, GBD07, KA09, KHS07, SZ08, YM09, YK09]. Game-Theoretic [KHS07, SZ08]. Games [CHL09]. gamma [Chu96]. Gang [WF03, ZFMS03]. Gang-Scheduling [ZFMS03]. Garbage [CRN09, KMW95, MJ06, RKHM06, SNI02a, SNI02b, HM92, IT93]. Gathering [LRS02, MY07, YKP08, ZS09]. Gaussian [ABK98, PH96]. Gemini [CFB02]. General [Agr99, BF04, CM95, CCY96, DP01, FF98, HMR94, PK95a, RS97b, SM97, YJHG06]. Generalization [PZLS01]. Generalized [Chu95, DFKS01, EAK95, FE97, HPT04, HCYD01, JHK97, LKKS05, LL06a, LL06b, MC95, OC93, PM96, UEA95, WCY95, CA93, FC91, ME92, ME93, SKF94, SB94a, ZL96]. Generated [TEF07]. Generating [BI95, MQ97, MM96]. Generation [FI95, GAK03, LF03, TG99, ZSMF01, Fos91, MCH^{+}90, SSG91]. Generational [MJ06]. Generic [PABD^{+}99]. Genetic [CFW98, CFR99, WYJ^{+}04, ZWM99, ZT01, ZW02, HAR94]. Genetic-Algorithms [ZW02]. Genome [WZZ09]. Genomic [JTP^{+}08, MDL06, SA09]. Geo [WLHB08]. Geo-Forwarding [WLHB08]. Geographic [XLPH06]. Geographical [CW06, FG06b, SvVB05]. Geometric [ALW^{+}03, CL09, KH97b, LW09c, Che95a]. Geometry [LOSW99, wJNPS97, ZA92]. Given [CM95]. Givens [MBM98]. Global [BNBH^{+}95, BCL09, BDD^{+}96, CLJ^{+}04, DGFHR03, DvdMK09, HHM^{+}00, Ksh03, LT97, MD97, MNS97, NX95, OXL06, PC05, TAKB06, TLM04, XL04, GG94a, KM91, jTM97]. Global-Scale [DvdMK09]. Global-State-Triggered [CLJ^{+}04]. Globally [AJF96]. Globus [CSR^{+}09]. Gnutella [ZH06]. Gnutella-Like [ZH06]. Goal [CV08]. Goal-Oriented [CV08]. Going [PW95]. Gossip [ST99a, ZBM09, HJB^{+}09]. Gossip-Based [HJB^{+}09]. Gossiping [Gon03, JSR98, LZ02, Rav07, LR93]. GpH [ATML08]. GPUs [LSVMW07]. Graceful [YJ97b, HW91]. Gradient [GVV09, GKS95, LCN^{+}07].

Gradient-Based [GVV09]. Gradually [LWN98]. Grain [RH04, Sun02]. Grained [AFAGR97, KL01, Ksh03, PKJ97, RH00, ZF07, DAF95]. grammars [KG92]. Granularity [FI95, GY93, MKH91]. Graph [AAD97, ACT^{+}97, BT98, CLB08, Che96, CL97, JJ07, MD97, MS03, OR97, PPP04, PWW00, RRG07, SBS98, TF01, THT^{+}97, TCS97, WMN99, EG93, FA94, LB94, Lat94, MS92, MJ94, Rao96, RJ90, VS96, WC90, YW93]. Graph-Based [TF01]. graph-level [EG93]. Graphics [TSP^{+}08]. Graphs [BDL95, BKS03, COP00, CS97a, CTS96, CH08, CKC08, DD01, DNSC09, GZ09, HY97, HCH99, yH02, Hsi03, HC97, ISAZM09, JK99, KA96, LKK02, LC99, LC01, SWC95, WY07, YV98, ZMM04, dBL98, Cor92, DT94, GY93, Lee91, LR93, LH94, PAM94, Sch91, SS94, VJ93, WY94, YC96]. Gravitational [HJB^{+}09]. Gray [MQ97, ZL96]. greater [HM90]. Greedy [XLPH06]. Grid [DvdMK09, LSZ09, PF08, RD09, WH95, WBO^{+}01, BcFGM08, BWC^{+}03, SVM07, VVR07]. Grid-Structured [WH95]. Grids [AMY09, CCD^{+}09, KA09, SZ08, XZNX08, CC93b, EF96, ATML08, BA07, BGJ06, DVV07, KHS07]. Group [AKNR^{+}04, AMP07, DS03a, DS03b, FB01b, Jou03, KKM08, KM01, LNYY03, LL07, MFLX01, SJd^{+}09, XP07, YW04]. Group-Based [SJd^{+}09]. Grouping [CH08, TKP00]. Groups [STW00]. Growth [GZ09]. Growth-Restricted [GZ09]. GTS [HPH08]. Guarantee [Ram99, XP05]. Guaranteed [DZHG04, KS01, LWXS06, SL01a]. Guaranteeing [MGA^{+}09]. Guarantees [ASB02, FZGC06, HH08, KCK^{+}06, LLA^{+}06, NK08]. GUARDS [PABD^{+}99]. Guest [CRS06, PP05, ACM08, GZ03, ON02, WA99, Zha03, ZH99a]. Guided [ZMRS08].

H [MKY^{+}09]. H-Tree [MKY^{+}09]. Hamiltonian [HCH99, LC01, Wan08].

Hamiltonicity [HL09b, Fu05]. Handle [XCZ04]. Handling [BCQD07, MRLD01, SP03, TCLY07, XRR00, YD94b]. Hard [BMR99, GMM97, HS99b, WMWL08]. Hard-Real-Time [BMR99]. Hardware [CY00b, CMDP09, DDS95, DS96, LLS06, LNO^{+}00, RSV90, XL08, ZY07]. Hardware-Algorithms [LNO^{+}00]. Hardware-Based [CMDP09, DS96]. Hardwired [SH95a]. Harmonic [ZX04, ZCSY08]. HARP [DFD93]. Hartley [AD95, ZA92]. HARTS [SH96, ZS95a]. Hash [HCY97, RHM09, TP95, OL92, WYTD93]. Hashing [DPH08, MD97]. Hazard [Mic04]. HBA [ZJWX08]. Hector [RRFH98]. Height [YCTW07]. Hellinger [SWWJ08]. Helper [LJLS09]. Herd [CB03]. Hereditary [yH02, Hsi03]. HERO [ZLZN09]. Heterogeneity [AD08, LP07, SGL06, WX07]. Heterogeneity-Aware [SGL06]. Heterogeneous [AAD08, AJMJS03, Ano04c, AA09, BA04, BDvD98, BBC^{+}04, BBRR01, BLR03, BLMR05, BGJ06, BP06, CYW08, CF00, CRS06, DR98, DÖ02, GVV09, GLQL09, HC97, KA06, KLH07, KSME08, LZ08, LXL08, SZXS05, TSAL97, TS98, THW02, VM04, XQ08, ZCLC06, ZSLW92, CR94, SL93a]. Heuristic [AMS97, CHC09, PK95a, PK95b, YF97, MS93, SL93a]. Heuristics [EDO06, HÖ00, JSWB97, KA06, TTB^{+}00, GDI93]. Hexagonal [DS05, NSZ02, YL96]. Hidden [JTP^{+}08]. Hide [LLY05]. Hiding [MLW06, SL09]. Hierarchical [CS08, DC95, GD95, HS97, HLL09, JLDC05, KW08, MB94, PAM94, Raj05, RJ05, SFP03, WTCY95, WCR09, YP98, CAB93, CPA93, KP92, ME92, ME93, MS94b, ZY95, Zia93]. Hierarchically [HZ96, SS07, ZH98]. Hierarchy [PHP03, LK94]. High [AGGD04, ATML08, AS96, AAB06, Ano05c, Ano09c, BGMZ97, CE95, CBD^{+}01, CS05, EBS02, FZGC06, FG06a, FLP^{+}07, GRS99, GMCB01, HDF07, HNY02, LLGS09, LBS05, MLW06, MJ98, MNM04, MDL06, ON06, OC05, PGBI03, RK08, RJ96, SS08, SkLC^{+}03, SD00a, SSP02, TCLY07, TGV08, TF96a, WOT^{+}07, WCCR^{+}97, ZMP07, Ant94, AB91b, WS93]. High-Bandwidth [BGMZ97]. High-Latency [GRS99]. High-Level [ATML08, MLW06, RJ96]. High-Performance [AGGD04, AAB06, Ano09c, EBS02, FG06a, FLP^{+}07, GMCB01, HDF07, LLGS09, MDL06, ON06, OC05, PGBI03, RK08, SkLC^{+}03, SD00a, SSP02, TGV08, ZMP07, WS93]. High-Speed [CBD^{+}01, FZGC06, MNM04, Ant94]. High-Throughput [WCCR^{+}97]. Highly [AGGD05, CB00, DAA00, DB08, GKK97, HK94, WL00, WLR93]. HiPER [MBW02]. HIPIQS [SSP02]. HLA [SF08]. HLA-Based [SF08]. Hoc [ALW^{+}03, Ano04d, BK09, BMPP06, BS08, CLW03, CKWC08, DW04a, DW04b, DW06, DPH08, DAMK06, DB08, GJDA06, GYS05, GY07, GS03, ISRS06, JJ07, LCWW03, LWS04, LH06a, LWC^{+}09, LJW^{+}07, NO00b, OSRS06a, OSRS06b, PDH06, SLFW06, TR06, WY07, WO04, Wu02, WCDY06, WD06, WYD07, XP05, YWD08, Yi09, ZL07b]. HOL [MGA^{+}09]. hold [HC92]. Holes [WCD08]. Home [TAKB06]. Homogeneous [Aro00, DNSC09, LS97, LJW05, TGV08, XQ08]. Honeycomb [PK01, Sto97]. Hop [CLW03, DZ04, LJW^{+}07, Liu08, MBW02, NO00a, RHM09, WWWA09, XP05]. Hop-by-Hop [MBW02]. Host [CN02, CN04, Rob04, SF07]. Host-Client [CN02, CN04, Rob04]. Hot [BRS97, LC95, NS95a, OKSA01, WSNA95, ZYC95]. Hot-Potato [NS95a]. Hotspot [YM09]. HPF [JB01, UZCZ97, vDSP96]. HRing [ZCSY08]. Hull [BGO^{+}96, HNO98a]. Hybrid [ADG06, Che01, CJLN09, CKC08, FV09, LP07, LOSW99, MMSM06, SE98, SvAS04, SL01a, SZ04, SJPS01, SS00, WO04, WYWZ08, LHS92]. Hydrodynamic [HC99b]. Hydrothermal [dSF03]. Hyper [GP93, LSBS98, THT^{+}97]. Hyper-Bus [THT^{+}97]. Hyper-deBruijn [GP93]. Hyper-Systolic [LSBS98]. Hyperchannel [CWYZ09]. Hypercube [AD95, lCL95, Che07, CC98, FYS05, FMG02, GVGD95, HS97, KP96, KC98, Lan95, LHP05, LWN98, MR06, PKL06, RTS95, SP95, SV97, WL97, Xia01, dCVGG02, AOB93, BJS90, CS90, DK92, GDJ94, HB92, IS90, JR93, KDL91, KLDR94, KP92, MB94, Nas93, OL92, PGDS94, RS91b, RB90, RJ90, SRT94, SF92b, YW93, YZW94, YN90, ZA93, Zia94]. Hypercube-Connected [AD95]. Hypercube-Derived [WL97]. Hypercube-Like [PKL06]. Hypercubes [Ano99h, Avr99, CCP95, CT97, DPS96a, DPS96b, DCF95, GP99b, HÖ00, HK95, HWKH01, JHK97, KLS00, OKSA01, SR98, SLH97, TW98, TK96b, TC98, YR96, dBL98, AM91, CL93, CC93b, DT94, EAL91, Fid92, KK93a, KS94, KP92, KSA94, LS94b, ÖD96, PGFS94, RS90, ST93, TR93, UEA95, VB93]. Hypercycle [DD95]. Hypercycle-based [DD95]. Hyperedges [LH05]. Hypergraph [BA07, ÇA99, GW06]. Hypergraph-Partitioning-Based [BA07, ÇA99]. Hypernet [HC99a]. Hyperthreaded [SL06]. Hypertool [WG90].

I/O [Bor00, HJH02, JSWB97, KKCB02a, KKCB02b, Kan01, KB03, kLCC^{+}06, OPZ99, RB90, TR04, VV99, YZC08]. I/O-Centric [HJH02]. IBA [KYD^{+}07]. IBM [BGBP01, HXA96, MS94a, MF01b]. IC [CMR07]. IC-Scheduling [CMR07]. ID [BRTM09]. Identical [JR03]. Identification [Che96, CT97, FHBJ97, MLSS07]. Identifier [LQZ09]. Identifier-to-Locator [LQZ09]. Identifying [HP03]. Identity [BRTM09, PZZ09, YK99]. Identity-Based [BRTM09]. Idle [RH00]. IEEE [BCG04, HPH08, JASA08, MGZN07, MSM06, NK08, XL04, XLW^{+}06, ZL07b]. II [KCN90b, LL06b, LPD05, OSRS06b, PK95b, RK94b, YK96b]. Image [BA07, EAF00, JS93, LHS03, SKB04, WS00, WCH^{+}08, Aln94a, CL94, GO93]. Image-Space-Parallel [BA07]. Images [EAF00]. Immucube [PG07]. Immune [SSZ06, ZS95a]. Impact [BIWK00, CH04b, CTF09, CY00a, DMKJ96, Lil94, SG94, SCL05, SSP00, VSD01, XLPH06, ZSMF01, DI95]. Implementation [ATG92, ACT^{+}97, BRSS08, BGBP01, BDD^{+}96, Din06, EBS04, FVR03, JTP^{+}08, JLF03, LAS04, MNM04, MR94, ON06, Pak07, QS03, SKJ07, SBF00, SOM05, TSP^{+}08, WR04, WMXZ06, XUAS99, XL08, YK92, vDSP96, Ahu93, AIK91, HK91, LKG92, LH93, LA93, SMBT90, SMJ92]. Implementations [kLCC^{+}06, PKJ97, PG01, GO93]. Implementing [AGWFH97, BA90, FG01, SSP00]. Implications [CGM^{+}07, HWWX99]. important [KLDR94]. Imposed [PDH06]. Improve [Kin06, SRD04, TT94]. Improved [BKS03, CWCC07, KYD^{+}07, Kla98, Li03, LLS06, LH06b, PZLS01, PPP04, SRT94, KKP91]. Improvement [KA06, LYW08]. Improving [BA04, CK08, GYS05, KK04, KCRB03, KA05, LY93a, LLX06, MOFD05, PH05, SF07, TJ07, TSG09, GS91]. In-Kernel [LBS05]. In-Network [DLS09]. In-Order [WSB09]. In-Situ [MCL^{+}07]. Incentive [TJ08, XZNX08]. Incentive-Based [XZNX08]. including [MM96]. Incomplete [CTS96, CT97, LB94, TK96b, SCD97]. Incorrectly [SCL05]. Increased [PPD03]. increasing [MKH91]. Incremental [OR97, SW96, WYJ^{+}04, YN00]. incrementally [LB94]. Independence [Gen00]. Independent [AAD08, BFL^{+}01, HP07, LH03, PG01, YCTW07, BA90, RK94a, RK94b]. Index [Ano97a, Ano98a, Ano99b, Ano01e, Ano02a, Ano03b, Ano04a, Ano07a, Ano08a, Ano08d, Ano09d, BQF99, Din01, EHJ94, Ano05b]. Index-Based [BQF99]. Indexed [BAH01]. Indexing [ZH07a]. Induced [BBH05, HMR99, TKW98, Tsa03]. Inexpensive [HNY02]. Inference [BBH05, HM98, JTC08]. InfiniBand [ASD04, BC06, BCQD07, LK07, NYD09, LBS05]. InfoBeacons [SC07]. Information [AAS03, Dah00, FRGL09, LW09a, LJW^{+}07, LC04, MZA02, Mit00, SC07, Xia01, ZB09, BFP96, Sin92, SL93c]. Information-Flow [AAS03]. information-structure [Sin92]. Informed [TM06]. Infrastructure [AJMJS03, KIBW99]. Infrastructures [GZ03, SCW07]. Inherent [AH06]. Inherently [PK95a, PK95b, PN93]. Initialization [CLW03, NO00a, NO00b, Rav07, OW91]. Injective [LF03]. Injector [CLJ^{+}04]. Injured [TW98]. Input [CCQ^{+}05, GCCC^{+}04, HS08, MR02, SV97, SSP02]. Input-Buffered [CCQ^{+}05]. Input-Queued [HS08]. Input/Output [GCCC^{+}04, MR02]. Insertion [PK99a]. Inside-Out [SyFL99]. Installment [CWCC07]. Instruction [AGWFH97, AF05, CF01, CC95, EP05, PSGD05, WB98, WSB09, XUAS99]. Instruction-Level [EP05]. instructions [BG90]. Integer [KBC^{+}01, PW95, SK95, TG99]. Integrated [ASS95, BcFGM08, CH07, CG02a, CG02b, RNKZ03, SKCL09, Sol02, SPF99, VKS^{+}09, ZFMS03, GH93]. Integrating [GAL01, TCC05]. Integration [AGGD04, HYP02, LBS05, Mha09]. Intel [FBD96]. Intelligent [SX03, WCBX06]. Intensive [EK95, KKC^{+}05, ON06, OXL06, XCZ04]. Interaction [HC97, JS98, LJCL08].

Interactions [WL08a]. Interactive [KMT91, RNR^{+}03, dB98]. Interagent [MX03]. Interconnect [BB05]. Interconnected [QM97]. Interconnection [CFB02, CL97, DC98, DAA97b, DD98, FR96, cFC98, GS95, HSWB07, HP03, Kop96, Lai00, LKK02, LR97, LSC95, LWN98, LK04, PR05a, PKL06, RO99, SS96, SPS98, SP07, SDFV96, SCL00, VDS99, WL97, WP00, WL00, XP07, YN00, YFJ^{+}01, AV94, Aga91, BDS94, CAB93, CI92, CO94, Chu96, HC92, Hsu93, KP92, LS94a, LC94, MB94, MR92, MJ94, MD96, Sch91, SL93a, VS96, YM95, Zia94]. interconnection-constrained [SL93a]. Interconnections [FG06a]. Interconnects [ADG^{+}08, HP06, PSGD05, YW03b, YW05a, ZY04, ZY06]. Interest [CLY08b]. Interface [DHN95, DFKS01, WOT^{+}07]. Interfaces [ZLKK07]. Interference [BPT03, WWLS08, YY95]. Interference-Aware [WWLS08]. Interleaved [HDF07, LS94b, SL94]. Interleaving [CY92, KHY09]. Interlocking [OZ96]. Intermediaries [KYB08]. Internet [GSS06, HY07, LKKS05, LA06, LQZ09, Sun02, SX03, TC07, YWF^{+}09, ZYKG07]. Internet-Based [Sun02]. Internet-Scale [ZYKG07]. Interoverlay [LJLN07]. Interpreters [AGWFH97]. Interpreting [Dah00]. Interprocedural [Agr98, Agr99, CHJL04, CY00a, HK91]. Interprocess [KB03, RSV90, TB94]. Interprocessor [KL99, PH04, SO95, GR90]. intertask [SS94]. Interval [FCF00]. Intervals [RRRM09, OSZ92]. Introduction [ACM08, ABC01b, Bhu09a, ON02, WA99, Yew06, ZH99a]. Intrusion [RNKZ03, MRW92]. Invalidation [TCO01]. Inverted [JO95]. investigating [LH94]. Invisible [YWF^{+}09]. invocation [BA90]. IP [ADG06, GS08, GWYS08, LBC03, SX03, WS03, WMXZ06, XZG09]. iPAK [MCL^{+}07]. IPC [SS08]. IPS [MCH^{+}90]. IPS-2 [MCH^{+}90]. Irregular [HT06, JKA07, KP01, LCB00, LSRT06, PSC^{+}95, PH02, QNR99, SD00a, SD00b, SKPS01, TW00, UZCZ97]. Irregularities [HP03]. Isolated [ZS95a]. Isomorphism [Che96, HWSH00, WMN99]. Isotach [RWW97]. ISPs [LJCL08]. Issue [AGWFH97, Ano97d, Ano97b, Ano97c, Ano98c, Ano98b, Ano01b, Ano01c, Ano01d, Ano02b, Ano03c, Ano04c, Ano04d, Ano05c, Ano08c, Ano09c, Ano09b, DF99, Ano99g, Ano07c]. Issues [AS96, TL05, VMXQ04, ZWM99, LY93b]. Items [OPZ99]. Iteration [GAK03]. Iterations [KGKL08]. Iterative [BCVCV05, BCVC05, BG90, KA06, Lee95, LRRV04, YF97, dLCK^{+}05, AH91, AC92, EG93, Pan93]. Iterative-Improvement-Based [KA06].

Jacobi [KGKL08]. January [Ano99g]. Java [CKK^{+}04, CS03, MJ06, SM02, YLL^{+}07]. Java-Enabled [CKK^{+}04]. JEWEL [LKG92]. Job [AAB^{+}00, AM06, CV08, CB03, DvdMK09, FFPF05, GBD07, KLDR94, LC91b, SP98, ZA93]. Jobs [BGJ06, HJS^{+}06, KC98, XCZ02, XCZ04, XQ08, KGM96, KS93]. Join [Che01, CST02, CY96c, HY01, LR96, LMT98, TP95, CY92, KS93, NM92, OL92, TRS90, WYTD93, WDY93]. Joins [HCY97, SY93]. Joint [BB05, KA09, LWXS06, SKJ07, WWLS08]. JSQ [LR96]. Just [YLL^{+}07]. Just-in-Time [YLL^{+}07].

k-ary [SG94]. k-Dimensional [CWCC07]. k-splitting [XB93]. Kernel [LBS05, MS94a, ABDZ94]. Kernels [NN96]. Kestrel [DDD^{+}05]. Key [AKNR^{+}04, EP05, MCL^{+}07, STW00, XH08, YGE06, YG08]. Knapsack [AR97]. Knots [BT98, MS03]. Knowledge [LHL^{+}08, TLM04, YG08, MLL92]. Known [XCZ02].

Labeled [WCL97, WY94]. Labeling [BBH05, Aln94a, D’H92]. laboratory [BEK^{+}93]. ladders [PN93]. Lambda [BcFGM08]. Lamport [BCBzC92, JK99]. LAN [LJZA04, LWY96]. Language [ATML08, ABJ^{+}93, Pak07, GR94, JWC94, NSD93]. language/compiler [NSD93]. Languages [Ano97d, Ano97b, Ano97c, BT00, CE95, PG01, WMB96, MR94]. LANs [BCG04, NK08, XLW^{+}06]. LAPI [BGBP01]. Large [Agr99, AM99, BG09, CY00b, CASM07, DS03a, EDO06, FT97, GMCB01, GP99b, HL09a, HS98b, HZ97, KMG03, KCW09, LCGC07, LC95, LK04, MY07, MA01, MMJ03, MDL06, OXL06, PM02, RD98, SkLC^{+}03, ST99a, SGL06, VVR07, XHYL05, XCZ04, ZJWX08, dB98, CO95, CTC93, EA93, OS94a, SG93, YTB92]. Large-Scale [BG09, CY00b, EDO06, GMCB01, HL09a, KMG03, KCW09, LCGC07, LC95, LK04, MY07, MA01, MMJ03, SkLC^{+}03, VVR07, XHYL05, SG93]. LARPBS [CPhX04]. Latency [Agr99, GRS99, KK03a, MROD07, PBA03, QM97, LNP94]. Latency-Aware [MROD07]. Latency-Tolerance [PBA03]. Latin [KP93b]. LaTTe [YLL^{+}07]. Lattice [TG99]. Lattices [FHBJ97]. Lawler [GRT97]. Layer [XZL05, ZCXF09]. Layered [LSRT06, ZL07a]. Layout [BG02, KK04, PHP03, CAB93]. Layouts [CLPT02, CL00, KCS^{+}99, LC96a]. Lazy [MKH91, SNI02a, SNI02b]. Leader [DB08, DIM97, NO02, Sin96, YK99, AAG94]. Leadership [MRT06]. Leading [OB00]. Leakage [ZB09]. Leapfrog [WHC03]. Learning [MR02]. Length [BBD00, VB93]. Lengths [FJL07]. Level [AGGD05, ATML08, ANKA99, CB05, DCF95, EP05, GY95b, HC99a, MLW06, RJ96, SKB04, SLT03, SZ04, WZP^{+}03, XRY09, ZCL04, BGM94, EG93, Lar93, ME92, ME93]. Levels [Wu00]. Leveraging [BRTM09]. LFSR [CCSC09]. Library [BBC^{+}95, LB00a]. LID [NYD09]. Life [SZ03a]. Lifetime [LWJ06, TX08, ZS09]. Lifetime-Constrained [TX08]. Lifting [TSP^{+}08]. LightFlood [JGZW08]. Lightweight [CY06, EBS04, ZBM09]. Like [BK09, LYW08, PKL06, XZNX08, ZH06, Pan93]. Limited [AS00, AM06, CBM^{+}07, FHA06, GY09, LSW04, PH04, ZY04, ZY06, FHRT93]. Limits [Aga91]. Linda [BS95, GT02]. Line [ANKA99, Bir93]. Linear [AAD08, CHC04, DSO02, Gre98, HWKH01, HCD97, KCS^{+}99, KBC^{+}01, KBD08, LPZ98, LLL09, MBM98, PK99a, VM04, WNKS96, WHW05, ZL08, ZLP09, AC93, EHJ94, IA95, KST94, Lin93, NJ94, O’H91, Pan93, ZL96]. Linear-Complement [HWKH01]. Linearization [MF96]. linearly [GDJ94]. Lines [NE01]. Link [CWLR09, hKY08, Sin96, THH08, TCS97, WWLS08, YW03b]. Link-Disjoint [YW03b]. Link-State [THH08]. Linked [LWN98]. Links [Add97, BV05, LWC^{+}09, SCY98, SX08, Wu02]. List [Ano99a, Ano00a, Ano01a, Ano03a, Ano04e, Ano05a, Ano06, Ano07b, Ano08b, Ano09a, FT97, HS98b, PKJ97, WL08a, RJ90]. List-Based [FT97, HS98b, WL08a]. lists [SH95b]. Little [CC99]. Live [BSS09, DF09, GLQL09, LJLN07]. Lived [STY09]. livelock [GPBS94, PGDS94]. livelock-free [GPBS94, PGDS94]. LMSR [SKK01]. Load [BCVC05, BCCP04, Bar98, BBR07, CWCC07, CT08, CHHC06, CK02, Dah00, DPS96a, DPS96b, DHB01, DP02, DHP^{+}07, DB06, DvdMK09, DW03, GZ06, GZ09, GO93, GKK05, GB06, HC99b, JJ09, LRRV04, LL06a, LL06b, Li03, LC99, LJW05, Mit01, PH05, SS08, SVM07, SX07, SH96, SRL98, SZ08, TP95, Tse09, WT98, Wu97b, ZRS^{+}05, ZMRS08, ZH05, ZT01, AT07, Bok93, GT93, GDI93, KK92, LY94, LK94, SH93, SH94, WLR93]. Load-Balanced [CHHC06, GZ06]. Load-Balancing [GZ09, LRRV04, LC99, SX07, ZT01]. load-dependent [AT07]. load-sharing [GDI93]. Loads [BCL^{+}05, CG08, VM04, YvdRC05]. Local [BT98, CBD^{+}01, DAMK06, HT07, KM01, KAY^{+}06, LWS04, MD97, PC05, WSG01, Xia01, PAM94]. Local-Spin [KM01]. Locality [CW06, HT06, KK04, KCRK00, KBC^{+}01, KCRB03, MA97, PLT00, SX07, TSG09, VKS^{+}09, YZZ00, ZH99b]. Locality-Aware [SX07]. Locality-Conscious [VKS^{+}09]. Localization [HCHM09, KS08b, KSP09, SRZF04, TN08, WWWA09, XCZ08, YWF^{+}09]. Localized [Ano04d, BMPP06, DW04a, GY07, LCWW03, LSW04, LH06a, MGZN07, OSRS06a, OSRS06b, SLFW06, SL01b, ZPY06]. Localizing [CS96]. Locating [DS02]. Location [CSR^{+}09, FCF00, LXL^{+}05, PM02, SL09, SZ03b, WLHB08, XPL04, XTL08, YGE06, BA90]. Location-Aware [YGE06]. Locator [LQZ09]. Lock [GPST09, HM92, JH97, Mic04, And90]. Lock-Free [GPST09, Mic04, HM92]. Locking [Sun02]. logarithm [MM96]. Logarithmic [EF95, WYD07]. Logging [ADG06, GS08]. LogGP [Ian97]. Logic [LLJ^{+}93, LNOZ03, PG01, RSP02, RJ99, ClW91, CR90, RK94a, RK94b]. Logical [FMG02]. LogP [DCSM96]. LoGPC [MF01a]. LOMARC [SL06]. Long [Kuc01, SX08]. Longest [WY07, LL94]. Lookahead [SL06, LL90]. Lookup [CHHC06]. Lookups [FRGL09, Tze06]. Loop [COS00, DY05, FLVG95, GMG96, Lar93, OD93, RJ96, SL01a, WL91, DR94, Gup92, LK90, Lil94, ML94, SKF94, SC91, SC93, TN93a, WW92]. Loop-Free [SL01a, SC93]. Loop-level [Lar93]. Loops [AKN95, CY96a, CY99, HCF03, Lee95, MA97, RSP02, RR02, RP99, TKP00, XC01, AH91, D’H92, GMG96, KM91, KS91, ST91, Uht92, WW92, YJZ97]. Loosely [LJLS09, XL96, TKT92]. Loss [KS01, WLL^{+}07]. Losses [MSM09]. Lossless [MNM04]. Low [FPGAD08, GvG06, GMCB01, KA99, LNP94, NE01, OC05, PS96c, RvG02, SKJ07, SKB04, TF96a, THW02, WWL06, WCCR^{+}97, XXZ03, YV98, dBL98, AB91b, BL91, Kum92, MS93, NZ95]. Low-Bandwidth [NE01]. Low-Complexity [KA99, THW02]. Low-Cost [GvG06, GMCB01, OC05, PS96c, RvG02, WWL06, XXZ03, BL91]. Low-Degree [YV98]. Low-latency [LNP94]. Low-Level [SKB04]. Low-Memory [WCCR^{+}97]. Lower [GW96a, JR94, SF92a, SRT94]. Lower-Dimensional [GW96a]. [lozenge]P [FMR07]. [lozenge]S [FMR07]. LRED [WLL^{+}07]. LRPD [RP99]. LRU [LWY96]. LRU-Based [LWY96]. LU [FJY98, KGKL08].

m [ME92, HZ97, SZ95a]. m-level [ME92]. Machine [Bor00, Cha96, RK94a, RK94b, SKB04, YL96, AT07, FC91, MR92, SR94, AS92, SM02]. Machine-Based [SKB04]. Machines [DA98, PKJ97, RvG02, SZ95b, TN08, YF97, YD95, GD94, LC91a, NSD^{+}91, RS91a, TB93]. Macro [YV98, AM93, PAM94]. macro-dataflow [AM93]. Macro-Star [YV98]. Macroscopic [LJW05]. MAD [NN96]. MAGIC [GD94]. Main [TP95]. Maintenance [LXL08, LBS01, TSK06]. Making [NE93]. Malicious [MSM09]. Manage [KKGS01]. Management [BIWK00, lCL95, CY06, HDRS00, KHY09, KSME08, KMW08, LLS06, LP07, Ram99, SF08, SBK02a, SBK02b, SJd^{+}09, SY07, SYC03, SRD08, SZ03b, SSsLY03, TC04a, TC06, VV99, XPL04, XZL05, YGE06, YG08, ZCL04, ZJWX08, JS90, LEH92, NSD93, RST95, TT94]. Managing [MZT08, RD98, US04, SB94b, WYTD93, WDY93].

Manchester [BG90]. Manets [AMH08, LW09c, STY09, WLHB08, WCR09]. Many [Ano09b, BRS97, CC97, PKL06, KST94, RWF94]. Many-to-Many [BRS97, PKL06]. Map [KS08b]. Mapping [AB07, AB03, BB05, CM95, CSR07, DPS96a, DPS96b, EAK97, HWKH01, HCYD01, HW08, LK90, LRRV04, LQZ09, RRG07, YLL^{+}07, CC93b, CA93, IS90, KN95, MS94a, SF92a, ST91, SA94, Zia93]. Mapping/Interconnect [BB05]. Mappings [LF03, DS94]. Mar [ME93]. Margin [HY07]. marked [WY94]. Marker [HM98]. Market [FLZ09, XZNX08, MLL92]. Market-Like [XZNX08]. market-propagation [MLL92]. Marking [ADG06, GS08, PC07, XZG09]. Markov [HN93, JTP^{+}08, LL96, MMSM06]. Martini [WOT^{+}07]. Maskable [WL97]. MasPar [ACT^{+}97]. Massively [CFW98, JTP^{+}08, LWN98, NGL94, YFJ^{+}01, GMG96, HISS94, LC91a, MB94, RJ94]. Master [BBC^{+}04, BLR03, KA06]. Master-Slave [BBC^{+}04, BLR03, KA06]. Match [DP02]. Matching [ACT^{+}97, BM00b, DÖ02, HL09b, Sto96, PDC94]. Matchmaking [SL06]. Mathematical [TTB^{+}00]. Matrices [BOPZ04, Che96, FLVG95, HCYL06]. Matrix [AAD97, BBRR01, BW96, ÇA99, Cha96, CLPT02, GKK97, LKHL03, LPZ98, Li07, PM96, Sah00a, SR98, THH96, TC95a, TC95b, ZP07, DFD93, ME95]. Max [HS08, HPT04]. Max-Min [HS08, HPT04]. Maximal [LH03, LWJ06]. Maximally [CXP09]. Maximization [LJCL08]. Maximize [HP07, ZS09, WL91]. Maximizing [JGZW08, SM97]. Maximum [BC95, CT97, KGKL08, LDG04, TYK99]. Measure [HT07]. Measured [WB98]. Measurement [DI95, KK03b, WLL^{+}07, HB92, LKG92, MRW92, MCH^{+}90, TV92]. Measurement-Based [KK03b, DI95]. measurements [LEH92]. measures [OC93]. Measuring [AMSK04]. Mechanism [BÖ98, GG09, MY07, RLD03, WS03, WXLZ06, CR94, Geh93, GD94]. Mechanisms [BLD05, CG08, Lop02, ZSMF01]. Media [BV05, CZLM09, ILL07, KSWR03, LL02, SBK02a, SBK02b, TJ07, WL08a, XHYL05, YK09, ZL07a]. Median [WH01, WH03b, XB93]. MediaPort [AOK09]. Mediator [SGB08]. Mediator-Free [SGB08]. MediaWorm [YKDV02]. Medium [JGA08, LJZA04]. Meet [HYP02]. Membership [DS03b, FB01b, MMSA94, YK96b]. Memories [CSR07, BC92, GS91]. Memory [AD98, AGGD04, AAS03, AKN95, Agr98, ADD^{+}02, BCdSFL09, BIWK00, BGMZ97, Bor00, CLS05, Cha96, CH04b, CH07, CH95, CKC08, CSR07, DDS95, DS96, DA98, DKKS04, Deb96, DMKJ96, FT97, FJY98, GAL01, GPST09, GP99a, GMR98, Hol98, HS98b, JR96, JVYA05, KH04, KL01, KHY09, KA05, Lee97, LT97, Li07, LC99, LCL03, Lop02, LBC03, MS94b, MA01, McK98, Mic04, MP97, NN96, OXL06, PAM95, PH96, Par01, PHP03, PH04, PD00, PPBSA97, Qad03, QD05, RvG02, RSB97, SCL05, SW96, SLT03, SLEV03, SNI02a, SNI02b, SZ95b, TD01, TF96a, TP95, WH95, WCCR^{+}97, XCZ02, XCZ04, YY95, YF97, YL97, ZYC95, AH93, AM93, ABJ^{+}93, BIA^{+}97, CF94, DC95, DF97, Don91, Geh93, GH93, Gup92, Har91, HE92, IT93, IC92, Kop94, KCPT96, LEH92, LY93a, Lil94, LH94]. memory [ML94, MR92, NSD^{+}91, PLW96, PAM94, RS91a, RP94, SST94, SL93c, SA93, TMTH96, VGGD94, WFP90, YJZ97, ZLE91, ZSLW92]. Memory-Mapping [CSR07]. Merge [HY05, HNO98c, LB95, WDY93]. merging [Wen96, XB93]. Mesh [BM00b, CT02, Chu95, EF96, EW97, FHA06, FZVT98, GG95, wJPP97, KY98, KyK09, LSF^{+}09, LOSW99, LWLN97, MDSS09, MBM98, NO97, PZLS01, PC96, RS98, SV97, SP98, SS01, TW00, TKP00, WS98, WS00, Wu00, WHC03, YK98, YYS97, dCVGG02, AV94, Cap92, CCCS90, CT94, CS92, GG94b, wJNPS97, LC91b, LMN94, OS94b, SC94, SP93, jTM97]. Mesh-Connected [Chu95, GG95, LWLN97, MBM98, PZLS01, TKP00, Wu00, EF96, CCCS90, GG94b, SP93]. Mesh/Relay [FHA06]. Meshes [Aro00, BBG^{+}95, BGO^{+}96, BGO^{+}98, BGOS97, BGOS98, BNO^{+}01, yCM98, CWCC07, CC01, CH01, CST02, CC99, CCJ02, DHB01, GN96, HNO98a, JSR98, KY98, LS96, LZ02, LC95, LC96b, Li03, LRTZ96, NO98, RS97b, SKK01, ST99a, SY98, SY00, SJPS01, TW98, YW02, BLO^{+}94, BGO^{+}97, EF96, LS94b, MS93, NS95b, PGFS94, UEA95]. Meshes/Tori [LZ02]. mess [RFDS97]. Message [AS99, Bhu06b, BHK^{+}97, CBDW96, DDY99, DFKS01, DHN96, EBS04, FYP07, Gon08, HK98, Hol98, LMN95, MF01a, MRT09, PSK99, RRG07, SRT96, SWC95, SP03, WCLF95, WP00, YC95, vDSP96, ATG92, AMAM94, BR91, BR94, IC92, WG90, YK92]. message-based [YK92]. Message-Dependent [SP03]. Message-Passing [BHK^{+}97, CBDW96, DHN96, HK98, MF01a, MRT09, WCLF95, vDSP96, ATG92, AMAM94, WG90]. Messages [BNH99, BBD00, CJPW06, JGZW08, Kuc01, NSU97, VJA97, WL97, XJZZ00, KGMB94, KH93]. Metadata [ZJWX08]. Metaheuristic [LZ08]. Metaheuristic-Based [LZ08]. Method [yCM98, GS03, HY05, LZ08, LC01, MROD07, PK95a, PK95b, RS97b, SM97, SZ04, TKP00, WZZ09, WHC03, XP07, MM96, SC91, SMJ92, WCS92]. Methodical [KK92]. Methodology [CM95, GBC^{+}07, HP06, LLY05, LP96, LLA^{+}06, LPD05, RRRM09, SRD04, XL08]. Methods [CWCC07, CS95, GKS95, HKM^{+}94, JTP^{+}08, Li03, LC99, MT97, PD99, PSC^{+}95, THH96, YLW07, CF94, DR94]. Metric [BBH05]. Metric-Induced [BBH05]. Metrics [FDC00, LRS02]. Microarchitectures [PSGD05]. Microprocessors [KET06, MC95, WB98]. Middleware [AJMJS03, Ano02b, CS03, FVR03, GZ03, KSC03, RNR^{+}03, TS08, WCH^{+}08, YK03, ZJ03]. Midimew [LC96a]. Migrant [DR98]. Migration [GS03, HY96, ZFMS03, GT93, SW92]. MIMD [BCJ90, CG02a, CG02b, HQL^{+}91, KE90, OD93]. Min [CZLM09, HS08, HPT04, DMTB93, QM94, YD95, ZYC95]. MIN-based [DMTB93, ZYC95]. Min-Cost [CZLM09]. Minigrids [LJW05]. Minima [NO98]. Minimal [DAA00, TC04a, TC04b, Wu00, YD95, Cap92, GPBS94, PGFS94, SC92, SC94]. Minimization [OS02, SWH98, ZKB08]. Minimized [HS98a, KP99]. Minimizing [AMS97, DÖ02, JGZW08, LB00b, TSAL97, TYK99, WCS92, YW93]. Minimum [BBD00, BSCB09, CH09, GW06, GY07, KPK09, KWL^{+}09, LS96, LW09a, LDG04, LL98, SY98, Yi09]. Minimum-Cost [LW09a]. Mining [BS08, CL09, DB06, JZ04]. Minislotted [CLW03]. MINs [VM99]. Mirroring [HJH02]. Mismatch [HLH09, Liu08]. Mitigating [SL09]. Mitosis [MGQS^{+}08]. Mix [FYJ^{+}09]. Mixed [DP01, SCY98, VKS^{+}09, KA94]. mixed-mode [KA94]. Mixed-Parallel [VKS^{+}09]. Mobile [ABS01, Ano01b, Ano01c, Ano01d, BHJ02, CS01b, CS02a, CKK^{+}04, DB08, DS02, GJDA06, GYS05, GY07, GS03, HL08, JLS02, LJW^{+}07, LW09b, MZT08, MX03, NOS99, NSZ02, ON02, PS08, PC05, PS96c, SFP03, SWH98, SZ03a, SZ03b, SSsLY03, TR06, TT01, WDCK04, WO04, WT08, WD06, WYD07, YWD08, ZW02, dLCK^{+}05]. Mobility [AD08, CBM^{+}07, FCF00, MZT08, TM06, WD06]. Mobility-Sensitive [WD06]. Möbius [Fan98, PN93]. Modality [Ksh03]. Mode [Gon08, WYWZ08, KA94]. Model [AMH08, BNBH^{+}95, BNH99, BSCB09, BES06, BP06, BDD^{+}96, CPhX04, Chi98, Chi00, CF99b, Fan02a, Fan02b, FB01a, GT02, GFG^{+}99, Gre98, HY99, HC09, JR96, JKA07, KL01, KS08a, KPR05, LSZ09, LMN95, MZA02, NOZ02, OKSA01, Qad03, Qua01, RMO^{+}95, RRG07, RJ05, SK02, SSS06, SE98, TS98, TTB^{+}00, TPL96, TNPK01, WH03a, WP00, XHYL05, YJ97a, YY95, ZB09, AAG94, AIK91, Bok93, ClW91, DK92, DMTB93, DI95, LH94, MS94b, NJ94, TV92, VGGD94]. Model-Based [BES06]. Modeled [WB98, OSZ92]. Modeling [DS05, FYJ^{+}09, GB00, HM90, KHS07, LYW08, LJW05, MF01a, PF96, SSP^{+}09, Sob96, SvAS04, TR04, vG03, BCBzC92, KCN90a, LEH92, ZY95]. Modelling [MAJ^{+}07]. Models [AAS03, AJMJS03, Ano04c, BDvD98, BA07, BC92, CRS06, CH95, CG02a, CG02b, GY95b, Lee06, LC04, MS99a, PD00, WSC97, WF06, AH93, CO95, Ost90, SH93]. Modes [SCY96, MP91]. modifications [DI95]. Modified [LK04, Chu96]. Modular [AM95, LF03, MF96, WCR09, ZP07, AM91, YZW94]. modularity [SM94]. Modules [DCF95]. Modulo [PP95, ZLAV04]. Molecular [DB06, SGTP08]. money [And90]. Monitoring [MG09, ZBM09, HKM^{+}94, OSS93]. Monitors [YWF^{+}09]. Monotonic [BMR99, LDG04]. Monte [You93]. Movement [AYA09, YLW07]. Movement-Assisted [AYA09, YLW07]. Moving [QD05, XCZ08]. MPEG [KS01]. MPI [BGBP01, kLCC^{+}06, NE01, WC09]. MPI-LAPI [BGBP01]. MPLS [THH08]. MPP [HWWX99]. MPPs [HK98]. MPSoCs [CK08]. mRACER [RE09]. MST [LWS04]. Mtool [GH93]. Multi [CWCC07, CGM^{+}07, FO05]. Multi-Core [CGM^{+}07]. Multi-DSP [FO05].

Multi-Installment [CWCC07]. Multiaccess [CS95, CS97b]. Multiagent [CK02]. multiattribute [GD94]. Multibus [Add97]. Multicast [ABS01, BRS07, BCR98, CHA07, CGK04, CSC07, CJHG08, CC98, CH98, CMDP09, CXN06, DPH08, Dua95b, FIMR01, GG09, GY07, GS03, GKG06, HÖ00, Jia95, JZXX99, KP99, KP01, LCGC07, LW09a, LN93, Mha09, RMC95, SH97, SPS98, SPC^{+}02, TJ07, XGN97, XH08, YMP08, YW99, YW03a, YL07, YL08, YWY08, ZCLC06, ZL07a, ZLP09, LMN94, MXEN94]. Multicasting [CFKR98, Gon03, Gon08, SKPS01, TPL96, VM99]. Multicasts [KWOA05, SS00]. Multichannel [LWN98]. Multiclass [CGL07, GBD07, KK03a, TT94]. Multicluster [BE07, DNSC09]. Multiclusters [HJS^{+}06]. Multicoloring [WH95]. Multicomputer [lCL95, CYY00, HSWB07, LCRW98, CF94, DA93, HB92, KS93, LN93, OS94a, OL92, RS91b, RFDS97, SF92b]. Multicomputers [AD95, CC98, GVGD95, KY98, Lan95, LC99, LCL03, LWLN97, RSB97, SP95, SP98, Ste96, TD01, TW00, TWH99, Wu98, Wu00, Xia01, XL96, dB98, dCVGG02, Bok93, CS90, CS94, GDJ94, GB92, LMN94, SA94]. Multicore [MCG08, SJPL08, TSG09]. Multidestination [PSK99, SSP00]. Multidimensional [AfAGR00, AA00, CW02a, DP02, DD98, Din01, FHBJ97, LCL03, MMSM06, PS96a, SS01, YW02, Aln94b, LK90]. Multidomain [SS07]. Multigrid [MT97]. Multigroup [TSJ07]. Multihop [DSY99, GP03, JGA08, MY07, SCP99, YYY09, ZL07b, KSF94]. Multilayer [AB03, NJ94]. Multilayered [LC02a]. Multilevel [JLF03, WT08]. Multimedia [BHJ02, BSS09, EKOAW02, GB06, HDRS00, LSCZ07, LA04, MEKOT03, SD04, CCQ^{+}05]. multimicroprocessor [VGGD94]. Multimode [MZ05]. Multinode [VB93].

Multioverlay [WLL08]. multipartite [FD94]. Multiparty [CL09, GWYS08]. Multipath [MDSS09, Sob96, WSNA95]. Multiple [AV96, AM06, AKSS04, BNH99, BBG^{+}95, BNO^{+}01, CF01, CHK07, Chu95, EAK97, JR03, JGA08, JO95, KP99, KH97a, LKK02, LJZA04, LL96, LSF^{+}09, PZLS01, PM02, RC95, SLH97, SS00, TH01, VB96, YY95, AN94, AIK91, BLO^{+}94, CCCS90, LG94, LS94c, SB94a, ST93]. Multiple-Beam [LJZA04]. Multiple-Bus [KH97a, TH01]. Multiple-Edge-Fault [SLH97]. multiple-fault [SB94a]. multiplexed [QM94]. Multiplexing [QM97]. Multiplication [BBRR01, ÇA99, CLPT02, wJPP97, LPZ98, Sah00a, SR98, TC95a, TC95b, ZP07]. multiply [ZL96]. multiply-twisted [ZL96]. Multiport [BNBH^{+}95, BNH99, BHK^{+}97, SPS98, jTM97]. Multiprocessing [LMT98, Sar93]. Multiprocessor [AK99b, AM95, Bak05, BÖ98, BKS03, BP96, BCL09, BJM^{+}05, BA97, CRN09, CFR99, FG06a, GY95a, GMM97, GVV09, HT07, JL99, JH97, KWH02, LLTW08, LJLS09, LT97, Li08, LDG04, LBC03, MM98a, MM98b, MJ06, NN96, PAM95, PM96, PPR95, QM97, SH95a, SO95, SJM09, SMJ92, SSZ06, VDS99, WMWL08, WM95, WYJ^{+}04, YJ97a, YJ97b, ZMC03, AC92, BIA^{+}97, Bir93, BC92, BEK^{+}93, CD94, CV92, CAB93, Cor92, DC95, EG93, GD94, GH93, Gup92, HAR94, IT93, IC92, JR94, LS94c, Lil94, MS94a, ME92, ME93, MLS94, QM94, RSS90, SRS93, ST91, SL93b, SL93c, TV92, VJ94, ZL96]. Multiprocessors [AGGD04, AGGD05, AKN95, BB05, BGMZ97, CS08, CW00, CY00b, CH95, CKC08, CY96c, DDS95, DS96, DD95, DMKJ96, FT97, GAL01, GP99a, GMR98, HS98b, KKC^{+}05, KL01, KB06, KA96, KA99, LP96, LLH^{+}01, LK04, LL98, MA01, McK98, PNZ^{+}02, PD00, PGBI03, Qad03, QD05, RTS95, WH95, WHC03, YL97, AOB93, ABJ^{+}93, And90, BJS90, CS92, DMTB93, Gab90, HM92, JF94, Kop94, KE90, KCPT96, LS94a, MS94b, ML94, Pad91, PAM94, RB90, SS90, SG93, SS94, TRS90, WW92, WFP90, YTB92, YW93, YD94a]. Multiprogrammed [YL97, SST94]. Multiquery [WTCY95]. Multiresource [SL06]. Multiround [YvdRC05]. Multisensor [SvVB05]. Multiserver [CGL07]. Multisignature [vdMDM07]. Multisite [SRD08]. Multiskewing [Deb96]. Multistage [BIWK00, LKK95, LSC95, RO99, SPS98, Sob96, TZ97, Tze04, WL97, XGN97, YW00, YW01, YW04, BIA^{+}97, CI92, HC92, LC94, MD96, YM95, YA93]. Multistage-Based [Tze04]. Multistep [dB98]. multistride [Har91]. multisystem [DY93]. Multitarget [PPBSA97]. Multithreaded [BKI06, BF04, CH95, CMBAN08, GMR98, LLS06, LPE^{+}99, MGQS^{+}08, SCL05, Aga92]. Multithreading [KET06, MB07]. Multitoroidal [ADG^{+}08]. Multiunit [XL08]. Multiway [LB95, MC95, Wen96]. Mutable [CS01b, CS02a]. Mutual [AMP07, CS01a, CH09, FT97, HL08, HY05, HS98b, JK99, Jou03, KKM08, KM01, LK00, TYK99, XXZ03, BCBzC92, HMR94, IK93, NLM90, Sin92]. MVSS [MR03]. Myrinet [FLMD02a, FLMD02b].

n [OC93, SG94]. n-cube [OC93, SG94]. NAD [SD04]. NAD-Based [SD04]. name [KM91]. name-space [KM91]. Narrow [MBW02]. NAS [KHS07]. NAS/PSA [KHS07]. Native [EBS02]. Natural [TS08]. Near [KLS00, YW02]. Near-Optimal [KLS00, YW02]. Nearest [KP96, LS96, NO97, WHW05]. Nearly [CC97]. Necessary [Dua95a, Dua96, NX95, VS11]. Negative [CH04b]. negligible [SS94]. Negotiation [JJ09]. Negotiation-Based [JJ09]. Neighbor [NO97, SSZ02, Sto04, WHW05].

Neighborhood [JJ07]. Neighbors [LS96]. nested [LK90, ST91, SC91, WW92]. nests [DR94]. net [CTC93, SMBT90, STMD96, VGGD94, NE01]. Net-dbx [NE01]. NETRA [CPA93]. Nets [JK99, BCBzC92, WF94]. Network [Ano04d, ABC01b, AB03, BBH05, BA97, BIWK00, Bok93, CFB02, CH04a, CHK07, CHL09, CS95, CJHG08, CZLM09, DC98, DS03a, DS05, DLS09, DR98, DLPP05, DCF95, EK95, FYS05, FV09, Fu05, GKK05, GBC^{+}07, GS95, HY04, HSWB07, HY99, HH08, HGC05, HH95, HW08, JTC08, KSWR03, KPBD09, LCRW98, LB95, LR93, LXN07, LWN98, LK04, LPD05, MKR00, MZT08, MKY^{+}09, MF01a, NT09, OPZ99, Pak07, PPD03, Pre99, PDH06, Ros02, Sah00a, Sah00b, SS96, SF08, SF95, SC07, SYC03, SSRV99, Sol02, SP05, Ste96, SSsLY03, TTB^{+}00, TZ97, THT^{+}97, TWH99, TF96b, US04, VB96, WCY95, WSNA95, Wan98, WOT^{+}07, WF06, WLL08, YW99, YFJ^{+}01, YWD08, ZL07a, ZS09, ZN04, ZLKK07, Aga91, AN94, Aln94a, Aln95, CV92, Chu96, KP92, LB94, LK94, MS94a, MR92, MJ94, PGDS94, PN93]. network [SSG91, WS93, SL09]. Network-Attached [MKR00]. Network-Based [Ste96]. Network-Coding-Based [CJHG08]. Network-Partitioning [TWH99]. Network-Supported [ZL07a]. Networked [BES06, CG08, KMW08]. Networks [AYA09, ABC^{+}01a, AB99, AV96, AS00, ALW^{+}03, AD08, AD09, AA00, Ano98b, Ano01b, Ano01c, Ano01d, Ano03c, AA09, BÖ98, BK09, BRS07, BRSS08, BBS^{+}09, BLD05, BSCB09, BCL^{+}05, BWS^{+}05, BRSR08, BC06, BM00a, BPT03, BHL^{+}07, BS08, BC95, BBR07, CLW03, ÇF99a, CHA07, CYW08, CDV^{+}06, CLB08, CBD^{+}01, CCC05, CBM^{+}07, CL97, CC97, CY06, CPX06, CSC07, CH08, CLY08b, CJLN09, CHC09, CTF09, CXP09, CJHG08, CKWC08, CS02b, CS97b, CFKR98, CMDP09, CNT05, DW04a, DW04b, DW06, DSY99, DPH08, DZ04, DAA97b, DAA97a, DAA00, DAA02, DAMK06, DLS09, DB08, DY05, DD98, DWX09, Dua95a, Dua95b, Dua96, Dua97, EF95, EAK95, EAK97, EKOAW02, FHA06, FCF00, FR96, FE97, FF98, FLMD02a, FLMD02b, FG06b, cFC98, FYJ^{+}09, GZ06, GY95a, GLY07, GRY07, GD95, GLS07, GJDA06, GP03]. Networks [GBC^{+}07, GY09, GYS05, GY07, GS03, GSS06, HÖD99, HS97, HS99a, HÖ99, HSLA05, HCHM09, HL09a, HP03, HTPS02, HYP02, HPT04, HLL09, HLH09, HL09b, HC09, HW97, HCD97, HZ96, HC99a, HC97, HWSH00, ISRS06, JL99, JGA08, JJ07, JLS02, JN08, JASA08, JKA07, KZ96, KZN07, KP99, KP01, KPK09, KWL^{+}09, KyK09, Kla98, KAY^{+}06, Kop96, KWH03, KS01, KS08b, Lai00, LKK02, LC96a, LKK95, LO95a, LW95b, LS97, LDCO08, LMS04, LL06a, LL06b, LCWW03, LWS04, LH06a, LSF^{+}09, LWC^{+}09, LCN^{+}07, LR97, LMN95, LLWC09, LRS02, LSC95, LWXS06, LH06b, LJW^{+}07, LWP07, LW09b, LZ05, LSRT06, MGZN07, MCL^{+}07, MY07, MEKOT03, MZA02, MMSM06, MRLD01, MR06, MTK06, MAJ^{+}07, NOS99, NO00a, NO00b, NOZ01, NO02, NGM97, NYD09, NSZ02, ON02, OSRS06a, OSRS06b, PHKC09, PSK99, PK01, PR05b, PR05a, PC96, PKL06, PP05, PS96b, PF96]. Networks [PW99, PG07, QNR99, RO99, RRX09, RGL05, Rav07, RLW^{+}07, Res97, RWW97, RE09, RMC95, RLD03, RH00, RH04, SKS02, SJd^{+}09, SRZF04, SO95, SJM09, SCP99, SX07, SD00a, SD00b, SPS98, SKPS01, Sob96, SY97, SC05, SLFW06, SP07, SGL06, SS07, Sto97, SL01a, SL01b, SSZ02, Sto04, SZ03b, SS01, SDFV96, SCL00, SCL01, SOM05, TX08, TXL08, THH08, TLM04, TR06, TN08, jTM96, TPL96, TLGP97, TH01, TSJ07, VDS99, VM04, VS11, WY07, WL97, WO04, WWL06, WCH^{+}08, WT08, WLZ08, WWLS08, WWWA09, WP00, WL00, WA99, Wu02, WCDY06, WD06, WYD07, WLZN07, WCD08, XXZ03, XPL04, XP05, XP07, XCZ08, XGN97, XTL08, YK99, YK98, YN00, YW00, YW01, YW03a, YW04, YW05b, YWD08, YLW07, YV98, Yi09, YGE06, YYY09, YJHG06, YKP08, YG08, YP98, ZCLC06, ZF07, ZS09, ZCXF09]. Networks [ZL07b, ZH98, ZPY06, ZKB08, ZL08, ZLP09, ZB09, ZL05, AAG94, AV94, Aln94b, Ant94, BR91, BR94, BFP96, BGM94, BIA^{+}97, BCH94, CAB93, CI92, CO94, Cor92, DA93, DGB^{+}96, DS94, Dua93, FD94, Fid92, GP93, GPBS94, HC92, HK94, JR93, KSF94, LS94a, LC94, LN93, MXEN94, MD96, NJ94, Nic92, NLM90, OC93, ÖD96, Pad91, PGFS94, RS94, RWF94, RFDS97, Sch91, SG94, SB94a, SC93, SR91, SCD97, Tak93, TH93, jTM97, UEA95, VS96, YK96a, YK96b, YC93, YM95, YN90, YA93, ZS95b, Zia94]. Networks-on-Chips [KAY^{+}06]. Neural [AB03, EAK97, Pre99, NJ94]. Next [ZSMF01]. NFS [BB08]. No [NO00a, GR90]. NoC [BJM^{+}05]. Node [BRTM09, KP99, RGL05, STY09, TCS97, WCD08, YW03b, YW05b, jTM97]. Node-Disjoint [YW03b, YW05b]. Nodes [BFL^{+}01, Fu05, GP99b, JHK97, LJZA04, SX08]. NODUP [CYW94]. Nomadic [KL02]. Non [CSC07, HJS^{+}06, PNZ^{+}02, PB96, KGM96, SS94]. Non-Cache-Coherent [PNZ^{+}02]. Non-DHT [CSC07]. non-negligible [SS94]. Non-Real-Time [HJS^{+}06, KGM96]. Non-Uniform [PB96]. Nonblocking [LZ05, QS03, SO95, YW03a, AB91a]. Nonclairvoyant [HHL08]. Noncombining [ST99a]. Noncontiguous [LWLN97]. Nonconvex [CC01]. Noncubic [SP95]. Nondominated [BI95, HY97, HY05, KH98]. Noninstantaneous [CGL07]. Nonlinear [BE98, KP09, CARW93, SC91]. Nonmigratory [LLTW08]. Nonuniform [CY96a, Kop96, WCD08, AM93].

Nonunimodular [FLVG95]. normalization [Omi90]. Notation [CF95]. Note [Bhu06a, Bhu07a, Bhu07b, Bhu08, Bhu09b, Bhu09c, CH98, HGC05, SCY96, Yew03, Yew04a, Yew04b, Yew05a, Yew05b]. Nothing [RD98]. Notice [Ano02c]. Novel [ADG06, BS08, CN02, CN04, Deb96, KL02, LM06, LZ08, Rob04, SKJ07, SX03, TH93, THH08, XL08]. NOWs [AA09]. NRMI [TS08]. NUCA [HKS^{+}07]. NULL [KH93]. NUMA [AGGD05, BIWK00, DMKJ96, LEH92, PGBI03, ZY95]. Number [BM00b, CH09, GP99b, PP95, UKY98, Tho93, YG94]. Numbers [YK99, NS95b]. numeric [HB92, Lar93].

O [WSB09, Bor00, JSWB97, KKCB02a, KKCB02b, Kan01, KB03, kLCC^{+}06, OPZ99, RB90, TR04, VV99, WSB09, YZC08]. O-Centric [HJH02]. O-O-O [WSB09]. OBIWAN [FVR03]. Object [GMS09, JLDC05, LSCZ07, RS08, RLW^{+}07, TF01, Tse09, XRR00, XTL08, YK03, SM94]. Object-Tracking [XTL08]. Objective [LSZ09]. Objectives [LKK02]. Objects [AM99, KMW95, LA04, Mic04, MTK06, IA95]. Oblique [ABRY03]. Observations [ZT01, ZW02]. Obtain [MRT06, BR91]. Occurrence [JK99]. OCGRR [GRY07]. OCI [LNYY03]. OCI-Based [LNYY03]. octrees [IA95]. Odd [Chi00, LH01, RS90]. Odd-Even [Chi00]. Off [FHA06, FLP^{+}07, QCC99, SP07]. Offloading [CKK^{+}04, SF08]. Offs [CKK^{+}04, DZH05, GZ09, MYA01, ZCXF09, DF97]. Offset [LCRW98]. OLAP [LA06]. Old [Mit00]. Omega [PW95, BR91, BR94]. On-Chip [AGGD04, Ano03c, HP06, KKC^{+}05, MKY^{+}09, PSGD05, PP05]. On-Demand [CZLM09, ILL07, JGA08, SKS02, WL08a, XTL06]. On-Line [ANKA99, Bir93]. On-the-Fly [KS06, PK00]. On/Off [SP07]. One [AJF96, CC97, FMR07, LWJ06, RHM09, XP05, KST94]. One-Directional [AJF96]. One-Hop [RHM09, XP05]. One-Shot [FMR07]. One-to-Many [CC97]. Online [CHL09, EDO06, HKL00, HHL08, TSRS07, Tse09, ZLZN09, ZBM09]. onto [EAK97, HÖ99, IS90, KB06, SS94, TKP00]. OPAM [BS96]. Open [BCL^{+}05, YLL^{+}07, DFD93]. Open-Source [YLL^{+}07]. OpenMP [ACD^{+}09, MM07]. Operand [BWS^{+}05, SS08]. Operand-Load-Based [SS08]. Operated [NK08]. Operating [LBS05, VGGD94]. Operation [HY01, HY05, Ian97, KST94]. Operational [LL07, SS09]. Operationally [KS94]. Operations [Agr99, BNBH^{+}95, Bar98, BDD^{+}96, GY07, JSWB97, LCL03, Sah00b, SCL05, THH96, WS98, MR92]. Operator [RSP02]. Operators [ZMP07]. Opportunistic [CWYZ09]. Opportunities [CW02a]. Opportunity [AAB^{+}00, KB03]. optic [AAG94]. Optical [CFB02, CWYZ09, DS03a, FR96, GP03, HSWB07, LWN98, LK04, MR06, MAJ^{+}07, RS97a, Sah00a, Sah00b, SCP99, WL00, WH01, YW01, YW05a, YJHG06, ZY04, ZY06]. Optically [QM97]. Optics [LCRW98]. Optimal [Aln94b, AR97, ABRY03, ADD^{+}02, BFP96, BBG^{+}95, BGO^{+}96, BGO^{+}98, BGM94, BGOS97, BNO^{+}01, CS01a, CC93a, CCP95, CGK04, CYW94, CC97, CC95, CNNS94, CXN06, DA98, DPS96a, DPS96b, DP02, Deb96, DS05, DY05, DD01, DD95, Din01, EK95, FLJ05, FJL07, FCF00, FI95, GW96a, GRS99, GAG96, HNO98b, HNO98c, HK95, HS02, HTPS02, HWKH01, HH95, HZ96, ISRS06, JR93, JR03, wJPP97, JLDC05, JVYA05, JEG07, KDW01, KZ96, KCS^{+}99, KR00, KLS00, LC96a, LC95, LS97, LT97, LHSML95, MC93, MS92, MG09, NO97, OW91, OSZ92, OZ96, RA04, Rav07, Res97, RMC95, Ros02, SK02, SP93, SWC95, ST99a, TCC07, TLGP97, TH01, WKS01, WMN99, WL08b, XGN97, YMP08, YW00, YW01, YW02, YL08, YXW03, ZY04, ZL96, AGE94, BGO^{+}97, Fid92, Fu97, JR94, LK94, LA93, SB94b]. optimal [Uht92]. Optimality [LC02a, XU01]. Optimally [BSS09]. Optimistic [QS03, VJA97]. Optimization [BCG04, HKL00, KA09, KM02, LKKS05, LSZ09, LJLN07, McK98, PC05, SKB04, SKCL09, SCO^{+}07, TM06, TKVD02, TK96a, XP05, ZCXF09, AT07]. Optimizations [CE95, KK04, KKCB02a, KKCB02b, KBC^{+}01]. Optimized [BV05, CFKR98]. Optimizing [AMY09, AKSS04, COS00, GSS96, HCYL06, KKC^{+}05, KCRK00, LA04, MGDZ07, PPP04, SRL98, WSB09, XLW^{+}06, ZXZ^{+}09, AC93]. Optimum [Bar98]. Optional [Sun02]. Optoelectronic [WS98, WS00]. Order [BC99, FIMR01, WSB09]. ordered [GDJ94]. Ordering [AJF96, CH98, EBS04, Jia95, SH97, Var93]. Orders [KSP09, HMW93]. ordinary [GP92]. organization [DC95]. Organizing [CDV^{+}06, SH95b]. Orientation [UKY98]. Oriented [CV08, GMS09, HL09a, KCK^{+}06, LP96, RNR^{+}03, YZC08, dBL98, MN92]. Orthogonal [HJH02, Sch91]. Oscillation [hKY08]. other [Fid92, PGFS94]. OTIS [CXP09, DAA02, RS98, WS98, WS00]. OTIS-Mesh [RS98, WS98, WS00]. OTIS-Networks [DAA02]. Out-of-Core [DW03, KCRK00, LRG99]. Outerplanarity [KR00]. Output [FZGC06, GCCC^{+}04, MLW06, MR02]. Overall [COS00, YJHG06]. Overflow [SFP03]. Overhead [BG02, CC99, FPGAD08, KB03, PF08, SRT96, Kum92, LLJ^{+}93, NZ95, ZLE91]. Overheads [SSRV99]. Overlapping [kLCC^{+}06, YYY09]. Overlay [AOK09, BRS07, BRSS08, BBR07, CLB08, CSC07, CXN06, GY09, LCGC07, LLSZ08, LXN07, PDH06, SL09, TJ07, TSJ07, WCBX06, WL08a, YMP08, YL07, ZCLC06, ZL08, ZLP09, ZCSY08]. Overlays [BK09, FRGL09, MG09, PZZ09]. Overload [Ram99]. Oversubscribed [TTB^{+}00]. Overview [LLY07]. Ownership [JB01].

P [HK98, SK02]. P-3PC [SK02]. P2P [BSS09, BRTM09, CSC07, CLY08b, CT08, CZLM09, HL08, LXN07, LZTY09, WXLZ06, WX07, WLL08, YM09, ZYKG07, ZH05, ZH06, ZH07c, ZCSY08]. P2P-Based [LZTY09, ZH07c]. P2Ps [LHL^{+}08]. Packaging [BP96]. Packet [ADG06, AH06, DHN95, DZH05, FR96, GR06, GS08, GG95, HPT04, KSP02, LMS04, LL06a, LL06b, LLY07, LSC95, MSM09, PC07, PF96, RS97b, SX03, Tze06, WR04, WLL^{+}07, XZG09, MS93, PGFS94]. Packet-Based [LL06a]. Packet-Switched [LSC95]. Packet-Switching [LL06a, LL06b]. Packets [LZ02, ST99a, VB93]. packing [BW94]. Packings [dBL98]. Page [DYJ97, Bir93]. page-parallel [Bir93]. Pages [HZ97]. Pair [WHW05]. Paired [WF03]. Pairwise [MCL^{+}07, MDL06, TC94]. pancake [BFP96]. PAPADS [Ano07c, ACM08]. Papers [Ano97d, Ano97b, Ano97c, Ano98c, Ano01b, Ano01c, Ano01d, Ano02b, Ano04b, Ano04c, Ano04d, Ano05c, Ano07c, Ano08c, Ano09c, Ano09b, Ano98b, Ano99c, Ano99d, Ano99e, Ano03c]. Paradigm [BLR03, JKR01, OC05, WSC97, ZL05, MN92]. Paradigms [OB00]. Paragon [FBD96]. Paralex [DGB^{+}96]. Parallel [AKN95, AK98, ACM08, AM90, AFAGR97, AJMJS03, AfAGR00, ATML08, ACT^{+}97, Aln95, AGL^{+}98, AM06, ABK98, AKSS04, Ano97d, Ano97b, Ano97c, Ano02a, ABDZ94, AH06, ADD^{+}02, AIK91, BT00, BCVCV05, BBC^{+}95, BDvD98, BJS90, BKB96, BA07, BAH01, BA97, BP06, COP00, CdMB05, ÇA99, CARW93, CFB02, CC93b, Cha96, CH07, Che95b, Che96, CC97, CFW98, Che01, CW02b, CPhX04, CV08, CY96c, CB00, CJPW06, CN02, CN04, CSR07, DPS96a, DPS96b, DHB01, DGB^{+}96, Deb96, DHN95, DDD^{+}05, DHN96, Din01, DL02, DCSM96, DNSC09, FE97, FHBJ97, FDC00, FFPF05, FA94, FBD96, FI95, FARH02, GMRC07, GRS99, GCCC^{+}04, GvG06, GY95b, GKS95, GSS96, GKK97, HM98, HNO98b, HCF03, HCY97, yH02, Hsi03, HLV94, HH95, HXA96, IA95, JTP^{+}08, JZ04, JVYA05, KABK03, KHWT95, KL01]. Parallel [KG92, KPR05, KA99, LB00a, LH93, LO95a, LC95, LL96, Lee97, LKHL03, LHS03, LM06, LCB96, LPZ98, Li07, LP07, LT00, LBS01, LC99, kLCC^{+}06, LOSW99, LLH^{+}01, LCL03, LNOZ03, LSBS98, LS06, LRTZ96, LWN98, LL94, LZ05, LMT98, MR02, MD97, MJ98, MT97, MNM04, MS99b, NZ95, NLW99, Nas93, NL02, NKP^{+}96, OHRW99, OXL06, OR97, PR05a, PKJ97, PWW00, PG01, PK95a, PK95b, Pre99, PH02, QCC99, Qua01, QS03, RL98, Raj05, RA04, RK93, RR02, Rob04, SA09, SKB04, SZ02, SF09, SW96, SSP00, SSRV99, Soh95, SCO^{+}07, SP03, SCP02, SPF99, SZ04, SOM05, TSP^{+}08, TP95, Var01, VV99, VB95, VKS^{+}09, WCL97, Wan98, WKS01, Wan04, WHM09, WL00, WCF91, WDY93, WTCY95, WHL95, WDY98, WMB96, Wu97b, XQ08, XB93, YFJ^{+}01, YDW^{+}09, YFM98, YZC08, ZFMS03, ZY07, ZH98]. Parallel [ZH99b, ZCO98, ZWM99, dSF03, vG03, vDSP96, AOB93, AH91, ADM92, Aln94a, AN93, AC93, BS95, BW94, Bir93, BCJ90, CA93, CCCS90, ClW91, CWL92, DM93, Don91, DFD93, Efe92, GO93, GR90, GMG96, GS91, GK93, HISS94, Har91, HQL^{+}91, HN93, HE92, HB92, HK93, IT93, JS90, KK94, KMT91, KCN90a, KCN90b, KM91, KGS94, KSA94, Lee93, LC91a, LNP94, Lil94, LL90, MS91, ML90, MB94, MM96, ME95, MCH^{+}90, MKH91, MTSDA93, NSD93, Nic92, NGL94, OSS93, OW91, OSZ92, Omi90, PLW96, RK94a, RK94b, Rao96, RJ94, SP93, SST94, SL94, SW95, SR94, SMJ92, Tak93, TB93, TN93b, Tze93, WW92, WCS92, Wen96, WLR93, WYTD93, WM93, YJZ97, YG94, YD94a, You93, YC96, ZLE91, KP93b]. parallel-acting [MM96]. Parallel-Pipeline [KPR05]. Parallel-Systems [SF09]. Parallelepiped [RR02]. Parallelepiped-Shaped [RR02]. Parallelism [AGWFH97, KCRK00, MA97, MA01, PAM95, PS96a, RSP02, RSB97, TSG09, WHL95, GP92, Lar93, MR94, RM90, WL91]. Parallelization [CL05, EHP98, Gre98, KP09, MSH00, OB00, PPBSA97, RP99, SJKC06, XC01, YR06, JWC94, KKP91, NE93, TN93a]. Parallelized [DHN96, TMTH96]. Parallelizing [ASS95, AK99b, FS00, FO05, HN90, HCYL06, Lee95, BE92, CS94, CL94, GB92, LYZ90, SLY90]. Parameter [XL04]. Parameterized [CWLR09]. Parameters [ZSMF01]. ParaScope [KMT91]. parentheses [PDC94]. parentheses-matching [PDC94]. Parenthesis [Sto96]. Parity [Par95]. Parsing [NLW99]. Part [DLPP05, LPD05, OSRS06b, PK95a, PK95b, RK94a, RK94b, YK96a, YK96b]. Partial [Agr98, DP02, FJY98, LSW04, RLW^{+}07, ZH07a, You93]. Partition [HY04, RL98]. partitionable [CPA93, JS90, LC91b, NSD^{+}91, WS93]. Partitioned [BC99, DS03a, MR06, RJ94, Sah00a, Sah00b]. Partitioners [SCP02]. Partitioning [AKN95, BA07, BR94, ÇA99, Cha96, CM95, COS00, CT02, D’H92, DWX09, IB95, JO95, LC02b, MROD07, OR97, PB96, RR02, ST91, SvVB05, TKP00, TWH99, Tze06, AH91, GB92, Gup92, LC91b]. Passing [BHK^{+}97, CBDW96, DFKS01, DHN96, HK98, Hol98, MF01a, MRT09, PSK99, RRG07, WCLF95, vDSP96, ATG92, AMAM94, WG90]. Passive [DS03a, GP99a, MR06, Sah00a, Sah00b]. Patch [KSP09]. Patch-and-Stitch [KSP09]. Path [FLJ05, FH97, GZ06, HSWB07, Hol98, KL99, KA96, PKL06, QM97, SM03, THT^{+}97, ZH98, BR91, CWL92, SCD97]. Path/Flooding [SL01a]. Paths [FJL07, LC01, PSK99, SX08, UFS96, YW03b, YW05b, GPBS94, KGMB94, TR93]. Pattern [DKKS04, LS06, SDFV96]. Pattern-Based [LS06]. Patterned [YY95]. Patterns [AMS97, Aro00, GS95, MR02, TW00, BR94]. Payment [TJ08]. Payment-Based [TJ08]. PC [JZ04]. PCBN [WS93]. PCS [FCF00, WOT^{+}07]. PDE [WH95]. PDF [Ano00b, Ano00c, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k]. PE [Kop94]. PE/memory [Kop94]. Peer [CW06, CLY08a, CJLN09, CHC09, CHHC06, CGM05, DF09, FRGJ07, FRGL09, GN06, GWYS08, GY09, GLQL09, GSS06, HL09a, HH08, HLL09, HLH09, HCC06, JGZW08, LXL08, LYW08, LLSZ08, LLWC09, LXL^{+}05, LLX06, MTK06, PDH06, RGL05, SC07, SX07, SGL06, STW00, TJ08, TXL08, WL08b, XXZ03, YK09, ZH07a, ZF07, ZXZ^{+}09, ZH07b, ZKB08]. Peer-to-Peer [CW06, CLY08a, CJLN09, CHC09, CHHC06, CGM05, DF09, FRGJ07, FRGL09, GN06, GWYS08, GY09, GLQL09, GSS06, HL09a, HH08, HLL09, HLH09, HCC06, JGZW08, LXL08, LYW08, LLSZ08, LLWC09, LXL^{+}05, LLX06, MTK06, PDH06, RGL05, SC07, SX07, SGL06, TJ08, TXL08, WL08b, XXZ03, YK09, ZH07a, ZF07, ZXZ^{+}09, ZH07b, ZKB08]. PeerCluster [HCC06]. peerTalk [GWYS08]. Percolation [AD09]. Perfect [HHM^{+}00, PR05b, PR05a, BE92, EHP98]. Performability [NGB^{+}05]. Performance [AD98, ASB02, AFM02, Abr97, AGGD04, AV94, Aga92, AC92, AS92, AMAM94, AS96, AAB06, AA00, Ano05c, Ano09c, BT00, BDvD98, BKB96, BMPP06, BIA^{+}97, BIWK00, BE92, BCG04, BCR98, BDS94, CE95, CLB08, CGK04, CY95, CK08, CLY08b, CTF09, CS95, CV08, CY00a, CY00b, CH95, CML05, CS03, CG02a, CG02b, CMBAN08, DW04b, DY93, DP06, Din06, Don91, EBS02, Fei05, FDC00, FLMD02a, FLMD02b, FG06a, FLP^{+}07, FYJ^{+}09, GB00, GvG06, GMCB01, GKS95, HDF07, HC92, HB92, HNY02, HK93, HWWX99, ICT93, JF94, KHY09, KL99, KYD^{+}07, KA05, KWOA05, KS93, LG94, LJZA04, LB00a, LP96, LSZ09, LY94, Li08, LLGS09, LT00, LR97, LBS05, LY93b, LC04, LMT98, MKR00, MS91, ME92, MRW92, MMSM06, MSM06, MD96, MCG08, MOFD05, MDL06, NJ94, NGM97, OHRW99, ON06]. Performance [OC05, Pak07, PR05b, PHP03, PPP04, PPR95, PGBI03, QNR99, RK08, RBSP02, SRD04, SG93, SFP03, SkLC^{+}03, SD00a, SSP02, SvAS04, SZ95b, SM02, SMH02, TSG09, TGV08, TM97, TL05, Tho06, THW02, TZ97, TKVD02, TK96b, VSD01, VMXQ04, Var93, VR05, WSC97, WB98, WOT^{+}07, WF06, XC04, XTL06, YW98, YD94b, ZYC95, ZMRS08, ZCXF09, ZH06, ZBM09, ZMP07, ZWM99, dBL98, vG03, Aga91, And90, DF97, DI95, DAF95, EAL91, EMS90, GH93, GS91, HKM^{+}94, LLJ^{+}93, ML90, RS94, SMS93, SF92b, WS93, YC93, ME93]. Performance-Based [AA00, KL99]. Performance-Driven [CML05]. Performance-Effective [THW02]. Performance-Guided [ZMRS08]. performance-memory [DF97]. Performance-Oriented [dBL98]. Performance-per-Watt [KHY09]. Performing [Lai00]. Perimeter [CS05]. Perimeter-Based [CS05]. Period [SC94]. Period-processor-time-minimal [SC94]. Periodic [JR03, MLW06, Ram95, SA94]. Periodically [Ano99f, PK99b]. Periods [RH00]. Permutation [CST02, DZ04, NOZ01, NS95a, SBF00, SyFL99, WMN99, MS93, RWF94, YC96].

Permutation-Based [CST02]. Permutations [Lai00, YW03b, YW05b]. Persistent [Lop02]. Personalized [FYP07, SS01, TG96, YW00, YW01, RWF94]. perspective [MTSDA93]. Perturbation [CL09, MRW92]. Pervasive [KKS07]. pessimistic [SB94b]. PET [CL94]. Petersen [ÖD96]. Petri [BCBzC92, CTC93, JK99, SMBT90, STMD96, VGGD94, WF94]. pf [BE92]. pFusion [ZYKG07]. Phase [Agr99, Her00, HY07, HLH04, LH01]. Phased [KKC03]. Phenomena [JN08]. Photonic [LZ05]. Physical [Ano08c, TGV08]. PI [HY07]. Pica [WCCR^{+}97]. pin [Fid92]. pin-optimal [Fid92]. Pipeline [KPR05, SS08, SM03, YKS03, AN94, EMS90]. Pipeline-Based [YKS03]. Pipelined [DSO02, HÖ99, KCN90a, KCN90b, LPZ98, Li03, RJ96, SDDY00, WHW05, ZMP07, CNNS94, JR93, SG94]. Pipelines [FDC00]. Pipelining [AB94, BLMR05, CDR98, GAG96, KL01, ANN95]. Pivoting [FJY98]. Placement [Agr99, BRSR08, KDW01, KM02, LSCZ07, Par95, RC95, RSG06, TX05, TC06, TCC07, Tse05, BJS90]. plane [SA93]. Planning [SKCL09, SZ03a, dSF03]. Platform [Ano04c, CRS06, FVR03]. Platforms [BBC^{+}04, BBRR01, BLMR05, BCL09, CF00, DNSC09, HK06, LSZ09, MTSDA93]. Player [CHL09]. PMC [HC09]. Podality [BGOS97]. Podality-Based [BGOS97]. Point [DSY99, HÖ99, SK02, ZP07, Cor92]. Point-to-Point [DSY99, HÖ99, SK02, Cor92]. Pointer [CHJL04, CAZ04]. Pointer-Based [CAZ04]. Pointers [Mic04]. Points [HNO98b, HNO98a]. Poisson [SZ04]. Policies [BRSR08, BIWK00, BE07, CV08, CYD98, DYJ97, LA06, DY93]. Policing [RH04]. Policy [BCdSFL09, LR96, LG09, SRD08]. Policy-Enforced [BCdSFL09]. Polling [Res97]. Polymorphic [Mar93, TC07]. Polynomial [BSCB09, CF94]. Port [HÖ00, HK95, KLS00, jTM96, YW02]. portability [ABJ^{+}93, AN93]. Portable [AGL^{+}98, BBC^{+}95, DR98, LB00a, Gab90]. Positions [Qua01]. possible [HMW93]. Postal [BNBH^{+}95, BDD^{+}96]. Posteriori [KGKL08]. Potato [BRS97, NS95a]. Potential [CV08, MTL95, SP05]. Power [ACM08, Ano07c, CMBAN08, FMR07, KGKL08, Li08, MGDZ07, MB07, Mit01, MCG08, PS08, SP07, SL01b, ZMM04, MM96, WT92]. Power-Aware [ACM08, Ano07c, Li08, PS08, SP07, SL01b, ZMM04]. Power-Performance [CMBAN08]. Power/Performance/Thermal [MCG08]. PowerPC [AAS03]. Powers [Li07]. PowerTrust [ZH07b]. pp [RFDS97]. pp-mess-sim [RFDS97]. Practical [AFAGR97, DDV^{+}07, FB01b, GS08, KA99, Ste96, WT98, Gab90, TN93b]. Practically [GLV06]. PRAM [Che95a, HNO98c, PDC94, WH03a]. Precedence [BKS03, BBD00, HÖ99, Ram95, AMAM94, SS94]. Precedence-Constrained [HÖ99, AMAM94]. Precedence-Related [Ram95]. Precedent [LT00]. precise [CT94]. Precomputation [MGQS^{+}08]. Preconditioned [GKS95]. Predicate [CK96, DL02, MSG07]. Predicates [Ksh03, GW94, GW96b]. predict [DI95]. Predictability [MF01b]. Predictable [HS99b, KSWR03]. Predicting [ML90, XC04]. Prediction [CMBAN08, Din06, DF99, GvG06, GDI93, LT00, SMS93, TAKB06]. Prediction-Based [CMBAN08, GDI93]. Predictions [TEF07]. Predictor [TAKB06]. Preface [OSRS06a, OSRS06b]. Prefetching [COS00, DDS95, DS96, KE90, LJLS09, SLT03, TCC05, TR04, TKVD02, VV99, Lil94]. Prefix [BM00b, Chu95, LNO^{+}00, LNOZ03, Tak93].

Presence [DHP^{+}07, NT09, OKSA01, Sin96, SCY98, VRKL96]. PReSENt [KyK09]. Presentation [GT02]. Preservation [CGM05]. Preserving [CL09, JBW^{+}08]. PRESS [CB05]. Prevention [LSC95]. Pricing [GBD07, HYP02]. Primary [ZJ99]. Primary-Backup [ZJ99]. primitives [JWC94]. Principle [XU01]. Prioritized [JH97, TTB^{+}00]. Priority [BÖ98, DPS96a, DPS96b, LLL09, WMWL08, EG93, Nic92, OW91]. Priority-Driven [BÖ98]. Privacy [CL09, LZR09]. Privacy-Aware [LZR09]. Privacy-Preserving [CL09]. Private [ZZMN07, WFP90]. Privatization [RP99]. Proactively [vdMDM07]. Probabilistic [Arv94, CHJL04, KMG03, KCK^{+}06, ZZMN07, LS94c]. Probability [DÖ02, HY99, MAJ^{+}07, RO99]. Problem [AK99b, BSCB09, BNO^{+}01, CT08, CKWC08, FH03, Gre98, HTPS02, HLH09, yH02, NO97, PPBSA97, PK95a, PK95b, TC04a, THT^{+}97, TKVD02, WLZ08, YK99, CWL92, FD94, LL94]. Problem-Solving [PK95a, PK95b]. Problems [BCL^{+}05, CB00, DMR01, FMR07, Gon08, HH95, IB95, LLY07, PLT00, RL98, SK02, SKB04, THT^{+}97, UZCZ97, WKS01, WHW05, O’H91, OSZ92, RJ90, SW95, WC90, YK96b]. Process [DTE07, GM09, JBW^{+}08, SvVB05, GT93]. Processes [BCdSFL09, CLB08, CF95, LPD05, MRT09, WM93]. Processing [BDvD98, CFB02, DHB01, DW03, HXA96, KYB08, KKC03, LB00a, RGK09, SKB04, TSP^{+}08, WS00, YKS03, YYX^{+}09, ZPY06, dSF03, BCJ90, CY92, DFD93, GDJ94, HK93, KK93b, LHS92, Lee93, LY93b, MLL92, MTSDA93, RS94, SST94, SMJ92, Tho93, YD94b]. Processor [BBC^{+}04, Bar98, BE07, CBE93, CW00, CYY00, CC95, CML05, DDD^{+}05, DD95, EP05, GW96a, GWL97, GR06, HK06, HWKH01, HCYD01, HW08, KN95, KBD08, LKHL03, LKKS05, LPZ98, LHSML95, LWLN97, MGQS^{+}08, MMSA94, OC05, PPR99, RTS95, SVP08, SP95, TKP00, UKY98, VM04, VKS^{+}09, WSC97, WF06, WDY98, Wu97b, WHC03, YK99, YL96, YL97, ZCO98, ZWM99, AB94, AN94, Cap92, CD94, CNNS94, GR94, GM94, KDL91, KLDR94, Mar93, ML94, SC92, SC94, SST94, SF92a, SL93a, SMS93, SL93c, SA93, WC90, WW92, YW93]. processor-cache [SL93c]. processor-time-minimal [Cap92, SC92]. Processors [AF05, BLR03, BF04, DF99, GY95b, HTPS02, HC97, JR03, LPE^{+}99, MBM98, SF08, SJPL08, SCY98, WSB09, Aga92, Aln94a, Aln95, HK93, YG94]. Produce [TK96a]. Product [DAA97b, DAA00, FE97, HC09, KWH03, Li07]. production [ATG92, AG96]. Products [EF95, LKHL03]. Profiles [RMO^{+}95]. Profiling [Hol98]. Program [Abr97, AK98, AN93, KP09, BCBzC92, MS94a, MCH^{+}90, RM90, TRS90]. Programmable [ZLKK07]. Programming [AAD08, AJMJS03, AGL^{+}98, BM00a, CdMB05, JZ04, KBC^{+}01, LCB96, OB00, PG01, PW95, RNR^{+}03, SK95, TSG09, YYX^{+}09, BS95, CR90, HQL^{+}91, HLV94, KMT91, WG90]. Programming-Based [AAD08]. Programs [CF00, DHN96, FO05, GSS96, Hol98, KA99, LRG99, LMT98, MF01a, NE01, OXL06, PH02, WNKS96, WBO^{+}01, ZH99b, ADM92, Bil94, BE92, ClW91, CR90, Fos91, Gab90, GW94, GW96b, GP92, HN90, Lar93, LC91a, LNP94, MKH91, RS94, RK94a, RK94b, SLY90]. Progress [WWWA09]. Progressive [SP03, ZZMN07]. Promoting [AD08]. PROMPT [HRG00]. proof [CG08]. Propagation [CH98, DYJ97, Jia95, SH97, SOM05, TLGP97, MLL92, Rao96]. Propagations [HM98]. Properties [Abr97, CSH00, DAA02, DS05, DCF95, EAL91, EAK95, HC99a, Pre99, Sto97, Tsa03, DT94, Ost90]. Property [SyFL99, BR91, LC94]. Proportional [FLZ09, LLY04, PC07, ZX04]. Proportional-Delay [LLY04]. Proportional-Share [FLZ09]. Protecting [WZP^{+}03]. Protection [WS03, WLZ08, WFS09, XRY09]. Protein [TAKB06]. Proteins [FARH02]. Protocol [ABS01, CBD^{+}01, CHHC06, DZ04, EBS04, FPGAD08, GCCC^{+}04, Gen00, GP99a, GJDA06, HRG00, HSLA05, HJB^{+}09, Jia95, JZXX99, KL02, LDCO08, LLY07, LC02a, LW09c, LK04, MEKOT03, MZA02, MTK06, PK00, RE09, SH97, SPC^{+}02, TF96a, WO04, XJZZ00, YWY08, YK03, ZMMS08, ZL07b, ZKB08, AB91a, KP93a, LG90, YTB92]. Protocol-Centric [PK00]. Protocols [AEA97, AK99a, Ano04d, BRSS08, BBS^{+}09, BMPP06, CH04a, CFKR98, DW04b, FRGJ07, GY95a, GKG06, ISRS06, MLSS07, NOS99, NO00a, NO00b, NO02, OSRS06a, OSRS06b, PD95, PDH06, SRT96, TKW98, Tsa03, TT01, WCR09, XXZ03, MSMA90]. prototype [DM93, LLJ^{+}93]. Provably [HHL08]. Provenance [GM09, JBW^{+}08]. Provenance-Preserving [JBW^{+}08]. Provide [MAS08]. Providing [FZGC06, RAHM05]. Provision [CLY08a, MGA^{+}09]. Provisioning [EKOAW02, WMXZ06]. Proxies [CC03, JLDC05, LA06, TCC05]. Proximate [HNO98b]. Proximity [ZH05]. Proximity-Aware [ZH05]. Proxy [ILL07]. Proxy-Client [ILL07]. Pruned [XP07]. Pruning [CB00, DW04b, MD97, SG93]. pruning-cache [SG93]. PSA [KHS07]. PSCR [GP99a]. Pseudo [LHL^{+}08]. Publish [ZH07c]. Publish/Subscribe [ZH07c]. Pull [KLH07]. Push [KLH07]. Push-Pull [KLH07]. Puzzles [ACT06]. Pyramid [PH96, DS94, JS93]. pyramids [GM94].

QoS [ASD04, CCQ^{+}05, CWYZ09, CS02b, EKOAW02, FHA06, HSH^{+}99, HYP02, KK03b, MAS^{+}07, MGA^{+}09, NK08, RGK09, RSG06, SJKC06, TX05, WMXZ06, XHYL05, XP05, YKDV02, ZPY06]. QoS-Aware [RGK09, TX05]. QoS-Constrained [ZPY06]. QoS-Enhanced [KK03b]. QoS-Provisioning [WMXZ06]. QoS-Sensitive [CS02b]. Quadratic [CHC04]. Quality [CHL09, HH08, KSC03, LLX06, MAS08, RAHM05, ZB09]. Quantifying [HP03, NGB^{+}05]. Quantitative [Bor00]. Quantization [JR03]. Quasi [CCSC09, MS99a]. Quasi-Aggregate [CCSC09]. Quasi-Synchronous [MS99a]. Quasidynamic [KK04]. Quasiregular [LH06b]. Queries [AKSS04, DP02, JN08, LG09, LA06, SC07, XTL08]. Query [BNO^{+}01, LLX06, SKCL09, CY92, LY93b, WCS92]. Querying [DLS09, PS03, BGO^{+}97]. Question [SMH02]. Question/Answering [SMH02]. Queue [hKY08, LR96, RMO^{+}95, DC95]. Queued [HS08]. queueing [Nic92]. Queues [Che01, DPS96a, DPS96b, OW91]. Queuing [AH06, FHA06, FZGC06, PF96, SV97, SSP02, TH06]. Quiescence [DTE07]. Quiver [RS08]. Quorum [AEA97, AMPR01, AMP07, CS01a, CY95, Jou03, MTK06, NW98, TYK99, YC95, AB91a, Fu97]. Quorum-Based [AEA97, AMP07, CS01a, Jou03, MTK06, TYK99]. Quorums [KKM08].

Race [PK00]. Radars [KKC03, KCK^{+}06]. Radio [DZ04, NOS99, NO00a, NOZ01, NO02, Rav07]. Radius [ISRS06, TF96b]. RAID [HJH02]. RAID5 [Tho06, TM97]. RAIN [BFL^{+}01]. Random [BGJ06, CH08, LKK02, LLL09, LWXS06, Rav07, VB96, RS94, You93]. Randomization [JS98]. Randomized [AS00, CPX06, FRGJ07, Mit01, NO00b, RS98, UFS96, YJ97a, BL91]. Randomly [CH08, VB93]. Range [CST02, SPF99, WWWA09, ZY04, ZY06]. Range-Free [WWWA09]. Range-Join [CST02]. Ranking [PKJ97, SS96, RJ90]. RAPID [HNY02]. RAPID-Cache [HNY02]. rasterizer [Bir93]. Rate [BMR99, EKOAW02, GAG96, HY07, HPT04, JASA08, LDG04, SS08]. Rate-Based [EKOAW02]. Rate-Monotonic [BMR99]. Rate-Optimal [GAG96]. Rateless [WL08b]. Rates [HJB^{+}09]. Rather [TEF07]. Ratio [GZ09, KS01, WLL^{+}07]. Raw [MYA01]. Rayleigh [Gre98]. RDT [Tsa03]. Reaching [KA94, TYK99, WYWZ08]. Read [KDW01]. Reading [KST94]. Real [AS99, Ano98c, AA09, BÖ98, BMR99, CRN09, CS97b, CS03, EDO06, FWDC^{+}00, GMM97, HS99a, HRG00, HJS^{+}06, HSH^{+}99, HS99b, KSF94, KGM97, KMW08, KWH02, KKC03, KS01, KS03, KgCS04, LL07, LHSML95, LWK05, MZ05, MM98a, MM98b, ME95, PABD^{+}99, Ram99, SJPL08, SCK00, SR99, TL05, VMXQ04, WJLK07, WCH^{+}08, WMWL08, XZG09, XP05, XQ08, YW98, ZS95a, ZS98, ZMC03, ZMM04, ZLZN09, ZJ99, CD94, KGM96, RSS90, SRS93, SH93, SH94, SA94, SMS93]. Real-Time [AS99, Ano98c, AA09, BÖ98, CRN09, CS97b, CS03, EDO06, FWDC^{+}00, GMM97, HS99a, HRG00, HJS^{+}06, HSH^{+}99, HS99b, KGM97, KMW08, KWH02, KKC03, KS01, KS03, KgCS04, LL07, LHSML95, LWK05, MZ05, MM98a, MM98b, PABD^{+}99, Ram99, SJPL08, SCK00, SR99, WJLK07, WCH^{+}08, WMWL08, XP05, XQ08, YW98, ZS95a, ZS98, ZMC03, ZMM04, ZLZN09, ZJ99, KSF94, CD94, KGM96, RSS90, SRS93, SH93, SH94, SA94, SMS93]. Realistic [Ano04c, CRS06, LR97, SSS06, WLZN07]. Realizability [SyFL99]. Realizable [GLV06]. Reallocation [Tse09]. Rearrangeable [ÇF99a]. Reassignment [CT08]. Recirculating [ZY06]. reclaiming [SRS93]. Reclamation [GPST09, Mic04, WCLF95, ZMC03]. Recognition [CW00, GR94, YC96]. Recognition-Complete [CW00]. Recognizing [KH98, PWW00]. Recomputing [YDW^{+}09]. Reconciliation [ACT06]. Reconfigurable [BM00a, BM00b, BA97, BGOS98, BNO^{+}01, DSO02, EW97, FZVT98, HNO98a, HTPS02, wJPP97, LS96, LPZ98, LO95b, NO97, NO98, PS08, RS97a, RJ99, SGTP08, WHW05, WH01, YZW94, ZP07, Aln94a, Aln95, wJNPS97, MR92, WC90]. Reconfiguration [Ano99h, Avr99, CBD^{+}01, DLPP05, KZ96, LHSML95, LPD05, PPD03, QM94, Tze93, YR96, MS94a]. Reconstruction [Sto96, CL94]. Recorded [LL98]. Recording [GM09]. Recoverable [MP97]. Recovery [CY96b, DYJ97, LL02, MGDZ07, PS96c, SNI02a, SNI02b, VJA97, ZLKK07, JF94, KK93a, KP93a, TKT92, WFP90]. Recurrence [BAH01]. Recurrences [WNKS96]. Recurrent [GWL97]. Recursion [ZL05]. Recursion-Based [ZL05]. Recursive [CLPT02, Fu05, HCD97, HGC05, LRG99, PH02, SCL00, TC04a, YFJ^{+}01, HN90, SCD97]. Recycling [WRB09]. ReDAL [DDV^{+}07]. Redirection [CCY03, RK08]. Redistribution [CHB98, CJPW06, DDP^{+}98, GAL01, HCYD01, HCYL06, KM02, PPR99, PD99, TCR96, KN95]. Reduce [Ian97, SJKC06, AH91, ME95]. Reduce-Scatter [Ian97]. reduced [Zia94]. Reducing [KCRB03, hKY08, Kop94, QM97, RJ05, WSNA95, YCTW07]. Reduction [KB03, MR92, PP95, RP99, SS00, YR06, ZMP07, LA93, STMD96]. Reductive [CMR07]. Redundancy [Agr98, LW95b, SWC95]. Redundant [CY99, JGZW08, MB07, KGMB94, KS91]. Refactoring [ZJ03]. Reference [GPST09, HE92]. References [CHC04]. Reflected [MQ97]. Regeneration [DHP^{+}07]. Regeneration-Theory [DHP^{+}07]. Region [GLS07]. Register [LPE^{+}99, YLL^{+}07, ZLAV04]. Registers [CH09]. Regular [Ano99f, CCC05, CM95, HC09, MDSS09, PK99b, PLT00, SK02, SKB04, TC95a, GMG96, HK91, MS91]. Regularity [LCB00]. Regularization [TC95a]. Regularly [Lai00, YY95]. Regulating [SP07]. Reinforcement [ZCO98]. Reinforcement-Based [ZCO98]. Related [BBG^{+}95, PR05a, Ram95, THT^{+}97, WKS01, JR93, KSA94, WC90]. Relational [RL98, Omi90]. relations [YA93]. Relationship [HY96, LW95b]. Relationships [MT97]. Relaxed [PD00]. Relaxing [HM95]. Relay [FHA06]. Release [VM04]. Reliability [yCM98, CH92, CI92, GB00, GYS05, SJ99, SR91, SRT94]. Reliable [ABS01, BFL^{+}01, DHN95, GPST09, GKG06, HNY02, KMG03, LWC^{+}09, MN92, RE09, RHM09, ST99b, XXZ03, ZF07, HK94, LS94b]. Relocation [TS98]. Remapping [BA07, YXW03]. Remote [JKR01, LWY96, LWY93, Tho93]. Removal [KS91]. Rendering [BA07, LLH^{+}01]. Reordering [LLY07]. Repair [Her00]. Repartitioning [SKK01]. Repeated [GG94a]. Replacement [CC03]. Replica [AMY09, BRSR08, CSR^{+}09, MMJ03, SRT96, TX05, TC06, TCC07]. Replicas [KDW01, QR07]. replicate [SY93]. Replicated [HZ97, KSC03, PM02, RSG06, Tos07, AB91a, RST95, SB94b, TT94]. Replication [AJ95, BKI06, CCD^{+}09, DvdMK09, FG01, GLV06, HY96, JLDC05, LTZS06, LWY93, LSCZ07, MBTPV06, SYC03, TC04b, WC09, ZJ99, TT94]. Replication-Based [WC09]. Reporting [SZ03a]. Representation [Abr97, CDV^{+}06, EBS02]. represented [IA95]. Reputation [SCW07, ZF07, ZH07b]. Reputation-Based [SCW07]. Request [CCY03, CB03, DDV^{+}07, LS94a, RK08]. Requests [JR03, TTB^{+}00]. Requirement [KPR05]. Requirements [HYP02, SSRV99, Uht92, GO93, MS93, SMS93]. Rerouting [NSZ02, SDDY00]. Rescheduling [SSZ06]. Research [RRX09]. Reservation [CS02b, SP05, XLW^{+}06, ZMMS08]. Reservation-Based [SP05, ZMMS08]. Reservations [RRX09]. Reshuffle [Din01]. Residue [BM00b, PP95]. Resilience [TJ07]. Resilient [AOK09, CWLR09, CC93a, DAA00, NLM90, SX07, WL08b, YK09, LW95a]. Resistant [BSS09]. Resolution [GFG^{+}99, SP05, WP00, XRR00]. Resolving [HLH09]. Resource [ANN95, AOK09, AMSK04, BCR98, CXN06, CNT05, DP06, Din06, GAG96, HWWX99, JJ09, KZN07, KSME08, KyK09, KCW09, KPR05, LJCL08, MEKOT03, RC95, RK08, RH04, SKJ07, SBK02a, SBK02b, SRS93, SRD08, TF96b, VVR07, XCZ02, XL08, XQ08, YMP08, ZWX06, PJC93]. Resource-Aware [VVR07]. Resource-Constrained [GAG96, ANN95]. Resources [BcFGM08, DP01, FLZ09, GKK05, SJKC06]. Respect [SLH97]. Respective [FMR07]. Response [CN04, KA09, LLTW08, LLX06, Var01, TRS90, WCS92]. Responsive [LAV03, Sun02, WLL^{+}07]. Restart [CLS04]. Restoration [AYA09, FCF00, MAJ^{+}07]. Restoration-Based [MAJ^{+}07]. Restricted [FZVT98, GZ09, NO97, CCJ02]. Restructuring [CK08, DKKS04]. Results [BCL^{+}05, CCY96, FCF00, Fei05]. Retiming [CDR98, CS97a, PS96a]. Retrieval [LC04, MZA02, SC07, ZYKG07]. Retry [CF01]. Reuse [PDH06]. Revenue [LJCL08]. Reversible [LF03]. Reviewers [Ano99a, Ano00a, Ano01a, Ano03a, Ano04e, Ano05a, Ano06, Ano07b, Ano08b, Ano09a]. Rewriting [SF07]. RFID [MLSS07]. RH [Zia94]. RHiNET [KWOA05]. RHiNET-2 [KWOA05]. Riding [LYW08]. Right [SF09]. Ring [ABC^{+}01a, BK09, CC93a, LW95b, TCS97, UKY98, ZYC95, ZY95]. Ring-Based [ZYC95]. Ring-Connected [LW95b]. Ring-Like [BK09]. Rings [Ano99f, HGC05, HLH04, KY97, LH01, PK99b, SCL00, YCTW07, VB93]. RIPS [SW96]. Ritz [Gre98]. RLE [EAF00]. RLE-Compressed [EAF00]. Robin [KSP02, LMS04, ZY07]. Robust [AKNR^{+}04, CPX06, EVW07, WLL^{+}07, ZH07b, LY94]. Robustness [AMSK04, PR05b]. Role [CHC09]. Role-Based [CHC09]. Rollback [CY96b, TKT92, TKW98]. Rollback-Recovery [CY96b]. Rolling [AT01]. Root [Fei05, CF94, LH93]. Rotation [SY97]. Rotations [MBM98]. Rotator [Cor92]. Round [KSP02, LMS04, ZY07]. Round-Robin [ZY07]. Rounds [Gen00]. Routable [YW00, YW03b]. Route [PDH06, SCK00]. Routed [BP98, CFKR98, FR96, FF98, HÖ00, HK95, KLS00, LMN95, RMC95, SS07, SCL01, jTM96, TG96, TPL96, TLGP97, TWH99, XGN97, ZL05, MXEN94, jTM97]. Router [CCQ^{+}05, DSY99, MBW02, PGBI03, SDFV96, WHM09, YKDV02]. Routers [BC99, Chi98, HDF07, LBC03, Tze04, Tze06, WS03, WFS09]. Routes [MAJ^{+}07, WZP^{+}03]. Routing [AM95, AS00, Ano98b, Aro00, BcFGM08, BRS07, BC06, BHL^{+}07, BC96, BCR98, BRS97, BC95, ÇF99a, CC97, CC01, Chi00, CKWC08, DSY99, DDY99, DS03a, DZ04, DS05, DY05, Dua95a, Dua95b, Dua96, Dua97, DP01, FYS05, FG06a, FG06b, GZ06, GZ09, GY95a, GN96, GJDA06, GO97, GG95, GYS05, GS03, HÖD99, HW97, HH95, HZ96, JZXX99, JKA07, KP01, KLC97, Kuc01, KPC09, Lan95, LO95a, LC96b, LW09a, LWC^{+}09, LMN95, LW09b, LW09c, LZ05, LSRT06, NOZ01, NOZ02, NS95a, NSZ02, OKSA01, PHKC09, PSK99, PDH06, PG07, QNR99, RS97a, RS98, RRRM09, RE09, RS97b, RLD03, Sah00a, SRD04, SyFL99, SC07, SCP99, SD00a, SGL06, SL01a, SL01b, SSsLY03, THH08, TW98, TL06, TR06, TW00, Tze06, UFS96, VDS99, VSD01, VB96, VS11, WL97, WO04, WWLS08, WMN99, Wu98, WA99, Wu00]. Routing [Wu02, WYD07, Xia01, XLPH06, XJZZ00, YW99, YW03b, YW05b, YWY08, AV94, BR91, CS90, DA93, Dua93, GPBS94, LMN94, MS93, MC93, OS94b, PGDS94, PGFS94, SC93, ST93, SCD97]. Routings [KWOA05]. Row [LC96b, NO98, SP93]. Row-Column [LC96b]. row/column [SP93]. Rows [BOPZ04]. RS [BGBP01]. RS/6000 [BGBP01]. rStream [WL08b]. Rule [WMB96]. Rules [BS08, JZ04]. Run [LCB00, LLY05, RP99, RRFH98, WBO^{+}01, RWF94]. Run-Time [LCB00, RP99, RRFH98, RWF94]. Running [AV96, HS98a]. Runtime [ASS95, BCG04, HW08, LP07, LLGS09, MF01b, PSC^{+}95, SW96, TEF07, Wu97b, XC01, YZZ00].

S [HK98]. S-to-P [HK98]. SACCS [WDCK04]. Safe [Mic04]. Safety [Kin06, SJ99, Wu98, Wu00, Xia01]. Sampling [GLY07]. SAMR [SCP02]. Sapphire [BES06]. SARA [JASA08]. Satisfying [TTB^{+}00]. saturation [SS90]. Scalability [AF05, BG02, DF09, GKS95, JW00, KW08, LZTY09, SR94, GK93]. ScalaBLAST [ON06]. Scalable [AGGD04, AGGD05, Add97, AK99a, AGL^{+}98, AAB^{+}00, BBC^{+}95, BS96, CCSC09, CF08, CPhX04, CHHC06, CYD98, CMDP09, DPH08, DZHG04, FBD96, FMG02, GWL97, GKK97, GKG06, HK98, HDF07, JTC08, KSWR03, KSA94, LCGC07, LXL08, kLCC^{+}06, LJLN07, LNO^{+}00, LXN07, LW09b, LWN98, LQZ09, MD97, MG09, MJ06, ON06, PAM95, PKJ97, PG07, RS08, SH95a, SYC03, Tze04, Tze06, WDCK04, WL00, WH03b, YN00, ZL07b, ZH07b, ZP07, GP93, KCPT96, LB94, MB92]. Scalar [BWS^{+}05, GS91]. Scale [BB05, BG09, CY00b, DvdMK09, EDO06, GMCB01, GY09, HL09a, KMG03, KCW09, LCGC07, LC95, LK04, MY07, MA01, MMJ03, SkLC^{+}03, VVR07, WHM09, XHYL05, ZYKG07, SG93, YTB92]. Scale-Free [GY09]. Scaling [FZVT98, HWWX99, KSME08, SGL06, WZZ09]. SCALLOP [CHHC06]. Scan [YLW07, Yi09]. Scan-Based [YLW07]. Scatter [Ian97]. Scatternet [LSW04]. Scatternets [TSK06]. Schedulability [AA09, Bak05, BCL09]. Schedule [LDCO08, SC94]. Scheduler [CC95, SKJ07]. Schedulers [BCF^{+}08, SF09]. Schedules [BOC09, COS00, Ros02, JR94]. Scheduling [AS99, AK98, AK99b, AAD08, AM06, ABK98, Ano04c, BA04, BcFGM08, BBC^{+}04, BKS03, BBD00, BCL^{+}05, BCL09, BMR99, BOC09, BE07, CCQ^{+}05, CG08, CRS06, CS08, CS97a, CV08, CRN09, CYY00, CKC08, CJPW06, CMR07, CFR99, DA98, DDP^{+}98, DWX09, DÖ02, DVV07, DZH05, DMKJ96, DNSC09, EK95, EDO06, FYP07, FFPF05, FH03, GRY07, GKK05, GMM97, GVV09, GRT97, HKL00, HHL08, HS08, JSWB97, KSP02, KGM96, KA06, KB06, KLH07, KA96, KC98, LLTW08, LKHL03, LZ08, LMS04, Li08, LWJ06, LWXS06, LDG04, MLS94, MM98a, MM98b, Mha09, MF01b, PAM95, PM96, RvG02, RRX09, Ram95, RLW^{+}07, RJ96, RBSP02, SD04, SS94, SJPL08, SZ02, SZXS05, SP98, SM03, SW96, SS05, SSS06, SP05, SCW07, SS00, SSZ06, TSAL97, TGV08, TD01, TTB^{+}00, THW02, VRKL96, VM04]. Scheduling [VVR07, VKS^{+}09, WR04, WWLS08, WSB09, WMWL08, WF03, WTCY95, Wu97b, WSG01, WYJ^{+}04, XU01, XZNX08, YG94, YF97, YKS03, YvdRC05, ZLAV04, ZSMF01, ZFMS03, ZY04, ZX04, ZMC03, ZMM04, ZCO98, ZWM99, AM93, AMAM94, DR94, EG93, Fos91, HAR94, KLDR94, KS93, LC91b, Lil94, ML94, OD93, PLW96, RSS90, SL93a, SL93b, SL93c, TN93b, YJZ97, ZLE91, ZA93]. Scheme [BHJ02, BG09, CCSC09, lCL95, CC01, CC98, CC99, CL05, DS05, DWX09, EKOAW02, FYP07, FT97, FI95, HCHM09, HS98b, HPH08, KCD07, LCL03, LJW^{+}07, MCL^{+}07, PAM95, PK99a, SJd^{+}09, SFP03, SZ95a, TS98, TJ08, TD01, WDCK04, WX07, XTL08, YYS97, YGE06, YG08, vdMDM07, AM91, CA93, HMR94, JS90, KDL91, LHS92, LC91b, MB92, SB94b, TH93, TN93b, YK92]. Schemes [AJ95, ADG06, CSR07, DF99, GBD07, HS99a, HW97, JO95, PSGD05, PPD03, SS96, Tos07, TYK99, VB96, WT08, CYW94, CO94, RJ94, SL94, SH93, ST93]. Schur [ME95]. Scientific [CH04b, CMBAN08, HT06, MLW06, NKP^{+}96, PF08, SkLC^{+}03]. Scope [JGZW08]. Search [AfAGR00, CW06, Che95b, CLY08b, CJLN09, CBDW96, JTP^{+}08, JGZW08, KLH07, LLSZ08, LLWC09, MD97, PWW00, RBSP02, SVP08, WX07, WZZ09, WSG01, ZYKG07, ZH07a, ZH06, AM90, CS90, KK94]. Search-Based [KLH07]. Searching [MTK06]. Second [ZCL04, MCH^{+}90]. Second-Level [ZCL04]. Secondary [WRB09]. Secret [NW98]. Section [ACM08, AAB06, ABC01b, CRS06, GZ03, IT07, ON02, OSRS06a, OSRS06b, PP05, SR99, Zha03, HK91]. Sections [HK06]. Secure [AKNR^{+}04, GBC^{+}07, HCHM09, KYB08, Lee06, MMJ03, STY09, SGB08, WCBX06, ZZMN07, vdMDM07]. Securing [PZZ09]. Security [BHL^{+}07, KPC09, LAV03, LK07, SF07, XQ08, Zha03]. Security-Aware [XQ08]. Segments [CW02b]. Selectable [HJB^{+}09]. Selecting [Qua01]. Selection [AFAGR97, AMY09, BW96, CH04a, CB03, GS03, KCW09, NSU97, RS97a, RS98, SCK00, WH03b, YK09, YR06, ZF07, BLO^{+}94]. Selective [CKC08, LA93]. Self [CDV^{+}06, DW04b, DAMK06, DB08, DIM97, DS03b, FG06a, KY97, Kar01, LH03, MS99b, SP07, TLM04, TH06, TNPK01, TK96a, UKY98, WLZ08, YW99, YW00, YW03b, Fos91, SH95b, TN93b]. Self-Control [TK96a]. Self-Optimization [TK96a]. Self-Organizing [CDV^{+}06, SH95b]. Self-Pruning [DW04b]. Self-Regulating [SP07]. Self-Routable [YW00, YW03b]. Self-Routing [FG06a, YW99]. self-scheduling [Fos91, TN93b]. Self-Stabilizing [DAMK06, DB08, DIM97, DS03b, KY97, Kar01, LH03, TH06, TNPK01, UKY98]. Self-Synchronization [MS99b]. Self-Tested [MS99b]. Self-Tuned [TLM04]. Selfish [KHS07, LTZS06, LSB^{+}07, LW09a]. Semi [ABRY03]. Semi-Oblique [ABRY03]. semijoins [CY92]. SenCar [MY07]. Sensing [CLW03, FG06b, RLW^{+}07, XLPH06]. Sensing-Covered [FG06b]. Sensitive [CS02b, WD06, YK03]. Sensor [AYA09, AD08, AD09, BK09, BBS^{+}09, BS08, CHA07, CYW08, CBM^{+}07, CY06, CPX06, CH08, CTF09, DLS09, DWX09, GLY07, GBC^{+}07, HSLA05, HCHM09, ISRS06, JN08, KZN07, KPK09, KS08b, LDCO08, LCN^{+}07, LRS02, LWJ06, LWXS06, LH06b, LWP07, MGZN07, MCL^{+}07, MY07, MZT08, MZA02, RLW^{+}07, RE09, SKS02, SJd^{+}09, SRZF04, TX08, TN08, WT08, WLZ08, WWWA09, WLZN07, WCD08, XCZ08, XTL08, YLW07, Yi09, YGE06, YYY09, YKP08, YG08, ZS09, ZPY06]. Sensor-Actuator [RE09]. Sensors [LWJ06]. separable [SP93]. Separating [BOPZ04]. Separation [BPT03]. Sequence [JTP^{+}08, LSVMW07, MQ97, RA04, YFM98, CY92]. Sequence-Search [JTP^{+}08]. Sequences [CCSC09, MDL06]. Sequencing [Bar98, BGM94]. Sequential [BGJ06, CHJ^{+}07, DDS95, DS96, Qad03, QCC99, SZ02, HMW93]. serializable [AG96]. Series [DL02, LCN^{+}07, TR04, ZCSY08, MM96]. Series-Parallel [DL02]. Server [ASB02, AFM02, CB05, CT08, CGL07, CYD98, DDV^{+}07, GB06, LLY04, QR07, RSG06, RJ05, SBK02a, SBK02b, VR05, CR94, ICT93]. server-based [CR94]. Servers [GB00, GMCB01, KK03a, KCD07, LL02, LKKS05, LLA^{+}06, RAHM05, RNKZ03, SD04, Tse05, WZP^{+}03, ZRS^{+}05, ZX04, ZWX06, KGM96]. Service [AOK09, AMH08, CPM07, DHN95, DAMK06, DS03b, FZGC06, GMS09, KKS07, KSC03, LJLN07, LLA^{+}06, LZTY09, MAS08, PS08, RAHM05, RE09, SY07, SL09, SS07, TJ08, YWY08, ZF07, ZX04, ZWX06, ZZMN07, ZJ99, AT07, CR94, CSR^{+}09]. Service-Centric [YWY08]. Service-Driven [RE09]. Services [AK99a, CLY08a, DZHG04, GRY07, KSC03, KSWR03, LAS04, NGB^{+}05, RS08, RD09, SYC03, WZZ09, ZH07c]. Session [ZWX06]. Session-Based [ZWX06]. Set [AMP07, BSCB09, DW04a, DMR01, DP01, LH03, WM95, Wu02, WCDY06]. Set-Associative [WM95]. Sets [JB01, KWL^{+}09, OZ96, PPR99, RD98, SSZ02, Sto04, Wan04]. SFC [SCP02]. Shape [GDK09, HS02]. Shaped [RR02]. Sharc [US04]. Share [FLZ09]. Shared [AD98, AGGD04, AAS03, AKN95, Bor00, Cha96, CH04b, DDS95, DS96, FB01a, FT97, GP99a, GMR98, Hol98, HS98b, KH04, KL01, KA05, LP96, LT97, LBC03, MA01, McK98, MP97, PC05, PPBSA97, Qad03, QD05, RGK09, RD98, SLEV03, SNI02a, SNI02b, SZ95b, TF96a, US04, VGGD94, WH95, YL97, ZYC95, AH93, ABJ^{+}93, And90, BIA^{+}97, CR90, DC95, Don91, Geh93, GH93, Gup92, IT93, IC92, KCPT96, Lil94, ML94, SL93c, WFP90, YJZ97, ZLE91, ZSLW92]. Shared-Bus [GP99a, LP96]. Shared-Memory [AGGD04, AKN95, DDS95, DS96, FT97, GP99a, Hol98, HS98b, KL01, LT97, MA01, McK98, PPBSA97, Qad03, QD05, SLEV03, WH95, YL97, ZYC95, AH93, DC95, Gup92, IT93, KCPT96, ML94, SL93c, YJZ97]. shared-money [And90]. Shared-Nothing [RD98]. Sharing [BCdSFL09, DYJ97, GG09, GP99a, HKS^{+}07, KCRB03, KA06, KyK09, LKKS05, LL06a, LL06b, LYW08, MTL95, NW98, RS08, SH96, VR05, WX07, DY93, GDI93, HK93, KK92, LY94, SH93, SH94]. shift [LO95b]. shifts [RS90]. Short [GZ06, STY09, KGMB94]. Short-Lived [STY09]. Short-Path [GZ06]. Shorter [UFS96]. Shortest [FH97, LR96, ZH98, SCD97, TR93]. Shot [FMR07]. Shrinking [JL99, JS93, SKF94]. Shuffle [FG06a, BCH94, Pad91]. shuffle-exchange [BCH94, Pad91]. Side [TCC05]. Signal [HXA96, KKC03, DFD93]. Signature [CCSC09, TC07]. Signature-Based [TC07]. Signatures [NW98]. sim [RFDS97]. SIMD [AGWFH97, AS96, BCJ90, CFW98, KK94, Nas93, NSD^{+}91, NSD93, PH96, RS90, SR98, SW95]. SIMD/MIMD [BCJ90]. SIMD/SPMD [NSD^{+}91, NSD93]. Similarity [WZZ09]. Simple [BAH01, COP00, EW97, Hsi03, KM01, KAY^{+}06, SC93]. SimpleFit [MYA01]. Simulated [CFW98, HM95, LL96, Soh95, BJS90, EG93, NZ95, WCF91]. Simulation [BT00, BG09, CCP95, DHN96, FZVT98, GY95b, NL02, QCC99, Qua01, QS03, SSP^{+}09, SF09, SE98, TK96b, WHL95, XC04, HN93, HE92, HB92, Kum92, KH93, LL90, Nic92, RB90, ZL96]. Simulations [MLW06, Sah00b, SF08, SGTP08, NGL94, PGFS94]. Simulator [PPR95, RFDS97]. Simultaneous [LPE^{+}99, FC91]. sine [MM96]. Single [CLW03, DZ04, GBD07, GS08, NO00a, SL01a, BGM94, Rao96]. single-fault [Rao96]. Single-Hop [CLW03, DZ04, NO00a]. single-level [BGM94]. Single-Packet [GS08]. Single-Path [SL01a]. Single-Path/Flooding [SL01a]. Single/Multiclass [GBD07]. sites [Tho93]. Situ [MCL^{+}07]. Size [DS03a, LZ02, LH01, OPZ99]. sized [Pad91]. Sizes [Sca99, YA93]. skew [EA93, WYTD93, WDY93]. Skip [WL08a]. Slack [MZ05, ZMC03]. Slave [BBC^{+}04, BLR03, KA06]. Sleep [DWX09]. Slices [MGQS^{+}08]. Slicing [MSG07]. sliding [SA93]. Slowdown [FB01a]. Small [HLL09, LLSZ08, YM95]. Small-World [HLL09]. Small-World-Based [LLSZ08]. Smaller [KP96, UKY98]. Smart [CB03, JGA08]. Smoothing [KgCS04]. SMP [YZZ00]. SMPs [LK04]. SMT [BG02, WSB09]. SNAP [DM93, MLL92]. SNAP-1 [DM93]. Snapshot [LCN^{+}07]. Snapshots [HMR99, NX95]. Snooping [LK04, SPC^{+}02]. SOCNs [WL00]. Soft [HJS^{+}06, KGM97, KgCS04, CD94, KGM96]. Software [CDR98, CL05, EBS04, GAG96, JJ09, KIBW99, KABK03, KA05, LPE^{+}99, LBC03, MBTPV06, PBA03, SDDY00, WDY98, XGN97, ZLKK07, ANN95, WF94]. Software-Based [SDDY00, ZLKK07]. Software-Directed [LPE^{+}99]. Solution [BSCB09, Che01, LC99, Liu08, CARW93, You93]. Solution-Adaptive [LC99]. Solutions [Bar98, BAH01, CCQ^{+}05, LLY07, Sto96, KST94]. solvable [YK96a]. Solve [CHC04, FMR07]. Solvent [FARH02]. Solvers [SZ04, WH95]. Solving [KBD08, Liu08, MSG07, MBM98, NCV05, PK95a, PK95b, THT^{+}97, O’H91, RJ90]. Some [Lee06, THT^{+}97, TC95b, O’H91, WC90]. SORD [AOK09]. Sort [LB00b, OPZ99, AOB93, WDY93]. Sorted [Che95b, HNO98a]. Sorter [PK99a]. Sorting [BGO^{+}98, CS92, DSO02, DCSM96, FE97, HW97, LB95, NS95b, OPZ99, RS97a, RS98, CO94, GG94b, Lin93, MN92, XB93]. Source [CTF09, GYS05, MM07, XZG09, YLL^{+}07, CSC07]. Source-Code-Correlated [MM07]. SP [BGBP01]. SP2 [HXA96, MF01b]. Space [AB07, BA07, CDV^{+}06, CL05, KABK03, KYD^{+}07, LB00a, LP07, MCG08, RA04, SP07, WCLF95, KM91]. Space-Time [LB00a, LP07]. Spacefilling [PB96]. Spaces [BCdSFL09, GAK03]. Spam [CWLR09, LZR09]. Spam-Resilient [CWLR09]. Spanners [ALW^{+}03]. Spanning [Ano99h, Avr99, CTS96, DPN09, EVW07, KPK09, KWH03, LS96, LWN98, YCTW07, GM94]. spare [AM91]. Sparing [TM97, Tho06]. Sparse [BW96, ÇA99, FJY98, GKK97, UZCZ97]. Sparse-Matrix [ÇA99]. Spatial [JN08, KCRB03, WDY98]. Spatiotemporal [HSLA05, LWP07]. Special [ACM08, AAB06, Ano97d, Ano97b, Ano97c, Ano98c, Ano98b, Ano01b, Ano01c, Ano01d, Ano02b, Ano03c, Ano04c, Ano04d, Ano05c, Ano07c, Ano08c, Ano09c, Ano09b, ABC01b, CRS06, GZ03, IT07, ON02, OSRS06a, OSRS06b, PP05, SR99, Zha03]. Specific [BJM^{+}05, GW96a, HP06, Pak07, PHKC09, Pre99, BGO^{+}97]. Specification [FB01b, GCCC^{+}04]. Specified [PSC^{+}95]. Specifying [HW91, SPC^{+}02]. Speculated [SCL05]. Speculation [KA05]. Speculative [BF04, CL05, CASM07, KL01, MGQS^{+}08, RP99, Soh95, TKVD02, VGSS01, MR94, WCF91]. Speed [CBD^{+}01, Chi98, FZGC06, Li08, MNM04, ZMC03, Ant94]. speedup [KH93]. Speedy [Tze06]. SPIFFI [FBD96]. Spilling [CHJ^{+}07]. Spin [JH97, KM01, LLS06, And90, ZLE91]. Spline [GM97]. Split [Agr99, SS08, SM03]. Split-Path [SM03]. Split-Phase [Agr99]. Splitting [MLSS07, XB93]. SPMD [CG02a, CG02b, NSD^{+}91, NSD93]. Spoken [GR94]. Spot [LC95, OKSA01, ZYC95]. Spots [WSNA95]. SPP1000 [AD98]. Spreading [JL99]. Square [BGO^{+}96, LZ02, LH93]. square-root [LH93]. Squares [KP93b]. SSL [KCD07]. SSW [LLSZ08]. Stability [FMG02]. Stabilizing [DAMK06, DB08, DIM97, DS03b, KY97, Kar01, LH03, TH06, TNPK01, UKY98]. Stable [Kin06, PK99a]. Stack [SSP^{+}09, WM95]. Stage [BOC09, HK93]. Staggered [Vai99]. Stale [Dah00]. Stamp [XC01, Var93]. Stampede [RNR^{+}03]. STAP [HWWX99]. Star [AAD97, BDL95, BCL^{+}05, CTS96, CC97, ISAZM09, SS96, SBS98, SWC95, TCS97, YV98, dBL98, BFP96, DT94, FA94, LB94, Lat94, MS92, MJ94]. Stars [DS03a, MR06, Sah00a, Sah00b]. Start [CLS04, SY98]. Start-Up [CLS04, SY98]. Starting [PK92]. State [CLJ^{+}04, LZ08, SKB04, SNI02a, SNI02b, THH08, TK96a, YL08, MS94b]. Stateless [DZHG04]. States [Lai00, UKY98]. Static [CD94, GvG06, KBC^{+}01, LWC^{+}09, NLW99, PP96, RWF94, SS00, WLZ08, WWLS08, LK94, SB94b]. Static-Dynamic [SS00]. Stations [XLW^{+}06]. Statistical [BES06, CGK04, CS97b, KS03, LLY05, RD98]. Statistically [KS01]. Statistics [Yi09]. steady [MS94b]. steady-state [MS94b]. Stealing [Ros02, RH04]. Steering [PSGD05]. Step [TC95a]. Stitch [KSP09]. Stitching [KS08b, KSP09]. Stochastic [BHL^{+}07, CLB08, GvG06, TS98, YXW03, BCBzC92, KS93, JASA08]. Stop [CD08]. Stopping [DGFHR03]. Storage [AMS97, CGM05, FRGJ07, HNY02, KDW01, Kin06, LLL09, MKR00, MR03, MJ98, MJRS06, Var01, XTL08, ZBJ^{+}05, ZJWX08]. Storages [XRY09]. Store [Dua96, WYD07]. Store-and-Forward [Dua96]. Store-Carry-Forward [WYD07]. Stored [LAV03]. Strategies [BBC^{+}04, GB00, GKK05, GLV06, LLGS09, MD97, SP95, TCO01, TX08, VVR07, WLR93, BL91, CV92, LY94, Lil94]. Strategy [BKS03, CG08, CW00, CPM07, DP02, GBD07, KKGS01, MTL95, YL97, AGE94, HC92, SC93]. Strategy-proof [CG08]. Stream [GN06, RNR^{+}03, RGK09, SKCL09, YY95, YYX^{+}09].

Stream-Oriented [RNR^{+}03]. Streaming [BSS09, CZLM09, DF09, ILL07, LJLN07, LSVMW07, PS03, TJ07, TJ08, WL08a, WLL08, WL08b, YM09, YK09, ZL07a, ZXZ^{+}09, ZX04]. Streams [BHJ02, CW02a, CH07]. Stretch [GZ09]. Stride [DS96]. String [ACT06, BM00b]. Striping [HJH02]. Strong [HC09, JS98, Kar01, GW96b]. Structure [BW96, DPN09, JJ07, TAKB06, Sin92]. Structured [ASS95, BRTM09, CT08, HY01, HZ96, LP07, PB96, PDH06, PZZ09, SX07, WH95, ZCSY08, Bil94]. Structures [CAZ04, CSR07, DB06, HLL09, HALT95, PR05a, EA93, GDJ94, HN90, LHS92, MS91]. Structuring [SM94, AN93]. Studies [ZWM99]. Study [AD98, CY00b, CGL07, Fei05, LS06, MTM02, NN96, SSRV99, DT94, DI95, EMS90, KH93, LY94, SLY90]. Studying [CKK^{+}04]. Style [GKG06, CR90]. Subarray [Par01]. Subcube [lCL95, CT97]. sublinear [KST94]. Submesh [yCM98, CH01, CC99, KY98]. submeshes [CT94]. Subnets [WYWZ08]. Subnetworks [ASD04]. Suboptimal [DD95]. Subscribe [ZH07c]. Subscript [SK95]. subsequence [LL94]. substitutional [TC94]. Substrate [HKS^{+}07]. Subsystem [LP96]. Subsystem-Oriented [LP96]. Subtasks [TSAL97]. Successive [Gre98, PF96]. Sufficient [Dua95a, Dua96, NX95, VS11]. SUIF [MSH00]. Suite [RE09]. Summation [DS03a]. Sums [BM00b, LNO^{+}00, LNOZ03]. Super [JZ04]. Super-Programming [JZ04]. Supercomputer [Ste96, TAKB06]. Supercomputers [ADG^{+}08, WNKS96]. Supernode [GDK09, HS98a, HS02]. Superpeer [XZL05]. Superposition [PF96]. Superscalar [CC95, DF99, WB98]. Support [CCQ^{+}05, CASM07, DZHG04, GBD07, LCB00, LNYY03, MAS^{+}07, MFLX01, MX03, PSC^{+}95, RH04, SYC03, SKPS01, SSZ06, TN08, RSV90]. Supported [ZL07a]. Supporting [BS95, DR98, SY07, SZ95a, YDW^{+}09, YMG03, ZN04]. Surface [FARH02]. Surfaces [AB07, GM97]. Surveillance [LWJ06]. Survey [MP97]. Survivable [THH08]. Sustained [NK08]. Swapped [CXP09]. Sweep [GRS99]. Switch [KP01, Lai00, MGA^{+}09, NGM97, SSP00, SSP02, YA93]. Switch-Based [KP01, NGM97, SSP00]. Switched [FYP07, HÖD99, LSC95, PC96, PS96b, SJM09, VM99, WR04, Bok93, HC92]. Switches [AH06, HS08, Mha09, QNR99, TC93]. Switching [DSY99, FZGC06, HDF07, LMS04, LL06a, LL06b, LZ05, MAS08, SO95, SV97, TZ97, Tze04, YW04, YJHG06, LO95b]. Symbiotic [HY96]. Symbolic [BE98, FS00, KP09, TNPK01, vG03, Lar93]. Symmetric [CS08, EP05, LK04, SY93, TC93, HK94]. Symmetric-Key [EP05]. Symmetrical [ÇF99a, HCYL06]. Symmetries [JK99]. Synchronization [BHJ02, CY99, Che01, CS95, CS96, CLS04, FR96, Gup92, HM95, HLH04, LH01, MX03, MS99b, NL02, OS02, SH95a, SC05, SCL01, YK98, ZL07b, dB98, Arv94, OS94a, TB94]. synchronized [AC92, RS94, TKT92]. Synchronous [AV96, FR96, FH03, LL96, MS99a, PN95, SZ95a, XL96, XC04, YXW03, ZS95b, AAG94, MS91]. synchrony [RPW93]. syndromes [LS94c]. Synthesis [BB05, BJM^{+}05, GW96a, RJ96, VJ93, UEA95]. Synthesize [LKK02]. Synthesizing [AGWFH97, LRG99, SC91, CTC93]. System [ANKA99, AM06, AMP07, BM00b, CLJ^{+}04, CBE93, CT07, CF99b, DSO02, DR98, FBD96, FI95, GWYS08, HM98, HCC06, ILL07, JTP^{+}08, KGM97, LM06, LPZ98, LBS05, Lop02, MJ98, MNM04, MX03, MRT09, NN96, PH96, Par01, PC05, PS03, RMO^{+}95, SFP03, SSRV99, SC05, SMH02, SSZ06, TSAL97, TEF07, WHW05, WMXZ06, XZG09, XL08, YYX^{+}09, ZSMF01, ZF07, ZH07b, Bil94, BCJ90, CV92, DI95, GH93, KS93, LKG92, LC91b, ME93, MCH^{+}90, TV92, Tze93, VGGD94, YD94b]. System-Generated [TEF07]. System-Level [ANKA99]. System-on-a-Chip [LM06]. System-on-Chip [XL08]. systematic [UEA95]. Systems [AS99, ASB02, AJ95, AAD08, AJMJS03, AM95, AMPR01, ABS01, Ano98c, Ano07c, Ano08c, ADD^{+}02, BQF99, BDvD98, BGBP01, BKS03, BBD00, BP96, BP98, BMR99, BJM^{+}05, BHJ02, BG09, BHK^{+}97, BE07, BRTM09, CW06, CS98, CS01a, CS01b, CS02a, CCY03, CG08, lCL95, CT02, CT08, CY96b, CRN09, CYY00, CGL07, CLS04, CYD98, DYJ97, DMR01, DHN95, DHP^{+}07, Din06, DL02, EAK97, EBS04, FRGJ07, FH97, FZGC06, FG06a, FO05, GCCC^{+}04, GMM97, GBD07, GVV09, HL08, HHM^{+}00, HSH^{+}99, HCD97, HT07, HNY02, HW08, JL99, JNGS06, JO95, JJ09, JW00, KHM05, KMG03, KL99, KLH07, KSME08, KCW09, KH97a, Kum92, KMW08, KBD08, KK03b, KC98, LKHL03, Lee06, LZ08, LJLS09, LT97, LLS06, Li07, LXL08, LYW08, LLL09, LXL^{+}05, LLX06, LWJ06, LS06, LWK05, LC02b, MKR00, MZ05, MM98a]. Systems [MM98b, MMJ03, MG09, MOFD05, MROD07, MP97, MS99b, MJ06, PAM95, PR05a, Par95, PBA03, PP95, PABD^{+}99, PS96c, PPR95, QM97, RGK09, SO95, SZXS05, SJM09, SF09, SSP00, SCO^{+}07, SP03, SJ99, SvVB05, SPF99, Sun02, SZ04, SS09, SR99, TF01, TT01, TF96b, Var01, VV99, VVR07, WCLF95, WXLZ06, WCBX06, WJLK07, WL00, WMWL08, WDY98, XHYL05, XL08, XB98, XRR00, YJ97a, YJ97b, YW98, Zha03, ZS98, ZMC03, ZMM04, ZH05, ZH06, ZJWX08, ZP07, ZCO98, ZWM99, dSF03, vG03, vDSP96, ATG92, AC92, AMAM94, AG96, Arv94, CARW93, CR94, CO95, CH92, CTC93, CYW94, CPA93, CT94, DC95, EMS90, Fu97, GMG96, Gup92, Har91, HK93, IK93, ICT93, IC92, KP93a, KK93b, KE90, LS94c, ME92, MB94, MSMA90, MMSA94, OSS93, OS94a, Pan93, RSS90, Rao96, RJ94, SST94, SRS93]. systems [ST91, SH93, SH94, SM94, Sin92, SW92, TKT92, VJ93, VJ94, WC90, WS93, WM93, WG90, YJZ97, YK92, ZLE91, Zia93, Ano02a]. Systems-on-Chip [BJM^{+}05]. Systolic [CW02a, EAF00, LSBS98, MF96, SH95b, BW94, Cap92, IS90, LK90, SC92]. systolic-based [BW94].

Table

[Ano00b, Ano00c, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, RBSP02, Tze06]. Tables [RHM09]. Tag [MLSS07]. Tag-Splitting [MLSS07]. TARA [KZN07]. Target [LWJ06]. Targets [KK03a]. Task [AS99, AK98, Ano09b, CKC08, CCD^{+}09, CYD98, DNSC09, FH03, GvG06, HKL00, HÖ99, HW08, HC97, JR03, JL99, JJ09, KA96, Lat94, LS97, LKHL03, Lee06, Li08, PLW96, RvG02, RSB97, RRG07, SS05, SSS06, SJ99, TGV08, THW02, YF97, YYS97, YD95, CO95, DC95, DK92, GY93, MKH91, SS94, SW92]. task-based [DK92]. Tasking [BBC^{+}04, SMBT90, STMD96]. Tasks [AAD08, ACD^{+}09, BA04, BCF^{+}08, CFR99, EK95, GMM97, HP07, KA06, LWK05, PH05, Ram95, Ros02, SJPL08, GO93, KK93a, YG94]. TCP [LLY07, FYJ^{+}09, WFS09]. TDMA [CLS04, LDCO08, WWLS08]. Team [BKB96]. Technique [CY96b, CHB98, CB00, CN02, CN04, Deb96, DDV^{+}07, GAK03, HCYD01, KA09, KHY09, KAY^{+}06, KA96, MZ05, MAS^{+}07, PF96, Rob04, SX03, TL06, CTC93, KGS94, MKH91, RM90, SL93b, TN93a, TC94]. Techniques [Ano04c, BB05, CRS06, JZXX99, KB06, NZP03, PP96, PBA03, PK04, SC07, SJM09, SZ03a, CS94, GS91, GB92, KN95, RS91a]. Technological [BP96]. Technology [BBR07]. template [SSG91]. template-based [SSG91]. Templates [ADD^{+}02]. Temporal [CW06]. Teng [YYX^{+}09]. Termination [DTE07, LT97, TT01, XL96, LW95a]. Test [FI95, PW95, RP99, HISS94, KKP91, PKK93, WT92, KKP91]. Test&Set [ST99b]. Testbed [NN96, VDS99]. Tested [MS99b]. Testing [BE98, HALT95, KR00, LC94, Pak07]. tests [Uht92]. Text [HM98]. Textured [HH95]. Their [HCD97, LW95b, SSP00, UZCZ97, WMN99]. theorem [WY94]. Theoretic [BHL^{+}07, KHS07, SZ08, YM09, YK09]. Theoretical [ASB02, KA09, TKW98]. Theory [CMR07, DHP^{+}07, DD98, Dua95b, Dua97, DP01, DLPP05, FF98, GBD07, IK93, LL06a, Dua93, WL91]. Theory-Based [GBD07]. Thermal [CGM^{+}07, MCG08, TGV08]. Thermal-Aware [TGV08]. Thing [SF09]. Thread [KL01, OC05, SLT03]. Threads [CASM07, DR98, HS99b, LJLS09]. Threat [YWF^{+}09]. Threats [ISAZM09]. Three [AD09, LCRW98, LHS03, MBTPV06, OB00, SZ03a]. Three-Dimensional [AD09, LCRW98, LHS03]. Three-Tier [MBTPV06]. Threshold [CGL07, vdMDM07]. Threshold-Based [CGL07]. Threshold-Multisignature [vdMDM07]. Throttling [TCLY07]. Through-Wafer [LCRW98]. Throughput [GLS07, HP07, WCCR^{+}97, ZXZ^{+}09]. Thwarting [CPM07]. Tier [MBTPV06]. Tiered [DTE07]. Tight [HK06, VV99]. Tighter [CL00, RO99]. Tightly [ADG^{+}08]. Tiled [GAK03, HCF03]. Tiles [RR02]. Tiling [ABRY03, JLF03, PHP03]. Time [AS99, ASS95, AMS97, Ano98c, AA09, AT01, BÖ98, BSCB09, BMR99, BM00a, BBG^{+}95, BGO^{+}98, BGOS97, BGO^{+}97, BGOS98, CF00, CRN09, CS97b, CKC08, CS03, CNT05, DÖ02, EDO06, FWDC^{+}00, FB01a, FLP^{+}07, GMM97, HS99a, HRG00, HNO98a, HNO98c, HJS^{+}06, HSH^{+}99, HS98a, HS02, HCF03, HS99b, KABK03, KHM05, KGM97, KA09, KMW08, KWH02, KKC03, KS01, KS03, KgCS04, KA99, LCB00, LLTW08, LB00a, LP07, LL07, LCN^{+}07, LHSML95, LA04, LWK05, LL98, MZ05, MM98a, MM98b, MT97, MRT06, MTL95, OS02, OZ96, PABD^{+}99, QCC99, Qua01, RA04, Ram99, RMO^{+}95, RP99, RRFH98, SJPL08, SCK00, ST99a, SE98, Sto96, SR99, TSAL97, TR04, Var01, WH03a, WR04, WJLK07, WCH^{+}08, WMWL08, XU01, XP05, XQ08, XC01, XTL06, YLL^{+}07, YW98, ZS95a, ZS98, ZMC03, ZMM04, ZLZN09, ZJ99]. time [AH91, ADM92, Aln94a, Aln95, Cap92, CD94, GG94b, GS91, HN93, JR94, wJNPS97, KSF94, KGM96, QM94, RSS90, RS91a, RWF94, Sar93, SC92, SC94, SF92a, SRS93, SH93, SH94, SA94, SL93a, SMS93, Var93, WC90, WCS92, DF97, GT93]. Time- [BGO^{+}98, OZ96]. Time-Aware [CNT05]. Time-Constrained [KHM05]. time-cost [Sar93]. Time-Critical [XTL06]. Time-Free [MRT06]. Time-Optimal [BBG^{+}95, BGOS97, ST99a, BGO^{+}97]. Time-Shared [FB01a]. time-stamp [Var93]. Time-Utility [WR04]. Timed [CF99b, Ost90]. Timeliness [HV07]. Timeliness-Accuracy [HV07]. Timeout [EBS04]. Timeout-Based [EBS04]. Timer [MRT06]. Timer-Based [MRT06]. Times [VM04, RS94, TRS90]. Timestamped [RKHM06]. timestamps [MB92]. Timing [KS08a, KCK^{+}06]. TMR [EMS90, EBS04]. Toeplitz [Pan93]. Toeplitz-like [Pan93]. Token [IKOY02, KY97, KKM08, HMR94]. token-and [HMR94]. Token-Based [KKM08]. TokenCMP [FPGAD08]. Tolerance [BHL^{+}07, CD08, FPGAD08, GMM97, HÖD99, KIBW99, KH97a, PBA03, SyFL99, SLH97, WC09, WMWL08, BP94, MN92, OC93, RJ94, SB94a, TC94]. Tolerant [AB99, AM95, Ano98b, BMR99, BC99, CYW08, lCL95, CC01, CXP09, CC98, CCD^{+}09, DDY99, DY05, Dua97, FIMR01, GY95a, GN96, GMCB01, HY99, HDF07, JZXX99, KH04, KLC97, Lan95, LDCO08, LH06a, LHSML95, MM98b, MJRS06, MBM98, PG07, RO99, RRRM09, SCP99, SDDY00, SNI02a, SNI02b, THH96, TL06, TCS97, TH01, VDS99, Wu98, WA99, Wu00, Xia01, YJ97a, YJ97b, YDW^{+}09, ZS98, dB98, AM91, BS95, BCH94, CL93, CS90, Chu96, FD94, KK93a, LG90, OS94a, OS94b, RST95, SM94, TB94, Tze93, VJ93, VJ94, WF94, YZW94]. Tolerate [Par95]. Tolerating [HY04, RCS01]. Tool [SRD08, Gab90]. Toolkit [Din06, SMBT90]. tools [HKM^{+}94]. Top [WZP^{+}03, KDL91]. top-down [KDL91]. Top-Level [WZP^{+}03]. Topological [CSH00, DAA02, DS05, Sto97, DT94, YA93]. Topologies [BS96, BBH05, BSS09, GY09, KWOA05, MDSS09, VB96]. Topology [Ano04d, BCQD07, CYW08, CTF09, CJHG08, DWX09, EVW07, GVGD95, HLH09, JJ07, JTC08, KZN07, LCRW98, LWS04, LH06a, LH06b, Liu08, MGZN07, NT09, OSRS06a, OSRS06b, RHM09, SD00a, SD00b, SLFW06, SGL06, SCL00, TL06, WD06, Cor92, Hsu93, MB94]. Topology-Aware [KZN07]. Topology-Flexible [TL06]. Tori [CH01, JSR98, LZ02, ST99a, SY98, TW98, YW02, UEA95]. Toroidal [AB99]. Torus [AB03, CYY00, GVGD95, PC96, PS96b, RMC95, SBS98, SS01, jTM96, TG96, TLGP97, YFJ^{+}01, GPBS94]. Total [CH98, DD98, DD01, FIMR01, HS98a, Jia95, SH97]. TPDS [Ano08d, Ano09d]. Trace [LLY05, LZTY09, PPR95, HE92, HB92, NGL94]. Trace-Driven [LZTY09, PPR95, HE92, NGL94].

Traceback [ADG06, GS08, SX03, XZG09]. Traceback-Based [SX03]. traces [HMW93, HE92]. Tracing [JBW^{+}08]. Trackability [TKW98]. Tracking [LH93, NSZ02, PPBSA97, XTL08, ZLZN09, AIK91]. Trade [CKK^{+}04, DZH05, FHA06, FLP^{+}07, GZ09, MYA01, QCC99, ZCXF09, DF97]. Trade-Off [FLP^{+}07, QCC99]. Trade-Offs [DZH05, GZ09, MYA01, ZCXF09, DF97]. tradeoffs [Aga92, DAF95]. Traffic [Aro00, BÖ98, CCQ^{+}05, HY07, Kop96, KPBD09, KgCS04, LKKS05, MSM06, OKSA01, RJ05, SY07, SZ95a, TSAL97, TK96b, XP05, AH91, CV92, Kop94]. Trail [QNR99]. Training [BBS^{+}09]. Transaction [QR07, ZMMS08, Tho93, YD94b]. Transactions [FG01, Ano02a]. Transcoding [CC03]. Transfer [KAY^{+}06, MS99b]. Transfers [EDO06, FV09, RRX09]. Transform [AD95, CPhX04, LHS03, TSP^{+}08, WH03a]. Transformation [BW96, FLVG95, HS98a, LL07, SS09, EHJ94, SC91, WL91]. Transformations [RJ96, D’H92, GMG96, SKF94, WW92]. transforms [Aln94b, ABDZ94, FA94, ZA92]. Transient [Her00, MGDZ07, KK93b]. Transient-Fault [MGDZ07]. Transition [LZ08, Ost90]. Transitive [TC95b, SC92, WC90]. Translation [QD05]. Transmission [BG09, ISRS06, LLY07, WCH^{+}08, RS94]. Transmissions [GG09, XL04, KGMB94]. Transparent [JLDC05, LSCZ07]. Transport [KS01, WS03, ZL07a]. Transport-Aware [WS03]. transpose [SH95b]. Transposition [RBSP02]. Transposition-Table-Driven [RBSP02]. Transputer [Add97]. Transversal [HY05]. Trapezoid [TN93b]. Traversals [Sto96]. Tree [ADD^{+}02, BCL^{+}05, BRSR08, CY95, CMDP09, DPN09, EVW07, GRS99, HY01, HH08, JZXX99, LC99, MKY^{+}09, Sto96, TC04a, VM99, WCL97, Wan98, WKS01, YK98, YC95, BGM94, Bil94, HMR94, KK94, LK94, SS90]. Tree-Based [HH08, LC99, MKY^{+}09, VM99, YK98, HMR94]. Tree-Sweep [GRS99]. Trees [AfAGR00, Ano99h, Avr99, Bar98, CCP95, CTS96, CH98, CBDW96, GRT97, Jia95, KDW01, KPK09, KWH03, LS96, LC96b, PWW00, RRRM09, SH97, Wan04, YR96, YCTW07, CL93, EF96, GM94]. Triangulation [LCWW03, LSW04]. Triangulations [BGOS98]. Tridiagonal [SZ04]. Triggered [CLJ^{+}04]. trigonometric [ABDZ94]. Trip [TPL96]. Trip-Based [TPL96]. TROP [THH08]. True [RLD03]. Trust [CHC09, LHL^{+}08, SJd^{+}09]. Trusted [ZH07b]. TTL [TCC07, TXL08]. TTL-Based [TCC07, TXL08]. Tunable [BBC^{+}95, YKP08]. Tuned [TLM04]. Tuning [ZBM09]. Tuple [BCdSFL09]. Turn [Chi00, JKA07]. Twin [AS00]. Twins [CDV^{+}06]. Twisted [FJL07, ZL96]. Two [AGGD05, BOC09, yCM98, CC99, DCF95, GG95, HC99a, Liu08, Mit01, Par95, SS96, SZ04, TC95a, WO04, ZWX06, BDS94, CV92, HK93, LC91b, ME95]. Two-Dimensional [yCM98, CC99, ZWX06, LC91b]. Two-Hop [Liu08]. Two-Level [AGGD05, DCF95, HC99a, SZ04]. Two-Stage [BOC09, HK93]. Two-Step [TC95a]. two-zero [ME95]. Two-Zone [WO04]. Type [CN02, CN04, Rob04]. Types [GT02].

Ubiquitous [RD09, YK03]. uCast [CHA07]. UCSC [DDD^{+}05]. unbounded [YG94]. Uncoordinated [QR07, WCLF95]. Understanding [CGM^{+}07]. Undirected [PWW00]. Unexpectedly [XCZ04]. Unfair [KY97]. Unfolding [CS97a]. Unicast [GP99b, LO95a, MXEN94, Mha09, SLFW06]. Unicast-based [MXEN94]. Unidirectional [HLH04, Wu02]. unification [RM90]. Unified [CHA07, FS00, GM97, GSS96, KCRK00, KCRB03, PK01, Yi09, AH93, DK92]. Uniform [DIM97, HLH04, KY97, LH01, NO02, O’H91, PB96, RMO^{+}95, Bil94, DR94, SF92a]. uniformization [HN93, TN93a]. Unifying [AC93]. unimodular [D’H92]. Unit [BSCB09, MC95]. Units [RSP02, TSP^{+}08]. UNITY [CR90]. UNITY-style [CR90]. Universal [AM99, GO97]. Unknown [GKK05, XCZ02]. Unordered [PWW00]. Unreliable [BV05, LWC^{+}09, SCW07]. unstable [GW94, GW96b]. Unstructured [BA07, CLY08b, GY09, HLH09, KK94, LLWC09, LXL^{+}05, OB00, SGL06, TXL08]. unused [KK93b]. Up* [SRD04]. Up*/Down* [SRD04]. Up-Down [KP01]. Update [FCF00, TC04b, LG94]. Upgradable [PABD^{+}99]. Uplink [KL02, MSM06]. Upper [CW02b, JR94]. Use [CT02, LSF^{+}09, SD00b, SSZ06, SS90]. Useful [Mit00]. User [CB05, CLY08b, HJB^{+}09, MF01b, PSC^{+}95, SLT03, TEF07]. User-Level [CB05, SLT03]. User-Selectable [HJB^{+}09]. User-Specified [PSC^{+}95]. Users [NSZ02]. Using [ACT06, AKNR^{+}04, AD09, BWC^{+}03, BR91, BCdSFL09, BDD^{+}96, CHC04, CWCC07, COS00, CH98, CCJ02, CHJ^{+}07, DW06, DP01, FLVG95, FMG02, GSS06, HKL00, HM98, Jia95, JK99, KGKL08, KBC^{+}01, KSP02, KSME08, KCW09, Kin06, KLS00, KAY^{+}06, KBD08, KET06, LCRW98, LRG99, Li03, LRS02, LJW^{+}07, LL98, MZT08, MZA02, MMSM06, ML94, MSG07, MQ97, OHRW99, OPZ99, OB00, OC05, PS96a, PDH06, QNR99, Ram99, RJ05, Sah00a, SWWJ08, SC07, SH97, SPS98, SSP02, SRL98, SY97, SP05, SL93c, TEF07, Tse09, TG99, TK96a, WSNA95, WLL^{+}07, WWWA09, WHM09, WF94, Wu98, Wu00, WHC03, WCDY06, Xia01, XCZ08, XB98, YN00, YL96, YG08, YZC08, ZFMS03, ZMC03, ZMP07, ZT01, ZW02, dLCK^{+}05, BCBzC92, DA93, GS08, HN93, HC92, KMT91, LS94c, LC91b, MS94b, SC91, SSG91, SMJ92, TKT92, WCF91, WFP90, ZL96]. Utility [CNT05, WR04]. Utility-Based [CNT05]. Utilization [CCJ02, LDG04, LWK05, MF01b, WJLK07, LY93a]. Utilizing [OXL06, SF07].

V256 [MS94a]. Valid [RJ96]. validated [TV92]. Value [AS00, RCS01]. Values [KP96, LL98]. Variable [AGWFH97]. variables [KST94]. Variation [BRS07]. Variational [Gre98]. Various [FJL07]. VCR [HL09a, WL08a]. VCR-Oriented [HL09a]. Vector [ÇA99, MS99b, NCV05, TN08, WNKS96, WH01, YY95, Har91, PKK93]. vectorization [KKP91]. Vectors [Wu98]. Vehicle [ZLZN09]. Velocity [SFP03]. Velocity-Based [SFP03]. Verification [HCHM09, JK99, PD95, PD00]. Verifying [CLS05, Qad03, SPC^{+}02]. versatile [GP93, Zia94]. Versioning [VGSS01]. versus [BCF^{+}08, TB93, TSP^{+}08]. VI [ZBJ^{+}05]. VI-Attached [ZBJ^{+}05]. Via [JS98, CS97a, CMR07, JBW^{+}08, KH93, NW98, TSG09, YXW03, ZZMN07]. Victor [MS94a]. Video [GB00, GLQL09, HL09a, KS01, LZTY09, XL04, YKS03]. Video-on-Demand [HL09a, LZTY09]. Vienna [UZCZ97]. Vienna-Fortran [UZCZ97]. Vienna-Fortran/HPF [UZCZ97]. Virtual [Cha96, CH04a, Dal92, GN96, KY98, KW08, Lee93, LW09c, LC02b, MOFD05, MROD07, MP97, SD00b, SZ95b, SM02, TPL96, VSD01, WYWZ08, DA93]. Virtual-channel [Dal92]. Virtual-Force-Based [LW09c]. Visibility [BBG^{+}95]. Visibility-Related [BBG^{+}95]. Vision [BA97, RJ99, CPA93]. Visual [Abr97, ADM92]. VLIW [AB94, CF01, MC95, OC05]. VLSI [Aln94b, AR97, BGO^{+}98, HALT95, TC93, ZA92]. VLSI-Optimal [BGO^{+}98]. VMMP [Gab90]. VMNet [WLZN07]. VOD [GMCB01]. Voice [WMXZ06, XL04, GWYS08]. Voice-over-IP [GWYS08]. VoIP [SWWJ08]. Vol [Ano02a]. Voltage [KSME08, Li08, ZMC03]. Voltage/Speed [ZMC03]. Volume [BA07]. Voronoi [AD08, EW97]. Vortex [HSWB07]. Voting [SB94b, XB98]. VOVO [HL09a].

Wafer [LCRW98]. Wait [Kuc01, FHRT93]. wait-depth [FHRT93]. Wait-Free [Kuc01]. Waiting [RMO^{+}95]. walk [You93]. Warp [AT01, CF00, QCC99, Qua01, SE98, DF97, GT93]. Warp-Based [QCC99]. Warshall [MF96]. Warshall-Floyd [MF96]. Watershed [GMRC07]. Watt [KHY09]. Wavefront [MA01, SKK01]. Wavelength [ZY04, ZY06]. Wavelet [TSP^{+}08]. WDM [GP03, SCP99, YW05a, ZY04, ZY06]. Weak [Kar01, GW94]. Web [ASB02, CCY03, CWLR09, CYD98, GB06, JLDC05, KK03a, KCD07, LLY04, LA04, LLA^{+}06, NE01, RK08, RAHM05, Ros03, RNKZ03, TC04b, TCC05, TSRS07, Tse05, ZRS^{+}05]. Web-Based [NE01]. Web-Computing [Ros03]. Web-Server [CYD98]. Weighted [DY05]. Weld [OC05]. Well [BDL95]. Well-Behaved [BDL95]. which [YK99]. Wide [DS02, DF99, Ant94]. Wide-Issue [DF99]. Wider [HTPS02]. Window [JN08]. Wire [EBS02]. Wireless [AYA09, ALW^{+}03, AD08, AD09, Ano01b, Ano01c, Ano01d, BK09, BBS^{+}09, BSCB09, BPT03, BCG04, BHJ02, BS08, CYW08, CPX06, CH08, CTF09, CKWC08, CNT05, DW04a, DW06, DPH08, DAMK06, DLS09, DWX09, EKOAW02, GZ06, GBC^{+}07, HSLA05, HCHM09, JGA08, JJ07, KPK09, KWL^{+}09, KyK09, KS08b, LJZA04, LDCO08, LCWW03, LWS04, LH06a, LSF^{+}09, LWC^{+}09, LWXS06, LWP07, MCL^{+}07, MEKOT03, MZA02, MMSM06, MTM02, NK08, ON02, SKS02, SJd^{+}09, SLFW06, SL01a, SL01b, SSZ02, Sto04, TCO01, TX08, TN08, WY07, WWL06, WT08, WLZ08, WWLS08, WWWA09, Wu02, WLZN07, WCD08, XLW^{+}06, XCZ08, YLW07, Yi09, YYY09, YG08, ZCXF09]. within [LCB00, NSD^{+}91]. without [Fu05, GN96, SWC95, VJA97]. WK [Fu05, SCD97]. WK-Recursive [Fu05, SCD97]. Word [CF01]. Work [ÇF99a, HNO98c, RBSP02, XU01]. Work-Efficient [ÇF99a]. Work-Time [HNO98c, XU01]. Workflow [LSZ09]. Workflows [PF08]. Workload [Ros02, ZSMF01, ZRS^{+}05]. Workload-Aware [ZRS^{+}05]. Workloads [CV08, MF01b, NKP^{+}96, PB96, YZC08]. Workstation [GKK05, LLH^{+}01]. Workstations [AA09, CdMB05, EK95, FB01a, JL99, Ros02, RH00, RH04, SD00a, SD00b, SOM05, DGB^{+}96, SSG91]. World [HLL09, LLSZ08]. Worm [JBW^{+}08, RS97b]. Wormhole [BP98, BLD05, BC96, BCR98, Chi98, Dua95a, Dua95b, Dua97, FF98, GN96, GO97, HÖD99, HÖ00, HK95, KP99, KLS00, LMS04, LMN95, MRLD01, NCV05, NGM97, OKSA01, PSK99, RMC95, RLD03, SCL01, jTM96, TG96, TPL96, TLGP97, TWH99, VM99, VS11, XGN97, ZL05, Dua93, LMN94, MXEN94, jTM97]. Wormhole-Routed [BP98, FF98, HÖ00, HK95, KLS00, LMN95, RMC95, SCL01, jTM96, TG96, TPL96, TLGP97, TWH99, XGN97, MXEN94, jTM97]. Wormhole-Switched [HÖD99, VM99]. Worms [SSP00, TC07]. Worst [GRT97, TSJ07]. Worst-Case [TSJ07]. Wraparound [SV97]. Wrapped [HWSH00, WMN99]. Write [BB08, HNY02, KDW01]. Write-Enabled [BB08]. Writing [WBO^{+}01]. WSN [KSP09].

X [GM94]. X-trees [GM94]. XML [CF08]. XNet [CF08].

Yama [MJ06].

Zero [LHL^{+}08, ME95]. Zero-Knowledge [LHL^{+}08]. Zone [WO04].

References

Anderson:2000:PBC

[AA00] J. R. Anderson and S. Abraham. Performance-based constraints for multidimensional networks. IEEE Transactions on Parallel and Distributed Systems, 11(1):21–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                                   http://dlib.computer.org/
                                                   td/books/td2000/pdf/l0021.
                                                   pdf;   http://www.computer.
                                                   org/tpds/td2000/l0021abs.
                                                   htm.

Auluck:2009:ESR

[AA09] Nitin Auluck and Dharma P. Agrawal. Enhancing the schedulability of real-time heterogeneous networks of workstations (NOWs). IEEE Transactions on Parallel and Distributed Systems, 20(11):1586–1599, November 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Amir:2000:OCA

[AAB^{+}00] Y. Amir, B. Awerbuch, A. Barak, R. S. Borgstrom, and A. Keren. An opportunity cost approach for job assignment in a scalable computing cluster. IEEE Transactions on Parallel and Distributed Systems, 11(7):760– ??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0760.
pdf;   http://www.computer.
org/tpds/td2000/l0760abs.
htm.

Aluru:2006:ESS

[AAB06] Srinivas Aluru, Nancy M. Amato, and David A. Bader. Editorial: Special section on highperformance computational biology. IEEE Transactions on Parallel and Distributed Systems, 17 (8):737–739, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/08/
l0737.pdf.

Al-Ayyoub:1997:MDS

[AAD97] A.-E. Al-Ayyoub and K. Day.

Matrix decomposition on the star graph. IEEE Transactions on Parallel and Distributed Systems, 8(8):803–812, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0803.
pdf;   http://www.computer.
org/tpds/td1997/l0803abs.
htm.

Al-Azzoni:2008:LPB

[AAD08] Issam Al-Azzoni and Douglas G.

Down. Linear programmingbased affinity scheduling of independent tasks on heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 19(12): 1671–1682, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abu-Amara:1994:NMA

[AAG94] Hosame Abu-Amara and Vidya Gummadi. New model and algorithms for leader election in synchronous fiber-optic networks. IEEE Transactions on Parallel and Distributed Systems, 5(8): 891–896, August 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Adir:2003:IFM

[AAS03] Allon Adir, Hagit Attiya, and Gil Shurek. Information-flow models for shared memory with an application to the PowerPC architecture. IEEE Transactions on Parallel and Distributed Systems, 14(5):502–515, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/05/l0502abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/05/l0502.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/05/l0502.
                                        pdf.

Agrawal:1991:NQC

[AB91a] Divyakant Agrawal and Arthur J. Bernstein. A nonblocking quorum consensus protocol for replicated data. IEEE Transactions on Parallel and Distributed Systems, 2(2):171–179, April 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aravena:1991:CLC

[AB91b] Jorge L. Aravena and Abdulkader O. Barbir. A class of low complexity high concurrence algorithms. IEEE Transactions on Parallel and Distributed Systems, 2(4):495–502, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abnous:1994:PBV

[AB94] Arthur Abnous and Nader Bagherzadeh. Pipelining and bypassing in a VLIW processor. IEEE Transactions on Parallel and Distributed Systems, 5 (6):658–664, June 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

AlMohammad:1999:FTC

[AB99] B. F. A. AlMohammad and B. Bose. Fault-tolerant communication algorithms in toroidal networks. IEEE Transactions on Parallel and Distributed Systems, 10(10):976–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0976.
pdf;   http://www.computer.
org/tpds/td1999/l0976abs.
htm.

Ayoubi:2003:EMA

[AB03] Rafic A. Ayoubi and Magdy A. Bayoumi. Efficient mapping algorithm of multilayer neural network on torus architecture. IEEE Transactions on Parallel and Distributed Systems, 14(9):932–943, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0932abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0932.
pdf.

Ahmed:2007:MSF

[AB07] Masood Ahmed and Shahid Bokhari. Mapping with space filling surfaces. IEEE Transactions on Parallel and Distributed Systems, 18(9):1258– 1269, September 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aiello:2001:ARN

[ABC^{+}01a] W. Aiello, S. N. Bhatt, F. R. K. Chung, A. L. Rosenberg, and R. K. Sitaraman. Augmented ring networks. IEEE Transactions on Parallel and Distributed Systems, 12(6):598–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0598.
pdf;   http://www.computer.
org/tpds/td2001/l0598abs.
htm.

Avresky:2001:ISS

[ABC01b] D. R. Avresky, J. Bruck, and D. E. Culler. Introduction to the special section on dependable network computing. IEEE Transactions on Parallel and Distributed Systems, 12(2):97–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0097.
pdf.

Arguello:1994:PAF

[ABDZ94] F. Arguello, J. D. Bruguera, R. Doallo, and E. L. Zapata. Parallel architecture for fast transforms with trigonometric kernel. IEEE Transactions on Parallel and Distributed Systems, 5(10):1091–1099, October 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alaghband:1993:LPA

[ABJ^{+}93] Gita Alaghband, Muhammad S.

Benten, Ruediger Jakob, Harry F. Jordan, and Aruna V. Ramanan. Language portability across shared memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 4(9):1064–1072, September 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Amoura:1998:SAP

[ABK98] A. K. Amoura, E. Bampis, and J.-C. König. Scheduling algorithms for parallel Gaussian elimination with communication costs. IEEE Transactions on Parallel and Distributed Systems, 9(7):679–??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0679.
pdf;   http://www.computer.
org/tpds/td1998/l0679abs.
htm.

Abrams:1997:EDP

[Abr97] Marc Abrams. An example of deriving performance properties from a visual representation of program execution. IEEE Transactions on Parallel and Distributed Systems, 8(6):658– 666, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0658.
pdf;   http://www.computer.
org/tpds/td1997/l0658abs.
htm.

Andonov:2003:OSO

[ABRY03] Rumen Andonov, Stefan Balev, Sanjay Rajopadhye, and Nicola Yanev. Optimal semi-oblique tiling. IEEE Transactions on Parallel and Distributed Systems, 14(9):944–960, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0944abs.
htm;   http://csdl.computer.

org/dl/trans/td/2003/09/l0944.
pdf.

Anastasi:2001:RMP

[ABS01] Giuseppe Anastasi, Alberto Bartoli, and Francesco Spadoni. A reliable multicast protocol for distributed mobile systems: Design and evaluation. IEEE Transactions on Parallel and Distributed Systems, 12(10):1009–1022, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1009.
pdf;   http://www.computer.
org/tpds/td2001/l1009abs.
htm.

Agrawal:1992:PAS

[AC92] V. D. Agrawal and S. T.

Chakradhar. Performance analysis of synchronized iterative algorithms on multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 3(6):739–746, November 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Angelaccio:1993:UOP

[AC93] Michele Angelaccio and Michele Colajanni. Unifying and optimizing parallel linear algebra algorithms. IEEE Transactions on Parallel and Distributed Systems, 4(12):1382–1397, December 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ayguade:2009:DOT

[ACD^{+}09] Eduard Ayguade, Nawal Copty, Alejandro Duran, Jay Hoeflinger, Yuan Lin, Federico Massaioli, Xavier Teruel, Priya Unnikrishnan, and Guansong Zhang. The design of OpenMP tasks. IEEE Transactions on Parallel and Distributed Systems, 20(3):404–418, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ahmad:2008:GEI

[ACM08] Ishfaq Ahmad, Kirk W. Cameron, and Rami Melhem. Guest Editor’s introduction: Special section on power-aware parallel and distributed computing (PA-PADS). IEEE Transactions on Parallel and Distributed Systems, 19(11):1441–1444, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Allen:1997:PAG

[ACT^{+}97] Robert Allen, Luigi Cinque, Steven Tanimoto, Linda Shapiro, and Dean Yasuda. A parallel algorithm for graph matching and its MasPar implementation. IEEE Transactions on Parallel and Distributed Systems, 8(5):490–501, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0490.
                                        pdf;   http://www.computer.

org/tpds/td1997/l0490abs.
htm.

Agarwal:2006:BES

[ACT06] Sachin Agarwal, Vikas Chauhan, and Ari Trachtenberg. Bandwidth efficient string reconciliation using puzzles. IEEE Transactions on Parallel and Distributed Systems, 17(11):1217– 1225, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aykanat:1995:EFH

[AD95] Cevdet Aykanat and Argun Dervis. Efficient fast Hartley transform algorithms for hypercube-connected multicomputers. IEEE Transactions on Parallel and Distributed Systems, 6(6):561–577, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0561abs.htm.

Abandah:1998:CDS

[AD98] G. A. Abandah and E. S.

Davidson. Characterizing distributed shared memory performance: a case study of the Convex SPP1000. IEEE Transactions on Parallel and Distributed Systems, 9(2):206–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0206.
pdf;   http://www.computer.
org/tpds/td1998/l0206abs.
htm.

Ammari:2008:PHM

[AD08] Habib M. Ammari and Sajal K. Das. Promoting heterogeneity, mobility, and energyaware Voronoi diagram in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(7):995– 1008, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Ammari:2009:CDC

[AD09] Habib M. Ammari and Sajal K. Das. Critical density for coverage and connectivity in threedimensional wireless sensor networks using continuum percolation. IEEE Transactions on Parallel and Distributed Systems, 20 (6):872–885, June 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Adda:1997:SMC

[Add97] Mahamed Adda. A scalable multibus configuration for connecting transputer links. IEEE Transactions on Parallel and Distributed Systems, 8(3):245– 253, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0245.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0245abs.
                                        htm.

Auletta:2002:OTA

[ADD^{+}02] Vincenzo Auletta, Sajal K. Das, Amelia De Vivo, M. Cristina Pinotti, and Vittorio Scarano. Optimal tree access by elementary and composite templates in parallel memory systems. IEEE Transactions on Parallel and Distributed Systems, 13(4):399– 412, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0399.
pdf;   http://www.computer.
org/tpds/td2002/l0399abs.
htm.

Al-Duwairi:2006:NHS

[ADG06] Basheer Al-Duwairi and Manimaran Govindarasu. Novel hybrid schemes employing packet marking and logging for IP traceback. IEEE Transactions on Parallel and Distributed Systems, 17(5):403–418, May 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aridor:2008:MIT

[ADG^{+}08] Yariv Aridor, Tamar Domany, Oleg Goldshmidt, Yevgeny Kliteynik, Edi Shmueli, and Jose E. Moreira. Multitoroidal interconnects for tightly coupled supercomputers. IEEE Transactions on Parallel and Distributed Systems, 19(1):52– 65, January 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.

                                        org/comp/trans/td/2008/01/
                                        ttd2008010052s.zip.

Abrams:1992:CVA

[ADM92] M. Abrams, N. Doraswamy, and A. Mathur. Chitra: visual analysis of parallel and distributed programs in the time, event, and frequency domains. IEEE Transactions on Parallel and Distributed Systems, 3(6): 672–685, November 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agrawal:1997:AQB

[AEA97] Divyakant Agrawal, Ömer Egecioğlu, and Amr El Abbadi. Analysis of quorum-based protocols for distributed (k + 1)-exclusion. IEEE Transactions on Parallel and Distributed Systems, 8(5):533–537, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0533.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0533abs.
                                        htm.

Aggarwal:2005:SAI

[AF05] Aneesh Aggarwal and Manoj Franklin. Scalability aspects of instruction distribution algorithms for clustered processors. IEEE Transactions on Parallel and Distributed Systems, 16(10): 944–955, October 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Al-Furiah:1997:PAS

[AFAGR97] I. Al-Furiah, S. Aluru, S. Goil, and S. Ranka. Practical algorithms for selection on coarse-grained parallel computers. IEEE Transactions on Parallel and Distributed Systems, 8(8):813–824, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0813.
pdf;   http://www.computer.
org/tpds/td1997/l0813abs.
htm.

Al-furaih:2000:PCM

[AfAGR00] I. Al-furaih, S. Aluru, S. Goil, and S. Ranka. Parallel construction of multidimensional binary search trees. IEEE Transactions on Parallel and Distributed Systems, 11(2):136–??, February 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0136.
pdf;   http://www.computer.
org/tpds/td2000/l0136abs.
htm.

Abdul-Fatah:2002:PCB

[AFM02] Istabrak Abdul-Fatah and Shikharesh Majumdar. Performance of CORBA-based client-server architectures. IEEE Transactions on Parallel and Distributed Systems, 13(2):111–127, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0111.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0111abs.
                                        htm.

Amaral:1996:CAS

[AG96] Jose Nelson Amaral and Joydeep Ghosh. Concurrent architecture for serializable production systems. IEEE Transactions on Parallel and Distributed Systems, 7(12):1265–1280, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agarwal:1991:LIN

[Aga91] Anant Agarwal. Limits on interconnection network performance. IEEE Transactions on Parallel and Distributed Systems, 2(4):398–412, October 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agarwal:1992:PTM

[Aga92] Anant Agarwal. Performance tradeoffs in multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 3(5):525–539, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abdel-Ghaffar:1994:OSC

[AGE94] Khaled A. S. Abdel-Ghaffar and Amr El Abbadi. An optimal strategy for comparing file copies. IEEE Transactions on Parallel and Distributed Systems, 5(1):87–93, January 1994.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Acacio:2004:AHP

[AGGD04] Manuel E. Acacio, José González, José M. García, and José Duato. An architecture for highperformance scalable sharedmemory multiprocessors exploiting on-chip integration. IEEE Transactions on Parallel and Distributed Systems, 15 (8):755–768, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/08/l0755.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/08/l0755.pdf.

Acacio:2005:TLD

[AGGD05] Manuel E. Acacio, José González, José M. García, and José Duato. A two-level directory architecture for highly scalable cc-NUMA multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 16(1):67–79, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0067abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0067.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0067.
pdf.

Alverson:1998:APS

[AGL^{+}98] G. A. Alverson, W. G. Griswold, C. Lin, D. Notkin, and L. Snyder. Abstractions for portable, scalable parallel programming. IEEE Transactions on Parallel and Distributed Systems, 9(1):71–86, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0071.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0071abs.
                                        htm.

Agrawal:1998:IPR

[Agr98] G. Agrawal. Interprocedural partial redundancy elimination with application to distributed memory compilation. IEEE Transactions on Parallel and Distributed Systems, 9(7):609– ??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0609.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0609abs.
                                        htm.

Agrawal:1999:GIF

[Agr99] G. Agrawal. A general interprocedural framework for placement of split-phase large latency operations. IEEE Transactions on Parallel and Distributed Systems, 10(4):394–??, April 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0394.
pdf;   http://www.computer.
org/tpds/td1999/l0394abs.
htm.

Abu-Ghazaleh:1997:SVI

[AGWFH97] Nael B. Abu-Ghazaleh, Philip A.

Wilsey, Xianzhi Fan, and Debra A. Hensgen. Synthesizing variable instruction issue interpreters for implementing functional parallelism on SIMD computers. IEEE Transactions on Parallel and Distributed Systems, 8(4):412–423, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0412.
pdf;   http://www.computer.
org/tpds/td1997/l0412abs.
htm.

Abraham:1991:CTP

[AH91] Santosh G. Abraham and David E. Hudak. Compile-time partitioning of iterative parallel loops to reduce cache coherency traffic. IEEE Transactions on Parallel and Distributed Systems, 2(3):318–328, July 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Adve:1993:UFF

[AH93] Sarita V. Adve and Mark D.

Hill. A unified formalization of four shared-memory models. IEEE Transactions on Parallel and Distributed Systems, 4 (6):613–624, June 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Attiya:2006:IQD

[AH06] Hagit Attiya and David Hay. The inherent queuing delay of parallel packet switches. IEEE Transactions on Parallel and Distributed Systems, 17(9):1048– 1056, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Ahuja:1993:IFC

[Ahu93] Mohan Ahuja. An implementation of F-channels. IEEE Transactions on Parallel and Distributed Systems, 4(6):658–667, June 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Averbuch:1991:PIM

[AIK91] Amir Averbuch, Samuel Itzikowitz, and Tal Kapon. Parallel implementation of multiple model tracking algorithms. IEEE Transactions on Parallel and Distributed Systems, 2(2):242– 252, April 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agrawal:1995:CBR

[AJ95] Gagan Agrawal and Pankaj Jalote. Coding-based replication schemes for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 6 (3):240–251, March 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0240abs.htm.

Ammann:1996:GCE

[AJF96] Paul Ammann, Sushil Jajodia, and Phyllis G. Frankl. Globally consistent event ordering in one-directional distributed environments. IEEE Transactions on Parallel and Distributed Systems, 7(6):665–670, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0665abs.htm.

Al-Jaroodi:2003:MIP

[AJMJS03] Jameela Al-Jaroodi, Nader Mohamed, Hong Jiang, and David Swanson. Middleware infrastructure for parallel and distributed programming models in heterogeneous systems. IEEE Transactions on Parallel and Distributed Systems, 14(11):1100–1111, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/11/l1100abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/11/l1100.
pdf.

Ahmad:1998:ETD

[AK98] I. Ahmad and Y.-K. Kwok.

On exploiting task duplication in parallel program scheduling. IEEE Transactions on Parallel and Distributed Systems, 9(9):872–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0872.
pdf;   http://www.computer.
org/tpds/td1998/l0872abs.
htm.

Ahamad:1999:SCP

[AK99a] M. Ahamad and R. Kordale. Scalable consistency protocols for distributed services. IEEE Transactions on Parallel and Distributed Systems, 10(9):888–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0888.
pdf;   http://www.computer.
org/tpds/td1999/l0888abs.
htm.

Ahmad:1999:PMS

[AK99b] I. Ahmad and Y.-K. Kwok. On parallelizing the multiprocessor scheduling problem. IEEE Transactions on Parallel and Distributed Systems, 10(4):414– ??, April 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0414.
pdf;   http://www.computer.
org/tpds/td1999/l0414abs.
htm.

Agarwal:1995:APP

[AKN95] Anant Agarwal, David A.

Kranz, and Venkat Natarajan. Automatic partitioning of parallel loops and data arrays for distributed shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6(9):943–962, September 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0943abs.htm.

Amir:2004:SGC

[AKNR^{+}04] Yair Amir, Yongdae Kim, Cristina Nita-Rotaru, John L. Schultz, Jonathan Stanton, and Gene Tsudik. Secure group communication using robust contributory key agreement. IEEE Transactions on Parallel and Distributed Systems, 15(5):468–480, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/05/l0468abs.
htm;   http://csdl.computer.
org/comp/trans/td/2004/05/
L0468s.pdf;   http://csdl.
computer.org/dl/trans/td/
2004/05/l0468.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/05/l0468.pdf.

Andrade:2004:OEM

[AKSS04] Henrique Andrade, Tahsin Kurc, Alan Sussman, and Joel Saltz. Optimizing the execution of multiple data analysis queries on parallel and distributed environments. IEEE Transactions on Parallel and Distributed Systems, 15(6):520–532, June 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/06/l0520.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/06/l0520.pdf.

Alnuweiri:1994:CTP

[Aln94a] Hussein M. Alnuweiri. Constanttime parallel algorithms for image labeling on a reconfigurable network of processors. IEEE Transactions on Parallel and Distributed Systems, 5(3):320– 326, March 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alnuweiri:1994:OVN

[Aln94b] Hussein M. Alnuweiri. Optimal VLSI networks for multidimensional transforms. IEEE Transactions on Parallel and Distributed Systems, 5(7):763–769, July 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alnuweiri:1995:PCT

[Aln95] Hussein M. Alnuweiri. Parallel constant-time connectivity algorithms on a reconfigurable network of processors. IEEE Transactions on Parallel and Distributed Systems, 6(1):105–110, January 1995. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alzoubi:2003:GSW

[ALW^{+}03] Khaled Alzoubi, Xiang-Yang Li, Yu Wang, Peng-Jun Wan, and Ophir Frieder. Geometric spanners for wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 14(4):408–421, April 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0408abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0408.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0408.
pdf.

Akl:1990:PBS

[AM90] Selim G. Akl and Henk Meijer.

Parallel binary search. IEEE Transactions on Parallel and Distributed Systems, 1(2):247– 250, April 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alam:1991:EMS

[AM91] M. Sultan Alam and Rami G.

Melhem. An efficient modular spare allocation scheme and its application to fault tolerant binary hypercubes. IEEE Transactions on Parallel and Distributed Systems, 2(1):117–126, January 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Al-Mouhamed:1993:AMD

[AM93] Mayez Al-Mouhamed. Analysis of macro-dataflow dynamic scheduling on nonuniform memory access architectures. IEEE Transactions on Parallel and Distributed Systems, 4(8):875– 888, August 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alam:1995:RMF

[AM95] M. Sultan Alam and Rami G. Melhem. Routing in modular fault-tolerant multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 6(11):1206– 1220, November 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1206abs.htm.

Anderson:1999:UCL

[AM99] J. H. Anderson and M. Moir. Universal constructions for large objects. IEEE Transactions on Parallel and Distributed Systems, 10(12):317–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1317.
pdf;   http://www.computer.
org/tpds/td1999/l1317abs.
htm.

Amoroso:2006:MJS

[AM06] Alessandro Amoroso and Keith Marzullo. Multiple job scheduling in a connection-limited data parallel system. IEEE Transactions on Parallel and Distributed Systems, 17(2):125–134, February 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Al-Mouhamed:1994:PES

[AMAM94] Mayez Al-Mouhamed and Adel Al-Maasarani. Performance evaluation of scheduling precedenceconstrained computations on message-passing systems. IEEE Transactions on Parallel and Distributed Systems, 5(12):1317– 1321, December 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Artail:2008:DDS

[AMH08] Hassan Artail, Khaleel W.

Mershad, and Hicham Hamze. DSDM: a distributed service discovery model for manets. IEEE Transactions on Parallel and Distributed Systems, 19(9):1224– 1236, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Atreya:2007:QBG

[AMP07] Ranganath Atreya, Neeraj Mittal, and Sathya Peri. A quorumbased group mutual exclusion algorithm for a distributed system with dynamic group set. IEEE Transactions on Parallel and Distributed Systems, 18(10): 1345–1360, October 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alvisi:2001:FDB

[AMPR01] Lorenzo Alvisi, Dahlia Malkhi, Evelyn Pierce, and Michael K.

Reiter. Fault detection for Byzantine quorum systems. IEEE Transactions on Parallel and Distributed Systems, 12(9): 996–1007, September 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l0996.
pdf;   http://www.computer.
org/tpds/td2001/l0996abs.
htm.

Al-Mouhamed:1997:HSM

[AMS97] Mayez A. Al-Mouhamed and Steven S. Seiden. A heuristic storage for minimizing access time of arbitrary data patterns. IEEE Transactions on Parallel and Distributed Systems, 8(4):441–447, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0441.
pdf;   http://www.computer.
org/tpds/td1997/l0441abs.
htm.

Ali:2004:MRR

[AMSK04] Shoukat Ali, Anthony A. Maciejewski, Howard Jay Siegel, and Jong-Kook Kim. Measuring the robustness of a resource allocation. IEEE Transactions on Parallel and Distributed Systems, 15(7):630–641, July 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/07/l0630.htm;   http:

//csdl.computer.org/dl/trans/
td/2004/07/l0630.pdf.

Al-Mistarihi:2009:FOR

[AMY09] Husni Hamad E. Al-Mistarihi and Chan Huah Yong. On fairness, optimizing replica selection in data grids. IEEE Transactions on Parallel and Distributed Systems, 20(8):1102–1111, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alverson:1993:PSE

[AN93] Gail A. Alverson and David Notkin. Program structuring for effective parallel portability. IEEE Transactions on Parallel and Distributed Systems, 4 (9):1041–1059, September 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agrawal:1994:CNF

[AN94] Prathima Agrawal and Antony Ng. Computing network flow on a multiple processor pipeline. IEEE Transactions on Parallel and Distributed Systems, 5 (6):653–658, June 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anderson:1990:PSL

[And90] Thomas E. Anderson. The performance of spin lock alternatives for shared-money multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 1(1):6–16, January 1990.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alkhalifa:1999:DES

[ANKA99] Z. Alkhalifa, V. S. S. Nair, N. Krishnamurthy, and J. A. Abraham. Design and evaluation of system-level checks for on-line control flow error detection. IEEE Transactions on Parallel and Distributed Systems, 10(6):627–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0627.
pdf;   http://www.computer.
org/tpds/td1999/l0627abs.
htm.

Aiken:1995:RCS

[ANN95] Alexander Aiken, Alexandru Nicolau, and Steven Novack. Resource-constrained software pipelining. IEEE Transactions on Parallel and Distributed Systems, 6(12):1248–1270, December 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:1997:AI

[Ano97a]  Anonymous.  Annual in-

dex. IEEE Transactions on Parallel and Distributed Systems, 8(12):1301–??, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1301.
pdf.

Anonymous:1997:CPSb

[Ano97b] Anonymous. Call for papers for special issue on compilers and languages for parallel and distributed computers. IEEE Transactions on Parallel and Distributed Systems, 8(10):1087– ??, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1087.
pdf.

Anonymous:1997:CPSc

[Ano97c]  Anonymous.  Call for pa-

pers for special issue on compilers and languages for parallel and distributed computers. IEEE Transactions on Parallel and Distributed Systems, 8(11):1184–??, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1184.
pdf.

Anonymous:1997:CPSa

[Ano97d] Anonymous. Call for papers: Special issue on compilers and languages for parallel and distributed computers. IEEE Transactions on Parallel and Distributed Systems, 8 (9):??, September 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:1998:AI

[Ano98a]  Anonymous.  Annual in-

dex. IEEE Transactions on Parallel and Distributed Systems, 9(12):269–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1269.
pdf.

Anonymous:1998:CPSb

[Ano98b] Anonymous. Call for papers for special issue on challenges in designing fault-tolerant routing in networks. IEEE Transactions on Parallel and Distributed Systems, 9(7):704–??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0704.
pdf.

Anonymous:1998:CPSa

[Ano98c] Anonymous. Call for papers for special issue on dependable real-time systems. IEEE Transactions on Parallel and Distributed Systems, 9(3):320–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0320.
pdf.

Anonymous:1999:RL

[Ano99a]  Anonymous.  1998 review-

ers list. IEEE Transactions on Parallel and Distributed Systems, 10(1):4–??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0004.
pdf.

Anonymous:1999:AI

[Ano99b] Anonymous. 1999 annual index. IEEE Transactions on Parallel and Distributed Systems, 10(12):333–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1333.
pdf.

Anonymous:1999:CPa

[Ano99c]  Anonymous.  Call for pa-

pers. IEEE Transactions on Parallel and Distributed Systems, 10(8):864–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0864.
pdf.

Anonymous:1999:CPb

[Ano99d] Anonymous. Call for papers. IEEE Transactions on Parallel and Distributed Systems, 10(9):958–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0958.
pdf.

Anonymous:1999:CPc

[Ano99e]  Anonymous.  Call for pa-

pers. IEEE Transactions on Parallel and Distributed Systems, 10(10):103–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l1103.
                                        pdf.

Anonymous:1999:CPR

[Ano99f]  Anonymous.  Correction to

“Periodically Regular Chordal Rings”. IEEE Transactions on Parallel and Distributed Systems, 10(7):767–??, July 1999. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/td/books/td1999/
                                        pdf/l0767.pdf. See [PK99b].

Anonymous:1999:CEJ

[Ano99g]  Anonymous.  Correction to

the Editorial in the January 1999 issue. IEEE Transactions on Parallel and Distributed Systems, 10(3):336–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0336.

pdf. See [Sta99].

Anonymous:1999:ECE

[Ano99h] Anonymous. Erratum: Correction to “Embedding and Reconfiguration of Spanning Trees in Faulty Hypercubes”. IEEE Transactions on Parallel and Distributed Systems, 10(10):102– ??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1102.

pdf. See [Avr99].

Anonymous:2000:RL

[Ano00a]  Anonymous.  1999 review-

ers list. IEEE Transactions on Parallel and Distributed Systems, 11(1):3–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0003.
pdf.

Anonymous:2000:TCPa

[Ano00b] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 11(9):??, September 2000. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td2000/
l9c1.pdf.

Anonymous:2000:TCPb

[Ano00c] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 11(11):??, November 2000. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td2000/
lyc1.pdf.

Anonymous:2001:RL

[Ano01a]  Anonymous.  2000 review-

ers list. IEEE Transactions on Parallel and Distributed Systems, 12(1):3–??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0003.
                                        pdf.

Anonymous:2001:CPSa

[Ano01b] Anonymous. Call for papers for special issue on mobile computing and wireless networks. IEEE Transactions on Parallel and Distributed Systems, 12(2): 238–??, February 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td2001/l0238.pdf.

Anonymous:2001:CPSb

[Ano01c] Anonymous. Call for papers for special issue on mobile computing and wireless networks. IEEE Transactions on Parallel and Distributed Systems, 12 (3):336–??, March 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td2001/l0336.pdf.

Anonymous:2001:CPSc

[Ano01d] Anonymous. Call for papers for special issue on mobile computing and wireless networks. IEEE Transactions on Parallel and Distributed Systems, 12 (4):432–??, April 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td2001/l0432.pdf.

Anonymous:2001:I

[Ano01e] Anonymous. Index. IEEE Transactions on Parallel and Distributed Systems, 12(12): 1332–1341, December 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l1332.
pdf;   http://www.computer.
org/tpds/td2001/l1332abs.
htm.

Anonymous:2001:TCPa

[Ano01f] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 12(1):??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1c1.
pdf.

Anonymous:2001:TCPb

[Ano01g] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 12(2):??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l2c1.
pdf.

Anonymous:2001:TCPc

[Ano01h] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 12(3):??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l3c1.
                                        pdf.

Anonymous:2001:TCPd

[Ano01i] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 12(4):??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l4c1.
                                        pdf.

Anonymous:2001:TCPe

[Ano01j] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 12(5):??, May 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l5c1.
                                        pdf.

Anonymous:2001:TCPf

[Ano01k] Anonymous. Table of contents in PDF. IEEE Transactions on Parallel and Distributed Systems, 12(6):??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l6c1.
pdf.

Anonymous:2002:ITP

[Ano02a] Anonymous. 2002 index Transactions on Parallel and Distributed Systems, vol. 13. IEEE Transactions on Parallel and Distributed Systems, 13(12): 1333–1342, December 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2002/12/
l1333.pdf;   http://csdl.
computer.org/dl/trans/td/
2002/12/l1333.htm.

Anonymous:2002:CPS

[Ano02b] Anonymous. Call for papers for special issue on middleware. IEEE Transactions on Parallel and Distributed Systems, 13(6):??, June 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td2002/l0672.pdf.

Anonymous:2002:NE

[Ano02c]  Anonymous.  Notice of er-

rata. IEEE Transactions on Parallel and Distributed Systems, 13(10):1084, October 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l1084.
pdf;   http://www.computer.

                                        org/tpds/td2002/l1084abs.
                                        htm.

Anonymous:2003:RL

[Ano03a] Anonymous. 2002 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 14 (1):94–96, January 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/01/
                                        l0094.pdf;   http://csdl.
                                        computer.org/dl/trans/td/
                                        2003/01/l0094.htm.

Anonymous:2003:I

[Ano03b] Anonymous. 2003 index. IEEE Transactions on Parallel and Distributed Systems, 14(12): 1286–1296, December 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/12/
                                        l1286.pdf.

Anonymous:2003:CPS

[Ano03c]  Anonymous.  Call for pa-

pers for special issue on onchip networks. IEEE Transactions on Parallel and Distributed Systems, 14(11):1200– ??, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/11/
                                        l1200.pdf.

Anonymous:2004:AI

[Ano04a]  Anonymous.  Annual index.

IEEE Transactions on Parallel and Distributed Systems, 15(12): 1138–1148, December 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/12/
l1138.pdf;   http://csdl.
computer.org/dl/trans/td/
2004/12/l1138.htm.

Anonymous:2004:CP

[Ano04b] Anonymous. Call for papers.

IEEE Transactions on Parallel and Distributed Systems, 15(12): 1149–??, December 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/12/
l1149.pdf.

Anonymous:2004:CPSa

[Ano04c] Anonymous. Call for papers for special issue on algorithm design and scheduling techniques (realistic platform models) for heterogeneous clusters. IEEE Transactions on Parallel and Distributed Systems, 15 (10):960–??, October 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/10/
l0960.pdf.

Anonymous:2004:CPSb

[Ano04d] Anonymous. Call for papers for special issue on localized communication and topology protocols for ad hoc network. IEEE Transactions on Parallel and Distributed Systems, 15(11):

1056–??, November 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2004/11/
                                        l1056.pdf.

Anonymous:2004:RL

[Ano04e] Anonymous. Reviewers list. IEEE Transactions on Parallel and Distributed Systems, 15 (1):93–96, January 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2004/01/
                                        l0093.pdf.

Anonymous:2005:RL

[Ano05a] Anonymous. 2004 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 16 (1):93–96, January 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2005/01/
                                        l0093.pdf.

Anonymous:2005:AAI

[Ano05b] Anonymous. Additions to 2004 Annual Index. IEEE Transactions on Parallel and Distributed Systems, 16(3):286– 287, March 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2005/03/
                                        l0286.pdf.

Anonymous:2005:CPS

[Ano05c] Anonymous. Call for papers for special issue on high performance computational biology. IEEE Transactions on Parallel and Distributed Systems, 16(3):288, March 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2005/03/
l0288.pdf.

Anonymous:2006:RL

[Ano06] Anonymous. 2005 reviewers list.

IEEE Transactions on Parallel and Distributed Systems, 17 (1):92–96, January 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/;
http://csdl.computer.org/
comp/trans/td/2006/01/l0092.
pdf.

Anonymous:2007:AI

[Ano07a] Anonymous. 2006 annual index.

IEEE Transactions on Parallel and Distributed Systems, 18(1): 130–142, January 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/;
http://csdl.computer.org/
comp/trans/td/2007/01/l0130.
pdf.

Anonymous:2007:RL

[Ano07b] Anonymous. 2006 reviewers list.

IEEE Transactions on Parallel and Distributed Systems, 18(1): 126–129, January 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2007/01/
l0126.pdf.

Anonymous:2007:CPS

[Ano07c] Anonymous. Call for papers: Special issue on power-aware and distributed systems (PA-PADS). IEEE Transactions on Parallel and Distributed Systems, 18(7):1024, July 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/;
 http://csdl.computer.org/
comp/trans/td/2007/07/l1024.
pdf.

Anonymous:2008:AI

[Ano08a] Anonymous. 2007 annual index. IEEE Transactions on Parallel and Distributed Systems, 19(1): 126–144, January 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/;
 http://csdl.computer.org/
comp/trans/td/2008/01/ttd2008010126.
pdf.

Anonymous:2008:RL

[Ano08b] Anonymous. 2007 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 19(1): 121–125, January 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2008/01/
ttd2008010121.pdf.

Anonymous:2008:CPS

[Ano08c] Anonymous. Call for papers for a special issue on distributed cyber-physical systems. IEEE Transactions on Parallel and Distributed Systems, 19(8): 1150–1151, August 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/;
 http://csdl.computer.org/
comp/trans/td/2008/08/ttd2008081150.
pdf.

Anonymous:2008:TAI

[Ano08d] Anonymous. TPDS 2008 annual index. IEEE Transactions on Parallel and Distributed Systems, 19(12):INDEX:1–INDEX:16, December 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2009:RL

[Ano09a] Anonymous. 2008 reviewers list.

IEEE Transactions on Parallel and Distributed Systems, 20(1): 140–144, January 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2009:CPSb

[Ano09b] Anonymous. Call for papers for a special issue on many task computing. IEEE Transactions on Parallel and Distributed Systems, 20(9):1392, September 2009. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2009:CPSa

[Ano09c]  Anonymous.  Call for pa-

pers for special issue on highperformance computing with accelerator. IEEE Transactions on Parallel and Distributed Systems, 20(6):912, June 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2009:TAI

[Ano09d] Anonymous. TPDS 2009 annual index. IEEE Transactions on Parallel and Distributed Systems, 20(12):INDEX:1–INDEX:17, December 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Antonio:1994:CCH

[Ant94] John K. Antonio. Concurrent communication in highspeed wide area networks. IEEE Transactions on Parallel and Distributed Systems, 5(3):264– 273, March 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abali:1993:BPS

[AOB93] Bulent Abali, Fusun Ozguner, and Abdulla Bataineh. Balanced parallel sort on hypercube multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 4(5):572–581, May 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Al-Oqily:2009:SFR

[AOK09] Ibrahim Al-Oqily and Ahmed Karmouch. SORD: a faultresilient service overlay for MediaPort resource discovery. IEEE Transactions on Parallel and Distributed Systems, 20(8):1112– 1125, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Andonov:1997:KVA

[AR97] Rumen Andonov and Sanjay Rajopadhye. Knapsack on VLSI: from algorithm to optimal circuit. IEEE Transactions on Parallel and Distributed Systems, 8(6):545–561, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0545.
pdf;   http://www.computer.
org/tpds/td1997/l0545abs.
htm.

Arazi:2008:CED

[Ara08]  Benjamin Arazi.  Computa-

tionally efficient distributed and delegated certification. IEEE Transactions on Parallel and Distributed Systems, 19(9):1167– 1174, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aronson:2000:HRH

[Aro00] L. D. Aronson. Homogeneous routing for homogeneous traffic patterns on meshes. IEEE Transactions on Parallel and Distributed Systems, 11(8):781– ??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0781.
pdf;   http://www.computer.
org/tpds/td2000/l0781abs.
htm.

Arvind:1994:PCS

[Arv94] K. Arvind. Probabilistic clock synchronization in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 5(5):474–487, May 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ahluwalia:1992:PAC

[AS92] A. K. Ahluwalia and M. Singhal. Performance analysis of the communication architecture of the Connection Machine. IEEE Transactions on Parallel and Distributed Systems, 3(6):728– 738, November 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Allen:1996:IDH

[AS96] James D. Allen and David E. Schimmel. Issues in the design of high performance SIMD architectures. IEEE Transactions on Parallel and Distributed Systems, 7(8):818–829, August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0818abs.htm.

Abdelzaher:1999:CTM

[AS99] T. F. Abdelzaher and K. G.

Shin. Combined task and message scheduling in distributed real-time systems. IEEE Transactions on Parallel and Distributed Systems, 10(11):179–??, November 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1179.
pdf;   http://www.computer.
org/tpds/td1999/l1179abs.
htm.

Alleyne:2000:ETN

[AS00] B. D. Alleyne and I. D. Scherson. On evil twin networks and the value of limited randomized routing. IEEE Transactions on Parallel and Distributed Systems, 11(9):910–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0910.
pdf;   http://www.computer.
org/tpds/td2000/l0910abs.
htm.

Abdelzaher:2002:PGW

[ASB02] Tarek F. Abdelzaher, Kang G.

Shin, and Nina Bhatti. Performance guarantees for Web server end-systems: a controltheoretical approach. IEEE Transactions on Parallel and Distributed Systems, 13(1):80– 96, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0080.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0080abs.
                                        htm.

Alfaro:2004:QIS

[ASD04] Francisco J. Alfaro, José L. Sánchez, and José Duato. QoS in InfiniBand subnetworks. IEEE Transactions on Parallel and Distributed Systems, 15(9):810– 823, September 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/09/l0810.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/09/l0810.pdf.

Agrawal:1995:IRC

[ASS95] Gagan Agrawal, Alan Sussman, and Joel Saltz. An integrated runtime and compile-time approach for parallelizing structured and block structured applications. IEEE Transactions on Parallel and Distributed Systems, 6(7):747–754, July 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0747abs.htm.

Avril:2001:RBC

[AT01] Hervé Avril and Carl Tropper. On rolling back and checkpointing in time warp. IEEE Transactions on Parallel and Distributed Systems, 12(11):1105– 1121, November 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l1105.
pdf;   http://www.computer.
org/tpds/td2001/l1105abs.
htm.

Ang:2007:AOS

[AT07] Chee-Wei Ang and Chen-Khong Tham. Analysis and optimization of service availability in a HA cluster with loaddependent machine availability. IEEE Transactions on Parallel and Distributed Systems, 18 (9):1307–1319, September 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Acharya:1992:IPS

[ATG92] Anurag Acharya, Milind Tambe, and Anoop Gupta. Implementation of production systems on message-passing computers. IEEE Transactions on Parallel and Distributed Systems, 3 (4):477–487, July 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

AlZain:2008:EHL

[ATML08] Abdallah D. Al Zain, Phil W.

Trinder, Greg J. Michaelson, and Hans-Wolfgang Loidl. Evaluating a high-level parallel language (GpH) for computational GRIDs. IEEE Transactions on Parallel and Distributed Systems, 19(2):219–233, February 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Adve:1994:PAM

[AV94] Vikram S. Adve and Mary K. Vernon. Performance analysis of mesh interconnection networks with deterministic routing. IEEE Transactions on Parallel and Distributed Systems, 5 (3):225–246, March 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ali:1996:EBR

[AV96] Arshad Ali and Ramachandran Vaidyanathan. Exact bounds on running ASCEND/DESCEND and FAN-IN algorithms on synchronous multiple bus networks. IEEE Transactions on Parallel and Distributed Systems, 7(8): 783–790, August 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0783abs.htm.

Avresky:1999:ERS

[Avr99] D. R. Avresky. Embedding and reconfiguration of spanning trees in faulty hypercubes. IEEE Transactions on Parallel and Distributed Systems, 10(3):211– ??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0211.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0211abs.

htm. See erratum [Ano99h].

Abbasi:2009:MAC

[AYA09] Ameer Ahmed Abbasi, Mohamed Younis, and Kemal Akkaya. Movement-assisted connectivity restoration in wireless sensor and actor networks. IEEE Transactions on Parallel and Distributed Systems, 20(9):1366– 1379, September 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Black:1990:ILI

[BA90] Andrew P. Black and Yeshayahu Artsy. Implementing location independent invocation. IEEE Transactions on Parallel and Distributed Systems, 1(1):107– 119, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhandarkar:1997:PCV

[BA97] Suchendra M. Bhandarkar and Hamid R. Arabnia. Parallel computer vision on a reconfigurable multiprocessor network. IEEE Transactions on Parallel and Distributed Systems, 8(3):292–309, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0292.
pdf;   http://www.computer.
org/tpds/td1997/l0292abs.
htm.

Bajaj:2004:IST

[BA04] Rashmi Bajaj and Dharma P.

Agrawal. Improving scheduling of tasks in a heterogeneous environment. IEEE Transactions on Parallel and Distributed Systems, 15(2):107–118, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0107abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0107.
pdf.

Barla:2007:HPB

[BA07] Berkant Barla and Cambazoglu Cevdet Aykanat. Hypergraph-partitioning-based remapping models for image-space-parallel direct volume rendering of unstructured Grids. IEEE Transactions on Parallel and Distributed Systems, 18(1):3–16, January 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ben-Asher:2001:PSS

[BAH01] Y. Ben-Asher and G. Haber. Parallel solutions of simple indexed recurrence equations. IEEE Transactions on Parallel and Distributed Systems, 12 (1):22–??, January 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l0022.
pdf;   http://www.computer.
org/tpds/td2001/l0022abs.
htm.

Baker:2005:AES

[Bak05] Theodore P. Baker. An analysis of EDF schedulability on a multiprocessor. IEEE Transactions on Parallel and Distributed Systems, 16(8):760–768, August 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Barlas:1998:CAO

[Bar98] G. D. Barlas. Collection-aware optimum sequencing of operations and closed-form solutions for the distribution of a divisible load on arbitrary processor trees. IEEE Transactions on Parallel and Distributed Systems, 9(5):429–??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0429.
pdf;   http://www.computer.
org/tpds/td1998/l0429abs.
htm.

Bambha:2005:JAM

[BB05] Neal K. Bambha and Shuvra S.

Bhattacharyya. Joint application mapping/interconnect synthesis techniques for embedded chip-scale multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 16(2): 99–112, February 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl2.computer.
org/comp/trans/td/2005/02/
l0099supp.pdf.

Batsakis:2008:NCW

[BB08] Alexandros Batsakis and Randal Burns. NFS–CD: Write-enabled cooperative caching in NFS. IEEE Transactions on Parallel and Distributed Systems, 19(3):

323–333, March 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bala:1995:CPT

[BBC^{+}95] Vasanth Bala, Jehoshua Bruck, Robert Cypher, Pablo Elustondo, Alex Ho, Ching-Tien Ho, Shlomo Kipnis, and Marc Snir. CCL: a portable and tunable collective communication library for scalable parallel computers. IEEE Transactions on Parallel and Distributed Systems, 6(2): 154–164, February 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0154abs.htm.

Banino:2004:SSM

[BBC^{+}04] Cyril Banino, Olivier Beaumont, Larry Carter, Jeanne Ferrante, Arnaud Legrand, and Yves Robert. Scheduling strategies for master-slave tasking on heterogeneous processor platforms. IEEE Transactions on Parallel and Distributed Systems, 15(4):319–330, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0319abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0319.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0319.
pdf.

Barcaccia:2000:CML

[BBD00] P. Barcaccia, M. A. Bonuccelli, and M. Di Ianni. Complexity of minimum length scheduling for precedence constrained messages in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 11(10):90– ??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1090.
pdf;   http://www.computer.
org/tpds/td2000/l1090abs.
htm.

Bhagavathi:1995:TOV

[BBG^{+}95] Dharmavani Bhagavathi, Venkata V.

Bokka, Himabindu Gurla, Stephan Olariu, James L. Schwing, Ivan Stojmenovic, and Jingyuan Zhang. Time-optimal visibilityrelated algorithms on meshes with multiple broadcasting. IEEE Transactions on Parallel and Distributed Systems, 6 (7):687–703, July 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0687abs.htm.

Bestavros:2005:ILM

[BBH05] Azer Bestavros, John W. Byers, and Khaled A. Harfoush. Inference and labeling of metricinduced network topologies. IEEE Transactions on Parallel and Distributed Systems, 16(11): 1053–1065, November 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bridgewater:2007:BON

[BBR07] Jesse Bridgewater, P. Oscar Boykin, and Vwani Roychowdhury. Balanced overlay networks (BON): An overlay technology for decentralized load balancing. IEEE Transactions on Parallel and Distributed Systems, 18(8): 1122–1133, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beaumont:2001:MMH

[BBRR01] Olivier Beaumont, Vincent Boudet, Fabrice Rastello, and Yves Robert. Matrix multiplication on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 12(10):1033–1051, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l1033.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1033abs.
                                        htm.

Barsi:2009:ACT

[BBS^{+}09] Ferruccio Barsi, Alan A. Bertossi, Francesco Betti Sorbelli, Roberto Ciotti, Stephan Olariu, and M. Cristina Pinotti. Asynchronous corona training protocols in wireless sensor and actor networks. IEEE Transactions on Parallel and Distributed Systems, 20(8):1216–1230, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bucher:1992:MAD

[BC92] I. Y. Bucher and D. A. Calahan. Models of access delays in multiprocessor memories. IEEE Transactions on Parallel and Distributed Systems, 3 (3):270–280, May 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Brassil:1995:BMD

[BC95] Jack T. Brassil and Rene L.

Cruz. Bounds on maximum delay in networks with deflection routing. IEEE Transactions on Parallel and Distributed Systems, 6(7):724–732, July 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0724abs.htm.

Boppana:1996:FDD

[BC96] Rajendra V. Boppana and Suresh Chalasani. A framework for designing deadlock-free wormhole routing algorithms. IEEE Transactions on Parallel and Distributed Systems, 7(2): 169–183, February 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0169abs.htm.

Boppana:1999:FTC

[BC99] R. V. Boppana and S. Chalasani. Fault-tolerant communication with partitioned dimension-order routers. IEEE Transactions on Parallel and Distributed Systems, 10(10):26– ??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1026.
pdf;   http://www.computer.
org/tpds/td1999/l1026abs.
htm.

Bermudez:2006:FRC

[BC06] Aurelio Bermudez and Rafael Casado. Fast routing computation on InfiniBand networks. IEEE Transactions on Parallel and Distributed Systems, 17(3): 215–226, March 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Balbo:1992:EME

[BCBzC92] Gianfranco Balbo, Giovanni Chiola, Steven C. Bruell, and Po zung Chen. An example of modeling and evaluation of a concurrent program using colored stochastic Petri nets: Lamport’s fast mutual exclusion algorithm. IEEE Transactions on Parallel and Distributed Systems, 3(2):221–240, March 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Barker:2004:LBF

[BCCP04] Kevin Barker, Andrey Chernikov, Nikos Chrisochoides, and Keshav Pingali. A load balancing framework for adaptive and asynchronous applications. IEEE Transactions on Parallel and Distributed Systems, 15(2):183–192, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0183abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0183.
pdf.

Bessani:2009:SMB

[BCdSFL09] Alysson Neves Bessani, Miguel Correia, Joni da Silva Fraga, and Lau Cheuk Lung. Sharing memory between Byzantine processes using policy-enforced tuple spaces. IEEE Transactions on Parallel and Distributed Systems, 20(3):419–432, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beaumont:2008:CVD

[BCF^{+}08] Olivier Beaumont, Larry Carter, Jeanne Ferrante, Arnaud Legrand, Loris Marchal, and Yves Robert. Centralized versus distributed schedulers for bag-of-tasks applications. IEEE Transactions on Parallel and Distributed Systems, 19(5):698–709, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Banerjee:2008:AIR

[BcFGM08] Amitabha Banerjee, Wu chun Feng, Dipak Ghosal, and Biswanath Mukherjee. Algorithms for integrated routing and scheduling for aggregating data from distributed resources on a Lambda Grid. IEEE Transactions on Parallel and Distributed Systems, 19(1):24–34, January 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bononi:2004:ROI

[BCG04] Luciano Bononi, Marco Conti, and Enrico Gregori. Runtime optimization of IEEE 802.11 wireless LANs performance. IEEE Transactions on Parallel and Distributed Systems, 15(1):66–80, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/01/l0066abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/01/l0066.
                                        pdf.

Bruck:1994:FTB

[BCH94] Jehoshua Bruck, Robert Cypher, and Ching-Tien Ho. Faulttolerant de Bruijn and shuffleexchange networks. IEEE Transactions on Parallel and Distributed Systems, 5(5):548– 553, May 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bronson:1990:EAD

[BCJ90] Edward C. Bronson, Thomas L. Casavant, and Leah H. Jamieson. Experimental application-driven architecture analysis of an SIMD/MIMD parallel processing system. IEEE Transactions on Parallel and Distributed Systems, 1(2):195–205, April 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beaumont:2005:SDL

[BCL^{+}05] Olivier Beaumont, Henri Casanova, Arnaud Legrand, Yves Robert, and Yang Yang. Scheduling divisible loads on star and tree networks: Results and open problems. IEEE Transactions on Parallel and Distributed Systems, 16(3):207–218, March 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bertogna:2009:SAG

[BCL09] Marko Bertogna, Michele Cirinei, and Giuseppe Lipari. Schedulability analysis of global scheduling algorithms on multiprocessor platforms. IEEE Transactions on Parallel and Distributed Systems, 20(4):553–566, April 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bermudez:2007:HTC

[BCQD07] Aurelio Bermúdez, Rafael Casado, Francisco J. Quiles, and José Duato. Handling topology changes in InfiniBand. IEEE Transactions on Parallel and Distributed Systems, 18(2):172– 185, February 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2007/02/
l0172s.pdf.

Boppana:1998:RDP

[BCR98] R. V. Boppana, S. Chalasani, and C. S. Raghavendra. Resource deadlocks and performance of wormhole multicast routing algorithms. IEEE Transactions on Parallel and Distributed Systems, 9(6):535–??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0535.
pdf;   http://www.computer.
org/tpds/td1998/l0535abs.
htm.

Bahi:2005:DLB

[BCVC05] Jacques M. Bahi, Sylvain Contassot-Vivier, and Raphael Couturier. Dynamic load balancing and efficient load estimators for asynchronous iterative algorithms. IEEE Transactions on Parallel and Distributed Systems, 16(4):289–299, April 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bahi:2005:DCD

[BCVCV05] Jacques M. Bahi, Sylvain Contassot-Vivier, Raphaël Couturier, and Flavien Vernier. A decentralized convergence detection algorithm for asynchronous parallel iterative algorithms. IEEE Transactions on Parallel and Distributed Systems, 16(1):4–13, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0004abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0004.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0004.
pdf.

Bruck:1996:DIB

[BDD^{+}96] Jehoshua Bruck, Luc De Coster, Natalie Dewulf, Ching-Tien Ho, and Rudy Lauwereins. On the design and implementation of broadcast and global combine operations using the postal model. IEEE Transactions on Parallel and Distributed Systems, 7(3):256–265, March 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0256abs.htm.

Bagherzadeh:1995:WBE

[BDL95] Nader Bagherzadeh, Martin Dowd, and Shahram Latifi. A well-behaved enumeration of star graphs. IEEE Transactions on Parallel and Distributed Systems, 6(5):531–535, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0531abs.htm.

Bucci:1994:PAT

[BDS94] Giacomo Bucci, Alberto Del Bimbo, and Simone Santini. Performance analysis of two different algorithms for Ethernet-FDDI interconnection. IEEE Transactions on Parallel and Distributed Systems, 5(6):614– 629, June 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Balsamo:1998:BPM

[BDvD98] S. Balsamo, L. Donatiello, and N. M. van Dijk. Bound performance models of heterogeneous parallel processing systems. IEEE Transactions on Parallel and Distributed Systems, 9(10):41–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1041.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1041abs.
                                        htm.

Blume:1992:PAP

[BE92] W. Blume and R. Eigenmann. Performance analysis pf parallelizing compilers on the Perfect Benchmarks programs. IEEE Transactions on Parallel and Distributed Systems, 3(6):643– 656, November 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Blume:1998:NSD

[BE98] W. Blume and R. Eigenmann. Nonlinear and symbolic data dependence testing. IEEE Transactions on Parallel and Distributed Systems, 9(12):180–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/

td/books/td1998/pdf/l1180.
pdf;   http://www.computer.
org/tpds/td1998/l1180abs.
htm.

Bucur:2007:SPP

[BE07] Anca I. D. Bucur and Dick H. J.

Epema. Scheduling policies for processor coallocation in multicluster systems. IEEE Transactions on Parallel and Distributed Systems, 18(7):958–972, July 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Burkhart:1993:ML

[BEK^{+}93] Helmar Burkhart, Rudolf Eigenmann, Heinz Kindlimann, Michael Moser, and Heinz Scholian. M^{3} multiprocessor laboratory. IEEE Transactions on Parallel and Distributed Systems, 4(5):507– 519, May 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bose:2006:SSC

[BES06] Abhijit Bose, Mohamed El Gendy, and Kang G. Shin. Sapphire: Statistical characterization and model-based adaptation of networked applications. IEEE Transactions on Parallel and Distributed Systems, 17(12): 1512–1525, December 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhowmik:2004:GCF

[BF04] Anasua Bhowmik and Manoj Franklin. A general compiler framework for speculative multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 15 (8):713–724, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/08/l0713.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/08/l0713.pdf.

Bohossian:2001:CRR

[BFL^{+}01] V. Bohossian, C. C. Fan, P. S. LeMahieu, M. D. Riedel, L. Xu, and J. Bruck. Computing in the RAIN: a reliable array of independent nodes. IEEE Transactions on Parallel and Distributed Systems, 12(2):99–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0099.
pdf;   http://www.computer.
org/tpds/td2001/l0099abs.
htm.

Berthome:1996:OID

[BFP96] Pascal Berthome, Afonso Ferreira, and Stephane Perennes. Optimal information dissemination in star and pancake networks. IEEE Transactions on Parallel and Distributed Systems, 7(12):1292–1300, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bohm:1990:IIM

[BG90] A. P. Wim Bohm and John R. Gurd. Iterative instructions in the Manchester Dataflow Computer. IEEE Transactions on Parallel and Distributed Systems, 1(2):129–139, April 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Burns:2002:SLO

[BG02] James Burns and Jean-Luc Gaudiot. SMT layout overhead and scalability. IEEE Transactions on Parallel and Distributed Systems, 13(2):142–155, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0142.
pdf;   http://www.computer.
org/tpds/td2001/l0142abs.
htm.

Boukerche:2009:EAT

[BG09] Azzedine Boukerche and Yunfeng Gu. An efficient adaptive transmission control scheme for large-scale distributed simulation systems. IEEE Transactions on Parallel and Distributed Systems, 20(2):246–260, February 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Banikazemi:2001:MLE

[BGBP01] Mohammad Banikazemi, Rama K.

Govindaraju, Robert Blackmore, and Dhabaleswar K. Panda. MPI-LAPI: An efficient implementation of MPI for IBM RS/6000 SP systems. IEEE Transactions on Parallel and Distributed Systems, 12(10):1081–1093, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1081.
pdf;   http://www.computer.
org/tpds/td2001/l1081abs.
htm.

Berten:2006:DSJ

[BGJ06] Vandy Berten, Joel Goossens, and Emmanuel Jeannot. On the distribution of sequential jobs in random brokering for heterogeneous computational Grids. IEEE Transactions on Parallel and Distributed Systems, 17(2): 113–124, February 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bharadwaj:1994:OSA

[BGM94] V. Bharadwaj, D. Ghose, and V. Mani. Optimal sequencing and arrangement in distributed single-level tree networks with communication delays. IEEE Transactions on Parallel and Distributed Systems, 5(9):968– 976, September 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Blelloch:1997:AMB

[BGMZ97] G. E. Blelloch, P. B. Gibbons, Y. Matias, and M. Zagha. Accounting for memory bank contention and delay in high-bandwidth multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 8(9):943–958, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0943.
pdf;   http://www.computer.
org/tpds/td1997/l0943abs.
htm.

Bhagavathi:1996:SMO

[BGO^{+}96] Dharmavani Bhagavathi, Himabindu Gurla, Stephan Olariu, James L. Schwing, and Jingyuan Zhang. Square meshes are not optimal for convex hull computation. IEEE Transactions on Parallel and Distributed Systems, 7 (6):545–554, June 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0545abs.htm.

Bokka:1997:TOD

[BGO^{+}97] Venkatavasu Bokka, Himabindu Gurla, Stephan Olariu, James L. Schwing, and Larry Wilson. Time-optimal domain-specific querying on enhanced meshes. IEEE Transactions on Parallel and Distributed Systems, 8 (1):13–24, January 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhagavathi:1998:TVO

[BGO^{+}98] D. Bhagavathi, H. Gurla, S. Olariu, J. L. Schwing, L. Wilson, and J. Zhang. Time- and VLSI-optimal sorting on enhanced meshes. IEEE Transactions on Parallel and Distributed Systems, 9(10):929–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0929.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0929abs.
                                        htm.

Bokka:1997:PBT

[BGOS97] V. Bokka, H. Gurla, S. Olariu, and J. L. Schwing. Podalitybased time-optimal computations on enhanced meshes. IEEE Transactions on Parallel and Distributed Systems, 8(10):1019– 1035, October 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td1997/pdf/l1019.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1019abs.
                                        htm.

Bokka:1998:CTA

[BGOS98] V. Bokka, H. Gurla, S. Olariu, and J. L. Schwing. Constanttime algorithms for constrained triangulations on reconfigurable meshes. IEEE Transactions on Parallel and Distributed Systems, 9(11):57–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1057.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1057abs.
                                        htm.

Boukerche:2002:ESS

[BHJ02] Azzedine Boukerche, Sungbum Hong, and Tom Jacob. An efficient synchronization scheme of multimedia streams in wireless and mobile systems. IEEE Transactions on Parallel and Distributed Systems, 13(9):911–923, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0911.
pdf;   http://www.computer.
org/tpds/td2002/l0911abs.
htm.

Bruck:1997:EAA

[BHK^{+}97] J. Bruck, Ching-Tien Ho, S. Kipnis, E. Upfal, and D. Weathersby. Efficient algorithms for all-to-all communications in multiport message-passing systems. IEEE Transactions on Parallel and Distributed Systems, 8(11):1143–1156, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1143.
pdf;   http://www.computer.
org/tpds/td1997/l1143abs.
htm.

Bohacek:2007:GTS

[BHL^{+}07] Stephan Bohacek, Joao Hespanha, Junsoo Lee, Chansook Lim, and Katia Obraczka. Game theoretic stochastic routing for fault tolerance and security in computer networks. IEEE Transactions on Parallel and Distributed Systems, 18(9):1227– 1240, September 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhuyan:2006:EN

[Bhu06a] Laxmi N. Bhuyan. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 17 (5):401–402, May 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2006/05/
                                        l0401.pdf.

Bhuyan:2006:EMN

[Bhu06b] Laxmi Narayan Bhuyan. Editorial: a message from the new Editor-in-Chief. IEEE Transactions on Parallel and Distributed Systems, 17(1):2, January 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2006/01/l0002.pdf.

Bhuyan:2007:ENa

[Bhu07a] Laxmi N. Bhuyan. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 18 (1):1–2, January 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2007/01/
                                        l0001.pdf.

Bhuyan:2007:ENb

[Bhu07b] Laxmi N. Bhuyan. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 18(9):1185–1187, September 2007. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2007/09/l1185.pdf.

Bhuyan:2008:EN

[Bhu08] Laxmi N. Bhuyan. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 19 (1):1–3, January 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2008/01/
ttd2008010001.pdf.

Bhuyan:2009:EEF

[Bhu09a] Laxmi N. Bhuyan. Editorial:

EIC farewell and new EIC introduction. IEEE Transactions on Parallel and Distributed Systems, 20(12):1713, December 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhuyan:2009:ENa

[Bhu09b] Laxmi N. Bhuyan. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 20(2): 145–146, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhuyan:2009:ENb

[Bhu09c] Laxmi N. Bhuyan. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 20(8): 1073–1074, August 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bioch:1995:GAN

[BI95] Jan C. Bioch and Toshihide Ibaraki. Generating and approximating nondominated coteries. IEEE Transactions on Parallel and Distributed Systems, 6(9): 905–914, September 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0905abs.htm.

Bhuyan:1997:PMB

[BIA^{+}97] Laxmi N. Bhuyan, Ravi R. Iyer, Tahsin Askar, Ashwini K. Nanda, and Mohan Kumar. Performance of multistage bus networks for a distributed shared memory multiprocessor. IEEE Transactions on Parallel and Distributed Systems, 8(1):82–95, January 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Billionnet:1994:ATS

[Bil94] Alain Billionnet. Allocating tree structured programs in a distributed system with uniform communication costs. IEEE Transactions on Parallel and Distributed Systems, 5(4):445– 448, April 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Birk:1993:LCD

[Bir93] Yitzhak Birk. On-line control and deadlock-avoidance in a page-parallel multiprocessor rasterizer. IEEE Transactions on Parallel and Distributed Systems, 4(2):155–163, February 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhuyan:2000:ICN

[BIWK00] L. N. Bhuyan, R. Iyer, H. Wang, and A. Kumar. Impact of CC-NUMA memory management policies on the application performance of multistage network. IEEE Transactions on Parallel and Distributed Systems, 11(3):230–??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0230.
pdf;   http://www.computer.
org/tpds/td2000/l0230abs.
htm.

Bertozzi:2005:NSF

[BJM^{+}05] Davide Bertozzi, Antoine Jalabert, Srinivasan Murali, Rutuparna Tamhankar, Stergios Stergiou, Luca Benini, and Giovanni De Micheli. NoC synthesis flow for customized domain specific multiprocessor systems-on-chip. IEEE Transactions on Parallel and Distributed Systems, 16(2): 113–129, February 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Banerjee:1990:PSA

[BJS90] Prithviraj Banerjee, Mark Howard Jones, and Jeff S. Sargent. Parallel simulated annealing algorithms for cell placement on hypercube multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 1(1):91– 106, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Banerjee:2009:BRL

[BK09] Amit Banerjee and Chung-Ta King. Building ring-like overlays on wireless ad hoc and sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(11):1553–1566, November 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Baran:1996:PAT

[BKB96] Benjamin Barán, Eugenius Kaszkurewicz, and Amit Bhaya. Parallel asynchronous team algorithms: Convergence and performance analysis. IEEE Transactions on Parallel and Distributed Systems, 7(7):677– 688, July 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0677abs.
                                        htm.

Basile:2006:ARM

[BKI06] Claudio Basile, Zbigniew Kalbarczyk, and Ravishankar K. Iyer. Active replication of multithreaded applications. IEEE Transactions on Parallel and Distributed Systems, 17(5):448– 465, May 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.

org/comp/trans/td/2006/05/
l0448s.pdf.

Bansal:2003:IDS

[BKS03] Savina Bansal, Padam Kumar, and Kuldip Singh. An improved duplication strategy for scheduling precedence constrained graphs in multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 14(6):533–544, June 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/06/l0533abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/06/l0533.
pdf.

Barbara:1991:CRS

[BL91] Daniel Barbara and Richard J.

Lipton. A class of randomized strategies for low-cost comparison of file copies. IEEE Transactions on Parallel and Distributed Systems, 2(2):160–170, April 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Baydal:2005:FMC

[BLD05] Elvira Baydal, Pedro Lopez, and Jose Duato. A family of mechanisms for congestion control in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 16(9): 772–784, September 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beaumont:2005:PBH

[BLMR05] Olivier Beaumont, Arnaud Legrand, Loris Marchal, and Yves Robert. Pipelining broadcasts on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 16(4):300–313, April 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhagavathi:1994:FSA

[BLO^{+}94] D. Bhagavathi, P. J. Looges, S. Olariu, J. L. Schwing, and J. Zhang. Fast selection algorithm for meshes with multiple broadcasting. IEEE Transactions on Parallel and Distributed Systems, 5(7):772–778, July 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beaumont:2003:MSP

[BLR03] Olivier Beaumont, Arnaud Legrand, and Yves Robert. The master-slave paradigm with heterogeneous processors. IEEE Transactions on Parallel and Distributed Systems, 14(9):897–908, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/09/l0897abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/09/l0897.
                                        pdf.

Bertossi:2000:CTD

[BM00a] A. A. Bertossi and A. Mei.

Constant time dynamic programming on directed reconfigurable networks. IEEE Transactions on Parallel and Distributed Systems, 11(6):529–??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0529.
pdf;   http://www.computer.
org/tpds/td2000/l0529abs.
htm.

Bertossi:2000:RNS

[BM00b] A. A. Bertossi and A. Mei. A residue number system on reconfigurable mesh with applications to prefix sums and approximate string matching. IEEE Transactions on Parallel and Distributed Systems, 11(11):186–??, November 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1186.
pdf;   http://www.computer.
org/tpds/td2000/l1186abs.
htm.

Basagni:2006:LPA

[BMPP06] Stefano Basagni, Michele Mastrogiovanni, Alessandro Panconesi, and Chiara Petrioli. Localized protocols for ad hoc clustering and backbone formation: a performance comparison. IEEE Transactions on Parallel and Distributed Systems, 17 (4):292–306, April 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bertossi:1999:FTR

[BMR99] A. A. Bertossi, L. V. Mancini, and F. Rossini. Fault-tolerant rate-monotonic first-fit scheduling in hard-real-time systems. IEEE Transactions on Parallel and Distributed Systems, 10(9):934–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0934.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0934abs.
                                        htm.

Bar-Noy:1995:CGC

[BNBH^{+}95] Amotz Bar-Noy, Jehoshua Bruck, Ching-Tien Ho, Shlomo Kipnis, and Baruch Schieber. Computing global combine operations in the multiport postal model. IEEE Transactions on Parallel and Distributed Systems, 6(8): 896–900, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0896abs.htm.

Bar-Noy:1999:BMM

[BNH99] A. Bar-Noy and C.-T. Ho. Broadcasting multiple messages in the multiport model. IEEE Transactions on Parallel and Distributed Systems, 10(5):500– ??, May 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0500.
pdf;   http://www.computer.
org/tpds/td1999/l0500abs.
htm.

Bokka:2001:OAM

[BNO^{+}01] Venkatavasu Bokka, Koji Nakano, Stephen Olariu, James L. Schwing, and Larry Wilson. Optimal algorithms for the multiple query problem on reconfigurable meshes, with applications. IEEE Transactions on Parallel and Distributed Systems, 12(9):875–887, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0875.
pdf;   http://www.computer.
org/tpds/td2001/l0875abs.
htm.

Balakrishnan:1998:PDF

[BÖ98] S. Balakrishnan and F. Özgüner.

A priority-driven flow control mechanism for real-time traffic in multiprocessor networks. IEEE Transactions on Parallel and Distributed Systems, 9 (7):664–??, July 1998. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1998/pdf/l0664.
pdf;   http://www.computer.
org/tpds/td1998/l0664abs.
htm.

Bozdag:2009:CST

[BOC09] Doruk Bozdağ, Fusun Ozguner, and Umit V. Catalyurek. Compaction of schedules and a twostage approach for duplicationbased DAG scheduling. IEEE Transactions on Parallel and Distributed Systems, 20(6):857– 871, June 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bokhari:1993:NFM

[Bok93] Shahid H. Bokhari. Network flow model for load balancing in circuit-switched multicomputers. IEEE Transactions on Parallel and Distributed Systems, 4 (6):649–657, June 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bertossi:2004:CMS

[BOPZ04] Alan A. Bertossi, Stephan Olariu, M. Cristina Pinotti, and Si-Qing Zheng. Classifying matrices separating rows and columns. IEEE Transactions on Parallel and Distributed Systems, 15(7):654–665, July 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/07/l0654.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/07/l0654.pdf.

Bordawekar:2000:QCA

[Bor00]  R. Bordawekar.  Quantita-

tive characterization and analysis of the I/O behavior of a commercial distributed-shared-memory machine. IEEE Transactions on Parallel and Distributed Systems, 11(5):509–??, May 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0509.
pdf;   http://www.computer.
org/tpds/td2000/l0509abs.
htm.

Blough:1994:ACF

[BP94] Douglas M. Blough and Andrzej Pelc. Almost certain fault diagnosis through algorithm-based fault tolerance. IEEE Transactions on Parallel and Distributed Systems, 5(5):532–539, May 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Basak:1996:DCM

[BP96] Debashis Basak and Dhabaleswar K. Panda. Designing clustered multiprocessor systems under packaging and technological advancements. IEEE Transactions on Parallel and Distributed Systems, 7(9):962– 978, September 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0962abs.htm.

Basak:1998:ACC

[BP98] D. Basak and D. K. Panda. Alleviating consumption channel bottleneck in wormhole-routed k-ary n-cube systems. IEEE Transactions on Parallel and Distributed Systems, 9(5):481– ??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0481.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0481abs.
                                        htm.

Bosque:2006:PCM

[BP06] Jose Luis Bosque and Luis Pastor. A parallel computational model for heterogeneous clusters. IEEE Transactions on Parallel and Distributed Systems, 17 (12):1390–1400, December 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bertossi:2003:CAS

[BPT03] Alan A. Bertossi, Cristina M. Pinotti, and Richard B. Tan. Channel assignment with separation for interference avoidance in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 14(3):222–235, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/03/l0222abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/03/l0222.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/03/l0222.
                                        pdf.

Baldoni:1999:IBC

[BQF99] R. Baldoni, F. Quaglia, and P. Fornara. An index-based checkpointing algorithm for autonomous distributed systems. IEEE Transactions on Parallel and Distributed Systems, 10(2):181–??, February 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0181.
pdf;   http://www.computer.
org/tpds/td1999/l0181abs.
htm.

Bernhard:1991:UDP

[BR91] P. J. Bernhard and D. J.

Rosenkrantz. Using the dual path property of omega networks to obtain conflict-free message routing. IEEE Transactions on Parallel and Distributed Systems, 2(4):503–507, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bernhard:1994:PMP

[BR94] P. J. Bernhard and D. J.

Rosenkrantz. Partitioning message patterns for bundled omega networks. IEEE Transactions on Parallel and Distributed Systems, 5(4):353–363, April 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Borodin:1997:DMM

[BRS97] Allan Borodin, Yuval Rabani, and Baruch Schieber. Deterministic many-to-many hot potato routing. IEEE Transactions on Parallel and Distributed Systems, 8(6):587–596, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0587.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0587abs.
                                        htm.

Banik:2007:MRD

[BRS07] Shankar M. Banik, Sridhar Radhakrishnan, and Chandra N. Sekharan. Multicast routing with delay and delay variation constraints for collaborative applications on overlay networks. IEEE Transactions on Parallel and Distributed Systems, 18(3): 421–431, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Benoit:2008:RPA

[BRSR08] Anne Benoit, Veronika Rehn-Sonigo, and Yves Robert. Replica placement and access policies in tree networks. IEEE Transactions on Parallel and Distributed Systems, 19(12): 1614–1627, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Banik:2008:IDF

[BRSS08] Shankar M. Banik, Sridhar Radhakrishnan, Venkatesh Sarangan, and Chandra N. Sekharan. Implementation of distributed floor control protocols on overlay networks. IEEE Transactions on Parallel and Distributed Systems, 19(8):1057–1070, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Butler:2009:LIB

[BRTM09] Kevin R. B. Butler, Sunam Ryu, Patrick Traynor, and Patrick D. McDaniel. Leveraging identitybased cryptography for node ID assignment in structured P2P systems. IEEE Transactions on Parallel and Distributed Systems, 20(12):1803–1815, December 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bakken:1995:SFT

[BS95] David E. Bakken and Richard D.

Schlichting. Supporting faulttolerant parallel programming in Linda. IEEE Transactions on Parallel and Distributed Systems, 6(3):287–302, March 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0287abs.htm.

Barak:1996:ECC

[BS96] Amnon Barak and Eugen Schenfeld. Embedding classical communication topologies in the scalable OPAM architecture. IEEE Transactions on Parallel and Distributed Systems, 7(9): 979–992, September 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0979abs.htm.

Boukerche:2008:NAM

[BS08] Azzedine Boukerche and Samer Samarah. A novel algorithm for mining association rules in wireless ad hoc sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19 (7):865–877, July 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Baysan:2009:PTS

[BSCB09] Mehmet Baysan, Kamil Sarac, Ramaswamy Chandrasekaran, and Sergey Bereg. A polynomial time solution to minimum forwarding set problem in wireless networks under unit disk coverage model. IEEE Transactions on Parallel and Distributed Systems, 20(7):913–924, July 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Brinkmeier:2009:ORP

[BSS09] Michael Brinkmeier, Gunter Schafer, and Thorsten Strufe. Optimally DoS resistant P2P topologies for live multimedia streaming. IEEE Transactions on Parallel and Distributed Systems, 20(6):831–844, June 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Boukerche:1998:DGA

[BT98] A. Boukerche and C. Tropper. A distributed graph algorithm for the detection of local cycles and knots. IEEE Transactions on Parallel and Distributed Systems, 9(8):748–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0748.
pdf;   http://www.computer.
org/tpds/td1998/l0748abs.
htm.

Bagrodia:2000:PEC

[BT00] R. L. Bagrodia and M. Takai.

Performance evaluation of conservative algorithms in parallel simulation languages. IEEE Transactions on Parallel and Distributed Systems, 11(4):395– ??, April 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0395.
pdf;   http://www.computer.
org/tpds/td2000/l0395abs.
htm.

Barlas:2005:ODD

[BV05] Gerassimos Barlas and Bharadwaj Veeravalli. Optimized distributed delivery of continuousmedia documents over unreliable communication links. IEEE Transactions on Parallel and Distributed Systems, 16(10):982– 994, October 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Berkey:1994:SBP

[BW94] Judith O. Berkey and Pearl Y.

Wang. A systolic-based parallel bin packing algorithm. IEEE Transactions on Parallel and Distributed Systems, 5(7):769– 772, July 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bik:1996:ADS

[BW96] Aart J. C. Bik and Harry A. G. Wijshoff. Automatic data structure selection and transformation for sparse matrix computations. IEEE Transactions on Parallel and Distributed Systems, 7(2):109–126, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0109abs.htm.

Berman:2003:ACG

[BWC^{+}03] Francine Berman, Richard Wolski, Henri Casanova, Walfredo Cirne, Holly Dail, Marcio Faerman, Silvia Figueira, Jim Hayes, Graziano Obertelli, Jennifer Schopf, Gary Shao, Shava Smallen, Neil Spring, Alan Su, and Dmitrii Zagorodnov. Adaptive computing on the Grid using AppLeS. IEEE Transactions on Parallel and Distributed Systems, 14(4):369–382, April 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0369abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0369.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0369.
pdf.

Bedford:2005:SON

[BWS^{+}05] Michael Bedford, Taylor Walter, Lee Saman, P. Amarasinghe, and Anant Agarwal. Scalar operand networks. IEEE Transactions on Parallel and Distributed Systems, 16(2):145– 162, February 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl2.computer.
org/comp/trans/td/2005/02/
l0145supp.pdf.

Chaudhary:1993:GSM

[CA93] V. Chaudhary and J. K. Aggarwal. A generalized scheme for mapping parallel algorithms. IEEE Transactions on Parallel and Distributed Systems, 4(3): 328–346, March 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Catalyurek:1999:HPB

[ÇA99] Ü. V. Çatalyürek and C. Aykanat.

Hypergraph-partitioning-based decomposition for parallel sparsematrix vector multiplication. IEEE Transactions on Parallel and Distributed Systems, 10 (7):673–??, July 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l0673.
pdf;   http://www.computer.
org/tpds/td1999/l0673abs.
htm.

Chen:1993:DNC

[CAB93] Chienhua Chen, Dharma P.

Agrawal, and J. Richard Burke. dBCube: a new class of hierarchical multiprocessor interconnection networks with area efficient layout. IEEE Transactions on Parallel and Distributed Systems, 4(12):1332–1344, December 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cappello:1992:PTM

[Cap92] Peter Cappello. A processortime-minimal systolic array for cubical mesh algorithms. IEEE Transactions on Parallel and Distributed Systems, 3(1):4–13, January 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chakraborty:1993:PCE

[CARW93] Amal Chakraborty, Donald C. S. Allison, Clvin J. Ribbens, and Layne T. Watson. Parallel complexity of embedding algorithms for the solution of systems of nonlinear equations. IEEE Transactions on Parallel and Distributed Systems, 4 (4):458–465, April 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Colohan:2007:CSL

[CASM07] Christopher B. Colohan, Anastasia Ailamaki, Gregory Steffan, and Todd C. Mowry. CMP support for large and dependent speculative threads. IEEE Transactions on Parallel and Distributed Systems, 18(8):1041– 1054, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Corbera:2004:FCD

[CAZ04] Francisco Corbera, Rafael Asenjo, and Emilio L. Zapata. A framework to capture dynamic data structures in pointer-based codes. IEEE Transactions on Parallel and Distributed Systems, 15(2):151–166, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0151abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0151.
pdf.

Cohen:2000:PPT

[CB00] N. Cohen and J. Brassil. A parallel pruning technique for highly asymmetric assignment problems. IEEE Transactions on Parallel and Distributed Systems, 11(6):550–??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0550.
pdf;   http://www.computer.
org/tpds/td2000/l0550abs.
htm.

Cirne:2003:WHS

[CB03] Walfredo Cirne and Francine Berman. When the herd is smart: Aggregate behavior in the selection of job request. IEEE Transactions on Parallel and Distributed Systems, 14(2):181–192, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/02/l0181abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0181.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0181.
pdf.

Carrera:2005:PCS

[CB05] Enrique V. Carrera and Ricardo Bianchini. PRESS: a clustered server based on user-level communication. IEEE Transactions on Parallel and Distributed Systems, 16(5):385–395, May 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Casado:2001:PDF

[CBD^{+}01] R. Casado, A. Bermúdez, J. Duato, F. J. Quiles, and J. L. Sánchez. A protocol for deadlock-free dynamic reconfiguration in high-speed local area networks. IEEE Transactions on Parallel and Distributed Systems, 12(2):115–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0115.
pdf;   http://www.computer.
org/tpds/td2001/l0115abs.
htm.

Colbrook:1996:AST

[CBDW96] Adrian Colbrook, Eric A. Brewer, Chrysanthos N. Dellarocas, and William E. Weihl. Algorithms for search trees on message-passing architectures. IEEE Transactions on Parallel and Distributed Systems, 7(2): 97–108, February 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0097abs.htm.

Chen:1993:EAS

[CBE93] Yinong Chen, Winfried Bucken, and Klaus Echtle. Efficient algorithms for system diagnosis with both processor and comparator faults. IEEE Transactions on Parallel and Distributed Systems, 4(4):371–381, April 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chellappan:2007:MLF

[CBM^{+}07] Sriram Chellappan, Xiaole Bai, Bin Ma, Dong Xuan, and Changqing Xu. Mobility limited flip-based sensor networks deployment. IEEE Transactions on Parallel and Distributed Systems, 18(2):199–211, February 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chan:1993:ORD

[CC93a] M. Y. Chan and F. Y. L.

Chin. Optimal resilient distributed algorithms for ring election. IEEE Transactions on Parallel and Distributed Systems, 4 (4):475–480, April 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chan:1993:PAE

[CC93b] M. Y. Chan and Francis Chin. Parallel algorithm for an efficient mapping of grids in hypercubes. IEEE Transactions on Parallel and Distributed Systems, 4(8): 933–946, August 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chou:1995:OIS

[CC95] Hong-Chich Chou and Chung-Ping Chung. An optimal instruction scheduler for superscalar processor. IEEE Transactions on Parallel and Distributed Systems, 6(3):303–313, March 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0303abs.htm.

Chen:1997:NOO

[CC97] Chi-Chang Chen and Jianer Chen. Nearly optimal oneto-many parallel routing in star networks. IEEE Transactions on Parallel and Distributed Systems, 8(12):1196–1202, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1196.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1196abs.
                                        htm.

Chiu:1998:EFT

[CC98] G.-M. Chiu and K.-S. Chen.

Efficient fault-tolerant multicast scheme for hypercube multicomputers. IEEE Transactions on Parallel and Distributed Systems, 9(10):952–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0952.
pdf;   http://www.computer.
org/tpds/td1998/l0952abs.
htm.

Chiu:1999:ESA

[CC99] G.-M. Chiu and S.-K. Chen.

An efficient submesh allocation scheme for two-dimensional meshes with little overhead. IEEE Transactions on Parallel and Distributed Systems, 10 (5):471–??, May 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l0471.
pdf;   http://www.computer.
org/tpds/td1999/l0471abs.
htm.

Chen:2001:FTR

[CC01] C.-L. Chen and G.-M. Chiu. A fault-tolerant routing scheme for meshes with nonconvex faults. IEEE Transactions on Parallel and Distributed Systems, 12 (5):467–??, May 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l0467.

                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0467abs.
                                        htm.

Chang:2003:EAE

[CC03] Cheng-Yue Chang and Ming-Syan Chen. On exploring aggregate effect for efficient cache replacement in transcoding proxies. IEEE Transactions on Parallel and Distributed Systems, 14(6):611–624, June 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/06/l0611abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/06/l0611.
                                        pdf.

Chang:2005:DRN

[CCC05] Guey-Yun Chang, Gerard J. Chang, and Gen-Huey Chen. Diagnosabilities of regular networks. IEEE Transactions on Parallel and Distributed Systems, 16(4):314–323, April 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1990:DEP

[CCCS90] Yen-Cheng Chen, Wen-Tsuen Chen, Gen-Huey Chen, and Jang-Ping Sheu. Designing efficient parallel algorithms on mesh-connected computers with multiple broadcasting. IEEE Transactions on Parallel and Distributed Systems, 1(2):241– 246, April 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chtepen:2009:ATC

[CCD^{+}09] Maria Chtepen, Filip H. A.

Claeys, Bart Dhoedt, Filip De Turck, Piet Demeester, and Peter A. Vanrolleghem. Adaptive task checkpointing and replication: Toward efficient faulttolerant grids. IEEE Transactions on Parallel and Distributed Systems, 20(2):180–190, February 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chuang:2002:BBU

[CCJ02] Po-Jen Chuang, Juei-Tang Chen, and Yue-Tsuen Jiang. Balancing buffer utilization in meshes using a ‘Restricted Area’ concept. IEEE Transactions on Parallel and Distributed Systems, 13(8):814–827, August 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0814.
pdf;   http://www.computer.
org/tpds/td2002/l0814abs.
htm.

Chan:1995:OSF

[CCP95] Bethany M. Y. Chan, Francis Y. L. Chin, and Chung-Keung Poon. Optimal simulation of full binary trees on faulty hypercubes. IEEE Transactions on Parallel and Distributed Systems, 6(3):269–286, March 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.

                                        computer.org/tpds/td1995/
                                        l0269abs.htm.

Caminero:2005:TSS

[CCQ^{+}05] Blanca Caminero, Carmen Carrion, Francisco J. Quiles, Jose Duato, and Sudhakar Yalamanchili. Traffic scheduling solutions with QoS support for an input-buffered MultiMedia router. IEEE Transactions on Parallel and Distributed Systems, 16(11):1009–1021, November 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chakrabarti:2009:ESQ

[CCSC09] Saikat Chakrabarti, Santosh Chandrasekhar, Mukesh Singhal, and Kenneth L. Calvert. An efficient and scalable quasiaggregate signature scheme based on LFSR sequences. IEEE Transactions on Parallel and Distributed Systems, 20(7):1059– 1072, July 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1996:GRA

[CCY96] Ming-Syan Chen, Jeng-Chun Chen, and Philip S. Yu. On general results for all-to-all broadcast. IEEE Transactions on Parallel and Distributed Systems, 7 (4):363–370, April 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0363abs.htm.

Cardellini:2003:RRA

[CCY03] Valeria Cardellini, Michele Colajanni, and Philip S. Yu. Request redirection algorithms for distributed Web systems. IEEE Transactions on Parallel and Distributed Systems, 14(4):355–368, April 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0355abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0355.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0355.
pdf.

Carlson:1994:SPA

[CD94] Brian M. Carlson and Lawrence W.

Dowdy. Static processor allocation in a soft real-time multiprocessor environment. IEEE Transactions on Parallel and Distributed Systems, 5(3):316– 320, March 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2008:ABF

[CD08] Zizhong Chen and Jack Dongarra. Algorithm-based fault tolerance for fail-stop failures. IEEE Transactions on Parallel and Distributed Systems, 19(12): 1628–1641, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Canto:2005:PDP

[CdMB05] Sebastián Dormido Canto, Ángel P. de Madrid, and Sebastián Dormido Bencomo. Parallel dynamic programming on clusters of workstations. IEEE Transactions on Parallel and Distributed Systems, 16(9):785–798, September 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Calland:1998:CRA

[CDR98] P.-Y. Calland, A. Darte, and Y. Robert. Circuit retiming applied to decomposed software pipelining. IEEE Transactions on Parallel and Distributed Systems, 9(1):24–35, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0024.
pdf;   http://www.computer.
org/tpds/td1998/l0024abs.
htm.

Carneiro:2006:TDA

[CDV^{+}06] Aline Carneiro, Viana Marcelo Dias de Amorim, Yannis Viniotis, Serge Fdida, and José Ferreira de Rezende. Twins: a dual addressing space representation for self-organizing networks. IEEE Transactions on Parallel and Distributed Systems, 17(12):1468–1481, December 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cann:1995:AAO

[CE95] David C. Cann and Paraskevas Evripidou. Advanced array optimizations for high performance functional languages. IEEE Transactions on Parallel and Distributed Systems, 6(3):229– 239, March 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0229abs.
htm.

Cosnard:1994:AAP

[CF94] Michel Cosnard and Pierre Fraigniaud. Analysis of asynchronous polynomial root finding methods on a distributed memory multicomputer. IEEE Transactions on Parallel and Distributed Systems, 5(6):639– 648, June 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chandy:1995:NDC

[CF95] K. Mani Chandy and Ian Foster. A notation for deterministic cooperating processes. IEEE Transactions on Parallel and Distributed Systems, 6(8): 863–871, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0863abs.htm.

Cam:1999:WER

[ÇF99a] H. Çam and J. A. B. Fortes.

Work-efficient routing algorithms for rearrangeable symmetrical networks. IEEE Transactions on Parallel and Distributed Systems, 10(7):733–??, July 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0733.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0733abs.
                                        htm.

Cristian:1999:TAD

[CF99b] F. Cristian and C. Fetzer. The timed asynchronous distributed system model. IEEE Transactions on Parallel and Distributed Systems, 10(6):642–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0642.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0642abs.
                                        htm.

Carothers:2000:EET

[CF00] C. D. Carothers and R. M. Fujimoto. Efficient execution of time warp programs on heterogeneous, NOW platforms. IEEE Transactions on Parallel and Distributed Systems, 11(3):299– ??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0299.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0299abs.
                                        htm.

Chen:2001:CAM

[CF01] S.-K. Chen and W. K. Fuchs. Compiler-assisted multiple instruction word retry for VLIW architectures. IEEE Transactions on Parallel and Distributed Systems, 12(12):1293–1304, December 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1293.
pdf;   http://www.computer.
org/tpds/td2001/l1293abs.
htm.

Chand:2008:SDX

[CF08] Raphael Chand and Pascal Felber. Scalable distribution of XML content with XNet. IEEE Transactions on Parallel and Distributed Systems, 19(4):447– 461, April 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chamberlain:2002:GOI

[CFB02] Roger D. Chamberlain, Mark A.

Franklin, and Ch’ng Shi Baw. Gemini: An optical interconnection network for parallel processing. IEEE Transactions on Parallel and Distributed Systems, 13(10):1038–1055, October 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l1038.
pdf;   http://www.computer.
org/tpds/td2002/l1038abs.
htm.

Fu:1998:CCN

[cFC98] A. W c. Fu and S.-C. Chau.

Cyclic-cubes: a new family of interconnection networks of even fixed-degrees. IEEE Transactions on Parallel and Distributed Systems, 9(12):253–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1253.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1253abs.
                                        htm.

Cohen:1998:OBM

[CFKR98] J. Cohen, P. Fraigniaud, J.-C. König, and A. Raspaud. Optimized broadcasting and multicasting protocols in cut-through routed networks. IEEE Transactions on Parallel and Distributed Systems, 9(8):788–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0788.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0788abs.
                                        htm.

Correa:1999:SMT

[CFR99] R. C. Corrêa, A. Ferreira, and P. Rebreyend. Scheduling multiprocessor tasks with genetic algorithms. IEEE Transactions on Parallel and Distributed Systems, 10(8):825–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0825.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0825abs.
                                        htm.

Chen:1998:PGS

[CFW98] H. Chen, N. S. Flann, and D. W. Watson. Parallel genetic simulated annealing: a massively parallel SIMD algorithm. IEEE Transactions on Parallel and Distributed Systems, 9(2):126–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0126.
pdf;   http://www.computer.
org/tpds/td1998/l0126abs.
htm.

Cremonesi:2002:IPMa

[CG02a] Paolo Cremonesi and Claudio Gennaro. Integrated performance models for SPMD applications and MIMD architectures. IEEE Transactions on Parallel and Distributed Systems, 13(7):745–757, July 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0745.
pdf;   http://www.computer.
org/tpds/td2002/l0745abs.

htm. Editor’s Note: This paper unfortunately contains some errors which led to the paper being reprinted in the December 2002 issue. Please see IEEE Transactions on Parallel and Distributed Systems, vol. 13, no. 12, December 2002, pp. 1320–1332 for the correct paper.

Cremonesi:2002:IPMb

[CG02b] Paolo Cremonesi and Claudio Gennaro. Integrated performance models for SPMD applications and MIMD architectures. IEEE Transactions on Parallel and Distributed Systems, 13(12):1320– 1332, December 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2002/12/
                                        l1320.pdf;   http://csdl.
                                        computer.org/comp/trans/td/
                                        2002/12/l1320abs.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2002/12/l1320.htm.

Carroll:2008:SPM

[CG08] Thomas E. Carroll and Daniel Grosu. Strategy-proof mechanisms for scheduling divisible loads in bus-networked distributed systems. IEEE Transactions on Parallel and Distributed Systems, 19(8):1124– 1135, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chandra:2004:POC

[CGK04] Punit Chandra, Pranav Gambhire, and Ajay D. Kshemkalyani. Performance of the optimal causal multicast algorithm: a statistical analysis. IEEE Transactions on Parallel and Distributed Systems, 15(1):40– 52, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2004/01/
                                        l0040abs.htm;   http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/01/l0040.pdf.

Chou:2007:MMT

[CGL07] Cheng-Fu Chou, Leana Golubchik, and John Chi-Shing Lui. Multiclass multiserver threshold-based systems: a study of noninstantaneous server activation. IEEE Transactions on Parallel and Distributed Systems, 18(1):96–110, January 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cooper:2005:PPD

[CGM05] Brian F. Cooper and Hector Garcia-Molina. Peer-to-peer data preservation through storage auctions. IEEE Transactions on Parallel and Distributed Systems, 16(3):246–257, March 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chaparro:2007:UTI

[CGM^{+}07] Pedro Chaparro, José González, Grigorios Magklis, Qiong Cai, and Antonio González. Understanding the thermal implications of multi-core architectures. IEEE Transactions on Parallel and Distributed Systems, 18(8): 1055–1065, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1992:RAD

[CH92] Deng-Jyi Chen and Tien-Hsiang Huang. Reliability analysis of distributed systems based on a fast reliability algorithm. IEEE Transactions on Parallel and Distributed Systems, 3(2):139– 154, March 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chong:1995:PAF

[CH95] Yong-Kim Chong and Kai Hwang. Performance analysis of four memory consistency models for multithreaded multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6 (10):1085–1099, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1085abs.htm.

Chiu:1998:NTO

[CH98] G.-M. Chiu and C.-M. Hsiao. A note on total ordering multicast using propagation trees. IEEE Transactions on Parallel and Distributed Systems, 9(2):217–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0217.
pdf;   http://www.computer.
org/tpds/td1998/l0217abs.
htm.

Chen:2001:SDF

[CH01] H.-L. Chen and S.-H. Hu. Submesh determination in faulty tori and meshes. IEEE Transactions on Parallel and Distributed Systems, 12(3):272–??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0272.
pdf;   http://www.computer.
org/tpds/td2001/l0272abs.
htm.

Chaudhuri:2004:EVN

[CH04a] Mainak Chaudhuri and Mark Heinrich. Exploring virtual network selection algorithms in DSM cache coherence protocols. IEEE Transactions on Parallel and Distributed Systems, 15(8): 699–712, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/08/l0699.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/08/l0699.pdf.

Chaudhuri:2004:INA

[CH04b] Mainak Chaudhuri and Mark Heinrich. The impact of negative acknowledgments in shared memory scientific applications. IEEE Transactions on Parallel and Distributed Systems, 15(2):134–150, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0134abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0134.
pdf.

Chaudhuri:2007:IMC

[CH07] Mainak Chaudhuri and Mark Heinrich. Integrated memory controllers with parallel coherence streams. IEEE Transactions on Parallel and Distributed Systems, 18(8):1159–1173, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2008:ADR

[CH08] Jen-Yeu Chen and Jianghai Hu. Analysis of distributed random grouping for aggregate computation on wireless sensor networks with randomly changing graphs. IEEE Transactions on Parallel and Distributed Systems, 19(8): 1136–1149, August 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2009:BBM

[CH09] Sheng-Hsiung Chen and Ting-Lu Huang. Bounded-bypass mutual exclusion with minimum number of registers. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1726–1737, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Charny:1996:MPV

[Cha96] Benjamin Charny. Matrix partitioning on a virtual shared memory parallel machine. IEEE Transactions on Parallel and Distributed Systems, 7(4):343– 355, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0343abs.
                                        htm.

Cao:2007:UUC

[CHA07] Qing Cao, Tian He, and Tarek Abdelzaher. uCast: Unified connectionless multicast for energy efficient content distribution in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(2):240–250, February 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chung:1998:BCC

[CHB98] Y.-C. Chung, C.-H. Hsu, and S.- W. Bai. A basic-cycle calculation technique for efficient dynamic data redistribution. IEEE Transactions on Parallel and Distributed Systems, 9(4):359– ??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0359.
pdf;   http://www.computer.
org/tpds/td1998/l0359abs.
htm.

Chang:2004:UEL

[CHC04] Weng-Long Chang, Jih-Woei Huang, and Chih-Ping Chu. Using elementary linear algebra to solve data alignment for arrays with linear or quadratic references. IEEE Transactions on Parallel and Distributed Systems, 15(1):28–39, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/01/l0028abs.
htm;   http://csdl.computer.

                                        org/dl/trans/td/2004/01/l0028.
                                        pdf.

Chen:2009:HDR

[CHC09] Ke Chen, Kai Hwang, and Gang Chen. Heuristic discovery of role-based trust chains in peerto-peer networks. IEEE Transactions on Parallel and Distributed Systems, 20(1):83–96, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1995:EGA

[Che95a] Danny Z. Chen. Efficient geometric algorithms on the EREW PRAM. IEEE Transactions on Parallel and Distributed Systems, 6(1):41–47, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1995:EPB

[Che95b] Danny Z. Chen. Efficient parallel binary search on sorted arrays, with applications. IEEE Transactions on Parallel and Distributed Systems, 6(4):440– 445, April 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1995/l0440abs.
                                        htm.

Chen:1996:GII

[Che96] Lin Chen. Graph isomorphism and identification matrices: Parallel algorithms. IEEE Transactions on Parallel and Distributed Systems, 7(3):308–319, March 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0308abs.htm.

Chen:2001:HSF

[Che01] Ray Jinzhu Chen. A hybrid solution of fork/join synchronization in parallel queues. IEEE Transactions on Parallel and Distributed Systems, 12(8):829– 845, August 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0829.
pdf;   http://www.computer.
org/tpds/td2001/l0829abs.
htm.

Chen:2007:CEH

[Che07] Yu-Wei Chen. A comment on “The Exchanged Hypercube”. IEEE Transactions on Parallel and Distributed Systems, 18 (4):576, April 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://bell.computer.
org/dlcomments/. See [LHP05].

Chou:2006:SSL

[CHHC06] Jerry C.-Y. Chou, Tai-Yi Huang, Kuang-Li Huang, and Tsung-Yen Chen. SCALLOP: a scalable and load-balanced peer-to-peer lookup protocol. IEEE Transactions on Parallel and Distributed Systems, 17(5):419–433, May 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chien:1998:CSM

[Chi98] A. A. Chien. A cost and speed model for k-ary n-cube wormhole routers. IEEE Transactions on Parallel and Distributed Systems, 9(2):150–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0150.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0150abs.
                                        htm.

Chiu:2000:OET

[Chi00]  G.-M. Chiu.  The odd-

even turn model for adaptive routing. IEEE Transactions on Parallel and Distributed Systems, 11(7):729–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0729.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0729abs.
                                        htm.

Congy:2007:ASA

[CHJ^{+}07] Jason Congy, Guoling Hany, Ashok Jagannathan, Glenn Reinmany, and Krzysztof Rutkowski. Accelerating sequential applications on CMPs using core spilling. IEEE Transactions on Parallel and Distributed Systems, 18(8):1094–1107, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2004:IPP

[CHJL04] Peng-Sheng Chen, Yuan-Shin Hwang, Roy Dz-Ching Ju, and Jenq Kuen Lee. Interprocedural probabilistic pointer analysis. IEEE Transactions on Parallel and Distributed Systems, 15(10):893–907, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/10/l0893abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0893.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0893.
pdf.

Chen:2007:CDD

[CHK07] Yu Chen, Kai Hwang, and Wei-Shinn Ku. Collaborative detection of DDoS attacks over multiple network domains. IEEE Transactions on Parallel and Distributed Systems, 18(12): 1649–1662, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2009:ENQ

[CHL09] Kuan-Ta Chen, Polly Huang, and Chin-Laung Lei. Effect of network quality on player departure behavior in online games. IEEE Transactions on Parallel and Distributed Systems, 20 (5):593–606, May 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chung:1995:PCG

[Chu95] Kuo-Liang Chung. Prefix computations on a generalized meshconnected computer with multiple buses. IEEE Transactions on Parallel and Distributed Systems, 6(2):196–199, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1995/
                                        l0196abs.htm.

Chuang:1996:CFT

[Chu96] Po-Jen Chuang. CGIN: a fault tolerant modified gamma interconnection network. IEEE Transactions on Parallel and Distributed Systems, 7(12):1301– 1306, December 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:1992:RCM

[CI92] Xian Cheng and Oliver C. Ibe. Reliability of a class of multistage interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 3(2): 241–246, March 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chi:2008:TDN

[CJHG08] Kaikai Chi, Xiaohong Jiang, Susumu Horiguchi, and Minyi Guo. Topology design of network-coding-based multicast networks. IEEE Transactions on Parallel and Distributed Systems, 19(5):627–640, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2009:DAH

[CJLN09] Hanhua Chen, Hai Jin, Yunhao Liu, and Lionel M. Ni. Difficultyaware hybrid search in peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 20(1):71–82, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cohen:2006:MSP

[CJPW06] Johanne Cohen, Emmanuel Jeannot, Nicolas Padoy, and Frederic Wagner. Messages scheduling for parallel data redistribution between clusters. IEEE Transactions on Parallel and Distributed Systems, 17(10): 1163–1175, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chiou:1996:EDE

[CK96] Hsien-Kuang Chiou and Willard Korfhage. Enhancing distributed event predicate detection algorithms. IEEE Transactions on Parallel and Distributed Systems, 7(7):673–676, July 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0673abs.htm.

Chow:2002:LBD

[CK02] Ka-Po Chow and Yu-Kwong Kwok. On load balancing for distributed multiagent computing. IEEE Transactions on Parallel and Distributed Systems, 13(8):787–801, August 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0787.
pdf;   http://www.computer.
org/tpds/td2002/l0787abs.
htm.

Chen:2008:CDC

[CK08] Guilin Chen and Mahmut Kandemir. Compiler-directed code restructuring for improving performance of MPSoCs. IEEE Transactions on Parallel and Distributed Systems, 19(9):1201– 1214, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Choudhury:2008:HSD

[CKC08] Pravanjan Choudhury, Rajeev Kumar, and P. P. Chakrabarti. Hybrid scheduling of dynamic task graphs with selective duplication for multiprocessors under memory and time constraints. IEEE Transactions on Parallel and Distributed Systems, 19 (7):967–980, July 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2004:SET

[CKK^{+}04] Guangyu Chen, Byung-Tae Kang, Mahmut Kandemir, Narayanan Vijaykrishnan, Mary Jane Irwin, and Rajarathnam Chandramouli. Studying energy trade offs in offloading computation/compilation in Javaenabled mobile devices. IEEE Transactions on Parallel and Distributed Systems, 15(9):795– 809, September 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/09/l0795.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/09/l0795.pdf.

Chiu:2008:BCR

[CKWC08] Chun-Yuan Chiu, Yu-Liang Kuo, Eric Hsiao-Kuang Wu, and Gen-Huey Chen. Bandwidthconstrained routing problem in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 19(1):4–14, January 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chan:1993:FTE

[CL93] M. Y. Chan and S.-J. Lee. Faulttolerant embedding of complete binary trees in hypercubes. IEEE Transactions on Parallel and Distributed Systems, 4 (3):277–288, March 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1994:PEA

[CL94] Chung-Ming Chen and Soo-Young Lee. On parallelizing the EM algorithm for PET image reconstruction. IEEE Transactions on Parallel and Distributed Systems, 5(8):860–873, August 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1997:CNF

[CL97] Guihai Chen and F. C. M. Lau. Comments on “A New Family of Cayley Graph Interconnection Networks of Constant Degree Four”. IEEE Transactions on Parallel and Distributed Systems, 8(12):1299–1300, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1299.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1299abs.

htm. See [VS96].

Chen:2000:TLC

[CL00] G. Chen and F. C. M. Lau. Tighter layouts of the cubeconnected cycles. IEEE Transactions on Parallel and Distributed Systems, 11(2):182–??, February 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0182.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0182abs.
                                        htm.

Cintra:2005:DSE

[CL05] Marcelo Cintra and Diego R. Llanos. Design space exploration of a software speculative parallelization scheme. IEEE Transactions on Parallel and Distributed Systems, 16(6):562–576, June 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2009:PPM

[CL09] Keke Chen and Ling Liu.

Privacy-preserving multiparty collaborative mining with geometric data perturbation. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1764–1776, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Carra:2008:SGP

[CLB08] Damiano Carra, Renato Lo Cigno, and Ernst W. Biersack. Stochastic graph processes for performance evaluation of content delivery applications in overlay networks. IEEE Transactions on Parallel and Distributed Systems, 19(2):247–261, February 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chandra:2004:GST

[CLJ^{+}04] Ramesh Chandra, Ryan M.

Lefever, Kaustubh R. Joshi, Michel Cukier, and William H. Sanders. A global-state-triggered fault injector for distributed system evaluation. IEEE Transactions on Parallel and Distributed Systems, 15 (7):593–605, July 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/07/l0593.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/07/l0593.pdf;   http:
//csdl2.computer.org/comp/

                                        trans/td/2004/07/l0593supp.
                                        pdf.

Chatterjee:2002:RAL

[CLPT02] Siddhartha Chatterjee, Alvin R. Lebeck, Praveen K. Patnala, and Mithuna Thottethodi. Recursive array layouts and fast matrix multiplication. IEEE Transactions on Parallel and Distributed Systems, 13(11): 1105–1123, November 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2002/11/l1105abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1105.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1105.
                                        pdf.

Claesson:2004:ETS

[CLS04] Vilgot Claesson, Henrik Lönn, and Neeraj Suri. An efficient TDMA start-up and restart synchronization approach for distributed embedded systems. IEEE Transactions on Parallel and Distributed Systems, 15 (8):725–739, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/08/l0725.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/08/l0725.pdf.

Cantin:2005:CVM

[CLS05] Jason F. Cantin, Mikko H. Lipasti, and James E. Smith. The complexity of verifying memory coherence and consistency. IEEE Transactions on Parallel and Distributed Systems, 16 (7):663–671, July 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1991:PEM

[ClW91] Albert C. Chen and Chuan lin Wu. A parallel execution model of logic programs. IEEE Transactions on Parallel and Distributed Systems, 2(1):79–92, January 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cai:2003:DIA

[CLW03] Zhijun Cai, Mi Lu, and Xiaodong Wang. Distributed initialization algorithms for single-hop ad hoc networks with minislotted carrier sensing. IEEE Transactions on Parallel and Distributed Systems, 14(5):516–528, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/05/l0516abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0516.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0516.
pdf.

Chen:2008:CSP

[CLY08a] Gang Chen, Chor Ping Low, and Zhonghua Yang. Coordinated services provision in peer-to-peer environments. IEEE Transactions on Parallel and Distributed Systems, 19(4):433–446, April 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2008:ESP

[CLY08b] Gang Chen, Chor Ping Low, and Zhonghua Yang. Enhancing search performance in unstructured P2P networks based on users’ common interest. IEEE Transactions on Parallel and Distributed Systems, 19(6):821– 836, June 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1995:GMP

[CM95] Xian Chen and Graham M. Megson. A general methodology of partitioning and mapping for given regular arrays. IEEE Transactions on Parallel and Distributed Systems, 6 (10):1100–1107, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1100abs.htm.

Curtis-Maury:2008:PBP

[CMBAN08] Matthew Curtis-Maury, Filip Blagojevic, Christos D. Antonopoulos, and Dimitrios S. Nikolopoulos. Prediction-based powerperformance adaptation of multithreaded scientific codes. IEEE Transactions on Parallel and Distributed Systems, 19(10): 1396–1410, October 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Coll:2009:ESH

[CMDP09] Salvador Coll, Francisco J.

Mora, Jose Duato, and Fabrizio Petrini. Efficient and scalable hardware-based multicast in fattree networks. IEEE Transactions on Parallel and Distributed Systems, 20(9):1285– 1298, September 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Corbalan:2005:PDP

[CML05] Julita Corbalan, Xavier Martorell, and Jesus Labarta. Performance-driven processor allocation. IEEE Transactions on Parallel and Distributed Systems, 16(7):599–611, July 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cordasco:2007:AIS

[CMR07] Gennaro Cordasco, Grzegorz Malewicz, and Arnold Rosenberg. Advances in IC-scheduling theory: Scheduling expansive and reductive dags and scheduling dags via duality. IEEE Transactions on Parallel and Distributed Systems, 18(11): 1607–1617, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Comino:2002:NDD

[CN02] Nicholas Comino and V. Lakshmi Narasimhan. A novel data distribution technique for host-client type parallel applications. IEEE Transactions on Parallel and Distributed Systems, 13(2):97–110, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0097.
pdf;   http://www.computer.
org/tpds/td2001/l0097abs.

htm. See comment [Rob04] and response [CN04].

Comino:2004:RCN

[CN04] Nick Comino and V. Lakshmi Narasimhan. Response to “Comment on ‘A Novel Data Distribution Technique for Host-Client Type Parallel Applications” ’. IEEE Transactions on Parallel and Distributed Systems, 15(6):576–??, June 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0576.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0576.pdf. See

[CN02, Rob04].

Choudhary:1994:OPA

[CNNS94] Alok N. Choudhary, Bhagirath Narahari, David M. Nicol, and Rahul Simha. Optimal processor assignment for a class of pipelined computations. IEEE Transactions on Parallel and Distributed Systems, 5(4):439– 445, April 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Curescu:2005:TAU

[CNT05] Calin Curescu and Simin Nadjm-Tehrani. Time-aware utilitybased resource allocation in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 16(7):624–636, July 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chien:1994:ABS

[CO94] Minze V. Chien and A. Yavuz Oruc. Adaptive binary sorting schemes and associated interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 5(6):561–572, June 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:1995:DTA

[CO95] Hua Wu David Chang and William J. B. Oldham. Dynamic task allocation models for large distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 6(12): 1301–1315, December 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Calamoneri:2000:SPA

[COP00] T. Calamoneri, S. Olariu, and R. Petreschi. A simple parallel algorithm to draw cubic graphs. IEEE Transactions on Parallel and Distributed Systems, 11(10):9–??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1009.
pdf;   http://www.computer.
org/tpds/td2000/l1009abs.
htm.

Corbett:1992:RGE

[Cor92]  Peter F. Corbett.  Rotator

graphs: An efficient topology for point-to-point multiprocessor networks. IEEE Transactions on Parallel and Distributed Systems, 3(5):622–626, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2000:OOL

[COS00] F. Chen, T. W. O’Neil, and E. H.-M. Sha. Optimizing overall loop schedules using prefetching and partitioning. IEEE Transactions on Parallel and Distributed Systems, 11(6):604– ??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0604.
pdf;   http://www.computer.
org/tpds/td2000/l0604abs.
htm.

Choudhary:1993:NHP

[CPA93] Alok N. Choudhary, Janak H. Patel, and Narendra Ahuja. NE-TRA: a hierarchical and partitionable architecture for computer vision systems. IEEE Transactions on Parallel and Distributed Systems, 4(10):1092– 1104, October 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2004:SEP

[CPhX04] Ling Chen, Yi Pan, and Xiao hua Xu. Scalable and efficient parallel algorithms for Euclidean distance transform on the LARPBS model. IEEE Transactions on Parallel and Distributed Systems, 15(11):975–982, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/11/l0975abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l0975.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l0975.
pdf.

Chen:2007:DCS

[CPM07] Ruiliang Chen, Jung-Min Park, and Randolph Marchany. A divide-and-conquer strategy for thwarting distributed denial-of-service attacks. IEEE Transactions on Parallel and Distributed Systems, 18(5):577–588, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2006:RCA

[CPX06] Jen-Yeu Chen, Gopal Pandurangan, and Dongyan Xu. Robust computation of aggregates in wireless sensor networks: Distributed randomized algorithms and analysis. IEEE Transactions on Parallel and Distributed Systems, 17(9):987–1000, September 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cunningham:1990:USP

[CR90] H. Conrad Cunningham and Gruia-Catalin Roman. UNITYstyle programming logic for shared dataspace programs. IEEE Transactions on Parallel and Distributed Systems, 1 (3):365–376, July 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:1994:SAM

[CR94] Rong N. Chang and Chinya V. Ravishankar. A service acquisition mechanism for server-based heterogeneous distributed systems. IEEE Transactions on Parallel and Distributed Systems, 5(2):154–169, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cho:2009:GCS

[CRN09] Hyeonjoong Cho, Binoy Ravindran, and Chewoo Na. Garbage collector scheduling in dynamic, multiprocessor real-time systems. IEEE Transactions on Parallel and Distributed Systems, 20(6):845–856, June 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Casanova:2006:GES

[CRS06] Henri Casanova, Yves Robert, and H. J. Siegel. Guest editorial: Special section on algorithm design and scheduling techniques (realistic platform models) for heterogeneous clusters. IEEE Transactions on Parallel and Distributed Systems, 17 (2):97–98, February 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/02/
l0097.pdf.

Chen:1990:DFS

[CS90] Ming-Syan Chen and Kang G.

Shin. Depth-first search approach for fault-tolerant routing in hypercube multicomputers. IEEE Transactions on Parallel and Distributed Systems, 1 (2):152–159, April 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Corbett:1992:SMC

[CS92] Peter F. Corbett and Isaac D.

Scherson. Sorting in mesh connected multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 3(5):626– 632, September 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1994:CFD

[CS94] Tzung-Shi Chen and Jang-Ping Sheu. Communication-free data allocation techniques for parallelizing compilers on multicomputers. IEEE Transactions on Parallel and Distributed Systems, 5(9):924–938, September 1994. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheung:1995:PBS

[CS95] Shun Yan Cheung and Vaidy S. Sunderam. Performance of barrier synchronization methods in a multiaccess network. IEEE Transactions on Parallel and Distributed Systems, 6(8): 890–895, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0890abs.htm.

Choy:1996:LFD

[CS96] Manhoi Choy and Ambuj K. Singh. Localizing failures in distributed synchronization. IEEE Transactions on Parallel and Distributed Systems, 7 (7):705–716, July 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0705abs.htm.

Chao:1997:SDF

[CS97a] Liang-Fang Chao and E. Hsing-Mean Sha. Scheduling dataflow graphs via retiming and unfolding. IEEE Transactions on Parallel and Distributed Systems, 8(12):1259–1267, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1259.
                                        pdf;   http://www.computer.

org/tpds/td1997/l1259abs.
htm.

Chou:1997:SRT

[CS97b] Chih-Che Chou and K. G. Shin.

Statistical real-time channels on multiaccess bus networks. IEEE Transactions on Parallel and Distributed Systems, 8(8):769– 780, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0769.
pdf;   http://www.computer.
org/tpds/td1997/l0769abs.
htm.

Cao:1998:CCD

[CS98] G. Cao and M. Singhal. On coordinated checkpointing in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 9(12):213–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1213.
pdf;   http://www.computer.
org/tpds/td1998/l1213abs.
htm.

Cao:2001:DOQ

[CS01a] G. Cao and M. Singhal. A delayoptimal quorum-based mutual exclusion algorithm for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 12(12):1256– 1268, December 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.

                                        org/td/books/td2001/pdf/l1256.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1256abs.
                                        htm.

Cao:2001:MCN

[CS01b] G. Cao and M. Singhal. Mutable checkpoints: a new checkpointing approach for mobile computing systems. IEEE Transactions on Parallel and Distributed Systems, 12(2):157–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0157.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0157abs.

htm. See correction [CS02a].

Cao:2002:CMC

[CS02a] G. Cao and M. Singhal. Correction to “Mutable Checkpoints: A New Checkpointing Approach for Mobile Computing Systems”. IEEE Transactions on Parallel and Distributed Systems, 13(5):527, May 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td2002/l0527.pdf.

See [CS01b].

Choi:2002:ABR

[CS02b] Sunghyun Choi and Kang G. Shin. Adaptive bandwidth reservation and admission control in QoS-sensitive cellular networks. IEEE Transactions on Parallel and Distributed Systems, 13(9):882–897, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0882.
pdf;   http://www.computer.
org/tpds/td2002/l0882abs.
htm.

Corsaro:2003:DPR

[CS03] Angelo Corsaro and Douglas C.

Schmidt. The design and performance of real-time Java middleware. IEEE Transactions on Parallel and Distributed Systems, 14(11):1155–1167, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/11/l1155abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/11/l1155.
pdf.

Chen:2005:PBD

[CS05] Shigang Chen and Qingguo Song. Perimeter-based defense against high bandwidth DDoS attacks. IEEE Transactions on Parallel and Distributed Systems, 16(6):526–537, June 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chandra:2008:HSS

[CS08] Abhishek Chandra and Prashant Shenoy. Hierarchical scheduling for symmetric multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 19 (3):418–431, March 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/;
                                         http://csdl.computer.org/
                                        comp/trans/td/2008/03/ttd2008030418s.
                                        pdf.

Chen:2007:ASC

[CSC07] Shiping Chen, Baile Shi, and Shigang Chen. ACOM: Anysource capacity-constrained overlay multicast in non-DHT P2P networks. IEEE Transactions on Parallel and Distributed Systems, 18(9):1188–1201, September 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2000:ECT

[CSH00] C.-P. Chang, T.-Y. Sung, and L.-H. Hsu. Edge congestion and topological properties of crossed cubes. IEEE Transactions on Parallel and Distributed Systems, 11(1):64–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0064.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0064abs.
                                        htm.

Cordasco:2007:BCM

[CSR07] Gennaro Cordasco, Vittorio Scarano, and Arnold L. Rosenberg. Bounded-collision memorymapping schemes for data structures with applications to parallel memories. IEEE Transactions on Parallel and Distributed Systems, 18(7):973–982, July 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chervenak:2009:GRL

[CSR^{+}09] Ann L. Chervenak, Robert Schuler, Matei Ripeanu, Muhammad Ali Amer, Shishir Bharathi, Ian Foster, Adriana Iamnitchi, and Carl Kesselman. The Globus Replica Location Service: Design and experience. IEEE Transactions on Parallel and Distributed Systems, 20 (9):1260–1272, September 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2002:PBR

[CST02] Shao Dong Chen, Hong Shen, and Rodney Topor. Permutationbased range-join algorithms on N-dimensional meshes. IEEE Transactions on Parallel and Distributed Systems, 13(4):413– 431, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0413.
pdf;   http://www.computer.
org/tpds/td2002/l0413abs.
htm.

Chuang:1994:APS

[CT94] Po-Jen Chuang and Nian-Feng Tzeng. Allocating precise submeshes in mesh connected systems. IEEE Transactions on Parallel and Distributed Systems, 5(2):211–217, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1997:BEB

[CT97] Hsing-Lung Chen and Nian-Feng Tzeng. A Boolean expression-based approach for maximum incomplete subcube identification in faulty hypercubes. IEEE Transactions on Parallel and Distributed Systems, 8(11):1171–1183, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1171.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1171abs.
                                        htm.

Chen:2002:MPE

[CT02] Jian Chen and Valerie E. Taylor. Mesh partitioning for efficient use of distributed systems. IEEE Transactions on Parallel and Distributed Systems, 13(1):67– 79, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0067.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0067abs.
                                        htm.

Chen:2007:DDA

[CT07] Shigang Chen and Yong Tang. DAW: a distributed antiworm system. IEEE Transactions on Parallel and Distributed Systems, 18(7):893–906, July 2007.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2008:SRP

[CT08] Chyouhwa Chen and Kun-Cheng Tsai. The server reassignment problem for load balancing in structured P2P systems. IEEE Transactions on Parallel and Distributed Systems, 19(2): 234–246, February 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1993:DAP

[CTC93] Yigang Chen, W. T. Tsai, and Daniel Chao. Dependency analysis — a Petri net based technique for synthesizing large concurrent systems. IEEE Transactions on Parallel and Distributed Systems, 4(4):414–426, April 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2009:ITP

[CTF09] Hongbin Chen, Chi K. Tse, and Jiuchao Feng. Impact of topology on performance and energy efficiency in wireless sensor networks for source extraction. IEEE Transactions on Parallel and Distributed Systems, 20 (6):886–897, June 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1996:BST

[CTS96] Tzung-Shi Chen, Yu-Chee Tseng, and Jang-Ping Sheu. Balanced spanning trees in complete and incomplete star graphs. IEEE Transactions on Parallel and Distributed Systems, 7 (7):717–723, July 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0717abs.htm.

Chalasani:1992:ETT

[CV92] S. Chalasani and A. Varma. Evaluation of two traffic distribution strategies for a dualnetwork multiprocessor system. IEEE Transactions on Parallel and Distributed Systems, 3 (3):375–384, May 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chiang:2008:DPP

[CV08] Su-Hui Chiang and Sangsuree Vasupongayya. Design and potential performance of goaloriented job scheduling policies for parallel computer workloads. IEEE Transactions on Parallel and Distributed Systems, 19(12): 1642–1656, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2000:ERC

[CW00] P.-J. Chen and C.-M. Wu. An efficient recognition-complete processor allocation strategy for k-ary n-cube multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 11 (5):485–??, May 2000. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2000/pdf/l0485.
pdf;   http://www.computer.
org/tpds/td2000/l0485abs.
htm.

Chai:2002:SOM

[CW02a] Sek M. Chai and Scott Wills.

Systolic opportunities for multidimensional data streams. IEEE Transactions on Parallel and Distributed Systems, 13(4):388– 398, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0388.
pdf;   http://www.computer.
org/tpds/td2002/l0388abs.
htm.

Chen:2002:CUE

[CW02b] Wei Chen and Koichi Wada. On computing the upper envelope of segments in parallel. IEEE Transactions on Parallel and Distributed Systems, 13(1):5–13, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0005.
pdf;   http://www.computer.
org/tpds/td2001/l0005abs.
htm.

Cai:2006:EGT

[CW06] Hailong Cai and Jun Wang. Exploiting geographical and temporal locality to boost search efficiency in peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 17(10): 1189–1203, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2007:IMD

[CWCC07] Yeim-Kuan Chang, Jia-Hwa Wu, Chi-Yeh Chen, and Chih-Ping Chu. Improved methods for divisible load distribution on kdimensional meshes using multiinstallment. IEEE Transactions on Parallel and Distributed Systems, 18(11):1618–1629, November 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1992:PCA

[CWL92] Gen-Huey Chen, Biing-Feng Wang, and Chi-Jen Lu. On the parallel computation of the algebraic path problem. IEEE Transactions on Parallel and Distributed Systems, 3(2):251– 256, March 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Caverlee:2009:PAS

[CWLR09] James Caverlee, Steve Webb, Ling Liu, and William B. Rouse. A parameterized approach to spam-resilient link analysis of the Web. IEEE Transactions on Parallel and Distributed Systems, 20(10):1422–1438, October 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2009:OOH

[CWYZ09] Jing Chen, Jianping Wang, Hui Yu, and Si-Qing Zheng. Opportunistic optical hyperchannel and its distributed QoS assuring access control. IEEE Transactions on Parallel and Distributed Systems, 20(11):1626– 1640, November 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2006:ORA

[CXN06] Yi Cui, Yuan Xue, and Klara Nahrstedt. Optimal resource allocation in overlay multicast. IEEE Transactions on Parallel and Distributed Systems, 17(8): 808–823, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See comments [YMP08].

Chen:2009:SON

[CXP09] Weidong Chen, Wenjun Xiao, and Behrooz Parhami. Swapped (OTIS) networks built of connected basis networks are maximally fault tolerant. IEEE Transactions on Parallel and Distributed Systems, 20(3):361– 366, March 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1992:IJS

[CY92] Ming-Syan Chen and Philip S.

Yu. Interleaving a join sequence with semijoins in distributed query processing. IEEE Transactions on Parallel and Distributed Systems, 3(5):611–621, September 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:1995:PCT

[CY95] Her-Kun Chang and Shyan-Ming Yuan. Performance characterization of the tree quorum algorithm. IEEE Transactions on Parallel and Distributed Systems, 6(6):658–662, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0658abs.htm.

Chen:1996:EEN

[CY96a] Ding-Kai Chen and Pen-Chung Yew. On effective execution of nonuniform DOACROSS loops. IEEE Transactions on Parallel and Distributed Systems, 7 (5):463–476, May 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0463abs.htm.

Chiu:1996:ERR

[CY96b] Ge-Ming Chiu and Cheng-Ru Young. Efficient rollbackrecovery technique in distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 7(6):565– 577, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0565abs.
                                        htm.

Chung:1996:PDJ

[CY96c] Soon M. Chung and Jaerheen Yang. A parallel distributive join algorithm for cube-connected multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 7(2):127–137, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0127abs.htm.

Chen:1999:RSE

[CY99] D.-K. Chen and P.-C. Yew. Redundant synchronization elimination for DOACROSS loops. IEEE Transactions on Parallel and Distributed Systems, 10 (5):459–??, May 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l0459.
pdf;   http://www.computer.
org/tpds/td1999/l0459abs.
htm.

Choi:2000:CAC

[CY00a] L. Choi and P.-C. Yew. Compiler analysis for cache coherence: Interprocedural array data-flow analysis and its impact on cache performance. IEEE Transactions on Parallel and Distributed Systems, 11(9):879–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0879.
pdf;   http://www.computer.
org/tpds/td2000/l0879abs.
htm.

Choi:2000:HCD

[CY00b] L. Choi and P.-C. Yew. Hardware and compiler-directed cache coherence in large-scale multiprocessors: Design considerations and performance study. IEEE Transactions on Parallel and Distributed Systems, 11 (4):375–??, April 2000. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2000/pdf/l0375.
pdf;   http://www.computer.
org/tpds/td2000/l0375abs.
htm.

Chen:2006:CAB

[CY06] Shigang Chen and Na Yang. Congestion avoidance based on lightweight buffer management in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 17(9):934–946, September 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Colajanni:1998:ATA

[CYD98] M. Colajanni, P. S. Yu, and D. M. Dias. Analysis of task assignment policies in scalable distributed Webserver systems. IEEE Transactions on Parallel and Distributed Systems, 9(6):585–??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0585.
pdf;   http://www.computer.
org/tpds/td1998/l0585abs.
htm.

Chen:1994:ONA

[CYW94] Ming-Syan Chen, Philip S. Yu, and Kun-Lung Wu. Optimal NODUP all-to-all broadcast schemes in distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 5(12):1275–1285, December 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cardei:2008:AFT

[CYW08] Mihaela Cardei, Shuhui Yang, and Jie Wu. Algorithms for fault-tolerant topology in heterogeneous wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(4):545–558, April 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Choo:2000:PSA

[CYY00] H. Choo, S.-M. Yoo, and H. Y.

Youn. Processor scheduling and allocation for 3D torus multicomputer systems. IEEE Transactions on Parallel and Distributed Systems, 11(5):475– ??, May 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0475.
pdf;   http://www.computer.
org/tpds/td2000/l0475abs.
htm.

Chu:2009:ABD

[CZLM09] Xiaowen Chu, Kaiyong Zhao, Zongpeng Li, and Anirban Mahanti. Auction-based ondemand P2P min-cost media streaming with network coding. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1816–1829, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dally:1993:DFA

[DA93] William J. Dally and Hiromichi Aoki. Deadlock-free adaptive routing in multicomputer networks using virtual channels. IEEE Transactions on Parallel and Distributed Systems, 4 (4):466–475, April 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Darbha:1998:OSA

[DA98] S. Darbha and D. P. Agrawal. Optimal scheduling algorithm for distributed-memory machines. IEEE Transactions on Parallel and Distributed Systems, 9(1):87–95, January 1998. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://dlib.
computer.org/td/books/td1998/
pdf/l0087.pdf;   http://www.
computer.org/tpds/td1998/
l0087abs.htm.

Day:1997:FDA

[DAA97a] K. Day and A. E. Al-Ayyoub. Fault diameter of k-ary ncube networks. IEEE Transactions on Parallel and Distributed Systems, 8(9):903–907, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0903.
pdf;   http://www.computer.
org/tpds/td1997/l0903abs.
htm.

Day:1997:CPI

[DAA97b] Khaled Day and Abdel-Elah Al-Ayyoub. The cross product of interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 8(2):109–118, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0109.
pdf;   http://www.computer.
org/tpds/td1997/l0109abs.
htm.

Day:2000:MFD

[DAA00] K. Day and A.-E. Al-Ayyoub.

Minimal fault diameter for highly resilient product networks. IEEE Transactions on Parallel and Distributed Systems, 11(9):926–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0926.
pdf;   http://www.computer.
org/tpds/td2000/l0926abs.
htm.

Day:2002:TPO

[DAA02] Khaled Day and Abdel-Elah Al-Ayyoub. Topological properties of OTIS-networks. IEEE Transactions on Parallel and Distributed Systems, 13(4):359– 366, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0359.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0359abs.
                                        htm.

Dubey:1995:EPT

[DAF95] Pradeep K. Dubey, George B. Adams III, and Michael J. Flynn. Evaluating performance tradeoffs between fine-grained and coarse-grained alternatives. IEEE Transactions on Parallel and Distributed Systems, 6 (1):17–27, January 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dahlin:2000:ISL

[Dah00] M. Dahlin. Interpreting stale load information. IEEE Transactions on Parallel and Distributed Systems, 11(10):33–??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l1033.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l1033abs.
                                        htm.

Dally:1992:VCF

[Dal92]  William J. Dally.  Virtual-

channel flow control. IEEE Transactions on Parallel and Distributed Systems, 3(2):194– 205, March 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Demirbas:2006:FLS

[DAMK06] Murat Demirbas, Anish Arora, Vineet Mittal, and Vinodkrishnan Kulathumani. A fault-local self-stabilizing clustering service for wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 17(9): 912–922, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

deAzevedo:1998:MIC

[dB98] M. M. de Azevedo and D. M.

Blough. Multistep interactive convergence: An efficient approach to the fault-tolerant clock synchronization of large multicomputers. IEEE Transactions on Parallel and Distributed Systems, 9(12):195–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1195.
pdf;   http://www.computer.
org/tpds/td1998/l1195abs.
htm.

DiFatta:2006:DLB

[DB06] Giuseppe Di Fatta and Michael R.

Berthold. Dynamic load balancing for the distributed mining of molecular structures. IEEE Transactions on Parallel and Distributed Systems, 17(8):773– 785, August 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Derhab:2008:SSL

[DB08] Abdelouahid Derhab and Nadjib Badache. A self-stabilizing leader election algorithm in highly dynamic ad hoc mobile networks. IEEE Transactions on Parallel and Distributed Systems, 19(7):926–939, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

deAzevedo:1998:LEP

[dBL98] M. M. de Azevedo, N. Bagherzadeh, and S. Latifi. Low expansion packings and embeddings of hypercubes into star graphs: a performance-oriented approach. IEEE Transactions on Parallel and Distributed Systems, 9(3):261–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0261.
pdf;   http://www.computer.
org/tpds/td1998/l0261abs.
htm.

Dandamudi:1995:HTQ

[DC95] Sivarama P. Dandamudi and Philip S. P. Cheng. Hierarchical task queue organization for shared-memory multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 6(1):1–16, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:1998:FDB

[DC98] S. Das(Bit) and A. Chaudhuri. Fault diagnosis in a Benes interconnection network. IEEE Transactions on Parallel and Distributed Systems, 9(7):700– ??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0700.
pdf;   http://www.computer.
org/tpds/td1998/l0700abs.
htm.

Duh:1995:APN

[DCF95] Dyi-Rong Duh, Gen-Huey Chen, and Jywe-Fei Fang. Algorithms and properties of a new two-level network with folded hypercubes as basic modules. IEEE Transactions on Parallel and Distributed Systems, 6 (7):714–723, July 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0714abs.htm.

Dusseau:1996:FPS

[DCSM96] Andrea C. Dusseau, David E.

Culler, Klaus Erik Schauser, and Richard P. Martin. Fast parallel sorting under LogP: Experience with the CM-5. IEEE Transactions on Parallel and Distributed Systems, 7(8):791–805, August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0791abs.htm.

deCerio:2002:HAM

[dCVGG02] Luis Díaz de Cerio, Miguel Valero-García, and Antonio González. Hypercube algorithms on mesh connected multicomputers. IEEE Transactions on Parallel and Distributed Systems, 13(12):1247–1260, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1247abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1247.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1247.
pdf.

Dimopoulos:1995:OSP

[DD95] Nikitas J. Dimopoulos and Vassilios V. Dimakopoulos. Optimal and suboptimal processor allocation for hypercycle-based multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6(2):175–185, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0175abs.htm.

Dimakopoulos:1998:TTE

[DD98] V. V. Dimakopoulos and N. J. Dimopoulos. A theory for total exchange in multidimensional interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 9(7):639– ??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0639.
pdf;   http://www.computer.
org/tpds/td1998/l0639abs.
htm.

Dimakopoulos:2001:OTE

[DD01] Vassilios V. Dimakopoulos and Nikitas J. Dimopoulos. Optimal total exchange in Cayley graphs. IEEE Transactions on Parallel and Distributed Systems, 12(11):1162–1168, November 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1162.
pdf;   http://www.computer.
org/tpds/td2001/l1162abs.
htm.

DiBlas:2005:UKP

[DDD^{+}05] Andrea Di Blas, David M.

Dahle, Mark Diekhans, Leslie Grate, Jeffrey Hirschberg, Kevin Karplus, Hansjörg Keller, Mark Kendrick, Francisco J. Mesa-Martinez, David Pease, Eric Rice, Angela Schultz, Don Speck, and Richard Hughey. The UCSC Kestrel parallel processor. IEEE Transactions on Parallel and Distributed Systems, 16(1):80–92, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0080abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0080.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0080.
pdf.

Desprez:1998:SBC

[DDP^{+}98] F. Desprez, J. Dongarra, A. Petitet, C. Randriamaro, and Y. Robert. Scheduling block-cyclic array redistribution. IEEE Transactions on Parallel and Distributed Systems, 9(2):192–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0192.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0192abs.
                                        htm.

Dahlgren:1995:SHP

[DDS95] Fredrik Dahlgren, Michel Dubois, and Per Stenstrom. Sequential hardware prefetching in shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6 (7):733–746, July 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0733abs.htm.

Dutta:2007:REP

[DDV^{+}07] Kaushik Dutta, Anindya Datta, Debra VanderMeer, Helen Thomas, and Krithi Ramamritham. ReDAL: An efficient and practical request distribution technique for application server clusters. IEEE Transactions on Parallel and Distributed Systems, 18(11): 1516–1528, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dao:1999:DCM

[DDY99] B. V. Dao, J. Duato, and S. Yalamanchili. Dynamically configurable message flow control for fault-tolerant routing. IEEE Transactions on Parallel and Distributed Systems, 10(1):7–??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0007.
pdf;   http://www.computer.
org/tpds/td1999/l0007abs.
htm.

Deb:1996:MNT

[Deb96]  Ashoke Deb.  Multiskewing

— a novel technique for optimal parallel memory access. IEEE Transactions on Parallel and Distributed Systems, 7 (6):595–604, June 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0595abs.htm.

Das:1997:EEP

[DF97] Samir R. Das and Richard M.

Fujimoto. An empirical evaluation of performance-memory trade-offs in Time Warp. IEEE Transactions on Parallel and Distributed Systems, 8(2):210– 224, February 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l0210.
pdf;   http://www.computer.
org/tpds/td1997/l0210abs.
htm.

Dutta:1999:CFP

[DF99] S. Dutta and M. Franklin. Control flow prediction schemes for wide-issue superscalar processors. IEEE Transactions on Parallel and Distributed Systems, 10(4):346–??, April 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td1999/pdf/l0346.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0346abs.
                                        htm.

Dan:2009:DAS

[DF09] Gyorgy Dan and Viktoria Fodor. Delay asymptotics and scalability for peer-to-peer live streaming. IEEE Transactions on Parallel and Distributed Systems, 20 (10):1499–1511, October 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dowling:1993:HOA

[DFD93] Eric M. Dowling, Zuqiang Fu, and Ron S. Drafz. HARP: An open architecture for parallel matrix and signal processing. IEEE Transactions on Parallel and Distributed Systems, 4(10): 1081–1091, October 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Demaine:2001:GCM

[DFKS01] E. D. Demaine, I. Foster, C. Kesselman, and M. Snir. Generalized communicators in the message passing interface. IEEE Transactions on Parallel and Distributed Systems, 12(6):610– ??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0610.
pdf;   http://www.computer.
org/tpds/td2001/l0610abs.
htm.

Davoli:1996:PCN

[DGB^{+}96] Renzo Davoli, Luigi-Alberto Gianchini, Ozalp Babaoğlu, Alessandro Amoroso, and Lorenzo Alvisi. Parallel computing in networks of workstations with Paralex. IEEE Transactions on Parallel and Distributed Systems, 7(4):371–384, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0371abs.htm.

Delporte-Gallet:2003:ESG

[DGFHR03] Carole Delporte-Gallet, Hugues Fauconnier, Jean-Michel Hélary, and Michel Raynal. Early stopping in global data computation. IEEE Transactions on Parallel and Distributed Systems, 14(9):909–921, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0909abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0909.
pdf.

DHollander:1992:PLL

[D’H92] Erik H. D’Hollander. Partitioning and labeling of loops by unimodular transformations. IEEE Transactions on Parallel and Distributed Systems, 3 (4):465–476, July 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:2001:PPA

[DHB01] S. K. Das, D. J. Harvey, and R. Biswas. Parallel processing of adaptive meshes with load balancing. IEEE Transactions on Parallel and Distributed Systems, 12(12):1269–1280, December 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l1269.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1269abs.
                                        htm.

Debbage:1995:IRP

[DHN95] Mark Debbage, Mark B. Hill, and Denis A. Nicole. An interface to a reliable packet delivery service for parallel systems. IEEE Transactions on Parallel and Distributed Systems, 6 (4):400–411, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0400abs.htm.

Dickens:1996:PDE

[DHN96] Phillip M. Dickens, Philip Heidelberger, and David M. Nicol. Parallelized direct execution simulation of message-passing parallel programs. IEEE Transactions on Parallel and Distributed Systems, 7(10):1090– 1105, October 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l1090abs.htm.

Dhakal:2007:DLB

[DHP^{+}07] Sagar Dhakal, Majeed M. Hayat, Jorge E. Pezoa, Cundong Yang, and David A. Bader. Dynamic load balancing in distributed systems in the presence of delays: a regeneration-theory approach. IEEE Transactions on Parallel and Distributed Systems, 18(4):485–497, April 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dimpsey:1995:MBM

[DI95] Robert T. Dimpsey and Ravishankar K. Iyer. Measurementbased model to predict the performance impact of system modifications: a case study. IEEE Transactions on Parallel and Distributed Systems, 6(1):28–40, January 1995. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dolev:1997:UDS

[DIM97] Shlomi Dolev, Amos Israeli, and Shlomo Moran. Uniform dynamic self-stabilizing leader election. IEEE Transactions on Parallel and Distributed Systems, 8(4):424–440, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0424.
pdf;   http://www.computer.
org/tpds/td1997/l0424abs.
htm.

Ding:2001:OIR

[Din01] C. H. Q. Ding. An optimal index reshuffle algorithm for multidimensional arrays and its applications for parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 12(3):306–??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0306.
pdf;   http://www.computer.
org/tpds/td2001/l0306abs.
htm.

Dinda:2006:DIP

[Din06] Peter A. Dinda. Design, implementation, and performance of an extensible toolkit for resource prediction in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 17(2):160–173, February 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:1992:UTB

[DK92] C. R. Das and J. Kim. A unified task-based dependability model for hypercube computers. IEEE Transactions on Parallel and Distributed Systems, 3 (3):312–324, May 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DeLaLuz:2004:APR

[DKKS04] Victor De La Luz, Ismail Kadayif, Mahmut Kandemir, and Uger Sezer. Access pattern restructuring for memory energy. IEEE Transactions on Parallel and Distributed Systems, 15(4):289–303, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0289abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0289.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0289.
pdf.

Dumais:2002:DPD

[DL02] Guy Dumais and Hon F. Li. Distributed predicate detection in series-parallel systems. IEEE Transactions on Parallel and Distributed Systems, 13(4):373– 387, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0373.
pdf;   http://www.computer.
org/tpds/td2002/l0373abs.
htm.

deLara:2005:IAM

[dLCK^{+}05] Eyal de Lara, Yogesh Chopra, Rajnish Kumar, Nilesh Vaghela, Dan S. Wallach, and Willy Zwaenepoel. Iterative adaptation for mobile clients using existing APIs. IEEE Transactions on Parallel and Distributed Systems, 16(10):966–981, October 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duato:2005:PTD

[DLPP05] Jose Duato, Olav Lysne, Ruoming Pang, and Timothy M. Pinkston. Part I: a theory for deadlock-free dynamic network reconfiguration. IEEE Transactions on Parallel and Distributed Systems, 16(5):412–427, May 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Demirbas:2009:NQF

[DLS09] Murat Demirbas, Xuming Lu, and Puneet Singla. An innetwork querying framework for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(8): 1202–1215, August 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DeMara:1993:SPA

[DM93] Ronald F. DeMara and Dan I. Moldovan. The SNAP-1 parallel AI prototype. IEEE Transactions on Parallel and Distributed Systems, 4(8):841–854, August 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Durand:1996:IMC

[DMKJ96] Dannie Durand, Thierry Montaut, Lionel Kervella, and William Jalby. Impact of memory contention on dynamic scheduling on NUMA multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 7(11):1201–1214, November 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//neumann.computer.org/td/
books/td1996/pdf/l1201.pdf;
 http://www.computer.org/
tpds/td1996/l1201abs.htm.

DePrisco:2001:SCP

[DMR01] R. De Prisco, D. Malkhi, and M. Reiter. On k-set consensus problems in asynchronous systems. IEEE Transactions on Parallel and Distributed Systems, 12(1):7–??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0007.
pdf;   http://www.computer.
org/tpds/td2001/l0007abs.
htm.

Das:1993:AMM

[DMTB93] Chita R. Das, Prasant Mohapatra, Lei Tien, and Laxmi N. Bhuyan. An availability model for MIN-based multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 4(10): 1118–1129, October 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dutot:2009:SPT

[DNSC09] Pierre-Francois Dutot, Tchimou N’Takpe, Frederic Suter, and Henri Casanova. Scheduling parallel task graphs on (almost) homogeneous multicluster platforms. IEEE Transactions on Parallel and Distributed Systems, 20(7):940–952, July 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dogan:2002:MSA

[DÖ02] A. Dogan and F. Ösgünger. Matching and scheduling algorithms for minimizing execution time and failure probability of applications in heterogeneous computing. IEEE Transactions on Parallel and Distributed Systems, 13(3):308–323, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0308.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0308abs.
                                        htm.

Donovan:1991:PSM

[Don91] Kevin Donovan. Performance of shared memory in a parallel computer. IEEE Transactions on Parallel and Distributed Systems, 2(2):253–256, April 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duato:2001:GTD

[DP01] J. Duato and T. M. Pinkston. A general theory for deadlock-free adaptive routing using a mixed set of resources. IEEE Transactions on Parallel and Distributed Systems, 12(12):1219–1235, December 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1219.
pdf;   http://www.computer.
org/tpds/td2001/l1219abs.
htm.

Das:2002:LBO

[DP02] Sajal K. Das and Cristina M.

Pinotti. Load balanced and optimal disk allocation strategy for partial match queries on multidimensional files. IEEE Transactions on Parallel and Distributed Systems, 13(12):1211–1219, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1211abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1211.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1211.
pdf.

Dimakopoulos:2006:PFB

[DP06] Vassilios V. Dimakopoulos and Evaggelia Pitoura. On the performance of flooding-based resource discovery. IEEE Transactions on Parallel and Distributed Systems, 17(11):1242– 1252, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:2008:DHS

[DPH08] Saumitra M. Das, Himabindu Pucha, and Y. Charlie Hu. Distributed hashing for scalable multicast in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 19(3):347–362, March 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dahan:2009:DST

[DPN09] Sylvain Dahan, Laurent Philippe, and Jean-Marc Nicod. The distributed spanning tree structure. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1738–1751, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:1996:COL

[DPS96a] S. K. Das, M. C. Pinotti, and F. Sarkar. Correction to “Optimal and Load Balanced Mapping of Parallel Priority Queues in Hypercubes”. IEEE Transactions on Parallel and Distributed Systems, 7(8):896–??, August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1996/

l0896abs.htm. See [DPS96b].

Das:1996:OLB

[DPS96b] Sajal K. Das, M. Cristina Pinotti, and Falguni Sarkar. Optimal and load balanced mapping of parallel priority queues in hypercubes. IEEE Transactions on Parallel and Distributed Systems, 7(6):555–564, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/

l0555abs.htm. See correction [DPS96a].

Darte:1994:CMS

[DR94] Alain Darte and Yves Robert.

Constructive methods for scheduling uniform loop nests. IEEE Transactions on Parallel and Distributed Systems, 5(8):814– 822, August 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dimitrov:1998:APT

[DR98] B. Dimitrov and V. Rego.

Arachne: a portable threads system supporting migrant threads on heterogeneous network farms. IEEE Transactions on Parallel and Distributed Systems, 9(5):459–??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0459.
pdf;   http://www.computer.
org/tpds/td1998/l0459abs.
htm.

Dingle:1994:EMP

[DS94] Adair Dingle and I. Hal Sudborough. Efficient mappings of pyramid networks. IEEE Transactions on Parallel and Distributed Systems, 5(10):1009– 1017, October 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dahlgren:1996:EHB

[DS96] Fredrik Dahlgren and Per Stenström. Evaluation of hardwarebased stride and sequential prefetching in shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 7(4):385–398, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0385abs.htm.

DiStefano:2002:LMA

[DS02] Antonella Di Stefano and Corrado Santoro. Locating mobile agents in a wide distributed environment. IEEE Transactions on Parallel and Distributed Systems, 13(8):844–864, August 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0844.
pdf;   http://www.computer.
org/tpds/td2002/l0844abs.
htm.

Datta:2003:SRP

[DS03a] Amitava Datta and Subbiah Soundaralakshmi. Summation and routing on a partitioned optical passive stars network with large group size. IEEE Transactions on Parallel and Distributed Systems, 14(12):1275–1285, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1275abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1275.
pdf.

Dolev:2003:CAS

[DS03b] Shlomi Dolev and Elad Schiller.

Communication adaptive selfstabilizing group membership service. IEEE Transactions on Parallel and Distributed Systems, 14(7):709–720, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/07/l0709abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/07/l0709.
pdf.

Decayeux:2005:HNM

[DS05] Catherine Decayeux and David Seme. 3D hexagonal network: Modeling, topological properties, addressing scheme, and optimal routing algorithm. IEEE Transactions on Parallel and Distributed Systems, 16(9):875– 884, September 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

daSilva:2003:PPA

[dSF03] Edson Luiz da Silva and Erlon Cristian Finardi. Parallel processing applied to the planning of hydrothermal systems. IEEE Transactions on Parallel and Distributed Systems, 14(8):721–729, August 2003. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0721abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/08/l0721.
pdf.

Datta:2002:FSA

[DSO02] A. Datta, S. Soundaralakshmi, and R. Owens. Fast sorting algorithms on a linear array with a reconfigurable pipelined bus system. IEEE Transactions on Parallel and Distributed Systems, 13(3):212–222, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0212.
pdf;   http://www.computer.
org/tpds/td2002/l0212abs.
htm.

Daniel:1999:RAF

[DSY99] S. W. Daniel, K. G. Shin, and S. K. Yun. A router architecture for flexible routing and switching in multihop point-to-point networks. IEEE Transactions on Parallel and Distributed Systems, 10(1):62–??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0062.
pdf;   http://www.computer.
org/tpds/td1999/l0062abs.
htm.

Day:1994:CST

[DT94] Khaled Day and Anand Tripathi. A comparative study of topological properties of hypercubes and star graphs. IEEE Transactions on Parallel and Distributed Systems, 5(1):31–38, January 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DeMara:2007:TAD

[DTE07] Ronald F. DeMara, Yili Tseng, and Abdel Ejnioui. Tiered algorithm for distributed process quiescence and termination detection. IEEE Transactions on Parallel and Distributed Systems, 18(11):1529–1538, November 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duato:1993:NTD

[Dua93] Jose Duato. New theory of deadlock-free adaptive routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 4(12):1320– 1331, December 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [Dua95a] and comment [VS11].

Duato:1995:NSC

[Dua95a] Jose Duato. A necessary and sufficient condition for deadlockfree adaptive routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 6(10):1055– 1067, October 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/

l1055abs.htm. See [Dua93] and comment [VS11].

Duato:1995:TDF

[Dua95b] Jose Duato. A theory of deadlock-free adaptive multicast routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 6(9): 976–987, September 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0976abs.htm.

Duato:1996:NSC

[Dua96] Jose Duato. A necessary and sufficient condition for deadlockfree routing in cut-through and store-and-forward networks. IEEE Transactions on Parallel and Distributed Systems, 7(8): 841–854, August 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0841abs.htm.

Duato:1997:TFT

[Dua97] J. Duato. A theory of faulttolerant routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 8(8):790–802, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0790.
pdf;   http://www.computer.

org/tpds/td1997/l0790abs.
htm.

Dobber:2009:DLB

[DvdMK09] Menno Dobber, Rob van der Mei, and Ger Koole. Dynamic load balancing and job replication in a global-scale grid environment: a comparison. IEEE Transactions on Parallel and Distributed Systems, 20(2):207– 218, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Doulamis:2007:FSA

[DVV07] Nikolaos Doulamis, Emmanouel Varvarigos, and Theodora Varvarigou. Fair scheduling algorithms in Grids. IEEE Transactions on Parallel and Distributed Systems, 18(11):1630– 1648, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Drozdowski:2003:CDL

[DW03] Maciej Drozdowski and Pawel Wolniewicz. Out-of-core divisible load processing. IEEE Transactions on Parallel and Distributed Systems, 14(10): 1048–1056, October 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2003/10/
l1048abs.htm;   http://csdl.
computer.org/dl/trans/td/
2003/10/l1048.pdf.

Dai:2004:ELA

[DW04a] Fei Dai and Jie Wu. An extended localized algorithm for connected dominating set formation in ad hoc wireless networks. IEEE Transactions on Parallel and Distributed Systems, 15(10):908–920, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/10/l0908abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0908.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0908.
pdf.

Dai:2004:PAB

[DW04b] Fei Dai and Jie Wu. Performance analysis of broadcast protocols in ad hoc networks based on self-pruning. IEEE Transactions on Parallel and Distributed Systems, 15(11):1027–1040, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/11/l1027abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1027.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1027.
pdf.

Dai:2006:EBA

[DW06] Fei Dai and Jie Wu. Efficient broadcasting in ad hoc wireless networks using directional antennas. IEEE Transactions on Parallel and Distributed Systems, 17(4):335–347, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ding:2009:APS

[DWX09] Yong Ding, Chen Wang, and Li Xiao. An adaptive partitioning scheme for sleep scheduling and topology control in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(9):1352– 1365, September 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dan:1993:PAB

[DY93] A. Dan and P. S. Yu. Performance analysis of buffer coherency policies in a multisystem data sharing environment. IEEE Transactions on Parallel and Distributed Systems, 4(3): 289–305, March 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dharmasena:2005:OFT

[DY05] Hettiehe P. Dharmasena and Xin Yan. An optimal faulttolerant routing algorithm for weighted bidirectional doubleloop networks. IEEE Transactions on Parallel and Distributed Systems, 16(9):841–852, September 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dan:1997:RAD

[DYJ97] Asit Dan, Philip S. Yu, and Anant Jhingran. Recovery analysis of data sharing systems under deferred dirty page propagation policies. IEEE Transactions on Parallel and Distributed Systems, 8(7):695–711, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0695.
pdf;   http://www.computer.
org/tpds/td1997/l0695abs.
htm.

Datta:2004:EEP

[DZ04] Amitava Datta and Albert Y. Zomaya. An energy-efficient permutation routing protocol for single-hop radio networks. IEEE Transactions on Parallel and Distributed Systems, 15(4):331–338, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0331abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0331.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0331.
pdf.

Duan:2005:FTO

[DZH05] Zhenhai Duan, Zhi-Li Zhang, and Yiwei Thomas Hou. Fundamental trade-offs in aggregate packet scheduling. IEEE Transactions on Parallel and Distributed Systems, 16(12):1166– 1177, December 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2004:CSB

[DZHG04] Zhenhai Duan, Zhi-Li Zhang, Yiwei Thomas Hou, and Lixin Gao. A core stateless bandwidth broker architecture for scalable support of guaranteed services. IEEE Transactions on Parallel and Distributed Systems, 15(2):167–182, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0167abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0167.
pdf.

El-Amaway:1993:CAL

[EA93] A. El-Amaway. Clocking arbitrarily large computing structures under constant skew bound. IEEE Transactions on Parallel and Distributed Systems, 4(3):241–255, March 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ercal:2000:SID

[EAF00] F. Ercal, M. Allen, and H. Feng. A systolic image difference algorithm for RLEcompressed images. IEEE Transactions on Parallel and Distributed Systems, 11(5):433– ??, May 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0433.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0433abs.
                                        htm.

El-Amawy:1995:PGB

[EAK95] Ahmed El-Amawy and Priyalal Kulasinghe. Properties of generalized branch and combine clock networks. IEEE Transactions on Parallel and Distributed Systems, 6(5):541–546, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0541abs.htm.

El-Amawy:1997:AMF

[EAK97] Ahmed El-Amawy and Priyalal Kulasinghe. Algorithmic mapping of feedforward neural networks onto multiple bus systems. IEEE Transactions on Parallel and Distributed Systems, 8(2):130–136, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0130.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0130abs.
                                        htm.

El-Amawy:1991:PPF

[EAL91] Ahmed El-Amawy and Shahram Latifi. Properties and performance of folded hypercubes. IEEE Transactions on Parallel and Distributed Systems, 2 (1):31–42, January 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Eisenhauer:2002:NDR

[EBS02] Greg Eisenhauer, Fabián E. Bustamante, and Karsten Schwan. Native data representation: An efficient wire format for highperformance distributed computing. IEEE Transactions on Parallel and Distributed Systems, 13(12):1234–1246, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1234abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1234.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1234.
pdf.

Ezhilchelvan:2004:TBM

[EBS04] Paul D. Ezhilchelvan, Francisco V. Brasileiro, and Neil A. Speirs. A timeout-based message ordering protocol for a lightweight software implementation of TMR systems. IEEE Transactions on Parallel and Distributed Systems, 15(1):53– 65, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/01/
extras/L0053s.pdf;   http:
//csdl.computer.org/comp/
trans/td/2004/01/l0053abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/01/l0053.
pdf.

Eltayeb:2006:CSE

[EDO06] Mohammed S. Eltayeb, Atakan Dogan, and Fusun Ozguner. Concurrent scheduling: Efficient heuristics for online largescale data transfers in distributed real-time environments. IEEE Transactions on Parallel and Distributed Systems, 17(11): 1348–1359, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Efe:1995:PNL

[EF95] Kemal Efe and Antonio Fernandez. Products of networks with logarithmic diameter and fixed degree. IEEE Transactions on Parallel and Distributed Systems, 6(9):963–975, September 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1995/
                                        l0963abs.htm.

Efe:1996:MCT

[EF96] Kemal Efe and Antonio Fernandez. Mesh-connected trees: a bridge between grids and meshes of trees. IEEE Transactions on Parallel and Distributed Systems, 7(12):1281–1291, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Efe:1992:CCA

[Efe92] Kemal Efe. The crossed cube architecture for parallel computation. IEEE Transactions on Parallel and Distributed Systems, 3(5):513–524, September 1992.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Evripidou:1993:BSI

[EG93] Paraskevas Evripidou and Jean-Luc Gaudiot. Block scheduling of iterative algorithms and graph-level priority scheduling in a simulated data-flow multiprocessor. IEEE Transactions on Parallel and Distributed Systems, 4(4):398–413, April 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Edelman:1994:ITA

[EHJ94] Alan Edelman, Steve Heller, and S. Lennart Johnsson. Index transformation algorithms in a linear algebra framework. IEEE Transactions on Parallel and Distributed Systems, 5(12): 1302–1309, December 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Eigenmann:1998:APP

[EHP98] R. Eigenmann, J. Hoeflinger, and D. Padua. On the automatic parallelization of the Perfect Benchmarks. IEEE Transactions on Parallel and Distributed Systems, 9(1):5–23, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0005.
pdf;   http://www.computer.
org/tpds/td1998/l0005abs.
htm.

Efe:1995:OSC

[EK95] Kemal Efe and Venkatesh Krishnamoorthy. Optimal scheduling of compute-intensive tasks on a network of workstations. IEEE Transactions on Parallel and Distributed Systems, 6 (6):668–673, June 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0668abs.htm.

El-Kadi:2002:RBB

[EKOAW02] Mona El-Kadi, Stephan Olariu, and Hussein Abdel-Wahab. A rate-based borrowing scheme for QoS provisioning in multimedia wireless networks. IEEE Transactions on Parallel and Distributed Systems, 13(2):156–166, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0156.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0156abs.
                                        htm.

Ezhilchelvan:1990:PES

[EMS90] Paul D. Ezhilchelvan, Isi Mitrani, and Santosh K. Shrivastava. A performance evaluation study of pipeline TMR systems. IEEE Transactions on Parallel and Distributed Systems, 1(4): 442–456, October 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Elbirt:2005:ILD

[EP05] Adam J. Elbirt and Christof Paar. An instruction-level distributed processor for symmetrickey cryptography. IEEE Transactions on Parallel and Distributed Systems, 16(5):468–480, May 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

England:2007:RST

[EVW07] Darin England, Bharadwaj Veeravalli, and Jon B. Weissman. A robust spanning tree topology for data collection and dissemination in distributed environments. IEEE Transactions on Parallel and Distributed Systems, 18(5):608–620, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

ElGindy:1997:SVD

[EW97] H. ElGindy and L. Wetherall. A simple Voronoi diagram algorithm for a reconfigurable mesh. IEEE Transactions on Parallel and Distributed Systems, 8(11):1133–1142, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1133.
pdf;   http://www.computer.
org/tpds/td1997/l1133abs.
htm.

Fragopoulou:1994:PAC

[FA94] Paraskevi Fragopoulou and Selim G. Akl. Parallel algorithm for computing Fourier transforms on the star graph. IEEE Transactions on Parallel and Distributed Systems, 5(5):525– 531, May 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:1998:DMC

[Fan98] J. Fan. Diagnosability of the Möbius cubes. IEEE Transactions on Parallel and Distributed Systems, 9(9):923–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0923.
pdf;   http://www.computer.
org/tpds/td1998/l0923abs.
htm.

Fan:2002:DCCa

[Fan02a] Jianxi Fan. Diagnosability of crossed cubes under the comparison diagnosis model. IEEE Transactions on Parallel and Distributed Systems, 13(7):687– 692, July 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://computer.org/tpds/
td2002/l1099.pdf;   http:/
/dlib.computer.org/td/books/
td2002/pdf/l0687.pdf;   http:
//www.computer.org/tpds/td2002/

l0687abs.htm. Editor’s Note: This paper unfortunately contains some errors which led to the paper being reprinted in the October 2002 issue. Please see IEEE Transactions on Parallel and Distributed Systems, vol.

13, no. 10, October 2002, pp. 1099-1104 for the correct paper.

Fan:2002:DCCb

[Fan02b] Jianxi Fan. Diagnosability of crossed cubes under the comparison diagnosis model. IEEE Transactions on Parallel and Distributed Systems, 13(10): 1099–1104, October 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2002/pdf/l1099.
pdf;   http://www.computer.
org/tpds/td2002/l1099abs.
htm.

Futamura:2002:EPA

[FARH02] Natsuhiko Futamura, Srinivas Aluru, Desh Ranjan, and Bhanu Hariharan. Efficient parallel algorithms for solvent accessible surface area of proteins. IEEE Transactions on Parallel and Distributed Systems, 13(6):544– 555, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0544.
pdf;   http://www.computer.
org/tpds/td2002/l0544abs.
htm.

Figueira:2001:SMA

[FB01a] S. M. Figueira and F. Berman.

A slowdown model for applications executing on time-shared clusters of workstations. IEEE Transactions on Parallel and Distributed Systems, 12(6):653– ??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0653.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0653abs.
                                        htm.

Franceschetti:2001:GMA

[FB01b] Massimo Franceschetti and Jehoshua Bruck. A group membership algorithm with a practical specification. IEEE Transactions on Parallel and Distributed Systems, 12(11):1190– 1200, November 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td2001/pdf/l1190.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1190abs.
                                        htm.

Freedman:1996:SSP

[FBD96] Craig S. Freedman, Josef Burger, and David J. De-Witt. SPIFFI: a scalable parallel file system for the Intel Paragon. IEEE Transactions on Parallel and Distributed Systems, 7(11):1185–1200, November 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //neumann.computer.org/td/
                                        books/td1996/pdf/l1185.pdf;
                                         http://www.computer.org/
                                        tpds/td1996/l1185abs.htm.

Fan:1991:GSA

[FC91] Zhenqiang Fan and Kam-Hoi Cheng. A generalized simultaneous access dictionary machine.

IEEE Transactions on Parallel and Distributed Systems, 2 (2):149–159, April 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2000:ARO

[FCF00] Y. Fang, I. Chlamtac, and H.-B. Fei. Analytical results for optimal choice of location update interval for mobility database failure restoration in PCS networks. IEEE Transactions on Parallel and Distributed Systems, 11(6):615–??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0615.
pdf;   http://www.computer.
org/tpds/td2000/l0615abs.
htm.

Farrag:1994:FTE

[FD94] Abdel Aziz Farrag and Robert J.

Dawson. Fault-tolerant extension problem for complete multipartite networks. IEEE Transactions on Parallel and Distributed Systems, 5(2):205–210, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fleury:2000:PME

[FDC00] M. Fleury, A. C. Downton, and A. F. Clark. Performance metrics for embedded parallel pipelines. IEEE Transactions on Parallel and Distributed Systems, 11(11):164–??, November 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l1164.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l1164abs.
                                        htm.

Fernandez:1997:GAP

[FE97] A. Fernández and K. Efe. Generalized algorithm for parallel sorting on product networks. IEEE Transactions on Parallel and Distributed Systems, 8(12):1211–1225, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1211.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1211abs.
                                        htm.

Feitelson:2005:EAR

[Fei05] Dror G. Feitelson. Experimental analysis of the root causes of performance evaluation results: a backfilling case study. IEEE Transactions on Parallel and Distributed Systems, 16(2): 175–182, February 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fleury:1998:GTD

[FF98] E. Fleury and P. Fraigniaud. A general theory for deadlock avoidance in wormholerouted networks. IEEE Transactions on Parallel and Distributed Systems, 9(7):626–??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0626.
pdf;   http://www.computer.
org/tpds/td1998/l0626abs.
htm.

Frachtenberg:2005:APJ

[FFPF05] Eitan Frachtenberg, Dror G.

Feitelson, Fabrizio Petrini, and Juan Fernandez. Adaptive parallel job scheduling with flexible coscheduling. IEEE Transactions on Parallel and Distributed Systems, 16(11):1066– 1077, November 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Frolund:2001:ITA

[FG01] S. Frølund and R. Guerraoui.

Implementing E-transactions with asynchronous replication. IEEE Transactions on Parallel and Distributed Systems, 12(2):133–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0133.
pdf;   http://www.computer.
org/tpds/td2001/l0133abs.
htm.

Francalanci:2006:HPS

[FG06a] Chiara Francalanci and Paolo Giacomazzi. High-performance self-routing algorithm for multiprocessor systems with shuffle interconnections. IEEE Transactions on Parallel and Distributed Systems, 17(1):38–50, January 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Frey:2006:GCB

[FG06b] Hannes Frey and Daniel Gorgen. Geographical cluster-based routing in sensing-covered networks. IEEE Transactions on Parallel and Distributed Systems, 17(9): 899–911, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fiduccia:1997:ECS

[FH97] C. M. Fiduccia and P. J. Hedrick. Edge congestion of shortest path systems for allto-all communication. IEEE Transactions on Parallel and Distributed Systems, 8(10):1043– 1054, October 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td1997/pdf/l1043.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1043abs.
                                        htm.

Fujimoto:2003:ABS

[FH03] Noriyuki Fujimoto and Kenichi Hagihara. On approximation of the bulk synchronous task scheduling problem. IEEE Transactions on Parallel and Distributed Systems, 14(11): 1191–1199, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/11/l1191abs.
                                        htm;   http://csdl.computer.

org/dl/trans/td/2003/11/l1191.
pdf.

Fallahi:2006:QET

[FHA06] Afshin Fallahi, Ekram Hossain, and Attahiru S. Alfa. QoS and energy trade off in distributed energy-limited mesh/ relay networks: a queuing analysis. IEEE Transactions on Parallel and Distributed Systems, 17 (6):576–592, June 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Fink:1997:PCI

[FHBJ97] S. J. Fink, C. Huston, S. B.

Baden, and K. Jansen. Parallel cluster identification for multidimensional lattices. IEEE Transactions on Parallel and Distributed Systems, 8(11):1089– 1097, November 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l1089.
pdf;   http://www.computer.
org/tpds/td1997/l1089abs.
htm.

Franaszek:1993:DCC

[FHRT93] Peter A. Franaszek, Jayant R.

Haritsa, John T. Robinson, and Alexander Thomasian. Distributed concurrency control based on limited wait-depth. IEEE Transactions on Parallel and Distributed Systems, 4(11): 1246–1264, November 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fujiwara:1995:OGS

[FI95] Hideo Fujiwara and Tomoo Inoue. Optimal granularity and scheme of parallel test generation in a distributed system. IEEE Transactions on Parallel and Distributed Systems, 6 (7):677–686, July 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0677abs.htm.

Fiduccia:1992:BHO

[Fid92] Charles M. Fiduccia. Bused hypercubes and other pin-optimal networks. IEEE Transactions on Parallel and Distributed Systems, 3(1):14–24, January 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fritzke:2001:CBF

[FIMR01] U. Fritzke, Jr., P. Ingels, A. Mostefaoui, and M. Raynal. Consensus-based faulttolerant total order multicast. IEEE Transactions on Parallel and Distributed Systems, 12(2):147–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0147.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0147abs.
                                        htm.

Fan:2007:OEP

[FJL07] Jianxi Fan, Xiaohua Jia, and Xiaola Lin. Optimal embeddings of paths with various lengths in twisted cubes. IEEE Transactions on Parallel and Distributed Systems, 18(4):511–521, April 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:1998:ESL

[FJY98] C. Fu, X. Jiao, and T. Yang.

Efficient sparse LU factorization with partial pivoting on distributed memory architectures. IEEE Transactions on Parallel and Distributed Systems, 9(2):109–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0109.
pdf;   http://www.computer.
org/tpds/td1998/l0109abs.
htm.

Fan:2005:OPE

[FLJ05] Jianxi Fan, Xiaola Lin, and Xiaohua Jia. Optimal path embedding in crossed cubes. IEEE Transactions on Parallel and Distributed Systems, 16(12): 1190–1200, December 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Flich:2002:BPMa

[FLMD02a] José Flich, Pedro López, M. P.

Malumbres, and Jose Duato. Boosting the performance of Myrinet networks. IEEE Transactions on Parallel and Distributed Systems, 13(7):693–709, July 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0693.
pdf;   http://www.computer.
org/tpds/td2002/l0693abs.

htm. Editor’s Note: This paper unfortunately contains some errors which led to the paper being reprinted in the November 2002 issue. Please see IEEE Transactions on Parallel and Distributed Systems, vol. 13, no. 11, November 2002, pp. 1166–1182 for the correct paper.

Flich:2002:BPMb

[FLMD02b] José Flich, Pedro López, M. P. Malumbres, and Jose Duato. Boosting the performance of Myrinet networks. IEEE Transactions on Parallel and Distributed Systems, 13(11):1166–1182, November 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/11/l1166abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/11/l1166.
htm;   http://csdl.computer.
org/dl/trans/td/2002/11/l1166.
pdf.

Freeh:2007:AET

[FLP^{+}07] Vincent W. Freeh, David K. Lowenthal, Feng Pan, Nandini Kappiah, Rob Springer, Barry L. Rountree, and Mark E. Femal. Analyzing the energytime trade-off in high-performance computing applications. IEEE Transactions on Parallel and Distributed Systems, 18(6):835– 848, June 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fernandez:1995:LTU

[FLVG95] Agustin Fernandez, Jose M.

Llaberia, and Miguel Valero-Garcia. Loop transformation using nonunimodular matrices. IEEE Transactions on Parallel and Distributed Systems, 6(8): 832–840, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0832abs.htm.

Feldman:2009:PSA

[FLZ09] Michal Feldman, Kevin Lai, and Li Zhang. The proportionalshare allocation market for computational resources. IEEE Transactions on Parallel and Distributed Systems, 20(8):1075– 1088, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Friedman:2002:SSD

[FMG02] Roy Friedman, Shiri Manor, and Katherine Guo. Scalable stability detection using logical hypercube. IEEE Transactions on Parallel and Distributed Systems, 13(9):972–984, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

                                        td/books/td2002/pdf/l0972.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0972abs.
                                        htm.

Friedman:2007:RPL

[FMR07] Roy Friedman, Achour Mostéfaoui, and Michel Raynal. On the respective power of [lozenge]P and [lozenge]S to solve oneshot agreement problems. IEEE Transactions on Parallel and Distributed Systems, 18(5):589– 597, May 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Franke:2005:CCA

[FO05] Bjorn Franke and Michael F. P. O’Boyle. A complete compiler approach to auto-parallelizing C programs for multi-DSP systems. IEEE Transactions on Parallel and Distributed Systems, 16(3):234–245, March 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Foster:1991:AGS

[Fos91] Ian Foster. Automatic generation of self-scheduling programs. IEEE Transactions on Parallel and Distributed Systems, 2 (1):68–78, January 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fernandez-Pascual:2008:ETC

[FPGAD08] Ricardo Fernández-Pascual, José M. García, Manuel E. Acacio, and José Duato. Extending the TokenCMP cache coherence protocol for low overhead fault tolerance in CMP architectures. IEEE Transactions on Parallel and Distributed Systems, 19(8): 1044–1056, August 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Feehrer:1996:PSS

[FR96] John R. Feehrer and Lars H.

Ramfelt. Packet synchronization for synchronous optical deflection-routed interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 7(6):605–611, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0605abs.htm.

Ferreira:2007:RPD

[FRGJ07] Ronaldo A. Ferreira, Murali K.

Ramanathan, Ananth Grama, and Suresh Jagannathan. Randomized protocols for duplicate elimination in peer-to-peer storage systems. IEEE Transactions on Parallel and Distributed Systems, 18(5):686–696, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fonseca:2009:FIL

[FRGL09] Pedro Fonseca, Rodrigo Rodrigues, Anjali Gupta, and Barbara Liskov. Full-information lookups for peer-to-peer overlays. IEEE Transactions on Parallel and Distributed Systems, 20 (9):1339–1351, September 2009.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fahringer:2000:USE

[FS00] T. Fahringer and B. Scholz. A unified symbolic evaluation framework for parallelizing compilers. IEEE Transactions on Parallel and Distributed Systems, 11(11):105–??, November 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l1105.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l1105abs.
                                        htm.

Fu:1997:CLB

[FT97] Shiwa S. Fu and Nian-Feng Tzeng. A circular list-based mutual exclusion scheme for large shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 8(6):628–639, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0628.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0628abs.

htm. See comments [HS98b].

Fu:1997:DOQ

[Fu97]  Ada Waichee Fu.  Delay-

optimal quorum consensus for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 8(1):59–69, January 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2005:HWR

[Fu05] Jung-Sheng Fu. Hamiltonicity of the WK-recursive network with and without faulty nodes. IEEE Transactions on Parallel and Distributed Systems, 16(9): 853–865, September 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2009:HNA

[FV09] Xiuduan Fang and Malathi Veeraraghavan. A hybrid network architecture for file transfers. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1714–1725, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ferreira:2003:ODI

[FVR03] Paulo Ferreira, Luís Veiga, and Carlos Ribeiro. OBI-WAN: Design and implementation of a middleware platform. IEEE Transactions on Parallel and Distributed Systems, 14(11):1086–1099, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/11/l1086abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/11/l1086.
pdf.

Fay-Wolfe:2000:RTC

[FWDC^{+}00] V. Fay-Wolfe, L. C. DiPippo, G. Cooper, R. Johnston, P. Kortmann, and B. Thuraisingham. Real-time CORBA. IEEE Transactions on Parallel and Distributed Systems, 11(10):73– ??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1073.
pdf;   http://www.computer.
org/tpds/td2000/l1073abs.
htm.

Fu:2009:TPF

[FYJ^{+}09] Xinwen Fu, Wei Yu, Shu Jiang, Steve Graham, and Yong Guan. TCP performance in flow-based mix networks: Modeling and analysis. IEEE Transactions on Parallel and Distributed Systems, 20(5):695–709, May 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Faraj:2007:MSS

[FYP07] Ahmad Faraj, Xin Yuan, and Pitch Patarasuk. A message scheduling scheme for all-to-all personalized communication on Ethernet switched clusters. IEEE Transactions on Parallel and Distributed Systems, 18(2): 264–276, February 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2005:RAB

[FYS05] Lee-Juan Fan, Chang-Biau Yang, and Shyue-Horng Shiau.

Routing algorithms on the busbased hypercube network. IEEE Transactions on Parallel and Distributed Systems, 16(4):335– 348, April 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Firoiu:2006:PSG

[FZGC06] Victor Firoiu, Xiaohui Zhang, Emre Gunduzhan, and Nicolas Christin. Providing service guarantees in high-speed switching systems with feedback output queuing. IEEE Transactions on Parallel and Distributed Systems, 17(5):434–447, May 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fernandez-Zepeda:1998:SSF

[FZVT98] J. A. Fernández-Zepeda, R. Vaidyanathan, and J. L. Trahan. Scaling simulation of the fusing-restricted reconfigurable mesh. IEEE Transactions on Parallel and Distributed Systems, 9(9):861–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0861.
pdf;   http://www.computer.
org/tpds/td1998/l0861abs.
htm.

Gabber:1990:VPT

[Gab90] Eran Gabber. VMMP: a practical tool for the development of portable and efficient programs for multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 1(3):304– 317, July 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Govindarajan:1996:FRC

[GAG96] R. Govindarajan, Erik R. Altman, and Guang R. Gao. A framework for resourceconstrained rate-optimal software pipelining. IEEE Transactions on Parallel and Distributed Systems, 7(11):1133– 1149, November 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://neumann.
computer.org/td/books/td1996/
pdf/l1133.pdf;   http://www.
computer.org/tpds/td1996/
l1133abs.htm.

Goumas:2003:ECG

[GAK03] Georgios Goumas, Maria Athanasaki, and Nectarios Koziris. An efficient code generation technique for tiled iteration spaces. IEEE Transactions on Parallel and Distributed Systems, 14 (10):1021–1034, October 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/10/l1021abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/10/l1021.
pdf.

Garcia:2001:FID

[GAL01] J. Garcia, E. Ayguadé, and J. Labarta. A framework for integrating data alignment, distribution, and redistribution in distributed memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 12(4):416–??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0416.
pdf;   http://www.computer.
org/tpds/td2001/l0416abs.
htm.

Gupta:1992:DAD

[GB92] Manish Gupta and Prithviraj Banerjee. Demonstration of automatic data partitioning techniques for parallelizing compilers on multicomputers. IEEE Transactions on Parallel and Distributed Systems, 3(2):179– 193, March 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gafsi:2000:MPC

[GB00] J. Gafsi and E. W. Biersack. Modeling and performance comparison of reliability strategies for distributed video servers. IEEE Transactions on Parallel and Distributed Systems, 11(4):412–??, April 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0412.
pdf;   http://www.computer.
org/tpds/td2000/l0412abs.
htm.

Guo:2006:LBC

[GB06] Jiani Guo and Laxmi Narayan Bhuyan. Load balancing in a cluster-based Web server for multimedia applications. IEEE Transactions on Parallel and Distributed Systems, 17(11): 1321–1334, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2007:NDM

[GBC^{+}07] Wenjun Gu, Xiaole Bai, Sriram Chellappan, Dong Xuan, and Weijia Jia. Network decoupling: a methodology for secure communications in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(12):1784–1796, December 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ghosh:2007:GTB

[GBD07] Preetam Ghosh, Kalyan Basu, and Sajal K. Das. A game theory-based pricing strategy to support single/multiclass job allocation schemes for bandwidthconstrained distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 18(3):289– 306, March 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2007/03/
                                        l0289s.pdf.

Garcia-Carballeira:2004:ACC

[GCCC^{+}04] Felix Garcia-Carballeira, Jesus Carretero, Alejandro Calderon, Jose M. Perez, and Jose D. Garcia. An adaptive cache coherence protocol specification for parallel input/output systems. IEEE Transactions on Parallel and Distributed Systems, 15 (6):533–545, June 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0533.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0533.pdf.

Ghandeharizadeh:1994:MMD

[GD94] Shahram Ghandeharizadeh and David J. DeWitt. MAGIC: a multiattribute declustering mechanism for multiprocessor database machines. IEEE Transactions on Parallel and Distributed Systems, 5(5):509– 524, May 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ghose:1995:HCN

[GD95] Kanad Ghose and Kiran R. Desai. Hierarchical cubic networks. IEEE Transactions on Parallel and Distributed Systems, 6 (4):427–435, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0427abs.htm. See comments

[YP98].

Goswami:1993:PBD

[GDI93] Kumar K. Goswami, Murthy Devarakonda, and Ravishankar K. Iyer. Prediction-based dynamic load-sharing heuristics. IEEE Transactions on Parallel and Distributed Systems, 4(6):638– 648, June 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ghosh:1994:CPL

[GDJ94] Joydeep Ghosh, Sajal K. Das, and Ajita John. Concurrent processing of linearly ordered data structures on hypercube multicomputers. IEEE Transactions on Parallel and Distributed Systems, 5(9):898–911, September 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goumas:2009:CAS

[GDK09] Georgios Goumas, Nikolaos Drosinos, and Nectarios Koziris. Communication-aware supernode shape. IEEE Transactions on Parallel and Distributed Systems, 20(4):498–511, April 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gehani:1993:CSM

[Geh93] Narain H. Gehani. Capsules: a shared memory access mechanism for Concurrent C/C++. IEEE Transactions on Parallel and Distributed Systems, 4 (7):795–811, July 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gennaro:2000:PAI

[Gen00] R. Gennaro. A protocol to achieve independence in constant rounds. IEEE Transactions on Parallel and Distributed Systems, 11(7):636–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0636.
pdf;   http://www.computer.
org/tpds/td2000/l0636abs.
htm.

GonzalezdeMendivil:1999:DDR

[GFG^{+}99] J. R. González de Mendívil, F. Fariña, J. R. Garitagoitia, C. F. Alastruey, and J. M. Bernabeu-Auban. A distributed deadlock resolution algorithm for the AND model. IEEE Transactions on Parallel and Distributed Systems, 10(5):433– ??, May 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0433.
pdf;   http://www.computer.
org/tpds/td1999/l0433abs.
htm.

Garg:1994:RCG

[GG94a] Vijay K. Garg and Joydeep Ghosh. Repeated computation of global functions in a distributed environment. IEEE Transactions on Parallel and Distributed Systems, 5(8):823– 834, August 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:1994:AAT

[GG94b] Qian Ping Gu and Jun Gu. Algorithms and average time bounds of sorting on a mesh-connected computer. IEEE Transactions on Parallel and Distributed Systems, 5(3):308–315, March 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:1995:TPR

[GG95] Qian-Ping Gu and Jun Gu. Two packet routing algorithms on a mesh-connected computer. IEEE Transactions on Parallel and Distributed Systems, 6 (4):436–440, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0436abs.htm.

Garg:2009:FDM

[GG09] Nandan Garg and Daniel Grosu. A faithful distributed mechanism for sharing the cost of multicast transmissions. IEEE Transactions on Parallel and Distributed Systems, 20(8):1089– 1101, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goldberg:1993:MIS

[GH93] Aaron J. Goldberg and John L. Hennessy. Mtool: An integrated system for performance debugging shared memory multiprocessor applications. IEEE Transactions on Parallel and Distributed Systems, 4(1):28–40, January 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gossain:2006:DED

[GJDA06] Hrishikesh Gossain, Tarun Joshi, Carlos De Morais Cordeiro, and Dharma P. Agrawal. DRP: An efficient directional routing protocol for mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 17(12): 1438–1541, December 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:1993:SFP

[GK93] Anshul Gupta and Vipin Kumar. The scalability of FFT on parallel computers. IEEE Transactions on Parallel and Distributed Systems, 4(8):922–932, August 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:2006:EAE

[GKG06] Indranil Gupta, Anne-Marie Kermarrec, and Ayalvadi J. Ganesh. Efficient and adaptive epidemic-style protocols for reliable and scalable multicast. IEEE Transactions on Parallel and Distributed Systems, 17 (7):593–605, July 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:1997:HSP

[GKK97] Anshul Gupta, George Karypis, and Vipin Kumar. Highly scalable parallel algorithms for sparse matrix factorization. IEEE Transactions on Parallel and Distributed Systems, 8 (5):502–520, May 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l0502.
pdf;   http://www.computer.
org/tpds/td1997/l0502abs.
htm.

Ghose:2005:ADL

[GKK05] Debasish Ghose, Hyoung Joong Kim, and Tae Hoon Kim. Adaptive divisible load scheduling strategies for workstation clusters with unknown network resources. IEEE Transactions on Parallel and Distributed Systems, 16(10):897–907, October 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:1995:PSP

[GKS95] Anshul Gupta, Vipin Kumar, and Ahmed Sameh. Performance and scalability of preconditioned conjugate gradient methods on parallel computers. IEEE Transactions on Parallel and Distributed Systems, 6 (5):455–469, May 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0455abs.htm.

Guo:2009:PPL

[GLQL09] Hui Guo, Kwok-Tung Lo, Yi Qian, and Jiang Li. Peerto-peer live video distribution under heterogeneous bandwidth constraints. IEEE Transactions on Parallel and Distributed Systems, 20(2):233–245, February 2009. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Giaccone:2007:TRF

[GLS07] Paolo Giaccone, Emilio Leonardi, and Devavrat Shah. Throughput region of finite-buffered networks. IEEE Transactions on Parallel and Distributed Systems, 18(2):251–263, February 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2006:PRE

[GLV06] Xin Gu, Wujuan Lin, and Bharadwaj Veeravalli. Practically realizable efficient data allocation and replication strategies for distributed databases with buffer constraints. IEEE Transactions on Parallel and Distributed Systems, 17(9):1001– 1013, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/09/
l1001s.pdf.

Gedik:2007:AAS

[GLY07] Bugra Gedik, Ling Liu, and Philip S. Yu. ASAP: An adaptive sampling approach to data collection in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(12): 1766–1783, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:1994:EBX

[GM94] Zicheng Guo and Rami G. Melhem. Embedding binary Xtrees and pyramids in processor arrays with spanning buses. IEEE Transactions on Parallel and Distributed Systems, 5 (6):664–672, June 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gopi:1997:UAC

[GM97] M. Gopi and S. Manohar. A unified architecture for the computation of B-spline curves and surfaces. IEEE Transactions on Parallel and Distributed Systems, 8(12):1275–1287, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1275.
pdf;   http://www.computer.
org/tpds/td1997/l1275abs.
htm.

Groth:2009:RPD

[GM09] Paul Groth and Luc Moreau. Recording process documentation for provenance. IEEE Transactions on Parallel and Distributed Systems, 20(9):1246– 1259, September 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Golubchik:2001:DFT

[GMCB01] L. Golubchik, R. R. Muntz, C.- F. Chou, and S. Berson. Design of fault-tolerant large-scale VOD servers: With emphasis on highperformance and low-cost. IEEE Transactions on Parallel and Distributed Systems, 12(4):363– ??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0363.
pdf;   http://www.computer.
org/tpds/td2001/l0363abs.
htm.

Gong:1996:LTF

[GMG96] Chun Gong, Rami Melhem, and Rajiv Gupta. Loop transformations for fault detection in regular loops on massively parallel systems. IEEE Transactions on Parallel and Distributed Systems, 7(12):1238–1249, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ghosh:1997:FTT

[GMM97] Sunondo Ghosh, Rami Melhem, and Daniel Mossé. Faulttolerance through scheduling of aperiodic tasks in hard real-time multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 8(3):272– 284, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0272.
pdf;   http://www.computer.
org/tpds/td1997/l0272abs.
htm.

Gomez:1998:CAM

[GMR98] J. C. Gomez, E. Mascarenhas, and V. Rego. The CLAM approach to multithreaded communication on shared memory multiprocessors: Design and experiments. IEEE Transactions on Parallel and Distributed Systems, 9(1):36–49, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0036.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0036abs.
                                        htm.

Galilee:2007:PAW

[GMRC07] Bruno Galilée, Franck Mamalet, Marc Renaudin, and Pierre-Yves Coulon. Parallel asynchronous watershed algorithmarchitecture. IEEE Transactions on Parallel and Distributed Systems, 18(1):44–56, January 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Georgiou:2009:DCD

[GMS09] Chryssis Georgiou, Peter M. Musial, and Alexander A. Shvartsman. Developing a consistent domain-oriented distributed object service. IEEE Transactions on Parallel and Distributed Systems, 20(11): 1567–1585, November 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Glass:1996:FTW

[GN96] Christopher J. Glass and Lionel M. Ni. Fault-tolerant wormhole routing in meshes without virtual channels. IEEE Transactions on Parallel and Distributed Systems, 7(6):620–636, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0620abs.htm.

Gu:2006:CSA

[GN06] Xiaohui Gu and Klara Nahrstedt. On composing stream applications in peer-to-peer environments. IEEE Transactions on Parallel and Distributed Systems, 17(8):824–837, August 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gerogiannis:1993:LBR

[GO93] Dimitris Gerogiannis and Stelios C. Orphanoudakis. Load balancing requirements in parallel implementations of image feature extraction tasks. IEEE Transactions on Parallel and Distributed Systems, 4(9):994– 1013, September 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Greenberg:1997:UWR

[GO97] Ronald I. Greenberg and Hyeong-Cheol Oh. Universal wormhole routing. IEEE Transactions on Parallel and Distributed Systems, 8(3):254–262, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0254.
pdf;   http://www.computer.
org/tpds/td1997/l0254abs.
htm.

Gonzalez:2003:EAG

[Gon03] Teofilo F. Gonzalez. An efficient algorithm for gossiping in the multicasting communication environment. IEEE Transactions on Parallel and Distributed Systems, 14(7):701–708, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/07/l0701abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/07/l0701.
                                        pdf.

Gonzalez:2008:CDM

[Gon08] Teofilo F. Gonzalez. Continuous delivery message dissemination problems under the multicasting communication mode. IEEE Transactions on Parallel and Distributed Systems, 19(8): 1034–1043, August 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Girkar:1992:AEF

[GP92] Milind Girkar and Constantine D. Polychronopoulos. Automatic extraction of functional parallelism from ordinary programs. IEEE Transactions on Parallel and Distributed Systems, 3(2):166–178, March 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ganesan:1993:HDN

[GP93] Elango Ganesan and Dhiraj K. Pradhan. Hyper-deBruijn networks: scalable versatile architecture. IEEE Transactions on Parallel and Distributed Systems, 4(9):962–978, September 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Giorgi:1999:PCP

[GP99a] R. Giorgi and C. A. Prete.

PSCR: a coherence protocol for eliminating passive sharing in shared-bus shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 10(7):742–??, July 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0742.
pdf;   http://www.computer.
org/tpds/td1999/l0742abs.
htm.

Gu:1999:UHL

[GP99b] Q.-P. Gu and S. Peng. Unicast in hypercubes with large number of faulty nodes. IEEE Transactions on Parallel and Distributed Systems, 10(10):964–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0964.
pdf;   http://www.computer.
org/tpds/td1999/l0964abs.
htm.

Gu:2003:MAA

[GP03] Qian-Ping Gu and Shietung Peng. Multihop all-to-all broadcast on WDM optical networks. IEEE Transactions on Parallel and Distributed Systems, 14(5):477–486, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/05/l0477abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/05/l0477.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/05/l0477.
                                        pdf.

Gravano:1994:ADL

[GPBS94] Luis Gravano, Gustavo D. Pifarre, Pablo E. Berman, and Jorge L. C. Sanz. Adaptive deadlock-and livelock-free routing with all minimal paths in torus networks. IEEE Transactions on Parallel and Distributed Systems, 5(12):1233–1251, December 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gidenstam:2009:ERL

[GPST09] Anders Gidenstam, Marina Papatriantafilou, Hakan Sundell, and Philippas Tsigas. Efficient and reliable lock-free memory reclamation based on reference counting. IEEE Transactions on Parallel and Distributed Systems, 20(8):1173–1187, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gertner:1990:PAD

[GR90] Izidor Gertner and Martin Rofheart. A parallel algorithm for 2-D DFT computation with no interprocessor communication. IEEE Transactions on Parallel and Distributed Systems, 1 (3):377–382, July 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Glinski:1994:SLR

[GR94] Stephen Glinski and David Roe.

Spoken language recognition on a DSP array processor. IEEE Transactions on Parallel and Distributed Systems, 5(7):697– 703, July 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gebali:2006:PAA

[GR06] Fayez Gebali and A. N. M. Ehtesham Rafiq. Processor array architectures for deep packet classification. IEEE Transactions on Parallel and Distributed Systems, 17(3):241–252, March 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Greer:1998:PMS

[Gre98] J. C. Greer. Parallelization model for successive approximations to the Rayleigh–Ritz linear variational problem. IEEE Transactions on Parallel and Distributed Systems, 9(10):938– ??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0938.
pdf;   http://www.computer.

org/tpds/td1998/l0938abs.
htm.

Gao:1999:OCT

[GRS99] L. Gao, A. L. Rosenberg, and R. K. Sitaraman. Optimal clustering of tree-sweep computations for high-latency parallel environments. IEEE Transactions on Parallel and Distributed Systems, 10(8):813–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0813.
pdf;   http://www.computer.
org/tpds/td1999/l0813abs.
htm.

Guinand:1997:WCA

[GRT97] F. Guinand, C. Rapine, and D. Trystram. Worst case analysis of Lawler’s algorithm for scheduling trees with communication delays. IEEE Transactions on Parallel and Distributed Systems, 8(10):1085–1086, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1085.
pdf;   http://www.computer.
org/tpds/td1997/l1085abs.
htm.

Ghaffar:2007:ONS

[GRY07] Akbar Ghaffar, Pour Rahbar, and Oliver Yang. OCGRR: a new scheduling algorithm for differentiated services networks. IEEE Transactions on Parallel and Distributed Systems, 18 (5):697–710, May 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:1991:CTT

[GS91] Rajiv Gupta and Mary Lou Soffa. Compile-time techniques for improving scalar access performance in parallel memories. IEEE Transactions on Parallel and Distributed Systems, 2 (2):138–148, April 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:1995:AEC

[GS95] Vipul Gupta and Eugen Schenfeld. Annealed embeddings of communication patterns in an interconnection cached network. IEEE Transactions on Parallel and Distributed Systems, 6 (11):1153–1167, November 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1153abs.htm.

Gupta:2003:ACS

[GS03] S. K. S. Gupta and P. K. Srimani. Adaptive core selection and migration method for multicast routing in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 14(1):27–38, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/01/l0027abs.
htm;   http://csdl.computer.

                                        org/dl/trans/td/2003/01/l0027.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/01/l0027.
                                        pdf.

Gong:2008:MPA

[GS08] Chao Gong and Kamil Sarac. A more practical approach for single-packet IP traceback using packet logging and marking. IEEE Transactions on Parallel and Distributed Systems, 19(10): 1310–1324, October 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:1996:UFO

[GSS96] Manish Gupta, Edith Schonberg, and Harini Srinivasan. A unified framework for optimizing communication in data-parallel programs. IEEE Transactions on Parallel and Distributed Systems, 7(7):689–704, July 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0689abs.htm.

Gupta:2006:CAI

[GSS06] Rohit Gupta, Varun Sekhri, and Arun K. Somani. CompuP2P: An architecture for Internet computing using peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 17(11):1306– 1320, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Glazer:1993:PML

[GT93] D. W. Glazer and C. Tropper. On process migration and load balancing in Time Warp. IEEE Transactions on Parallel and Distributed Systems, 4(3): 318–327, March 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gibaud:2002:CDB

[GT02] Alain Gibaud and Philippe Thomin. Communications directed by bound types in Linda: Presentation and formal model. IEEE Transactions on Parallel and Distributed Systems, 13(8):828–843, August 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0828.
pdf;   http://www.computer.
org/tpds/td2002/l0828abs.
htm.

Gupta:1992:SCC

[Gup92] Rajiv Gupta. Synchronization and communication costs of loop partitioning on shared-memory multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 3(4):505– 512, July 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gautama:2006:LCS

[GvG06] Hasyim Gautama and Arjan J. C. van Gemund. Lowcost static performance prediction of parallel stochastic task compositions. IEEE Transactions on Parallel and Distributed Systems, 17(1):78–91, January 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gonzalez:1995:EAH

[GVGD95] Antonio Gonzalez, Miguel Valero-Garcia, and Luis Diaz de Cerio. Executing algorithms with hypercube topology on torus multicomputers. IEEE Transactions on Parallel and Distributed Systems, 6(8):803–814, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0803abs.htm.

Goh:2009:DFE

[GVV09] Lee Kee Goh, Bharadwaj Veeravalli, and Sivakumar Viswanathan. Design of fast and efficient energy-aware gradient-based scheduling algorithms heterogeneous embedded multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 20(1):1–12, January 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Garg:1994:DWU

[GW94] Vijay K. Garg and Brian Waldecker. Detection of weak unstable predicates in distributed programs. IEEE Transactions on Parallel and Distributed Systems, 5(3):299–307, March 1994. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Ganapathy:1996:OSA

[GW96a] Kumar N. Ganapathy and Benjamin W. Wah. Optimal synthesis of algorithm-specific lowerdimensional processor arrays. IEEE Transactions on Parallel and Distributed Systems, 7 (3):274–287, March 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0274abs.htm.

Garg:1996:DSU

[GW96b] Vijay K. Garg and Brian Waldecker. Detection of strong unstable predicates in distributed programs. IEEE Transactions on Parallel and Distributed Systems, 7(12):1323– 1333, December 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2006:EAM

[GW06] Qian-Ping Gu and Yong Wang.

Efficient algorithms for minimum congestion hypergraph embedding in a cycle. IEEE Transactions on Parallel and Distributed Systems, 17(3):205–214, March 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ganapathy:1997:DSP

[GWL97] K. N. Ganapathy, B. W. Wah, and Chien-Wei Li. Designing a scalable processor array for recurrent computations. IEEE Transactions on Parallel and Distributed Systems, 8(8):840– 856, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0840.
pdf;   http://www.computer.
org/tpds/td1997/l0840abs.
htm.

Gu:2008:PPP

[GWYS08] Xiaohui Gu, Zhen Wen, Philip S. Yu, and Zon-Yin Shae. peerTalk: a peer-to-peer multiparty Voiceover-IP system. IEEE Transactions on Parallel and Distributed Systems, 19(4):515–528, April 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gerasoulis:1993:GCD

[GY93] Apostolos Gerasoulis and Tao Yang. On the granularity and clustering of directed acyclic task graphs. IEEE Transactions on Parallel and Distributed Systems, 4(6):686–701, June 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gaughan:1995:FFT

[GY95a] Patrick T. Gaughan and Sudhakar Yalamanchili. A family of fault-tolerant routing protocols for direct multiprocessor networks. IEEE Transactions on Parallel and Distributed Systems, 6(5):482–497, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0482abs.htm.

Ghosh:1995:ADA

[GY95b] Sumit Ghosh and Meng-Lin Yu.

An asynchronous distributed approach for the simulation of behavior-level models on parallel processors. IEEE Transactions on Parallel and Distributed Systems, 6(6):639–652, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0639abs.htm.

Guo:2007:LOD

[GY07] Song Guo and Oliver Yang.

Localized operations for distributed minimum energy multicast algorithm in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 18(2):186–198, February 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guclu:2009:LSF

[GY09] Hasan Guclu and Murat Yuksel. Limited scale-free overlay topologies for unstructured peerto-peer networks. IEEE Transactions on Parallel and Distributed Systems, 20(5):667–679, May 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2005:ISR

[GYS05] Song Guo, Oliver Yang, and Yantai Shu. Improving source routing reliability in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 16(4):362–373, April 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guerraoui:2003:GES

[GZ03] Rachid Guerraoui and Willy Zwaenepoel. Guest Editorial: Special section on middleware infrastructures. IEEE Transactions on Parallel and Distributed Systems, 14(11):1057– ??, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/11/
                                        l1057.pdf.

Gao:2006:LBS

[GZ06] Jie Gao and Li Zhang. Loadbalanced short-path routing in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 17(4):377–388, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2009:TOB

[GZ09] Jie Gao and Li Zhang. Tradeoffs between stretch factor and load-balancing ratio in routing on growth-restricted graphs. IEEE Transactions on Parallel and Distributed Systems, 20(2):

171–179, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:1995:DAC

[HALT95] Kaiyuan Huang, Vinod K. Agarwal, Laurence LaForge, and K. Thulasiraman. A diagnosis algorithm for constant degree structures and its application to VLSI circuit testing. IEEE Transactions on Parallel and Distributed Systems, 6 (4):363–372, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0363abs.htm.

Harper:1991:BMV

[Har91] David T. Harper III. Block, multistride vector, and FFT accesses in parallel memory systems. IEEE Transactions on Parallel and Distributed Systems, 2(1):43–51, January 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:1994:GAM

[HAR94] Edwin S. H. Hou, Nirwan Ansari, and Hong Ren. A genetic algorithm for multiprocessor scheduling. IEEE Transactions on Parallel and Distributed Systems, 5(2):113–120, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:1992:PMT

[HB92] Jiun-Ming Hsu and Prithviraj Banerjee. Performance measurement and trace driven simulation of parallel CAD and numeric applications on a hypercube multicomputer. IEEE Transactions on Parallel and Distributed Systems, 3(4):451– 464, July 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsiao:1992:PEC

[HC92] Shuo-Hsien Hsiao and C. Y. Roger Chen. Performance evaluation of circuit switched multistage interconnection networks using a hold strategy. IEEE Transactions on Parallel and Distributed Systems, 3(5):632–640, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hui:1997:ATI

[HC97] Chi-Chung Hui and S. T. Chanson. Allocating task interaction graphs to processors in heterogeneous networks. IEEE Transactions on Parallel and Distributed Systems, 8(9):908–925, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0908.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0908abs.
                                        htm.

Huang:1999:CPT

[HC99a] H.-L. Huang and G.-H. Chen.

Combinatorial properties of twolevel hypernet networks. IEEE Transactions on Parallel and Distributed Systems, 10(11):192– ??, November 1999. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l1192.
pdf;   http://www.computer.
org/tpds/td1999/l1192abs.
htm.

Hui:1999:HLB

[HC99b] C.-C. Hui and S. T. Chanson. Hydrodynamic load balancing. IEEE Transactions on Parallel and Distributed Systems, 10(11):118–??, November 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1118.
pdf;   http://www.computer.
org/tpds/td1999/l1118abs.
htm.

Hsieh:2009:SDR

[HC09] Sun-Yuan Hsieh and Tsung-Yen Chuang. The strong diagnosability of regular networks and product networks under the PMC model. IEEE Transactions on Parallel and Distributed Systems, 20(3):367–378, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2006:PCB

[HCC06] Xin-Mao Huang, Cheng-Yue Chang, and Ming-Syan Chen. PeerCluster: a cluster-based peer-to-peer system. IEEE Transactions on Parallel and Distributed Systems, 17(10): 1110–1123, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:1997:LRN

[HCD97] Wen-Jing Hsu, Moon Jung Chung, and Amitabha Das. Linear recursive networks and their applications in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 8(7):673–680, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0673.
pdf;   http://www.computer.
org/tpds/td1997/l0673abs.
htm.

Hogstedt:2003:PET

[HCF03] Karin Högstedt, Larry Carter, and Jeanne Ferrante. On the parallel execution time of tiled loops. IEEE Transactions on Parallel and Distributed Systems, 14(3):307–321, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0307abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0307.
htm;   http://csdl.computer.

org/dl/trans/td/2003/03/l0307.
pdf.

Hsieh:1999:FFH

[HCH99] S.-Y. Hsieh, G.-H. Chen, and C.-W. Ho. Fault-free Hamiltonian cycles in faulty arrangement graphs. IEEE Transactions on Parallel and Distributed Systems, 10(3):223–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0223.
pdf;   http://www.computer.
org/tpds/td1999/l0223abs.
htm.

He:2009:DVE

[HCHM09] Daojing He, Lin Cui, Hejiao Huang, and Maode Ma. Design and verification of enhanced secure localization scheme in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(7):1050– 1058, July 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsiao:1997:PEH

[HCY97] Hui-I Hsiao, Ming-Syan Chen, and P. S. Yu. Parallel execution of hash joins in parallel databases. IEEE Transactions on Parallel and Distributed Systems, 8(8):872–883, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0872.

pdf;   http://www.computer.
org/tpds/td1997/l0872abs.
htm.

Hsu:2001:GPM

[HCYD01] Ching-Hsien Hsu, Yeh-Ching Chung, Don-Lin Yang, and Chyi-Ren Dow. A generalized processor mapping technique for array redistribution. IEEE Transactions on Parallel and Distributed Systems, 12(7):743– 757, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0743.
pdf;   http://www.computer.
org/tpds/td2001/l0743abs.
htm.

Hsu:2006:OCD

[HCYL06] Ching-Hsien Hsu, Ming-Hao Chen, Chao-Tung Yang, and Kuan-Ching Li. Optimizing communications of dynamic data redistribution on symmetrical matrices in parallelizing compilers. IEEE Transactions on Parallel and Distributed Systems, 17(11):1226–1241, November 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2007:FTI

[HDF07] Rongsen He and José G. Delgado-Frias. Fault tolerant interleaved switching fabrics for scalable high-performance routers. IEEE Transactions on Parallel and Distributed Systems, 18(12):1727–1739, December 2007. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Hagin:2000:DMA

[HDRS00] A. Hagin, G. Dermler, K. Rothermel, and G. Shchemelev. Distributed multimedia application configuration management. IEEE Transactions on Parallel and Distributed Systems, 11(7):669–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0669.
pdf;   http://www.computer.
org/tpds/td2000/l0669abs.
htm.

Holliday:1992:AMR

[HE92] Mark A. Holliday and Carla S.

Ellis. Accuracy of memory reference traces of parallel computations in trace-driven simulation. IEEE Transactions on Parallel and Distributed Systems, 3 (1):97–109, January 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Herman:2000:PCT

[Her00] T. Herman. Phase clocks for transient fault repair. IEEE Transactions on Parallel and Distributed Systems, 11(10):48– ??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1048.
pdf;   http://www.computer.
org/tpds/td2000/l1048abs.
htm.

Hu:2005:NRC

[HGC05] Honghui Hu, Naijie Gu, and Jing Cao. A note on recursive cube of rings network. IEEE Transactions on Parallel and Distributed Systems, 16(10):1007–1008, October 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:1995:ECP

[HH95] Garng M. Huang and Wen-Lin Hsieh. Exact convergence of a parallel textured algorithm for data network optimal routing problems. IEEE Transactions on Parallel and Distributed Systems, 6(11):1132– 1146, November 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l1132abs.htm.

Hsiao:2008:TBP

[HH08] Hung-Chang Hsiao and Chih-Peng He. A tree-based peerto-peer network with quality guarantees. IEEE Transactions on Parallel and Distributed Systems, 19(8):1099–1110, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2008:PEO

[HHL08] Yuxiong He, Wen-Jing Hsu, and Charles E. Leiserson. Provably efficient online nonclairvoyant adaptive scheduling. IEEE Transactions on Parallel and Distributed Systems, 19(9):1263– 1279, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Helary:2000:CGF

[HHM^{+}00] J.-M. Hélary, M. Hurfin, A. Mostefaoui, M. Raynal, and F. Tronel. Computing global functions in asynchronous distributed systems with perfect failure detectors. IEEE Transactions on Parallel and Distributed Systems, 11(9):897–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0897.
pdf;   http://www.computer.
org/tpds/td2000/l0897abs.
htm.

Hancu:1994:CTA

[HISS94] Marius V. A. Hancu, Kazuhiko Iwasaki, Yuji Sato, and Mamoru Sugie. Concurrent test architecture for massively parallel computers and its error detection capability. IEEE Transactions on Parallel and Distributed Systems, 5(11):1169–1184, November 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hopkinson:2009:AGG

[HJB^{+}09] Kenneth Hopkinson, Kate Jenkins, Kenneth Birman, James Thorp, Gregory Toussaint, and Manu Parashar. Adaptive gravitational Gossip: a Gossipbased communication protocol with user-selectable rates. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1830–1843, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hwang:2002:OSM

[HJH02] Kai Hwang, Hai Jin, and Roy S. C. Ho. Orthogonal striping and mirroring in distributed RAID for I/O-centric cluster computing. IEEE Transactions on Parallel and Distributed Systems, 13(1):26–44, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0026.
pdf;   http://www.computer.
org/tpds/td2001/l0026abs.
htm.

He:2006:ANR

[HJS^{+}06] Ligang He, Stephen A. Jarvis, Daniel P. Spooner, Hong Jiang, Donna N. Dillenberger, and Graham R. Nudd. Allocating nonreal-time and soft real-time jobs in multiclusters. IEEE Transactions on Parallel and Distributed Systems, 17(2):99–112, February 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Havlak:1991:IIB

[HK91] Paul Havlak and Ken Kennedy. An implementation of interprocedural bounded regular section analysis. IEEE Transactions on Parallel and Distributed Systems, 2(3):350–360, July 1991.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:1993:PED

[HK93] Jau-Hsiung Huang and L. Kleinrock. Performance evaluation of dynamic sharing of processors in two-stage parallel processing systems. IEEE Transactions on Parallel and Distributed Systems, 4(3):306–317, March 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huisman:1994:HRS

[HK94] Leendert M. Huisman and Sandip Kundu. Highly reliable symmetric networks. IEEE Transactions on Parallel and Distributed Systems, 5(1):94–97, January 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ho:1995:OBA

[HK95] Ching-Tien Ho and Ming-Yang Kao. Optimal broadcast in all-port wormhole-routed hypercubes. IEEE Transactions on Parallel and Distributed Systems, 6(2):200–204, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0200abs.htm.

Hambrusch:1998:SPB

[HK98] S. E. Hambrusch and A. A.

Khokhar. Scalable S-to-P broadcasting on messagepassing MPPs. IEEE Transactions on Parallel and Distributed Systems, 9(8):758–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0758.
pdf;   http://www.computer.
org/tpds/td1998/l0758abs.
htm.

Higham:2006:TBC

[HK06] Lisa Higham and Jalal Kawash. Tight bounds for critical sections in processor consistent platforms. IEEE Transactions on Parallel and Distributed Systems, 17(10):1072– 1083, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/10/
l1072s.pdf.

Hamidzadeh:2000:DTS

[HKL00] B. Hamidzadeh, L. Y. Kit, and D. J. Lilja. Dynamic task scheduling using online optimization. IEEE Transactions on Parallel and Distributed Systems, 11(11):151–??, November 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1151.
pdf;   http://www.computer.
org/tpds/td2000/l1151abs.
htm.

Hofmann:1994:DPM

[HKM^{+}94] R. Hofmann, R. Klar, B. Mohr, A. Quick, and M. Siegle. Distributed performance monitoring: Methods, tools, and applications. IEEE Transactions on Parallel and Distributed Systems, 5(6):585–598, June 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huh:2007:NSF

[HKS^{+}07] Jaehyuk Huh, Changkyu Kim, Hazim Shafi, Lixin Zhang, Doug Burger, and Stephen W. Keckler. A NUCA substrate for flexible CMP cache sharing. IEEE Transactions on Parallel and Distributed Systems, 18(8): 1028–1040, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2008:RQO

[hKY08] Jong hwan Kim and Ikjun Yeom.

Reducing queue oscillation at a congested link. IEEE Transactions on Parallel and Distributed Systems, 19(3):394–407, March 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2008:MAM

[HL08] Jinsong Han and Yunhao Liu.

Mutual anonymity for mobile P2P systems. IEEE Transactions on Parallel and Distributed Systems, 19(8):1009–1019, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2009:VVO

[HL09a] Yuan He and Yunhao Liu. VOVO: VCR-oriented video-on-demand in large-scale peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 20(4):528–539, April 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsieh:2009:CEF

[HL09b] Sun-Yuan Hsieh and Chia-Wei Lee. Conditional edge-fault hamiltonicity of matching composition networks. IEEE Transactions on Parallel and Distributed Systems, 20(4):581–592, April 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2004:APS

[HLH04] Shing-Tsaan Huang, Tzong-Jye Liu, and Su-Shen Hung. Asynchronous phase synchronization in uniform unidirectional rings. IEEE Transactions on Parallel and Distributed Systems, 15(4):378–384, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/04/l0378abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/04/l0378.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/04/l0378.
                                        pdf.

Hsiao:2009:RTM

[HLH09] Hung-Chang Hsiao, Hao Liao, and Cheng-Chyun Huang. Resolving the topology mismatch problem in unstructured peerto-peer networks. IEEE Transactions on Parallel and Distributed Systems, 20(11):1668– 1681, November 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsiao:2009:BSW

[HLL09] Hung-Chang Hsiao, Yung-Chih Lin, and Hao Liao. Building small-world peer-to-peer networks based on hierarchical structures. IEEE Transactions on Parallel and Distributed Systems, 20(7):1023– 1037, July 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:1994:PDP

[HLV94] Shou-Hsuan Stephen Huang, Hongfei Liu, and Venkatraman Viswanathan. Parallel dynamic programming. IEEE Transactions on Parallel and Distributed Systems, 5(3):326–328, March 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Helmbold:1990:MGT

[HM90] David P. Helmbold and Charles E.

McDowell. Modeling speedup(n) greater than n. IEEE Transactions on Parallel and Distributed Systems, 1(2):250–256, April 1990. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Herlihy:1992:LFG

[HM92] M. P. Herlihy and J. E. B. Moss. Lock-free garbage collection for multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 3(3):304–311, May 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hong:1995:RSD

[HM95] Chul-Eui Hong and Bruce M. McMillin. Relaxing synchronization in distributed simulated annealing. IEEE Transactions on Parallel and Distributed Systems, 6(2):189–195, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0189abs.htm.

Harabagiu:1998:PST

[HM98] S. M. Harabagiu and D. I. Moldovan. A parallel system for text inference using marker propagations. IEEE Transactions on Parallel and Distributed Systems, 9(8):729–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0729.
pdf;   http://www.computer.
org/tpds/td1998/l0729abs.
htm.

Helary:1994:GST

[HMR94] J.-M. Helary, A. Mostefaoui, and M. Raynal. General scheme for token-and tree-based distributed mutual exclusion algorithms. IEEE Transactions on Parallel and Distributed Systems, 5(11):1185–1196, November 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Helary:1999:CID

[HMR99] J.-M. Hélary, A. Mostefaoui, and M. Raynal. Communicationinduced determination of consistent snapshots. IEEE Transactions on Parallel and Distributed Systems, 10(9):865–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0865.
pdf;   http://www.computer.
org/tpds/td1999/l0865abs.
htm.

Helmbold:1993:DPE

[HMW93] David P. Helmbold, Charles E.

McDowell, and Jian-Zhong Wang. Determining possible event orders by analyzing sequential traces. IEEE Transactions on Parallel and Distributed Systems, 4(7):827–840, July 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hendren:1990:PPR

[HN90] Laurie J. Hendren and Alexandru Nicolau. Parallelizing programs with recursive data structures. IEEE Transactions on Parallel and Distributed Systems, 1(1):35–47, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Heidelberger:1993:CPS

[HN93] Philip Heidelberger and David M. Nicol. Conservative parallel simulation of continuous time Markov chains using uniformization. IEEE Transactions on Parallel and Distributed Systems, 4 (8):906–921, August 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hayashi:1998:TAC

[HNO98a] T. Hayashi, K. Nakano, and S. Olariu. An O((log log n)^{2}) time algorithm to compute the convex hull of sorted points on reconfigurable meshes. IEEE Transactions on Parallel and Distributed Systems, 9(12):167– ??, December 1998. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1998/pdf/l1167.
pdf;   http://www.computer.
org/tpds/td1998/l1167abs.
htm.

Hayashi:1998:OPA

[HNO98b] T. Hayashi, K. Nakano, and S. Olariu. Optimal parallel algorithms for finding proximate points, with applications. IEEE Transactions on Parallel and Distributed Systems, 9(12):153–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1153.
pdf;   http://www.computer.
org/tpds/td1998/l1153abs.
htm.

Hayashi:1998:WTO

[HNO98c] T. Hayashi, K. Nakano, and S. Olariu. Work-time optimal k-merge algorithms on the PRAM. IEEE Transactions on Parallel and Distributed Systems, 9(3):275–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0275.
pdf;   http://www.computer.
org/tpds/td1998/l0275abs.
htm.

Hu:2002:RCR

[HNY02] Y. Hu, T. Nightingale, and Q. Yang. RAPID-cache — a reliable and inexpensive write cache for high performance storage systems. IEEE Transactions on Parallel and Distributed Systems, 13(3):290–307, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0290.
pdf;   http://www.computer.
org/tpds/td2002/l0290abs.
htm.

Hary:1999:PCT

[HÖ99] S. L. Hary and F. Özgüner. Precedence-constrained task allocation onto point-to-point networks for pipelined execution. IEEE Transactions on Parallel and Distributed Systems, 10(8):838–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0838.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0838abs.
                                        htm.

Halwan:2000:EHA

[HÖ00] V. Halwan and F. Özgüner. Efficient heuristics for all-port multicast in wormhole-routed hypercubes. IEEE Transactions on Parallel and Distributed Systems, 11(1):81–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0081.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0081abs.
                                        htm.

Halwan:1999:RWS

[HÖD99] V. Halwan, F. Özgüner, and A. Dogan. Routing in wormholeswitched clustered networks with applications to fault tolerance. IEEE Transactions on Parallel and Distributed Systems, 10(10):1–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1001.
pdf;   http://www.computer.
org/tpds/td1999/l1001abs.
htm.

Hollingsworth:1998:CPP

[Hol98]  J. K. Hollingsworth.  Crit-

ical path profiling of message passing and shared-memory programs. IEEE Transactions on Parallel and Distributed Systems, 9(10):29–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1029.
pdf;   http://www.computer.
org/tpds/td1998/l1029abs.
htm.

Ho:2003:CAI

[HP03] Wai Hong Ho and Timothy Mark Pinkston. A clustering approach for identifying and quantifying irregularities in interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 14(12):1222–1239, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1222abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1222.
pdf.

Ho:2006:DME

[HP06] Wai Hong Ho and Timothy Mark Pinkston. A design methodology for efficient application-specific on-chip interconnects. IEEE Transactions on Parallel and Distributed Systems, 17(2):174– 190, February 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/;
                                         http://csdl.computer.org/
                                        comp/trans/td/2006/02/l0174s.
                                        pdf.

Hong:2007:AAI

[HP07] Bo Hong and Viktor Prasanna. Adaptive allocation of independent tasks to maximize throughput. IEEE Transactions on Parallel and Distributed Systems, 18 (10):1420–1435, October 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2008:AGA

[HPH08] Yu-Kai Huang, Ai-Chun Pang, and Hui-Nien Hung. An adaptive GTS allocation scheme for IEEE 802.15.4. IEEE Transactions on Parallel and Distributed Systems, 19(5):641–651, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:2004:GMM

[HPT04] Y. Thomas Hou, Shivendra S. Panwar, and Henry H.-Y. Tzeng. On generalized max-min rate allocation and distributed convergence algorithm for packet networks. IEEE Transactions on Parallel and Distributed Systems, 15(5):401–416, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/05/l0401abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0401.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0401.
pdf.

Hatcher:1991:DPP

[HQL^{+}91] Philip J. Hatcher, Michael J.

Quinn, Anthony J. Lapadula, Bradley K. Seevers, Ray J. Anderson, and Robert R. Jones. Data-parallel programming on MIMD computers. IEEE Transactions on Parallel and Distributed Systems, 2(3):377–383, July 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Haritsa:2000:PRT

[HRG00] J. R. Haritsa, K. Ramamritham, and R. Gupta. The PROMPT real-time commit protocol. IEEE Transactions on Parallel and Distributed Systems, 11(2):160–??, February 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0160.
pdf;   http://www.computer.
org/tpds/td2000/l0160abs.
htm.

Hamdi:1997:EHH

[HS97] M. Hamdi and S. W. Song.

Embedding hierarchical hypercube networks into the hypercube. IEEE Transactions on Parallel and Distributed Systems, 8(9):897–902, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0897.
pdf;   http://www.computer.
org/tpds/td1997/l0897abs.
htm.

Hodzic:1998:STM

[HS98a] E. Hodzic and W. Shang. On supernode transformation with minimized total running time. IEEE Transactions on Parallel and Distributed Systems, 9(5):417–??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0417.
pdf;   http://www.computer.
org/tpds/td1998/l0417abs.
htm.

Huang:1998:CCL

[HS98b] T.-L. Huang and C. H. Shann. A comment on “A Circular List-Based Mutual Exclusion Scheme for Large Shared-Memory Multiprocessors”. IEEE Transactions on Parallel and Distributed Systems, 9(4):415–??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0415.
pdf;   http://www.computer.

org/tpds/td1998/l0415abs.

htm. See [FT97].

Han:1999:EEB

[HS99a] S. Han and K. G. Shin. Experimental evaluation of behaviorbased failure-detection schemes in real-time communication networks. IEEE Transactions on Parallel and Distributed Systems, 10(6):613–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0613.
pdf;   http://www.computer.
org/tpds/td1999/l0613abs.
htm.

Humphrey:1999:PTD

[HS99b] M. Humphrey and J. A.

Stankovic. Predictable threads for dynamic, hard real-time environments. IEEE Transactions on Parallel and Distributed Systems, 10(3):281–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0281.
pdf;   http://www.computer.
org/tpds/td1999/l0281abs.
htm.

Hodzic:2002:TOS

[HS02] Edin Hodzic and Weijia Shang.

On time optimal supernode shape. IEEE Transactions on Parallel and Distributed Systems, 13(12):1220–1233, December 2002. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2002/12/l1220abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/12/l1220.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/12/l1220.
                                        pdf.

Hosaagrahara:2008:MMF

[HS08] Madhusudan Hosaagrahara and Harish Sethu. Max-min fair scheduling in input-queued switches. IEEE Transactions on Parallel and Distributed Systems, 19(4):462–475, April 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hiltunen:1999:RTD

[HSH^{+}99] M. A. Hiltunen, R. D. Schlichting, X. Han, M. M. Cardozo, and R. Das. Real-time dependable channels: Customizing QoS attributes for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 10(6):600–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0600.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0600abs.
                                        htm.

Hsieh:2003:SFP

[Hsi03] Sun-Yuan Hsieh. A simple and fast parallel coloring algorithm for distance-hereditary graphs. IEEE Transactions on Parallel and Distributed Systems, 14(12):1201–1208, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1201abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1201.
pdf.

He:2005:SCP

[HSLA05] Tian He, John A. Stankovic, Chenyang Lu, and Tarek F. Abdelzaher. A spatiotemporal communication protocol for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 16(10):995–1006, October 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:1993:FCN

[Hsu93] Wen-Jing Hsu. Fibonacci cubes:

a new interconnection topology. IEEE Transactions on Parallel and Distributed Systems, 4(1):3– 12, January 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hawkins:2007:DVA

[HSWB07] Cory Hawkins, Benjamin A.

Small, D. Scott Wills, and Keren Bergman. The Data Vortex, an all optical path multicomputer interconnection network. IEEE Transactions on Parallel and Distributed Systems, 18(3): 409–420, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2006:ELI

[HT06] Hwansoo Han and Chau-Wen Tseng. Exploiting locality for irregular scientific codes. IEEE Transactions on Parallel and Distributed Systems, 17(7):606– 618, July 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:2007:LDM

[HT07] Guo-Huang Hsu and Jimmy J. M. Tan. A local diagnosability measure for multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 18(5):598–607, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Horng:2002:OAC

[HTPS02] Shi-Jinn Horng, Horng-Ren Tsai, Yi Pan, and Jennifer Seitzer. Optimal algorithms for the channel-assignment problem on a reconfigurable array of processors with wider bus networks. IEEE Transactions on Parallel and Distributed Systems, 13(11):1124–1138, November 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2002/11/l1124abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1124.
                                        htm;   http://csdl.computer.

org/dl/trans/td/2002/11/l1124.
pdf.

Han:2007:TAB

[HV07] Qi Han and Nalini Venkatasubramanian. Timeliness-accuracy balanced collection of dynamic context data. IEEE Transactions on Parallel and Distributed Systems, 18(2):158–171, February 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Herlihy:1991:SGD

[HW91] Maurice P. Herlihy and Jeannette M. Wing. Specifying graceful degradation. IEEE Transactions on Parallel and Distributed Systems, 2(1):93–104, January 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:1997:ERS

[HW97] D. F. Hsu and D. S. L. Wei.

Efficient routing and sorting schemes for de Bruijn networks. IEEE Transactions on Parallel and Distributed Systems, 8(11):1157–1170, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1157.
pdf;   http://www.computer.
org/tpds/td1997/l1157abs.
htm.

Huang:2008:EDT

[HW08] Xin Huang and Tilman Wolf.

Evaluating dynamic task mapping in network processor runtime systems. IEEE Transactions on Parallel and Distributed Systems, 19(8):1086–1098, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:2001:OPM

[HWKH01] Y. Hou, C.-M. Wang, C.-Y. Ku, and L.-H. Hsu. Optimal processor mapping for linear-complement communication on hypercubes. IEEE Transactions on Parallel and Distributed Systems, 12(5):514– ??, May 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0514.
pdf;   http://www.computer.
org/tpds/td2001/l0514abs.
htm.

Hung:2000:IBC

[HWSH00] C.-N. Hung, J.-J. Wang, T.- Y. Sung, and L.-H. Hsu. On the isomorphism between cycliccubes and wrapped butterfly networks. IEEE Transactions on Parallel and Distributed Systems, 11(8):864–??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0864.
pdf;   http://www.computer.
org/tpds/td2000/l0864abs.
htm.

Hwang:1999:RSE

[HWWX99] K. Hwang, C. Wang, C.-L.

Wang, and Z. Xu. Resource scaling effects on MPP performance: The STAP benchmark implications. IEEE Transactions on Parallel and Distributed Systems, 10(5):509–??, May 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0509.
pdf;   http://www.computer.
org/tpds/td1999/l0509abs.
htm.

Hwang:1996:BEI

[HXA96] Kai Hwang, Zhiwei Xu, and Masahiro Arakawa. Benchmark evaluation of the IBM SP2 for parallel signal processing. IEEE Transactions on Parallel and Distributed Systems, 7(5):522– 536 (or 522–535??), May 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0522abs.htm.

Hurley:1996:FMF

[HY96] Richard T. Hurley and Soon Aun Yeap. File migration and file replication: a symbiotic relationship. IEEE Transactions on Parallel and Distributed Systems, 7(6):578–586, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0578abs.htm.

Harada:1997:NCG

[HY97] Takashi Harada and Masafumi Yamashita. Nondominated coteries on graphs. IEEE Transactions on Parallel and Distributed Systems, 8(6):667–672, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0667.
pdf;   http://www.computer.
org/tpds/td1997/l0667abs.
htm.

Haynos:1999:AMB

[HY99] M. P. Haynos and Y. Yang. An analytical model on the blocking probability of a faulttolerant network. IEEE Transactions on Parallel and Distributed Systems, 10(10):40–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1040.
pdf;   http://www.computer.
org/tpds/td1999/l1040abs.
htm.

Harada:2001:CJO

[HY01] Takashi Harada and Masafumi Yamashita. Coterie join operation and tree structured kcoteries. IEEE Transactions on Parallel and Distributed Systems, 12(9):865–874, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0865.
pdf;   http://www.computer.
org/tpds/td2001/l0865abs.
htm.

Harada:2004:CTN

[HY04] Takashi Harada and Masafumi Yamashita. k-coteries for tolerating network 2-partition. IEEE Transactions on Parallel and Distributed Systems, 15 (7):666–672, July 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/07/l0666.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/07/l0666.pdf.

Harada:2005:TMO

[HY05] Takashi Harada and Masafumi Yamashita. Transversal merge operation: a nondominated coterie construction method for distributed mutual exclusion. IEEE Transactions on Parallel and Distributed Systems, 16(2): 183–192, February 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hong:2007:DAP

[HY07] Yang Hong and Oliver W. W.

Yang. Design of adaptive PI rate controller for best-effort traffic in the Internet based on phase margin. IEEE Transactions on Parallel and Distributed Systems, 18 (4):550–561, April 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:2002:IPC

[HYP02] Jiongkuan Hou, Jie Yang, and Symeon Papavassiliou. Integration of pricing with call admission control to meet QoS requirements in cellular networks. IEEE Transactions on Parallel and Distributed Systems, 13(9):898–910, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0898.
pdf;   http://www.computer.
org/tpds/td2002/l0898abs.
htm.

Huang:1996:NAO

[HZ96] Garng M. Huang and Shan Zhu. A new HAD algorithm for optimal routing of hierarchically structured data networks. IEEE Transactions on Parallel and Distributed Systems, 7(9):939– 953, September 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0939abs.htm.

Hwang:1997:DCP

[HZ97] F. K. Hwang and Wenan Zang. Detecting corrupted pages in M replicated large files. IEEE Transactions on Parallel and Distributed Systems, 8(12):1241–1245, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1241.
pdf;   http://www.computer.
org/tpds/td1997/l1241abs.
htm.

Ibaroudene:1995:PDO

[IA95] Djaffer Ibaroudene and Raj Acharya. Parallel display of objects represented by linear octrees. IEEE Transactions on Parallel and Distributed Systems, 6(1):79–85, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iannello:1997:EAR

[Ian97] G. Iannello. Efficient algorithms for the reduce-scatter operation in LogGP. IEEE Transactions on Parallel and Distributed Systems, 8(9):970–982, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0970.
pdf;   http://www.computer.
org/tpds/td1997/l0970abs.
htm.

Iqbal:1995:EAC

[IB95] M. Ashraf Iqbal and Shahid H.

Bokhari. Efficient algorithms for a class of partitioning problems. IEEE Transactions on Parallel and Distributed Systems, 6(2): 170–175, February 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0170abs.htm.

Islam:1992:DCS

[IC92] N. Islam and R. H. Campbell. Design considerations for shared memory multiprocessor message systems. IEEE Transactions on Parallel and Distributed Systems, 3(6):702–711, November 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ibe:1993:PEC

[ICT93] Oliver C. Ibe, Hoon Choi, and Kishor S. Trivedi. Performance evaluation of client-server systems. IEEE Transactions on Parallel and Distributed Systems, 4(11):1217–1229, November 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ibaraki:1993:TCM

[IK93] Toshihide Ibaraki and Tiko Kameda. Theory of coteries: mutual exclusion in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 4(7):779–794, July 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ikeda:2002:FCT

[IKOY02] Satoshi Ikeda, Izumi Kubo, Norihiro Okumoto, and Masafumi Yamashita. Fair circulation of a token. IEEE Transactions on Parallel and Distributed Systems, 13(4):367–372, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0367.
pdf;   http://www.computer.

org/tpds/td2002/l0367abs.
htm.

Ip:2007:CAC

[ILL07] Alan T. S. Ip, Jiangchuan Liu, and John Chi-Shing Lui. CO-PACC: An architecture of cooperative proxy-client caching system for on-demand media streaming. IEEE Transactions on Parallel and Distributed Systems, 18(1):70–83, January 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ibarra:1990:MSA

[IS90] Oscar H. Ibarra and Stephen M.

Sohn. On mapping systolic algorithms onto the hypercube. IEEE Transactions on Parallel and Distributed Systems, 1 (1):48–63, January 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Imani:2009:DTS

[ISAZM09] Navid Imani, Hamid Sarbazi-Azad, Albert Y. Zomaya, and Parya Moinzadeh. Detecting threats in star graphs. IEEE Transactions on Parallel and Distributed Systems, 20(4):474– 483, April 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ingelrest:2006:OTR

[ISRS06] Francois Ingelrest, David Simplot-Ryl, and Ivan Stojmenovic. Optimal transmission radius for energy efficient broadcasting protocols in ad hoc and sensor networks. IEEE Transactions on Parallel and Distributed Systems, 17(6):536–547, June 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Imai:1993:EPC

[IT93] Akira Imai and Evan Tick. Evaluation of parallel copying garbage collection on a sharedmemory multiprocessor. IEEE Transactions on Parallel and Distributed Systems, 4(9):1030– 1040, September 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iyer:2007:ESS

[IT07] Ravi Iyer and Dean M. Tullsen. Editorial: Special section on CMP architectures. IEEE Transactions on Parallel and Distributed Systems, 18(8):1025– 1027, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2007/08/
l1025.pdf.

Joshi:2008:SSA

[JASA08] Tarun Joshi, Disha Ahuja, Damanjit Singh, and Dharma P. Agrawal. SARA: Stochastic Automata Rate Adaptation for IEEE 802.11 networks. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1579–1590, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Joisha:2001:ECO

[JB01] Pramod G. Joisha and Prithviraj Banerjee. The efficient computation of ownership sets in HPF. IEEE Transactions on Parallel and Distributed Systems, 12(8):769–788, August 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0769.
pdf;   http://www.computer.
org/tpds/td2001/l0769abs.
htm.

Jiang:2008:TWB

[JBW^{+}08] Xuxian Jiang, Florian Buchholz, Aaron Walters, Dongyan Xu, Yi-Min Wang, and Eugene H. Spafford. Tracing worm break-in and contaminations via process coloring: a provenance-preserving approach. IEEE Transactions on Parallel and Distributed Systems, 19(7):890–902, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jung:2007:ODC

[JEG07] Eunjin (EJ) Jung, Ehab S.

Elmallah, and Mohamed G. Gouda. Optimal dispersal of certificate chains. IEEE Transactions on Parallel and Distributed Systems, 18(4):474–484, April 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Janssens:1994:PCB

[JF94] Bob Janssens and W. Kent Fuchs. Performance of cachebased error recovery in multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 5(10):1033–1043, October 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jain:2008:DMA

[JGA08] Vivek Jain, Anurag Gupta, and Dharma P. Agrawal. Ondemand medium access in multihop wireless networks with multiple beam smart antennas. IEEE Transactions on Parallel and Distributed Systems, 19 (4):489–502, April 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2008:LMR

[JGZW08] Song Jiang, Lei Guo, Xiaodong Zhang, and Haodong Wang. LightFlood: Minimizing redundant messages and maximizing scope of peer-to-peer search. IEEE Transactions on Parallel and Distributed Systems, 19 (5):601–614, May 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Johnson:1997:PMS

[JH97] T. Johnson and K. Harathi. A prioritized multiprocessor spin lock. IEEE Transactions on Parallel and Distributed Systems, 8(9):926–933, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

td/books/td1997/pdf/l0926.
pdf;   http://www.computer.
org/tpds/td1997/l0926abs.
htm.

Jiang:1997:EGF

[JHK97] Feng-Shr Jiang, Shi-Jinn Horng, and Tzong-Wann Kao. Embedding of generalized Fibonacci cubes in hypercubes with faulty nodes. IEEE Transactions on Parallel and Distributed Systems, 8(7):727–737, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0727.
pdf;   http://www.computer.
org/tpds/td1997/l0727abs.
htm.

Jia:1995:TOM

[Jia95] Xiaohua Jia. A total ordering multicast protocol using propagation trees. IEEE Transactions on Parallel and Distributed Systems, 6(6):617–627, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0617abs.htm. See comment

[SH97].

Jeng:2007:NGA

[JJ07] Andy An-Kai Jeng and Rong-Hong Jan. The r-neighborhood graph: An adjustable structure for topology control in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 18(4):536–549, April 2007. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2009:CRN

[JJ09] Yichuan Jiang and Jiuchuan Jiang. Contextual resource negotiation-based task allocation and load balancing in complex software systems. IEEE Transactions on Parallel and Distributed Systems, 20(5):641– 653, May 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jorgensen:1999:CAV

[JK99] J. B. Jørgensen and L. M. Kristensen. Computer aided verification of Lamport’s fast mutual exclusion algorithm using colored Petri nets and occurrence graphs with symmetries. IEEE Transactions on Parallel and Distributed Systems, 10(7):714– ??, July 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0714.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0714abs.
                                        htm.

Jouraku:2007:EDD

[JKA07] Akiya Jouraku, Michihiro Koibuchi, and Hideharu Amano. An effective design of deadlock-free routing algorithms based on 2D turn model for irregular networks. IEEE Transactions on Parallel and Distributed Systems, 18 (3):320–333, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Johnson:2001:DPA

[JKR01] Binu K. Johnson, R. Karthikeyan, and D. Janaki Ram. DP: a paradigm for anonymous remote computation and communication for cluster computing. IEEE Transactions on Parallel and Distributed Systems, 12(10):1052–1065, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1052.
pdf;   http://www.computer.
org/tpds/td2001/l1052abs.
htm.

Jacob:1999:TSS

[JL99] J. C. Jacob and S.-Y. Lee. Task spreading and shrinking on multiprocessor systems and networks of workstations. IEEE Transactions on Parallel and Distributed Systems, 10(10):82– ??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1082.
pdf;   http://www.computer.
org/tpds/td1999/l1082abs.
htm.

Jia:2005:ORD

[JLDC05] Xiaohua Jia, Deying Li, Hongwei Du, and Jinli Cao. On optimal replication of data object at hierarchical and transparent Web proxies. IEEE Transactions on Parallel and Distributed Systems, 16(8):673–685, August 2005. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Jimenez:2003:CEI

[JLF03] Marta Jiménez, José M. Llabería, and Agustín Fernández. A costeffective implementation of multilevel tiling. IEEE Transactions on Parallel and Distributed Systems, 14(10):1006–1020, October 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/10/l1006abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/10/l1006.
                                        pdf.

Jiang:2002:DDC

[JLS02] Jianping Jiang, Ten-Hwang Lai, and Neelam Soundarajan. On distributed dynamic channel allocation in mobile cellular networks. IEEE Transactions on Parallel and Distributed Systems, 13(10):1024–1037, October 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l1024.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l1024abs.
                                        htm.

Jin:2008:TSW

[JN08] Guang Jin and Silvia Nittel. Toward spatial window queries over continuous phenomena in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(4):559–571, April 2008.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jafari:2006:AEF

[JNGS06] Roozbeh Jafari, Hyduke Noshadi, Soheil Ghiasi, and Majid Sarrafzadeh. Adaptive electrocardiogram feature extraction on distributed embedded systems. IEEE Transactions on Parallel and Distributed Systems, 17 (8):797–807, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jeong:1995:IFP

[JO95] Byeong-Soo Jeong and Edward Omiecinski. Inverted file partitioning schemes in multiple disk systems. IEEE Transactions on Parallel and Distributed Systems, 6(2):142–153, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0142abs.htm.

Joung:2003:QBA

[Jou03]  Yuh-Jzer Joung.  Quorum-

based algorithms for group mutual exclusion. IEEE Transactions on Parallel and Distributed Systems, 14(5):463–476, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/05/l0463abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0463.
htm;   http://csdl.computer.

                                        org/dl/trans/td/2003/05/l0463.
                                        pdf.

JaJa:1993:OAP

[JR93] Joseph JáJá and Kwan Woo Ryu. Optimal algorithms on the pipelined hypercube and related networks. IEEE Transactions on Parallel and Distributed Systems, 4(5):582–591, May 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jain:1994:LUB

[JR94] Kamal Kumar Jain and V. Rajaraman. Lower and upper bounds on time for multiprocessor optimal schedules. IEEE Transactions on Parallel and Distributed Systems, 5(8):879– 886, August 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

JaJa:1996:BDM

[JR96] Joseph F. JáJá and Kwan Woo Ryu. The block distributed memory model. IEEE Transactions on Parallel and Distributed Systems, 7(8):830–840, August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1996/
                                        l0830abs.htm.

Jackson:2003:OQP

[JR03] Laura E. Jackson and George N. Rouskas. Optimal quantization of periodic task requests on multiple identical processors. IEEE Transactions on Parallel and Distributed Systems, 14(8):795–806, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0795abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/08/l0795.
pdf.

Jeng:1990:DMS

[JS90] Menkae Jeng and Howard Jay Siegel. Distributed management scheme for partitionable parallel computers. IEEE Transactions on Parallel and Distributed Systems, 1(1):120–126, January 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jenq:1993:ISE

[JS93] Jing-Fu Jenq and Sartaj Sahni.

Image shrinking and expanding on a pyramid. IEEE Transactions on Parallel and Distributed Systems, 4(11):1291– 1296, November 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Joung:1998:SIF

[JS98] Y.-J. Joung and S. A. Smolka.

Strong interaction fairness via randomization. IEEE Transactions on Parallel and Distributed Systems, 9(2):137–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0137.

                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0137abs.
                                        htm.

Juurlink:1998:GMT

[JSR98] B. H. H. Juurlink, J. F. Sibeyn, and P. S. Rao. Gossiping on meshes and tori. IEEE Transactions on Parallel and Distributed Systems, 9(6):513–??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0513.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0513abs.
                                        htm.

Jain:1997:HSO

[JSWB97] Ravi Jain, Kiran Somalwar, John Werth, and J. C. Browne. Heuristics for scheduling I/O operations. IEEE Transactions on Parallel and Distributed Systems, 8(3):310–320, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0310.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0310abs.
                                        htm.

Jin:2008:SEE

[JTC08] Xing Jin, Wanqing Tu, and S.- H. Gary Chan. Scalable and efficient end-to-end network topology inference. IEEE Transactions on Parallel and Distributed Systems, 19(6):837–850, June 2008. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsai:1996:BAA

[jTM96] Yih jia Tsai and Philip K.

McKinley. A broadcast algorithm for all-port wormholerouted torus networks. IEEE Transactions on Parallel and Distributed Systems, 7(8):876– 885, August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1996/l0876abs.
htm.

Tsai:1997:EDN

[jTM97] Yih jia Tsai and Philip K.

McKinley. An extended dominating node approach to broadcast and global combine in multiport wormhole-routed mesh networks. IEEE Transactions on Parallel and Distributed Systems, 8(1):41–58, January 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2008:EPI

[JTP^{+}08] Karl Jiang, Øystein Thorsen, Amanda Peters, Brian Smith, and Carlos P. Sosa. An efficient parallel implementation of the hidden Markov methods for genomic sequence-search on a massively parallel system. IEEE Transactions on Parallel and Distributed Systems, 19 (1):15–23, January 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2005:CMO

[JVYA05] Ruoming Jin, Karthikeyan Vaidyanathan, Ge Yang, and Gagan Agrawal. Communication and memory optimal parallel data cube construction. IEEE Transactions on Parallel and Distributed Systems, 16(12): 1105–1119, December 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jogalekar:2000:ESD

[JW00] P. Jogalekar and M. Woodside. Evaluating the scalability of distributed systems. IEEE Transactions on Parallel and Distributed Systems, 11(6):589– ??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0589.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0589abs.
                                        htm.

Ju:1994:CFP

[JWC94] Dz-Ching R. Ju, Chuan-Lin Wu, and Paul Carini. Classification, fusion and parallelization of array language primitives. IEEE Transactions on Parallel and Distributed Systems, 5(10): 1113–1120, October 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jia:2004:DAC

[JXT^{+}04] Weijia Jia, Dong Xuan, Wanqing Tu, Lidong Lin, and Wei Zhao. Distributed admission control for anycast flows.

IEEE Transactions on Parallel and Distributed Systems, 15 (8):673–686, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/08/l0673.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/08/l0673.pdf.

Jin:2004:SPA

[JZ04] Dejiang Jin and Sotirios G. Ziavras. A super-programming approach for mining association rules in parallel on PC clusters. IEEE Transactions on Parallel and Distributed Systems, 15(9):783–794, September 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/09/l0783.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/09/l0783.pdf.

Jia:1999:EFT

[JZXX99] W. Jia, W. Zhao, D. Xuan, and G. Xu. An efficient fault-tolerant multicast routing protocol with core-based tree techniques. IEEE Transactions on Parallel and Distributed Systems, 10(10):984–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0984.
pdf;   http://www.computer.
org/tpds/td1999/l0984abs.
htm.

Kieckhafer:1994:RAA

[KA94] R. M. Kieckhafer and M. H. Azadmanesh. Reaching approximate agreement with mixedmode faults. IEEE Transactions on Parallel and Distributed Systems, 5(1):53–63, January 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kwok:1996:DCP

[KA96] Yu-Kwong Kwok and Ishfaq Ahmad. Dynamic criticalpath scheduling: An effective technique for allocating task graphs to multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 7 (5):506–521, May 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0506abs.htm.

Kwok:1999:FPL

[KA99] Y.-K. Kwok and I. Ahmad. FASTEST: a practical lowcomplexity algorithm for compiletime assignment of parallel programs to multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 10(2):147–??, February 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0147.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0147abs.
                                        htm.

Kistler:2005:IPS

[KA05] Michael Kistler and Lorenzo Alvisi. Improving the performance of software distributed shared memory with speculation. IEEE Transactions on Parallel and Distributed Systems, 16(9):885–896, September 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kaya:2006:IIB

[KA06] Kamer Kaya and Cevdet Aykanat.

Iterative-improvement-based heuristics for adaptive scheduling of tasks sharing files on heterogeneous master-slave environments. IEEE Transactions on Parallel and Distributed Systems, 17(8):883– 896, August 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://bell.computer.
org/dlcomments/.

Khan:2009:CGT

[KA09] Samee U. Khan and Ishfaq Ahmad. A cooperative game theoretical technique for joint optimization of energy consumption and response time in computational grids. IEEE Transactions on Parallel and Distributed Systems, 20(3):346–360, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kalyanaraman:2003:STE

[KABK03] Anantharaman Kalyanaraman, Srinivas Aluru, Volker Brendel, and Suresh Kothari. Space and time efficient parallel algorithms and software for EST clustering. IEEE Transactions on Parallel and Distributed Systems, 14(12):1209–1221, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1209abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1209.
pdf.

Kandemir:2001:CDC

[Kan01]  M. Kandemir.  Compiler-

directed collective-I/O. IEEE Transactions on Parallel and Distributed Systems, 12(12): 1318–1331, December 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l1318.
pdf;   http://www.computer.
org/tpds/td2001/l1318abs.
htm.

Karaata:2001:SSS

[Kar01] M. H. Karaata. Self-stabilizing strong fairness under weak fairness. IEEE Transactions on Parallel and Distributed Systems, 12(4):337–??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0337.
pdf;   http://www.computer.
org/tpds/td2001/l0337abs.
htm.

Koibuchi:2006:SDT

[KAY^{+}06] Michihiro Koibuchi, Kenichiro Anjo, Yutaka Yamada, Akiya Jouraku, and Hideharu Amano. A simple data transfer technique using local address for networkson-chips. IEEE Transactions on Parallel and Distributed Systems, 17(12):1425–1437, December 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Keren:2003:OCA

[KB03] Arie Keren and Amnon Barak.

Opportunity cost algorithms for reduction of I/O and interprocess communication overhead in a computing cluster. IEEE Transactions on Parallel and Distributed Systems, 14(1):39–50, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/01/l0039abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0039.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0039.
pdf.

Kianzad:2006:ETC

[KB06] Vida Kianzad and Shuvra S.

Bhattacharyya. Efficient techniques for clustering and scheduling onto embedded multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 17 (7):667–680, July 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kandemir:2001:SDL

[KBC^{+}01] Mahmut Kandemir, Prithviraj Banerjee, Alok Choudhary, J. Ramanujam, and Eduard Ayguadé. Static and dynamic locality optimizations using integer linear programming. IEEE Transactions on Parallel and Distributed Systems, 12(9):922–941, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0922.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0922abs.
                                        htm.

Kurzak:2008:SSL

[KBD08] Jakub Kurzak, Alfredo Buttari, and Jack Dongarra. Solving systems of linear equations on the CELL processor using Cholesky factorization. IEEE Transactions on Parallel and Distributed Systems, 19(9):1175– 1186, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kwon:1998:ASJ

[KC98] O.-H. Kwon and K.-Y. Chwa. An algorithm for scheduling jobs in hypercube systems. IEEE Transactions on Parallel and Distributed Systems, 9(9):856–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0856.

pdf;   http://www.computer.
org/tpds/td1998/l0856abs.
htm.

Kim:2007:SBE

[KCD07] Jin-Ha Kim, Gyu Sang Choi, and Chita R. Das. An SSL back-end forwarding scheme in cluster-based Web servers. IEEE Transactions on Parallel and Distributed Systems, 18(7):946– 957, July 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kuo:2006:COR

[KCK^{+}06] Chin-Fu Kuo, Ya-Shu Chen, Tei-Wei Kuo, Phone Lin, and Cheng Chang. Component-oriented radars with probabilistic timing guarantees. IEEE Transactions on Parallel and Distributed Systems, 17(7):723–735, July 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

King:1990:PDPa

[KCN90a] Chung-Ta King, Wen-Hwa Chou, and Lionel M. Ni. Pipelined data parallel algorithms — I: Concept and modeling. IEEE Transactions on Parallel and Distributed Systems, 1 (4):470–485, October 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

King:1990:PDPb

[KCN90b] Chung-Ta King, Wen-Hwa Chou, and Lionel M. Ni.

Pipelined data parallel algorithms — II: Design. IEEE Transactions on Parallel and Distributed Systems, 1(4):486– 499, October 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koufaty:1996:DFS

[KCPT96] David A. Koufaty, Xiangfeng Chen, David K. Poulsen, and Josep Torrellas. Data forwarding in scalable shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 7(12):1250–1264, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kandemir:2003:RFS

[KCRB03] Mahmut Kandemir, Alok Choudhary, J. Ramanujam, and Prith Banerjee. Reducing false sharing and improving spatial locality in a unified compilation framework. IEEE Transactions on Parallel and Distributed Systems, 14(4):337–354, April 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0337abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0337.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0337.
pdf.

Kandemir:2000:UFO

[KCRK00] M. Kandemir, A. Choudhary, J. Ramanujam, and M. A. Kandaswamy. A unified framework for optimizing locality, parallelism, and communication in out-of-core computations. IEEE Transactions on Parallel and Distributed Systems, 11(7):648– ??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0648.
pdf;   http://www.computer.
org/tpds/td2000/l0648abs.
htm.

Kandemir:1999:LAF

[KCS^{+}99] M. Kandemir, A. Choudhary, N. Shenoy, P. Banerjee, and J. Ramanujam. A linear algebra framework for automatic determination of optimal data layouts. IEEE Transactions on Parallel and Distributed Systems, 10(2):115–??, February 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0115.
pdf;   http://www.computer.
org/tpds/td1999/l0115abs.
htm.

Kim:2009:UDA

[KCW09] Jinoh Kim, Abhishek Chandra, and Jon B. Weissman. Using data accessibility for resource selection in large-scale distributed systems. IEEE Transactions on Parallel and Distributed Systems, 20(6):788–801, June 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:1991:TPA

[KDL91] Jong Kim, Chita R. Das, and Woei Lin. A top-down processor allocation scheme for hypercube computers. IEEE Transactions on Parallel and Distributed Systems, 2(1):20–30, January 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kalpakis:2001:OPR

[KDW01] K. Kalpakis, K. Dasgupta, and O. Wolfson. Optimal placement of replicas in trees with read, write, and storage costs. IEEE Transactions on Parallel and Distributed Systems, 12(6):628– ??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0628.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0628abs.
                                        htm.

Kotz:1990:PFS

[KE90] David F. Kotz and Carla Schlatter Ellis. Prefetching in file systems for MIMD multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 1 (2):218–230, April 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kyriacou:2006:DDM

[KET06] Costas Kyriacou, Paraskevas Evripidou, and Pedro Trancoso. Data-driven multithreading using conventional microprocessors. IEEE Transactions on Parallel and Distributed Systems, 17 (10):1176–1188, October 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Klaiber:1992:PEA

[KG92] Alexander Klaiber and Maya Gokhale. Parallel evaluation of attribute grammars. IEEE Transactions on Parallel and Distributed Systems, 3(2):206– 220, March 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kweon:2004:SRT

[KgCS04] Seok-Kyu Kweon, Min gyu Cho, and Kang G. Shin. Soft realtime communication over Ethernet with adaptive traffic smoothing. IEEE Transactions on Parallel and Distributed Systems, 15(10):946–959, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/10/l0946abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0946.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0946.
pdf.

Kamal:2008:PCA

[KGKL08] Zill-E-Huma Kamal, Ajay Gupta, Ashfaq A. Khokhar, and Leszek Lilien. Power consumption analysis of maximum a posteriori classifier using LU decomposition and Jacobi iterations. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1473–1483, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kao:1996:SSR

[KGM96] Benjamin Kao and Hector Garcia-Molina. Scheduling soft real-time jobs over dual nonreal-time servers. IEEE Transactions on Parallel and Distributed Systems, 7(1):56–68, January 1996. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kao:1997:DAD

[KGM97] B. Kao and H. Garcia-Molina. Deadline assignment in a distributed soft real-time system. IEEE Transactions on Parallel and Distributed Systems, 8(12):1268–1274, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1268.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1268abs.
                                        htm.

Kao:1994:ATS

[KGMB94] Ben Kao, Hector Garcia-Molina, and Daniel Barbara. Aggressive transmissions of short messages over redundant paths. IEEE Transactions on Parallel and Distributed Systems, 5(1): 102–109, January 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kramer:1994:CDT

[KGS94] Robert Kramer, Rajiv Gupta, and Mary Lou Soffa. Combining DAG: a technique for parallel data flow analysis. IEEE Transactions on Parallel and Distributed Systems, 5(8):805–813, August 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:1993:SAS

[KH93] D. Kumar and S. Harous. A study of achievable speedup in distributed simulation via NULL messages. IEEE Transactions on Parallel and Distributed Systems, 4(3):347–354, March 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ku:1997:CFT

[KH97a] Hung-Kuei Ku and John P.

Hayes. Connective fault tolerance in multiple-bus systems. IEEE Transactions on Parallel and Distributed Systems, 8 (6):574–586, June 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l0574.
pdf;   http://www.computer.
org/tpds/td1997/l0574abs.
htm.

Kuo:1997:GAC

[KH97b] Yu-Chen Kuo and Shing-Tsaan Huang. A geometric approach for constructing coteries and k-coteries. IEEE Transactions on Parallel and Distributed Systems, 8(4):402–411, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0402.
pdf;   http://www.computer.
org/tpds/td1997/l0402abs.
htm.

Kuo:1998:RNC

[KH98] Y.-C. Kuo and S.-T. Huang. Recognizing nondominated coteries and wr-coteries by availability. IEEE Transactions on Parallel and Distributed Systems, 9(8):721–728, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0721.
pdf;   http://www.computer.
org/tpds/td1998/l0721abs.
htm.

Katsinis:2004:FTD

[KH04] Constantine Katsinis and Diana Hecht. Fault-tolerant distributed shared memory on a broadcast-based architecture. IEEE Transactions on Parallel and Distributed Systems, 15(12):1082–1092, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/12/l1082abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1082.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1082.
pdf.

Kandasamy:2005:TCF

[KHM05] Nagarajan Kandasamy, John P.

Hayes, and Brian T. Murray. Time-constrained failure diagnosis in distributed embedded systems: Application to actuator diagnosis. IEEE Transactions on Parallel and Distributed Systems, 16(3):258–270, March 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kwok:2007:SGG

[KHS07] Yu-Kwong Kwok, Kai Hwang, and ShanShan Song. Selfish Grids: Game-theoretic modeling and NAS/PSA benchmark evaluation. IEEE Transactions on Parallel and Distributed Systems, 18(5):621–636, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kao:1995:DEP

[KHWT95] Tzong-Wann Kao, Shi-Jinn Horng, Yue-Li Wang, and Horng-Ren Tsai. Designing efficient parallel algorithms on CRAP. IEEE Transactions on Parallel and Distributed Systems, 6(5):554–560, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0554abs.htm.

Khargharia:2009:AIT

[KHY09] Bithika Khargharia, Salim Hariri, and Mazin S. Yousif.

An adaptive interleaving technique for memory performanceper-watt management. IEEE Transactions on Parallel and Distributed Systems, 20(7):1011– 1022, July 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kalbarczyk:1999:CSI

[KIBW99] Z. T. Kalbarczyk, R. K. Iyer, S. Bagchi, and K. Whisnant. Chameleon: a software infrastructure for adaptive fault tolerance. IEEE Transactions on Parallel and Distributed Systems, 10(6):560–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0560.
pdf;   http://www.computer.
org/tpds/td1999/l0560abs.
htm.

Kiniwa:2006:HIS

[Kin06] Jun Kiniwa. How to improve safety under convergence using stable storage. IEEE Transactions on Parallel and Distributed Systems, 17(4):389–398, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kremien:1992:MAA

[KK92] O. Kremien and J. Kramer. Methodical analysis of adaptive load sharing algorithms. IEEE Transactions on Parallel and Distributed Systems, 3(6): 747–760, November 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:1993:DRB

[KK93a] K. H. (Kane) Kim and A. Kavianpour. Distributed recovery block approach to fault-tolerant execution of application tasks in hypercubes. IEEE Transactions on Parallel and Distributed Systems, 4(1):104–111, January 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kleinrock:1993:CUP

[KK93b] Leonard Kleinrock and Williard Korfhage. Collecting unused processing capacity: an analysis of transient distributed systems. IEEE Transactions on Parallel and Distributed Systems, 4 (5):535–546, May 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Karypis:1994:UTS

[KK94] George Karypis and Vipin Kumar. Unstructured tree search on SIMD parallel computers. IEEE Transactions on Parallel and Distributed Systems, 5(10): 1057–1072, October 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kanodia:2003:ELT

[KK03a] Vikram Kanodia and Edward W. Knightly. Ensuring latency targets in multiclass Web servers. IEEE Transactions on Parallel and Distributed Systems, 14(1):84–93, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/01/l0084abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0084.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0084.
pdf.

Kuzmanovic:2003:MBC

[KK03b] Aleksandar Kuzmanovic and Edward W. Knightly. Measurementbased characterization and classification of QoS-enhanced systems. IEEE Transactions on Parallel and Distributed Systems, 14(7):671–685, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/07/l0671abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/07/l0671.
pdf.

Kadayif:2004:QLO

[KK04] Ismail Kadayif and Mahmut Kandemir. Quasidynamic layout optimizations for improving data locality. IEEE Transactions on Parallel and Distributed Systems, 15(11):996–1011, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/11/l0996abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l0996.
htm;   http://csdl.computer.

org/dl/trans/td/2004/11/l0996.
pdf.

Kuo:2003:RTD

[KKC03] Chin-Fu Kuo, Tei-Wei Kuo, and Cheng Chang. Real-time digital signal processing of phased array radars. IEEE Transactions on Parallel and Distributed Systems, 14(5):433–446, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/05/l0433abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0433.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0433.
pdf.

Kadayif:2005:OAI

[KKC^{+}05] Ismail Kadayif, Mahmut Kandemir, Guilin Chen, Ozcan Ozturk, Mustafa Karakoy, and Ugur Sezer. Optimizing arrayintensive applications for onchip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 16(5):396– 411, May 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kandaswamy:2002:EEOa

[KKCB02a] Meenakshi A. Kandaswamy, Mahmut Kandemir, Alok Choudhary, and David Bernholdt. An experimental evaluation of I/O optimizations on different applications. IEEE Transactions on Parallel and Distributed Systems, 13(7):728–744, July 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0728.
pdf;   http://www.computer.
org/tpds/td2002/l0728abs.

htm. Editor’s Note: This paper unfortunately contains some errors which led to the paper being reprinted in the December 2002 issue. Please see IEEE Transactions on Parallel and Distributed Systems, vol. 13, no. 12, December 2002, pp. 1303–1319 for the correct paper.

Kandaswamy:2002:EEOb

[KKCB02b] Meenakshi A. Kandaswamy, Mahmut Kandemir, Alok Choudhary, and David Bernholdt. An experimental evaluation of I/O optimizations on different applications. IEEE Transactions on Parallel and Distributed Systems, 13(12):1303– 1319, December 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2002/12/
l1303.pdf;   http://csdl.
computer.org/comp/trans/td/
2002/12/l1303abs.htm;   http:
//csdl.computer.org/dl/trans/
td/2002/12/l1303.htm.

Kahol:2001:SMC

[KKGS01] Anurag Kahol, Sumit Khurana, Sandeep K. S. Gupta, and Pradip K. Srimani. A strategy to manage cache consistency in a disconnected distributed environment. IEEE Transactions on Parallel and Distributed Systems, 12(7):686– 700, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0686.
pdf;   http://www.computer.
org/tpds/td2001/l0686abs.
htm.

Kakugawa:2008:TBD

[KKM08] Hirotsugu Kakugawa, Sayaka Kamei, and Toshimitsu Masuzawa. A token-based distributed group mutual exclusion algorithm with quorums. IEEE Transactions on Parallel and Distributed Systems, 19(9):1153– 1166, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kong:1991:TID

[KKP91] Xiangyun Kong, David Klappholz, and Kleanthis Psarris. The I Test: An improved dependence test for automatic parallelization and vectorization. IEEE Transactions on Parallel and Distributed Systems, 2 (3):342–349, July 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kalasapur:2007:DSC

[KKS07] Swaroop Kalasapur, Mohan Kumar, and Behrooz A. Shirazi. Dynamic service composition in pervasive computing. IEEE Transactions on Parallel and Distributed Systems, 18(7):907– 918, July 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:1999:PBP

[KL99] J. S. Kim and D. J. Lilja. Performance-based path determination for interprocessor communication in distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 10(3):316–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0316.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0316abs.
                                        htm.

Kazi:2001:CGT

[KL01] Iffat H. Kazi and David J. Lilja. Coarse-grained thread pipelining: a speculative parallel execution model for shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 12(9):952–966, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0952.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0952abs.
                                        htm.

Kwok:2002:NCA

[KL02] Yu-Kwong Kwok and Vincent K. N. Lau. A novel channeladaptive uplink access control protocol for nomadic computing. IEEE Transactions on Parallel and Distributed Systems, 13(11):1150–1165, November 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/11/l1150abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/11/l1150.
htm;   http://csdl.computer.
org/dl/trans/td/2002/11/l1150.
pdf.

Klasing:1998:ICC

[Kla98] R. Klasing. Improved compressions of cube-connected cycles networks. IEEE Transactions on Parallel and Distributed Systems, 9(8):803–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0803.
pdf;   http://www.computer.
org/tpds/td1998/l0803abs.
htm.

Kim:1997:CRF

[KLC97] Jae H. Kim, Ziqiang Liu, and Andrew A. Chien. Compressionless routing: a framework for adaptive and faulttolerant routing. IEEE Transactions on Parallel and Distributed Systems, 8(3):229–244, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0229.
pdf;   http://www.computer.
org/tpds/td1997/l0229abs.
htm.

Liao:2006:SDI

[kLCC^{+}06] Wei keng Liao, Kenin Coloma, Alok Choudhary, Lee Ward, Eric Russell, and Neil Pundit. Scalable design and implementations for MPI parallel overlapping I/O. IEEE Transactions on Parallel and Distributed Systems, 17(11):1264–1276, November 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Krueger:1994:JSM

[KLDR94] Phillip Krueger, Ten-Hwang Lai, and Vibha A. Dixit-Radiya. Job scheduling is more important than processor allocation for hypercube computers. IEEE Transactions on Parallel and Distributed Systems, 5(5):488– 497, May 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2007:PPD

[KLH07] Sang Cheol Kim, Sunggu Lee, and Jaegyoon Hahm. Push-Pull: Deterministic search-based DAG scheduling for heterogeneous cluster systems. IEEE Transactions on Parallel and Distributed Systems, 18(11):1489– 1502, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ko:2000:NOB

[KLS00] H. Ko, S. Latifi, and P. K. Srimani. Near-optimal broadcast in all-port wormholerouted hypercubes using errorcorrecting codes. IEEE Transactions on Parallel and Distributed Systems, 11(3):247–??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0247.
pdf;   http://www.computer.
org/tpds/td2000/l0247abs.
htm.

Koelbel:1991:CGN

[KM91] Charles Koelbel and Piyush Mehrotra. Compiling global name-space parallel loops for distributed execution. IEEE Transactions on Parallel and Distributed Systems, 2(4):440– 451, October 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Keane:2001:SLS

[KM01] Patrick Keane and Mark Moir.

A simple local-spin group mutual exclusion algorithm. IEEE Transactions on Parallel and Distributed Systems, 12(7):673– 685, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0673.
pdf;   http://www.computer.
org/tpds/td2001/l0673abs.
htm.

Knoop:2002:DAP

[KM02] Jens Knoop and Eduard Mehofer.

Distribution assignment placement: Effective optimization of redistribution costs. IEEE Transactions on Parallel and Distributed Systems, 13(6):628– 647, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0628.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0628abs.
                                        htm.

Kermarrec:2003:PRD

[KMG03] Anne-Marie Kermarrec, Laurent Massoulié, and Ayalvadi J. Ganesh. Probabilistic reliable dissemination in largescale systems. IEEE Transactions on Parallel and Distributed Systems, 14(3):248–258, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/03/l0248abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/03/l0248.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/03/l0248.
                                        pdf.

Kennedy:1991:IPP

[KMT91] Ken Kennedy, Kathryn S. McKinley, and Chau-Wen Tseng. Interactive parallel programming using the ParaScope Editor. IEEE Transactions on Parallel and Distributed Systems, 2 (3):329–341, July 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kafura:1995:CDG

[KMW95] Dennis Kafura, Manibrata Mukherji, and Douglas M.

Washabaugh. Concurrent and distributed garbage collection of active objects. IEEE Transactions on Parallel and Distributed Systems, 6(4):337–350, April 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0337abs.htm.

Kumar:2008:EEE

[KMW08] G. Sudha Anil Kumar, Govindarasu Manimaran, and Zhengdao Wang. End-to-end energy management in networked real-time embedded systems. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1498–1510, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kalns:1995:PMT

[KN95] Edgar T. Kalns and Lionel M.

Ni. Processor mapping techniques toward efficient data redistribution. IEEE Transactions on Parallel and Distributed Systems, 6(12):1234–1247, December 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koppelman:1994:RPM

[Kop94] David M. Koppelman. Reducing PE/memory traffic in multiprocessors by the difference coding of memory addresses. IEEE Transactions on Parallel and Distributed Systems, 5(11):1156– 1168, November 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koppelman:1996:FIN

[Kop96]  David M. Koppelman.  A

family of interconnection networks for nonuniform traffic. IEEE Transactions on Parallel and Distributed Systems, 7 (5):486–492, May 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0486abs.htm.

Kumar:1992:EHH

[KP92] J. Mohan Kumar and L. M. Patnaik. Extended hypercube: a hierarchical interconnection network of hypercubes. IEEE Transactions on Parallel and Distributed Systems, 3(1):45–57, January 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:1993:EPC

[KP93a] Junguk L. Kim and Taesoon Park. Efficient protocol for checkpointing recovery in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 4(8):955–960, August 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:1993:LSP

[KP93b] Kichul Kim and Victor K. Prasanna. Latin squares for Parallel Array Access. IEEE Transactions on Parallel and Distributed Systems, 4(4):361–370, April 1993. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Kravets:1996:ANS

[KP96] Dina Kravets and C. Greg Plaxton. All nearest smaller values on the hypercube. IEEE Transactions on Parallel and Distributed Systems, 7(5):456– 462, May 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1996/l0456abs.
htm.

Kesavan:1999:MMM

[KP99] R. Kesavan and D. K. Panda.

Multiple multicast with minimized node contention on wormhole k-ary n-cube networks. IEEE Transactions on Parallel and Distributed Systems, 10 (4):371–??, April 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l0371.
pdf;   http://www.computer.
org/tpds/td1999/l0371abs.
htm.

Kesavan:2001:EMI

[KP01] Ram Kesavan and Dhabaleswar K.

Panda. Efficient multicast on irregular switch-based cutthrough networks with updown routing. IEEE Transactions on Parallel and Distributed Systems, 12(8):808–828, August 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0808.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0808abs.
                                        htm.

Kyriakopoulos:2009:NSA

[KP09]  Konstantinos  Kyriakopoulos

and Kleanthis Psarris. Nonlinear symbolic analysis for advanced program parallelization. IEEE Transactions on Parallel and Distributed Systems, 20 (5):623–640, May 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kundu:2009:AFA

[KPBD09] Sumantra R. Kundu, Sourav Pal, Kalyan Basu, and Sajal K. Das. An architectural framework for accurate characterization of network traffic. IEEE Transactions on Parallel and Distributed Systems, 20(1):111–123, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kuo:2009:DRS

[KPC09] Chin-Fu Kuo, Ai-Chun Pang, and Sheng-Kun Chan. Dynamic routing with security considerations. IEEE Transactions on Parallel and Distributed Systems, 20(1):48–58, January 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khan:2009:DAC

[KPK09] Maleq Khan, Gopal Pandurangan, and V. S. Anil Kumar.

Distributed algorithms for constructing approximate minimum spanning trees in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(1):124–139, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kuntraruk:2005:ARR

[KPR05] Jirada Kuntraruk, William M.

Pottenger, and Andrew M. Ross. Application resource requirement estimation in a parallelpipeline model of execution. IEEE Transactions on Parallel and Distributed Systems, 16(12): 1154–1165, December 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kazmierczak:2000:ODE

[KR00] A. Kazmierczak and S. Radhakrishnan. An optimal distributed ear decomposition algorithm with applications of biconnectivity and outerplanarity testing. IEEE Transactions on Parallel and Distributed Systems, 11(2):110–??, February 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0110.
pdf;   http://www.computer.
org/tpds/td2000/l0110abs.
htm.

Krothapalli:1991:RRD

[KS91] V. P. Krothapalli and P. Sadayappan. Removal of redundant dependences in DOACROSS loops with constant dependences. IEEE Transactions on Parallel and Distributed Systems, 2(3):281–289, July 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:1993:PAS

[KS93] Anurag Kumar and Rajeev Shorey. Performance analysis and scheduling of stochastic fork-join jobs in a multicomputer system. IEEE Transactions on Parallel and Distributed Systems, 4(10):1147–1164, October 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:1994:OEF

[KS94] Jong Kim and Kang G. Shin. Operationally enhanced folded hypercubes. IEEE Transactions on Parallel and Distributed Systems, 5(12):1310–1316, December 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kweon:2001:RTT

[KS01] S.-K. Kweon and K. G. Shin. Real-time transport of MPEG video with a statistically guaranteed loss ratio in ATM networks. IEEE Transactions on Parallel and Distributed Systems, 12(4):387–??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0387.

pdf;   http://www.computer.
org/tpds/td2001/l0387abs.
htm.

Kweon:2003:SRT

[KS03] Seok-Kyu Kweon and Kang G.

Shin. Statistical real-time communication over Ethernet. IEEE Transactions on Parallel and Distributed Systems, 14(3):322–335, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0322abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0322.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0322.
pdf.

Krintz:2006:AFC

[KS06] Chandra Krintz and Sezgin Sucu. Adaptive on-the-fly compression. IEEE Transactions on Parallel and Distributed Systems, 17(1):15–24, January 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Keidar:2008:HCT

[KS08a] Idit Keidar and Alexander Shraer. How to choose a timing model. IEEE Transactions on Parallel and Distributed Systems, 19(10):1367–1380, October 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kwon:2008:LTM

[KS08b] Oh-Heum Kwon and Ha-Joo Song. Localization through map stitching in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(1):93–105, January 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:1994:SPF

[KSA94] Vipin Kumar, Shashi Shekhar, and Minesh B. Amin. Scalable parallel formulation of the backpropagation algorithm for hypercubes and related architectures. IEEE Transactions on Parallel and Distributed Systems, 5(10):1073–1090, October 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Krishnamurthy:2003:AQS

[KSC03] Sudha Krishnamurthy, William H. Sanders, and Michel Cukier. An adaptive quality of service aware middleware for replicated services. IEEE Transactions on Parallel and Distributed Systems, 14(11):1112–1125, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/11/l1112abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/11/l1112.
                                        pdf.

Kandlur:1994:RTC

[KSF94] Dilip D. Kandlur, Kang G. Shin, and Domenico Ferrari. Realtime communication in multihop networks. IEEE Transactions on Parallel and Distributed Systems, 5(10):1044–1056, October 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kshemkalyani:2003:FGM

[Ksh03]  Ajay D. Kshemkalyani.  A

fine-grained modality classification for global predicates. IEEE Transactions on Parallel and Distributed Systems, 14(8):807–816, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0807abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/08/l0807.
pdf.

Kim:2008:DRM

[KSME08] Jong-Kook Kim, Howard Jay Siegel, Anthony A. Maciejewski, and Rudolf Eigenmann. Dynamic resource management in energy constrained heterogeneous computing systems using voltage scaling. IEEE Transactions on Parallel and Distributed Systems, 19(11):1445– 1457, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kanhere:2002:FEP

[KSP02] S. S. Kanhere, H. Sethu, and A. B. Parekh. Fair and efficient packet scheduling using elastic round robin. IEEE Transactions on Parallel and Distributed Systems, 13(3):324–336, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0324.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0324abs.
                                        htm.

Kwon:2009:ESO

[KSP09] Oh-Heum Kwon, Ha-Joo Song, and Sangjoon Park. The effects of stitching orders in patchand-stitch WSN localization algorithms. IEEE Transactions on Parallel and Distributed Systems, 20(9):1380–1391, September 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kirousis:1994:RMV

[KST94] Lefteris M. Kirousis, Paul Spirakis, and Philippas Tsigas. Reading many variables in one atomic operation: solutions with linear or sublinear complexity. IEEE Transactions on Parallel and Distributed Systems, 5 (7):688–696, July 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Krishnamurthy:2003:NCS

[KSWR03] Raj Krishnamurthy, Karsten Schwan, Richard West, and Marcel-Catalin Rosu. On network CoProcessors for scalable, predictable media services. IEEE Transactions on Parallel and Distributed Systems, 14(7):655–670, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/07/l0655abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/07/l0655.
pdf.

Kucera:2001:WFD

[Kuc01] L. Kucera. Wait-free deflection routing of long messages. IEEE Transactions on Parallel and Distributed Systems, 12(5):476– ??, May 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0476.
pdf;   http://www.computer.
org/tpds/td2001/l0476abs.
htm.

Kumar:1992:SLD

[Kum92] Devendra Kumar. Systems with low distributed simulation overhead. IEEE Transactions on Parallel and Distributed Systems, 3(2):155–165, March 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kuo:2001:CA

[Kuo01] Yu-Chen Kuo. Composite karbiters. IEEE Transactions on Parallel and Distributed Systems, 12(11):1134–1145, November 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l1134.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1134abs.
                                        htm.

Kwok:2008:SAH

[KW08] Michael Kwok and Johnny W. Wong. Scalability analysis of the hierarchical architecture for distributed virtual environments. IEEE Transactions on Parallel and Distributed Systems, 19(3): 408–417, March 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kuo:2002:RTC

[KWH02] Tei-Wei Kuo, Jun Wu, and Hsin-Chia Hsih. Real-time concurrency control in a multiprocessor environment. IEEE Transactions on Parallel and Distributed Systems, 13(6):659– 671, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0659.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0659abs.
                                        htm.

Ku:2003:CED

[KWH03] Shan-Chyun Ku, Biing-Feng Wang, and Ting-Kai Hung. Constructing edge-disjoint spanning trees in product networks. IEEE Transactions on Parallel and Distributed Systems, 14(3):213–221, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0213abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0213.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0213.
pdf.

Kim:2009:CMC

[KWL^{+}09] Donghyun Kim, Yiwei Wu, Yingshu Li, Feng Zou, and Ding-Zhu Du. Constructing minimum connected dominating sets with bounded diameters in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 20(2):147–157, February 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koibuchi:2005:PED

[KWOA05] Michihiro Koibuchi, Konosuke Watanabe, Tomohiro Otsuka, and Hideharu Amano. Performance evaluation of deterministic routings, multicasts, and topologies on RHiNET-2 cluster. IEEE Transactions on Parallel and Distributed Systems, 16 (8):747–759, August 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kakugawa:1997:USS

[KY97] H. Kakugawa and Masafumi Yamashita. Uniform and selfstabilizing token rings allowing unfair daemon. IEEE Transactions on Parallel and Distributed Systems, 8(2):154–163, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0154.
pdf;   http://www.computer.
org/tpds/td1997/l0154abs.
htm.

Kim:1998:SAM

[KY98] G. Kim and H. Yoon. On submesh allocation for mesh multicomputers: a best-fit allocation and a virtual submesh allocation for faulty meshes. IEEE Transactions on Parallel and Distributed Systems, 9(2):175–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0175.
pdf;   http://www.computer.
org/tpds/td1998/l0175abs.
htm.

Koglin:2008:ESC

[KYB08] Yunhua Koglin, Danfeng Yao, and Elisa Bertino. Efficient and secure content processing and distribution by cooperative intermediaries. IEEE Transactions on Parallel and Distributed Systems, 19(5):615–626, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2007:EID

[KYD^{+}07] Eun Jung Kim, Ki Hwan Yum, Chita R. Das, Mazin Yousif, and José Duato. Exploring IBA design space for improved performance. IEEE Transactions on Parallel and Distributed Systems, 18(4):498–510, April 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2009:PCF

[KyK09] Hwangnam Kim and E yong Kim. PReSENt: a collaboration framework for resource sharing in wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 20(3):289– 302, March 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kamat:1996:EOR

[KZ96] Sanjay Kamat and Wei Zhao.

An efficient optimal reconfiguration algorithm for FDDI-based networks. IEEE Transactions on Parallel and Distributed Systems, 7(4):411–424, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0411abs.htm.

Kang:2007:TTA

[KZN07] Jaewon Kang, Yanyong Zhang, and Badri Nath. TARA: Topology-aware resource adaptation to alleviate congestion in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(7):919–931, July 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lindon:1993:OIB

[LA93] L. F. Lindon and S. G. Akl. An optimal implementation of broadcasting with selective reduction. IEEE Transactions on Parallel and Distributed Systems, 4(3):256–269, March 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Loukopoulos:2004:ODT

[LA04] Thanasis Loukopoulos and Ishfaq Ahmad. Optimizing download time of embedded multimedia objects for Web browsing. IEEE Transactions on Parallel and Distributed Systems, 15(10):934–945, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/10/l0934abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/10/l0934.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/10/l0934.
                                        pdf.

Loukopoulos:2006:PCO

[LA06] Thanasis Loukopoulos and Ishfaq Ahmad. Policies for caching OLAP queries in Internet proxies. IEEE Transactions on Parallel and Distributed Systems, 17 (10):1124–1135, October 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lai:2000:PPI

[Lai00] W. K. Lai. Performing permutations on interconnection networks by regularly changing switch states. IEEE Transactions on Parallel and Distributed Systems, 11(8):829–??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0829.
pdf;   http://www.computer.
org/tpds/td2000/l0829abs.
htm.

Lan:1995:AFT

[Lan95]  Youran Lan.  An adaptive

fault-tolerant routing algorithm for hypercube multicomputers. IEEE Transactions on Parallel and Distributed Systems, 6 (11):1147–1152, November 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1147abs.htm.

Larus:1993:LLP

[Lar93] James R. Larus. Loop-level parallelism in numeric and symbolic programs. IEEE Transactions on Parallel and Distributed Systems, 4(7):812–826, July 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2004:DIE

[LAS04] Ying Lu, Tarek F. Abdelzaher, and Avneesh Saxena. Design, implementation, and evaluation of differentiated caching services. IEEE Transactions on Parallel and Distributed Systems, 15(5):440–452, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/05/l0440abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0440.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0440.
pdf.

Latifi:1994:TAS

[Lat94] Shahram Latifi. Task allocation in the star graph. IEEE Transactions on Parallel and Distributed Systems, 5(11):1220– 1224, November 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lakshmanan:2003:RSS

[LAV03] Subramanian Lakshmanan, Mustaque Ahamad, and H. Venkateswaran. Responsive security for stored data. IEEE Transactions on Parallel and Distributed Systems, 14(9):818–828, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0818abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0818.
pdf.

Lawrie:1995:E

[Law95] Duncan Lawrie. Editorial. IEEE Transactions on Parallel and Distributed Systems, 6(3):??, March 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lawrie:1997:E

[Law97] D. H. Lawrie. Editorial. IEEE Transactions on Parallel and Distributed Systems, 8(3):225– ??, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0225.
pdf.

Latifi:1994:ISI

[LB94] Shahram Latifi and Nader Bagherzadeh. Incomplete star: an incrementally scalable network based on the star graph. IEEE Transactions on Parallel and Distributed Systems, 5 (1):97–102, January 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1995:MMS

[LB95] De-Lei Lee and Kenneth E.

Batcher. A multiway merge sorting network. IEEE Transactions on Parallel and Distributed Systems, 6(2):211–215, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0211abs.htm.

Lebak:2000:DPE

[LB00a] J. M. Lebak and A. W. Bojanczyk. Design and performance evaluation of a portable parallel library for space-time adaptive processing. IEEE Transactions on Parallel and Distributed Systems, 11(3):287– ??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0287.
pdf;   http://www.computer.
org/tpds/td2000/l0287abs.
htm.

Lee:2000:MCB

[LB00b] J.-D. Lee and K. E. Batcher. Minimizing communication in the bitonic sort. IEEE Transactions on Parallel and Distributed Systems, 11(5):459–??, May 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0459.
pdf;   http://www.computer.
org/tpds/td2000/l0459abs.
htm.

Luo:2003:SMM

[LBC03] Yan Luo, Laxmi Narayan Bhuyan, and Xi Chen. Shared memory multiprocessor architectures for software IP routers. IEEE Transactions on Parallel and Distributed Systems, 14(12):1240–1249, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1240abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1240.
pdf.

Liang:2001:FDM

[LBS01] Weifa Liang, Richard P. Brent, and Hong Shen. Fully dynamic maintenance of k-connectivity in parallel. IEEE Transactions on Parallel and Distributed Systems, 12(8):846–864, August 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0846.
pdf;   http://www.computer.
org/tpds/td2001/l0846abs.
htm.

Liss:2005:KIO

[LBS05] Liran Liss, Yitzhak Birk, and Assaf Schuster. In-kernel integration of operating system and Infiniband functions for high performance computing clusters: a DSM example. IEEE Transactions on Parallel and Distributed Systems, 16(9):830–840, September 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:1991:CCE

[LC91a] Jingke Li and Marina Chen.

Compiling communication-efficient programs for massively parallel machines. IEEE Transactions on Parallel and Distributed Systems, 2(3):361–376, July 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:1991:JSP

[LC91b] Keqin Li and Kam-Hoi Cheng.

Job scheduling in a partitionable mesh using a twodimensional buddy system partitioning scheme. IEEE Transactions on Parallel and Distributed Systems, 2(4):413–422, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1994:TDF

[LC94] Tsern-Huei Lee and Jin-Jye Chou. Testing the dynamic full access property of a class of multistage interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 5(11):1206–1210, November 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1995:OHS

[LC95] Soo-Young Lee and Chung-Ming Chen. Optimal hot spot allocation on meshes for largescale data-parallel algorithms. IEEE Transactions on Parallel and Distributed Systems, 6(8): 788–802, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0788abs.htm.

Lau:1996:OLM

[LC96a] Francis C. M. Lau and Guihai Chen. Optimal layouts of midimew networks. IEEE Transactions on Parallel and Distributed Systems, 7(9):954– 961, September 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0954abs.htm.

Lee:1996:ECB

[LC96b] Sang-Kyu Lee and Hyeong-Ah Choi. Embedding of complete binary trees into meshes with rowcolumn routing. IEEE Transactions on Parallel and Distributed Systems, 7(5):493–497, May 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0493abs.htm.

Liao:1999:TBP

[LC99] C.-J. Liao and Y.-C. Chung.

Tree-based parallel load-balancing methods for solution-adaptive finite element graphs on distributed memory multicomputers. IEEE Transactions on Parallel and Distributed Systems, 10(4):360–??, April 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l0360.
pdf;   http://www.computer.
org/tpds/td1999/l0360abs.
htm.

Lo:2001:EHP

[LC01] R.-S. Lo and G.-H. Chen. Embedding Hamiltonian paths in faulty arrangement graphs with the backtracking method. IEEE Transactions on Parallel and Distributed Systems, 12(2):209– ??, February 2001. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td2001/pdf/l0209.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0209abs.
                                        htm.

Lin:2002:AOM

[LC02a] Cheng-Ru Lin and Ming-Syan Chen. On the asymptotical optimality of multilayered decentralized consensus protocol. IEEE Transactions on Parallel and Distributed Systems, 13(8):769– 786, August 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0769.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0769abs.
                                        htm.

Lui:2002:EPA

[LC02b] J. C. S. Lui and M. F. Chan. An efficient partitioning algorithm for distributed virtual environment systems. IEEE Transactions on Parallel and Distributed Systems, 13(3):193–211, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0193.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0193abs.
                                        htm.

Losee:2004:IRD

[LC04] Robert M. Losee and Lewis Church, Jr. Information retrieval with distributed databases: Analytic models of performance. IEEE Transactions on Parallel and Distributed Systems, 15(1):18–27, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/01/l0018abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/01/l0018.
pdf.

Lewandowski:1996:AAP

[LCB96] Gary Lewandowski, Anne Condon, and Eric Bach. Asynchronous analysis of parallel dynamic programming algorithms. IEEE Transactions on Parallel and Distributed Systems, 7 (4):425–438, April 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0425abs.htm.

Lain:2000:CRT

[LCB00] A. Lain, D. R. Chakrabarti, and P. Banerjee. Compiler and run-time support for exploiting regularity within irregular applications. IEEE Transactions on Parallel and Distributed Systems, 11(2):119–??, February 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0119.
pdf;   http://www.computer.
org/tpds/td2000/l0119abs.
htm.

Lao:2007:SOM

[LCGC07] Li Lao, Jun-Hong Cui, Mario Gerla, and Shigang Chen. A scalable overlay multicast architecture for large-scale applications. IEEE Transactions on Parallel and Distributed Systems, 18(4):449–459, April 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:1995:FTD

[lCL95] Yi long Chen and Jyh-Charn Liu. A fault-tolerant distributed subcube management scheme for hypercube multicomputer systems. IEEE Transactions on Parallel and Distributed Systems, 6(7):766–772, July 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0766abs.htm.

Lin:2003:EDP

[LCL03] Chun-Yuan Lin, Yeh-Ching Chung, and Jen-Shiuh Liu. Efficient data parallel algorithms for multidimensional array operations based on the EKMR scheme for distributed memory multicomputers. IEEE Transactions on Parallel and Distributed Systems, 14(7):625–639, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/07/l0625abs.
                                        htm;   http://csdl.computer.

org/dl/trans/td/2003/07/l0625.
pdf.

Lian:2007:GBD

[LCN^{+}07] Jie Lian, Lei Chen, Kshirasagar Naik, Yunhao Liu, and Gordon B. Agnew. Gradient boundary detection for time series snapshot construction in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(10):1462–1475, October 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lacy:1998:OCT

[LCRW98] W. S. Lacy, J. L. Cruz-Rivera, and D. S. Wills. The offset cube: a three-dimensional multicomputer network topology using through-wafer optics. IEEE Transactions on Parallel and Distributed Systems, 9(9):893–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0893.
pdf;   http://www.computer.
org/tpds/td1998/l0893abs.
htm.

Li:2003:LDT

[LCWW03] Xiang-Yang Li, Gruia Calinescu, Peng-Jun Wan, and Yu Wang. Localized Delaunay triangulation with application in ad hoc wireless networks. IEEE Transactions on Parallel and Distributed Systems, 14(10):1035–1047, October 2003. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/10/l1035abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/10/l1035.
                                        pdf.

Lee:2008:FFS

[LDCO08] Winnie Louis Lee, Amitava Datta, and Rachel Cardell-Oliver. FlexiTP: a flexibleschedule-based TDMA protocol for fault-tolerant and energyefficient wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(6):851–864, June 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Lopez:2004:MMU

[LDG04] José M. López, José L. Díaz, and Daniel F. García. Minimum and maximum utilization bounds for multiprocessor rate monotonic scheduling. IEEE Transactions on Parallel and Distributed Systems, 15(7):642–653, July 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/07/l0642.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/07/l0642.pdf.

Lee:1991:CDG

[Lee91] Edward Ashford Lee. Consistency in dataflow graphs. IEEE Transactions on Parallel and Distributed Systems, 2(2):223– 235, April 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1993:VBA

[Lee93] K. C. Lee. Virtual bus architecture for dynamic parallel processing. IEEE Transactions on Parallel and Distributed Systems, 4(2):121–130, February 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1995:PIL

[Lee95] Gyungho Lee. Parallelizing iterative loops with conditional branching. IEEE Transactions on Parallel and Distributed Systems, 6(2):185–189, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0185abs.htm.

Lee:1997:EAD

[Lee97] PeiZong Lee. Efficient algorithms for data distribution on distributed memory parallel computers. IEEE Transactions on Parallel and Distributed Systems, 8(8):825–839, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0825.
pdf;   http://www.computer.
org/tpds/td1997/l0825abs.
htm.

Lee:2006:SCS

[Lee06] Daniel C. Lee. Some compartmentalized secure task assignment models for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 17(12):1414–1424, December 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

LaRowe:1992:ENM

[LEH92] R. P. LaRowe, Jr., C. S. Ellis, and M. A. Holliday. Evaluation of NUMA memory management through modeling and measurements. IEEE Transactions on Parallel and Distributed Systems, 3(6):686–701, November 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2003:GIR

[LF03] Hyuk Jae Lee and José A. B. Fortes. Generation of injective and reversible modular mappings. IEEE Transactions on Parallel and Distributed Systems, 14(1):1–12, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/01/l0001abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/01/l0001.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/01/l0001.
                                        pdf.

Luan:1990:FTP

[LG90] Shyh-Wei Luan and Virgil D.

Gligor. A fault-tolerant protocol for atomic broadcast. IEEE Transactions on Parallel and Distributed Systems, 1(3):271– 285, July 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lakshman:1994:PEE

[LG94] T. V. Lakshman and Dipak Ghosal. Performance evaluation of an efficient multiple copy update algorithm. IEEE Transactions on Parallel and Distributed Systems, 5(2):217–224, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2008:DFD

[LG08] Alex X. Liu and Mohamed G.

Gouda. Diverse firewall design. IEEE Transactions on Parallel and Distributed Systems, 19 (9):1237–1251, September 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2009:FPQ

[LG09] Alex X. Liu and Mohamed G.

Gouda. Firewall policy queries. IEEE Transactions on Parallel and Distributed Systems, 20 (6):766–777, June 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1993:PIE

[LH93] Edward K. B. Lee and Simon Haykin. Parallel implementation of the extended square-root covariance filter for tracking applications. IEEE Transactions on Parallel and Distributed Systems, 4(4):446–457, April 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Linder:1994:AGM

[LH94] Daniel H. Linder and James C. Harden. Access graphs: a model for investigating memory consistency. IEEE Transactions on Parallel and Distributed Systems, 5(1):39–52, January 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2001:PSA

[LH01] T.-J. Liu and S.-T. Huang. Phase synchronization on asynchronous uniform rings with odd size. IEEE Transactions on Parallel and Distributed Systems, 12(6):638–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0638.
pdf;   http://www.computer.
org/tpds/td2001/l0638abs.
htm.

Lin:2003:EFC

[LH03] Ji-Cherng Lin and Tetz C. Huang. An efficient faultcontaining self-stabilizing algorithm for finding a maximal independent set. IEEE Transactions on Parallel and Distributed Systems, 14(8):742–754, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0742abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/08/l0742.
pdf.

Lee:2005:AAE

[LH05] SingLing Lee and Hann-Jang Ho. A 1.5 approximation algorithm for embedding hyperedges in a cycle. IEEE Transactions on Parallel and Distributed Systems, 16(6):481–488, June 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2006:LFT

[LH06a] Ning Li and Jennifer C. Hou.

Localized fault-tolerant topology control in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 17(4):307–320, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2006:TQS

[LH06b] Xiaowen Liu and Martin Haenggi.

Toward quasiregular sensor networks: Topology control algorithms for improved energy efficiency. IEEE Transactions on Parallel and Distributed Systems, 17(9):975–986, September 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2008:PTZ

[LHL^{+}08] Li Lu, Jinsong Han, Yunhao Liu, Lei Hu, Jin-Peng Huai, Lionel Ni, and Jian Ma. Pseudo trust: Zero-knowledge authentication in anonymous P2Ps. IEEE Transactions on Parallel and Distributed Systems, 19(10): 1325–1337, October 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Loh:2005:EH

[LHP05] Peter K. K. Loh, Wen Jing Hsu, and Yi Pan. The exchanged hypercube. IEEE Transactions on Parallel and Distributed Systems, 16(9):866–874, September 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See comment [Che07].

Lee:1992:HSP

[LHS92] Ben Lee, A. R. Hurson, and Behrooz Shirazi. A hybrid scheme for processing data structures in a dataflow environment. IEEE Transactions on Parallel and Distributed Systems, 3(1):83–96, January 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2003:PCE

[LHS03] Yu-Hua Lee, Shi-Jinn Horng, and Jennifer Seitzer. Parallel computation of the Euclidean distance transform on a three-dimensional image array. IEEE Transactions on Parallel and Distributed Systems, 14(3):203–212, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0203abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0203.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0203.
pdf.

Libeskind-Hadas:1995:ORA

[LHSML95] Ran Libeskind-Hadas, Nimish Shrivastava, Rami G. Melhem, and C. L. Liu. Optimal reconfiguration algorithms for real-time fault-tolerant processor arrays. IEEE Transactions on Parallel and Distributed Systems, 6 (5):498–510, May 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0498abs.htm.

Li:2003:IMD

[Li03] Keqin Li. Improved methods for divisible load distribution on k-dimensional meshes using pipelined communications. IEEE Transactions on Parallel and Distributed Systems, 14(12):1250–1261, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1250abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1250.
pdf.

Li:2007:APA

[Li07] Keqin Li. Analysis of parallel algorithms for matrix chain product and matrix powers on distributed memory systems. IEEE Transactions on Parallel and Distributed Systems, 18(7):865– 878, July 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2008:PAP

[Li08] Keqin Li. Performance analysis of power-aware task scheduling algorithms on multiprocessor computers with dynamic voltage and speed. IEEE Transactions on Parallel and Distributed Systems, 19(11):1484–1497, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lilja:1994:IPL

[Lil94] David J. Lilja. Impact of parallel loop scheduling strategies on prefetching in a shared memory multiprocessor. IEEE Transactions on Parallel and Distributed Systems, 5(6):573–584, June 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:1993:BSL

[Lin93] Yen-Chun Lin. On balancing sorting on a linear array. IEEE Transactions on Parallel and Distributed Systems, 4(5):566– 571, May 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2008:THS

[Liu08] Yunhao Liu. A two-hop solution to solving topology mismatch. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1591–1600, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2008:IID

[LJCL08] Sam C. M. Lee, Joe W. J.

Jiang, Dah-Ming Chiu Chiu, and John C. S. Lui. Interaction of ISPs: Distributed resource allocation and revenue maximization. IEEE Transactions on Parallel and Distributed Systems, 19(2):204–218, February 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2007:SLS

[LJLN07] Xiaofei Liao, Hai Jin, Yunhao Liu, and Lionel M. Ni. Scalable live streaming service based on interoverlay optimization. IEEE Transactions on Parallel and Distributed Systems, 18 (12):1663–1674, December 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2009:PHT

[LJLS09] Jaejin Lee, Changhee Jung, Daeseob Lim, and Yan Solihin. Prefetching with helper threads for loosely coupled multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 20(9):1309–1324, September 2009. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2005:ABL

[LJW05] Jiming Liu, Xiaolong Jin, and Yuanshi Wang. Agent-based load balancing on homogeneous minigrids: Macroscopic modeling and characterization. IEEE Transactions on Parallel and Distributed Systems, 16(7):586– 598, July 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2007:DEF

[LJW^{+}07] Hai Liu, Xiaohua Jia, Peng-Jun Wan, Xinxin Liu, and Frances F. Yao. A distributed and efficient flooding scheme using 1hop information in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 18(5):658–671, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lal:2004:PEM

[LJZA04] Dhananjay Lal, Vivek Jain, Qing-An Zeng, and Dharma P. Agrawal. Performance evaluation of medium access control for multiple-beam antenna nodes in a wireless LAN. IEEE Transactions on Parallel and Distributed Systems, 15(12):1117–1129, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/12/l1117abs.
htm;   http://csdl.computer.

org/dl/trans/td/2004/12/l1117.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1117.
pdf.

Lee:1990:MNL

[LK90] Pei-Zong Lee and Zvi Meir Kedem. Mapping nested loop algorithms into multidimensional systolic arrays. IEEE Transactions on Parallel and Distributed Systems, 1(1):64–76, January 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:1994:DAO

[LK94] Jie Li and Hisao Kameda. Decomposition algorithm for optimal static load balancing in tree hierarchy network configurations. IEEE Transactions on Parallel and Distributed Systems, 5(5):540–548, May 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lodha:2000:FDM

[LK00] S. Lodha and A. Kshemkalyani.

A fair distributed mutual exclusion algorithm. IEEE Transactions on Parallel and Distributed Systems, 11(6):537–??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0537.
pdf;   http://www.computer.
org/tpds/td2000/l0537abs.
htm.

Louri:2004:OIN

[LK04] Ahmed Louri and Avinash Karanth Kodi. An optical interconnection network and a modified snooping protocol for the design of large-scale symmetric multiprocessors (SMPs). IEEE Transactions on Parallel and Distributed Systems, 15(12):1093–1104, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/12/l1093abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/12/l1093.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/12/l1093.
                                        pdf.

Lee:2007:CFE

[LK07] Manhee Lee and Eun Jung Kim. A comprehensive framework for enhancing security in InfiniBand architecture. IEEE Transactions on Parallel and Distributed Systems, 18(10):1393–1406, October 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lange:1992:JDI

[LKG92] F. Lange, R. Kroeger, and M. Gergeleit. JEWEL: design and implementation of a distributed measurement system. IEEE Transactions on Parallel and Distributed Systems, 3(6): 657–671, November 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2003:PAT

[LKHL03] Heejo Lee, Jong Kim, Sung Je Hong, and Sunggu Lee. Processor allocation and task scheduling of matrix chain products on parallel systems. IEEE Transactions on Parallel and Distributed Systems, 14(4):394–407, April 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0394abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0394.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0394.
pdf.

Lee:1995:AFB

[LKK95] Gyungho Lee, Byung-Chang Kang, and Richard Y. Kain. Analysis of finite buffered multistage combining networks. IEEE Transactions on Parallel and Distributed Systems, 6 (7):760–766, July 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0760abs.htm.

Lakamraju:2002:FRG

[LKK02] Vijay Lakamraju, Israel Koren, and C. M. Krishna. Filtering random graphs to synthesize interconnection networks with multiple objectives. IEEE Transactions on Parallel and Distributed Systems, 13(11):1139–1149, November 2002. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2002/11/l1139abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1139.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1139.
                                        pdf.

Lee:2005:BOI

[LKKS05] Ju Yong Lee, Sunggon Kim, Deokseong Kim, and Dan Keun Sung. Bandwidth optimization for Internet traffic in generalized processor sharing servers. IEEE Transactions on Parallel and Distributed Systems, 16 (4):324–334, April 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:1990:ELP

[LL90] Yi-Bing Lin and Edward D. Lazowska. Exploiting lookahead in parallel simulation. IEEE Transactions on Parallel and Distributed Systems, 1(4):457–469, October 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:1994:PAL

[LL94] Mi Lu and Hua Lin. Parallel algorithms for the longest common subsequence problem. IEEE Transactions on Parallel and Distributed Systems, 5(8): 835–848, August 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1996:SAP

[LL96] Soo-Young Lee and Kyung Geun Lee. Synchronous and asynchronous parallel simulated annealing with multiple Markov chains. IEEE Transactions on Parallel and Distributed Systems, 7(10):993–1008, October 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0993abs.htm.

Lundberg:1998:URV

[LL98] L. Lundberg and H. Lennerstad.

Using recorded values for bounding the minimum completion time in multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 9(4):346– ??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0346.
pdf;   http://www.computer.
org/tpds/td1998/l0346abs.
htm.

Lee:2002:ARD

[LL02] Jack Y. B. Lee and John C. S. Lui. Automatic recovery from disk failure in continuousmedia servers. IEEE Transactions on Parallel and Distributed Systems, 13(5):499–515, May 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0499.
pdf;   http://www.computer.

                                        org/tpds/td2002/l0499abs.
                                        htm.

Leung:2006:GLSa

[LL06a] Ka-Cheong Leung and Victor O. K. Li. Generalized load sharing for packet-switching networks I: Theory and packetbased algorithm. IEEE Transactions on Parallel and Distributed Systems, 17(7):694–702, July 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leung:2006:GLSb

[LL06b] Ka-Cheong Leung and Victor O. K. Li. Generalized load sharing for packet-switching networks II: Flow-based algorithms. IEEE Transactions on Parallel and Distributed Systems, 17 (7):703–712, July 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2007:NOT

[LL07] Rui Li and Du Li. A new operational transformation framework for real-time group editors. IEEE Transactions on Parallel and Distributed Systems, 18 (3):307–319, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2006:FCA

[LLA^{+}06] Chenyang Lu, Ying Lu, Tarek F. Abdelzaher, John A. Stankovic, and Sang Hyuk Son. Feedback control architecture and design methodology for service delay guarantees in Web servers.

IEEE Transactions on Parallel and Distributed Systems, 17 (9):1014–1027, September 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2009:FAR

[LLGS09] Yawei Li, Zhiling Lan, Prashasta Gujrati, and Xian-He Sun. Fault-aware runtime strategies for high-performance computing. IEEE Transactions on Parallel and Distributed Systems, 20 (4):460–473, April 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2001:APR

[LLH^{+}01] W.-S. Lin, R. W. H. Lau, K. Hwang, X. Lin, and P. Y. S. Cheung. Adaptive parallel rendering on multiprocessors and workstation clusters. IEEE Transactions on Parallel and Distributed Systems, 12(3):241– ??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0241.
pdf;   http://www.computer.
org/tpds/td2001/l0241abs.
htm.

Lenoski:1993:DPL

[LLJ^{+}93] Daniel Lenoski, James Laudon, Truman Joe, David Nakahira, Luis Stevens, Anoop Gupta, and John Hennessy. The DASH prototype: Logic overhead and performance. IEEE Transactions on Parallel and Distributed Systems, 4(1):41–61, January 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2009:PRL

[LLL09] Yunfeng Lin, Ben Liang, and Baochun Li. Priority random linear codes in distributed storage systems. IEEE Transactions on Parallel and Distributed Systems, 20(11):1653–1667, November 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2006:SDH

[LLS06] Tong Li, Alvin R. Lebeck, and Daniel J. Sorin. Spin detection hardware for improved management of multithreaded systems. IEEE Transactions on Parallel and Distributed Systems, 17 (6):508–521, June 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2008:SSW

[LLSZ08] Mei Li, Wang-Chien Lee, Anand Sivasubramaniam, and Jing Zhao. SSW: a small-world-based overlay for peer-to-peer search. IEEE Transactions on Parallel and Distributed Systems, 19 (6):735–749, June 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lam:2008:NMS

[LLTW08] Tak-Wah Lam, Lap-Kei Lee, Isaac K. K. To, and Prudence W. H. Wong. Nonmigratory multiprocessor scheduling for response time and energy. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1527–1539, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2009:DSA

[LLWC09] Tsungnan Lin, Pochiang Lin, Hsinping Wang, and Chiahung Chen. Dynamic search algorithm in unstructured peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 20(5):654–666, May 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2006:IQR

[LLX06] Xiaomei Liu, Yunhao Liu, and Li Xiao. Improving query response delivery quality in peerto-peer systems. IEEE Transactions on Parallel and Distributed Systems, 17(11):1335– 1347, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2004:PDD

[LLY04] Sam C. M. Lee, John C. S.

Lui, and David K. Y. Yau. A proportional-delay DiffServenabled Web server: Admission control and dynamic adaptation. IEEE Transactions on Parallel and Distributed Systems, 15(5):385–400, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/

                                        trans/td/2004/05/l0385abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/05/l0385.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/05/l0385.
                                        pdf.

Law:2005:YCR

[LLY05] Terence K. T. Law, John C. S. Lui, and David K. Y. Yau. You can run, but you can’t hide: An effective statistical methodology to trace back DDoS attackers. IEEE Transactions on Parallel and Distributed Systems, 16(9): 799–813, September 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leung:2007:OPR

[LLY07] Ka-Cheong Leung, Victor O. K. Li, and Daiqin Yang. An overview of packet reordering in transmission control protocol (TCP): Problems, solutions, and challenges. IEEE Transactions on Parallel and Distributed Systems, 18(4):522–535, April 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2006:NPB

[LM06] Jaehwan John Lee and Vincent John Mooney. A novel O(n) parallel banker’s algorithm for system-on-a-chip. IEEE Transactions on Parallel and Distributed Systems, 17(12):1377– 1389, December 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:1994:DFM

[LMN94] Xiaola Lin, Philip K. McKinley, and Lionel M. Ni. Deadlockfree multicast wormhole routing in 2-D mesh multicomputers. IEEE Transactions on Parallel and Distributed Systems, 5 (8):793–804, August 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:1995:MFM

[LMN95] Xiaola Lin, Philip K. McKinley, and Lionel M. Ni. The message flow model for routing in wormhole-routed networks. IEEE Transactions on Parallel and Distributed Systems, 6 (7):755–760, July 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0755abs.htm.

Lenzini:2004:EBR

[LMS04] Luciano Lenzini, Enzo Mingozzi, and Giovanni Stea. Eligibility-based round robin for fair and efficient packet scheduling in wormhole switching networks. IEEE Transactions on Parallel and Distributed Systems, 15(3):244–256, March 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/03/l0244abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0244.
htm;   http://csdl.computer.

                                        org/dl/trans/td/2004/03/l0244.
                                        pdf.

Lui:1998:CPB

[LMT98] J. C. S. Lui, R. R. Muntz, and D. Towsley. Computing performance bounds of fork-join parallel programs under a multiprocessing environment. IEEE Transactions on Parallel and Distributed Systems, 9(3):295– ??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0295.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0295abs.
                                        htm.

Lin:1993:MCM

[LN93] Xiaola Lin and Lionel M. Ni. Multicast communication in multicomputer networks. IEEE Transactions on Parallel and Distributed Systems, 4(10):1105– 1117, October 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2000:SHA

[LNO^{+}00] R. Lin, K. Nakano, S. Olariu, M. C. Pinotti, J. L. Schwing, and A. Y. Zomaya. Scalable hardware-algorithms for binary prefix sums. IEEE Transactions on Parallel and Distributed Systems, 11(8):838–??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0838.

pdf;   http://www.computer.
org/tpds/td2000/l0838abs.
htm.

Lin:2003:EPP

[LNOZ03] Rong Lin, Koji Nakano, Stephan Olariu, and Albert Y. Zomaya. An efficient parallel prefix sums architecture with domino logic. IEEE Transactions on Parallel and Distributed Systems, 14(9):922–931, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0922abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0922.
pdf.

Li:1994:LLC

[LNP94] Kai Li, Jeffrey F. Naughton, and James S. Plank. Lowlatency, concurrent checkpointing for parallel programs. IEEE Transactions on Parallel and Distributed Systems, 5(8):874– 879, August 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2003:OBG

[LNYY03] Dongman Lee, Dukyun Nam, Hee Yong Youn, and Chansu Yu. OCI-based group communication support in CORBA. IEEE Transactions on Parallel and Distributed Systems, 14(11):1126–1139, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/11/l1126abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/11/l1126.
                                        pdf.

Lee:1995:FPA

[LO95a] Ching-Yi Lee and A. Yavuz Oruç. A fast parallel algorithm for routing unicast assignments in Benes networks. IEEE Transactions on Parallel and Distributed Systems, 6 (3):329–334, March 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0329abs.htm.

Lin:1995:RBS

[LO95b] Rong Lin and Stephan Olariu. Reconfigurable buses with shift switching: Concepts and applications. IEEE Transactions on Parallel and Distributed Systems, 6(1):93–102, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lopriore:2002:ACM

[Lop02]  Lanfranco Lopriore.  Access

control mechanisms in a distributed, persistent memory system. IEEE Transactions on Parallel and Distributed Systems, 13(10):1066–1083, October 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l1066.

pdf;   http://www.computer.
org/tpds/td2002/l1066abs.
htm.

Lin:1999:MHB

[LOSW99] R. Lin, S. Olariu, J. L. Schwing, and B.-F. Wang. The mesh with hybrid buses: An efficient parallel architecture for digital geometry. IEEE Transactions on Parallel and Distributed Systems, 10(3):266–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0266.
pdf;   http://www.computer.
org/tpds/td1999/l0266abs.
htm.

Lee:1996:SOP

[LP96] Chiung-San Lee and Tai-Ming Parng. A subsystem-oriented performance analysis methodology for shared-bus multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 7 (7):755–767, July 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0755abs.htm.

Li:2007:HRM

[LP07] Xiaolin Li and Manish Parashar.

Hybrid runtime management of space-time heterogeneity for parallel structured adaptive applications. IEEE Transactions on Parallel and Distributed Systems, 18(9):1202–1214, September 2007. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Lysne:2005:PIM

[LPD05] Olav Lysne, Timothy Mark Pinkston, and Jose Duato. Part II: a methodology for developing deadlock-free dynamic network reconfiguration processes. IEEE Transactions on Parallel and Distributed Systems, 16 (5):428–443, May 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lo:1999:SDR

[LPE^{+}99] J. L. Lo, S. S. Parekh, S. J. Eggers, H. M. Levy, and D. M. Tullsen. Software-directed register deallocation for simultaneous multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 10(9):922–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0922.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0922abs.
                                        htm.

Li:1998:FPE

[LPZ98] K. Li, Y. Pan, and S. Q. Zheng. Fast and processor efficient parallel matrix multiplication algorithms on a linear array with a reconfigurable pipelined bus system. IEEE Transactions on Parallel and Distributed Systems, 9(8):705– ??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0705.
pdf;   http://www.computer.
org/tpds/td1998/l0705abs.
htm.

Luo:2009:DBI

[LQZ09] Hongbin Luo, Yajuan Qin, and Hongke Zhang. A DHT-based identifier-to-locator mapping approach for a scalable Internet. IEEE Transactions on Parallel and Distributed Systems, 20(12): 1790–1802, December 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liestman:1993:NCE

[LR93] Arthur L. Liestman and Dana Richards. Network communication in edge-colored graphs: gossiping. IEEE Transactions on Parallel and Distributed Systems, 4(4):438–445, April 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:1996:AAJ

[LR96] Hwa-Chun Lin and C. S.

Raghavendra. An approximate analysis of the join the shortest queue (JSQ) policy. IEEE Transactions on Parallel and Distributed Systems, 7 (3):301–307, March 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0301abs.htm.

Ligon:1997:TMR

[LR97] Walter B. Ligon III and Umakishore Ramachandran. Toward a more realistic performance evaluation of interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 8(7):681–694, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0681.
pdf;   http://www.computer.
org/tpds/td1997/l0681abs.
htm.

Li:1999:SEC

[LRG99] Z. Li, J. H. Reif, and S. K. S. Gupta. Synthesizing efficient out-of-core programs for block recursive algorithms using blockcyclic data distributions. IEEE Transactions on Parallel and Distributed Systems, 10(3):297– ??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0297.
pdf;   http://www.computer.
org/tpds/td1999/l0297abs.
htm.

Legrand:2004:MLB

[LRRV04] Arnaud Legrand, Hélène Renard, Yves Robert, and Frédéric Vivien. Mapping and loadbalancing iterative computations. IEEE Transactions on Parallel and Distributed Systems, 15(6):546–558, June 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0546.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0546.pdf.

Lindsey:2002:DGA

[LRS02] Stephanie Lindsey, Cauligi Raghavendra, and Krishna M. Sivalingam. Data gathering algorithms in sensor networks using energy metrics. IEEE Transactions on Parallel and Distributed Systems, 13(9):924–935, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0924.
pdf;   http://www.computer.
org/tpds/td2002/l0924abs.
htm.

Lo:1996:PDC

[LRTZ96] Virginia Lo, Sanjay Rajopadhye, Jan Arne Telle, and Xiaoxiong Zhong. Parallel divide and conquer on meshes. IEEE Transactions on Parallel and Distributed Systems, 7 (10):1049–1058, October 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l1049abs.htm.

Lebeck:1994:RCM

[LS94a] Alvin R. Lebeck and Gurindar S.

Sohi. Request combining in multiprocessors with arbitrary interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 5(11):1140– 1155, November 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1994:IAA

[LS94b] Sunggu Lee and Kang G. Shin. Interleaved all-to-all reliable broadcast on meshes and hypercubes. IEEE Transactions on Parallel and Distributed Systems, 5(5):449–458, May 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1994:PDM

[LS94c] Sunggu Lee and Kang G. Shin. On probabilistic diagnosis of multiprocessor systems using multiple syndromes. IEEE Transactions on Parallel and Distributed Systems, 5(6):630– 638, June 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lai:1996:CEM

[LS96] Ten H. Lai and Ming-Jye Sheng. Constructing Euclidean minimum spanning trees and all nearest neighbors on reconfigurable meshes. IEEE Transactions on Parallel and Distributed Systems, 7(8):806–817, August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1996/
                                        l0806abs.htm.

Lee:1997:OTA

[LS97] Cheol-Hoon Lee and Kang G. Shin. Optimal task assignment in homogeneous networks. IEEE Transactions on Parallel and Distributed Systems, 8(2):119–129, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0119.
pdf;   http://www.computer.
org/tpds/td1997/l0119abs.
htm.

Liu:2006:PPB

[LS06] Weiguo Liu and Bertil Schmidt.

Parallel pattern-based systems for computational biology: a case study. IEEE Transactions on Parallel and Distributed Systems, 17(8):750–763, August 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Laoutaris:2007:DSC

[LSB^{+}07] Nikolaos Laoutaris, Georgios Smaragdakis, Azer Bestavros, Ibrahim Matta, and Ioannis Stavrakakis. Distributed selfish caching. IEEE Transactions on Parallel and Distributed Systems, 18(10):1361–1376, October 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lippert:1998:HSP

[LSBS98] T. Lippert, A. Seyfried, A. Bode, and K. Schilling. Hyper-systolic parallel computing. IEEE Transactions on Parallel and Distributed Systems, 9(2):97–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0097.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0097abs.
                                        htm.

Liu:1995:PCP

[LSC95] Jyh-Charn Liu, Kang G. Shin, and Charles C. Chang. Prevention of congestion in packetswitched multistage interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 6(5):535–541, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1995/
                                        l0535abs.htm.

Li:2007:MOP

[LSCZ07] Keqiu Li, Hong Shen, Francis Y. L. Chin, and Weishi Zhang. Multimedia object placement for transparent data replication. IEEE Transactions on Parallel and Distributed Systems, 18(2): 212–224, February 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2009:HEU

[LSF^{+}09] Pan Li, Nicola Scalabrino, Yuguang (Michael) Fang, Enrico Gregori, and Imrich Chlamtac. How to effectively use multiple channels in wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 20(11):1641–1652, November 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lysne:2006:LRI

[LSRT06] Olav Lysne, Tor Skeie, Sven-Arne Reinemo, and Ingebjorg Theiss. Layered routing in irregular networks. IEEE Transactions on Parallel and Distributed Systems, 17(1):51–65, January 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2007:SAB

[LSVMW07] Weiguo Liu, Bertil Schmidt, Gerrit Voss, and Wolfgang Muller-Wittig. Streaming algorithms for biological sequence alignment on GPUs. IEEE Transactions on Parallel and Distributed Systems, 18(9):1270– 1281, September 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2004:PDT

[LSW04] Xiang-Yang Li, Ivan Stojmenovic, and Yu Wang. Partial Delaunay triangulation and degree limited localized Bluetooth Scatternet formation. IEEE Transactions on Parallel and Distributed Systems, 15(4):350–361, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0350abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0350.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0350.
pdf.

Lee:2009:PDO

[LSZ09] Young Choon Lee, Riky Subrata, and Albert Y. Zomaya. On the performance of a dualobjective optimization model for workflow applications on grid platforms. IEEE Transactions on Parallel and Distributed Systems, 20(9):1273–1284, September 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leung:1997:OAG

[LT97] Ho-Fung Leung and Hing-Fung Ting. An optimal algorithm for global termination detection in shared-memory asynchronous multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 8(5):538– 543, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0538.
pdf;   http://www.computer.
org/tpds/td1997/l0538abs.
htm.

Liang:2000:PPP

[LT00] D.-R. Liang and S. K. Tripathi. On performance prediction of parallel computations with precedent constraints. IEEE Transactions on Parallel and Distributed Systems, 11(5):491– ??, May 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0491.
pdf;   http://www.computer.

org/tpds/td2000/l0491abs.
htm.

Laoutaris:2006:DSR

[LTZS06] Nikolaos Laoutaris, Orestis Telelis, Vassilios Zissimopoulos, and Ioannis Stavrakakis. Distributed selfish replication. IEEE Transactions on Parallel and Distributed Systems, 17(12): 1401–1413, December 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lai:1995:RAD

[LW95a] Ten-Hwang Lai and Li-Fen Wu.

An (N − 1)-resilient algorithm for distributed termination detection. IEEE Transactions on Parallel and Distributed Systems, 6(1):63–78, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:1995:RCN

[LW95b] Isaac Yi-Yuan Lee and Sheng-De Wang. Ring-connected networks and their relationship to cubical ring connected cycles and dynamic redundancy networks. IEEE Transactions on Parallel and Distributed Systems, 6(9): 988–996, September 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0988abs.htm.

Li:2009:EMC

[LW09a] Zongpeng Li and Carey Williamson.

Enforcing minimum-cost multicast routing against selfish information flows. IEEE Transactions on Parallel and Distributed Systems, 20(9):1299– 1308, September 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2009:SRC

[LW09b] Cong Liu and Jie Wu. Scalable routing in cyclic mobile networks. IEEE Transactions on Parallel and Distributed Systems, 20(9):1325–1338, September 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2009:VFB

[LW09c] Cong Liu and Jie Wu. Virtualforce-based geometric routing protocol in MANETs. IEEE Transactions on Parallel and Distributed Systems, 20(4):433– 445, April 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2009:REE

[LWC^{+}09] Xiang-Yang Li, Yu Wang, Haiming Chen, Xiaowen Chu, Yanwei Wu, and Yong Qi. Reliable and energy-efficient routing for static wireless ad hoc networks with unreliable links. IEEE Transactions on Parallel and Distributed Systems, 20(10):1408–1421, October 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2006:MLS

[LWJ06] Hai Liu, Pengjun Wan, and Xiaohua Jia. Maximal lifetime scheduling for sensor surveillance systems with K sensors to one target. IEEE Transactions on Parallel and Distributed Systems, 17(12):1526– 1536, December 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Lu:2005:FUC

[LWK05] Chenyang Lu, Xiaorui Wang, and Xenofon Koutsoukos. Feedback utilization control in distributed real-time systems with end-to-end tasks. IEEE Transactions on Parallel and Distributed Systems, 16(6):550– 561, June 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2005/06/
l0550s.pdf.

Lo:1997:NPA

[LWLN97] Virginia Lo, Kurt J. Windisch, Wanqian Liu, and Bill Nitzberg. Noncontiguous processor allocation algorithms for meshconnected multicomputers. IEEE Transactions on Parallel and Distributed Systems, 8(7):712– 726, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0712.
pdf;   http://www.computer.

                                        org/tpds/td1997/l0712abs.
                                        htm.

Louri:1998:SML

[LWN98] A. Louri, B. Weech, and C. Neocleous. A spanning multichannel linked hypercube: a gradually scalable optical interconnection network for massively parallel computing. IEEE Transactions on Parallel and Distributed Systems, 9(5):497–??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0497.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0497abs.
                                        htm.

Liu:2007:EED

[LWP07] Chong Liu, Kui Wu, and Jian Pei. An energy-efficient data collection framework for wireless sensor networks by exploiting spatiotemporal correlation. IEEE Transactions on Parallel and Distributed Systems, 18(7): 1010–1023, July 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2004:ALM

[LWS04] Xiang-Yang Li, Yu Wang, and Wen-Zhan Song. Applications of k-local MST for topology control and broadcasting in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 15(12):1057–1069, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/12/l1057abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1057.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1057.
pdf.

Liu:2006:RCG

[LWXS06] Chong Liu, Kui Wu, Yang Xiao, and Bo Sun. Random coverage with guaranteed connectivity: Joint scheduling for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 17(6):562–575, June 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leff:1993:RAR

[LWY93] Avraham Leff, Joel L. Wolf, and Philip S. Yu. Replication algorithms in a remote caching architecture. IEEE Transactions on Parallel and Distributed Systems, 4(11):1185–1204, November 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leff:1996:ELB

[LWY96] Avraham Leff, Joel L. Wolf, and Philip S. Yu. Efficient LRUbased buffering in a LAN remote caching architecture. IEEE Transactions on Parallel and Distributed Systems, 7(2):191– 206, February 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.

                                        computer.org/tpds/td1996/
                                        l0191abs.htm.

Liu:2005:LAU

[LXL^{+}05] Yunhao Liu, Li Xiao, Xiaomei Liu, Lionel M. Ni, and Xiaodong Zhang. Location awareness in unstructured peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 16(2):163–174, February 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2008:ESC

[LXL08] Zhenyu Li, Gaogang Xie, and Zhongcheng Li. Efficient and scalable consistency maintenance for heterogeneous peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 19(12):1695–1708, December 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2007:BSB

[LXN07] Yunhao Liu, Li Xiao, and Lionel Ni. Building a scalable bipartite P2P overlay network. IEEE Transactions on Parallel and Distributed Systems, 18 (9):1296–1306, September 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lilja:1993:IMU

[LY93a] David J. Lilja and Pen-Chung Yew. Improving memory utilization in cache coherence directories. IEEE Transactions on Parallel and Distributed Systems, 4 (10):1130–1146, October 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:1993:PID

[LY93b] Chengwen Liu and Clement Yu. Performance issues in distributed query processing. IEEE Transactions on Parallel and Distributed Systems, 4(8):889– 905, August 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leff:1994:PSR

[LY94] Avraham Leff and Philip S. Yu.

Performance study of robust distributed load sharing strategies. IEEE Transactions on Parallel and Distributed Systems, 5(12): 1286–1301, December 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2008:FRB

[LYW08] Minglu Li, Jiadi Yu, and Jie Wu. Free-riding on BitTorrentlike peer-to-peer file sharing systems: Modeling analysis and improvement. IEEE Transactions on Parallel and Distributed Systems, 19(7):954–966, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:1990:EDD

[LYZ90] Zhiyuan Li, Pen-Chung Yew, and Chuan-Qi Zhu. Efficient data dependence analysis for parallelizing compilers. IEEE Transactions on Parallel and Distributed Systems, 1(1):26–34, January 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lau:2002:FGS

[LZ02] Francis C. M. Lau and Shi-Heng Zhang. Fast gossiping in square meshes/tori with bounded-size packets. IEEE Transactions on Parallel and Distributed Systems, 13(4):349– 358, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0349.
pdf;   http://www.computer.
org/tpds/td2002/l0349abs.
htm.

Lu:2005:PRA

[LZ05] Enyue Lu and S. Q. Zheng. Parallel routing algorithms for nonblocking electronic and photonic switching networks. IEEE Transactions on Parallel and Distributed Systems, 16(8):702– 713, August 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2008:NST

[LZ08] Young-Choon Lee and Albert Zomaya. A novel state transition method for metaheuristicbased scheduling in heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 19(9):1215– 1223, September 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2009:PAC

[LZR09] Kang Li, Zhenyu Zhong, and Lakshmish Ramaswamy. Privacyaware collaborative spam filtering. IEEE Transactions on Parallel and Distributed Systems, 20 (5):725–739, May 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2009:TDA

[LZTY09] Jian-Guang Luo, Qian Zhang, Yun Tang, and Shi-Qiang Yang. A trace-driven approach to evaluate the scalability of P2P-based video-on-demand service. IEEE Transactions on Parallel and Distributed Systems, 20(1):59– 70, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Manjikian:1997:FLP

[MA97] N. Manjikian and Tarek S.

Abdelrahman. Fusion of loops for parallelism and locality. IEEE Transactions on Parallel and Distributed Systems, 8(2):193–209, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0193.
pdf;   http://www.computer.
org/tpds/td1997/l0193abs.
htm.

Manjikian:2001:EWP

[MA01] N. Manjikian and T. S. Abdelrahman. Exploiting wavefront parallelism on largescale shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 12(3):259–??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0259.
pdf;   http://www.computer.
org/tpds/td2001/l0259abs.
htm.

Mostafa:2007:RPM

[MAJ^{+}07] M. Mostafa, A. Azim, Xiaohong Jiang, Pin-Han Ho, Susumu Horiguchi, and Minyi Guo. Restoration probability modelling for active restorationbased optical networks with correlation among backup routes. IEEE Transactions on Parallel and Distributed Systems, 18(11): 1592–1606, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Maresca:1993:PPA

[Mar93] Massimo Maresca. Polymorphic processor arrays. IEEE Transactions on Parallel and Distributed Systems, 4(5):490–506, May 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martinez:2007:NCE

[MAS^{+}07] Alejandro Martínez, Francisco J. Alfaro, José L. Sánchez, Francisco J. Quiles, and José Duato. A new cost-effective technique for QoS support in clusters. IEEE Transactions on Parallel and Distributed Systems, 18 (12):1714–1726, December 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martinez:2008:FPQ

[MAS08] Raúl Martínez, Francisco J. Alfaro, and José L. Sánchez. A framework to provide quality of service over advanced switching. IEEE Transactions on Parallel and Distributed Systems, 19(8): 1111–1123, August 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Min:1992:DAS

[MB92] Sang Lyul Min and Jean-Loup Baer. Design and analysis of a scalable cache coherence scheme based on clocks and timestamps. IEEE Transactions on Parallel and Distributed Systems, 3 (1):25–44, January 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Malluhi:1994:HHN

[MB94] Qutaibah M. Malluhi and Magdy A. Bayoumi. Hierarchical hypercube: a new interconnection topology for massively parallel systems. IEEE Transactions on Parallel and Distributed Systems, 5(1):17–30, January 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Madan:2007:PEA

[MB07] Niti Madan and Rajeev Balasubramonian. Power efficient approaches to redundant multithreading. IEEE Transactions on Parallel and Distributed Systems, 18(8):1066–1079, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Murthy:1998:NAB

[MBM98] K. N. Balasubramanya Murthy, K. Bhuvaneswari, and C. Siva Ram Murthy. A new algorithm based on Givens rotations for solving linear equations on fault-tolerant mesh-connected processors. IEEE Transactions on Parallel and Distributed Systems, 9(8):825–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0825.
pdf;   http://www.computer.
org/tpds/td1998/l0825abs.
htm.

Marchetti:2006:FDT

[MBTPV06] Carlo Marchetti, Roberto Baldoni, Sara Tucci-Piergiovanni, and Antonino Virgillito. Fully distributed three-tier active software replication. IEEE Transactions on Parallel and Distributed Systems, 17(7):633–645, July 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

May:2002:HCN

[MBW02] Phil May, Santithorn Bunchua, and D. Scott Wills. HiPER: a compact narrow channel router with hop-by-hop error correction. IEEE Transactions on Parallel and Distributed Systems, 13(5):485–498, May 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0485.
pdf;   http://www.computer.
org/tpds/td2002/l0485abs.
htm.

Meliksetian:1993:ORA

[MC93] Dikran S. Meliksetian and C. Y. Roger Chen. Optimal routing algorithm and the diameter of the cube-connected cycles. IEEE Transactions on Parallel and Distributed Systems, 4(10): 1172–1178, October 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moon:1995:GMB

[MC95] Soo-Mook Moon and Scott D.

Carson. Generalized multiway branch unit for VLIW microprocessors. IEEE Transactions on Parallel and Distributed Systems, 6(8):850–862, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0850abs.htm.

Monchiero:2008:PPT

[MCG08] Matteo Monchiero, Ramon Canal, and Antonio González. Power/performance/thermal designspace exploration for multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 19(5):666–681, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Miller:1990:ISG

[MCH^{+}90] Barton P. Miller, Morgan Clark, Jeff Hollingsworth, Steven Kierstead, Sek-See Lim, and Timothy Torzewski. IPS-2: The second generation of a parallel program measurement system. IEEE Transactions on Parallel and Distributed Systems, 1 (2):206–217, April 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

McKinley:1998:COA

[McK98] K. S. McKinley. A compiler optimization algorithm for sharedmemory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 9(8):769– ??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0769.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0769abs.
                                        htm.

Ma:2007:ISP

[MCL^{+}07] Liran Ma, Xiuzhen Cheng, Fang Liu, Fengguang An, and Jose Rivera. iPAK: An in-situ pairwise key bootstrapping scheme for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(8): 1174–1184, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Mohapatra:1996:PAF

[MD96] Prasant Mohapatra and Chita R.

Das. Performance analysis of finite-buffered asynchronous multistage interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 7(1):18–25, January 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahapatra:1997:SGL

[MD97] Nihar R. Mahapatra and Shantanu Dutt. Scalable global and local hashing strategies for duplicate pruning in parallel A* graph search. IEEE Transactions on Parallel and Distributed Systems, 8(7):738– 756, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0738.
pdf;   http://www.computer.
org/tpds/td1997/l0738abs.
htm.

Mueller:2006:HPD

[MDL06] Christopher Mueller, Mehmet M.

Dalkilic, and Andrew Lumsdaine. High-performance direct pairwise comparison of large genomic sequences. IEEE Transactions on Parallel and Distributed Systems, 17(8):764–772, August 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mamidisetty:2009:MDR

[MDSS09] Kranthi K. Mamidisetty, Minlan Duan, Shivakumar Sastry, and P. S. Sastry. Multipath dissemination in regular mesh topologies. IEEE Transactions on Parallel and Distributed Systems, 20(8):1188–1201, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahgoub:1992:PAG

[ME92] Imadeldin O. Mahgoub and A. K. Elmagarmid. Performance analysis of a generalized class of m-level hierarchical multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 3(2):129–138, March 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahgoub:1993:CCP

[ME93] I. O. Mahgoub and A. K. Elmagarmid. Clarifications and corrections to ‘Performance analysis of a generalized class of mlevel hierarchical multiprocessor systems’ (Mar 1992 129–138). IEEE Transactions on Parallel and Distributed Systems, 4 (6):719–720, June 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mantharam:1995:RTZ

[ME95] Mythili Mantharam and P. J. Eberlein. Real two-zero algorithm: a parallel algorithm to reduce a real matrix to a real Schur form. IEEE Transactions on Parallel and Distributed Systems, 6(1):48–62, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Malla:2003:FRA

[MEKOT03] Anjlica Malla, Mona El-Kadi, Stephan Olariu, and Petia Todorova. A fair resource allocation protocol for multimedia wireless networks. IEEE Transactions on Parallel and Distributed Systems, 14(1):63– 71, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2003/01/
l0063abs.htm;   http://csdl.
computer.org/dl/trans/td/
2003/01/l0063.htm;   http:
//csdl.computer.org/dl/trans/
td/2003/01/l0063.pdf.

Menon:2005:AFD

[Men05] Syam Menon. Allocating fragments in distributed databases. IEEE Transactions on Parallel and Distributed Systems, 16 (7):577–585, July 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Myoupo:1996:MSL

[MF96] Jean Frédéric Myoupo and Anne Cécile Fabret. A modular systolic linearization of the Warshall-Floyd algorithm. IEEE Transactions on Parallel and Distributed Systems, 7 (5):449–455, May 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0449abs.htm.

Moritz:2001:LMN

[MF01a] C. A. Moritz and M. I. Frank. LoGPC: Modeling network contention in messagepassing programs. IEEE Transactions on Parallel and Distributed Systems, 12(4):404–??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0404.
pdf;   http://www.computer.
org/tpds/td2001/l0404abs.
htm.

Mualem:2001:UPW

[MF01b] A. W. Mu’alem and D. G. Feitelson. Utilization, predictability, workloads, and user runtime estimates in scheduling the IBM SP2 with backfilling. IEEE Transactions on Parallel and Distributed Systems, 12(6):529– ??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0529.
pdf;   http://www.computer.
org/tpds/td2001/l0529abs.
htm.

Mishra:2001:GCS

[MFLX01] S. Mishra, L. Fei, X. Lin, and G. Xing. On group communication support in CORBA. IEEE Transactions on Parallel and Distributed Systems, 12(2): 193–??, February 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l0193.
pdf;   http://www.computer.
org/tpds/td2001/l0193abs.
htm.

Morales:2009:AOS

[MG09] Ramses Morales and Indranil Gupta. AVMON: Optimal and scalable discovery of consistent availability monitoring overlays for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 20(4):446– 459, April 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martinez:2009:SAG

[MGA^{+}09] Alejandro Martinez, Pedro J.

Garcia, Francisco J. Alfaro, Jose L. Sanchez, Jose Flich, Francisco J. Quiles, and Jose Duato. A switch architecture guaranteeing QoS provision and HOL blocking elimination. IEEE Transactions on Parallel and Distributed Systems, 20(1):13– 24, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2007:ODC

[MGDZ07] Yi Ma, Hongliang Gao, Martin Dimitrov, and Huiyang Zhou. Optimizing dual-core execution for power efficiency and transient-fault recovery. IEEE Transactions on Parallel and Distributed Systems, 18(8):1080– 1093, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Madriles:2008:MSM

[MGQS^{+}08] Carlos Madriles, Carlos García-Quiñones, Jesús Sánchez, Pedro Marcuello, Antonio González, Dean M. Tullsen, Hong Wang, and John P. Shen. Mitosis: a speculative multithreaded processor based on precomputation slices. IEEE Transactions on Parallel and Distributed Systems, 19(7):914–925, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2007:EEL

[MGZN07] Jian Ma, Min Gao, Qian Zhang, and Lionel M. Ni. Energy-efficient localized topology control algorithms in IEEE 802.15.4-based sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18 (5):711–720, May 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Mhamdi:2009:IUM

[Mha09] Lotfi Mhamdi. On the integration of unicast and multicast cell scheduling in buffered crossbar switches. IEEE Transactions on Parallel and Distributed Systems, 20(6):818–830, June 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Michael:2004:HPS

[Mic04] Maged M. Michael. Hazard pointers: Safe memory reclamation for lock-free objects. IEEE Transactions on Parallel and Distributed Systems, 15 (6):491–504, June 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0491.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0491.pdf.

Mitzenmacher:2000:HUO

[Mit00] M. Mitzenmacher. How useful is old information? IEEE Transactions on Parallel and Distributed Systems, 11(1):6–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0006.
pdf;   http://www.computer.
org/tpds/td2000/l0006abs.
htm.

Mitzenmacher:2001:PTC

[Mit01] Michael Mitzenmacher. The power of two choices in randomized load balancing. IEEE Transactions on Parallel and Distributed Systems, 12(10): 1094–1104, October 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l1094.

                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1094abs.
                                        htm.

Misic:1994:CAS

[MJ94] Jelena Misic and Zoran Jovanovic. Communication aspects of the star graph interconnection network. IEEE Transactions on Parallel and Distributed Systems, 5(7):678–687, July 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Malluhi:1998:CHA

[MJ98] Q. M. Malluhi and W. E. Johnston. Coding for high availability of a distributed-parallel storage system. IEEE Transactions on Parallel and Distributed Systems, 9(12):237–??, December 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1237.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1237abs.
                                        htm.

Muthukumar:2006:YSG

[MJ06] R. M. Muthukumar and D. Janakiram. Yama: a scalable generational garbage collector for Java in multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 17(2): 148–159, February 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Marcelin-Jimenez:2006:CSF

[MJRS06] Ricardo Marcelin-Jimenez, Sergio Rajsbaum, and Brett Stevens. Cyclic storage for faulttolerant distributed executions. IEEE Transactions on Parallel and Distributed Systems, 17 (9):1028–1036, September 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mohr:1991:LTC

[MKH91] Eric Mohr, David A. Kranz, and Robert H. Halstead, Jr. Lazy task creation: a technique for increasing the granularity of parallel programs. IEEE Transactions on Parallel and Distributed Systems, 2(3):264–280, July 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2000:PES

[MKR00] G. Ma, A. Khaleel, and A. L. N. Reddy. Performance evaluation of storage systems based on networkattached disks. IEEE Transactions on Parallel and Distributed Systems, 11(9):956–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0956.
pdf;   http://www.computer.
org/tpds/td2000/l0956abs.
htm.

Matsutani:2009:FHT

[MKY^{+}09] Hiroki Matsutani, Michihiro Koibuchi, Yutaka Yamada, D. Frank Hsu, and Hideharu Amano. Fat H-tree: a costefficient tree-based on-chip network. IEEE Transactions on Parallel and Distributed Systems, 20(8):1126–1141, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mak:1990:PPP

[ML90] Victor W. Mak and Stephen F. Lundstrom. Predicting performance of parallel computations. IEEE Transactions on Parallel and Distributed Systems, 1 (3):257–270, July 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Markatos:1994:UPA

[ML94] Evangelos P. Markatos and Thomas J. LeBlanc. Using processor affinity in loop scheduling on shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 5 (4):379–400, April 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moldovan:1992:SMP

[MLL92] Dan Moldovan, Wing Lee, and Changhwa Lin. SNAP: a market-propagation architecture for knowledge processing. IEEE Transactions on Parallel and Distributed Systems, 3(4):397– 410, July 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Malloy:1994:SDA

[MLS94] Brian A. Malloy, Errol L. Lloyd, and Mary Lou Soffa. Scheduling DAG’s for asynchronous multiprocessor execution. IEEE Transactions on Parallel and Distributed Systems, 5(5):498– 508, May 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Myung:2007:TSA

[MLSS07] Jihoon Myung, Wonjun Lee, Jaideep Srivastava, and Timothy K. Shih. Tag-splitting: Adaptive collision arbitration protocols for RFID tag identification. IEEE Transactions on Parallel and Distributed Systems, 18(6):763–775, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2006:HLB

[MLW06] Xiaosong Ma, Jonghyun Lee, and Marianne Winslett. Highlevel buffering for hiding periodic output cost in scientific simulations. IEEE Transactions on Parallel and Distributed Systems, 17(3):193–204, March 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mandelbaum:1996:FEP

[MM96] David M. Mandelbaum and Stefanie G. Mandelbaum. Fast, efficient parallel-acting method of generating functions defined by power series, including logarithm, exponential, and sine, cosine. IEEE Transactions on Parallel and Distributed Systems, 7 (1):33–45, January 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Manimaran:1998:EDS

[MM98a] G. Manimaran and C. S. R. Murthy. An efficient dynamic scheduling algorithm for multiprocessor real-time systems. IEEE Transactions on Parallel and Distributed Systems, 9(3):312–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0312.
pdf;   http://www.computer.
org/tpds/td1998/l0312abs.
htm.

Manimaran:1998:FTD

[MM98b] G. Manimaran and C. Siva Ram Murthy. A fault-tolerant dynamic scheduling algorithm for multiprocessor real-time systems and its analysis. IEEE Transactions on Parallel and Distributed Systems, 9(11):137–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1137.
pdf;   http://www.computer.
org/tpds/td1998/l1137abs.
htm.

Marathe:2007:SCC

[MM07] Jaydeep Marathe and Frank Mueller. Source-code-correlated cache coherence characterization of OpenMP benchmarks. IEEE Transactions on Parallel and Distributed Systems, 18(6):818– 834, June 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mei:2003:SDF

[MMJ03] Alessandro Mei, Luigi V.

Mancini, and Sushil Jajodia. Secure dynamic fragment and replica allocation in large-scale distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 14(9):885–896, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0885abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0885.
pdf.

Moser:1994:PMA

[MMSA94] Louise E. Moser, P. M. Melliar-Smith, and Vivek Agrawala. Processor membership in asynchronous distributed systems. IEEE Transactions on Parallel and Distributed Systems, 5 (5):459–473, May 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Manoj:2006:UMM

[MMSM06] B. S. Manoj, V. Mythili, Ranganath C. Siva, and Ram Murthy. On using multidimensional Markov chains for performance evaluation of hybrid wireless networks. IEEE Transactions on Parallel and Distributed Systems, 17(12):1496–1511, December 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

McMillin:1992:RDS

[MN92] Bruce M. McMillin and Lionel M. Ni. Reliable distributed sorting through the application-oriented fault tolerance paradigm. IEEE Transactions on Parallel and Distributed Systems, 3(4):411–420, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Milward:2004:DIL

[MNM04] Mark Milward, José Luis Núñez, and David Mulvaney. Design and implementation of a lossless parallel high-speed data compression system. IEEE Transactions on Parallel and Distributed Systems, 15(6): 481–490, June 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0481.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0481.pdf.

Manivannan:1997:FCG

[MNS97] D. Manivannan, Robert H. B. Netzer, and Mukesh Singhal. Finding consistent global checkpoints in a distributed computation. IEEE Transactions on Parallel and Distributed Systems, 8(6):623–627, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0623.
pdf;   http://www.computer.
org/tpds/td1997/l0623abs.
htm.

Morillo:2005:IPD

[MOFD05] Pedro Morillo, Juan M. Orduna, Marcos Fernandez, and Jose Duato. Improving the performance of distributed virtual environment systems. IEEE Transactions on Parallel and Distributed Systems, 16(7):637–649, July 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meyer:1991:CDF

[MP91] Fred J. Meyer and Dhiraj K.

Pradhan. Consensus with dual failure modes. IEEE Transactions on Parallel and Distributed Systems, 2(2):214–222, April 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Morin:1997:SRD

[MP97] C. Morin and I. Puaut. A survey of recoverable distributed shared virtual memory systems. IEEE Transactions on Parallel and Distributed Systems, 8(9):959–969, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0959.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0959abs.
                                        htm.

Moore:1997:GEB

[MQ97] J. A. Moore and M. J. Quinn. Generating an efficient broadcast sequence using reflected gray codes. IEEE Transactions on Parallel and Distributed Systems, 8(11):1117–1122, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1117.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1117abs.
                                        htm.

Miguet:1992:ROD

[MR92] Serge Miguet and Yves Robert. Reduction operations on a distributed memory machine with a reconfigurable interconnection network. IEEE Transactions on Parallel and Distributed Systems, 3(4):500–505, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Murthy:1994:ISP

[MR94] P. V. R. Murthy and V. Rajaraman. Implementation of speculative parallelism in functional languages. IEEE Transactions on Parallel and Distributed Systems, 5(11):1197–1205, November 1994. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Madhyastha:2002:LCP

[MR02] Tara M. Madhyastha and Daniel A. Reed. Learning to classify parallel input/output access patterns. IEEE Transactions on Parallel and Distributed Systems, 13(8):802–813, August 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0802.
pdf;   http://www.computer.
org/tpds/td2002/l0802abs.
htm.

Ma:2003:MAS

[MR03] Xiaonan Ma and A. L. Narasimha Reddy. MVSS: An active storage architecture. IEEE Transactions on Parallel and Distributed Systems, 14(10):993–1005, October 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/10/l0993abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/10/l0993.
pdf.

Mei:2006:HCP

[MR06] Alessandro Mei and Romeo Rizzi. Hypercube computations on partitioned optical passive stars networks. IEEE Transactions on Parallel and Distributed Systems, 17(6):497–507, June 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martinez-Rubio:2001:CEA

[MRLD01] Juan M. Martínez-Rubio, Pedro López, and Jose Duato. A cost-effective approach to deadlock handling in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 12(7):716–729, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0716.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0716abs.
                                        htm.

Morillo:2007:LAP

[MROD07] P. Morillo, S. Rueda, J. M. Orduna, and J. Duato. A latencyaware partitioning method for distributed virtual environment systems. IEEE Transactions on Parallel and Distributed Systems, 18(9):1215–1226, September 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mostefaoui:2006:TFT

[MRT06] Achour Mostefaoui, Michel Raynal, and Corentin Travers. Time-free and timer-based assumptions can be combined to obtain eventual leadership. IEEE Transactions on Parallel and Distributed Systems, 17 (7):656–666, July 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mostefaoui:2009:FEP

[MRT09] Achour Mostefaoui, Michel Raynal, and Gilles Tredan. On the fly estimation of the processes that are alive in an asynchronous message-passing system. IEEE Transactions on Parallel and Distributed Systems, 20 (6):778–787, June 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Malony:1992:PMI

[MRW92] Allen D. Malony, Daniel A.

Reed, and Harry A. G. Wijshoff. Performance measurement intrusion and perturbation analysis. IEEE Transactions on Parallel and Distributed Systems, 3(4):433–450, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Madala:1991:PSP

[MS91] Sridhar Madala and James B.

Sinclair. Performance of synchronous parallel algorithms with regular structures. IEEE Transactions on Parallel and Distributed Systems, 2(1):105– 116, January 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mendia:1992:OBS

[MS92] Victor E. Mendia and Dilip Sarkar. Optimal broadcasting on the star graph. IEEE Transactions on Parallel and Distributed Systems, 3(4):389–396, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Makedon:1993:EHP

[MS93] F. Makedon and A. Symvonis. An efficient heuristic for permutation packet routing on meshes with low buffer requirements. IEEE Transactions on Parallel and Distributed Systems, 4(3): 270–276, March 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:1994:KEK

[MS94a] Eva Ma and Dennis G. Shea. E-kernel: An embedding kernel on the IBM Victor V256 multiprocessor for program mapping and network reconfiguration. IEEE Transactions on Parallel and Distributed Systems, 5(9):977–994, September 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahmud:1994:MBA

[MS94b] Syed Masud Mahmud and L. Tissa Samaratunga. Memory bandwidth analysis of hierarchical multiprocessors using model decomposition and steady-state flow analysis. IEEE Transactions on Parallel and Distributed Systems, 5(5):553–560, May 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Manivannan:1999:QSC

[MS99a] D. Manivannan and M. Singhal. Quasi-synchronous checkpointing: Models, characterization, and classification. IEEE Transactions on Parallel and Distributed Systems, 10(7):703– ??, July 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0703.
pdf;   http://www.computer.
org/tpds/td1999/l0703abs.
htm.

Mu:1999:VTS

[MS99b] F. Mu and C. Svensson. Vector transfer by self-tested selfsynchronization for parallel systems. IEEE Transactions on Parallel and Distributed Systems, 10(8):769–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0769.
pdf;   http://www.computer.
org/tpds/td1999/l0769abs.
htm.

Manivannan:2003:EDA

[MS03] D. Manivannan and Mukesh Singhal. An efficient distributed algorithm for detection of knots and cycles in a distributed graph. IEEE Transactions on Parallel and Distributed Systems, 14(10):961–972, October 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/10/l0961abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/10/l0961.
pdf.

Mittal:2007:SCS

[MSG07] Neeraj Mittal, Alper Sen, and Vijay K. Garg. Solving computation slicing using predicate detection. IEEE Transactions on Parallel and Distributed Systems, 18(12):1700– 1713, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2007/12/
                                        l1700s.pdf.

Moon:2000:EAP

[MSH00] S. Moon, B. So, and M. W. Hall. Evaluating automatic parallelization in SUIF. IEEE Transactions on Parallel and Distributed Systems, 11(1):36– ??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0036.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0036abs.
                                        htm.

Misic:2006:PBE

[MSM06] Jelena Misic, Shairmina Shafi, and Vojislav B. Misic. Performance of a beacon enabled IEEE 802.15.4 cluster with downlink and uplink traffic. IEEE Transactions on Parallel and Distributed Systems, 17(4):361–376, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mizrak:2009:DMP

[MSM09] Alper T. Mızrak, Stefan Savage, and Keith Marzullo. Detecting malicious packet losses. IEEE Transactions on Parallel and Distributed Systems, 20(2): 191–206, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Melliar-Smith:1990:BPD

[MSMA90] P. M. Melliar-Smith, Louise E.

Moser, and Vivek Agrawala. Broadcast protocols for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 1(1):17–25, January 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martin:1997:RBE

[MT97] Ignacio Martín and Francisco Tirado. Relationships between efficiency and execution time of full multigrid methods on parallel computers. IEEE Transactions on Parallel and Distributed Systems, 8(6):562–573, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0562.
pdf;   http://www.computer.
org/tpds/td1997/l0562abs.
htm.

Miura:2006:QBP

[MTK06] Ken Miura, Taro Tagawa, and Hirotsugu Kakugawa. A quorum-based protocol for searching objects in peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 17(1):25–37, January 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mounes-Toussi:1995:PCT

[MTL95] Farnaz Mounes-Toussi and David J. Lilja. The potential of compile-time analysis to adapt the cache coherence enforcement strategy to the data sharing characteristics. IEEE Transactions on Parallel and Distributed Systems, 6(5):470–481, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0470abs.htm.

McKinley:2002:SAF

[MTM02] Philip K. McKinley, Chiping Tang, and Arun P. Mani. A study of adaptive forward error correction for wireless collaborative computing. IEEE Transactions on Parallel and Distributed Systems, 13(9):936–947, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0936.
pdf;   http://www.computer.
org/tpds/td2002/l0936abs.
htm.

Mueller-Thuns:1993:BPP

[MTSDA93] Robert B. Mueller-Thuns, Daniel G. Saab, Robert F. Damiano, and Jacob A. Abraham. Benchmarking parallel processing platforms: an applications perspective. IEEE Transactions on Parallel and Distributed Systems, 4 (8):947–954, August 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mishra:2003:ICS

[MX03] Shivakant Mishra and Peng Xie.

Interagent communication and synchronization support in the DaAgent mobile agent-based computing system. IEEE Transactions on Parallel and Distributed Systems, 14(3):290–306, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0290abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0290.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0290.
pdf.

McKinley:1994:UBM

[MXEN94] Philip K. McKinley, Hong Xu, Abdol-Hossein Esfahanian, and Lionel M. Ni. Unicastbased multicast communication in wormhole-routed networks. IEEE Transactions on Parallel and Distributed Systems, 5(12): 1252–1265, December 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2007:SEE

[MY07] Ming Ma and Yuanyuan Yang.

SenCar: An energy-efficient data gathering mechanism for largescale multihop sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(10):

1476–1488, October 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Moritz:2001:SFA

[MYA01] Csaba Andras Moritz, Donald Yeung, and Anant Agarwal. SimpleFit: a framework for analyzing design trade-offs in raw architectures. IEEE Transactions on Parallel and Distributed Systems, 12(7):730– 742, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0730.
pdf;   http://www.computer.
org/tpds/td2001/l0730abs.
htm.

Mahapatra:2005:EES

[MZ05] Rabi N. Mahapatra and Wei Zhao. An energy-efficient slack distribution technique for multimode distributed real-time embedded systems. IEEE Transactions on Parallel and Distributed Systems, 16(7):650–662, July 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Manjeshwar:2002:AMI

[MZA02] Arati Manjeshwar, Qing-An Zeng, and Dharma P. Agrawal. An analytical model for information retrieval in wireless sensor networks using enhanced APTEEN protocol. IEEE Transactions on Parallel and Distributed Systems, 13(12):1290–1302, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1290abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1290.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1290.
pdf.

Ma:2008:MMM

[MZT08] Ke Ma, Yanyong Zhang, and Wade Trappe. Managing the mobility of a mobile sensor network using network dynamics. IEEE Transactions on Parallel and Distributed Systems, 19(1): 106–120, January 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nassimi:1993:PAC

[Nas93] David Nassimi. Parallel algorithms for the classes of ±2^{b} DE-SCEND and ASCEND computations on a SIMD hypercube. IEEE Transactions on Parallel and Distributed Systems, 4(12): 1372–1381, December 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Neves:2005:SVC

[NCV05] Nuno F. Neves, Miguel Correia, and Paulo Verissimo. Solving vector consensus with a wormhole. IEEE Transactions on Parallel and Distributed Systems, 16(12):1120–1131, December 2005. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Nakatani:1993:MCB

[NE93] Toshio Nakatani and Kemal Ebcioglu. Making compactionbased parallelization affordable. IEEE Transactions on Parallel and Distributed Systems, 4 (9):1014–1029, September 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Neophytou:2001:NDW

[NE01] Neophytos Neophytou and Paraskevas Evripidou. Net-dbx: a Web-based debugger of MPI programs over low-bandwidth lines. IEEE Transactions on Parallel and Distributed Systems, 12(9):986–995, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0986.
pdf;   http://www.computer.
org/tpds/td2001/l0986abs.
htm.

Nagaraja:2005:QPC

[NGB^{+}05] Kiran Nagaraja, Gustavo Gama, Ricardo Bianchini, Richard P. Martin, Wagner Meira, Jr., and Thu D. Nguyen. Quantifying the performability of cluster-based services. IEEE Transactions on Parallel and Distributed Systems, 16(5):456–467, May 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nicol:1994:MPA

[NGL94] David M. Nicol, Albert G.

Greenberg, and Boris D. Lubachevsky. Massively parallel algorithms for trace-driven cache simulations. IEEE Transactions on Parallel and Distributed Systems, 5 (8):849–859, August 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ni:1997:PES

[NGM97] Lionel M. Ni, Yadong Gui, and Sherry Moore. Performance evaluation of switchbased wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 8(5):462– 474, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0462.
pdf;   http://www.computer.
org/tpds/td1997/l0462abs.
htm.

Nicol:1992:CPS

[Nic92] D. Nicol. Conservative parallel simulation of priority class queueing networks. IEEE Transactions on Parallel and Distributed Systems, 3(3):294–303, May 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Naylor:1994:PMM

[NJ94] David Naylor and Simon Jones.

Performance model for multilayer neural networks in linear arrays. IEEE Transactions on Parallel and Distributed Systems, 5(12):1322–1328, December 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nafaa:2008:SQG

[NK08] Abdelhamid Nafaa and Adlen Ksentini. On sustained QoS guarantees in operated IEEE 802.11 wireless LANs. IEEE Transactions on Parallel and Distributed Systems, 19(8):1020– 1033, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nieuwejaar:1996:FAC

[NKP^{+}96] Nils Nieuwejaar, David Kotz, Apratim Purakayastha, Carla Schlatter Ellis, and Michael L. Best. File-access characteristics of parallel scientific workloads. IEEE Transactions on Parallel and Distributed Systems, 7 (10):1075–1089, October 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l1075abs.htm.

Nicol:2002:CSP

[NL02] David M. Nicol and Jason Liu. Composite synchronization in parallel discrete-event simulation. IEEE Transactions on Parallel and Distributed Systems, 13(5):433–446, May 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

td/books/td2002/pdf/l0433.
pdf;   http://www.computer.
org/tpds/td2002/l0433abs.
htm.

Nishio:1990:RME

[NLM90] Shojiro Nishio, Kin F. Li, and Eric G. Manning. Resilient mutual exclusion algorithm for computer networks. IEEE Transactions on Parallel and Distributed Systems, 1(3):344– 355, July 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nagumo:1999:PPA

[NLW99] H. Nagumo, M. Lu, and K. L.

Watson. Parallel parsing algorithms for static dictionary compression. IEEE Transactions on Parallel and Distributed Systems, 10(12):241–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1241.
pdf;   http://www.computer.
org/tpds/td1999/l1241abs.
htm.

Neilsen:1992:CJA

[NM92] Mitchell L. Neilsen and Masaaki Mizuno. Coterie join algorithm. IEEE Transactions on Parallel and Distributed Systems, 3(5): 582–590, September 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nanda:1996:MKE

[NN96] Arun K. Nanda and Lionel M.

Ni. MAD kernels: An experimental testbed to study multiprocessor memory system behavior. IEEE Transactions on Parallel and Distributed Systems, 7(2):207–217, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0207abs.htm.

Nakano:1997:OAA

[NO97] K. Nakano and S. Olariu. An optimal algorithm for the anglerestricted all nearest neighbor problem on the reconfigurable mesh, with applications. IEEE Transactions on Parallel and Distributed Systems, 8(9):983–990, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0983.
pdf;   http://www.computer.
org/tpds/td1997/l0983abs.
htm.

Nakano:1998:EAR

[NO98] K. Nakano and S. Olariu. An efficient algorithm for row minima computations on basic reconfigurable meshes. IEEE Transactions on Parallel and Distributed Systems, 9(6):561– ??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0561.
pdf;   http://www.computer.
org/tpds/td1998/l0561abs.
htm.

Nakano:2000:EEI

[NO00a] K. Nakano and S. Olariu.

Energy-efficient initialization protocols for single-hop radio networks with no collision detection. IEEE Transactions on Parallel and Distributed Systems, 11(8):851–??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0851.
pdf;   http://www.computer.
org/tpds/td2000/l0851abs.
htm.

Nakano:2000:RIP

[NO00b] K. Nakano and S. Olariu. Randomized initialization protocols for ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 11(7):749– ??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0749.
pdf;   http://www.computer.
org/tpds/td2000/l0749abs.
htm.

Nakano:2002:ULE

[NO02] Koji Nakano and Stephan Olariu. Uniform leader election protocols for radio networks. IEEE Transactions on Parallel and Distributed Systems, 13 (5):516–526, May 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2002/pdf/l0516.

                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0516abs.
                                        htm.

Nakano:1999:BEP

[NOS99] K. Nakano, S. Olariu, and J. L. Schwing. Broadcastefficient protocols for mobile radio networks. IEEE Transactions on Parallel and Distributed Systems, 10(12):276–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l1276.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l1276abs.
                                        htm.

Nakano:2001:EEP

[NOZ01] K. Nakano, S. Olariu, and A. Y. Zomaya. Energy-efficient permutation routing in radio networks. IEEE Transactions on Parallel and Distributed Systems, 12(6):544–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0544.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0544abs.
                                        htm.

Nakano:2002:EER

[NOZ02] Koji Nakano, Stephan Olariu, and Albert Y. Zomaya. Energyefficient routing in the broadcast communication model. IEEE Transactions on Parallel and Distributed Systems, 13(12):1201–1210, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1201abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1201.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1201.
pdf.

Newman:1995:HPA

[NS95a] Ilan Newman and Assaf Schuster. Hot-potato algorithms for permutation routing. IEEE Transactions on Parallel and Distributed Systems, 6(11):1168– 1176, November 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1168abs.htm.

Nigam:1995:SNM

[NS95b] Madhusudan Nigam and Sartaj Sahni. Sorting n^{2} numbers on n×n meshes. IEEE Transactions on Parallel and Distributed Systems, 6(12):1221–1225, December 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nichols:1991:EMF

[NSD^{+}91] Mark A. Nichols, Howard Jay Siegel, Henry G. Dietz, Russell W. Quong, and Wayne G. Nation. Eliminating memory fragmentation within partitionable SIMD/SPMD machines. IEEE Transactions on Parallel and Distributed Systems, 2 (3):290–303, July 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nichols:1993:DMC

[NSD93] Mark A. Nichols, Howard Jay Siegel, and Henry G. Dietz. Data management and control-flow aspects of an SIMD/SPMD parallel language/compiler. IEEE Transactions on Parallel and Distributed Systems, 4(2):222– 234, February 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Negro:1997:EDS

[NSU97] Alberto Negro, Nicola Santoro, and Jorge Urrutia. Efficient distributed selection with bounded messages. IEEE Transactions on Parallel and Distributed Systems, 8(4):397–401, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0397.
pdf;   http://www.computer.
org/tpds/td1997/l0397abs.
htm.

Nocetti:2002:ARH

[NSZ02] Fabian Garcia Nocetti, Ivan Stojmenovic, and Jingyuan Zhang. Addressing and routing in hexagonal networks with applications for tracking mobile users and connection rerouting in cellular networks. IEEE Transactions on Parallel and Distributed Systems, 13(9):963–971, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0963.
pdf;   http://www.computer.
org/tpds/td2002/l0963abs.
htm.

Nesterenko:2009:DNT

[NT09] Mikhail Nesterenko and Sebastien Tixeuil. Discovering network topology in the presence of Byzantine faults. IEEE Transactions on Parallel and Distributed Systems, 20(12):1777–1789, December 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Naor:1998:ACS

[NW98] M. Naor and A. Wool. Access control and signatures via quorum secret sharing. IEEE Transactions on Parallel and Distributed Systems, 9(9):909–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0909.
pdf;   http://www.computer.
org/tpds/td1998/l0909abs.
htm.

Netzer:1995:NSC

[NX95] Robert H. B. Netzer and Jian Xu. Necessary and sufficient conditions for consistent global snapshots. IEEE Transactions on Parallel and Distributed Systems, 6(2):165–169, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1995/
                                        l0165abs.htm.

Nienaber:2009:LAI

[NYD09] Wickus Nienaber, Xin Yuan, and Zhenhai Duan. LID assignment in InfiniBand networks. IEEE Transactions on Parallel and Distributed Systems, 20 (4):484–497, April 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nabhan:1995:PSA

[NZ95] Tarek M. Nabhan and Albert Y. Zomaya. Parallel simulated annealing algorithm with low communication overhead. IEEE Transactions on Parallel and Distributed Systems, 6(12):1226– 1233, December 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Navarro:2003:CTD

[NZP03] Angeles Navarro, Emilio Zapata, and David Padua. Compiler techniques for the distribution of data and computation. IEEE Transactions on Parallel and Distributed Systems, 14(6):545–562, June 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/06/l0545abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/06/l0545.
                                        pdf.

Oliker:2000:PDU

[OB00] L. Oliker and R. Biswas. Parallelization of a dynamic unstructured algorithm using three leading programming paradigms. IEEE Transactions on Parallel and Distributed Systems, 11(9):931–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0931.
pdf;   http://www.computer.
org/tpds/td2000/l0931abs.
htm.

Oh:1993:GMF

[OC93] A. Duksu Oh and Hyeong-Ah Choi. Generalized measures of fault tolerance in n-cube networks. IEEE Transactions on Parallel and Distributed Systems, 4(6):702–703, June 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ozer:2005:HPL

[OC05] Emre Ozer and Thomas M.

Conte. High-performance and low-cost dual-thread VLIW processor using Weld architecture paradigm. IEEE Transactions on Parallel and Distributed Systems, 16(12):1132–1142, December 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

OKeefe:1993:LCS

[OD93] Matthew T. O’Keefe and Henry G. Dietz. Loop coalescing and scheduling for barrier MIMD architectures. IEEE Transactions on Parallel and Distributed Systems, 4(9):1060– 1064, September 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ohring:1996:FPC

[ÖD96] Sabine Öhring and Sajal K. Das. Folded Petersen cube networks: new competitors for the hypercubes. IEEE Transactions on Parallel and Distributed Systems, 7(2):151–168, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1996/
                                        l0151abs.htm.

OHallaron:1991:UAS

[O’H91] David R. O’Hallaron. Uniform approach for solving some classical problems on a linear array. IEEE Transactions on Parallel and Distributed Systems, 2 (2):236–241, April 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Obreni:1999:UEE

[OHRW99] B. Obreni, M. C. Herbordt, A. L. Rosenberg, and C. C. Weems. Using emulations to enhance the performance of parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 10(10):67–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l1067.
                                        pdf;   http://www.computer.

org/tpds/td1999/l1067abs.
htm.

Ould-Khaoua:2001:AMA

[OKSA01] M. Ould-Khaoua and H. Sarbazi-Azad. An analytical model of adaptive wormhole routing in hypercubes in the presence of hot spot traffic. IEEE Transactions on Parallel and Distributed Systems, 12(3):283–??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0283.
pdf;   http://www.computer.
org/tpds/td2001/l0283abs.
htm.

Omiecinski:1992:AHJ

[OL92] E. Omiecinski and E. T. Lin.

The adaptive-hash join algorithm for a hypercube multicomputer. IEEE Transactions on Parallel and Distributed Systems, 3(3):334–349, May 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Omiecinski:1990:PAR

[Omi90] Edward R. Omiecinski. A parallel algorithm for relational database normalization. IEEE Transactions on Parallel and Distributed Systems, 1(4):415– 423, October 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olariu:2002:GEI

[ON02] Stephan Olariu and Koji Nakano.

Guest Editors’ introduction to special section on mobile computing and wireless networks. IEEE Transactions on Parallel and Distributed Systems, 13(9):865, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0865.
pdf;   http://www.computer.
org/tpds/td2002/l0865abs.
htm.

Oehmen:2006:SSI

[ON06] Christopher Oehmen and Jarek Nieplocha. ScalaBLAST: a scalable implementation of BLAST for high-performance dataintensive bioinformatics analysis. IEEE Transactions on Parallel and Distributed Systems, 17 (8):740–749, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olariu:1999:HSI

[OPZ99] S. Olariu, M. C. Pinotti, and S. Q. Zheng. How to sort N items using a sorting network of fixed I/O size. IEEE Transactions on Parallel and Distributed Systems, 10(5):487– ??, May 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0487.
pdf;   http://www.computer.
org/tpds/td1999/l0487abs.
htm.

Ou:1997:PIG

[OR97] Chao-Wei Ou and S. Ranka.

Parallel incremental graph partitioning. IEEE Transactions on Parallel and Distributed Systems, 8(8):884–896, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0884.
pdf;   http://www.computer.
org/tpds/td1997/l0884abs.
htm.

Olson:1994:FTC

[OS94a] Alan Olson and Kang G. Shin.

Fault-tolerant clock synchronization in large multicomputer systems. IEEE Transactions on Parallel and Distributed Systems, 5(9):912–923, September 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olson:1994:FTR

[OS94b] Alan Olson and Kang G. Shin.

Fault-tolerant routing in mesh architectures. IEEE Transactions on Parallel and Distributed Systems, 5(11):1225– 1232, November 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

OBoyle:2002:CTB

[OS02] Michael O’Boyle and Elena Stöhr. Compile time barrier synchronization minimization. IEEE Transactions on Parallel and Distributed Systems, 13 (6):529–543, June 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2002/pdf/l0529.
pdf;   http://www.computer.
org/tpds/td2002/l0529abs.
htm.

Olariu:2006:LCTa

[OSRS06a] Stephan Olariu, David Simplot-Ryl, and Ivan Stojmenovic. Localized communication and topology protocols for ad hoc networks: a preface to the special section. IEEE Transactions on Parallel and Distributed Systems, 17(4):289– 291, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/04/
l0289.pdf.

Olariu:2006:LCTb

[OSRS06b] Stephan Olariu, David Simplot-Ryl, and Ivan Stojmenovic. Localized communication and topology protocols for ad hoc networks: Part II: a preface to the special section. IEEE Transactions on Parallel and Distributed Systems, 17(9):897– 898, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/09/
l0897.pdf.

Ogle:1993:ADD

[OSS93] David M. Ogle, Karsten Schwan, and Richard Snodgrass. Applicationdependent dynamic monitoring of distributed and parallel systems. IEEE Transactions on Parallel and Distributed Systems, 4(7):762–778, July 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ostroff:1990:DPT

[Ost90] Jonathan S. Ostroff. Deciding properties of timed transition models. IEEE Transactions on Parallel and Distributed Systems, 1(2):170–183, April 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olariu:1992:OPA

[OSZ92] S. Olariu, J. L. Schwing, and J. Zhang. Optimal parallel algorithms for problems modeled by a family of intervals. IEEE Transactions on Parallel and Distributed Systems, 3(3):364– 374, May 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olariu:1991:OPI

[OW91] Stephan Olariu and Zhaofang Wen. Optimal parallel initialization algorithms for a class of priority queues. IEEE Transactions on Parallel and Distributed Systems, 2(4):423–429, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Oleszkiewicz:2006:EUG

[OXL06] John Oleszkiewicz, Li Xiao, and Yunhao Liu. Effectively utilizing global cluster memory for large data-intensive parallel programs. IEEE Transactions on Parallel and Distributed Systems, 17(1):66–77, January 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olariu:1996:TCO

[OZ96] Stephan Olariu and Albert Y. Zomaya. A time- and costoptimal algorithm for interlocking sets (with applications). IEEE Transactions on Parallel and Distributed Systems, 7 (10):1009–1025, October 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l1009abs.htm.

Powell:1999:GGU

[PABD^{+}99] D. Powell, J. Arlat, L. Beus-Dukic, A. Bondavalli, P. Coppola, A. Fantechi, E. Jenn, C. Rabéjac, and A. Wellings. GUARDS: a generic upgradable architecture for real-time dependable systems. IEEE Transactions on Parallel and Distributed Systems, 10(6):580– ??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0580.
pdf;   http://www.computer.
org/tpds/td1999/l0580abs.
htm.

Padmanabhan:1991:DAE

[Pad91] Krishnan Padmanabhan. Design and analysis of even-sized binary shuffle-exchange networks for multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 2(4):385–397, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pakin:2007:DID

[Pak07] Scott Pakin. The design and implementation of a domainspecific language for network performance testing. IEEE Transactions on Parallel and Distributed Systems, 18(10): 1436–1449, October 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Prasanna:1994:HCM

[PAM94] G. N. Srinivasa Prasanna, Anant Agarwal, and Bruce R. Musicus. Hierarchical compilation of macro dataflow graphs for multiprocessors with local memory. IEEE Transactions on Parallel and Distributed Systems, 5 (7):720–736, July 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pande:1995:SSS

[PAM95] Santosh Pande, Dharma P.

Agrawal, and Jon Mauney. A scalable scheduling scheme for functional parallelism on distributed memory multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 6(4):388–399, April 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.

                                        computer.org/tpds/td1995/
                                        l0388abs.htm.

Pan:1993:CIA

[Pan93] Victor Y. Pan. Concurrent iterative algorithm for Toeplitzlike linear systems. IEEE Transactions on Parallel and Distributed Systems, 4(5):592–600, May 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:1995:EPP

[Par95] Chan-Ik Park. Efficient placement of parity and data to tolerate two disk failures in disk array systems. IEEE Transactions on Parallel and Distributed Systems, 6(11):1177– 1184, November 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l1177abs.htm.

Park:2001:EBM

[Par01]  J. W. Park.  An efficient

buffer memory system for subarray access. IEEE Transactions on Parallel and Distributed Systems, 12(3):316–??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0316.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0316abs.
                                        htm.

Pilkington:1996:DPN

[PB96] John R. Pilkington and Scott B.

Baden. Dynamic partitioning of non-uniform structured workloads with spacefilling curves. IEEE Transactions on Parallel and Distributed Systems, 7(3): 288–300 (or 288–299??), March 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0288abs.htm.

Pinto:2003:CLT

[PBA03] Raquel Pinto, Ricardo Bianchini, and Claudio L. Amorim. Comparing latency-tolerance techniques for software DSM systems. IEEE Transactions on Parallel and Distributed Systems, 14(11):1180–1190, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/11/l1180abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/11/l1180.
pdf.

Park:1996:CSB

[PC96] Ju-Young L. Park and Hyeong-Ah Choi. Circuit-switched broadcasting in torus and mesh networks. IEEE Transactions on Parallel and Distributed Systems, 7(2):184–190, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0184abs.htm.

Peng:2005:SDA

[PC05] Wen-Chih Peng and Ming-Syan Chen. Shared data allocation in a mobile computing system: Exploring local and global optimization. IEEE Transactions on Parallel and Distributed Systems, 16(4):374–384, April 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2007:FBA

[PC07] Eun-Chan Park and Chong-Ho Choi. Feedback-based adaptive packet marking for proportional bandwidth allocation. IEEE Transactions on Parallel and Distributed Systems, 18(2): 225–239, February 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pong:1995:NAV

[PD95] Fong Pong and Michel Dubois. A new approach for the verification of cache coherence protocols. IEEE Transactions on Parallel and Distributed Systems, 6 (8):773–787, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0773abs.htm.

Petitet:1999:ARM

[PD99] A. P. Petitet and J. J. Dongarra. Algorithmic redistribution methods for block-cyclic decompositions. IEEE Transactions on Parallel and Distributed Systems, 10(12):201–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1201.
pdf;   http://www.computer.
org/tpds/td1999/l1201abs.
htm.

Pong:2000:FAV

[PD00] F. Pong and M. Dubois. Formal automatic verification of cache coherence in multiprocessors with relaxed memory models. IEEE Transactions on Parallel and Distributed Systems, 11(9):989–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0989.
pdf;   http://www.computer.
org/tpds/td2000/l0989abs.
htm.

Prasad:1994:EEP

[PDC94] Sushil K. Prasad, Sajal K. Das, and Calvin C.-Y. Chen. Efficient EREW PRAM algorithms for parentheses-matching. IEEE Transactions on Parallel and Distributed Systems, 5(9):995– 1008, September 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pucha:2006:IRR

[PDH06] Himabindu Pucha, Saumitra M.

Das, and Y. Charlie Hu. Imposed route reuse in ad hoc network routing protocols using structured peer-to-peer overlay routing. IEEE Transactions on Parallel and Distributed Systems, 17(12):1452–1467, December 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Picker:1996:SST

[PF96] Dan Picker and Ronald D. Fellman. Successive superposition: a technique for the exact modeling of deterministic packet queuing networks. IEEE Transactions on Parallel and Distributed Systems, 7(10):1106– 1120, October 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l1106abs.htm.

Prodan:2008:OAS

[PF08] Radu Prodan and Thomas Fahringer. Overhead analysis of scientific workflows in grid environments. IEEE Transactions on Parallel and Distributed Systems, 19(3):378–393, March 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pontelli:2001:BIP

[PG01] Enrico Pontelli and Gopal Gupta. Backtracking in independent and-parallel implementations of logic programming languages. IEEE Transactions on Parallel and Distributed Systems, 12(11):1169–1189, November 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1169.
pdf;   http://www.computer.
org/tpds/td2001/l1169abs.
htm.

Puente:2007:ISF

[PG07] Valentin Puente and José Ángel Gregorio. Immucube: Scalable fault-tolerant routing for k-ary n-cube networks. IEEE Transactions on Parallel and Distributed Systems, 18(6):776–788, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Puente:2003:DHP

[PGBI03] Valentín Puente, José-Ángel Gregorio, Ramón Beivide, and Cruz Izu. On the design of a high-performance adaptive router for CC-NUMA multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 14(5):487–501, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/05/l0487abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0487.
htm;   http://csdl.computer.
org/dl/trans/td/2003/05/l0487.
pdf.

Pifarre:1994:ADL

[PGDS94] Gustavo D. Pifarre, Luis Gravano, Gustavo Denicolay, and Jorge L. C. Sanz. Adaptive deadlock-and livelock-free routing in the hypercube network.

IEEE Transactions on Parallel and Distributed Systems, 5(11): 1121–1139, November 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pifarre:1994:FAM

[PGFS94] Gustavo D. Pifarre, Luis Gravano, Sergio A. Felperin, and Jorge L. C. Sanz. Fully adaptive minimal deadlock-free packet routing in hypercubes, meshes, and other networks: Algorithms and simulations. IEEE Transactions on Parallel and Distributed Systems, 5(3):247– 263, March 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:1996:EMS

[PH96] Jong Won Park and David T. Harper III. An efficient memory system for the SIMD construction of a Gaussian pyramid. IEEE Transactions on Parallel and Distributed Systems, 7 (8):855–860, August 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0855abs.htm.

Prechelt:2002:EPE

[PH02] Lutz Prechelt and Stefan U. Hänßgen. Efficient parallel execution of irregular recursive programs. IEEE Transactions on Parallel and Distributed Systems, 13(2):167–178, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0167.
pdf;   http://www.computer.
org/tpds/td2001/l0167abs.
htm.

Pinar:2004:ICL

[PH04] Ali Pinar and Bruce Hendrickson. Interprocessor communication with limited memory. IEEE Transactions on Parallel and Distributed Systems, 15 (7):606–616, July 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/07/l0606.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/07/l0606.pdf.

Pinar:2005:ILB

[PH05] Ali Pinar and Bruce Hendrickson. Improving load balance with flexibly assignable tasks. IEEE Transactions on Parallel and Distributed Systems, 16(10): 956–965, October 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Palesi:2009:ASR

[PHKC09] Maurizio Palesi, Rickard Holsmark, Shashi Kumar, and Vincenzo Catania. Application specific routing algorithms for networks on chip. IEEE Transactions on Parallel and Distributed Systems, 20(3):316–330, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2003:TBD

[PHP03] Neungsoo Park, Bo Hong, and Viktor K. Prasanna. Tiling, block data layout, and memory hierarchy performance. IEEE Transactions on Parallel and Distributed Systems, 14(7):640– 654, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/07/
                                        l0640abs.htm;   http://csdl.
                                        computer.org/dl/trans/td/
                                        2003/07/l0640.pdf.

Page:1993:FAD

[PJC93] Ivor Page, Tom Jacob, and Eric Chern. Fast algorithms for distributed resource allocation. IEEE Transactions on Parallel and Distributed Systems, 4(2): 188–197, February 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pilarski:1992:CDD

[PK92] Slawomir Pilarski and Tiko Kameda. Checkpointing for distributed databases: Starting from the basics. IEEE Transactions on Parallel and Distributed Systems, 3(5):602–610, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pramanick:1995:IPMa

[PK95a] Ira Pramanick and Jon G. Kuhl. An inherently parallel method for heuristic problemsolving: Part I: General framework. IEEE Transactions on Parallel and Distributed Systems, 6(10):1006–1015, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l1006abs.htm.

Pramanick:1995:IPMb

[PK95b] Ira Pramanick and Jon G. Kuhl.

An inherently parallel method for heuristic problem-solving: Part II: Example applications. IEEE Transactions on Parallel and Distributed Systems, 6 (10):1016–1028, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1016abs.htm.

Parhami:1999:DDC

[PK99a] B. Parhami and D.-M. Kwai.

Data-driven control scheme for linear arrays: Application to a stable insertion sorter. IEEE Transactions on Parallel and Distributed Systems, 10(1):23– ??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0023.
pdf;   http://www.computer.
org/tpds/td1999/l0023abs.
htm.

Parhami:1999:PRC

[PK99b] B. Parhami and D.-M. Kwai. Periodically regular chordal rings. IEEE Transactions on Parallel and Distributed Systems, 10 (6):658–??, June 1999. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td1999/pdf/l0658.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0658abs.

htm. See correction [Ano99f].

Perkovic:2000:PCA

[PK00] D. Perkovic and P. J. Keleher. A protocol-centric approach to on-the-fly race detection. IEEE Transactions on Parallel and Distributed Systems, 11(10):58– ??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l1058.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l1058abs.
                                        htm.

Parhami:2001:UFH

[PK01] B. Parhami and D.-M. Kwai. A unified formulation of honeycomb and diamond networks. IEEE Transactions on Parallel and Distributed Systems, 12(1):74–??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0074.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0074abs.
                                        htm.

Psarris:2004:EED

[PK04] Kleanthis Psarris and Konstantinos Kyriakopoulos. An experimental evaluation of data dependence analysis techniques.

IEEE Transactions on Parallel and Distributed Systems, 15(3):196–213, March 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/03/l0196abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0196.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0196.
pdf.

Patel:1997:SPI

[PKJ97] J. N. Patel, A. A. Khokhar, and L. H. Jamieson. Scalable parallel implementations of list ranking on fine-grained machines. IEEE Transactions on Parallel and Distributed Systems, 8(10):1006–1018, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1006.
pdf;   http://www.computer.
org/tpds/td1997/l1006abs.
htm.

Psarris:1993:DVT

[PKK93] Kleanthis Psarris, Xiangyun Kong, and David Klappholz. The direction vector I test. IEEE Transactions on Parallel and Distributed Systems, 4(11):1280– 1290, November 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2006:MMD

[PKL06] Jung-Heum Park, Hee-Chul Kim, and Hyeong-Seok Lim.

Many-to-many disjoint path covers in hypercube-like interconnection networks with faulty elements. IEEE Transactions on Parallel and Distributed Systems, 17(3):227– 240, March 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/03/
l0227s.pdf.

Plank:1998:DC

[PLP98] J. S. Plank, K. Li, and M. A. Puening. Diskless checkpointing. IEEE Transactions on Parallel and Distributed Systems, 9(10):972–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0972.
pdf;   http://www.computer.
org/tpds/td1998/l0972abs.
htm.

Prieto:2000:DLE

[PLT00] M. Prieto, I. M. Llorente, and F. Tirado. Data locality exploitation in the decomposition of regular domain problems. IEEE Transactions on Parallel and Distributed Systems, 11(11):141–??, November 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1141.
pdf;   http://www.computer.
org/tpds/td2000/l1141abs.
htm.

Palis:1996:TCS

[PLW96] Michael A. Palis, Jing-Chiou Liou, and David S. L. Wei. Task clustering and scheduling for distributed memory parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 7(1):46–55, January 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Prasanna:1996:GMS

[PM96] G. N. Srinivasa Prasanna and B. R. Musicus. Generalized multiprocessor scheduling and applications to matrix computations. IEEE Transactions on Parallel and Distributed Systems, 7 (6):650–664, June 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0650abs.htm.

Park:2002:ELD

[PM02] Changsik Park and John J.

Metzner. Efficient location of discrepancies in multiple replicated large files. IEEE Transactions on Parallel and Distributed Systems, 13(6):597–610, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0597.
pdf;   http://www.computer.
org/tpds/td2002/l0597abs.
htm.

Pritchard:1993:CCM

[PN93] David J. Pritchard and Denis A. Nicole. Cube connected Möbius ladders: an inherently deadlock-free fixed degree network. IEEE Transactions on Parallel and Distributed Systems, 4(1):111–117, January 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pratt:1995:DSC

[PN95] Gill A. Pratt and John Nguyen. Distributed synchronous clocking. IEEE Transactions on Parallel and Distributed Systems, 6 (3):314–328, March 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0314abs.htm.

Paek:2002:ACF

[PNZ^{+}02] Y. Paek, A. Navarro, E. Zapata, J. Hoeflinger, and D. Padua. An advanced compiler framework for non-cache-coherent multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 13(3):241–259, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0241.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0241abs.
                                        htm.

Posch:1995:MRR

[PP95] Karl C. Posch and Reinhard Posch. Modulo reduction in residue number systems. IEEE Transactions on Parallel and Distributed Systems, 6(5):449– 454, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0449abs.
htm.

Petersen:1996:SDE

[PP96] Paul M. Petersen and David A.

Padua. Static and dynamic evaluation of data dependence analysis techniques. IEEE Transactions on Parallel and Distributed Systems, 7(11): 1121–1132, November 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//neumann.computer.org/td/
books/td1996/pdf/l1121.pdf;
http://www.computer.org/
tpds/td1996/l1121abs.htm.

Peh:2005:GES

[PP05] Li-Shiuan Peh and Timothy Mark Pinkston. Guest editorial: Special section on onchip networks. IEEE Transactions on Parallel and Distributed Systems, 16(2):97–98, February 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/02/l0097.pdf.

Popp:1997:SMP

[PPBSA97] R. L. Popp, K. R. Pattipati, Y. Bar-Shalom, and R. A.

Ammar. Shared-memory parallelization of the data association problem in multitarget tracking. IEEE Transactions on Parallel and Distributed Systems, 8(10):993–1005, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0993.
pdf;   http://www.computer.
org/tpds/td1997/l0993abs.
htm.

Pinkston:2003:DFD

[PPD03] Timothy Mark Pinkston, Ruoming Pang, and José Duato. Deadlock-free dynamic reconfiguration schemes for increased network dependability. IEEE Transactions on Parallel and Distributed Systems, 14(8):780–794, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0780abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/08/l0780.
pdf.

Park:2004:OGA

[PPP04] Joon-Sang Park, Michael Penner, and Viktor K. Prasanna. Optimizing graph algorithms for improved cache performance. IEEE Transactions on Parallel and Distributed Systems, 15(9): 769–782, September 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.

org/comp/trans/td/2004/09/
L0769s.pdf;   http://csdl.
computer.org/dl/trans/td/
2004/09/l0769.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/09/l0769.pdf.

Prete:1995:TDS

[PPR95] Cosimo Antonio Prete, Gianpaolo Prina, and Luigi Ricciardi. A trace-driven simulator for performance evaluation of cachebased multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 6(9): 915–929, September 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0915abs.htm.

Park:1999:EAB

[PPR99] N. Park, V. K. Prasanna, and C. S. Raghavendra. Efficient algorithms for block-cyclic array redistribution between processor sets. IEEE Transactions on Parallel and Distributed Systems, 10(12):217–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1217.
pdf;   http://www.computer.
org/tpds/td1999/l1217abs.
htm.

Parhami:2005:PDN

[PR05a] Behrooz Parhami and Mikhail Rakov. Perfect difference networks and related interconnection structures for parallel and distributed systems. IEEE Transactions on Parallel and Distributed Systems, 16(8):714– 724, August 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parhami:2005:PAR

[PR05b] Behrooz Parhami and Mikhail A. Rakov. Performance, algorithmic, and robustness attributes of perfect difference networks. IEEE Transactions on Parallel and Distributed Systems, 16 (8):725–736, August 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Prechelt:1999:EDS

[Pre99] L. Prechelt. Exploiting domainspecific properties: Compiling parallel dynamic neural network algorithm into efficient code. IEEE Transactions on Parallel and Distributed Systems, 10(11):105–??, November 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l1105.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l1105abs.
                                        htm.

Passos:1996:AFP

[PS96a] Nelson Luiz Passos and Edwin Hsing-Mean Sha. Achieving full parallelism using multidimensional retiming. IEEE Transactions on Parallel and Distributed Systems, 7(11): 1150–1163, November 1996.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//neumann.computer.org/td/
books/td1996/pdf/l1150.pdf;
http://www.computer.org/
tpds/td1996/l1150abs.htm.

Peters:1996:CSB

[PS96b] Joseph G. Peters and Michel Syska. Circuit-switched broadcasting in torus networks. IEEE Transactions on Parallel and Distributed Systems, 7(3):246– 255, March 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1996/l0246abs.
htm.

Prakash:1996:LCC

[PS96c] Ravi Prakash and Mukesh Singhal. Low-cost checkpointing and failure recovery in mobile computing systems. IEEE Transactions on Parallel and Distributed Systems, 7(10):1035– 1048, October 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l1035abs.htm.

Plale:2003:DQS

[PS03] Beth Plale and Karsten Schwan.

Dynamic querying of streaming data with the dQUOB system. IEEE Transactions on Parallel and Distributed Systems, 14(4):422–432, April 2003. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0422abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0422.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0422.
pdf.

Park:2008:RSC

[PS08] Eunjeong Park and Heonshik Shin. Reconfigurable service composition and categorization for power-aware mobile computing. IEEE Transactions on Parallel and Distributed Systems, 19 (11):1553–1564, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ponnusamy:1995:RSC

[PSC^{+}95] Ravi Ponnusamy, Joel Saltz, Alok Choudhary, Yuan-Shin Hwang, and Geoffrey Fox. Runtime support and compilation methods for user-specified irregular data distributions. IEEE Transactions on Parallel and Distributed Systems, 6(8):815– 831, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0815abs.
htm.

Parcerisa:2005:CII

[PSGD05] Joan-Manuel Parcerisa, Julio Sahuquillo, Antonio Gonzalez, and Jose Duato. On-chip interconnects and instruction steering schemes for clustered microarchitectures. IEEE Transactions on Parallel and Distributed Systems, 16(2):130–144, February 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Panda:1999:MMP

[PSK99] D. K. Panda, S. Singal, and R. Kesavan. Multidestination message passing in wormhole kary n-cube networks with base routing conformed paths. IEEE Transactions on Parallel and Distributed Systems, 10(1):76– ??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0076.
pdf;   http://www.computer.
org/tpds/td1999/l0076abs.
htm.

Pugh:1995:GBI

[PW95] William Pugh and David Wonnacott. Going beyond integer programming with the omega test to eliminate false data dependences. IEEE Transactions on Parallel and Distributed Systems, 6(2):204–211, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0204abs.htm.

Pinkston:1999:CDA

[PW99] T. M. Pinkston and S. Warnakulasuriya. Characterization of deadlocks in k-ary ncube networks. IEEE Transactions on Parallel and Distributed Systems, 10(9):904–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0904.
pdf;   http://www.computer.
org/tpds/td1999/l0904abs.
htm.

Peng:2000:RUD

[PWW00] C.-H. Peng, B.-F. Wang, and J.-S. Wang. Recognizing unordered depth-first search trees of an undirected graph in parallel. IEEE Transactions on Parallel and Distributed Systems, 11(6):559–??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0559.
pdf;   http://www.computer.
org/tpds/td2000/l0559abs.
htm.

Pan:2001:IGM

[PZLS01] Y. Pan, S. Q. Zheng, K. Li, and H. Shen. An improved generalization of meshconnected computers with multiple buses. IEEE Transactions on Parallel and Distributed Systems, 12(3):293–??, March 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0293.
pdf;   http://www.computer.
org/tpds/td2001/l0293abs.
htm.

Puttaswamy:2009:SSO

[PZZ09] Krishna P. N. Puttaswamy, Haitao Zheng, and Ben Y. Zhao. Securing structured overlays against identity attacks. IEEE Transactions on Parallel and Distributed Systems, 20(10): 1487–1498, October 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qadeer:2003:VSC

[Qad03] Shaz Qadeer. Verifying sequential consistency on sharedmemory multiprocessors by model checking. IEEE Transactions on Parallel and Distributed Systems, 14(8):730–741, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0730abs.
htm;   http://csdl.computer.
org/comp/trans/td/2003/08/
l0730s.pdf;   http://csdl.
computer.org/dl/trans/td/
2003/08/l0730.pdf.

Quaglia:1999:TBS

[QCC99] F. Quaglia, V. Cortellessa, and B. Ciciani. Trade-off between sequential and time warp-based parallel simulation. IEEE Transactions on Parallel and Distributed Systems, 10(8):781–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0781.
pdf;   http://www.computer.

                                        org/tpds/td1999/l0781abs.
                                        htm.

Qiu:2005:MAT

[QD05] Xiaogang Qiu and Michel Dubois. Moving address translation closer to memory in distributed shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 16(7):612–623, July 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiao:1994:RTD

[QM94] Chunming Qiao and Rami Melhem. Reconfiguration with time division multiplexed MIN’s for multiprocessor communications. IEEE Transactions on Parallel and Distributed Systems, 5 (4):337–352, April 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiao:1997:RCL

[QM97] Chunming Qiao and Rami Melhem. Reducing communication latency with path multiplexing in optically interconnected multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 8(2):97–108, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0097.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0097abs.
                                        htm.

Qiao:1999:ATR

[QNR99] W. Qiao, L. M. Ni, and T. Rokicki. Adaptive-trail routing and performance evaluation in irregular networks using cutthrough switches. IEEE Transactions on Parallel and Distributed Systems, 10(11):138–??, November 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1138.
pdf;   http://www.computer.
org/tpds/td1999/l1138abs.
htm.

Quaglia:2007:ETA

[QR07] Francesco Quaglia and Paolo Romano. Ensuring e-transaction with asynchronous and uncoordinated application server replicas. IEEE Transactions on Parallel and Distributed Systems, 18 (3):364–378, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Quaglia:2003:NCO

[QS03] Francesco Quaglia and Andrea Santoro. Nonblocking checkpointing for optimistic parallel simulation: Description and an implementation. IEEE Transactions on Parallel and Distributed Systems, 14(6):593–610, June 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/06/l0593abs.
htm;   http://csdl.computer.

org/dl/trans/td/2003/06/l0593.
pdf.

Quaglia:2001:CMS

[Qua01] F. Quaglia. A cost model for selecting checkpoint positions in time warp parallel simulation. IEEE Transactions on Parallel and Distributed Systems, 12(4):346–??, April 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0346.
pdf;   http://www.computer.
org/tpds/td2001/l0346abs.
htm.

Rajko:2004:STO

[RA04] Stjepan Rajko and Srinivas Aluru. Space and time optimal parallel sequence alignments. IEEE Transactions on Parallel and Distributed Systems, 15(12):1070–1081, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/12/l1070abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1070.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1070.
pdf.

Rashid:2005:AAP

[RAHM05] Mohammad M. Rashid, Attahiru Sule Alfa, Ekram Hossain, and Muthucumaru Maheswaran. An analytical approach to providing controllable differentiated quality of service in Web servers. IEEE Transactions on Parallel and Distributed Systems, 16(11):1022–1033, November 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rajasekaran:2005:EPH

[Raj05] Sanguthevar Rajasekaran. Efficient parallel hierarchical clustering algorithms. IEEE Transactions on Parallel and Distributed Systems, 16(6):497–502, June 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramamritham:1995:ASP

[Ram95]  Krithi Ramamritham.  Al-

location and scheduling of precedence-related periodic tasks. IEEE Transactions on Parallel and Distributed Systems, 6 (4):412–420, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0412abs.htm.

Ramanathan:1999:OMR

[Ram99]  P. Ramanathan.  Overload

management in real-time control applications using (m, k)- firm guarantee. IEEE Transactions on Parallel and Distributed Systems, 10(6):549–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0549.
pdf;   http://www.computer.

                                        org/tpds/td1999/l0549abs.
                                        htm.

Rao:1996:PAS

[Rao96] Nageswara S. V. Rao. On parallel algorithms for single-fault diagnosis in fault propagation graph systems. IEEE Transactions on Parallel and Distributed Systems, 7(12):1217–1223, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ravelomanana:2007:OIG

[Rav07] Vlady Ravelomanana. Optimal initialization and gossiping algorithms for random radio networks. IEEE Transactions on Parallel and Distributed Systems, 18(1):17–28, January 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Reddy:1990:DAS

[RB90] A. L. Narasimha Reddy and Prithviraj Banerjee. Design, analysis, and simulation of I/O architectures for hypercube multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 1(2):140–151, April 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Romein:2002:PAT

[RBSP02] John W. Romein, Henri E. Bal, Jonathan Schaeffer, and Aske Plaat. A performance analysis of transposition-table-driven work scheduling in distributed search. IEEE Transactions on Parallel and Distributed Systems, 13(5):447–459, May 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0447.
pdf;   http://www.computer.
org/tpds/td2002/l0447abs.
htm.

Ramanathan:1995:RPM

[RC95] Parameswaran Ramanathan and Suresh Chalasani. Resource placement with multiple adjacency constraints in k-ary ncubes. IEEE Transactions on Parallel and Distributed Systems, 6(5):511–519, May 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0511abs.htm.

Ren:2001:AAT

[RCS01] Y. Ren, M. Cukier, and W. H.

Sanders. An adaptive algorithm for tolerating value faults and crash failures. IEEE Transactions on Parallel and Distributed Systems, 12(2):173–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0173.
pdf;   http://www.computer.
org/tpds/td2001/l0173abs.
htm.

Rigoutsos:1998:MSB

[RD98] I. Rigoutsos and A. Delis. Managing statistical behavior in large data sets in shared-nothing architectures. IEEE Transactions on Parallel and Distributed Systems, 9(11):73–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1073.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1073abs.
                                        htm.

Roy:2009:EAG

[RD09] Nirmalya Roy and Sajal K. Das. Enhancing availability of grid computational services to ubiquitous computing applications. IEEE Transactions on Parallel and Distributed Systems, 20 (7):953–967, July 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rezgui:2009:MRA

[RE09] Abdelmounaam Rezgui and Mohamed Eltoweissy. mRACER: a reliable adaptive service-driven efficient routing protocol suite for sensor-actuator networks. IEEE Transactions on Parallel and Distributed Systems, 20 (5):607–622, May 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rescigno:1997:OPC

[Res97]  Adele A. Rescigno.  Opti-

mal polling in communication networks. IEEE Transactions on Parallel and Distributed Systems, 8(5):449–461, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0449.
pdf;   http://www.computer.
org/tpds/td1997/l0449abs.
htm.

Rexford:1997:PMS

[RFDS97] J. Rexford, Wu-Chang Feng, J. Dolter, and K. G. Shin. ppmess-sim: a flexible and extensible simulator for evaluating multicomputer networks. IEEE Transactions on Parallel and Distributed Systems, 8(1):25–40, January 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Repantis:2009:QAS

[RGK09] Thomas Repantis, Xiaohui Gu, and Vana Kalogeraki. QoSaware shared component composition for distributed stream processing systems. IEEE Transactions on Parallel and Distributed Systems, 20(7):968–982, July 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramaswamy:2005:DAN

[RGL05] Lakshmish Ramaswamy, Bugra Gedik, and Ling Liu. A distributed approach to node clustering in decentralized peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 16(9):814–829, September 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ryu:2000:EFG

[RH00] K. D. Ryu and J. K. Hollingsworth. Exploiting fine-grained idle periods in networks of workstations. IEEE Transactions on Parallel and Distributed Systems, 11(7):683–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0683.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0683abs.
                                        htm.

Ryu:2004:RPS

[RH04] Kyung Dong Ryu and Jeffrey K. Hollingsworth. Resource policing to support fine-grain cycle stealing in networks of workstations. IEEE Transactions on Parallel and Distributed Systems, 15(10):878–892, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/10/l0878abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/10/l0878.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/10/l0878.
                                        pdf.

Risson:2009:TDR

[RHM09] John Risson, Aaron Harwood, and Tim Moors. Topology dissemination for reliable one-hop distributed hash tables. IEEE Transactions on Parallel and Distributed Systems, 20(5):680– 694, May 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ryu:1990:EAL

[RJ90] Kwan Woo Ryu and Joseph JáJá. Efficient algorithms for list ranking and for solving graph problems on the hypercube. IEEE Transactions on Parallel and Distributed Systems, 1(1):83–90, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rexford:1994:PES

[RJ94] Jennifer Rexford and Niraj K.

Jha. Partitioned encoding schemes for algorithm-based fault tolerance in massively parallel systems. IEEE Transactions on Parallel and Distributed Systems, 5(6):649–653, June 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rim:1996:VTN

[RJ96] Minjoong Rim and Rajiv Jain.

Valid transformations: a new class of loop transformations for high-level synthesis and pipelined scheduling applications. IEEE Transactions on Parallel and Distributed Systems, 7(4):399–410, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0399abs.htm.

Ratha:1999:CVA

[RJ99] N. K. Ratha and A. K. Jain. Computer vision algorithms on reconfigurable logic arrays. IEEE Transactions on Parallel and Distributed Systems, 10(1):29–??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0029.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0029abs.
                                        htm.

Rubio:2005:RSD

[RJ05] Juan Rubio and Lizy Kurian John. Reducing server data traffic using a hierarchical computation model. IEEE Transactions on Parallel and Distributed Systems, 16(10):933–943, October 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rao:1993:EPB

[RK93] V. Nageshwara Rao and Vipin Kumar. On the efficiency of parallel backtracking. IEEE Transactions on Parallel and Distributed Systems, 4(4):427–437, April 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramkumar:1994:MIPa

[RK94a] Balkrishna Ramkumar and Laxmikant V. Kale. Machine independent AND and OR parallel execution of logic programs: Part I — the binding environment. IEEE Transactions on Parallel and Distributed Systems, 5(2):170–180, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramkumar:1994:MIPb

[RK94b] Balkrishna Ramkumar and Laxmikant V. Kale. Machine independent AND and OR parallel execution of logic programs: Part II — compiled execution. IEEE Transactions on Parallel and Distributed Systems, 5(2): 181–192, February 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ranjan:2008:HPR

[RK08] Supranamaya Ranjan and Edward Knightly. High-performance resource allocation and request redirection algorithms for Web clusters. IEEE Transactions on Parallel and Distributed Systems, 19(9):1186–1200, September 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramachandran:2006:DGC

[RKHM06] Umakishore Ramachandran, Kathleen Knobe, Nissim Harel, and Hasnain A. Mandviwala. Distributed garbage collection algorithms for timestamped data. IEEE Transactions on Parallel and Distributed Systems, 17(10):1057–1071, October 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rajasekaran:1998:PAR

[RL98] S. Rajasekaran and I. Lee. Parallel algorithms for relational coarsest partition problems. IEEE Transactions on Parallel and Distributed Systems, 9 (7):687–??, July 1998. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1998/pdf/l0687.
pdf;   http://www.computer.
org/tpds/td1998/l0687abs.
htm.

Rubio:2003:FFC

[RLD03] Juan-Miguel Martínez Rubio, Pedro López, and José Duato. FC3D: Flow controlbased distributed deadlock detection mechanism for true fully adaptive routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 14(8):765–779, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/08/l0765abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/08/l0765.
pdf.

Ren:2007:DAS

[RLW^{+}07] Shansi Ren, Qun Li, Haining Wang, Xin Chen, and Xiaodong Zhang. Design and analysis of sensing scheduling algorithms under partial coverage for object detection in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(3):

334–350, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rego:1990:EPA

[RM90] Vernon Rego and Aditya P.

Mathur. Exploiting parallelism across program execution: a unification technique and its analysis. IEEE Transactions on Parallel and Distributed Systems, 1 (4):399–414, October 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Robinson:1995:OMC

[RMC95] David F. Robinson, Philip K.

McKinley, and Betty H. C. Cheng. Optimal multicast communication in wormhole-routed torus networks. IEEE Transactions on Parallel and Distributed Systems, 6(10):1029– 1042, October 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1029abs.htm.

Rao:1995:AWT

[RMO^{+}95] Nageswara S. V. Rao, Kurt J.

Maly, Stephan Olariu, Sudheer Dharanikota, Liping Zhang, and David E. Game. Average waiting time profiles of uniform distributed queue dual bus system model. IEEE Transactions on Parallel and Distributed Systems, 6(10):1068–1084, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l1068abs.htm.

Ryutov:2003:IAC

[RNKZ03] Tatyana Ryutov, Clifford Neuman, Dongho Kim, and Li Zhou. Integrated access control and intrusion detection for Web servers. IEEE Transactions on Parallel and Distributed Systems, 14(9):841–850, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0841abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0841.
pdf.

Ramachandran:2003:SCP

[RNR^{+}03] Umakishore Ramachandran, Rishiyur S. Nikhil, James M. Rehg, Yavor Angelov, Arnab Paul, Sameer Adhikari, Kenneth M. Mackenzie, Nissim Harel, and Kathleen Knobe. Stampede: a cluster programming middleware for interactive stream-oriented applications. IEEE Transactions on Parallel and Distributed Systems, 14(11): 1140–1154, November 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2003/11/
extras/l1140s.pdf;   http:
//csdl.computer.org/comp/
trans/td/2003/11/l1140abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/11/l1140.
pdf.

Rai:1999:TBF

[RO99] S. Rai and Y. C. Oh. Tighter bounds on full access probability in fault-tolerant multistage interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 10(3):328– ??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0328.
pdf;   http://www.computer.
org/tpds/td1999/l0328abs.
htm.

Robertazzi:2004:CND

[Rob04] Thomas G. Robertazzi. Comment on “A Novel Data Distribution Technique for Host-Client Type Parallel Applications”. IEEE Transactions on Parallel and Distributed Systems, 15(6):575–??, June 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0575.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0575.pdf. See

[CN02, CN04].

Rosenberg:2002:OSC

[Ros02] Arnold L. Rosenberg. Optimal schedules for cycle-stealing in a network of workstations with a bag-of-tasks workload. IEEE Transactions on Parallel and Distributed Systems, 13(2):179–191, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0179.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0179abs.
                                        htm.

Rosenberg:2003:AWC

[Ros03] Arnold L. Rosenberg. Accountable Web-computing. IEEE Transactions on Parallel and Distributed Systems, 14(2):97– 106, February 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/02/
                                        l0097abs.htm;   http://csdl.
                                        computer.org/dl/trans/td/
                                        2003/02/l0097.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2003/02/l0097.pdf.

Rogers:1994:CDM

[RP94] Anne Rogers and Keshav Pingali. Compiling for distributed memory architectures. IEEE Transactions on Parallel and Distributed Systems, 5(3):281– 298, March 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rauchwerger:1999:LTS

[RP99] L. Rauchwerger and D. A. Padua. The LRPD test: Speculative run-time parallelization of loops with privatization and reduction parallelization. IEEE Transactions on Parallel and Distributed Systems, 10(2):160–??, February 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0160.
pdf;   http://www.computer.
org/tpds/td1999/l0160abs.
htm.

Roman:1993:DSA

[RPW93] Gruia-Catalin Roman, Jerome Y.

Plun, and C. Donald Wilcox. Dynamic synchrony among atomic actions. IEEE Transactions on Parallel and Distributed Systems, 4(6):677–685, June 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rastello:2002:APP

[RR02] Fabrice Rastello and Yves Robert. Automatic partitioning of parallel loops with parallelepiped-shaped tiles. IEEE Transactions on Parallel and Distributed Systems, 13 (5):460–470, May 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2002/pdf/l0460.
pdf;   http://www.computer.
org/tpds/td2002/l0460abs.
htm.

Russ:1998:HDR

[RRFH98] S. H. Russ, J. Robinson, B. K.

Flachs, and B. Heckel. The Hector distributed run-time environment. IEEE Transactions on Parallel and Distributed Systems, 9(11):102–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1102.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1102abs.
                                        htm.

Roig:2007:NTG

[RRG07] Concepcio Roig, Ana Ripoll, and Fernando Guirado. A new task graph model for mapping message passing applications. IEEE Transactions on Parallel and Distributed Systems, 18(12): 1740–1753, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Requena:2009:FDF

[RRRM09] Crispin Gomez Requena, Maria Engracia Gomez Requena, Pedro Juan Lopez Rodriguez, and Jose Francisco Duato Marin. FT^{EI}: a dynamic fault-tolerant routing methodology for fat trees with exclusion intervals. IEEE Transactions on Parallel and Distributed Systems, 20 (6):802–817, June 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rajah:2009:ARS

[RRX09] Kannan Rajah, Sanjay Ranka, and Ye Xia. Advance reservations and scheduling for bulk transfers in research networks. IEEE Transactions on Parallel and Distributed Systems, 20(11): 1682–1697, November 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ranka:1990:OES

[RS90] Sanjay Ranka and Sartaj Sahni.

Odd even shifts in SIMD hypercubes. IEEE Transactions on Parallel and Distributed Systems, 1(1):77–82, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramanujam:1991:CTT

[RS91a] J. Ramanujam and P. Sadayappan. Compile-time techniques for data distribution in distributed memory machines. IEEE Transactions on Parallel and Distributed Systems, 2(4): 472–482, October 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ranka:1991:CHM

[RS91b] Sanjay Ranka and Sartaj Sahni.

Clustering on a hypercube multicomputer. IEEE Transactions on Parallel and Distributed Systems, 2(2):129–137, April 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rajsbaum:1994:PSP

[RS94] Sergio Rajsbaum and Moshe Sidi. On the performance of synchronized programs in distributed networks with random processing times and transmission delays. IEEE Transactions on Parallel and Distributed Systems, 5(9):939–950, September 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rajasekaran:1997:SSR

[RS97a] S. Rajasekaran and S. Sahni. Sorting, selection, and routing on the array with reconfigurable optical buses. IEEE Transactions on Parallel and Distributed Systems, 8(11):1123– 1132, November 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l1123.
pdf;   http://www.computer.
org/tpds/td1997/l1123abs.
htm.

Roberts:1997:GMD

[RS97b] A. Roberts and A. Symvonis. A general method for deflection worm routing on meshes based on packet routing algorithms. IEEE Transactions on Parallel and Distributed Systems, 8(8):781–789, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0781.
pdf;   http://www.computer.
org/tpds/td1997/l0781abs.
htm.

Rajasekaran:1998:RRS

[RS98] S. Rajasekaran and S. Sahni. Randomized routing, selection, and sorting on the OTISmesh. IEEE Transactions on Parallel and Distributed Systems, 9(9):833–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0833.
pdf;   http://www.computer.
org/tpds/td1998/l0833abs.
htm.

Reiter:2008:QCS

[RS08] Michael K. Reiter and Asad Samar. Quiver: Consistent and scalable object sharing for edge services. IEEE Transactions on Parallel and Distributed Systems, 19(7):878–889, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramaswamy:1997:FET

[RSB97] S. Ramaswamy, S. Sapatnekar, and P. Banerjee. A framework for exploiting task and data parallelism on distributed memory multicomputers. IEEE Transactions on Parallel and Distributed Systems, 8(11):1098–1116, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1098.
pdf;   http://www.computer.
org/tpds/td1997/l1098abs.
htm.

Rodolakis:2006:RSP

[RSG06] Georgios Rodolakis, Stavroula Siachalou, and Leonidas Georgiadis. Replicated server placement with QoS constraints. IEEE Transactions on Parallel and Distributed Systems, 17(10): 1151–1162, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramasubramanian:2002:ACL

[RSP02] Narasimhan Ramasubramanian, Ram Subramanian, and Santosh Pande. Automatic compilation of loops to exploit operator parallelism on configurable arithmetic logic units. IEEE Transactions on Parallel and Distributed Systems, 13(1):45–66, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0045.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0045abs.
                                        htm.

Ramamritham:1990:ESA

[RSS90] Krithi Ramamritham, John A. Stankovic, and Perng-Fei Shiah. Efficient scheduling algorithms for real-time multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 1(2):184–194, April 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rangarajan:1995:FTA

[RST95] Sampath Rangarajan, Sanjeev Setia, and Satish K. Tripathi. Fault-tolerant algorithm for replicated data management. IEEE Transactions on Parallel and Distributed Systems, 6(12): 1271–1282, December 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramachandran:1990:HSI

[RSV90]  Umakishore  Ramachandran,

Marvin Solomon, and Mary K. Vernon. Hardware support for interprocess communication. IEEE Transactions on Parallel and Distributed Systems, 1 (3):318–329, July 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rai:1995:PAH

[RTS95] Suresh Rai, Jerry L. Trahan, and Thomas Smailus. Processor allocation in hypercube multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6(6):606–616, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0606abs.htm.

Radulescu:2002:LCT

[RvG02] Andrei Radulescu and Arjan J. C. van Gemund. Low-cost task scheduling for distributedmemory machines. IEEE Transactions on Parallel and Distributed Systems, 13(6):597–610, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0648.
pdf;   http://www.computer.
org/tpds/td2002/l0648abs.
htm.

Ranka:1994:SRT

[RWF94] Sanjay Ranka, Jhy-Chun Wang, and Geoffrey C. Fox. Static and run-time algorithms for all-to-many personalized communication on permutation networks. IEEE Transactions on Parallel and Distributed Systems, 5(12): 1266–1274, December 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Reynolds:1997:IN

[RWW97] Paul F. Reynolds, Jr., Craig Williams, and Raymond R. Wagner, Jr. Isotach networks. IEEE Transactions on Parallel and Distributed Systems, 8 (4):337–348, April 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l0337.
pdf;   http://www.computer.
org/tpds/td1997/l0337abs.
htm.

Sunwoo:1993:SMP

[SA93] Myung Hoon Sunwoo and J. K. Aggarwal. A sliding memory plane array processor. IEEE Transactions on Parallel and Distributed Systems, 4(6):601– 612, June 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shukla:1994:FMP

[SA94] Shridhar B. Shukla and Dharma P. Agrawal. Framework for mapping periodic real-time applications on multicomputers. IEEE Transactions on Parallel and Distributed Systems, 5(7):778– 784, July 1994. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sarje:2009:PGA

[SA09] Abhinav Sarje and Srinivas Aluru. Parallel genomic alignments on the Cell Broadband Engine. IEEE Transactions on Parallel and Distributed Systems, 20(11):1600–1610, November 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sahni:2000:MMD

[Sah00a] S. Sahni. Matrix multiplication and data routing using a partitioned optical passive stars network. IEEE Transactions on Parallel and Distributed Systems, 11(7):720–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0720.
pdf;   http://www.computer.
org/tpds/td2000/l0720abs.
htm.

Sahni:2000:POP

[Sah00b]  S. Sahni.  The partitioned

optical passive stars network: Simulations and fundamental operations. IEEE Transactions on Parallel and Distributed Systems, 11(7):739–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0739.
pdf;   http://www.computer.
org/tpds/td2000/l0739abs.
htm.

Sarkar:1993:CTC

[Sar93] Dilip Sarkar. Cost and time-cost effectiveness of multiprocessing. IEEE Transactions on Parallel and Distributed Systems, 4 (6):704–712, June 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shih:1994:AMF

[SB94a] C. Jimmy Shih and Kenneth E. Batcher. Adding multiplefault tolerance to generalized cube networks. IEEE Transactions on Parallel and Distributed Systems, 5(8):785–792, August 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Spasojevic:1994:VOS

[SB94b] Mirjana Spasojevic and Piotr Berman. Voting as the optimal static pessimistic scheme for managing replicated data. IEEE Transactions on Parallel and Distributed Systems, 5 (1):64–73, January 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Subbiah:2004:DDD

[SB04] Arun Subbiah and Douglas M. Blough. Distributed diagnosis in dynamic fault environments. IEEE Transactions on Parallel and Distributed Systems, 15(5):453–467, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/

trans/td/2004/05/l0453abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0453.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0453.
pdf.

Seguel:2000:FDI

[SBF00] J. Seguel, D. Bollman, and J. Feo. A framework for the design and implementation of FFT permutation algorithms. IEEE Transactions on Parallel and Distributed Systems, 11(7):625– ??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0625.
pdf;   http://www.computer.
org/tpds/td2000/l0625abs.
htm.

Shahabi:2002:DRMa

[SBK02a] Cyrus Shahabi and Farnoush Banaei-Kashani. Decentralized resource management for a distributed continuous media server. IEEE Transactions on Parallel and Distributed Systems, 13(7):710–727, July 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0710.
pdf;   http://www.computer.
org/tpds/td2002/l0710abs.

htm. Editor’s Note: This paper unfortunately contains some errors which led to the paper being reprinted in the November 2002 issue. Please see IEEE Transactions on Parallel and Distributed Systems, vol. 13, no. 11, November 2002, pp. 1183–1200 for the correct paper.

Shahabi:2002:DRMb

[SBK02b] Cyrus Shahabi and Farnoush Banaei-Kashani. Decentralized resource management for a distributed continuous media server. IEEE Transactions on Parallel and Distributed Systems, 13(11):1183–1200, November 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2002/11/l1183abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1183.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2002/11/l1183.
                                        pdf.

Saikia:1998:ETS

[SBS98] D. K. Saikia, R. Badrinath, and R. K. Sen. Embedding torus on the star graph. IEEE Transactions on Parallel and Distributed Systems, 9(7):650– ??, July 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0650.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0650abs.
                                        htm.

Sheu:1991:SNL

[SC91] Jang-Ping Sheu and Chih-Yung Chang. Synthesizing nested loop algorithms using nonlinear transformation method. IEEE Transactions on Parallel and Distributed Systems, 2(3):304– 317, July 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scheiman:1992:PTM

[SC92] C. J. Scheiman and P. R. Cappello. A processor-time-minimal systolic array for transitive closure. IEEE Transactions on Parallel and Distributed Systems, 3(3):257–269, May 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shin:1993:SDL

[SC93] Kang G. Shin and Chih-Che Chou. Simple distributed loopfree routing strategy for computer communication networks. IEEE Transactions on Parallel and Distributed Systems, 4(12): 1308–1319, December 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scheiman:1994:PPT

[SC94] Chris Scheiman and Peter Cappello. Period-processor-time-minimal schedule for cubical mesh algorithms. IEEE Transactions on Parallel and Distributed Systems, 5(3):274–280, March 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2005:FBS

[SC05] Hyo Jung Song and Andrew A.

Chien. Feedback-based synchronization in system area networks for cluster computing. IEEE Transactions on Parallel and Distributed Systems, 16(10): 908–920, October 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Seshadri:2007:RQT

[SC07] Sangeetha Seshadri and Brian F. Cooper. Routing queries through a peer-to-peer InfoBeacons network using information retrieval techniques. IEEE Transactions on Parallel and Distributed Systems, 18(12): 1754–1765, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scarano:1999:SEF

[Sca99] V. Scarano. On the sizes of extended Fibonacci cubes. IEEE Transactions on Parallel and Distributed Systems, 10(7):764– ??, July 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0764.
pdf;   http://www.computer.
org/tpds/td1999/l0764abs.
htm.

Su:1997:SPR

[SCD97] Ming-Yang Su, Gen-Huey Chen, and Dyi-Rong Duh. A shortest path routing algorithm for incomplete WK-recursive networks. IEEE Transactions on Parallel and Distributed Systems, 8(4):367–379, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0367.
pdf;   http://www.computer.
org/tpds/td1997/l0367abs.
htm.

Scherson:1991:OGC

[Sch91] Isaac D. Scherson. Orthogonal graphs for the construction of a class of interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 2(1):3–19, January 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shin:2000:DRS

[SCK00] K. G. Shin, C.-C. Chou, and S.- K. Kweon. Distributed route selection for establishing realtime channels. IEEE Transactions on Parallel and Distributed Systems, 11(3):318–??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0318.
pdf;   http://www.computer.
org/tpds/td2000/l0318abs.
htm.

Sun:2000:RCR

[SCL00] Y. Sun, Paul Y. S. Cheung, and X. Lin. Recursive cube of rings: a new topology for interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 11(3):275– ??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0275.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0275abs.
                                        htm.

Sun:2001:BSW

[SCL01] Y. Sun, P. Y. S. Cheung, and X. Lin. Barrier synchronization on wormhole-routed networks. IEEE Transactions on Parallel and Distributed Systems, 12(6):583–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0583.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0583abs.
                                        htm.

Sendag:2005:IIS

[SCL05] Resit Sendag, Ying Chen, and David J. Lilja. The impact of incorrectly speculated memory operations in a multithreaded architecture. IEEE Transactions on Parallel and Distributed Systems, 16(3):271– 285, March 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Son:2007:CDE

[SCO^{+}07] Seung Woo Son, Guangyu Chen, Ozcan Ozturk, Mahmut Kandemir, and Alok Choudhary. Compiler-directed energy optimization for parallel disk based systems. IEEE Transactions on Parallel and Distributed Systems, 18(9):1241–1257, September 2007. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:1999:EFT

[SCP99] H. Shen, F. Chin, and Y. Pan.

Efficient fault-tolerant routing in multihop optical WDM networks. IEEE Transactions on Parallel and Distributed Systems, 10(10):12–??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1012.
pdf;   http://www.computer.
org/tpds/td1999/l1012abs.
htm.

Steensland:2002:ACC

[SCP02] Johan Steensland, Sumir Chandra, and Manish Parashar. An application-centric characterization of domain-based SFC partitioners for parallel SAMR. IEEE Transactions on Parallel and Distributed Systems, 13(12):1275–1289, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1275abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1275.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1275.
pdf.

Sonnek:2007:ARB

[SCW07] Jason Sonnek, Abhishek Chandra, and Jon Weissman. Adaptive reputation-based scheduling on unreliable distributed infrastructures. IEEE Transactions on Parallel and Distributed Systems, 18(11):1551–1564, November 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Siu:1996:NCD

[SCY96] Hin-Sing Siu, Yeh-Hao Chin, and Wei-Pang Yang. A note on consensus on dual failure modes. IEEE Transactions on Parallel and Distributed Systems, 7 (3):225–230, March 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0225abs.htm.

Siu:1998:BAP

[SCY98] H.-S. Siu, Y.-H. Chin, and W.- P. Yang. Byzantine agreement in the presence of mixed faults on processors and links. IEEE Transactions on Parallel and Distributed Systems, 9(4):335– ??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0335.
pdf;   http://www.computer.
org/tpds/td1998/l0335abs.
htm.

Silla:2000:HPR

[SD00a] F. Silla and J. Duato. Highperformance routing in networks of workstations with irregular topology. IEEE Transactions on Parallel and Distributed Systems, 11(7):699–??, July 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0699.
pdf;   http://www.computer.
org/tpds/td2000/l0699abs.
htm.

Silla:2000:UVC

[SD00b] F. Silla and J. Duato. On the use of virtual channels in networks of workstations with irregular topology. IEEE Transactions on Parallel and Distributed Systems, 11(8):813–??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0813.
pdf;   http://www.computer.
org/tpds/td2000/l0813abs.
htm.

Sarhan:2004:CSN

[SD04] Nabil J. Sarhan and Chita R.

Das. Caching and scheduling in NAD-based multimedia servers. IEEE Transactions on Parallel and Distributed Systems, 15(10):921–933, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/10/l0921abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0921.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0921.
pdf.

Suh:2000:SBR

[SDDY00] Y.-J. Suh, B. V. Dao, J. Duato, and S. Yalamanchili. Software-based rerouting for fault-tolerant pipelined communication. IEEE Transactions on Parallel and Distributed Systems, 11(3):193–??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0193.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0193abs.
                                        htm.

Summerville:1996:FBP

[SDFV96] Douglas H. Summerville, José G. Delgado-Frias, and Stamatis Vassiliadis. A flexible bitpattern associative router for interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 7(5):477– 485, May 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0477abs.
                                        htm.

Soliman:1998:AMH

[SE98] H. M. Soliman and A. S. Elmaghraby. An analytical model for hybrid checkpointing in time warp distributed simulation. IEEE Transactions on Parallel and Distributed Systems, 9(10):947–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0947.
pdf;   http://www.computer.
org/tpds/td1998/l0947abs.
htm.

Shang:1992:TMU

[SF92a] W. Shang and J. A. B. Fortes.

On time mapping of uniform dependence algorithms into lower dimensional processor arrays. IEEE Transactions on Parallel and Distributed Systems, 3 (3):350–363, May 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stunkel:1992:ACP

[SF92b] Craig B. Stunkel and W. Kent Fuchs. An analysis of cache performance for a hypercube multicomputer. IEEE Transactions on Parallel and Distributed Systems, 3(4):421–432, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Seo:1995:CBN

[SF95] Seung-Woo Seo and Tse-Yun Feng. The composite Banyan network. IEEE Transactions on Parallel and Distributed Systems, 6(10):1043–1054, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l1043abs.htm.

Song:2007:UBR

[SF07] Yougang Song and Brett D.

Fleisch. Utilizing binary rewriting for improving end-host security. IEEE Transactions on Parallel and Distributed Systems, 18 (12):1687–1699, December 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Santoro:2008:ODD

[SF08] Andrea Santoro and Richard M. Fujimoto. Offloading data distribution management to network processors in HLA-based distributed simulations. IEEE Transactions on Parallel and Distributed Systems, 19(3):289– 298, March 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shmueli:2009:SDP

[SF09] Edi Shmueli and Dror G. Feitelson. On simulation and design of parallel-systems schedulers: Are we doing the right thing? IEEE Transactions on Parallel and Distributed Systems, 20(7):983–996, July 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shan:2003:PEH

[SFP03] Wenhao Shan, Pingzhi Fan, and Yi Pan. Performance evaluation of a hierarchical cellular system with mobile velocitybased bidirectional call-overflow scheme. IEEE Transactions on Parallel and Distributed Systems, 14(1):72–83, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/01/l0072abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0072.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0072.
pdf.

Scott:1993:PPC

[SG93] Steven L. Scott and James R.

Goodman. Performance of pruning-cache directories for large-scale multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 4(5):520– 534, May 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scott:1994:IPC

[SG94] Steven L. Scott and James R.

Goodman. Impact of pipelined channels on k-ary n-cube networks. IEEE Transactions on Parallel and Distributed Systems, 5(1):2–16, January 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shehab:2008:SCM

[SGB08] Mohamed Shehab, Arif Ghafoor, and Elisa Bertino. Secure collaboration in a mediator-free distributed environment. IEEE Transactions on Parallel and Distributed Systems, 19(10): 1338–1351, October 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Srivatsa:2006:LSU

[SGL06] Mudhakar Srivatsa, Bugra Gedik, and Ling Liu. Large scaling unstructured peer-to-peer networks with heterogeneityaware topology and routing. IEEE Transactions on Parallel and Distributed Systems, 17(11): 1277–1293, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scrofano:2008:AMD

[SGTP08] Ronald Scrofano, Maya B. Gokhale, Frans Trouw, and Viktor K. Prasanna. Accelerating molecular dynamics simulations with reconfigurable computers. IEEE Transactions on Parallel and Distributed Systems, 19 (6):764–778, June 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shin:1993:AMA

[SH93] Kang G. Shin and Chao-Ju Hou. Analytic models of adaptive load sharing schemes in distributed real-time systems. IEEE Transactions on Parallel and Distributed Systems, 4(7):740–761, July 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shin:1994:DEE

[SH94] Kang G. Shin and Chao-Ju Hou. Design and evaluation of effective load sharing in distributed real-time systems. IEEE Transactions on Parallel and Distributed Systems, 5(7):704–719, July 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shang:1995:DHB

[SH95a] Shisheng Shang and Kai Hwang.

Distributed hardwired barrier synchronization for scalable multiprocessor clusters. IEEE Transactions on Parallel and Distributed Systems, 6(6):591– 605, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0591abs.
htm.

Stauffer:1995:SSO

[SH95b] L. M. Stauffer and D. S.

Hirschberg. Systolic selforganizing lists under transpose. IEEE Transactions on Parallel and Distributed Systems, 6(1): 102–105, January 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shin:1996:ELS

[SH96] Kang G. Shin and Chao-Ju Hou. Evaluation of load sharing in HARTS with consideration of its communication activities. IEEE Transactions on Parallel and Distributed Systems, 7 (7):724–739, July 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0724abs.htm.

Shieh:1997:CTO

[SH97] Shiuh-Pyng Shieh and Fu-Shen Ho. A comment on “A Total Ordering Multicast Protocol Using Propagation Trees”. IEEE Transactions on Parallel and Distributed Systems, 8(10):1084– ??, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1084.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1084abs.

htm. See [Jia95].

Singhal:1992:DIS

[Sin92] Mukesh Singhal. A dynamic information-structure mutual exclusion algorithm for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 3(1):121–125, January 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Singh:1996:LEP

[Sin96] Gurdip Singh. Leader election in the presence of link failures. IEEE Transactions on Parallel and Distributed Systems, 7 (3):231–236, March 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0231abs.htm.

Srinivasan:1999:SRD

[SJ99] S. Srinivasan and N. J. Jha. Safety and reliability driven task allocation in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 10(3):238–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0238.
pdf;   http://www.computer.
org/tpds/td1999/l0238abs.
htm.

Shaikh:2009:GBT

[SJd^{+}09] Riaz Ahmed Shaikh, Hassan Jameel, Brian J. d’Auriol, Heejo Lee, Sungyoung Lee, and Young-Jae Song. Group-based trust management scheme for clustered wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(11): 1698–1712, November 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sohail:2006:QDP

[SJKC06] Shaleeza Sohail, Sanjay K. Jha, Salil S. Kanhere, and Chun Tung Chou. QoS driven parallelization of resources to reduce file download delay. IEEE Transactions on Parallel and Distributed Systems, 17(10):1204– 1215, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Shao:2009:CTE

[SJM09] Shuyi Shao, Alex K. Jones, and Rami Melhem. Compiler techniques for efficient communications in circuit switched networks for multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 20(3):331–345, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Seo:2008:EES

[SJPL08] Euiseong Seo, Jinkyu Jeong, Seonyeong Park, and Joonwon Lee. Energy efficient scheduling of real-time tasks on multicore processors. IEEE Transactions on Parallel and Distributed Systems, 19(11):1540–1552, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sundar:2001:HAC

[SJPS01] N. S. Sundar, D. N. Jayasimha, D. K. Panda, and P. Sadayappan. Hybrid algorithms for complete exchange in 2D meshes. IEEE Transactions on Parallel and Distributed Systems, 12(12):1201–1218, December 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l1201.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l1201abs.
                                        htm.

Subhlok:1995:IPA

[SK95] Jaspal Subhlok and Ken Kennedy. Integer programming for array subscript analysis. IEEE Transactions on Parallel and Distributed Systems, 6(6):662– 668, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0662abs.
htm.

Seinstra:2002:PPP

[SK02] F. J. Seinstra and D. Koelma. P-3PC: a point-to-point communication model for automatic and optimal decomposition of regular domain problems. IEEE Transactions on Parallel and Distributed Systems, 13(7):758– 768, July 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0758.
pdf;   http://www.computer.
org/tpds/td2002/l0758abs.
htm.

Seinstra:2004:FSM

[SKB04] Frank J. Seinstra, Dennis Koelma, and Andrew D. Bagdanov. Finite state machinebased optimization of data parallel regular domain problems applied in low-level image processing. IEEE Transactions on Parallel and Distributed Systems, 15(10):865–877, October 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/10/l0865abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/10/l0865.
htm;   http://csdl.computer.

                                        org/dl/trans/td/2004/10/l0865.
                                        pdf.

Seshadri:2009:DSQ

[SKCL09] Sangeetha Seshadri, Vibhore Kumar, Brian Cooper, and Ling Liu. A distributed stream query optimization framework through integrated planning and deployment. IEEE Transactions on Parallel and Distributed Systems, 20(10):1439–1453, October 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shang:1994:LTG

[SKF94] Weijia Shang, Matthew T. Keefe, and Jose A. B. Fortes. On loop transformations for generalized cycle shrinking. IEEE Transactions on Parallel and Distributed Systems, 5(2):193– 204, February 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sabrina:2007:DAI

[SKJ07] Fariza Sabrina, Salil S. Kanhere, and Sanjay K. Jha. Design, analysis, and implementation of a novel low complexity scheduler for joint resource allocation. IEEE Transactions on Parallel and Distributed Systems, 18 (6):749–762, June 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Schloegel:2001:WDL

[SKK01] K. Schloegel, G. Karypis, and V. Kumar. Wavefront diffusion and LMSR: Algorithms for dynamic repartitioning of adaptive meshes. IEEE Transactions on Parallel and Distributed Systems, 12(5):451–??, May 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0451.
pdf;   http://www.computer.
org/tpds/td2001/l0451abs.
htm.

Shen:2003:HPA

[SkLC^{+}03] Xiaohui Shen, Wei keng Liao, Alok Choudhary, Gokhan Memik, and Mahmut Kandemir. A high-performance application data environment for largescale scientific computations. IEEE Transactions on Parallel and Distributed Systems, 14(12):1262–1274, December 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/12/l1262abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/12/l1262.
pdf.

Sivaram:2001:ASE

[SKPS01] R. Sivaram, R. Kesavan, D. K.

Panda, and C. B. Stunkel. Architectural support for efficient multicasting in irregular networks. IEEE Transactions on Parallel and Distributed Systems, 12(5):489–??, May 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

                                        td/books/td2001/pdf/l0489.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0489abs.
                                        htm.

Schurgers:2002:DDA

[SKS02] Curt Schurgers, Gautam Kulkarni, and Mani B. Srivastava. Distributed on-demand address assignment in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 13(10):1056–1065, October 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l1056.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l1056abs.
                                        htm.

Sih:1993:CTS

[SL93a] Gilbert C. Sih and Edward A. Lee. Compile-time scheduling heuristic for interconnectionconstrained heterogeneous processor architectures. IEEE Transactions on Parallel and Distributed Systems, 4(2):175– 187, February 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sih:1993:DNM

[SL93b] Gilbert C. Sih and Edward A. Lee. Declustering: a new multiprocessor scheduling technique. IEEE Transactions on Parallel and Distributed Systems, 4 (6):625–637, June 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Squillante:1993:UPC

[SL93c] Mark S. Squillante and Edward D. Lazowska. Using processor-cache affinity information in shared-memory multiprocessor scheduling. IEEE Transactions on Parallel and Distributed Systems, 4(2):131–143, February 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Seznec:1994:IPS

[SL94] Andre Seznec and Jacques Lenfant. Interleaved parallel schemes. IEEE Transactions on Parallel and Distributed Systems, 5(12):1329–1334, December 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2001:LFH

[SL01a] Ivan Stojmenovic and Xu Lin.

Loop-free hybrid single-path/ flooding routing algorithms with guaranteed delivery for wireless networks. IEEE Transactions on Parallel and Distributed Systems, 12(10):1023–1032, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1023.
pdf;   http://www.computer.
org/tpds/td2001/l1023abs.
htm.

Stojmenovic:2001:PAL

[SL01b] Ivan Stojmenovic and Xu Lin.

Power-aware localized routing in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 12(11):1122– 1133, November 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l1122.
pdf;   http://www.computer.
org/tpds/td2001/l1122abs.
htm.

Sodan:2006:LLM

[SL06] Angela C. Sodan and Lei Lan. LOMARC: Lookahead matchmaking for multiresource coscheduling on hyperthreaded CPUs. IEEE Transactions on Parallel and Distributed Systems, 17(11):1360–1375, November 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://bell.computer.org/
dlcomments/.

Srivatsa:2009:MDS

[SL09] Mudhakar Srivatsa and Ling Liu. Mitigating denial-of-service attacks on the Chord Overlay Network: a location hiding approach. IEEE Transactions on Parallel and Distributed Systems, 20(4):512–527, April 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sorin:2003:AES

[SLEV03] Daniel J. Sorin, Jonathan L. Lemon, Derek L. Eager, and Mary K. Vernon. Analytic evaluation of shared-memory architectures. IEEE Transactions on Parallel and Distributed Systems, 14(2):166–180, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/02/l0166abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0166.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0166.
pdf.

Song:2006:LTC

[SLFW06] Wen-Zhan Song, Xiang-Yang Li, Ophir Frieder, and WeiZhao Wang. Localized topology control for unicast and broadcast in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 17 (4):321–334, April 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sung:1997:MEF

[SLH97] Ting-Yi Sung, Men-Yang Lin, and Tung-Yang Ho. Multipleedge-fault tolerance with respect to hypercubes. IEEE Transactions on Parallel and Distributed Systems, 8(2):187–192, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0187.
pdf;   http://www.computer.
org/tpds/td1997/l0187abs.
htm.

Solihin:2003:CPU

[SLT03] Yan Solihin, Jaejin Lee, and Josep Torrellas. Correlation prefetching with a user-level memory thread. IEEE Transactions on Parallel and Distributed Systems, 14(6):563–580, June 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/06/l0563abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/06/l0563.
                                        pdf.

Shen:1990:ESF

[SLY90] Zhiyu Shen, Zhiyuan Li, and Pen-Chung Yew. Empirical study of Fortran programs for parallelizing compilers. IEEE Transactions on Parallel and Distributed Systems, 1(3):356– 364, July 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shrivastava:1994:SFT

[SM94] Santosh K. Shrivastava and Daniel L. McCue. Structuring fault-tolerant object systems for modularity in a distributed environment. IEEE Transactions on Parallel and Distributed Systems, 5(4):421–432, April 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Schollmeyer:1997:GMM

[SM97] Martina Schollmeyer and Bruce McMillin. A general method for maximizing the error-detecting ability of distributed algorithms. IEEE Transactions on Parallel and Distributed Systems, 8(2):164–172, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0164.
pdf;   http://www.computer.
org/tpds/td1997/l0164abs.
htm.

Surdeanu:2002:DPA

[SM02] Mihai Surdeanu and Dan Moldovan. Design and performance analysis of a distributed Java Virtual Machine. IEEE Transactions on Parallel and Distributed Systems, 13(6):611– 627, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0611.
pdf;   http://www.computer.
org/tpds/td2002/l0611abs.
htm.

Shim:2003:SPE

[SM03] SangMin Shim and Soo-Mook Moon. Split-path enhanced pipeline scheduling. IEEE Transactions on Parallel and Distributed Systems, 14(5):447– 462, May 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2003/05/
l0447abs.htm;   http://csdl.
computer.org/dl/trans/td/
2003/05/l0447.htm;   http:
//csdl.computer.org/dl/trans/
td/2003/05/l0447.pdf.

Shatz:1990:DIP

[SMBT90] Sol M. Shatz, Khanh Mai, Christopher Black, and Shengru Tu. Design and implementation of a Petri net based toolkit for Ada tasking analysis. IEEE Transactions on Parallel and Distributed Systems, 1 (4):424–441, October 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Surdeanu:2002:PAD

[SMH02] Mihai Surdeanu, Dan I. Moldovan, and Sanda M. Harabagiu. Performance analysis of a distributed question/answering system. IEEE Transactions on Parallel and Distributed Systems, 13(6):579–596, June 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/td/books/td2002/
                                        pdf/l0579.pdf;   http://www.
                                        computer.org/tpds/td2002/
                                        l0579abs.htm.

Sung:1992:MID

[SMJ92] Wonyong Sung, Sanjit K. Mitra, and Branko Jeren. Multiprocessor implementation of digital filtering algorithms using a parallel block processing method. IEEE Transactions on Parallel and Distributed Systems, 3(1): 110–120, January 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Som:1993:PPP

[SMS93] Sukhamoy Som, Roland R. Mielke, and John W. Stoughton.

Prediction of performance and processor requirements in realtime data flow architectures. IEEE Transactions on Parallel and Distributed Systems, 4(11): 1205–1216, November 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sultan:2002:LGCa

[SNI02a] Florin Sultan, Thu D. Nguyen, and Liviu Iftode. Lazy garbage collection of recovery state for fault-tolerant distributed shared memory. IEEE Transactions on Parallel and Distributed Systems, 13(7):673–686, July 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/computer.org/tpds/td2002/
l1085.pdf;   http://dlib.
computer.org/td/books/td2002/
pdf/l0673.pdf;   http://www.
computer.org/tpds/td2002/

l0673abs.htm. Editor’s Note: This paper unfortunately contains some errors which led to the paper being reprinted in the October 2002 issue. Please see IEEE Transactions on Parallel and Distributed Systems, vol. 13, no. 10, October 2002, pp. 1085-1098 for the correct paper.

Sultan:2002:LGCb

[SNI02b] Florin Sultan, Thu D. Nguyen, and Liviu Iftode. Lazy garbage collection of recovery state for fault-tolerant distributed shared memory. IEEE Transactions on Parallel and Distributed Systems, 13(10):1085–1098, October 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l1085.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l1085abs.
                                        htm.

Shao:1995:ENS

[SO95] Fong-Chih Shao and A. Yavuz Oruç. Efficient nonblocking switching networks for interprocessor communications in multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 6(2):132– 141, February 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0132abs.htm.

Sobalvarro:1996:AMM

[Sob96] Patrick G. Sobalvarro. Analytical modeling of multistage, multipath networks. IEEE Transactions on Parallel and Distributed Systems, 7(10):1059– 1064, October 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l1059abs.htm.

Sohn:1995:PAS

[Soh95] Andrew Sohn. Parallel Nary speculative computation of simulated annealing. IEEE Transactions on Parallel and Distributed Systems, 6(10):997– 1005, October 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0997abs.htm.

Solworth:2002:INB

[Sol02] Jon A. Solworth. Integrated network barriers. IEEE Transactions on Parallel and Distributed Systems, 13(4):337–348, April 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0337.
pdf;   http://www.computer.
org/tpds/td2002/l0337abs.
htm.

Suresh:2005:PIB

[SOM05] S. Suresh, S. N. Omkar, and V. Mani. Parallel implementation of back-propagation algorithm in networks of workstations. IEEE Transactions on Parallel and Distributed Systems, 16(1):24–34, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0024abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0024.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0024.
pdf.

Serrano:1993:OAA

[SP93] Mauricio J. Serrano and Behrooz Parhami. Optimal architectures and algorithms for meshconnected parallel computers with separable row/column buses. IEEE Transactions on Parallel and Distributed Systems, 4(10):1073–1080, October 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sharma:1995:PAH

[SP95] Debendra Das Sharma and Dhiraj K. Pradhan. Processor allocation in hypercube multicomputers: Fast and efficient strategies for cubic and noncubic allocation. IEEE Transactions on Parallel and Distributed Systems, 6(10):1108–1123, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1995/
                                        l1108abs.htm.

Sharma:1998:JSM

[SP98] D. Das Sharma and D. K. Pradhan. Job scheduling in mesh multicomputers. IEEE Transactions on Parallel and Distributed Systems, 9(1):57–70, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0057.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0057abs.
                                        htm.

Song:2003:PAH

[SP03] Yong Ho Song and Timothy Mark Pinkston. A progressive approach to handling message-dependent deadlock in parallel computer systems. IEEE Transactions on Parallel and Distributed Systems, 14(3):259–275, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0259abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0259.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0259.
pdf.

Song:2005:DRN

[SP05] Yong Ho Song and Timothy Mark Pinkston. Distributed resolution of network congestion and potential deadlock using reservation-based scheduling. IEEE Transactions on Parallel and Distributed Systems, 16 (8):686–701, August 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Soteriou:2007:EDS

[SP07] Vassos Soteriou and Li-Shiuan Peh. Exploring the design space of self-regulating power-aware on/off interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 18 (3):393–408, March 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2007/03/
l0393s.pdf.

Sorin:2002:SVB

[SPC^{+}02] Daniel J. Sorin, Manoj Plakal, Anne E. Condon, Mark D.

Hill, Milo M. K. Martin, and David A. Wood. Specifying and verifying a broadcast and a multicast snooping cache coherence protocol. IEEE Transactions on Parallel and Distributed Systems, 13(6):556–578, June 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0556.
pdf;   http://www.computer.
org/tpds/td2002/l0556abs.
htm.

Sun:1999:IRC

[SPF99] X.-H. Sun, M. Pantano, and T. Fahringer. Integrated range comparison for data-parallel compilation systems. IEEE Transactions on Parallel and Distributed Systems, 10(5):448– ??, May 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0448.
pdf;   http://www.computer.
org/tpds/td1999/l0448abs.
htm.

Sivaram:1998:EBM

[SPS98] R. Sivaram, D. K. Panda, and C. B. Stunkel. Efficient broadcast and multicast on multistage interconnection networks using multiport encoding. IEEE Transactions on Parallel and Distributed Systems, 9(10):4–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

td/books/td1998/pdf/l1004.
pdf;   http://www.computer.
org/tpds/td1998/l1004abs.
htm.

Soh:1991:CCA

[SR91] Sieteng Soh and Suresh Rai.

CAREL: Computer aided reliability evaluator for distributed computing networks. IEEE Transactions on Parallel and Distributed Systems, 2(2):199– 213, April 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:1994:SPA

[SR94] Xian-He Sun and Diane T.

Rover. Scalability of parallel algorithm-machine combinations. IEEE Transactions on Parallel and Distributed Systems, 5(6):599–613, June 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sengupta:1998:AAB

[SR98] A. Sengupta and C. S. Raghavendra. All-to-all broadcast and matrix multiplication in faulty SIMD hypercubes. IEEE Transactions on Parallel and Distributed Systems, 9(6):550–??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0550.
pdf;   http://www.computer.
org/tpds/td1998/l0550abs.
htm.

Suri:1999:ESS

[SR99] N. Suri and K. Ramamritham. Editorial: Special section on dependable realtime systems. IEEE Transactions on Parallel and Distributed Systems, 10(6):529–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0529.
pdf.

Sancho:2004:EMI

[SRD04] José C. Sancho, Antonio Robles, and José Duato. An effective methodology to improve the performance of the Up*/Down* routing algorithm. IEEE Transactions on Parallel and Distributed Systems, 15 (8):740–754, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/08/l0740.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/08/l0740.pdf;   http:
//csdl2.computer.org/comp/
trans/td/2004/08/l0740supp.
pdf.

Srivatsa:2008:PET

[SRD08] Mudhakar Srivatsa, Nithya Rajamani, and Murthy Devarakonda. A policy evaluation tool for multisite resource management. IEEE Transactions on Parallel and Distributed Systems, 19(10):1352–1366, October 2008. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Sohn:1998:OCC

[SRL98] J. Sohn, T. G. Robertazzi, and S. Luryi. Optimizing computing costs using divisible load analysis. IEEE Transactions on Parallel and Distributed Systems, 9(3):225–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0225.
pdf;   http://www.computer.
org/tpds/td1998/l0225abs.
htm.

Shen:1993:RRM

[SRS93] Chia Shen, Krithi Ramamritham, and John A. Stankovic. Resource reclaiming in multiprocessor real-time systems. IEEE Transactions on Parallel and Distributed Systems, 4(4):382– 397, April 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Soh:1994:ILB

[SRT94] Sieteng Soh, Suresh Rai, and Jerry L. Trahan. Improved lower bounds on the reliability of hypercube architectures. IEEE Transactions on Parallel and Distributed Systems, 5 (4):364–378, April 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Saha:1996:AAM

[SRT96] Debanjan Saha, Sampath Rangarajan, and Satish K. Tripathi.

An analysis of the average message overhead in replica control protocols. IEEE Transactions on Parallel and Distributed Systems, 7(10):1026–1034, October 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l1026abs.htm.

Shang:2004:LCS

[SRZF04] Yi Shang, Wheeler Ruml, Ying Zhang, and Markus Fromherz. Localization from connectivity in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 15(11):961–974, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/11/l0961abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l0961.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l0961.
pdf.

Scott:1990:UFM

[SS90] Steven L. Scott and Gurindar S. Sohi. The use of feedback in multiprocessors and its application to tree saturation control. IEEE Transactions on Parallel and Distributed Systems, 1(4): 385–398, October 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Selvakumar:1994:SPC

[SS94] S. Selvakumar and C. Siva Ram Murthy. Scheduling precedence constrained task graphs with non-negligible intertask communication onto multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 5(3): 328–336, March 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Saikia:1996:TRS

[SS96] Dilip K. Saikia and Ranjan K.

Sen. Two ranking schemes for efficient computation on the star interconnection network. IEEE Transactions on Parallel and Distributed Systems, 7 (4):321–327, April 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0321abs.htm.

Surma:2000:CRM

[SS00] D. R. Surma and E. H.-M.

Sha. Communication reduction in multiple multicasts based on hybrid static-dynamic scheduling. IEEE Transactions on Parallel and Distributed Systems, 11(9):865–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0865.
pdf;   http://www.computer.
org/tpds/td2000/l0865abs.
htm.

Suh:2001:AAP

[SS01] Y.-J. Suh and K. G. Shin. All-to-all personalized communication in multidimensional torus and mesh networks. IEEE Transactions on Parallel and Distributed Systems, 12(1):38–??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0038.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0038abs.
                                        htm.

Sinnen:2005:CCT

[SS05] Oliver Sinnen and Leonel A. Sousa. Communication contention in task scheduling. IEEE Transactions on Parallel and Distributed Systems, 16(6):503– 515, June 2005. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Steinder:2007:MDE

[SS07] Małgorzata Steinder and Adarshpal S. Sethi. Multidomain diagnosis of end-to-end service failures in hierarchically routed networks. IEEE Transactions on Parallel and Distributed Systems, 18(3):379–392, March 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sangireddy:2008:OLB

[SS08] Rama Sangireddy and Jatan Shah. Operand-load-based split pipeline architecture for high clock rate and commensurable IPC. IEEE Transactions on Parallel and Distributed Systems, 19(4):529–544, April 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2009:CBO

[SS09] David Sun and Chengzheng Sun. Context-based operational transformation in distributed collaborative editing systems. IEEE Transactions on Parallel and Distributed Systems, 20(10):1454–1470, October 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Singh:1991:TBA

[SSG91] Ajit Singh, Jonathan Schaeffer, and Mark Green. A templatebased approach to the generation of distributed applications using a network of workstations. IEEE Transactions on Parallel and Distributed Systems, 2 (1):52–67, January 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sivaram:2000:IMW

[SSP00] R. Sivaram, C. B. Stunkel, and D. K. Panda. Implementing multidestination worms in switch-based parallel systems: Architectural alternatives and their impact. IEEE Transactions on Parallel and Distributed Systems, 11(8):794–??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0794.
pdf;   http://www.computer.

org/tpds/td2000/l0794abs.
htm.

Sivaram:2002:HHP

[SSP02] Rajeev Sivaram, Craig B. Stunkel, and Dhabaleswar K. Panda. HIPIQS: a highperformance switch architecture using input queuing. IEEE Transactions on Parallel and Distributed Systems, 13(3):275– 289, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0275.
pdf;   http://www.computer.
org/tpds/td2002/l0275abs.
htm.

Shi:2009:MSS

[SSP^{+}09] Xudong Shi, Feiqi Su, Jih-Kwon Peir, Ye Xia, and Zhen Yang. Modeling and stack simulation of CMP cache capacity and accessibility. IEEE Transactions on Parallel and Distributed Systems, 20(12):1752–1763, December 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sivasubramaniam:1999:ADS

[SSRV99] A. Sivasubramaniam, A. Singla, U. Ramachandran, and H. Venkateswaran. An application-driven study of parallel system overheads and network bandwidth requirements. IEEE Transactions on Parallel and Distributed Systems, 10(3):193–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0193.
pdf;   http://www.computer.
org/tpds/td1999/l0193abs.
htm.

Sinnen:2006:TRT

[SSS06] Oliver Sinnen, Leonel Augusto Sousa, and Frode Eika Sandnes. Toward a realistic task scheduling model. IEEE Transactions on Parallel and Distributed Systems, 17(3):263–275, March 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sum:2003:AMA

[SSsLY03] John Sum, Hong Shen, Chi sing Leung, and G. Young. Analysis on a mobile agent-based algorithm for network routing and management. IEEE Transactions on Parallel and Distributed Systems, 14(3):193–202, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0193abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0193.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0193.
pdf.

Setia:1994:APA

[SST94] Sanjeev K. Setia, Mark S.

Squillante, and Satish K. Tripathi. Analysis of processor allocation in multiprogrammed, distributed-memory parallel processing systems.

IEEE Transactions on Parallel and Distributed Systems, 5 (4):401–420, April 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2002:DSN

[SSZ02] Ivan Stojmenovic, Mahtab Seddigh, and Jovisa Zunic. Dominating sets and neighbor elimination-based broadcasting algorithms in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 13(1):14–25, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0014.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0014abs.

htm. See comments and corrections [Sto04].

Swiecicka:2006:MSR

[SSZ06] Anna Swiecicka, Franciszek Seredynski, and Albert Y. Zomaya. Multiprocessor scheduling and rescheduling with use of cellular automata and artificial immune system support. IEEE Transactions on Parallel and Distributed Systems, 17 (3):253–262, March 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sheu:1991:PMN

[ST91] Jang-Ping Sheu and Tsu-Huei Tai. Partitioning and mapping nested loops on multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 2(4):430–439, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stamoulis:1993:ERS

[ST93] George D. Stamoulis and John N. Tsitsiklis. Efficient routing schemes for multiple broadcasts in hypercubes. IEEE Transactions on Parallel and Distributed Systems, 4(7):725– 739, July 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Soch:1999:TOG

[ST99a] M. Soch and P. Tvrdík. Timeoptimal gossip of large packets in noncombining 2D tori and meshes. IEEE Transactions on Parallel and Distributed Systems, 10(12):252–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1252.
pdf;   http://www.computer.
org/tpds/td1999/l1252abs.
htm.

Stomp:1999:CRT

[ST99b] F. Stomp and G. Taubenfeld. Constructing a reliable test&set bit. IEEE Transactions on Parallel and Distributed Systems, 10(3):252–??, March 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

                                        td/books/td1999/pdf/l0252.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0252abs.
                                        htm.

Stankovic:1998:E

[Sta98] John A. Stankovic. Editorial. IEEE Transactions on Parallel and Distributed Systems, 9(1):1– ??, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0001.
                                        pdf.

Stankovic:1999:E

[Sta99] J. A. Stankovic. Editorial. IEEE Transactions on Parallel and Distributed Systems, 10(1):1–??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0001.

pdf. See correction [Ano99g].

Stankovic:2000:E

[Sta00] J. A. Stankovic. Editorial. IEEE Transactions on Parallel and Distributed Systems, 11(1):1–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0001.
                                        pdf.

Stankovic:2001:E

[Sta01] J. A. Stankovic. Editorial. IEEE Transactions on Parallel and Distributed Systems, 12(1):1–??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0001.
pdf.

Stankovic:2002:E

[Sta02] J. A. Stankovic. Editorial. IEEE Transactions on Parallel and Distributed Systems, 13(1):1–3, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0001.
pdf;   http://www.computer.
org/tpds/td2001/l0001abs.
htm.

Steenkiste:1996:NBM

[Ste96]  Peter Steenkiste.  Network-

based multicomputers: a practical supercomputer architecture. IEEE Transactions on Parallel and Distributed Systems, 7(8): 861–875, August 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0861abs.htm.

Shatz:1996:APN

[STMD96] Sol M. Shatz, Shengru Tu, Tadao Murata, and Sastry Duri. Application of Petri net reduction for Ada tasking deadlock analysis. IEEE Transactions on Parallel and Distributed Systems, 7(12):1307–1322, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:1996:CTB

[Sto96] Ivan Stojmenovic. Constant time BSR solutions to parenthesis matching, tree decoding, and tree reconstruction from its traversals. IEEE Transactions on Parallel and Distributed Systems, 7(2):218–224, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0218abs.htm.

Stojmenovic:1997:HNT

[Sto97] I. Stojmenovic. Honeycomb networks: Topological properties and communication algorithms. IEEE Transactions on Parallel and Distributed Systems, 8(10):1036–1042, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1036.
pdf;   http://www.computer.
org/tpds/td1997/l1036abs.
htm.

Stojmenovic:2004:CCD

[Sto04]  Ivan Stojmenovic.  Com-

ments and corrections to “Dominating Sets and Neighbor Elimination-Based Broadcasting Algorithms in Wireless Networks”. IEEE Transactions on Parallel and Distributed Systems, 15(11):1054–1055, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/

trans/td/2004/11/l1054abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1054.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1054.

pdf. See [SSZ02].

Steiner:2000:KAD

[STW00] M. Steiner, G. Tsudik, and M. Waidner. Key agreement in dynamic peer groups. IEEE Transactions on Parallel and Distributed Systems, 11(8):769– ??, August 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0769.
pdf;   http://www.computer.
org/tpds/td2000/l0769abs.
htm.

Saxena:2009:ENA

[STY09] Nitesh Saxena, Gene Tsudik, and Jeong Hyun Yi. Efficient node admission and certificateless secure communication in short-lived MANETs. IEEE Transactions on Parallel and Distributed Systems, 20(2): 158–170, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2002:ORF

[Sun02] Chengzheng Sun. Optional and responsive fine-grain locking in Internet-based collaborative systems. IEEE Transactions on Parallel and Distributed Systems, 13(9):994–1008, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0994.
pdf;   http://www.computer.
org/tpds/td2002/l0994abs.
htm.

Sharma:1997:CSI

[SV97] Vishal Sharma and Emmanouel A. Varvarigos. Circuit switching with input queuing: An analysis for the d-dimensional wraparound mesh and the hypercube. IEEE Transactions on Parallel and Distributed Systems, 8(4):349–366, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0349.
pdf;   http://www.computer.
org/tpds/td1997/l0349abs.
htm.

Spinnato:2004:PMD

[SvAS04] Piero F. Spinnato, G. D. van Albada, and Peter M. A. Sloot. Performance modeling of distributed hybrid architectures. IEEE Transactions on Parallel and Distributed Systems, 15(1):81–92, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/01/l0081abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/01/l0081.
pdf.

Shah:2007:DAD

[SVM07] Ruchir Shah, Bhardwaj Veeravalli, and Manoj Misra. On the design of adaptive and decentralized load balancing algorithms with load estimation for computational Grid environments. IEEE Transactions on Parallel and Distributed Systems, 18(12): 1675–1686, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scarpazza:2008:EBF

[SVP08] Daniele Paolo Scarpazza, Oreste Villa, and Fabrizio Petrini. Efficient breadth-first search on the Cell/BE processor. IEEE Transactions on Parallel and Distributed Systems, 19(10):1381– 1395, October 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Storms:2005:PDA

[SvVB05] Patrick P. A. Storms, J. Bernard van Veelen, and Erik Boasson. A process distribution approach for multisensor data fusion systems based on geographical dataspace partitioning. IEEE Transactions on Parallel and Distributed Systems, 16(1):14–23, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0014abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0014.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0014.
pdf.

Suen:1992:ETM

[SW92] Tony T. Y. Suen and Johnny S. K. Wong. Efficient task migration algorithm for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 3(4):488–499, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shu:1995:APS

[SW95] Wei Shu and Min-You Wu. Asynchronous problems on SIMD parallel computers. IEEE Transactions on Parallel and Distributed Systems, 6(7):704– 713, July 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1995/l0704abs.
                                        htm.

Shu:1996:RIP

[SW96] Wei Shu and Min-You Wu. Runtime incremental parallel scheduling (RIPS) on distributed memory computers. IEEE Transactions on Parallel and Distributed Systems, 7 (6):637–649, June 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0637abs.htm.

Sheu:1995:OBA

[SWC95] Jang-Ping Sheu, Chao-Tsung Wu, and Tzung-Shi Chen. An optimal broadcasting algorithm without message redundancy in star graphs. IEEE Transactions on Parallel and Distributed Systems, 6(6):653–658, June 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0653abs.htm.

Sistla:1998:MCC

[SWH98] A. P. Sistla, O. Wolfson, and Y. Huang. Minimization of communication cost through caching in mobile environments. IEEE Transactions on Parallel and Distributed Systems, 9(4):378– ??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0378.
pdf;   http://www.computer.
org/tpds/td1998/l0378abs.
htm.

Sengar:2008:DVF

[SWWJ08] Hemant Sengar, Haining Wang, Duminda Wijesekera, and Sushil Jajodia. Detecting VoIP floods using the Hellinger distance. IEEE Transactions on Parallel and Distributed Systems, 19 (6):794–805, June 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sung:2003:ITB

[SX03] Minho Sung and Jun Xu.

IP traceback-based intelligent packet filtering: a novel technique for defending against Internet DDoS attacks. IEEE Transactions on Parallel and Distributed Systems, 14(9):861– 872, September 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2003/09/
                                        l0861abs.htm;   http://csdl.
                                        computer.org/dl/trans/td/
                                        2003/09/l0861.pdf.

Shen:2007:LAC

[SX07] Haiying Shen and Cheng-Zhong Xu. Locality-aware and churnresilient load-balancing algorithms in structured peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 18(6):849– 862, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Stewart:2008:ELP

[SX08] Iain A. Stewart and Yonghong Xiang. Embedding long paths in k-ary n-cubes with faulty nodes and links. IEEE Transactions on Parallel and Distributed Systems, 19(8):1071–1085, August 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stewart:2009:BBA

[SX09] Iain A. Stewart and Yonghong Xiang. Bipanconnectivity and bipancyclicity in k-ary n-cubes. IEEE Transactions on Parallel and Distributed Systems, 20 (1):25–33, January 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stamos:1993:SFR

[SY93] James W. Stamos and Honesty C. Young. Symmetric fragment and replicate algorithm for distributed joins. IEEE Transactions on Parallel and Distributed Systems, 4(12):1345–1354, December 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:1997:BNU

[SY97] Jian Song and O. Yang. Backbone networks using rotation counters. IEEE Transactions on Parallel and Distributed Systems, 8(12):1288–1298, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1288.
pdf;   http://www.computer.
org/tpds/td1997/l1288abs.
htm.

Suh:1998:AAC

[SY98] Y.-J. Suh and S. Yalamanchili. All-to-all communication with minimum start-up costs in 2D/3D tori and meshes. IEEE Transactions on Parallel and Distributed Systems, 9(5):442– ??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0442.
pdf;   http://www.computer.
org/tpds/td1998/l0442abs.
htm.

Suh:2000:CAC

[SY00] Y.-J. Suh and S. Yalamanchili. Configurable algorithms for complete exchange in 2D meshes. IEEE Transactions on Parallel and Distributed Systems, 11(4):337–??, April 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0337.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0337abs.
                                        htm.

Shan:2007:BMS

[SY07] Tong Shan and Oliver Yang. Bandwidth management for supporting differentiated service aware traffic engineering. IEEE Transactions on Parallel and Distributed Systems, 18(9):1320– 1331, September 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2003:CSR

[SYC03] Kai Shen, Tao Yang, and Lingkun Chu. Clustering support and replication management for scalable network services. IEEE Transactions on Parallel and Distributed Systems, 14(11):1168–1179, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/11/l1168abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/11/l1168.
                                        pdf.

Seo:1999:PRF

[SyFL99] S.-W. Seo, T. y. Feng, and H.-I. Lee. Permutation realizability and fault tolerance property of the inside-out routing algorithm. IEEE Transactions on Parallel and Distributed Systems, 10(9):946–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0946.
pdf;   http://www.computer.
org/tpds/td1999/l0946abs.
htm.

Shin:1995:FMS

[SZ95a] Kang G. Shin and Qin Zheng.

FDDI-M: a scheme to double FDDI’s ability of supporting synchronous traffic. IEEE Transactions on Parallel and Distributed Systems, 6(11):1125– 1131, November 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1125abs.htm.

Sun:1995:PCS

[SZ95b] Xian-He Sun and Jianping Zhu. Performance considerations of shared virtual memory machines. IEEE Transactions on Parallel and Distributed Systems, 6(11):1185– 1194, November 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.

computer.org/tpds/td1995/
l1185abs.htm.

Seredynski:2002:SPC

[SZ02] Franciszek Seredynski and Albert Y. Zomaya. Sequential and parallel cellular automatabased scheduling algorithms. IEEE Transactions on Parallel and Distributed Systems, 13(10): 1009–1023, October 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2002/pdf/l1009.
pdf;   http://www.computer.
org/tpds/td2002/l1009abs.
htm.

Subrata:2003:CTA

[SZ03a] Riky Subrata and Albert Y. Zomaya. A comparison of three artificial life techniques for reporting cell planning in mobile computing. IEEE Transactions on Parallel and Distributed Systems, 14(2):142–153, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/02/l0142abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0142.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0142.
pdf.

Subrata:2003:ECA

[SZ03b] Riky Subrata and Albert Y. Zomaya. Evolving cellular automata for location management in mobile computing networks. IEEE Transactions on Parallel and Distributed Systems, 14(1):13–26, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/01/l0013abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0013.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0013.
pdf.

Sun:2004:PTL

[SZ04] Xian-He Sun and Wu Zhang. A parallel two-level hybrid method for tridiagonal systems and its application to fast Poisson solvers. IEEE Transactions on Parallel and Distributed Systems, 15(2):97–106, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0097abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0097.
pdf.

Subrata:2008:GTA

[SZ08] Riky Subrata and Albert Y.

Zomaya. Game-theoretic approach for load balancing in computational grids. IEEE Transactions on Parallel and Distributed Systems, 19(1):66– 76, January 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shao:2005:EAS

[SZXS05] Zili Shao, Qingfeng Zhuge, Chun Xue, and Edwin H.-M. Sha. Efficient assignment and scheduling for heterogeneous DSP systems. IEEE Transactions on Parallel and Distributed Systems, 16 (6):516–525, June 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Takesue:1993:FPP

[Tak93] Masaru Takesue. Family of parallel prefix algorithms embedded in networks. IEEE Transactions on Parallel and Distributed Systems, 4(10):1179–1184, October 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Taufer:2006:PPS

[TAKB06] Michela Taufer, Chahm An, Andreas Kerstens, and Charles L.

                                        Brooks III. Predictor@Home:

a “Protein Structure Prediction Supercomputer” based on global computing. IEEE Transactions on Parallel and Distributed Systems, 17(8):786–796, August 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Teeuw:1993:CVD

[TB93] Wouter B. Teeuw and Henk M. Blanken. Control versus data flow in parallel database machines. IEEE Transactions on Parallel and Distributed Systems, 4(11):1265–1279, November 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsay:1994:FTA

[TB94] Yih-Kuen Tsay and Rajive L.

Bagrodia. Fault-tolerant algorithms for fair interprocess synchronization. IEEE Transactions on Parallel and Distributed Systems, 5(7):737–748, July 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tamir:1993:SCA

[TC93] Yuval Tamir and Hsin-Chou Chi.

Symmetric crossbar arbiters for VLSI communication switches. IEEE Transactions on Parallel and Distributed Systems, 4 (1):13–27, January 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tzeng:1994:PSF

[TC94] Nian-Feng Tzeng and Po-Jen Chuang. Pairwise substitutional fault tolerance technique for the cube-connected cycles architecture. IEEE Transactions on Parallel and Distributed Systems, 5(4):433–438, April 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsay:1995:DER

[TC95a] Jong-Chuang Tsay and Pen-Yuang Chang. Design of efficient regular arrays for matrix multiplication by two-step regularization. IEEE Transactions on Parallel and Distributed Systems, 6(2):215–222, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0215abs.htm.

Tsay:1995:SND

[TC95b] Jong-Chuang Tsay and Pen-Yuang Chang. Some new designs of 2-D array for matrix multiplication and transitive closure. IEEE Transactions on Parallel and Distributed Systems, 6 (4):351–362, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0351abs.htm.

Tzeng:1998:FCH

[TC98] Nian-Feng Tzeng and Hsing-Lung Chen. Fast compaction in hypercubes. IEEE Transactions on Parallel and Distributed Systems, 9(1):50–56, January 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0050.
pdf;   http://www.computer.
org/tpds/td1998/l0050abs.
htm.

Tang:2004:MCD

[TC04a] Xueyan Tang and Samuel T. Chanson. The minimal cost distribution tree problem for recursive expiration-based consistency management. IEEE Transactions on Parallel and Distributed Systems, 15(3):214– 227, March 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/03/
l0214abs.htm;   http://csdl.
computer.org/comp/trans/td/
2004/03/L0214s.pdf;   http:
//csdl.computer.org/dl/trans/
td/2004/03/l0214.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/03/l0214.pdf.

Tang:2004:MCR

[TC04b] Xueyan Tang and Samuel T.

Chanson. Minimal cost replication of dynamic Web contents under flat update delivery. IEEE Transactions on Parallel and Distributed Systems, 15(5):431–439, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/05/l0431abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0431.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0431.
pdf.

Tang:2006:ARP

[TC06] Xueyan Tang and Samuel T.

Chanson. Analysis of replica placement under expirationbased consistency management. IEEE Transactions on Parallel and Distributed Systems, 17(11): 1253–1263, November 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2007:ASB

[TC07] Yong Tang and Shigang Chen.

An automated signature-based approach against polymorphic Internet worms. IEEE Transactions on Parallel and Distributed Systems, 18(7):879–892, July 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Teng:2005:IWC

[TCC05] Wei-Guang Teng, Cheng-Yue Chang, and Ming-Syan Chen. Integrating Web caching and Web prefetching in client-side proxies. IEEE Transactions on Parallel and Distributed Systems, 16(5):444–455, May 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2007:ORP

[TCC07] Xueyan Tang, Huicheng Chi, and Samuel T. Chanson. Optimal replica placement under TTL-based consistency. IEEE Transactions on Parallel and Distributed Systems, 18(3):351– 363, March 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2007/03/
                                        l0351s.pdf.

Tan:2007:DTA

[TCLY07] Chee Wei Tan, Dah-Ming Chiu, John C. S. Lui, and David K. Y. Yau. A distributed throttling approach for handling high bandwidth aggregates. IEEE Transactions on Parallel and Distributed Systems, 18(7):983–995, July 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2001:ECI

[TCO01] Kian-Lee Tan, Jun Cai, and Beng Chin Ooi. An evaluation of cache invalidation strategies in wireless environments. IEEE Transactions on Parallel and Distributed Systems, 12(8):789– 807, August 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0789.
pdf;   http://www.computer.
org/tpds/td2001/l0789abs.
htm.

Thakur:1996:EAA

[TCR96] Rajeev Thakur, Alok Choudhary, and J. Ramanujam. Efficient algorithms for array redistribution. IEEE Transactions on Parallel and Distributed Systems, 7(6):587–594, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0587abs.htm.

Tseng:1997:FTR

[TCS97] Yu-Chee Tseng, Shu-Hui Chang, and Jang-Ping Sheu. Faulttolerant ring embedding in a star graph with both link and node failures. IEEE Transactions on Parallel and Distributed Systems, 8(12):1185–1195, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1185.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1185abs.
                                        htm.

Thanalapati:2001:EAS

[TD01] Thyagaraj Thanalapati and Sivarama Dandamudi. An efficient adaptive scheduling scheme for distributed memory multicomputers. IEEE Transactions on Parallel and Distributed Systems, 12(7):758–768, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0758.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0758abs.
                                        htm.

Tsafrir:2007:BUS

[TEF07] Dan Tsafrir, Yoav Etsion, and Dror G. Feitelson. Backfilling using system-generated predictions rather than user runtime estimates. IEEE Transactions on Parallel and Distributed Systems, 18(6):789–803, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Theel:1996:DCP

[TF96a] Oliver E. Theel and Brett D. Fleisch. A dynamic coherence protocol for distributed shared memory enforcing high data availability at low costs. IEEE Transactions on Parallel and Distributed Systems, 7(9):

915–930, September 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0915abs.htm.

Tzeng:1996:RAC

[TF96b] Nian-Feng Tzeng and Gui-Liang Feng. Resource allocation in cube network systems based on the covering radius. IEEE Transactions on Parallel and Distributed Systems, 7 (4):328–342, April 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0328abs.htm.

Tari:2001:CAD

[TF01] Z. Tari and A. Fry. Controlling aggregation in distributed object systems: a graph-based approach. IEEE Transactions on Parallel and Distributed Systems, 12(12):1236–1255, December 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1236.
pdf;   http://www.computer.
org/tpds/td2001/l1236abs.
htm.

Tseng:1996:AAP

[TG96] Yu-Chee Tseng and Sandeep K. S. Gupta. All-to-all personalized communication in a wormhole-routed torus. IEEE Transactions on Parallel and Distributed Systems, 7(5):498– 505, May 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0498abs.
                                        htm.

Tseng:1999:CGA

[TG99] E. H.-Y. Tseng and J.-L. Gaudiot. Communication generation for aligned and cyclic(K) distributions using integer lattice. IEEE Transactions on Parallel and Distributed Systems, 10(2):136–??, February 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0136.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0136abs.
                                        htm.

Tang:2008:EET

[TGV08] Qinghui Tang, Sandeep Kumar S. Gupta, and Georgios Varsamopoulos. Energy-efficient thermal-aware task scheduling for homogeneous highperformance computing data centers: a cyber-physical approach. IEEE Transactions on Parallel and Distributed Systems, 19(11):1458–1472, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tao:1993:NCE

[TH93] D. L. Tao and C. R. P. Hartmann. Novel concurrent error detection scheme for FFT networks. IEEE Transactions on Parallel and Distributed Systems, 4(2):198–221, February 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tu:2001:FOF

[TH01] H.-Y. Tu and L. W. Hawkes.

Families of optimal faulttolerant multiple-bus networks. IEEE Transactions on Parallel and Distributed Systems, 12 (1):60–??, January 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l0060.
pdf;   http://www.computer.
org/tpds/td2001/l0060abs.
htm.

Tirthapura:2006:SSD

[TH06] Srikanta Tirthapura and Maurice Herlihy. Self-stabilizing distributed queuing. IEEE Transactions on Parallel and Distributed Systems, 17(7):646–655, July 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tao:1996:NED

[THH96] D. L. Tao, C. R. P. Hartmann, and Yunghsing S. (Sam) Han. New encoding/decoding methods for designing fault-tolerant matrix operations. IEEE Transactions on Parallel and Distributed Systems, 7(9):931–938, September 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.

                                        org/tpds/td1996/l0931abs.
                                        htm.

Tapolcai:2008:TNA

[THH08] János Tapolcai, Pin-Han Ho, and Anwar Haque. TROP: a novel approximate link-state dissemination framework for dynamic survivable routing in MPLS networks. IEEE Transactions on Parallel and Distributed Systems, 19(3):311–322, March 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Thomasian:1993:DNR

[Tho93]  Alexander Thomasian.  De-

termining the number of remote sites accessed in distributed transaction processing. IEEE Transactions on Parallel and Distributed Systems, 4 (1):99–103, January 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Thomasian:2006:CRP

[Tho06] Alexander Thomasian. Comment on “RAID5 Performance with Distributed Sparing”. IEEE Transactions on Parallel and Distributed Systems, 17 (4):399–400, April 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/. See [TM97].

Tsai:1997:SAP

[THT^{+}97] Horng-Ren Tsai, Shi-Jinn Horng, Shun-Shan Tsai, Tzong-Wann Kao, and Shung-Shing Lee. Solving an algebraic path problem and some related graph problems on a hyper-bus broadcast network. IEEE Transactions on Parallel and Distributed Systems, 8(12):1226–1235, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1226.
pdf;   http://www.computer.
org/tpds/td1997/l1226abs.
htm.

Topcuoglu:2002:PEL

[THW02] H. Topcuoglu, S. Hariri, and M. Y. Wu. Performanceeffective and low-complexity task scheduling for heterogeneous computing. IEEE Transactions on Parallel and Distributed Systems, 13(3):260–274, March 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0260.
pdf;   http://www.computer.
org/tpds/td2002/l0260abs.
htm.

Tan:2007:IFR

[TJ07] Guang Tan and Stephen A.

Jarvis. Improving the fault resilience of overlay multicast for media streaming. IEEE Transactions on Parallel and Distributed Systems, 18(6):721–734, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2008:PBI

[TJ08] Guang Tan and Stephen A. Jarvis. A payment-based incentive and service differentiation scheme for peer-to-peer streaming broadcast. IEEE Transactions on Parallel and Distributed Systems, 19(7):940–953, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tung:1996:UFS

[TK96a] Brian Tung and Leonard Kleinrock. Using finite state automata to produce self-optimization and self-control. IEEE Transactions on Parallel and Distributed Systems, 7(4):439–448, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0439abs.htm.

Tzeng:1996:TAS

[TK96b] Nian-Feng Tzeng and Harish Kumar. Traffic analysis and simulation performance of incomplete hypercubes. IEEE Transactions on Parallel and Distributed Systems, 7(7):740– 754, July 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0740abs.
                                        htm.

Tsanakas:2000:CGM

[TKP00] P. Tsanakas, N. Koziris, and G. Papakonstantinou. Chain grouping: a method for partitioning loops onto meshconnected processor arrays. IEEE Transactions on Parallel and Distributed Systems, 11(9): 941–??, September 2000. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2000/pdf/l0941.
pdf;   http://www.computer.
org/tpds/td2000/l0941abs.
htm.

Tong:1992:RRD

[TKT92] Zhijun Tong, Richard Y. Kain, and W. T. Tsai. Rollback recovery in distributed systems using loosely synchronized clocks. IEEE Transactions on Parallel and Distributed Systems, 3(2): 246–251, March 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tuah:2002:POP

[TKVD02] Nor Jaidi Tuah, Mohan Kumar, Svetha Venkatesh, and Sajal K. Das. Performance optimization problem in speculative prefetching. IEEE Transactions on Parallel and Distributed Systems, 13(5):471–484, May 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0471.
pdf;   http://www.computer.
org/tpds/td2002/l0471abs.
htm.

Tsai:1998:TAC

[TKW98] J. Tsai, S.-Y. Kuo, and Y.-M.

Wang. Theoretical analysis for communication-induced checkpointing protocols with rollback dependency trackability. IEEE Transactions on Parallel and Distributed Systems, 9(10):963– ??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0963.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0963abs.
                                        htm.

Thomasian:2005:CIC

[TL05] Alexander Thomasian and Chang Liu. Comment on em Issues and Challenges in the Performance Analysis of Real Disk Arrays. IEEE Transactions on Parallel and Distributed Systems, 16(11):1103–1104, November 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Theiss:2006:FFT

[TL06] Ingebjorg Theiss and Olav Lysne. FRoots: a fault tolerant and topology-flexible routing technique. IEEE Transactions on Parallel and Distributed Systems, 17(10):1136–1150, October 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tseng:1997:BOC

[TLGP97] Yu-Chee Tseng, Ting-Hsien Lin, Sandeep K. S. Gupta, and Dhabaleswar K. Panda. Bandwidthoptimal complete exchange on wormhole-routed 2D/3D torus networks: a diagonalpropagation approach. IEEE Transactions on Parallel and Distributed Systems, 8(4):380– 396, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0380.
pdf;   http://www.computer.
org/tpds/td1997/l0380abs.
htm.

Thottethodi:2004:EGK

[TLM04] Mithuna Thottethodi, Alvin R.

Lebeck, and Shubhendu S. Mukherjee. Exploiting global knowledge to achieve self-tuned congestion control for k-ary ncube networks. IEEE Transactions on Parallel and Distributed Systems, 15(3):257–272, March 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/03/l0257abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0257.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0257.
pdf.

Thomasian:1997:RPD

[TM97] Alexander Thomasian and Jai Menon. RAID5 performance with distributed sparing. IEEE Transactions on Parallel and Distributed Systems, 8(6):640– 657, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

                                        td/books/td1997/pdf/l0640.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0640abs.

htm. See [Tho06].

Tang:2006:EOU

[TM06] Chiping Tang and Philip K. McKinley. Energy optimization under informed mobility. IEEE Transactions on Parallel and Distributed Systems, 17(9): 947–962, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Torrie:1996:CMB

[TMTH96] Evan Torrie, Margaret Martonosi, Chau-Wen Tseng, and Mary W. Hall. Characterizing the memory behavior of compilerparallelized applications. IEEE Transactions on Parallel and Distributed Systems, 7(12):1224– 1237, December 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tzen:1993:DUL

[TN93a] Ten H. Tzen and Lionel M. Ni. Dependence uniformization: a loop parallelization technique. IEEE Transactions on Parallel and Distributed Systems, 4 (5):547–558, May 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tzen:1993:TSS

[TN93b] Ten H. Tzen and Lionel M. Ni. Trapezoid self-scheduling. A practical scheduling scheme for parallel compilers. IEEE Transactions on Parallel and Distributed Systems, 4(1):87–98, January 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tran:2008:LWS

[TN08] Duc A. Tran and Thinh Nguyen.

Localization in wireless sensor networks based on support vector machines. IEEE Transactions on Parallel and Distributed Systems, 19(7):981–994, July 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsuchiya:2001:SMC

[TNPK01] T. Tsuchiya, S. Nagano, R. B.

Paidi, and T. Kikuno. Symbolic model checking for selfstabilizing algorithms. IEEE Transactions on Parallel and Distributed Systems, 12(1):81– ??, January 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0081.
pdf;   http://www.computer.
org/tpds/td2001/l0081abs.
htm.

Tosun:2007:ACR

[Tos07] Ali Saman Tosun. Analysis and comparison of replicated declustering schemes. IEEE Transactions on Parallel and Distributed Systems, 18(11):1578– 1591, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tout:1995:DLB

[TP95] Walid R. Tout and Sakti Pramanik. Distributed load balancing for parallel main memory hash join. IEEE Transactions on Parallel and Distributed Systems, 6(8):841–849, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1995/
                                        l0841abs.htm.

Tseng:1996:TBM

[TPL96] Yu-Chee Tseng, Dhabaleswar K. Panda, and Ten-Hwang Lai. A trip-based multicasting model in wormhole-routed networks with virtual channels. IEEE Transactions on Parallel and Distributed Systems, 7(2):138–150, February 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

                                        computer.org/tpds/td1996/
                                        l0138abs.htm.

Tien:1993:ABS

[TR93] Sing-Ban Tien and C. S. Raghavendra. Algorithms and bounds for shortest paths and diameter in faulty hypercubes. IEEE Transactions on Parallel and Distributed Systems, 4 (6):713–718, June 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tran:2004:AAT

[TR04] Nancy Tran and Daniel A. Reed. Automatic ARIMA time series modeling for adaptive I/O prefetching. IEEE Transactions on Parallel and Distributed Systems, 15(4):362–377, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0362abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0362.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0362.
pdf.

Tran:2006:CAR

[TR06] Duc A. Tran and Harish Raghavendra. Congestion adaptive routing in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 17(11):1294–1305, November 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Towsley:1990:AFJ

[TRS90] Don Towsley, C. Gary Rommel, and John A. Stankovic. Analysis of fork-join program response times on multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 1 (3):286–303, July 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:1998:SMH

[TS98] M. Tan and H. J. Siegel. A stochastic model for heterogeneous computing and its application in data relocation scheme development. IEEE Transactions on Parallel and Distributed Systems, 9(11):88–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l1088.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1088abs.
                                        htm.

Tilevich:2008:NNE

[TS08] Eli Tilevich and Yannis Smaragdakis. NRMI: Natural and efficient middleware. IEEE Transactions on Parallel and Distributed Systems, 19(2):174–187, February 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsai:2003:PRC

[Tsa03]  Jichiang Tsai.  On proper-

ties of RDT communicationinduced checkpointing protocols. IEEE Transactions on Parallel and Distributed Systems, 14(8):755–764, August 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/08/l0755abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/08/l0755.
                                        pdf.

Tan:1997:MAE

[TSAL97] Min Tan, H. J. Siegel, J. K. Antonio, and Y. A. Li. Minimizing the application execution time through scheduling of subtasks and communication traffic in a heterogeneous computing system. IEEE Transactions on Parallel and Distributed Systems, 8(8):857–871, August 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0857.
pdf;   http://www.computer.
org/tpds/td1997/l0857abs.
htm.

Tse:2005:AAD

[Tse05] Savio S. H. Tse. Approximate algorithms for document placement in distributed Web servers. IEEE Transactions on Parallel and Distributed Systems, 16 (6):489–496, June 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tse:2009:OBL

[Tse09] Savio S. H. Tse. Online bicriteria load balancing using object reallocation. IEEE Transactions on Parallel and Distributed Systems, 20(3):379–388, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2009:IPD

[TSG09] Guangming Tan, Ninghui Sun, and Guang R. Gao. Improving performance of dynamic programming via parallelism and locality on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 20(2): 261–274, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tu:2007:WCD

[TSJ07] Wanqing Tu, Cormac J. Sreenan, and Weijia Ji. Worst-case delay control in multigroup overlay networks. IEEE Transactions on Parallel and Distributed Systems, 18(10):1407–1419, October 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tekkalmaz:2006:DCM

[TSK06] Metin Tekkalmaz, Hasan Sozer, and Ibrahim Korpeoglu. Distributed construction and maintenance of bandwidth and energy efficient Bluetooth scatternets. IEEE Transactions on Parallel and Distributed Systems, 17(9):963–974, September 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tenllado:2008:PID

[TSP^{+}08] Christian Tenllado, Javier Setoain, Manuel Prieto, Luis Piñuel, and Francisco Tirado. Parallel implementation of the 2D discrete wavelet transform on graphics processing units: Filter bank versus lifting. IEEE Transactions on Parallel and Distributed Systems, 19(3):299– 310, March 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsai:2007:OWC

[TSRS07] Chang-Hao Tsai, Kang G. Shin, John Reumann, and Sharad Singhal. Online Web cluster capacity estimation and its application to energy conservation. IEEE Transactions on Parallel and Distributed Systems, 18 (7):932–945, July 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Triantafillou:1994:MRD

[TT94] Peter Triantafillou and David J.

Taylor. Multiclass replicated data management: exploiting replication to improve efficiency. IEEE Transactions on Parallel and Distributed Systems, 5(2): 121–138, February 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tseng:2001:TDP

[TT01] Y.-C. Tseng and C.-C. Tan.

Termination detection protocols for mobile distributed systems. IEEE Transactions on Parallel and Distributed Systems, 12(6):558–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0558.
pdf;   http://www.computer.
org/tpds/td2001/l0558abs.
htm.

Theys:2000:MMS

[TTB^{+}00] M. D. Theys, M. Tan, N. B.

Beck, H. J. Siegel, and M. Jurczyk. A mathematical model and scheduling heuristics for satisfying prioritized data requests in an oversubscribed communication network. IEEE Transactions on Parallel and Distributed Systems, 11(9):969–??, September 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2000/pdf/l0969.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0969abs.
                                        htm.

Tsuei:1992:MBD

[TV92] T.-F. Tsuei and M. K. Vernon. A multiprocessor bus design model validated by system measurement. IEEE Transactions on Parallel and Distributed Systems, 3(6):712–727, November 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tasi:1998:FAR

[TW98] M.-J. Tasi and S.-D. Wang. A fully adaptive routing algorithm for dynamically injured hypercubes, meshes, and tori. IEEE Transactions on Parallel and Distributed Systems, 9(2):163–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0163.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0163abs.
                                        htm.

Tsai:2000:ADF

[TW00] M.-J. Tsai and S.-D. Wang. Adaptive and deadlock-free routing for irregular faulty patterns in mesh multicomputers. IEEE Transactions on Parallel and Distributed Systems, 11(1):50–??, January 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0050.
pdf;   http://www.computer.
org/tpds/td2000/l0050abs.
htm.

Tseng:1999:EBW

[TWH99] Y.-C. Tseng, S.-Y. Wang, and C.-W. Ho. Efficient broadcasting in wormhole-routed multicomputers: a network-partitioning approach. IEEE Transactions on Parallel and Distributed Systems, 10(1):44–??, January 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0044.
pdf;   http://www.computer.
org/tpds/td1999/l0044abs.
htm.

Tang:2005:QAR

[TX05] Xueyan Tang and Jianliang Xu. QoS-aware replica placement for content distribution. IEEE Transactions on Parallel and Distributed Systems, 16(10): 921–932, October 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2008:ADC

[TX08] Xueyan Tang and Jianliang Xu.

Adaptive data collection strategies for lifetime-constrained wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(6):721– 734, June 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2008:ATB

[TXL08] Xueyan Tang, Jianliang Xu, and Wang-Chien Lee. Analysis of TTL-based consistency in unstructured peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 19(12):1683–1694, December 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsuchiya:1999:MMD

[TYK99] T. Tsuchiya, M. Yamaguchi, and T. Kikuno. Minimizing the maximum delay for reaching consensus in quorum-based mutual exclusion schemes. IEEE Transactions on Parallel and Distributed Systems, 10(4):337– ??, April 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0337.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0337abs.
                                        htm.

Torrellas:1997:PCM

[TZ97] Josep Torrellas and Zheng Zhang. The performance of the cedar multistage switching network. IEEE Transactions on Parallel and Distributed Systems, 8(4):321–336, April 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0321.
pdf;   http://www.computer.
org/tpds/td1997/l0321abs.
htm.

Tzeng:1993:RAF

[Tze93] Nian-Feng Tzeng. Reconfiguration and analysis of a faulttolerant circular butterfly parallel system. IEEE Transactions on Parallel and Distributed Systems, 4(8):855–863, August 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tzeng:2004:MBS

[Tze04] Nian-Feng Tzeng. Multistagebased switching fabrics for scalable routers. IEEE Transactions on Parallel and Distributed Systems, 15(4):304–318, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0304abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0304.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0304.
pdf.

Tzeng:2006:RTP

[Tze06] Nian-Feng Tzeng. Routing table partitioning for speedy packet lookups in scalable routers. IEEE Transactions on Parallel and Distributed Systems, 17 (5):481–494, May 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Umasankar:1995:GAS

[UEA95] Maheshwar Umasankar and Ahmed El-Amawy. Generalized algorithms for systematic synthesis of branch-and-combine clock networks for meshes, tori, and hypercubes. IEEE Transactions on Parallel and Distributed Systems, 6(12):1283–1300, December 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Upfal:1996:RRS

[UFS96] Eli Upfal, Sergio Felperin, and Marc Snir. Randomized routing with shorter paths. IEEE Transactions on Parallel and Distributed Systems, 7(4):356– 362, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1996/l0356abs.
                                        htm.

Uht:1992:ROE

[Uht92] Augustus K. Uht. Requirements for optimal execution of loops with tests. IEEE Transactions on Parallel and Distributed Systems, 3(5):573–581, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Umemoto:1998:SSR

[UKY98] N. Umemoto, H. Kakugawa, and M. Yamashita. A selfstabilizing ring orientation algorithm with a smaller number of processor states. IEEE Transactions on Parallel and Distributed Systems, 9(6):579– ??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0579.
pdf;   http://www.computer.
org/tpds/td1998/l0579abs.
htm.

Urgaonkar:2004:SMC

[US04] Bhuvan Urgaonkar and Prashant Shenoy. Sharc: Managing CPU and network bandwidth in shared clusters. IEEE Transactions on Parallel and Distributed Systems, 15(1):2–17, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/01/l0002abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/01/l0002.
pdf.

Ujaldon:1997:VFH

[UZCZ97] M. Ujaldon, E. L. Zapata, B. M. Chapman, and H. P. Zima. Vienna-Fortran/HPF extensions for sparse and irregular problems and their compilation. IEEE Transactions on Parallel and Distributed Systems, 8(10):1068–1083, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1068.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1068abs.
                                        htm.

Vaidya:1999:SCC

[Vai99] N. H. Vaidya. Staggered consistent checkpointing. IEEE Transactions on Parallel and Distributed Systems, 10(7):694– ??, July 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0694.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0694abs.
                                        htm.

Varma:1993:PET

[Var93] Subir Varma. Performance evaluation of the time-stamp ordering algorithm in a distributed database. IEEE Transactions on Parallel and Distributed Systems, 4(6):668–676, June 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Varki:2001:RTA

[Var01] Elizabeth Varki. Response time analysis of parallel computer and storage systems. IEEE Transactions on Parallel and Distributed Systems, 12(11):1146– 1161, November 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td2001/pdf/l1146.

pdf;   http://www.computer.
org/tpds/td2001/l1146abs.
htm.

Varvarigos:1993:MBH

[VB93] Emmanouel A. Varvarigos and Dimitri P. Bertsekas. Multinode broadcast in hypercubes and rings with randomly distributed length of packets. IEEE Transactions on Parallel and Distributed Systems, 4(2):144–154, February 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Varvarigos:1995:DBP

[VB95] Emmanouel A. Varvarigos and Dimitri P. Bertsekas. Dynamic broadcasting in parallel computing. IEEE Transactions on Parallel and Distributed Systems, 6(2):120–131, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0120abs.htm.

Varvarigos:1996:RSM

[VB96] Emmanouel A. Varvarigos and Ayan Banerjee. Routing schemes for multiple random broadcasts in arbitrary network topologies. IEEE Transactions on Parallel and Distributed Systems, 7(8): 886–895, August 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0886abs.htm.

vanderMerwe:2007:FDP

[vdMDM07] Johann van der Merwe, Dawoud S. Dawoud, and Stephen McDonald. A fully distributed proactively secure thresholdmultisignature scheme. IEEE Transactions on Parallel and Distributed Systems, 18(4):562– 575, April 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vaidya:1999:TEF

[VDS99] A. S. Vaidya, C. R. Das, and A. Sivasubramaniam. A testbed for evaluation of fault-tolerant routing in multiprocessor interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 10(10):52– ??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1052.
pdf;   http://www.computer.
org/tpds/td1999/l1052abs.
htm.

vanReeuwijk:1996:IFH

[vDSP96] Kees van Reeuwijk, Will Denissen, Henk J. Sips, and Edwin M. R. M. Paalvast. An implementation framework for HPF distributed arrays on messagepassing parallel computer systems. IEEE Transactions on Parallel and Distributed Systems, 7(9):897–914, September 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0897abs.htm.

vanGemund:2003:SPM

[vG03] Arjan J. C. van Gemund. Symbolic performance modeling of parallel systems. IEEE Transactions on Parallel and Distributed Systems, 14(2):154–165, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/02/l0154abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0154.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0154.
pdf.

Vallejo:1994:SMM

[VGGD94] F. Vallejo, J. A. Gregorio, M. Gonzalez Harbour, and J. M. Drake. Shared memory multimicroprocessor operating system with an extended Petri net model. IEEE Transactions on Parallel and Distributed Systems, 5(7):749–762, July 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vijaykumar:2001:SVC

[VGSS01] T. N. Vijaykumar, S. Gopal, J. E. Smith, and G. Sohi. Speculative versioning cache. IEEE Transactions on Parallel and Distributed Systems, 12(12): 1305–1317, December 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.

org/td/books/td2001/pdf/l1305.
pdf;   http://www.computer.
org/tpds/td2001/l1305abs.
htm.

Vinnakota:1993:SAB

[VJ93] Bapiraju Vinnakota and Niraj K. Jha. Synthesis of algorithm-based fault-tolerant systems from dependence graphs. IEEE Transactions on Parallel and Distributed Systems, 4 (8):864–874, August 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vinnakota:1994:DAB

[VJ94] Bapiraju Vinnakota and Niraj K. Jha. Design of algorithmbased fault-tolerant multiprocessor systems for concurrent error detection and fault diagnosis. IEEE Transactions on Parallel and Distributed Systems, 5(10): 1099–1106, October 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Venkatesan:1997:OCR

[VJA97] S. Venkatesan, Tony Tong-Ying Juang, and Sridhar Alagar. Optimistic crash recovery without changing application messages. IEEE Transactions on Parallel and Distributed Systems, 8(3):263–271, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0263.
pdf;   http://www.computer.
org/tpds/td1997/l0263abs.
htm.

Vydyanathan:2009:IAL

[VKS^{+}09] Naga Vydyanathan, Sriram Krishnamoorthy, Gerald M. Sabin, Umit V. Catalyurek, Tahsin Kurc, Ponnuswamy Sadayappan, and Joel H. Saltz. An integrated approach to localityconscious processor allocation and scheduling of mixed-parallel applications. IEEE Transactions on Parallel and Distributed Systems, 20(8):1158–1172, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Varavithya:1999:ATB

[VM99] V. Varavithya and P. Mohapatra. Asynchronous treebased multicasting in wormholeswitched MINs. IEEE Transactions on Parallel and Distributed Systems, 10(11):159–??, November 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1159.
pdf;   http://www.computer.
org/tpds/td1999/l1159abs.
htm.

Veeravalli:2004:SDL

[VM04] Bharadwaj Veeravalli and Wong Han Min. Scheduling divisible loads on heterogeneous linear daisy chain networks with arbitrary processor release times. IEEE Transactions on Parallel and Distributed Systems, 15(3):273–288, March 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/03/l0273abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/03/l0273.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/03/l0273.
                                        pdf.

Varki:2004:ICP

[VMXQ04] Elizabeth Varki, Arif Merchant, Jianzhang Xu, and Xiaozhou Qiu. Issues and challenges in the performance analysis of real disk arrays. IEEE Transactions on Parallel and Distributed Systems, 15(6):559–574, June 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/06/l0559.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/06/l0559.pdf.

Villela:2005:PAS

[VR05] Daniel Villela and Dan Rubenstein. Performance analysis of server sharing collectives for content distribution. IEEE Transactions on Parallel and Distributed Systems, 16(12):1178– 1189, December 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Varvarigou:1996:SFP

[VRKL96] Theodora A. Varvarigou, Vwani P. Roychowdhury, Thomas Kailath, and Eugene Lawler. Scheduling in and out forests in the presence of communication delays. IEEE Transactions on Parallel and Distributed Systems, 7 (10):1065–1074, October 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l1065abs.htm.

Vadapalli:1996:NFC

[VS96] Premkumar Vadapalli and Pradip K.

Srimani. A new family of Cayley graph interconnection networks of constant degree four. IEEE Transactions on Parallel and Distributed Systems, 7 (1):26–32, January 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See comments [CL97].

Verbeek:2011:CNS

[VS11] Freek Verbeek and Julien Schmaltz. A comment on “A Necessary and Sufficient Condition for Deadlock-Free Adaptive Routing in Wormhole Networks”. IEEE Transactions on Parallel and Distributed Systems, 22(10):1775–1776, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic). See [Dua93, Dua95a].

Vaidya:2001:IVC

[VSD01] A. S. Vaidya, A. Sivasubramaniam, and C. R. Das. Impact of virtual channels and adaptive routing on application performance. IEEE Transactions on Parallel and Distributed Systems, 12(2):223–??, February 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0223.
pdf;   http://www.computer.
org/tpds/td2001/l0223abs.
htm.

Varman:1999:TBP

[VV99] P. J. Varman and R. M. Verma. Tight bounds for prefetching and buffer management algorithms for parallel I/O systems. IEEE Transactions on Parallel and Distributed Systems, 10(12):262–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1262.
pdf;   http://www.computer.
org/tpds/td1999/l1262abs.
htm.

Viswanathan:2007:RAD

[VVR07] Sivakumar Viswanathan, Bharadwaj Veeravalli, and Thomas G. Robertazzi. Resource-aware distributed scheduling strategies for large-scale computational cluster/Grid systems. IEEE Transactions on Parallel and Distributed Systems, 18(10):1450– 1461, October 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2007/10/
l1450s.pdf.

Wu:1999:GEI

[WA99] J. Wu and D. P. Agrawal. Guest Editors’ introduction: Challenges in designing fault-tolerant routing in networks. IEEE Transactions on Parallel and Distributed Systems, 10(10):961– ??, October 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0961.
pdf.

Wang:1998:FTC

[Wan98] B.-F. Wang. Finding a k-tree core and a k-tree center of a tree network in parallel. IEEE Transactions on Parallel and Distributed Systems, 9(2):186–??, February 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0186.
pdf;   http://www.computer.
org/tpds/td1998/l0186abs.
htm.

Wang:2004:FDS

[Wan04] Biing-Feng Wang. Finding rdominating sets and p-centers of trees in parallel. IEEE Transactions on Parallel and Distributed Systems, 15(8):687– 398, August 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/08/l0687.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/08/l0687.pdf.

Wang:2008:EHC

[Wan08] Dajin Wang. On embedding Hamiltonian cycles in crossed cubes. IEEE Transactions on Parallel and Distributed Systems, 19(3):334–346, March 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wallace:1998:MMI

[WB98] S. Wallace and N. Bagherzadeh. Modeled and measured instruction fetching performance for superscalar microprocessors. IEEE Transactions on Parallel and Distributed Systems, 9(6):570– ??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0570.
pdf;   http://www.computer.
org/tpds/td1998/l0570abs.
htm.

Wolski:2001:WPR

[WBO^{+}01] Rich Wolski, John Brevik, Graziano Obertelli, Neil Spring, and Alan Su. Writing programs that run EveryWare on the computational grid. IEEE Transactions on Parallel and Distributed Systems, 12(10):1066–1080, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1066.
pdf;   http://www.computer.
org/tpds/td2001/l1066abs.
htm.

Wang:1990:CTA

[WC90] Biing-Feng Wang and Gen-Huey Chen. Constant time algorithms for the transitive closure and some related graph problems on processor arrays with reconfigurable bus systems. IEEE Transactions on Parallel and Distributed Systems, 1(4): 500–507, October 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Walters:2009:RBF

[WC09] John Paul Walters and Vipin Chaudhary. Replication-based fault tolerance for MPI applications. IEEE Transactions on Parallel and Distributed Systems, 20(7):997–1010, July 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2006:ESO

[WCBX06] Xun Wang, Sriram Chellappan, Phillip Boyer, and Dong Xuan. On the effectiveness of secure overlay forwarding systems under intelligent distributed DoS attacks. IEEE Transactions on Parallel and Distributed Systems, 17(7):619–632, July 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wills:1997:HTL

[WCCR^{+}97] D. S. Wills, H. H. Cat, J. Cruz-Rivera, W. S. Lacy, J. M. Baker, Jr., J. C. Eble, A. López-Lagunas, and M. Hopper. Highthroughput, low-memory applications on the Pica architecture. IEEE Transactions on Parallel and Distributed Systems, 8(10):1055–1067, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1055.
pdf;   http://www.computer.
org/tpds/td1997/l1055abs.
htm.

Wu:2008:AEH

[WCD08] Xiaobing Wu, Guihai Chen, and Sajal K. Das. Avoiding energy holes in wireless sensor networks with nonuniform node distribution. IEEE Transactions on Parallel and Distributed Systems, 19(5):710–720, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Wu:2006:EDS

[WCDY06] Jie Wu, Mihaela Cardei, Fei Dai, and Shuhui Yang. Extended dominating set and its applications in ad hoc networks using cooperative communication. IEEE Transactions on Parallel and Distributed Systems, 17(8): 851–864, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Witte:1991:PSA

[WCF91] Ellen E. Witte, Roger D. Chamberlain, and Mark A. Franklin. Parallel simulated annealing using speculative computation. IEEE Transactions on Parallel and Distributed Systems, 2(4): 483–494, October 1991. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2008:CBA

[WCH^{+}08] Xiaorui Wang, Ming Chen, Huang-Ming Huang, Venkita Subramonian, Chenyang Lu, and Christopher D. Gill. Controlbased adaptive middleware for real-time image transmission over bandwidth-constrained networks. IEEE Transactions on Parallel and Distributed Systems, 19(6):779–793, June 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:1997:PAC

[WCL97] Yue-Li Wang, Hon-Chan Chen, and Wei-Kai Liu. A parallel algorithm for constructing a labeled tree. IEEE Transactions on Parallel and Distributed Systems, 8(12):1236–1240, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1236.
pdf;   http://www.computer.
org/tpds/td1997/l1236abs.
htm.

Wang:1995:CSR

[WCLF95] Yi-Min Wang, Pi-Yu Chung, In-Jen Lin, and W. Kent Fuchs. Checkpoint space reclamation for uncoordinated checkpointing in message-passing systems. IEEE Transactions on Parallel and Distributed Systems, 6 (5):546–554, May 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.

                                        computer.org/tpds/td1995/
                                        l0546abs.htm.

Wu:2009:ECM

[WCR09] Weigang Wu, Jiannong Cao, and Michel Raynal. Eventual clusterer: a modular approach to designing hierarchical consensus protocols in MANETs. IEEE Transactions on Parallel and Distributed Systems, 20(6):753– 765, June 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:1992:PEM

[WCS92] C. Wang, A. L. P. Chen, and S.-C. Shyu. A parallel execution method for minimizing distributed query response time. IEEE Transactions on Parallel and Distributed Systems, 3 (3):325–333, May 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:1995:BAG

[WCY95] S. C. Wang, Y. H. Chin, and K. Q. Yan. Byzantine agreement in a generalized connected network. IEEE Transactions on Parallel and Distributed Systems, 6(4):420–427, April 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0420abs.htm.

Wu:2006:MST

[WD06] Jie Wu and Fei Dai. Mobilitysensitive topology control in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 17(6):522– 535, June 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2004:SAC

[WDCK04] Zhijun Wang, Sajal K. Das, Hao Che, and Mohan Kumar. A scalable asynchronous cache consistency scheme (SACCS) for mobile environments. IEEE Transactions on Parallel and Distributed Systems, 15(11):983– 995, November 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/11/
l0983abs.htm;   http://csdl.
computer.org/dl/trans/td/
2004/11/l0983.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/11/l0983.pdf.

Wolf:1993:PSM

[WDY93] Joel L. Wolf, Daniel M. Dias, and Philip S. Yu. Parallel sort merge join algorithm for managing data skew. IEEE Transactions on Parallel and Distributed Systems, 4(1):70–86, January 1993. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wong:1998:SAA

[WDY98] D. C. Wong, E. W. Davis, and J. O. Young. A software approach to avoiding spatial cache collisions in parallel processor systems. IEEE Transactions on Parallel and Distributed Systems, 9(6):601– ??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0601.
pdf;   http://www.computer.
org/tpds/td1998/l0601abs.
htm.

Wen:1996:MMP

[Wen96] Zhaofang Wen. Multiway merging in parallel. IEEE Transactions on Parallel and Distributed Systems, 7(1):11–17, January 1996. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:1994:UPN

[WF94] Jie Wu and Eduardo B. Fernandez. Using Petri nets for the design of conversation boundaries in fault-tolerant software. IEEE Transactions on Parallel and Distributed Systems, 5(10): 1106–1112, October 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wiseman:2003:PGS

[WF03] Yair Wiseman and Dror G. Feitelson. Paired gang scheduling. IEEE Transactions on Parallel and Distributed Systems, 14(6):581–592, June 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/06/l0581abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/06/l0581.
pdf.

Wolf:2006:PMN

[WF06] Tilman Wolf and Mark A.

Franklin. Performance models for network processor design. IEEE Transactions on Parallel and Distributed Systems, 17 (6):548–561, June 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:1990:ERS

[WFP90] Kun-Lung Wu, W. Kent Fuchs, and Janak H. Patel. Error recovery in shared memory multiprocessors using private caches. IEEE Transactions on Parallel and Distributed Systems, 1 (2):231–240, April 1990. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wen:2009:DBA

[WFS09] Shushan Wen, Yuguang (Michael)

Fang, and Hairong Sun. Differentiated bandwidth allocation with TCP protection in core routers. IEEE Transactions on Parallel and Distributed Systems, 20(1):34–47, January 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:1990:HPA

[WG90] Min-You Wu and Daniel D.

Gajski. Hypertool: a programming aid for message-passing systems. IEEE Transactions on Parallel and Distributed Systems, 1(3):330–343, July 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:1995:MGS

[WH95] Hwang-Cheng Wang and Kai Hwang. Multicoloring of gridstructured PDE solvers on shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6 (11):1195–1205, November 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l1195abs.htm.

Wolfson:1998:CAC

[WH98] O. Wolfson and Y. Huang. Competitive analysis of caching in distributed databases. IEEE Transactions on Parallel and Distributed Systems, 9(4):391– ??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0391.
pdf;   http://www.computer.
org/tpds/td1998/l0391abs.
htm.

Wu:2001:VMF

[WH01] C.-H. Wu and S.-J. Horng. L_{2} vector median filters on arrays with reconfigurable optical buses. IEEE Transactions on Parallel and Distributed Systems, 12(12):1281–1292, December 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l1281.
pdf;   http://www.computer.

org/tpds/td2001/l1281abs.
htm.

Wang:2003:TAE

[WH03a] Yuh-Rau Wang and Shi-Jinn Horng. An O(1) time algorithm for the 3D Euclidean distance transform on the CRCW PRAM model. IEEE Transactions on Parallel and Distributed Systems, 14(10):973–982, October 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/10/l0973abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/10/l0973.
pdf.

Wu:2003:FSS

[WH03b] Chin-Hsiung Wu and Shi-Jinn Horng. Fast and scalable selection algorithms with applications to median filtering. IEEE Transactions on Parallel and Distributed Systems, 14(10):983–992, October 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/10/l0983abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/10/l0983.
pdf.

Wu:2003:PAM

[WHC03] Fan Wu, Ching-Chi Hsu, and Li-Ping Chou. Processor allocation in the mesh multiprocessors using the leapfrog method. IEEE Transactions on Parallel and Distributed Systems, 14(3):276–289, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/03/l0276abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0276.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0276.
pdf.

Wong:1995:PAC

[WHL95] Yung-Chang Wong, Shu-Yuen Hwang, and Jason Yi-Bing Lin. A parallelism analyzer for conservative parallel simulation. IEEE Transactions on Parallel and Distributed Systems, 6 (6):628–638, June 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0628abs.htm.

Wang:2009:UPD

[WHM09] Feng Wang, Mounir Hamdi, and Jogesh K. Muppala. Using parallel DRAM to scale router buffers. IEEE Transactions on Parallel and Distributed Systems, 20(5):710–724, May 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2005:EAA

[WHW05] Yuh-Rau Wang, Shi-Jinn Horng, and Chin-Hsiung Wu. Efficient algorithms for the all nearest neighbor and closest pair problems on the linear array with a reconfigurable pipelined bus system. IEEE Transactions on Parallel and Distributed Systems, 16 (3):193–206, March 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2007:DDE

[WJLK07] Xiaorui Wang, Dong Jia, Chenyang Lu, and Xenofon Koutsoukos. DEUCON: Decentralized end-to-end utilization control for distributed real-time systems. IEEE Transactions on Parallel and Distributed Systems, 18(7):996–1009, July 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jang:1997:CTA

[wJNPS97] Ju wook Jang, Madhusudan Nigam, Viktor K. Prasanna, and Sartaj Sahni. Constant time algorithms for computational geometry on the reconfigurable mesh. IEEE Transactions on Parallel and Distributed Systems, 8(1):1–12, January 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jang:1997:OMA

[wJPP97] Ju wook Jang, Heonchul Park, and Viktor K. Prasanna. An optimal multiplication algorithm on reconfigurable mesh. IEEE Transactions on Parallel and Distributed Systems, 8(5):521– 532, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0521.
pdf;   http://www.computer.
org/tpds/td1997/l0521abs.
htm.

Wang:2001:COP

[WKS01] Biing-Feng Wang, Shan-Chyun Ku, and Keng-Hua Shii. Costoptimal parallel algorithms for the tree bisector and related problems. IEEE Transactions on Parallel and Distributed Systems, 12(9):888–898, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0888.
pdf;   http://www.computer.
org/tpds/td2001/l0888abs.
htm.

Wolf:1991:LTT

[WL91] Michael E. Wolf and Monica S. Lam. Loop transformation theory and an algorithm to maximize parallelism. IEEE Transactions on Parallel and Distributed Systems, 2(4):452–471, October 1991. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:1997:RMM

[WL97] Feng-Hsu Wang and Ferng-Ching Lin. On routing maskable messages in hypercubederived multistage interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 8(3):285–291, March 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0285.
pdf;   http://www.computer.
org/tpds/td1997/l0285abs.
htm.

Webb:2000:CHS

[WL00] B. Webb and A. Louri. A class of highly scalable optical crossbar-connected interconnection networks (SOCNs) for parallel computing systems. IEEE Transactions on Parallel and Distributed Systems, 11(5):444– ??, May 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0444.
pdf;   http://www.computer.
org/tpds/td2000/l0444abs.
htm.

Wang:2008:DSL

[WL08a] Dan Wang and Jiangchuan Liu.

A dynamic skip list-based overlay for on-demand media streaming with VCR interactions. IEEE Transactions on Parallel and Distributed Systems, 19 (4):503–514, April 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2008:RRO

[WL08b] Chuan Wu and Baochun Li.

rStream: Resilient and optimal peer-to-peer streaming with rateless codes. IEEE Transactions on Parallel and Distributed Systems, 19(1):77–92, January 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2008:AGF

[WLHB08] Xiaoxin Wu, Jun Liu, Xiaoyan Hong, and Elisa Bertino. Anonymous geo-forwarding in MANETs through location cloaking. IEEE Transactions on Parallel and Distributed Systems, 19(10):1297–1309, October 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2007:LRR

[WLL^{+}07] Chonggang Wang, Jiangchuan Liu, Bo Li, Kazem Sohraby, and Y. Thomas Hou. LRED: a robust and responsive AQM algorithm using packet loss ratio measurement. IEEE Transactions on Parallel and Distributed Systems, 18(1):29–43, January 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2008:DBA

[WLL08] Chuan Wu, Baochun Li, and Zongpeng Li. Dynamic bandwidth auctions in multioverlay P2P streaming with network coding. IEEE Transactions on Parallel and Distributed Systems, 19(6):806–820, June 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Willebeek-LeMair:1993:SDL

[WLR93] Marc H. Willebeek-LeMair and Anthony P. Reeves. Strategies for dynamic load balancing on highly parallel computers. IEEE Transactions on Parallel and Distributed Systems, 4(9): 979–993, September 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2008:EAS

[WLZ08] Yu Wang, Xiang-Yang Li, and Qian Zhang. Efficient algorithms for p-self-protection problem in static wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(10): 1426–1438, October 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2007:VRE

[WLZN07] Hejun Wu, Qiong Luo, Pei Zheng, and Lionel M. Ni. VM-Net: Realistic emulation of wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(2):277– 288, February 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Woodside:1993:FAP

[WM93] C. Murray Woodside and Gerald G. Monforton. Fast allocation of processes in distributed and parallel systems. IEEE Transactions on Parallel and Distributed Systems, 4(2): 164–174, February 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:1995:SEA

[WM95] Yuguang Wu and Richard Muntz. Stack evaluation of arbitrary set-associative multiprocessor caches. IEEE Transactions on Parallel and Distributed Systems, 6(9):930–942, September 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0930abs.
htm.

Wu:1996:DAP

[WMB96] Shiow-Yang Wu, Daniel P. Miranker, and James C. Browne. Decomposition abstraction in parallel rule languages. IEEE Transactions on Parallel and Distributed Systems, 7(11): 1164–1184, November 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//neumann.computer.org/td/
books/td1996/pdf/l1164.pdf;
 http://www.computer.org/
tpds/td1996/l1164abs.htm.

Wei:1999:IDF

[WMN99] D. S. L. Wei, F. P. Muga, and K. Naik. Isomorphism of degree four Cayley graph and wrapped butterfly and their optimal permutation routing algorithm. IEEE Transactions on Parallel and Distributed Systems, 10(12):290–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/

td/books/td1999/pdf/l1290.
pdf;   http://www.computer.
org/tpds/td1999/l1290abs.
htm.

Wei:2008:FPA

[WMWL08] Tongquan Wei, Piyush Mishra, Kaijie Wu, and Han Liang. Fixed-priority allocation and scheduling for energy-efficient fault tolerance in hard realtime multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1511–1526, November 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2006:DIQ

[WMXZ06] Shengquan Wang, Zhibin Mai, Dong Xuan, and Wei Zhao. Design and implementation of QoS-provisioning system for voice over IP. IEEE Transactions on Parallel and Distributed Systems, 17(3):276– 288, March 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://bell.computer.
org/dlcomments/.

Wang:1996:CPC

[WNKS96] Haigeng Wang, Alexandru Nicolau, Stephen Keung, and Kai-Yeung (Sunny) Siu. Computing programs containing band linear recurrences on vector supercomputers. IEEE Transactions on Parallel and Distributed Systems, 7(8):769–782 (or 769–781??), August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0769abs.htm.

Wang:2004:TZH

[WO04] Lan Wang and Stephan Olariu. A two-zone hybrid routing protocol for mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 15(12):1105–1116, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2004/12/l1105abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/12/l1105.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2004/12/l1105.
                                        pdf.

Watanabe:2007:MNI

[WOT^{+}07] Konosuke Watanabe, Tomohiro Otsuka, Junichiro Tsuchiya, Hiroaki Nishi, Junji Yamamoto, Noboru Tanabe, Tomohiro Kudoh, and Hideharu Amano. Martini: a network interface controller chip for high performance computing with distributed PCs. IEEE Transactions on Parallel and Distributed Systems, 18 (9):1282–1295, September 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Warnakulasuriya:2000:FMM

[WP00] S. Warnakulasuriya and T. M. Pinkston. A formal model of message blocking and deadlock resolution in interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 11(3):212–??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0212.
pdf;   http://www.computer.
org/tpds/td2000/l0212abs.
htm.

Wang:2004:TUF

[WR04] Jinggang Wang and Binoy Ravindran. Time-utility functiondriven switched Ethernet: Packet scheduling algorithm, implementation, and feasibility analysis. IEEE Transactions on Parallel and Distributed Systems, 15(2):119–133, February 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/02/l0119abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/02/l0119.
pdf.

Wei:2009:CSA

[WRB09] Qiang Wei, Matei Ripeanu, and Konstantin Beznosov. Cooperative secondary authorization recycling. IEEE Transactions on Parallel and Distributed Systems, 20(2):275–288, February 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Woo:1993:PHP

[WS93] Tai-Kuo Woo and Stanley Y. W.

Su. PCBN: a high-performance partitionable circular bus network for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 4(12): 1298–1307, December 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:1998:BOO

[WS98] C.-F. Wang and S. Sahni. Basic operations on the OTISmesh optoelectronic computer. IEEE Transactions on Parallel and Distributed Systems, 9(12): 226–??, December 1998. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td1998/pdf/l1226.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l1226abs.
                                        htm.

Wang:2000:IPO

[WS00] C.-F. Wang and S. Sahni. Image processing on the OTISmesh optoelectronic computer. IEEE Transactions on Parallel and Distributed Systems, 11 (2):97–??, February 2000. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.
                                        org/td/books/td2000/pdf/l0097.
                                        pdf;   http://www.computer.
                                        org/tpds/td2000/l0097abs.
                                        htm.

Wang:2003:TAI

[WS03] Haining Wang and Kang G. Shin. Transport-aware IP routers: a built-in protection mechanism to counter DDoS attacks. IEEE Transactions on Parallel and Distributed Systems, 14(9):873–884, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0873abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0873.
pdf.

Wang:2009:OIS

[WSB09] Hui Wang, Rama Sangireddy, and Sandeep Baldawa. Optimizing instruction scheduling through combined in-order and O-O-O execution in SMT processors. IEEE Transactions on Parallel and Distributed Systems, 20(3):389–403, March 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wagner:1997:PMP

[WSC97] Alan S. Wagner, Halsur V.

Sreekantaswamy, and Samuel T. Chanson. Performance models for the processor farm paradigm. IEEE Transactions on Parallel and Distributed Systems, 8(5):475–489, May 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0475.
pdf;   http://www.computer.
org/tpds/td1997/l0475abs.
htm.

Wu:2001:ELS

[WSG01] M.-Y. Wu, W. Shu, and J. Gu. Efficient local search for DAG scheduling. IEEE Transactions on Parallel and Distributed Systems, 12(6):617–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0617.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0617abs.
                                        htm.

Wang:1995:UMN

[WSNA95] Mu-Cheng Wang, Howard Jay Siegel, Mark A. Nichols, and Seth Abraham. Using a multipath network for reducing the effects of hot spots. IEEE Transactions on Parallel and Distributed Systems, 6(3):252– 268, March 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://www.computer.
                                        org/tpds/td1995/l0252abs.
                                        htm.

Wolfe:1992:PTD

[WT92] Michael Wolfe and Chau-Wen Tseng. The power test for data dependence. IEEE Transactions on Parallel and Distributed Systems, 3(5):591–601, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Watts:1998:PAD

[WT98] J. Watts and S. Taylor. A practical approach to dynamic load balancing. IEEE Transactions on Parallel and Distributed Systems, 9(3):235–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0235.
pdf;   http://www.computer.
org/tpds/td1998/l0235abs.
htm.

Wang:2008:DDS

[WT08] You-Chiun Wang and Yu-Chee Tseng. Distributed deployment schemes for mobile wireless sensor networks to ensure multilevel coverage. IEEE Transactions on Parallel and Distributed Systems, 19(9):1280–1294, September 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wolf:1995:HAP

[WTCY95] Joel L. Wolf, John Turek, Ming-Syan Chen, and Philip S. Yu. A hierarchical approach to parallel multiquery scheduling. IEEE Transactions on Parallel and Distributed Systems, 6 (6):578–590, June 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0578abs.htm.

Wu:1997:EFC

[Wu97a] Jie Wu. Extended Fibonacci cubes. IEEE Transactions on Parallel and Distributed Systems, 8(12):1203–1210, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1203.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1203abs.
                                        htm.

Wu:1997:RPS

[Wu97b] Min-You Wu. On runtime parallel scheduling for processor load balancing. IEEE Transactions on Parallel and Distributed Systems, 8(2):173–186, February 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0173.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0173abs.
                                        htm.

Wu:1998:AFT

[Wu98]  J. Wu.  Adaptive fault-

tolerant routing in cube-based multicomputers using safety vectors. IEEE Transactions on Parallel and Distributed Systems, 9(4):321–??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0321.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0321abs.
                                        htm.

Wu:2000:FTA

[Wu00] J. Wu. Fault-tolerant adaptive and minimal routing in mesh-connected multicomputers using extended safety levels. IEEE Transactions on Parallel and Distributed Systems, 11(2):149–??, February 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0149.
pdf;   http://www.computer.
org/tpds/td2000/l0149abs.
htm.

Wu:2002:EDS

[Wu02] Jie Wu. Extended dominatingset-based routing in ad hoc wireless networks with unidirectional links. IEEE Transactions on Parallel and Distributed Systems, 13(9):866–881, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0866.
pdf;   http://www.computer.
org/tpds/td2002/l0866abs.
htm.

Wang:1992:EPA

[WW92] Chien-Min Wang and Sheng-De Wang. Efficient processor assignment algorithms and loop transformations for executing nested parallel loops on multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 3 (1):71–82, January 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2006:EDL

[WWL06] Yu Wang, WeiZhao Wang, and Xiang-Yang Li. Efficient distributed low-cost backbone formation for wireless networks. IEEE Transactions on Parallel and Distributed Systems, 17 (7):681–693, July 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2008:IAJ

[WWLS08] Yu Wang, Weizhao Wang, Xiang-Yang Li, and Wen-Zhan Song. Interference-aware joint routing and TDMA link scheduling for static wireless networks. IEEE Transactions on Parallel and Distributed Systems, 19(12): 1709–1726, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2009:RFL

[WWWA09] Yun Wang, Xiaodong Wang, Demin Wang, and Dharma P. Agrawal. Range-free localization using expected hop progress in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(10): 1540–1552, October 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2007:EPS

[WX07] Chen Wang and Li Xiao. An effective P2P search scheme to exploit file sharing heterogeneity. IEEE Transactions on Parallel and Distributed Systems, 18(2): 145–157, February 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2006:DED

[WXLZ06] Chen Wang, Li Xiao, Yunhao Liu, and Pei Zheng. DiCAS: An efficient distributed caching mechanism for P2P systems. IEEE Transactions on Parallel and Distributed Systems, 17(10): 1097–1109, October 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wolfsthal:1994:ETL

[WY94] Yaron Wolfsthal and Michael Yoeli. An equivalence theorem for labeled marked graphs. IEEE Transactions on Parallel and Distributed Systems, 5(8): 886–891, August 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wan:2007:LEG

[WY07] Peng-Jun Wan and Chih-Wei Yi. On the longest edge of Gabriel graphs in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 18(1):111–125, January 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2007:LSC

[WYD07] Jie Wu, Shuhui Yang, and Fei Dai. Logarithmic store-carry-forward routing in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 18(6):735–748, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2004:IGA

[WYJ^{+}04] Annie S. Wu, Han Yu, Shiyuan Jin, Kuo-Chi Lin, and Guy Schiavone. An incremental genetic algorithm approach to multiprocessor scheduling. IEEE Transactions on Parallel and Distributed Systems, 15(9):824– 834, September 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/09/l0824.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/09/l0824.pdf.

Wolf:1993:PHJ

[WYTD93] Joel L. Wolf, Philip S. Yu, John Turek, and Daniel M. Dias. A parallel hash join algorithm for managing data skew. IEEE Transactions on Parallel and Distributed Systems, 4(12): 1355–1371, December 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2008:RAA

[WYWZ08] S.-C. Wang, K.-Q. Yan, S.-S. Wang, and G.-Y. Zheng. Reaching agreement among virtual subnets in hybrid failure mode. IEEE Transactions on Parallel and Distributed Systems, 19 (9):1252–1262, September 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2003:PBR

[WZP^{+}03] Lan Wang, Xiaoliang Zhao, Dan Pei, Randy Bush, Daniel Massey, and Lixia Zhang. Protecting BGP routes to top-level DNS servers. IEEE Transactions on Parallel and Distributed Systems, 14(9):851–860, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0851abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0851.
pdf.

Wang:2009:DMS

[WZZ09] C. Wang, B. B. Zhou, and Albert Y. Zomaya. A decentralized method for scaling up genome similarity search services. IEEE Transactions on Parallel and Distributed Systems, 20(3):303– 315, March 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiong:1993:PMS

[XB93] Renbing Xiong and Theodore Brown. Parallel median splitting and k-splitting with application to merging and sorting. IEEE Transactions on Parallel and Distributed Systems, 4 (5):559–565, May 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:1998:DVD

[XB98] L. Xu and J. Bruck. Deterministic voting in distributed systems using error-correcting codes. IEEE Transactions on Parallel and Distributed Systems, 9(8):813–??, August 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0813.
pdf;   http://www.computer.
org/tpds/td1998/l0813abs.
htm.

Xu:2001:TSA

[XC01] C.-Z. Xu and V. Chaudhary. Time stamp algorithms for runtime parallelization of DOACROSS loops with dynamic dependences. IEEE Transactions on Parallel and Distributed Systems, 12(5):433– ??, May 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0433.
pdf;   http://www.computer.
org/tpds/td2001/l0433abs.
htm.

Xu:2004:PPS

[XC04] Jinsheng Xu and Moon Jung Chung. Predicting the performance of synchronous discrete event simulation. IEEE Transactions on Parallel and Distributed Systems, 15(12): 1130–1137, December 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/12/l1130abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1130.
htm;   http://csdl.computer.
org/dl/trans/td/2004/12/l1130.
pdf.

Xiao:2002:DCR

[XCZ02] L. Xiao, S. Chen, and X. Zhang.

Dynamic cluster resource allocations for jobs with known and unknown memory demands. IEEE Transactions on Parallel and Distributed Systems, 13 (3):223–240, March 2002. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2002/pdf/l0223.
pdf;   http://www.computer.
org/tpds/td2002/l0223abs.
htm.

Xiao:2004:AMA

[XCZ04] Li Xiao, Songqing Chen, and Xiaodong Zhang. Adaptive memory allocations in clusters to handle unexpectedly large data-intensive jobs. IEEE Transactions on Parallel and Distributed Systems, 15(7): 577–592, July 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/07/l0577.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/07/l0577.pdf.

Xiao:2008:DLU

[XCZ08] Bin Xiao, Hekang Chen, and Shuigeng Zhou. Distributed localization using a moving beacon in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19 (5):587–600, May 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:1997:OSM

[XGN97] Hong Xu, Yadong Gui, and Lionel M. Ni. Optimal software multicast in wormhole-routed multistage networks. IEEE Transactions on Parallel and Distributed Systems, 8(6):597– 607, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0597.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0597abs.
                                        htm.

Xu:2008:CEM

[XH08] Lihao Xu and Cheng Huang. Computation-efficient multicast key distribution. IEEE Transactions on Parallel and Distributed Systems, 19(5):577–587, May 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2005:DAC

[XHYL05] Zhonghang Xia, Wei Hao, I-Ling Yen, and Peng Li. A distributed admission control model for QoS assurance in large-scale media delivery systems. IEEE Transactions on Parallel and Distributed Systems, 16(12):1143–1153, December 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiang:2001:FTR

[Xia01]  Dong Xiang.  Fault-tolerant

routing in hypercube multicomputers using local safety information. IEEE Transactions on Parallel and Distributed Systems, 12(9):942–951, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0942.
pdf;   http://www.computer.
org/tpds/td2001/l0942abs.
htm.

Xuan:2000:RPA

[XJZZ00] D. Xuan, W. Jia, W. Zhao, and H. Zhu. A routing protocol for anycast messages. IEEE Transactions on Parallel and Distributed Systems, 11(6):571– ??, June 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0571.
pdf;   http://www.computer.
org/tpds/td2000/l0571abs.
htm.

Xu:1996:ETD

[XL96] Chengzhong Xu and Francis C. M. Lau. Efficient termination detection for loosely synchronous applications in multicomputers. IEEE Transactions on Parallel and Distributed Systems, 7(5):537–544, May 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0537abs.htm.

Xiao:2004:VVT

[XL04] Yang Xiao and Haizhon Li.

Voice and video transmissions with global data parameter control for the IEEE 802.11e enhance distributed channel access. IEEE Transactions on Parallel and Distributed Systems, 15(11):1041–1053, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/11/l1041abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1041.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1041.
pdf.

Xiao:2008:NDD

[XL08] Xiang Xiao and Jaehwan John Lee. A novel O(1) deadlock detection methodology for multiunit resource systems and its hardware implementation for system-on-chip. IEEE Transactions on Parallel and Distributed Systems, 19(12):1657–1670, December 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xing:2006:ISC

[XLPH06] Guoliang Xing, Chenyang Lu, Robert Pless, and Qingfeng Huang. Impact of sensing coverage on greedy geographic routing algorithms. IEEE Transactions on Parallel and Distributed Systems, 17(4):348– 360, April 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/04/
l0348s.pdf.

Xiao:2006:OBC

[XLW^{+}06] Yang Xiao, Frank Haizhon Li, Kui Wu, Kin K. Leung, and Qiang Ni. On optimizing backoff counter reservation and classifying stations for the IEEE 802.11 distributed wireless LANs. IEEE Transactions on Parallel and Distributed Systems, 17(7):713– 722, July 2006. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2005:DQG

[XP05] Yang Xiao and Yi Pan. Differentiation, QoS guarantee, and optimization for real-time traffic over one-hop ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 16 (6):538–549, June 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2007:GCM

[XP07] Wenjun Xiao and Behrooz Parhami. A group construction method with applications to deriving pruned interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 18(5):637–643, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2004:DAL

[XPL04] Yang Xiao, Yi Pan, and Jie Li. Design and analysis of location management for 3G cellular networks. IEEE Transactions on Parallel and Distributed Systems, 15(4):339–349, April 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/04/l0339abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0339.
htm;   http://csdl.computer.
org/dl/trans/td/2004/04/l0339.
pdf.

Xie:2008:SAR

[XQ08] Tao Xie and Xiao Qin. Securityaware resource allocation for real-time parallel jobs on homogeneous and heterogeneous clusters. IEEE Transactions on Parallel and Distributed Systems, 19 (5):682–697, May 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2000:CEH

[XRR00] J. Xu, A. Romanovsky, and B. Randell. Concurrent exception handling and resolution in distributed object systems. IEEE Transactions on Parallel and Distributed Systems, 11(10):19–??, October 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1019.
pdf;   http://www.computer.
org/tpds/td2000/l1019abs.
htm.

Xiao:2009:CCD

[XRY09] Weijun Xiao, Jin Ren, and Qing Yang. A case for continuous data protection at block level in disk array storages. IEEE Transactions on Parallel and Distributed Systems, 20(6):898–911, June 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2006:TCD

[XTL06] Jianliang Xu, Xueyan Tang, and Wang-Chien Lee. Timecritical on-demand data broadcast: Algorithms, analysis, and performance evaluation. IEEE Transactions on Parallel and Distributed Systems, 17(1):3– 14, January 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2006/01/
l0003s.pdf.

Xu:2008:NSS

[XTL08] Jianliang Xu, Xueyan Tang, and Wang-Chien Lee. A new storage scheme for approximate location queries in object-tracking sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(2):262–275, February 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiang:2001:TBW

[XU01] Limin Xiang and Kazuo Ushijima. On time bounds, the work-time scheduling principle, and optimality for BSR. IEEE Transactions on Parallel and Distributed Systems, 12(9):912– 921, September 2001. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/l0912.
pdf;   http://www.computer.
org/tpds/td2001/l0912abs.
htm.

Xiang:1999:EIB

[XUAS99] L. Xiang, K. Ushijima, S. G. Akl, and I. Stojmenovic. An efficient implementation for the BROADCAST instruction of BSR+. IEEE Transactions on Parallel and Distributed Systems, 10(8):852–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0852.
pdf;   http://www.computer.
org/tpds/td1999/l0852abs.
htm.

Xiao:2003:LCR

[XXZ03] Li Xiao, Zhichen Xu, and Xiaodong Zhang. Low-cost and reliable mutual anonymity protocols in peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 14(9):829–840, September 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/09/l0829abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/09/l0829.
pdf.

Xiang:2009:FDP

[XZG09] Yang Xiang, Wanlei Zhou, and Minyi Guo. Flexible deterministic packet marking: An IP traceback system to find the real source of attacks. IEEE Transactions on Parallel and Distributed Systems, 20(4):567–580, April 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2005:DLM

[XZL05] Li Xiao, Zhenyun Zhuang, and Yunhao Liu. Dynamic layer management in superpeer architectures. IEEE Transactions on Parallel and Distributed Systems, 16(11):1078–1091, November 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2008:IBS

[XZNX08] Lijuan Xiao, Yanmin Zhu, Lionel M. Ni, and Zhiwei Xu. Incentive-based scheduling for market-like computational grids. IEEE Transactions on Parallel and Distributed Systems, 19 (7):903–913, July 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Youssef:1993:FTR

[YA93] Abdou Youssef and Bruce Arden. Functional and topological relations among Banyan multistage networks of differing switch sizes. IEEE Transactions on Parallel and Distributed Systems, 4(2):235–240, February 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Youn:1993:CPE

[YC93] Hee Young Youn and Calvin Ching-Yoen Chen. Comprehensive performance evaluation of crossbar networks. IEEE Transactions on Parallel and Distributed Systems, 4(5):481–489, May 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:1995:MCT

[YC95] Shyan-Ming Yuan and Her-Kun Chang. Message complexity of the tree quorum algorithm. IEEE Transactions on Parallel and Distributed Systems, 6(8): 887–890, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0887abs.htm.

Yu:1996:EPR

[YC96] Chang-Wu Yu and Gen-Huey Chen. Efficient parallel recognition algorithm for bipartitepermutation graphs. IEEE Transactions on Parallel and Distributed Systems, 7(1):3–10, January 1996. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:1998:EMA

[yCM98] C. y. Chang and P. Mohapatra. An efficient method for approximating submesh reliability of two-dimensional meshes. IEEE Transactions on Parallel and Distributed Systems, 9(11):115–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1115.
pdf;   http://www.computer.
org/tpds/td1998/l1115abs.
htm.

Yang:2007:RHI

[YCTW07] Jinn-Shyong Yang, Jou-Ming Chang, Shyue-Ming Tang, and Yue-Li Wang. Reducing the height of independent spanning trees in chordal rings. IEEE Transactions on Parallel and Distributed Systems, 18(5):644– 657, May 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:1994:EPB

[YD94a] Myung K. Yang and Chita R.

Das. Evaluation of a parallel branch-and-bound algorithm on a class of multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 5 (1):74–86, January 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:1994:PET

[YD94b] Philips S. Yu and Asit Dan.

Performance evaluation of transaction processing coupling architectures for handling system dynamics. IEEE Transactions on Parallel and Distributed Systems, 5(2):139–153, February 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:1995:DTA

[YD95] Chansu Yu and Chita R. Das. Disjoint task allocation algorithms for MIN machines with minimal conflicts. IEEE Transactions on Parallel and Distributed Systems, 6 (4):373–387, April 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1995/
                                        l0373abs.htm.

Yang:2009:FSF

[YDW^{+}09] Xuejun Yang, Yunfei Du, Panfeng Wang, Hongyi Fu, and Jia Jia. FTPA: Supporting fault-tolerant parallel computing through parallel recomputing. IEEE Transactions on Parallel and Distributed Systems, 20 (10):1471–1486, October 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yew:2002:E

[Yew02] P.-C. Yew. Editorial. IEEE Transactions on Parallel and Distributed Systems, 13(1):4–??, January 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0004.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0004abs.
                                        htm.

Yew:2003:EN

[Yew03] Pen-Chung Yew. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 14 (6):529–532, June 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2003/06/l0529.pdf.

Yew:2004:ENa

[Yew04a] Pen Yew. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 15(1):1– ??, January 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/01/
l0001.pdf.

Yew:2004:ENb

[Yew04b] Pen-Chung Yew. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 15 (3):193–195, March 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2004/03/
l0193.pdf;   http://csdl.
computer.org/dl/trans/td/
2004/03/l0193.htm.

Yew:2005:ENa

[Yew05a] Pen-Chung Yew. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 16 (1):1–3, January 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2005/01/
l0001.pdf;   http://csdl.

                                        computer.org/dl/trans/td/
                                        2005/01/l0001.htm.

Yew:2005:ENb

[Yew05b] Pen-Chung Yew. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 16(9): 769–771, September 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2005/09/
                                        l0769.pdf.

Yew:2006:EEF

[Yew06] Pen-Chung Yew. Editorial: EIC farewell and new EIC introduction. IEEE Transactions on Parallel and Distributed Systems, 17(1):1, January 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/trans/td/2006/01/
                                        l0001.pdf.

Yang:1997:HAS

[YF97] Tao Yang and Cong Fu. Heuristic algorithms for scheduling iterative task computations on distributed memory machines. IEEE Transactions on Parallel and Distributed Systems, 8(6):608–622, June 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l0608.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l0608abs.
                                        htm.

Yang:2001:RDT

[YFJ^{+}01] Yulu Yang, Akira Funahashi, Akiya Jouraku, Hiroaki Nishi, Hideharu Amano, and Toshinori Sueyoshi. Recursive diagonal torus: An interconnection network for massively parallel computers. IEEE Transactions on Parallel and Distributed Systems, 12(7):701–715, July 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0701.
pdf;   http://www.computer.
org/tpds/td2001/l0701abs.
htm.

Yao:1998:PCB

[YFM98] T. K. Yao, O. Frieder, and R. L. Martino. Parallel computation in biological sequence analysis. IEEE Transactions on Parallel and Distributed Systems, 9(3):283–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0283.
pdf;   http://www.computer.
org/tpds/td1998/l0283abs.
htm.

Yang:1994:DSP

[YG94] Tao Yang and Apostolos Gerasoulis. DSC: Scheduling parallel tasks on an unbounded number of processors. IEEE Transactions on Parallel and Distributed Systems, 5(9):951–967, September 1994. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2008:KMS

[YG08] Zhen Yu and Yong Guan. A key management scheme using deployment knowledge for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(10):1411–1425, October 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Younis:2006:LAC

[YGE06] Mohamed F. Younis, Kajaldeep Ghumman, and Mohamed Eltoweissy. Location-aware combinatorial key management scheme for clustered sensor networks. IEEE Transactions on Parallel and Distributed Systems, 17(8):865–882, August 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsieh:2002:EPA

[yH02]  Sun yuan Hsieh.  An ef-

ficient parallel algorithm for the efficient domination problem on distance-hereditary graphs. IEEE Transactions on Parallel and Distributed Systems, 13(9):985–993, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2002/pdf/l0985.
                                        pdf;   http://www.computer.
                                        org/tpds/td2002/l0985abs.
                                        htm.

Yi:2009:UAF

[Yi09] Chih-Wei Yi. A unified analytic framework based on minimum scan statistics for wireless ad hoc and sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20 (9):1233–1245, September 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yajnik:1997:ARD

[YJ97a] Shalini Yajnik and Niraj K. Jha.

Analysis and randomized design of algorithm-based fault tolerant multiprocessor systems under an extended model. IEEE Transactions on Parallel and Distributed Systems, 8(7):757– 768, July 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0757.
pdf;   http://www.computer.
org/tpds/td1997/l0757abs.
htm.

Yajnik:1997:GDA

[YJ97b] Shalini Yajnik and Niraj K.

Jha. Graceful degradation in algorithm-based fault tolerant multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 8(2):137– 153, February 1997. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/l0137.
pdf;   http://www.computer.

                                        org/tpds/td1997/l0137abs.
                                        htm.

Yu:2006:OBB

[YJHG06] Chen Yu, Xiaohong Jiang, Susumu Horiguchi, and Minyi Guo. Overall blocking behavior analysis of general banyanbased optical switching networks. IEEE Transactions on Parallel and Distributed Systems, 17(9):1037–1047, September 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yan:1997:ASP

[YJZ97] Yong Yan, Canming Jin, and Xiaodong Zhang. Adaptively scheduling parallel loops in distributed shared-memory systems. IEEE Transactions on Parallel and Distributed Systems, 8(1):70–81, January 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:1992:ICS

[YK92] Seung-Min Yang and K. H. Kim. Implementation of the conversation scheme in message-based distributed computer systems. IEEE Transactions on Parallel and Distributed Systems, 3(5): 555–572, September 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yamashita:1996:CANa

[YK96a] Masafumi Yamashita and Tsunehiko Kameda. Computing on anonymous networks: Part I: characterizing the solvable cases. IEEE Transactions on Parallel and Distributed Systems, 7 (1):69–89, January 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yamashita:1996:CANb

[YK96b] Masafumi Yamashita and Tsunehiko Kameda. Computing on anonymous networks: Part II: decision and membership problems. IEEE Transactions on Parallel and Distributed Systems, 7(1):90–96, January 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:1998:DTB

[YK98] J.-S. Yang and C.-T. King.

Designing tree-based barrier synchronization on 2D mesh networks. IEEE Transactions on Parallel and Distributed Systems, 9(6):526–??, June 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0526.
pdf;   http://www.computer.
org/tpds/td1998/l0526abs.
htm.

Yamashita:1999:LEP

[YK99] M. Yamashita and T. Kameda.

Leader election problem on networks in which processor identity numbers are not distinct. IEEE Transactions on Parallel and Distributed Systems, 10(9):878–??, September 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0878.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0878abs.
                                        htm.

Yau:2003:EEO

[YK03] Stephen S. Yau and Fariaz Karim. An energy-efficient object discovery protocol for context-sensitive middleware for ubiquitous computing. IEEE Transactions on Parallel and Distributed Systems, 14(11): 1074–1085, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/11/l1074abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/11/l1074.
                                        pdf.

Yeung:2009:GTP

[YK09] Mark Kai Ho Yeung and Yu-Kwong Kwok. On game theoretic peer selection for resilient peer-to-peer media streaming. IEEE Transactions on Parallel and Distributed Systems, 20(10): 1512–1525, October 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yum:2002:MQC

[YKDV02] Ki Hwan Yum, Eun Jung Kim, Chita R. Das, and Aniruddha S. Vaidya. MediaWorm: a QoS capable router architecture for clusters. IEEE Transactions on Parallel and Distributed Systems, 13(12):1261–1274, December 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2002/12/l1261abs.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1261.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/l1261.
pdf.

Yu:2008:DGT

[YKP08] Yang Yu, Bhaskar Krishnamachari, and Viktor K. Prasanna. Data gathering with tunable compression in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 19(2):276– 287, February 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://bell.computer.
org/dlcomments/.

Yang:2003:PBA

[YKS03] Mau-Tsuen Yang, Rangachar Kasturi, and Anand Sivasubramaniam. A pipelinebased approach for scheduling video processing algorithms on NOW. IEEE Transactions on Parallel and Distributed Systems, 14(2):119–130, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/02/l0119abs.
htm;   http://csdl.computer.

                                        org/dl/trans/td/2003/02/l0119.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/02/l0119.
                                        pdf.

Youn:1996:EDM

[YL96] Hee Yong Youn and Jae Young Lee. An efficient dictionary machine using hexagonal processor arrays. IEEE Transactions on Parallel and Distributed Systems, 7(3):266–273 (or 266–272??), March 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0266abs.htm.

Yue:1997:EPA

[YL97] K. K. Yue and D. J. Lilja. An effective processor allocation strategy for multiprogrammed shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 8(12):1246–1258, December 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1997/pdf/l1246.
                                        pdf;   http://www.computer.
                                        org/tpds/td1997/l1246abs.
                                        htm.

Yang:2007:BEO

[YL07] De-Nian Yang and Wanjiun Liao. On bandwidth-efficient overlay multicast. IEEE Transactions on Parallel and Distributed Systems, 18(11):1503– 1515, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2008:OSA

[YL08] De-Nian Yang and Wan-Jiun Liao. Optimal state allocation for multicast communications with explicit multicast forwarding. IEEE Transactions on Parallel and Distributed Systems, 19(4):476–488, April 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2007:ERM

[YLL^{+}07] Byung-Sun Yang, Junpyo Lee, SeungIl Lee, Seongbae Park, Yoo C. Chung, Suhyun Kim, Kemal Ebcioglu, Erik Altman, and Soo-Mook Moon. Efficient register mapping and allocation in LaTTe, an open-source Java just-in-time compiler. IEEE Transactions on Parallel and Distributed Systems, 18(1):57– 69, January 2007. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yangy:2007:SBM

[YLW07] Shuhui Yangy, Minglu Liz, and Jie Wu. Scan-based movement-assisted sensor deployment methods in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 18(8):1108–1121, August 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Youn:1995:MIN

[YM95] Hee Yong Youn and Youngsong Mun. On multistage interconnection networks with small clock cycles. IEEE Transactions on Parallel and Distributed Systems, 6(1):86–93, January 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2009:HAP

[YM09] Zhen Yang and Huadong Ma. Hotspot avoidance for P2P streaming distribution application: a game theoretic approach. IEEE Transactions on Parallel and Distributed Systems, 20(2): 219–232, February 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2003:ASC

[YMG03] Xin Yuan, Rami Melhem, and Rajiv Gupta. Algorithms for supporting compiled communication. IEEE Transactions on Parallel and Distributed Systems, 14(2):107–118, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/02/l0107abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/02/l0107.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/02/l0107.
                                        pdf.

Yan:2008:COR

[YMP08] Jinyao Yan, Martin May, and Bernhard Plattner. Comments on “Optimal Resource Allocation in Overlay Multicast”. IEEE Transactions on Parallel and Distributed Systems, 19 (4):572–573, April 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://bell.computer.
org/dlcomments/. See [CXN06].

Youssef:1990:BHN

[YN90] Abdou S. Youssef and Bhagirath Narahari. The Banyanhypercube networks. IEEE Transactions on Parallel and Distributed Systems, 1(2):160– 169, April 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2000:IDS

[YN00] M. Yang and L. M. Ni. Incremental design of scalable interconnection networks using basic building blocks. IEEE Transactions on Parallel and Distributed Systems, 11(11):126–??, November 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l1126.
pdf;   http://www.computer.
org/tpds/td2000/l1126abs.
htm.

Youssef:1993:PAR

[You93] A. Youssef. A parallel algorithm for random walk construction with application to the Monte Carlo solution of partial differential equations. IEEE Transactions on Parallel and Distributed Systems, 4(3):355–360, March 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yun:1998:CHC

[YP98] S. K. Yun and K. H. Park. Comments on “Hierarchical Cubic Networks”. IEEE Transactions on Parallel and Distributed Systems, 9(4):410–??, April 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1998/pdf/l0410.
                                        pdf;   http://www.computer.
                                        org/tpds/td1998/l0410abs.

htm. See [GD95].

Yang:1996:ERB

[YR96] Pei-Ji Yang and C. S. Raghavendra. Embedding and reconfiguration of binary trees in faulty hypercubes. IEEE Transactions on Parallel and Distributed Systems, 7(3):237–245, March 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic).  URL http://www.
                                        computer.org/tpds/td1996/
                                        l0237abs.htm.

Yu:2006:AAS

[YR06] Hao Yu and Lawrence Rauchwerger. An adaptive algorithm selection framework for reduction parallelization. IEEE Transactions on Parallel and Distributed Systems, 17(10):1084–1096, October 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:1992:DAC

[YTB92] Q. Yang, G. Thangadurai, and L. N. Bhuyan. Design of an adaptive cache coherence protocol for large scale multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 3 (3):281–293, May 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yeh:1998:MSN

[YV98] C.-H. Yeh and E. A. Varvarigos. Macro-star networks: Efficient low-degree alternatives to star graphs. IEEE Transactions on Parallel and Distributed Systems, 9(10):987–??, October 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0987.
pdf;   http://www.computer.
org/tpds/td1998/l0987abs.
htm.

Yang:2005:MAS

[YvdRC05] Yang Yang, Krijn van der Raadt, and Henri Casanova. Multiround algorithms for scheduling divisible loads. IEEE Transactions on Parallel and Distributed Systems, 16(11):1092– 1102, November 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:1993:NGA

[YW93] Qing Yang and Hong Wang.

New graph approach to minimizing processor fragmentation in hypercube multiprocessors.

IEEE Transactions on Parallel and Distributed Systems, 4(10): 1165–1171, October 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yen:1998:PER

[YW98] T.-Y. Yen and W. Wolf. Performance estimation for realtime distributed embedded systems. IEEE Transactions on Parallel and Distributed Systems, 9(11):125–??, November 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l1125.
pdf;   http://www.computer.
org/tpds/td1998/l1125abs.
htm.

Yang:1999:NSR

[YW99] Y. Yang and J. Wang. A new self-routing multicast network. IEEE Transactions on Parallel and Distributed Systems, 10(12):299–??, December 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l1299.
pdf;   http://www.computer.
org/tpds/td1999/l1299abs.
htm.

Yang:2000:OAA

[YW00] Y. Yang and J. Wang. Optimal all-to-all personalized exchange in self-routable multistage networks. IEEE Transactions on Parallel and Distributed Systems, 11(3):261–??, March 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0261.
pdf;   http://www.computer.
org/tpds/td2000/l0261abs.
htm.

Yang:2001:OAA

[YW01] Y. Yang and J. Wang. Optimal all-to-all personalized exchange in a class of optical multistage networks. IEEE Transactions on Parallel and Distributed Systems, 12(6):567–??, June 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0567.
pdf;   http://www.computer.
org/tpds/td2001/l0567abs.
htm.

Yang:2002:NOA

[YW02] Yuanyuan Yang and Jianchao Wang. Near-optimal all-to-all broadcast in multidimensional all-port meshes and tori. IEEE Transactions on Parallel and Distributed Systems, 13(2):128–141, February 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0128.
pdf;   http://www.computer.
org/tpds/td2001/l0128abs.
htm.

Yang:2003:NFM

[YW03a] Yuanyuan Yang and Jianchao Wang. Nonblocking k-fold multicast networks. IEEE Transactions on Parallel and Distributed Systems, 14(2):131–141, February 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/02/l0131abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0131.
htm;   http://csdl.computer.
org/dl/trans/td/2003/02/l0131.
pdf.

Yang:2003:RPL

[YW03b] Yuanyuan Yang and Jianchao Wang. Routing permutations with link-disjoint and node-disjoint paths in a class of self-routable interconnects. IEEE Transactions on Parallel and Distributed Systems, 14(4):383–393, April 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/04/l0383abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0383.
htm;   http://csdl.computer.
org/dl/trans/td/2003/04/l0383.
pdf.

Yang:2004:CMC

[YW04] Yuanyuan Yang and Jianchao Wang. A class of multistage conference switching networks for group communication. IEEE Transactions on Parallel and Distributed Systems, 15(3):228–243, March 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/03/l0228abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0228.
htm;   http://csdl.computer.
org/dl/trans/td/2004/03/l0228.
pdf.

Yang:2005:CED

[YW05a] Yuanyuan Yang and Jianchao Wang. Cost-effective designs of WDM optical interconnects. IEEE Transactions on Parallel and Distributed Systems, 16(1):51–66, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0051abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0051.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0051.
pdf.

Yang:2005:RPB

[YW05b] Yuanyuan Yang and Jianchao Wang. Routing permutations on baseline networks with nodedisjoint paths. IEEE Transactions on Parallel and Distributed Systems, 16(8):737–746, August 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2008:EDN

[YWD08] Shuhui Yang, Jie Wu, and Fei Dai. Efficient directional network backbone construction in mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 19(12): 1601–1613, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2009:ILA

[YWF^{+}09] Wei Yu, Xun Wang, Xinwen Fu, Dong Xuan, and Wei Zhao. An invisible localization attack to Internet threat monitors. IEEE Transactions on Parallel and Distributed Systems, 20(11): 1611–1625, November 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2008:SCM

[YWY08] Yuanyuan Yang, Jianchao Wang, and Min Yang. A service-centric multicast architecture and routing protocol. IEEE Transactions on Parallel and Distributed Systems, 19(1):35–51, January 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yin:2003:ORD

[YXW03] G. George Yin, Cheng-Zhong Xu, and Le Yi Wang. Optimal remapping in dynamic bulk synchronous computations via a stochastic control approach. IEEE Transactions on Parallel and Distributed Systems, 14(1):51–62, January 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/

trans/td/2003/01/l0051abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0051.
htm;   http://csdl.computer.
org/dl/trans/td/2003/01/l0051.
pdf.

Yang:1995:MIM

[YY95] Qing Yang and Tao Yang.

A memory interference model for regularly patterned multiple stream vector accesses. IEEE Transactions on Parallel and Distributed Systems, 6 (5):520–530, May 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0520abs.htm.

Yoo:1997:ETA

[YYS97] Seong-Moo Yoo, Hee Yong Youn, and Behrooz Shirazi. An efficient task allocation scheme for 2D mesh architectures. IEEE Transactions on Parallel and Distributed Systems, 8(9):934–942, September 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l0934.
pdf;   http://www.computer.
org/tpds/td1997/l0934abs.
htm.

Yang:2009:FTS

[YYX^{+}09] Xuejun Yang, Xiaobo Yan, Zuocheng Xing, Yu Deng, Jiang Jiang, Jing Du, and Ying Zhang. Fei Teng 64 stream processing system: Architecture, compiler, and programming. IEEE Transactions on Parallel and Distributed Systems, 20(8):1142– 1157, August 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Youssef:2009:OMC

[YYY09] Moustafa A. Youssef, Adel Youssef, and Mohamed F. Younis. Overlapping multihop clustering for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(12):1844–1856, December 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yue:2008:EOE

[YZC08] Jianhui Yue, Yifeng Zhu, and Zhao Cai. An energy-oriented evaluation of buffer cache algorithms using parallel I/O workloads. IEEE Transactions on Parallel and Distributed Systems, 19(11):1565–1578, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:1994:RMF

[YZW94] C. S. Yang, L.-P. Zu, and Y. N. Wu. Reconfigurable modular fault-tolerant hypercube architecture. IEEE Transactions on Parallel and Distributed Systems, 5(10):1018–1032, October 1994. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yan:2000:CRA

[YZZ00] Y. Yan, X. Zhang, and Z. Zhang.

Cacheminer: a runtime approach to exploit cache locality on SMP. IEEE Transactions on Parallel and Distributed Systems, 11(4):357–??, April 2000. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2000/pdf/l0357.
pdf;   http://www.computer.
org/tpds/td2000/l0357abs.
htm.

Zapata:1992:VCG

[ZA92] Emilio L. Zapata and F. Arguello. A VLSI constant geometry architecture for the fast Hartley and Fourier transforms. IEEE Transactions on Parallel and Distributed Systems, 3 (1):58–70, January 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:1993:JSH

[ZA93] Yahui Zhu and Mohan Ahuja.

On job scheduling on a hypercube. IEEE Transactions on Parallel and Distributed Systems, 4(1):62–69, January 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2009:ILM

[ZB09] Ye Zhu and Riccardo Bettati. Information leakage as a model for quality of anonymity networks. IEEE Transactions on Parallel and Distributed Systems, 20 (4):540–552, April 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2005:VAD

[ZBJ^{+}05] Yuanyuan Zhou, Angelos Bilas, Suresh Jagannathan, Dimitrios Xinidis, Cezary Dubnicki, and Kai Li. VI-attached database storage. IEEE Transactions on Parallel and Distributed Systems, 16(1):35–50, January 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2005/01/l0035abs.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0035.
htm;   http://csdl.computer.
org/dl/trans/td/2005/01/l0035.
pdf.

Zhu:2009:LOP

[ZBM09] Julie Wenbin Zhu, Patrick G. Bridges, and Arthur (Barney) B. Maccabe. Lightweight online performance monitoring and tuning with embedded gossip. IEEE Transactions on Parallel and Distributed Systems, 20 (7):1038–1049, July 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2004:SLB

[ZCL04] Yuanyuan Zhou, Zhifeng Chen, and Kai Li. Second-level buffer cache management. IEEE Transactions on Parallel and Distributed Systems, 15(6): 505–519, June 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/06/l0505.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/06/l0505.pdf.

Zhang:2006:CAM

[ZCLC06] Zhan Zhang, Shigang Chen, Yibei Ling, and Randy Chow. Capacity-aware multicast algorithms on heterogeneous overlay networks. IEEE Transactions on Parallel and Distributed Systems, 17(2):135–147, February 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zomaya:1998:FRB

[ZCO98] A. Y. Zomaya, M. Clements, and S. Olariu. A framework for reinforcement-based scheduling in parallel processor systems. IEEE Transactions on Parallel and Distributed Systems, 9(3):249–??, March 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0249.
pdf;   http://www.computer.
org/tpds/td1998/l0249abs.
htm.

Zhuge:2008:HSP

[ZCSY08] Hai Zhuge, Xue Chen, Xiaoping Sun, and Erlin Yao. HRing: a structured P2P overlay based on harmonic series. IEEE Transactions on Parallel and Distributed Systems, 19(2):145–158, February 2008. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2009:CCL

[ZCXF09] Xiaoying Zheng, Feng Chen, Ye Xia, and Yuguang (Michael) Fang. A class of cross-layer optimization algorithms for performance and complexity tradeoffs in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 20(10): 1393–1407, October 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2007:FGR

[ZF07] Yanchao Zhang and Yuguang Fang. A fine-grained reputation system for reliable service selection in peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 18(8): 1134–1145, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2003:IAP

[ZFMS03] Yanyong Zhang, Hubertus Franke, Jose Moreira, and Anand Sivasubramaniam. An integrated approach to parallel scheduling using gangscheduling, backfilling, and migration. IEEE Transactions on Parallel and Distributed Systems, 14(3):236–247, March 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/03/l0236abs.
                                        htm;   http://csdl.computer.

org/dl/trans/td/2003/03/l0236.
htm;   http://csdl.computer.
org/dl/trans/td/2003/03/l0236.
pdf.

Zhu:1998:NPD

[ZH98] S. Zhu and G. M. Huang. A new parallel and distributed shortest path algorithm for hierarchically clustered data networks. IEEE Transactions on Parallel and Distributed Systems, 9(9):841–??, September 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0841.
pdf;   http://www.computer.
org/tpds/td1998/l0841abs.
htm.

Zhu:1999:GEI

[ZH99a] Y. Zhu and L. Hendren. Guest Editors’ introduction. IEEE Transactions on Parallel and Distributed Systems, 10(2):97– ??, February 1999. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1999/pdf/l0097.
pdf.

Zhu:1999:LAP

[ZH99b] Y. Zhu and L. Hendren. Locality analysis for parallel C programs. IEEE Transactions on Parallel and Distributed Systems, 10(2):99–??, February 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0099.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0099abs.
                                        htm.

Zhu:2005:EPA

[ZH05] Yingwu Zhu and Yiming Hu. Efficient, proximity-aware load balancing for DHT-based P2P systems. IEEE Transactions on Parallel and Distributed Systems, 16(4):349–361, April 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2006:ESP

[ZH06] Yingwu Zhu and Yiming Hu. Enhancing search performance on Gnutella-like P2P systems. IEEE Transactions on Parallel and Distributed Systems, 17(12): 1482–1495, December 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2007:APP

[ZH07a] Rongmei Zhang and Y. Charlie Hu. Assisted peer-to-peer search with partial indexing. IEEE Transactions on Parallel and Distributed Systems, 18(8): 1146–1158, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2007:PRS

[ZH07b] Runfang Zhou and Kai Hwang. PowerTrust: a robust and scalable reputation system for trusted peer-to-peer computing. IEEE Transactions on Parallel and Distributed Systems, 18 (4):460–473, April 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2007:FPB

[ZH07c] Yingwu Zhu and Yiming Hu.

Ferry: a P2P-based architecture for content-based publish/ subscribe services. IEEE Transactions on Parallel and Distributed Systems, 18(5):672–685, May 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2003:GES

[Zha03] Wei Zhao. Guest Editorial:

Special section on security in distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 14(9): 817–??, September 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.
org/comp/trans/td/2003/09/
l0817.pdf.

Ziavras:1993:EMA

[Zia93] Sotirios G. Ziavras. Efficient mapping algorithms for a class of hierarchical systems. IEEE Transactions on Parallel and Distributed Systems, 4(11):1230– 1245, November 1993. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ziavras:1994:RVF

[Zia94] Sotirios G. Ziavras. RH: a versatile family of reduced hypercube interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 5(11): 1210–1220, November 1994. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zou:1999:RTP

[ZJ99] H. Zou and F. Jahanian. A real-time primary-backup replication service. IEEE Transactions on Parallel and Distributed Systems, 10(6):533–??, June 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1999/pdf/l0533.
pdf;   http://www.computer.
org/tpds/td1999/l0533abs.
htm.

Zhang:2003:RMA

[ZJ03] Charles Zhang and Hans-Arno Jacobsen. Refactoring middleware with aspects. IEEE Transactions on Parallel and Distributed Systems, 14(11): 1058–1073, November 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2003/11/l1058abs.
htm;   http://csdl.computer.
org/dl/trans/td/2003/11/l1058.
pdf.

Zhu:2008:HDM

[ZJWX08] Yifeng Zhu, Hong Jiang, Jun Wang, and Feng Xian. HBA: Distributed metadata management for large cluster-based storage systems. IEEE Transactions on Parallel and Distributed Systems, 19(6):750–763, June 2008.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2008:DDC

[ZKB08] Zhenzhou Zhu, Panos Kalnis, and Spiridon Bakiras. DCMP: a distributed cycle minimization protocol for peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 19(3):363–377, March 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:1996:OSL

[ZL96] S. Q. Zheng and Shahram Latifi. Optimal simulation of linear multiprocessor architectures on multiply-twisted cube using generalized Gray Codes. IEEE Transactions on Parallel and Distributed Systems, 7 (6):612–619, June 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1996/
l0612abs.htm.

Zhuang:2005:RBB

[ZL05] Xiaotong Zhuang and Vincenzo Liberatore. A recursion-based broadcast paradigm in wormhole routed networks. IEEE Transactions on Parallel and Distributed Systems, 16(11):1034– 1052, November 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2007:NSL

[ZL07a] Zaichen Zhang and Victor Li. Network-supported layered multicast transport control for streaming media. IEEE Transactions on Parallel and Distributed Systems, 18(9):1332– 1344, September 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Zhou:2007:ASC

[ZL07b] Dong Zhou and Ten H. Lai. An accurate and scalable clock synchronization protocol for IEEE 802.11-based multihop ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 18(12):1797– 1808, December 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Zhu:2008:ONL

[ZL08] Ying Zhu and Baochun Li. Overlay networks with linear capacity constraints. IEEE Transactions on Parallel and Distributed Systems, 19(2):159–173, February 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zalamea:2004:RCM

[ZLAV04] Javier Zalamea, Josep Llosa, Eduard Ayguadé, and Mateo Valero. Register constrained modulo scheduling. IEEE Transactions on Parallel and Distributed Systems, 15(5):417–430, May 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/05/l0417abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0417.
htm;   http://csdl.computer.
org/dl/trans/td/2004/05/l0417.
pdf.

Zahorjan:1991:ESD

[ZLE91] John Zahorjan, Edward D. Lazowska, and Derek L. Eager. The effect of scheduling discipline on spin overhead in shared memory parallel systems. IEEE Transactions on Parallel and Distributed Systems, 2(2):180–198, April 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2007:SBF

[ZLKK07] Yizheng Zhou, Vijay Lakamraju, Israel Koren, and C. M. Krishna. Software-based failure detection and recovery in programmable network interfaces. IEEE Transactions on Parallel and Distributed Systems, 18(11): 1539–1550, November 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2009:DMO

[ZLP09] Ying Zhu, Baochun Li, and Ken Qian Pu. Dynamic multicast in overlay networks with linear capacity constraints. IEEE Transactions on Parallel and Distributed Systems, 20(7):925– 939, July 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2009:HOR

[ZLZN09] Hongzi Zhu, Minglu Li, Yanmin Zhu, and Lionel M. Ni. HERO: Online real-time vehicle tracking. IEEE Transactions on Parallel and Distributed Systems, 20 (5):740–752, May 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2003:SDV

[ZMC03] Dakai Zhu, Rami Melhem, and Bruce R. Childers. Scheduling with dynamic voltage/speed adjustment using slack reclamation in multiprocessor realtime systems. IEEE Transactions on Parallel and Distributed Systems, 14(7):686–700, July 2003. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        trans/td/2003/07/l0686abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/td/2003/07/l0686.
                                        pdf.

Zhu:2004:PAS

[ZMM04] Dakai Zhu, Daniel Mossé, and Rami Melhem. Power-aware scheduling for AND/OR graphs in real-time systems. IEEE Transactions on Parallel and Distributed Systems, 15(9):849– 864, September 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/09/l0849.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/09/l0849.pdf.

Zhao:2008:RBE

[ZMMS08] Wenbing Zhao, Louise E.

Moser, and P. M. Melliar-Smith. A reservation-based extended transaction protocol. IEEE Transactions on Parallel and Distributed Systems, 19(2): 188–203, February 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhuo:2007:HPR

[ZMP07] Ling Zhuo, Gerald R. Morris, and Viktor K. Prasanna. Highperformance reduction circuits using deeply pipelined operators on FPGAs. IEEE Transactions on Parallel and Distributed Systems, 18(10):1377–1392, October 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2008:PGL

[ZMRS08] Qi Zhang, Ningfang Mi, Alma Riska, and Evgenia Smirni. Performance-guided load (Un)balancing under autocorrelated flows. IEEE Transactions on Parallel and Distributed Systems, 19 (5):652–665, May 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2004:ECA

[ZN04] Pei Zheng and Lionel M. Ni.

EMPOWER: a cluster architecture supporting network emulation. IEEE Transactions on Parallel and Distributed Systems, 15 (7):617–629, July 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/dl/trans/td/
2004/07/l0617.htm;   http:
//csdl.computer.org/dl/trans/
td/2004/07/l0617.pdf.

Zhuo:2007:SMA

[ZP07] Ling Zhuo and Viktor K. Prasanna. Scalable and modular algorithms for floating-point matrix multiplication on reconfigurable computing systems. IEEE Transactions on Parallel and Distributed Systems, 18 (4):433–448, April 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2006:ALQ

[ZPY06] Jin Zhu, Symeon Papavassiliou, and Jie Yang. Adaptive localized QoS-constrained data aggregation and processing in distributed sensor networks. IEEE Transactions on Parallel and Distributed Systems, 17(9): 923–933, September 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2005:WAL

[ZRS^{+}05] Qi Zhang, Alma Riska, Wei Sun, Evgenia Smirni, and Gianfranco Ciardo. Workloadaware load balancing for clustered Web servers. IEEE Transactions on Parallel and Distributed Systems, 16(3):219–233, March 2005. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:1995:EIF

[ZS95a] Qin Zheng and Kang G. Shin.

Establishment of isolated failure immune real-time channels in HARTS. IEEE Transactions on Parallel and Distributed Systems, 6(2):113–119, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1995/
l0113abs.htm.

Zheng:1995:SBA

[ZS95b] Qin Zheng and Kang G. Shin.

Synchronous bandwidth allocation in FDDI networks. IEEE Transactions on Parallel and Distributed Systems, 6(12):1332– 1338, December 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:1998:FTR

[ZS98] Q. Zheng and K. G. Shin.

Fault-tolerant real-time communication in distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 9(5):470–??, May 1998. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1998/pdf/l0470.
pdf;   http://www.computer.
org/tpds/td1998/l0470abs.
htm.

Zhang:2009:BEC

[ZS09] Haibo Zhang and Hong Shen. Balancing energy consumption to maximize network lifetime in data-gathering sensor networks. IEEE Transactions on Parallel and Distributed Systems, 20(10): 1526–1539, October 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:1992:HDS

[ZSLW92] Songnian Zhou, Michael Stumm, Kai Li, and David Wortman. Heterogeneous distributed shared memory. IEEE Transactions on Parallel and Distributed Systems, 3(5):540–554, September 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2001:IWS

[ZSMF01] Yanyong Zhang, Anand Sivasubramaniam, José Moreira, and Hubertus Franke. Impact of workload and system parameters on next generation cluster scheduling mechanisms. IEEE Transactions on Parallel and Distributed Systems, 12(9):967–985, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td2001/pdf/l0967.
                                        pdf;   http://www.computer.
                                        org/tpds/td2001/l0967abs.
                                        htm.

Zomaya:2001:OUG

[ZT01] Albert Y. Zomaya and Yee-Hwei Teh. Observations on using genetic algorithms for dynamic load-balancing. IEEE Transactions on Parallel and Distributed Systems, 12(9):899–911, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0899.
pdf;   http://www.computer.
org/tpds/td2001/l0899abs.
htm.

Zomaya:2002:OUG

[ZW02] Albert Y. Zomaya and Michael Wright. Observations on using genetic-algorithms for channel allocation in mobile computing. IEEE Transactions on Parallel and Distributed Systems, 13(9):948–962, September 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2002/pdf/l0948.
pdf;   http://www.computer.
org/tpds/td2002/l0948abs.
htm.

Zomaya:1999:GSP

[ZWM99] A. Y. Zomaya, C. Ward, and B. Macey. Genetic scheduling for parallel processor systems: Comparative studies and performance issues. IEEE Transactions on Parallel and Distributed Systems, 10(8):795–??, August 1999. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://dlib.computer.org/
                                        td/books/td1999/pdf/l0795.
                                        pdf;   http://www.computer.
                                        org/tpds/td1999/l0795abs.
                                        htm.

Zhou:2006:RAS

[ZWX06] Xiaobo Zhou, Jianbin Wei, and Cheng-Zhong Xu. Resource allocation for session-based twodimensional service differentiation on e-commerce servers. IEEE Transactions on Parallel and Distributed Systems, 17(8): 838–850, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2004:HPB

[ZX04] Xiaobo Zhou and Cheng-Zhong Xu. Harmonic proportional bandwidth allocation and scheduling for service differentiation on streaming servers. IEEE Transactions on Parallel and Distributed Systems, 15(9): 835–848, September 2004. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/trans/td/
                                        2004/09/l0835.htm;   http:
                                        //csdl.computer.org/dl/trans/
                                        td/2004/09/l0835.pdf.

Zhang:2009:OTD

[ZXZ^{+}09] Meng Zhang, Yongqiang Xiong, Qian Zhang, Lifeng Sun, and Shiqiang Yang. Optimizing the throughput of data-driven peerto-peer streaming. IEEE Transactions on Parallel and Distributed Systems, 20(1):97–110, January 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:1995:CME

[ZY95] Xiaodong Zhang and Yong Yan. Comparative modeling and evaluation of CC-NUMA and COMA on hierarchical ring architectures. IEEE Transactions on Parallel and Distributed Systems, 6(12):1316–1331, December 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2004:OSA

[ZY04] Zhenghao Zhang and Yuanyuan Yang. Optimal scheduling algorithms in WDM optical interconnects with limited range wavelength conversion capability. IEEE Transactions on Parallel and Distributed Systems, 15(11):1012–1026, November 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/comp/
trans/td/2004/11/l1012abs.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1012.
htm;   http://csdl.computer.
org/dl/trans/td/2004/11/l1012.
pdf.

Zhang:2006:WOI

[ZY06] Zhenghao Zhang and Yuanyuan Yang. WDM optical interconnects with recirculating buffering and limited range wavelength conversion. IEEE Transactions on Parallel and Distributed Systems, 17(5):466–480, May 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2007:AHC

[ZY07] Si Qing Zheng and Mei Yang. Algorithm-hardware codesign of fast parallel round-robin arbiters. IEEE Transactions on Parallel and Distributed Systems, 18(1):84–95, January 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:1995:CPE

[ZYC95] Xiaodong Zhang, Yong Yan, and Robert Castaneda. Comparative performance evaluation of hot spot contention between MINbased and ring-based sharedmemory architectures. IEEE Transactions on Parallel and Distributed Systems, 6(8):872– 886, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0872abs.
htm.

Zeinalipour-Yazti:2007:PPA

[ZYKG07] Demetrios Zeinalipour-Yazti, Vana Kalogeraki, and Dimitrios Gunopulos. pFusion: a P2P architecture for Internet-scale content-based search and retrieval. IEEE Transactions on Parallel and Distributed Systems, 18(6):804–817, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2007:PSS

[ZZMN07] Feng Zhu, Wei Zhu, Matt W.

Mutka, and Lionel M. Ni. Private and secure service discovery via progressive and probabilistic exposure. IEEE Transactions on Parallel and Distributed Systems, 18(11):1565–1577, November 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).